A Complete Bibliography of Publications in Fisheries Research (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 February 2025 Version 1.50

Title word cross-reference

1 [DJFU20]. 2 [CMRP20]. 3 [DMM^{+}21, DJFU20, MCMM20, WZL^{+}22]. 30 [HPPT24]. ^{14} [CPR^{+}24]. ^{15} [FRP22]. ^{86} [ZOS^{+}23]. ^{87} [ZOS^{+}23]. B_{0} [BBPT^{+}22]. δ^{18} [PD25, RWFT25]. F_{msy} [HH20]. i [BW20]. L_{50} [MBOCdAM24]. M [CWRR24, CH22]. × [HML^{+}20].

-crab [MPEBdR23]. -dependent [AGB^{+}24]. -ha [DJFU20]. -pass [DJFU20]. -years [HPPT24].

0.2 [CH22].

1 [WCN^{+}24]. 10-year [ZJJ^{+}25]. 19 [Apo25, FBM^{+}21, HJA^{+}21, PZL^{+}23, SCC^{+}22].

2 [CWRR24]. 2020’ [MBD^{+}21, Ano20a, Ano20b, Ano20c, Ano20p, Ano20q, Ano20s, Ano20r, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x]. 2021 [Ano21a, Ano21b, Ano21c, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, BWN^{+}21]. 2022 [Ano22a, Ano22b, Ano22c, Ano22p, Ano22r, Ano22t, Ano22s, Ano22u, Ano22v, Ano22w, Ano22x, Ano22q, Ano22y]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x]. 2024 [Ano24a, Ano24b, Ano24c, Ano24p, Ano24q, Ano24s, Ano24r, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x]. 2025 [Ano25d, Ano25e, Ano25f]. 218 [WGC^{+}21]. 219 [Tho21]. 221 [Ash20a]. 233 [VK21a]. 236 [YH21a]. 238 [OS21a]. 243 [KBPS22]. 255 [DBV23]. 273 [SIM^{+}24a].

45^{o} [TNDM23]. 47^{o} [TNDM23]. 48^{o} [MFM^{+}20]. 4Mediterranean [CPL^{+}25].

7 [LG21].

abalone [HSM21, NAV^{+}23, PRA^{+}23, VAVQGD^{+}20]. ability [CTS^{+}23, CFP22, FKW^{+}22]. able [DCL^{+}20, TCVG20]. Abundance [FGSD25, SRT^{+}20, AAPG21, AAT^{+}21, BGG^{+}22, BFA^{+}21, BHH21, CPL^{+}25, CMV21, CMBL21, DLKH22, DCR^{+}20, DHCS23, FSP22, FAK24, GMPD23, HPPT24, HCDF24, HC22c, JBČ^{+}22, KHL^{+}24, KSI20b, LD25, LCW23b, MTX^{+}20, MKS^{+}22, MKS^{+}21a, MYKO23, OOAF^{+}21, PFFdC22, PMC^{+}24, PSS^{+}21, RCVGMI22, RWT^{+}20, SS23, SPM^{+}24, SFYM24, TSC^{+}22, TMP20]. abundance-index [MYKO23]. abundances [CTCB22]. abundant [SSSF25]. acanthias [HM25]. accelerometer [RNP^{+}24]. access [DTSR22, KKCP20, VP22]. access-point [VP22]. according [KAB^{+}22]. accordion [YLS^{+}23]. accordion-shaped [YLS^{+}23]. account [HM25, MRG^{+}23, WZX^{+}20, WH23]. accountability [BGCCP22]. Accounting [Aks24, CMV21, CGC24, CHAY^{+}25, PBDM23, SJW^{+}22, DCK^{+}22]. accounts [CAAFH21]. accuracy [AYTM21, SFMA23]. accurate [CAZN24, FCMP23]. Aceh [AIM^{+}23, FMSA21]. Acetes [SZS^{+}24, WZL^{+}22]. achieved [BČD^{+}21, MAH^{+}22]. Achieving [DCS24, RUHM20]. acid [ASD^{+}22, PZG^{+}20, QMGRIU22]. acidification [PDF20, TLC^{+}22]. acids [CZ25, TNDM23]. Acoupa [SdOR^{+}23]. Acoustic [JND^{+}23, MKS^{+}21b, VMI21, WOG^{+}25, YMYH20, YAO^{+}23, BHNP22, BHD^{+}23, CCKL^{+}20, CBD^{+}22, ERS^{+}23, FGTA24, GLP^{+}20, JRW^{+}21, KWE^{+}21, KSS^{+}22, KBM23, LD25, PGD^{+}25, PCC^{+}23, RGG22, RGN^{+}20, WPB22]. acoustic-based [LD25]. acoustic-optic [JRW^{+}21]. acoustical [LPAE^{+}24]. acoustically [KBB^{+}21]. Acoustics [CTR^{+}21, KMO20, KBM23, Peñ21, TGG^{+}24]. acronotus [KFDE^{+}22]. across [BKC21, FBB20, KKLM24, MDC^{+}22b, ÖA21, PRK23, Ten22, TDI^{+}21]. action [KC22, WCGB22]. actions [HF20, SKBA23]. Active [GGL^{+}24, MLS^{+}21, PCC^{+}23, RB22, ŠBB^{+}22]. active-selection [RB22]. activities [Ham22, YMS21]. activity [BLC^{+}22, FGCB^{+}21, GGMRC^{+}22, HBE^{+}22, LNP25, MPM^{+}23, VAVQGD^{+}20, Žák21]. ActSel [RB22]. aculeatus [MHB^{+}23]. acutus [KCB^{+}24]. adaptation [MA20, SSG^{+}22]. Adapting [LMM^{+}24]. Adaptive [KMC^{+}23, SPC22, SPC^{+}23, JGU21]. added [Seu22]. additional [SSP24]. additive [LCL25, LK25]. Addressing [CGC24, BSAP22]. adequacy [MBOCdAM24]. adjacency [MLCMdS23]. adjacent [FFG^{+}20]. adopted [BBJ22, PSS^{+}20]. Adoption [COG22, RBG^{+}24]. Adriatic [AMSC20, CMD^{+}23, PDC^{+}23]. adult [DLZ^{+}25, EMR^{+}22, HMC^{+}23, LP23, MCHA21, PPC^{+}23b, SKJJ25, WW21b]. Advancement [MCS^{+}25]. Advances [GP21, DGMG^{+}22]. adversarial [MZ24]. advice [RMNB^{+}21]. Aegean [AHEV24, GCK^{+}21, ŞGK^{+}20, ÖÜÖG20]. aeglefinus [SKD^{+}20, WGFM21]. Aequipecten [MSV21, OOAF^{+}21, TLAM25]. aequispinus [SDV^{+}22]. Aerial [WSL21, DLP^{+}24]. affect [AUHK22, BÖN20, GKC21, GdSPL21, LLFL21, LCHB^{+}24, MOI23, NGDC25, WZS^{+}21, YiTM23, dLHER24]. affecting [CTM^{+}20, GW21, LS24, NMJ^{+}24, TZL^{+}24]. affects [BGM^{+}23, NVSG24, ŠBB^{+}22, SCS25]. AFORO [MCGL^{+}25]. Africa [NAS^{+}20, NNS^{+}22, DMM^{+}23, FCMP23, GMERCM^{+}24, RCH^{+}21, SSM^{+}23]. African [BSAP22, BCM^{+}21, CNE^{+}22, CAAFH21, GMERCM^{+}24, OCdMC24, SCN^{+}24]. after [ABK^{+}21, BTFL22, HUK^{+}23, JSKM20, MAH^{+}22]. Age [FHE^{+}24, GFDN^{+}22, MZZ^{+}21, NVSG24, SR21, SdOR^{+}23, SLW^{+}20, AT20, AAT^{+}21, ACS23, ATA^{+}24, ALRB^{+}20, BVR^{+}21, BTB^{+}21, BHB24, BGBM22, BHH21, BKHA21, BCOBB^{+}23, BRN^{+}20, BCM^{+}21, Cam23, CSB^{+}23, CTIC23, CSRL20, DCS24, ECK^{+}21, EVS^{+}23, FCKG^{+}22, dlTGPLC23, HW24, HKCW24, HMR^{+}24b, İTAD24, JCCAS^{+}21, JMS25, KFO20, LMP24, Lor22, MCS^{+}24, MMP^{+}24, NSRM22, PJSQ20, PUC^{+}23, PPC^{+}21, PJMP22, PTD^{+}21, PTL^{+}24, RdBAT^{+}23, SSV^{+}20, SFMA23, SDdMG^{+}20, SPD^{+}24, SF20, SBRM^{+}22, SWH24b, SM21, TDJ^{+}21, TNDM23, TSI^{+}21, VK21a, VK21b, VBB20, WPB^{+}20, WPGO21, WBA23, dSTMV20, vDJB^{+}23]. age-0 [BTB^{+}21, TNDM23]. age-at-length [HW24]. age-at-maturity [SWH24b]. age-based [AT20]. age-dependent [Lor22]. age-growth [BCM^{+}21]. Age-structured [MZZ^{+}21, PJMP22]. age-varying [SM21]. Ageing [MAH^{+}22, CPR^{+}24, HW24, NSRM22]. agent [DH20]. agents [HKKa^{+}25]. ages [DCS24]. aggregated [AAT^{+}21]. Aggregating [BDA^{+}20, AUM21, IKBL23, KNP^{+}20, SFC21, TFC^{+}20, ZZC^{+}21]. aggregation [AVB^{+}23, CQA^{+}24, FTH^{+}23, MFO21, MFR^{+}22b, RFMS^{+}21]. aggregations [BDA^{+}20, BHST^{+}21]. Aging [BWB^{+}24, HKCW24, SKW^{+}21]. ahead [BS20a]. AI [PNGGO^{+}22]. aid [AGB^{+}24]. aimed [FS20]. Air [BKR^{+}22, AMM^{+}22, CGB^{+}22, MSJ21]. air-breathing [AMM^{+}22]. air-exposure [CGB^{+}22]. Aires [DDA^{+}20]. al [BWN^{+}21, MBD^{+}21]. Alabama [BHNP22]. Alaska [AHB^{+}22, BWOR23, CGC24, CH21b, DB22, GSS^{+}23, GBWM22, GPT^{+}21, Hut22, JRW^{+}21, LADA^{+}22, SW24, SDV^{+}22, SNHM23, WBA23]. Alaskan [NGDC25]. albacares [AAFLL^{+}25, FFG^{+}20, PBPM^{+}23, PZG^{+}20, SF22a, ZZC^{+}21]. albacore [SML^{+}24]. albatross [GJSW22]. Albert [NCB^{+}23]. Alberta [FRP22]. Alboran [VMI21]. albula [KSV^{+}22, SHV20]. Alburnus [SPW^{+}22]. alchemy [Cop24]. alcocki [PCGG20]. ALert [CCKL^{+}20]. alerting [CCKL^{+}20]. Aleutian [KDF^{+}25, SDV^{+}22, SC20]. algal [Béc20]. algorithm [GC25, LCLM23, SZS^{+}24, XDX^{+}23, ZHC^{+}25]. algorithms [BGBM22, Tho19, Tho21, YAO^{+}23]. alien [BN21, Dik24, HKKa^{+}25]. alizarin [CHL^{+}20, LR21, SW20]. alletteratus [BNTK23]. Allocating [PPD^{+}25]. Allocation [KK21, DSJG20, PvZ22]. allometry [RdBAT^{+}23]. Allopatric [AdABW^{+}22]. allow [GRJW20]. along [BGG^{+}22, BBM^{+}24, BNL^{+}23, BJK24, CKD^{+}21, DKBF23, Dra22, dSRFFN^{+}20, FJJT^{+}21, GMRRG20, JvPOG25, MA20, MOA23, ML24, MRC24, NB24, OAM^{+}21, ÖSL^{+}23, PYX^{+}20, PCBL23, PASdCF23, SCN^{+}24, SCHSC21, VAVQGD^{+}20, VBL^{+}24, WBBG^{+}23]. alongshore [KNS^{+}22]. Alopias [AWVS21]. Alopiidae [MRC24]. Alosa [BHD^{+}23, HRH22]. alphabet [WSL^{+}24]. Alpine [ZOS^{+}23]. alpinus [LvCdGSL23]. alter [FGCB^{+}21]. alteration [TBÓ^{+}22]. altered [RFF^{+}22]. alternate [RB22]. alternative [CFB^{+}23, GZLCRG25, MPH21, MFJ^{+}24, PTD^{+}21, PBDM21, MMML24]. alters [XSS^{+}23]. always [SF20, vdHR23]. Amakusa [YTHM20]. Amazon [ASD^{+}22, dFBPL^{+}20, dSRFFN^{+}20, PASdCF23, PdAMdM^{+}23, PFdSBL25, SdFZFJ21]. Amazonian [SdOR^{+}23]. Amending [SSKS21, SRB^{+}25]. America [ACP^{+}23, SS24, ZCFG23, dSTMV20]. American [ANB^{+}24, BHD^{+}23, EWPB22, GLA^{+}20, GGL^{+}24, HSM^{+}24, JGG^{+}24, LPS^{+}25, LBP^{+}24, LCMS^{+}22, LYLC21, Mac22, MKS^{+}21a, RRH^{+}24, RRB24]. americanus [BAW^{+}24, CKM^{+}20, EWPB22, GGL^{+}24, HSM^{+}24, JGG^{+}24, LPS^{+}25, LBP^{+}24, MKS^{+}21a]. Americas [GdSPL21]. amidst [Apo25]. Ammodytes [SBL^{+}23]. among [AGS20, ALW^{+}21, BBJ22, DHH^{+}22, Det23, FTH^{+}23, FBW^{+}21, GQP^{+}25, HLCC22, KPS20, LOFS22, LDM^{+}24, MCC20, NDRR20, OMG^{+}23, PW25, SPC^{+}23, SF20, SHC21, ZLXL20]. amount [SKD^{+}20]. Amphioctopus [JZQZ20]. Amundsen [ZSDZ24]. Anadara [ZJY^{+}24]. anadromous [HC22c, JND^{+}23]. anadromy [RDR^{+}23, vP20]. anaesthesia [TRS^{+}24]. analyses [BÁP^{+}23, BRR^{+}21, CZ25, CPPK23, LJX^{+}20, Mur20, NVRG^{+}21, PBM^{+}23, QMGRIU22, SFMA23, WHCF22, WWT^{+}25]. ANalysis [WZX^{+}20, AAVÁM23, DSP22, LS24, LCL25, MCHA21, AUC25, Alv21, ACS23, Apo25, AOA^{+}22, ASJ^{+}22, BCG^{+}25, BHST^{+}21, BRGB^{+}23, CPM21, CEAL21, CCR24, CPB^{+}21, CSB^{+}23, CHPT20, CAGLT23, CdSLP21, CAYM^{+}23, DRSPTA20, DSP^{+}23, DHC^{+}20, EHBJ20, FHSC21, FM21, FGPVPPGG22, FKW^{+}22, GGTÁVT^{+}20, GPW^{+}20, HHJ^{+}22, HF20, HWW22, HLC^{+}25, HOQ20, HUK^{+}23, HRC23, HIF^{+}24, HCWH23, HLZ^{+}20, KM23, KMSJ^{+}25, KAC^{+}23, KPS20, KPWS21, LGD^{+}20, LVP22, LCW23b, MRC24, NSM^{+}21, NS23, OFF^{+}20, PJNGJ^{+}22, PDJ24, QSMG^{+}23, RCVGMI22, RPM^{+}21, SCGM^{+}21, SBC^{+}22, SK21a, SCN^{+}24, Seu22, SLW^{+}21, SYZ^{+}25, SF20, SF22b, SMLT24, TSG25, UVA20, VAVQGD^{+}20, VBB20, WQGS25, WW21a, WJN^{+}25, WMT^{+}20, WMSW22, YLX^{+}24, ZJY^{+}24, dAdSR^{+}20, vP20, GVK^{+}23]. analytic [BCOBB^{+}23]. analytical [Cop24]. Analyzing [SBJ^{+}20]. Anarhichas [GKM^{+}23]. Anatomical [LPG^{+}24]. anchoita [MMG^{+}24, DBL^{+}25]. anchoveta [MVLC^{+}20]. anchovetas [GPASM22]. anchovies [dHJTCE23]. anchovy [AUC25, BGLP21, CGM^{+}22, HLI^{+}20, LIA25, OCdMC24, PAY23, DBL^{+}25, TTYT24, VMI21]. ancylodon [AdABW^{+}22]. Andaman [CDAK23]. Angle [KSS^{+}22]. Angle-dependent [KSS^{+}22]. angled [AHL20, BLC^{+}22, BKR^{+}22, CGB^{+}22, CCC^{+}21, RPL^{+}24, TCL^{+}21]. Angler [FFG21, LYLC21, RTHB25, SGW^{+}22, ASB^{+}24, ALRA20, BHB^{+}22, BR22, BN21, BAC^{+}22, CMA^{+}22, CBN^{+}21, GAB^{+}22b, HJA^{+}21, HCDF24, KKCP20, KPK^{+}23, KMA22, LVP22, LWH^{+}23b, MMC^{+}24, PJOR20, SLF23, SKBA23, SCS25, TCL^{+}24, YTSS22, YiTM23, vdHBBR20, vdHR23]. angler-access [KKCP20]. angler-trips [HCDF24]. anglerfish [DCK^{+}22]. anglers [ALW^{+}21, BKM^{+}23c, FBR^{+}24, KWMA23, LOFS22, LTE^{+}23, MCC20, SLF23, SPE^{+}23]. angles [WKSF20]. Angling [DGMG^{+}22, AGNS^{+}21, BLE^{+}22, BALBC23, BAC^{+}22, CH21a, CGSL22, CBR^{+}23, HJMS20, HEG^{+}23b, JDP22, KM23, KC22, LPG^{+}24, Lya20a, SGW^{+}22, SFJ^{+}23, SL24, SCS25, Žák21, vdHC20]. Angolan [MPV^{+}24]. Anguilla [BÖN20, CGT^{+}23, DKBF23, LYLC21, OCBJG20, PRWK20, PMC^{+}24]. animals [MMBH23]. annual [AAPG21, AGS20, CGM^{+}22, DHC^{+}20, ECK^{+}21, HPPT24, LM22, MDMS21, MKS^{+}22, PZL^{+}23, PRRR23, SKD^{+}20, TBF^{+}21]. anomaly [AFR^{+}24]. Anoplopoma [CGC24]. answering [SBC^{+}23]. Antarctic [CZ25, CPB^{+}21, GMPD23, MZ24, ZSDZ24, ZZ22]. Antarctica [ZSDZ24]. anthropized [AMdC^{+}20]. anthropogenic [HSJ^{+}24, NA22, ZCM^{+}23]. anti [HBE^{+}22]. anti-theft [HBE^{+}22]. Antimora [KFO20]. Aomori [KM23]. Aotearoa [Mac25, RGP^{+}23]. AOTTP [APGG22, PGAG22]. Aplodinotus [CH21a]. apparent [AJB20]. Appendage [BHG^{+}24a]. Applicability [KM23]. Application [BTML20, CDAK23, DJFU20, HWMVM23, KBPS21, KBPS22, LK25, MPM22, MDMS21, NS23, ÖRS^{+}25, SV23, VCG^{+}23, ACP^{+}23, BCG^{+}25, CMP20, CMD^{+}23, CAAFH21, CH22, DHC^{+}20, DBGV^{+}22, GVK^{+}23, HC22a, HIMP23, HL20, HCDB22, HBMC21, MCS^{+}22, PDA^{+}24, SIM^{+}24a, SIM^{+}24b, SDBS21, WWF^{+}20, WSB22, YRTP20]. application-based [CH22]. applications [WST^{+}23]. applied [BJS^{+}22b, CGT^{+}23, KHK^{+}20, LCW^{+}23a, SLA24, SBBH24]. Applying [CRCAF^{+}22, HF20, LBD23, SBRM^{+}22, WH23, HIKM21, LGD^{+}20]. apportionment [DQMV21]. Approach [RMNB^{+}21, AHB^{+}22, AAZ20, BVR^{+}21, BaLK^{+}21, BSAP22, BHG^{+}24b, CPL^{+}25, Det21a, DSP22, DHCS23, FDB^{+}20, FBPCC^{+}21, GAB^{+}22a, GAW^{+}22, HPL^{+}24, HGHH25, HEGR24, HRH22, HCK^{+}21, JMS25, JCL^{+}21, KMA23, KBPS21, KBPS22, LNW20, LdV22, MCS^{+}24, MKS^{+}21a, MZSZVP^{+}23, NY23, NS23, NAV^{+}23, OGFW24, PF20, PPD^{+}25, PAY23, QMC^{+}22, SMA^{+}24, SJW^{+}22, SCC^{+}22, TSC^{+}22, TDI^{+}21, VCG^{+}23, VP22]. Approaches [HWMVM23, ASJ^{+}20, AVCA22, BMJ^{+}24, Bro24, CHGC25, FS20, HKG^{+}21, PBRT22, PJP20, QSMG^{+}23, SDV^{+}22, SXK^{+}24]. approaching [WGW^{+}23]. appropriate [Mac25]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a]. aquaculture [Apo25]. aquatic [FKS^{+}20]. arabesque [OMK24]. Arabian [AAM^{+}20, SSV^{+}20]. arapaima [WS20]. archipelagic [HHD^{+}20]. Archipelago [SAA23]. archival [NGDC25, NEBP^{+}23, OOM^{+}23, RNP^{+}24]. archived [DOB^{+}24]. Arctic [HP23, LvCdGSL23, PD25, SKD^{+}20, TEO25]. Arctocephalus [IAB20]. Area [GGMMMV^{+}20, RAE^{+}21, BAW^{+}24, BWR24, CTCB22, FHE^{+}24, HPL^{+}24, JAN23, LMT^{+}22, MFM^{+}20, MPP20, NOL23, OOAF^{+}21, PK24, TdL24, YMS21]. area-based [PK24]. areas [AVB^{+}23, AAFLL^{+}25, CQA^{+}24, DSP22, HPL^{+}24, LLS23, LRW^{+}24, NMS^{+}22, PIP^{+}22, SSS^{+}23, Spa24, YTSS22, BSA^{+}23, RGP^{+}23, SHS20]. arenaria [LCM^{+}23]. Argentina [AdlBCN25, DDA^{+}20, TLAM25]. Argentine [MDL^{+}21, PCF23, DBL^{+}25]. argus [BBR^{+}22, BBSM24, HMR^{+}24b, HMR24a, RBHM24a, RBHM24b]. Argyrosomus [BM24b, BAC^{+}22, HMS^{+}22, RTB^{+}21]. ariid [VBL^{+}24]. ARIS [JGU21]. Aristeidae [PCGG20]. Aristeus [PCGG20]. ark [TC24, ZJY^{+}24]. armatus [BJK24, CBTH20, JYH21, MHD^{+}21]. Arousa [OOAF^{+}21]. array [BWG^{+}21, CBD^{+}22, KWE^{+}21]. arrow [LPP^{+}20, MdCG20]. arrowtooth [Dra22]. Artemesia [SS24]. artificial [BTML20, BTFL22, BLFT23, BBM^{+}24, BHNP22, FGCB^{+}21, HPL^{+}24, JLYR24, LWH^{+}21, MOA23, NZP^{+}21, SKY^{+}24, SEM^{+}23a, SSG^{+}22, SSJ^{+}21, TCL^{+}21, WHR^{+}24, WASS20, WPB22]. artificially [SSD^{+}20b]. Artisanal [TTK^{+}25, AMM^{+}22, AdlBCN25, CTCB22, DEM^{+}23, MA20, MSD21, MRG^{+}23, NCB^{+}23, OPL21, OOAF^{+}21, PCCMOA^{+}24, PFdSBL25, SKST23, SPC^{+}25, dCHdMS^{+}23]. Ascaridoid [DCL^{+}20]. aspects [NK23]. aspera [Ten22]. assemblage [TKB^{+}21]. assemblages [CLB^{+}21, LYX^{+}21]. assertions [CBHS24]. assess [BWG^{+}21, BGCCP22, CSH^{+}21, CB20, DJFU20, EMJ^{+}22, EM23, GW21, HF20, LK25, MCPJET^{+}20, PD25, RBD^{+}22]. assessed [ECY21, MNPMM^{+}22]. Assessing [BBC^{+}25, BRGB^{+}23, BM24a, CSSB22, CC20, FM21, FCSA21, GC21, Hut22, LCM^{+}23, LMK^{+}23, LRGB25, MBOCdAM24, PGAG22, PRCF22, PDJ24, PSS^{+}21, RdBAT^{+}23, SMA^{+}24, SMKJ21, TSPK24, WKBMW24, BBHF25, LCW23b, OGFW24, PDD^{+}22, RBH^{+}24, SS23, SHS20, ZCM^{+}23]. Assessment [ACL^{+}20, BW20, BJS^{+}22a, Det23, FMLC^{+}22, LCG^{+}21, MCS^{+}25, OOM^{+}23, SM21, SKJJ25, WLZ^{+}21, WGC^{+}19, WGC^{+}21, WBD^{+}21, XWD^{+}21, AAM^{+}20, AT20, AAT^{+}21, AAZ20, BJS^{+}22b, BBG^{+}24, BMA^{+}24, Bro24, BAC^{+}22, CAZN24, CWRR24, Cad20, CMP20, CGBJ23, CC22, CPF20, CTIC23, CGC24, Cla22, CHM24, Cop24, CRCAF^{+}22, CFP21, DDDP21, FMD^{+}24, FDS^{+}23, FBALRR^{+}22, GVK^{+}23, GAW^{+}22, GML^{+}23, HIMP23, HC22b, HIKM21, HEG^{+}23a, HFKS20, HMP^{+}22, HCDB^{+}24, HW24, JC21, JMP^{+}21, KNO^{+}21, KNS^{+}22, KSL^{+}23, LDM^{+}24, LBD23, LBD24, MSD21, MTX^{+}20, MHL^{+}23a, MPSM25, MJC^{+}23, MOA23, MMP^{+}24, MRG^{+}23, MVDH24, MCH21, OAB^{+}23, OYOO21, PC21, PDE^{+}20, PCJH^{+}21, PTD^{+}21, Pun23, Pun24, RRH^{+}24, RRB24, Sat23, SMK^{+}24, SSP^{+}23, SRF^{+}24, SXM^{+}21, TAA^{+}20, Tho19, Tho21, TMSS^{+}23, TMDA22, TAK^{+}23, WWF^{+}20, WASS20, XMCC20, dLH23, vPDD^{+}25]. assessments [AYTM21, Bea21, CWC^{+}21, CHGC25, CHAY^{+}25, DDDP21, DMF^{+}21, DBV22, DBV23, FSP22, FCSA21, Fis25, GBC23a, HL20, HWMVM23, JDH22, KSO^{+}21, KDBOC25, LMP24, MLS^{+}21, MUF^{+}22, MVMdS^{+}24, Mon24, PPC^{+}23b, PJP20, PJMP22, PTD^{+}20, PTL^{+}24, RWB^{+}23, SECB21, SBZ^{+}21, TDJ^{+}21, WH23, WS24]. assignment [KPS20]. assimilation [KMA23]. assists [PNRS23]. associated [BWG^{+}21, BGCCP22, BPT^{+}20, HGC^{+}21, HdLHD22, HA23, MTX^{+}20, ML24, Mun24, PNAPH24, TRS^{+}24, WPB22, ZZC^{+}21]. Association [TFC^{+}20, SYZ^{+}25]. associations [BW20, FMMA20]. assumed [Žák21]. Assumptions [VP23, CFO23, SSP24, TGG^{+}24, TCJ^{+}21]. asymmetry [FGL22]. asymptotic [BHVB^{+}24]. asynchrony [CLB^{+}21]. at-vessel [BM23a]. Atheresthes [Dra22]. Atlantic [ABT^{+}24, AAR^{+}21, BNTK23, BFA^{+}21, CBJ24, CVM^{+}20, dSCCC^{+}22, IAB20, íKMíH^{+}22, MCC^{+}23, OBB^{+}20, PGAG22, SSP^{+}22, dAdCdO^{+}23, AdABW^{+}22, AAG22, AMM^{+}22, AGB^{+}20, AGS20, BGG^{+}22, BPT^{+}25, BÁP^{+}23, BHST^{+}21, BWB^{+}23, BUG^{+}24, BLHS20, CWRR24, CRL21, CKM^{+}20, CHT20, CSH^{+}21, CMTP^{+}21, CG21, CRF^{+}24, FMMA20, GGG^{+}22, GKC^{+}22, GC21, GKM^{+}23, HCCC21, HBC^{+}22, HM25, HC22b, HWW22, HOQ20, HUK^{+}23, IKBL23, KWC^{+}20, KRH^{+}24, íKMíH^{+}22, LMT^{+}22, LNW20, MSC^{+}24, MHL^{+}23b, MPC^{+}20, MMM^{+}20, NSM^{+}21, NEBP^{+}23, NFdSJO25, NBF20, NDRR20, ORdlG^{+}24, PMS^{+}23, PPM^{+}23, PBDM21, QMC^{+}22, RMRG22, RS21, RBG^{+}20, RSPE22, SS24, SSRC24, SKST23, SFMA23, SPG^{+}21, SCD^{+}22, SAA^{+}22, SBL^{+}23, SDC^{+}22, SGZD20, SGZD21, SGZH^{+}22, SBRM^{+}22, SWH24b, SXM^{+}21, SWIRF21, SMLT24, TFC^{+}20, TMDA22, UPBH^{+}20, WAA^{+}22, WCN^{+}24, WMT^{+}20]. atlanticus [BCOBB^{+}23, ECK^{+}21]. atricauda [AdABW^{+}22]. attachments [BWB^{+}23]. attempts [VRS^{+}22]. Attitudes [YMS21, FMLC^{+}22, SKBA23, vdHBBR20]. attract [GLP^{+}20]. attractant [SSG^{+}22]. attractants [UPBH^{+}20]. attractive [APGG22]. attractiveness [GJGW23]. attractor [AMHH21]. auction [HBW21, SGZD21]. augmentation [MCS^{+}24]. augments [CKK^{+}20]. August [Ano20b, Ano21b, Ano22b, Ano23b, Ano24b]. aurata [vBMP^{+}23]. auratus [TP24]. aurorubens [CSB^{+}23]. Australia [AGL^{+}24, BLFT23, BBM^{+}24, BBPT^{+}22, CSTdL25, CTR^{+}21, HSPC21, HdLHD22, HMS^{+}22, KBH22, KCB^{+}24, LRMH21, LZY24b, MSD21, MPSH25, MWTH23, ML24, PHP^{+}20, PBM^{+}23, SJH^{+}23, SSP^{+}23, SRF^{+}24, WCLN20]. Australian [KBPS22, BBJ22, BJK24, BM22, BM23b, BM23a, BKM23a, BKM23b, BM24a, BM24b, Bro25, BM25a, BM25b, BM25c, BCOBB^{+}23, HTSJ23, HPS^{+}24, JYH21, KBPS21, ODM20, PPC23a, vPDD^{+}25]. australis [BMA^{+}20, BMA^{+}20, IAB20, KBPS21, KBPS22, NHE^{+}23]. authentic [FNH^{+}21]. autochthonous [SSD^{+}20b]. autocorrelation [LJX^{+}20, SXM^{+}21]. automated [HHJ^{+}22, HL21, KD22]. Automatic [PÁEMC22, AFR^{+}24, LCLM23]. Automating [PPC^{+}21]. autonomic [JWL^{+}24]. autonomous [BWG^{+}21, WSUN^{+}23]. Autoregressive [PAY23, CBA^{+}24]. autumn [RMRG22]. availability [JRW^{+}21, KTFY22, KLS20, LPG^{+}24]. available [GRJW20]. Aveiro [MGB24]. average [CLM^{+}22, DET21b, Lya20a, SR20]. avoidance [CWM^{+}23, DKD^{+}21, ERS^{+}23]. Avoiding [KLBHK23]. awareness [GBC23a, RBG^{+}24]. away [BAM^{+}24]. ayraud [GFM^{+}23]. Azores [TMDA22].

Back [HPPT24, ECK^{+}21, SBRM^{+}22, SWH24b]. back-calculated [SBRM^{+}22, SWH24b]. back-calculation [ECK^{+}21]. backscatter [EBGE21, PGD^{+}25, TPD20, WPB22]. Bahamas [SGD^{+}21, TSC^{+}20]. Bahia [MPEBdR23, MFO21]. bairdi [HSSD^{+}21, Mur20, Mur21, NGDC25]. bait [AK23, BBC^{+}25, BUG^{+}24, FBB20, LNMA23, LSM^{+}23, LRGB25, NMJ^{+}24, SSRC24, SMKJ21]. baited [CSSB22]. baits [CBTH20, DMNV^{+}24, GJGW23, MFJ^{+}24]. Baja [VAVQGD^{+}20]. balanced [CZJ^{+}24, NNS^{+}22]. Balaton [SSJ^{+}21]. Balearic [SBT^{+}20]. Balistes [CPR^{+}24, PRRR23]. balloti [CLY^{+}22]. Baltic [FDS^{+}23, ÖSL^{+}23, ALW^{+}21, ARD^{+}23, BDR^{+}20, BLE^{+}22, BLK23, CCKL^{+}20, DAR^{+}23, DSNK^{+}22, DKBF23, EBN^{+}23, EHE^{+}23, EM23, EFM25, FGSD25, FTH^{+}23, GKKL24, HMM^{+}21, HML^{+}20, KLBHK23, KMA22, KHK^{+}20, KZT^{+}23, LWH^{+}23a, LNvD^{+}23, MINS21, MHH^{+}20, NVB^{+}23, OAB^{+}23, PNAPH24, PZG23, RDR^{+}23, RNKB23, RSPE22, SPC^{+}23, SKBA23, SF20, SMC^{+}24, SAA23, WBBG^{+}23]. bamboo [FTB^{+}21]. ban [FGPVPPGG22, PVC^{+}22, ZJJ^{+}25]. banana [HPD^{+}22, PHP^{+}20]. banded [FTB^{+}21]. bands [MB25]. Bangladesh [FM21, HS21, MRS^{+}25, ZHM23]. Bank [CWRR24, EWPB22, KPK^{+}23, BAW^{+}24, WGFM21, WCN^{+}24]. bank- [KPK^{+}23]. barbed [CCC^{+}22a]. barbel [RTHB25]. barbless [CCC^{+}22a]. barbs [CLD^{+}22]. Barbus [RTHB25, AMdC^{+}20]. barcoding [FMSA21, LNR^{+}21, PKRL21, RLQ^{+}20, WQGS25]. Barents [ASGG21, AIJ^{+}23, CAZN24, IB20, JBL^{+}22, JFS21, SV23, TRWH23]. barium [TLCD21]. barotrauma [SACS23]. barrier [HRC23]. bartramii [WHCF22]. Based [EYAO20, HF20, HFKS20, AAVÁM23, AMHR22, AT20, AUC25, AAZ20, AIW^{+}23, ASJ^{+}22, Bea21, BJS^{+}22a, BJS^{+}22b, BKT^{+}20, BWOR23, CBA^{+}24, CPM21, CSH^{+}21, CMIMS20, CFO23, CAAFH21, CH22, Cop24, dSCCC^{+}22, CB20, CRCAF^{+}22, CHAY^{+}25, DTSR22, DSP22, DMM^{+}21, DET21b, DMF^{+}21, DWS^{+}23, DHC^{+}20, DMS22, FOS^{+}21, FDS^{+}23, FAK24, Fu22, GFM^{+}23, HC22a, HPL^{+}24, HC22b, HH20, HTK^{+}24, HZZ^{+}20, HLZ^{+}20, İTAD24, JRW^{+}21, JFS21, KSV^{+}22, KSL^{+}23, KK22, KDF^{+}25, LVL25, LRMH21, LD25, LSZ^{+}23, LLC^{+}20, LCW^{+}23a, LCL25, LdV22, MCS^{+}24, Mac25, MPM22, MdCG20, MTX^{+}20, MKS^{+}21a, MZ24, MDS^{+}20, MRP^{+}23, NSM^{+}21, NLL^{+}25, NVRG^{+}21, OAM^{+}21, PPD^{+}25, PK24, PTD^{+}20, QMC^{+}22, RFMS^{+}21, RWFT25, RRH^{+}24, SB20, SLW^{+}21, SYZ^{+}25, SDV^{+}22, SXK^{+}24, SBZ^{+}21, SML^{+}24, SRB^{+}25, TSC^{+}22, TCJ^{+}21, TMSS^{+}23, VCG^{+}23, WZL^{+}22, WCN^{+}24, ZJY^{+}24]. Based [CBHS24, ZF21]. baseline [MHH^{+}20]. Basic [BČD^{+}21]. Basin [ACP^{+}23, FDdCS^{+}20, SCG^{+}24, TTC^{+}25, BQBW20, BKC21, MA20, PdAMdM^{+}23, ZJJ^{+}25]. basin-scale [SCG^{+}24]. basis [OB21]. basket [aLBK^{+}21]. Basque [CAH^{+}20]. Bass [OGFW24, CBN^{+}21, CG21, GSH22, HWW22, KSI20a, LCB^{+}21b, LRW^{+}24, MSW21, PZL^{+}23, RBS^{+}24, WMT^{+}20, CLD^{+}22]. bat [GRHHGM^{+}20]. bather [BC20, LSM^{+}23]. bather-protection [BC20]. bathymetric [BNL^{+}23, TYYK21]. batoids [MWTH23]. batten [MMML24]. Bay [DWLT21, HPL^{+}24, MDW^{+}21, MWJ^{+}24, MPEBdR23, RMRG22, SCGW24, HSM^{+}25, NGDC25, SYZ^{+}25, XCB^{+}21, AUC25, BMM^{+}24, BGLP21, BUG^{+}24, CFB^{+}23, CRCAF^{+}22, DSP^{+}23, FGSD25, Fre22, HS21, LMG^{+}24, PVA^{+}24, PCC^{+}23, RLO^{+}21, SSFL24]. Bayesian [KBPS22, BP20, CSB^{+}23, DWLT21, DHCS23, GPW^{+}20, GJSW22, KBPS21, KBH22, Mon24, NFAL^{+}22, OOAF^{+}21, SKW^{+}21, SYZ^{+}25, SFPR^{+}23, SH22b, TCL^{+}24, YIM^{+}20]. bays [BAF23, FTH^{+}23]. be [AWC^{+}23, BČD^{+}21, BAJB^{+}24, GSS^{+}23, MKC20, RNKB23, SMK^{+}24]. beak [ALRB^{+}20]. beaked [HP23, PVA^{+}24, PCC^{+}23]. beaks [FHSC21, WHCF22]. beam [UAB^{+}21, UAG^{+}23, VMFF^{+}20, vORP23]. beds [NHE^{+}23]. before [ZJJ^{+}25]. beginning [VMI21]. Behavior [AUM21, CMRP20, WW21b, BSKL^{+}22, DQK^{+}23, HL21, HJMS20, JDP22, JCL^{+}21, KPS20, KFH^{+}25, LP23, MZ24, MCC20, ÖRS^{+}25, RWT^{+}20, SGW^{+}22, WGW^{+}23, YSB^{+}21]. behavior-induced [DQK^{+}23]. Behavioral [JGG^{+}24, SSG^{+}22, CBD^{+}22, GAB^{+}22a, HGC^{+}21, KKCP20, LRGB25, RNP^{+}24, SGW^{+}20, SOS^{+}21]. behaviors [KLS20, LYLC21]. Behaviour [ŠBB^{+}22, AWC^{+}23, ALRA20, BAC^{+}22, GHAZ21, LCB^{+}21b, MRE^{+}24, TKB^{+}21, YiTM23, dLHER24]. Behavioural [GRG24, PAA^{+}24, CGB^{+}22]. behaviours [FMLC^{+}22]. behind [CFO23]. Beibu [SZXC25]. Belgian [UAB^{+}21, UAG^{+}23]. beliefs [SKBA23]. bellicosus [VDCMRF^{+}21]. Belo [dAdSR^{+}20]. benchmarks [Mun24]. beneath [TKB^{+}21]. benefit [GdSPL21, HSW25]. Benefits [AVCA22, FHHH20, Sch23, SR20, Hut22]. Bengal [HS21, DSP^{+}23]. Benguela [KNS^{+}22, WJN^{+}25]. benthic [CMV21, FHSC21, FOM21, GRG24, LCB^{+}21a, MMBH23, NVB^{+}23, TRS^{+}24]. Benthosema [KHGB25]. bentincki [MVLC^{+}20]. berglax [BHH21]. Bering [SFYM24, EYAO20, FJS^{+}25, HSSD^{+}21, KDF^{+}25, Mur20, Mur21, PDD^{+}22, PDA^{+}24, RL24, SC20, Szu22, Ten22]. Beringraja [HTK^{+}24]. bertheloti [GMERCM^{+}24]. Best [CGBJ23, HEG^{+}23b, BR22, PJMP22]. Bet [PvZ22]. Bet-hedging [PvZ22]. better [BMOC22, MKS^{+}22, MBB^{+}23, SF20, TLV23]. between [AYTM21, AMM^{+}22, ALRA20, AVCA22, Ash20a, Ash20b, ÁHGCVAI22, BW20, BQGV^{+}24, BBM^{+}24, BPT^{+}20, CHT20, CHGC25, DSP^{+}23, DKBF23, DHCS23, FGTA24, aFLpX^{+}21, FNKY20, HLMV24, HSPC21, HMR24a, KTFY22, KKCP20, KFI^{+}21, LHPR21, NOL23, NAS^{+}20, NTJN21, OGFW24, PASdCF23, RB22, RTB^{+}21, SLF23, SCG^{+}24, SYZ^{+}25, SOS^{+}21, SBL^{+}23, SFYM24, SSD^{+}20b, SCC^{+}22, TSC^{+}20, WWF^{+}20, YIM^{+}20, YMYH20, YiTM23, ZZ22, vDJB^{+}23, vdHC20]. Beyond [TDJ^{+}21]. bi [WKSF20]. bi-phasic [WKSF20]. Bias [WBA23, Det23, DLP^{+}24, DQMV21, FS20, KNO^{+}21, MMQ21, SS23, Sat23]. biases [LWH^{+}23b]. Bibliometric [SO25, NS23]. Big [KLS20, DPL20, EPHDB24, MTC^{+}22, SKS^{+}23]. Bigeye [AWVS21, LCDM^{+}24, LK25, MLCMdS23, SXMV^{+}21, TEO25, WAA^{+}22, WSL21, XMLCMV24, RCH^{+}21]. bighead [NTJN21]. Bight [CG21, KHE^{+}22, RWT^{+}20, SSS^{+}23]. bigmouth [ASDW24]. billfish [AGB^{+}24]. billfishes [CNDDAPMR21]. Bimini [SGD^{+}21]. bin [WZX^{+}20]. Bio [GTS^{+}21, BDR^{+}20]. bio-economic [BDR^{+}20]. Bio-economical [GTS^{+}21]. biochemical [GBO^{+}20, MCHA21, SWIRF21]. biodegradable [LNMA23]. biodiversity [LFdB^{+}21, NMS^{+}22, RMD^{+}25]. Bioeconomic [MZSZVP^{+}23, BaLK^{+}21, DQK^{+}23, Fre22, HPD^{+}22]. bioelectrical [CHPT20]. Biogeographic [SS24]. bioinvaded [MCK23]. Biological [AMSC20, MOM^{+}25, MBZSM20, Aks24, CLB^{+}21, DTSR22, DBV22, DBV23, HSM^{+}24, MBOCdAM24, MOI23, MZSZVP^{+}23, MCHA21, PRK23, RCVGMI22, SH22b, XSS^{+}23]. Biology [CSTdL25, Lor22, AHEV24, DD25, KSI^{+}22, KCB^{+}24, MCC^{+}23, PBPM^{+}23, PDC^{+}24, RAG22, SF22a]. biomarkers [HLG^{+}21]. Biomass [VTSI^{+}24, ALW^{+}21, BCPH22, CQA^{+}24, DLP^{+}24, FJA^{+}20, JBČ^{+}22, KTFY22, KHPB20, MSS^{+}21, MRP^{+}23, MMBH23, PDC^{+}23, PIP^{+}22, PTK^{+}20, PNGGO^{+}22, RWB^{+}23, dAGCR21]. biometric [SWH^{+}24a]. Biphasic [CMD^{+}23]. bird [WEH^{+}25]. Biscay [BMM^{+}24, AUC25, BGLP21, BUG^{+}24, CRCAF^{+}22, PVA^{+}24, PCC^{+}23, RLO^{+}21]. bite [LTE^{+}23]. bite-shortened [LTE^{+}23]. bites [CMA^{+}22]. Bitter [FJS^{+}25]. bivalves [dAGCR21]. black [CG21, HBMC21, PZL^{+}23, BGK22, ŞGK^{+}20]. blackfoot [NAV^{+}23]. blacknose [KFDE^{+}22]. blackspot [CRF^{+}24]. blacktip [BRGB^{+}23]. blacodes [BÁP^{+}23]. bleeding [LPG^{+}24, SGW^{+}22, UVA20]. Bloch [ORdlG^{+}24, SKST23]. blocked [vP20]. blonde [TLV23]. Blondes [TLV23]. Blood [FJA^{+}20, BKR^{+}22, BAC^{+}22]. blooms [Béc20]. Blue [SCGW24, SSG^{+}22, BJK24, BMA^{+}20, DMNV^{+}24, GPW^{+}20, JYH21, LZY24b, MHD^{+}21, MCC^{+}22, MCC^{+}23, MABR^{+}20, NFC20, PDC^{+}24, PBM^{+}23, SSFL24, SCD^{+}22, SKS^{+}23, WCGB22, WWO20]. blue-swimmer [BJK24]. bluefin [AGS20, ASO^{+}22, AIW^{+}23, BUG^{+}24, HBC^{+}22, KWC^{+}20, MCS^{+}24, NBF20, ORdlG^{+}24, RBG^{+}20, SPG^{+}21, SAW^{+}20, SBRM^{+}22, SWH24b]. Bluegill [HGC^{+}21]. Board [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25a, Ano25b, Ano25c]. boards [WGNM24]. boarfish [CSRL20]. boat [BM23a, BM25c, KPK^{+}23, LRMH21, VP22, ZHC^{+}25]. boat-angler [KPK^{+}23]. boat-based [LRMH21]. boat-fishing [VP22]. boating [AYTM21]. boats [KWW^{+}21]. bodied [KCB^{+}24, LTE^{+}23, WGC^{+}19, WGC^{+}21]. Body [DAR^{+}23, BHVB^{+}24, BSR^{+}22, FSP22, LCG22, LCHB^{+}24, Lya20a, SK22, SFYM24, SWIRF21, TP24]. BOFFFFs [EPHDB24]. bogaraveo [CRF^{+}24]. Bohai [DSJG20, HPL^{+}24, LSZ^{+}23]. boldness [KLS20]. Bomb [BKHA21, CPR^{+}24, CSB^{+}23]. Bomb-produced [BKHA21]. bone [FCKG^{+}22]. BONGO [FSP22]. bonito [ORdlG^{+}24, SKST23]. Boom [dLHER24]. bootstrap [HW24, MB25]. bootstrapped [SFMA23]. boreal [SHV20, VTSI^{+}24]. borealis [GC21, HKCW24, IJS^{+}22]. both [SDBS21]. Bottom [NMJ^{+}24, BMM^{+}24, BWR24, BLHS20, Bro25, CYBW22, DML^{+}20, GDVBB^{+}20, GGMRC^{+}22, HEGR24, HEG^{+}23a, IAB20, JLYR24, JBL^{+}22, JMP^{+}21, LWH^{+}21, MSC^{+}24, NBEI23, PTK^{+}20, PFGQ20, PCC^{+}23, RBM21, RNP^{+}24, TMN^{+}21, TTL^{+}20, TZL^{+}24]. bottom-opening [Bro25]. bottom-set [MSC^{+}24]. bottom-trawl [CYBW22, GDVBB^{+}20]. bottom-trawling [IAB20]. bottomfish [LG21]. boulders [HVMEO25]. boundaries [CGBJ23]. boweni [DOB^{+}24]. brachysoma [KPS20, KPWS21]. brachyura [TLV23]. brackish [ALW^{+}21, DAR^{+}23, RDR^{+}23, SPC^{+}23, WBBG^{+}23]. Brakoniecki [RdLSdB^{+}21]. Brasil [FDdCS^{+}20]. brasilianus [dAdCdO^{+}23]. brasiliensis [ACP^{+}23, PFFdC22, SSP^{+}22, SSS^{+}23, SDC^{+}22]. Brazil [Mac22, MFO21, MdCG20, MPEBdR23, MFR^{+}22b, NFAL^{+}22, SCHSC21, WCC24, dAdSR^{+}20, dS21]. Brazilian [SSS^{+}23, ASD^{+}22, CdSLP21, dSRFFN^{+}20, RdLSdB^{+}21, SdOR^{+}23, SSP^{+}22, SSS^{+}23, SdFZFJ21]. break [BQGV^{+}24]. breathing [AMM^{+}22]. Brexit [FGPVPPGG22]. bridge [ZCM^{+}23]. Bridging [GBWM22, LZY^{+}24c, SCG^{+}24]. brief [DD25]. brink [HPPT24]. Bristol [MHL^{+}23b]. British [RGN^{+}20, ZF21, BJS^{+}22a, BJS^{+}22b, MOB^{+}22]. broad [DHCS23]. broad-scale [DHCS23]. Broadband [WYM^{+}25, BPT^{+}20, DJFF23, KMO20, KBM23]. broodstocks [BJS^{+}22a, BJS^{+}22b]. broughtonii [ZJY^{+}24]. Brown [GTS^{+}21, MDS^{+}20, AGB^{+}20, CBJ24, FTB^{+}21, JAN23, MDJP24, ÖRS^{+}25, TBH^{+}22, VMFF^{+}20, ZDZR^{+}22]. brown-banded [FTB^{+}21]. Brownie [Fu22]. Brunswick [LCM^{+}23]. Buccinum [CB20, EHBJ20]. budget [HGC24]. Buenos [DDA^{+}20]. buffalo [ASDW24]. buffering [WXJ^{+}24]. buffers [HSSD^{+}21]. bug [LMG^{+}24]. Building [RRH^{+}24]. bull [GLP^{+}20, JSKM20]. bullet [BSAP22]. buoy [AWVS21, SH22b]. buoys [BDA^{+}20]. Burn [VK21a, VK21b]. burning [VK21a, VK21b]. business [JvPOG25]. businesses [Béc20]. Busycotypus [MFJ^{+}24, WPGO21]. buy [DDA^{+}20]. buyers [WSB24]. buyouts [MFR22a]. by-catch [SME^{+}24]. by-products [UPBH^{+}20]. Bycatch [AAM^{+}20, FMLSP20, MHL^{+}23b, PCK23, PCC^{+}23, AL22, BTR^{+}24, BWB^{+}23, BGCCP22, BM23b, BM24b, CJC^{+}20, CCKL^{+}20, CC20, FM21, FMD^{+}24, FDB^{+}20, GP25, GGTÁVT^{+}20, GM21, GJSW22, GMPD23, JLYR24, KBB^{+}21, KSS^{+}22, KMC20, LWH^{+}21, MB23, NLW^{+}22, OB21, VS23, PVA^{+}24, PFGQ20, PCF23, PSS^{+}21, RBHM24a, RB22, RL24, SSRC24, SOS^{+}21, SJBT20, SBB^{+}24, SMLT24, TFC^{+}20, TJS23, YSB^{+}21, YLS^{+}23]. bycatches [ÖSL^{+}23]. bycaught [CNDDAPMR21, DDCNMR21].

C [CPR^{+}24]. C&R [FNH^{+}21]. C. [Mur20, Mur21]. Côte [AAPG21]. Ca [ZOS^{+}23]. Cadiz [SRT^{+}20, GGMRC^{+}22]. cage [KDdOM^{+}22]. caged [WKSF20]. calcein [DH20]. calcium [LHH^{+}25]. calculated [SBRM^{+}22, SWH24b]. Calculating [TAK^{+}23, Det21a]. calculation [ECK^{+}21]. calculations [KSO^{+}21]. Caledonia [RFF^{+}22]. California [KHE^{+}22, AVACA^{+}23, DLP^{+}24, GGTÁVT^{+}20, LP23, MCS^{+}22, MSLMOC^{+}24, RWT^{+}20, SAW^{+}20, VAVQGD^{+}20]. californica [GRHHGM^{+}20]. californicus [MCS^{+}22]. call [CTCB22]. Callinectes [DMNV^{+}24, GZLCRG25, VDCMRF^{+}21, WCGB22]. Cambodia [CAGLT23]. camera [AYTM21, BWG^{+}21, BTML20, BWOR23, GAB^{+}22a, WGW^{+}23]. camera-based [BWOR23]. cameras [FBALRR^{+}22, RWT^{+}20, SEM^{+}23a, TRWH23, WSB22]. caml [MPP22]. Campaign [Mel23]. campechanus [BPT^{+}25, FGCB^{+}21, GAB^{+}22a, PJSQ20, SACS23, WSB22]. camtschaticus [AAH^{+}23, LYH^{+}21, MSJ21]. Can [AWC^{+}23, BMA^{+}20, CPM21, DQK^{+}23, ELM20, GSS^{+}23, JSKM20, SDdMG^{+}20, VKS^{+}25, BMOC22, BČD^{+}21, BAJB^{+}24, CMIMS20, MKC20]. Canada [ASWS^{+}21, BMSM22, BAM^{+}24, BJS^{+}22a, BJS^{+}22b, BH23, CRS23, FRP22, HJA^{+}21, LCM^{+}23, MOB^{+}22, WBD^{+}21]. Canadian [Bea21]. canaliculatus [MFJ^{+}24, WPGO21]. Cancer [CBJ24, GC21, HKCW24, MDS^{+}20, MDJP24, RPH20]. candidate [HLG^{+}21, RWB^{+}23]. cannibalism [BCRI21]. capabilities [LJB^{+}24]. capability [BJHS^{+}23]. capacity [CEAL21, KM23, LCL25, Mau22, SPC22]. CAPAM [MCS^{+}25, MPSM25]. Cape [Sch23]. capelin [ASGG21, FJJT^{+}21, FJHT^{+}22, JHB21, JFS21, SBJ^{+}20, SGKAR24]. capensis [KNS^{+}22]. capriscus [CPR^{+}24, PRRR23]. capture [AWVS21, BKM23b, CSSB22, CHM24, CBTH20, DJFU20, EMR^{+}22, GMK23, GFC^{+}22, HPPT24, HSL^{+}22, JND^{+}23, KLS20, MKS^{+}22, NS23, PMS^{+}23, QAS^{+}25, RB22, SFCG^{+}21, ŠBB^{+}22, SBT^{+}20]. capture-recapture [HPPT24]. capture-related [GMK23]. captured [CCC^{+}22a, LTE^{+}23, SAW^{+}20, SK22, SGH^{+}20]. Capturing [PDA^{+}24, RS21]. Caranx [DMM^{+}23, GFDN^{+}22, GFC^{+}22]. carbon [HTK^{+}24]. carbunculus [DOB^{+}24]. Carcharhinus [BRGB^{+}23, GLP^{+}20, KCB^{+}24, KFDE^{+}22, PCBL23, SASB24, SFCG^{+}21, SGH^{+}20]. carcharias [BBM^{+}24]. Carcharodon [BBM^{+}24]. Carcinoscorpius [HSM^{+}25]. Carcinus [FBB20, MOM^{+}25, OMG^{+}23, PFGQ20]. Caribbean [APBN22, AGNR^{+}21, BQGV^{+}24, BF25, CHB24, GAB^{+}24, HMR^{+}24b, HMR24a, NVRG^{+}21, RBHM24a, WSUN^{+}23]. Carlo [Alv21]. Carolina [BFA^{+}21, Sch23]. carp [ASDW24, CWM^{+}23, aFLpX^{+}21, SLW^{+}21, Žák21]. carpio [AK23, CWM^{+}23, Lya20a, NA22]. carps [Lya20a]. carrying [Mau22]. Case [HHD^{+}20, SSP^{+}23, AVMBEB22, AVCA22, ALW^{+}21, BQGV^{+}24, BBPT^{+}24, BSAP22, BCM^{+}21, CCR24, CFB^{+}23, CMIMS20, dSCCC^{+}22, CHL^{+}20, CRF^{+}24, DCK^{+}22, EBN^{+}23, EBdRPC24, FCKG^{+}22, FBM^{+}21, FDdCS^{+}20, GPW^{+}20, GM21, GPT^{+}21, HBW21, HSSD^{+}21, Hut22, HSW25, KC22, KFO20, LZY24b, LWH^{+}23a, LNW20, LKSiI22, MA20, MdCG20, MKC20, MPP20, MBZSM20, OMK24, OYOO21, PIP^{+}22, PHH^{+}23, PSS^{+}21, RKD24, RRSP^{+}24, SFMA23, SPC^{+}25, SKBA23, SGZD21, SGZH^{+}22, SMA^{+}24, SH22b, Szu22, TMN^{+}21, TRWH23, WCN^{+}24, WPB^{+}20, ZCM^{+}23, CAZN24]. cases [PJOR20]. Caspian [BTB^{+}21, RKN23, TAA^{+}20]. caspius [TAA^{+}20]. castable [DNLM23]. Catalan [SCGM^{+}21]. Catalonia [MDMS21]. cataractae [KSI20a]. Catch [CGSL22, GVK^{+}23, HCDB^{+}24, PPM^{+}23, SPE^{+}23, SGW^{+}22, vdHBBR20, AAT^{+}21, AGB^{+}20, AGNS^{+}21, BW20, BLC^{+}21, dFBPL^{+}20, BCS^{+}22, BWR24, BDM^{+}20, BHB^{+}22, BWB^{+}23, BR22, BSR^{+}22, BBC^{+}20, BLHS20, BN21, BALBC23, BKM^{+}23c, BAYR^{+}24, BAC^{+}22, CPM21, CH21a, CSH^{+}21, CCCM^{+}20, DHX^{+}23, DHH^{+}22, DLKH22, DWS^{+}23, DSJG20, DBGV^{+}22, EFM25, FBB20, FJA^{+}20, GFC^{+}22, GAB^{+}22b, HMY25, HSM21, HRC23, HEG^{+}23b, HCDB22, IJS^{+}22, JBL^{+}22, JYH21, JMP^{+}21, KHS^{+}20, KLBHK23, KMA22, KFH^{+}25, KHPB20, KFI^{+}21, LRMH21, LOFS22, LPG^{+}24, LZY24a, LCMS^{+}22, LCDM^{+}24, LWH^{+}23b, LSM^{+}23, LTT^{+}23, MLCMdS23, MTX^{+}20, MMML24, MRE^{+}24, MDC^{+}22b, NY23, NLW^{+}22, OWF^{+}23, PÁEMC22, PF20, PASdCF23, PBDM23, PRCF22, PTL^{+}24, RFMS^{+}21, RBHM24a, RTHB25, RRB24, SKST23, SLF23, SJH^{+}23, Sch23, SBJ^{+}20, SFJ^{+}23, SJ24, SOSK22, SAdC20, SCS25, SEM^{+}23b, SME^{+}24, TDJ^{+}21]. catch [TMP20, TNTN20, TCL^{+}24, UAG^{+}23, VBB20, WSF22, WS24, XMLCMV24, YTSS22, ZZH^{+}24, vORP23]. catch-and [BAC^{+}22]. catch-and-release [AGNS^{+}21, BCS^{+}22, BSR^{+}22, BALBC23, CH21a, GFC^{+}22, HEG^{+}23b, LPG^{+}24, YTSS22]. catch-at-age [PTL^{+}24, TDJ^{+}21, VBB20]. catch-based [DWS^{+}23]. catch-curve [KHPB20]. catch-only [DHX^{+}23, FJA^{+}20]. catch-per-unit [MTX^{+}20]. catch-per-unit-effort [DBGV^{+}22, HCDB22, TMP20, XMLCMV24]. catch-related [EFM25]. catchability [BCS^{+}22, GKC21, GLA^{+}20, HMC^{+}23, JRW^{+}21, MA20, YiTM23]. Catches [ÖSL^{+}23, AIJ^{+}23, APB^{+}20, BWN^{+}21, BGM^{+}23, BS20a, BM25a, CFO23, DEM^{+}23, EHE^{+}23, FMMC20, GMT24, KLNB^{+}24, KPWS21, MRG^{+}23, ML24, MBD^{+}21, NBD^{+}20, OCBJG20, PPD^{+}25, SKS^{+}23, SR20, SGD^{+}21].

Catching [MBB^{+}23, EDA^{+}22]. Catchment [CGT^{+}23]. Catchment-scale [CGT^{+}23]. catfish [BKR^{+}22, FBR^{+}24, NTJN21, NFC20, SCGW24]. catfishes [VBL^{+}22, VBL^{+}24]. caught [AHL20, AIW^{+}23, BCPH22, BRvL^{+}22, BLK23, CHM24, CBR^{+}23, EMJ^{+}22, EFM25, GAB^{+}24, LPG^{+}24, MVDH24, MBH^{+}22, SHH^{+}21, SSRC24, SXMV^{+}21, TPW23, YTH22]. causal [HLCC22]. cause [DQK^{+}23]. Caveats [AAT^{+}21]. Celtic [BMM^{+}24, BGM^{+}23, HG20, LCB^{+}21a, LRW^{+}24, MHL^{+}23b]. Center [MCS^{+}25]. Central [AMSC20, BNTK23, CPPK23, LP23, PDG^{+}22, SML^{+}24, ABBO20, BCRI21, CMD^{+}23, CGSL22, Lya20b, Lya22, MOA23, MVLC^{+}20, PVPN22, SPM^{+}24, VAVQGD^{+}20, VP23, AHEV24, LCMS^{+}22, PFdSBL25, TSS^{+}23]. Centropristis [CG21]. Centrostephanus [BWB^{+}24]. century [NNTM^{+}20, RPH20, SBJ^{+}20]. cephalopod [GP21]. Cephalopoda [GMERCM^{+}24, RdLSdB^{+}21, SSV^{+}20]. cephalopods [LPRB^{+}21]. certain [RNKB23]. Cestode [FRP22]. cetacean [KBB^{+}21]. cetaceans [CHT20]. cf [GBC^{+}23b]. chain [AOA^{+}22, BRvL^{+}22, CAGLT23, DSP22, dHJTCE23, RRSP^{+}24, SO25, Seu22, SCHSC21]. challengeri [MTS^{+}21]. challenges [YH21a, YH21b]. Chamelea [MPM22, ÖA21, PHV^{+}21, dAGCR21]. Champsocephalus [ZZ22]. Change [LWH^{+}23a, GGL^{+}24, GdSPL21, HJMS20, LMM^{+}24, MA20, MMC^{+}24, Mur20, NSQV22, NNTM^{+}20, PFdSBL25, SGW^{+}20, SB24, TCVG20]. changed [BTFL22]. Changes [BTC23, LYX^{+}21, vORP23, AIW^{+}23, BBPT^{+}24, BKM23b, CGM^{+}22, CAYM^{+}23, DPL20, FAK24, GBWM22, GBO^{+}20, HCCC21, HW21, HPS^{+}24, KHMC23, LLFL21, LK25, MMC^{+}24, MCPJET^{+}20, NNTM^{+}20, SSI^{+}23, SPM^{+}24, SAA23, YWC^{+}21]. changing [BM25b, GTS^{+}21, PVPN22, PDD^{+}22]. channel [BKR^{+}22, ABF^{+}21, LCB^{+}21a, MHL^{+}23b, NLS21]. channeled [MFJ^{+}24, WPGO21]. Channichthyidae [LGD^{+}20]. char [LvCdGSL23, YIM^{+}20]. Characiformes [AGNS^{+}21]. characterisation [AMHR22, BAYR^{+}24]. Characterising [CCGR20]. Characteristics [ECM^{+}24, CBN^{+}21, CBHS24, KNP^{+}20, KPUB22, LCW^{+}23a, LTT^{+}23, MKS^{+}21b, NFC20, PSSFS24, SOS^{+}21, SKBA23, SAHW22, TCL^{+}21, WWF^{+}20, WYM^{+}25, WSWL25]. Characterization [BNL^{+}23, FBR^{+}24, PGD^{+}25, CKD^{+}21, CH22, NSRM22]. Characterizing [RNP^{+}24, BDA^{+}20]. charr [PD25, YiTM23]. Charter [ÖÜÖG20]. Charybdis [TNTN20]. Chatham [LPP^{+}20, NAV^{+}23]. chemical [FJJT^{+}21, MJD^{+}21, MMM^{+}20, MHB^{+}23]. chemistry [AAR^{+}21, BTB^{+}21, FJJT^{+}21, HRH22, MZZ^{+}21, PYX^{+}20, RTB^{+}21, ZSDZ24, dAdCdO^{+}23]. Chesapeake [DWLT21, MDW^{+}21, MWJ^{+}24, SCGW24, SSFL24]. Chile [BCRI21, COG22, CQA^{+}24, ECY21, MVLC^{+}20, PAY23, PSS^{+}21, QMGRIU22, SCCAM21]. Chilean [HRC23, QAS^{+}25, RUHM20, YRTP20]. Chiloscyllium [FTB^{+}21]. China [FHSC21, LZC^{+}21, LZCC24, MYKO23, NLW^{+}22, SZXC25, WLZ^{+}21, XWD^{+}21, YTHM20, YH21a, ZZC^{+}21, ZZH^{+}24, BaLK^{+}21, DSJG20, FKW^{+}22, HPL^{+}24, LCW^{+}23a, MZZ^{+}21, RLQ^{+}20, WW21a, XCB^{+}21, XqRJ^{+}23, YCC22, YH21b, YLS^{+}23, ZLXL20]. chinensis [WZL^{+}22]. Chinese [JWL^{+}24, PYX^{+}20, XqRJ^{+}23]. Chinook [BJS^{+}22b, BKC21, CSDH^{+}23, CCC^{+}22b, IRJ^{+}22, LP23, WBA23]. Chioggia [SPC^{+}25]. Chionoecetes [AIJ^{+}23, BMSM22, BWN^{+}21, HLG^{+}21, HSSD^{+}21, MBD^{+}21, Mur20, Mur21, NGDC25, NBD^{+}20]. Chishui [LYX^{+}21]. chizothorax [HLC^{+}25]. choice [KMA22, PDD^{+}22, PNRS23, TGG^{+}24]. choices [PvZ22]. Chondrichthyes [SPM^{+}24, PTK^{+}20]. choosing [LDM^{+}24]. chronological [HMR^{+}24b, SDdMG^{+}20]. Chrysophrys [TP24]. chrysops [CG21]. chub [CMTP^{+}21, HGC24, WYM^{+}25]. chum [OKKW20, ACS23]. circle [FMD^{+}24]. cirratus [BRN^{+}20]. citizen [GAB^{+}20, HMS^{+}22, PJNGJ^{+}22, SGW^{+}22]. citizen-science [HMS^{+}22]. Cittarium [BQGV^{+}24]. clam [BBHF25, LCM^{+}23, MPM22, MMBH23, ÖA21, PUC^{+}23, PHV^{+}21, SCW21]. clams [TC24]. Clarias [NTJN21]. clarifying [PFFdC22]. class [LADA^{+}22]. classes [NSRM22]. classical [HIF^{+}24]. Classification [FGTA24, BGBM22, JMS25, LGD^{+}20, LCLM23, SKY^{+}24, SPD^{+}24, WSB24]. classify [HZZ^{+}20]. clavata [TLV23]. claw [GC21]. clawed [TSPK24]. clear [RWB^{+}23, TRWH23]. clients [FCML^{+}22]. Climate [ANB^{+}24, HJMS20, KHMC23, MA20, Mur20, GGL^{+}24, GdSPL21, HSJ^{+}24, LMM^{+}24, LADA^{+}22, MMC^{+}24, MJC^{+}23, SCCAM21, SB24, TRN^{+}23, TCVG20]. Climate-driven [KHMC23]. climate-stressed [LADA^{+}22]. close [Fis25, PTL^{+}24]. close-kin [Fis25, PTL^{+}24]. closed [Det23]. Closing [RNKB23]. closure [JWL^{+}24, LYX^{+}21]. closures [DML^{+}20, EHE^{+}23, HLMV24, HCDF24, WKBMW24]. Clupanodon [XqRJ^{+}23]. Clupea [íKMíH^{+}22, RS21, SMC^{+}24, WYM^{+}25]. clupeid [MNS^{+}20]. cluster [LCMS^{+}22]. clustering [FOS^{+}21]. Cnidaria [GKC^{+}22]. Co [EBN^{+}23, KMC^{+}23, KFI^{+}21, MBZSM20, NC20, RUHM20]. Co-designing [EBN^{+}23]. co-existence [KFI^{+}21]. co-management [KMC^{+}23, MBZSM20, NC20, RUHM20]. coast [ASD^{+}22, BGG^{+}22, BBM^{+}24, CKD^{+}21, CDAK23, CLY^{+}22, Dra22, dSRFFN^{+}20, GMRRG20, GHAZ21, LMT^{+}22, MWTH23, ML24, MRC24, MVDH24, MPV^{+}24, OAM^{+}21, ÖSL^{+}23, ODM20, PYX^{+}20, PCGG20, QMGRIU22, RKN23, SCGM^{+}21, SS24, SCN^{+}24, TMN^{+}21, VAVQGD^{+}20, WBBG^{+}23, BTC23, CRCAF^{+}22, JLYR24, LWH^{+}21, MPSH25, SdOR^{+}23, SS23]. coastal [ARD^{+}23, BR22, BLK23, CCR24, DAR^{+}23, DEM^{+}23, DSNK^{+}22, EHE^{+}23, FFG^{+}20, FDS^{+}23, FBALRR^{+}22, Fre22, FHE^{+}24, GSS^{+}23, GdSPL21, HWW22, HMS^{+}22, KFI^{+}21, LLK^{+}22, LCC25, LMK^{+}23, MKS^{+}21b, NVSG24, OAB^{+}23, OOAF^{+}21, RMD^{+}25, RSPE22, SBD^{+}22, SKBA23, SGD^{+}21, SGZH^{+}22, SME^{+}24, TSC^{+}20, TYYK21, TMDA22, TCVG20, WEH^{+}25, ZDF^{+}22]. coastline [BJK24, FJJT^{+}21, PCBL23]. coastwide [WMT^{+}20]. coatesi [KCB^{+}24]. cobia [SGW^{+}20]. cockles [TC24]. Cod [SMLT24, BWB^{+}23, BKM^{+}23c, CWRR24, CSH^{+}21, EM23, HMM^{+}21, HOQ20, JBL^{+}22, KLBHK23, KHK^{+}20, LNvD^{+}23, MINS21, MHH^{+}20, MB20, NEBP^{+}23, OOM^{+}23, PNAPH24, PDA^{+}24, RBD^{+}22, RBM21, RNP^{+}24, RNKB23, SJB^{+}20, SBL^{+}23, SGZD20, SGZD21, SGZH^{+}22, SF20, SF22b, SWIRF21, UPBH^{+}20, WCN^{+}24]. cod-end [RBM21]. coded [Bea21, SW20]. coded-wire [Bea21]. codend [BJHS^{+}23, BW23, BLHS20, BM22, BM24a, BM25b, GTS^{+}21, IB20, JBL^{+}22, LTT^{+}23, MINS21, MOB^{+}23, MBB^{+}23]. codends [OFF^{+}20, SHB^{+}23]. coho [GQP^{+}25, BJS^{+}22a, DBS^{+}21, LCHB^{+}24]. cohort [AGS20, CPM21, CTIC23]. COI [BQGV^{+}24]. coincidental [AAG22]. coindetii [OBB^{+}20]. Cold [UAB^{+}21]. colias [CMTP^{+}21, NVSG24, PFC^{+}21, SCN^{+}24]. collaboration [AVCA22, MVMdS^{+}24]. collaborations [vPDD^{+}25]. collaborative [SCC^{+}22]. collapse [AVACA^{+}23, BAJB^{+}24]. collapsed [MHD^{+}21]. Collating [DDDP21]. collected [DTSR22, GRJW20, RRH^{+}24]. collecting [GAB^{+}20]. Cololabis [HLZ^{+}20, LZL^{+}24]. Colombian [APBN22, AGNR^{+}21, GZLCRG25, GHAZ21]. color [AMHH21, Peñ21, SK22]. Columbia [BJS^{+}22a, BJS^{+}22b, MOB^{+}22, BKC21, RGN^{+}20, ZF21]. column [CKK^{+}20]. combinations [HF20]. combined [Bea21, GRJW20, OAM^{+}21]. Combining [DWS^{+}23, WHM23, PPC^{+}21, TSC^{+}22, WSF22]. come [ANB^{+}24]. comeback [NBF20]. Comfort [DBL^{+}25]. Comments [MBD^{+}21, BWN^{+}21]. Commercial [HGHH25, ACL^{+}20, AHEV24, ARD^{+}23, BDA^{+}20, DHH^{+}22, DC24, FOM21, GC25, GWMC21, HSM^{+}24, HMR24a, Hut22, JYH21, Ken21, KBB^{+}21, LCB^{+}21a, LPRB^{+}21, MTC^{+}22, MFR22a, MKC20, MBE^{+}20, MAA^{+}20, MMBH23, NOL23, PBDM21, RAE^{+}21, STH25, SWH^{+}24a, SCC^{+}22, SLW^{+}20, TJS23, WEH^{+}25]. commercial-recreational [ARD^{+}23]. commercialization [SCHSC21]. commercially [DSP^{+}23, EHBJ20, FMSA21, KPWS21, KK21, MRP^{+}23, MGC^{+}22, MTE^{+}20, NVRG^{+}21, PCK23, TLV23, WCGB22, YHST22]. commercially-important [PCK23]. Common [BRN^{+}20, RMNB^{+}21, ASDW24, CMD^{+}23, CHGC25, CWM^{+}23, EHBJ20, FGPVPPGG22, FMMC20, GMT24, LCD^{+}23, LDM^{+}24, Lya20a, MVLC^{+}20, PVA^{+}24, PIP^{+}22, PRF^{+}21, PCC^{+}23, SPC^{+}25, Žák21]. communication [CCKL^{+}20, MHB^{+}23]. communities [BCG^{+}25, HSL^{+}22, LWX^{+}20, PSS^{+}20, SBD^{+}22, TCVG20, ZJJ^{+}25]. community [GWGM24, KWMA23, LCFJ22, MOA23, MLS^{+}21, SWLH20, SZXC25, SAA23, WASS20]. community-governed [KWMA23]. Comparative [BN21, LGD^{+}20, CYBW22, GCK^{+}21, LCW23b, QSMG^{+}23, TLV23]. compare [CSH^{+}21]. compared [IB20, JBL^{+}22, MSW21, NVB^{+}23, PZG^{+}20]. Comparing [Bro25, HCWH23, MPP20, PDG^{+}22, SFMA23, SSM^{+}23, SAHW22, CHGC25, SMK^{+}24]. Comparison [Alv21, CB20, FBB20, FS20, HPL^{+}24, HSL^{+}22, KTFY22, SFYM24, TBF^{+}21, WBBG^{+}23, YIM^{+}20, CBA^{+}24, FGTA24, FG23, GLA^{+}20, LOFS22, MFR22a, MLS^{+}21, PGD^{+}25, SEM^{+}23a, WQGS25]. comparisons [NSRM22]. Competition [SGZD21, Mur20, OGFW24]. competitions [SRP^{+}22, Žák21]. Complementary [LVL25]. complements [PCK23]. complete [HGC24, LYX^{+}21]. complex [ABF^{+}21, CGC24, XDX^{+}23]. complexities [BCM^{+}21, MBZSM20]. complexity [FTH^{+}23, OCdMC24]. complexone [CHL^{+}20]. complexone-stained [CHL^{+}20]. compliance [BSAP22]. complicate [DLKH22]. component [LNvD^{+}23]. composition [ASD^{+}22, BLC^{+}21, BBC^{+}20, CHGC25, CNE^{+}22, DEM^{+}23, ECM^{+}24, FCSA21, GBO^{+}20, HWW22, HW24, LCMS^{+}22, LCDM^{+}24, LJX^{+}20, Lun25, MTX^{+}20, PAA^{+}24, PPM^{+}23, PKRL21, SJ24, SAA23, TNDM23, UAG^{+}23, WSB22]. compositional [TAK^{+}23]. compositions [SKST23]. comprehensive [AIM^{+}23, GGTÁVT^{+}20, HOQ20]. computational [SDBS21]. computationally [CTIC23]. computer [SKY^{+}24, SLA24]. Computing [DLP^{+}24, SPD^{+}24]. concept [WGNM24, YLS^{+}23]. conceptual [MVMdS^{+}24]. concerning [YTSS22]. concurrent [LM22, PHH^{+}23]. condemned [Pun23]. condition [ABT^{+}24, BSR^{+}22, CHPT20, CHM24, GKC^{+}22, LR21, LHPR21, MRC24, PNAPH24, SFYM24, SWIRF21, WCN^{+}24, vDJB^{+}23]. conditional [HW24, JND^{+}23, PAY23]. conditioned [VBB20]. conditions [GRG24, GAB^{+}22b, JHB21, KSI20b, MSW21, OB21, PPC^{+}23b, SSJ^{+}21, TNDM23, YMS21]. conductive [RBS^{+}24]. Conference [CSTdL25]. confidence [CAZN24, CLM^{+}22, MB25, RRH^{+}24]. configuration [JBL^{+}22, SKS^{+}23, ZHC^{+}25]. configurations [BM25b, LDS^{+}21, RB22]. confirming [DHC^{+}20]. conflicts [DSJG20]. confluence [LP23]. confluentus [JSKM20]. confounded [Szu22]. congeners [WJN^{+}25]. Conger [MZZ^{+}21]. Congruence [BQGV^{+}24]. connection [WSWL25]. connections [MCGL^{+}25]. connectivity [BRGB^{+}23, CPB^{+}21, GAW^{+}22, LNvD^{+}23, MSLMOC^{+}24, MPC^{+}20, PYX^{+}20, PCBL23, SXK^{+}24, ZLXL20]. Consequences [BSR^{+}22, FBPCC^{+}21, MJC^{+}23, PW25, PCJH^{+}21, SF20, AUHK22, Pun23]. consequent [MPSM25]. Conservation [KMC^{+}23, CGSL22, JZQZ20, LFdB^{+}21, MPV^{+}24, NOL23, NAS^{+}20, PKRL21, PGAG22, PK24, PBD25, PDC^{+}24, SKBA23]. conserve [JWL^{+}24]. considerations [HC22a, MAA^{+}20, Mun24]. considering [PDD^{+}22]. consistency [WSL^{+}24]. constant [DB22]. construction [CTS^{+}23]. Consumer [RSD23, ZDF^{+}22, DCL^{+}20]. Consumption [OKKW20, BGL^{+}22, BCRI21, CAGLT23, HLCC22, LMK^{+}23, Sch23, VA20, dCHdMS^{+}23]. consumption-mortality [VA20]. contact [AdABW^{+}22, MKH^{+}20]. contemporary [BBG^{+}24, BCM^{+}21]. content [BÁP^{+}23, BALBC23, Lun25, RWFT25, WWT^{+}25]. context [BHB^{+}22, DSS^{+}23, GLA21, GGL^{+}24, MHL^{+}23a]. context-dependent [DSS^{+}23]. contextual [GAB^{+}22b]. contiguous [PDG^{+}22]. continental [ABK^{+}21, BPT^{+}25, CBD^{+}22, FSS^{+}24, MFO21, Ten22]. contingent [JCL^{+}21]. continued [BMJ^{+}24]. continuity [RGG22]. continuous [LZW^{+}21, MGC^{+}22]. continuum [Cop24]. Contrast [KPS20]. Contrasting [CKDP^{+}20, JFS21]. contribute [BMOC22, CMIMS20, HCCC21, UAB^{+}21]. contributing [BLC^{+}21, YTSS22]. Contribution [SSS^{+}23, BÁP^{+}23, FKS^{+}20, NFdSJO25]. contributions [ZSWK25]. control [BJHS^{+}23, BBPT^{+}22, BBPT^{+}24, Che21, HSSD^{+}21, LLFL21, WHM23, WH23]. controlled [GRG24, PF20]. controlling [CHPT20]. controls [DRSPTA20, WY20]. conventional [BNTK23, JBL^{+}22, MCC^{+}22]. convergence [FKW^{+}22]. convolution [İTAD24]. convolutional [WZL^{+}22]. conwayi [FCMP23]. Cook [Har21]. cookbook [CWC^{+}21]. cooling [LTR20a]. cooperation [CMIMS20]. cooperative [BMJ^{+}24, HMS^{+}22]. cooperatives [EBdRPC24]. copepods [EYAO20]. coral [ASJ^{+}22, MTE^{+}20, RBHM24a, WCC24]. coralgroupers [SK22]. cordatus [ASD^{+}22, MPEBdR23]. core [TGG^{+}24]. Coregonus [KSV^{+}22, SHV20, WKSF20]. Cormorant [WEH^{+}25]. cormorants [ALW^{+}21, Lya20b]. correct [TAK^{+}23]. correcting [FS20, MMQ21]. corrections [RWFT25]. correlated [SCD^{+}22]. correlations [SCG^{+}24]. Correspondence [LHPR21]. corresponds [Žák21]. Corrigendum [Ash20a, DBV23, OS21a, SIM^{+}24a, Tho21, VK21a, WGC^{+}21, YH21a]. Corroborating [LRMH21]. corrugata [VAVQGD^{+}20]. Corsica [MPM^{+}23]. cortisol [BALBC23, CCC^{+}22b]. Corvina [AVACA^{+}23]. Cost [BH23, AYTM21, ANB^{+}24, Bea21, EVS^{+}23, HZZ^{+}20, HBMC21, KKC25, LCN^{+}20, PBDM23, PJOR20, TLAM25]. cost-effective [Bea21]. cost-effectiveness [KKC25]. Costa [OPL21]. Costly [KAC^{+}23]. costs [SGZD20, ZSWK25]. Cottidae [TEO25]. could [RNKB23]. count [PJOR20]. counted [KWW^{+}21]. Counter [TTYT24, MKH^{+}20]. counter-herding [MKH^{+}20]. counting [FWKR21, SZS^{+}24]. countries [RRSP^{+}24]. country [HLCC22]. counts [DGMG^{+}22, HL21]. coupled [BHB24, HIF^{+}24, JK20]. covariate [PMC^{+}24]. covariates [RRB24, SM21]. cover [BM22, PASdCF23]. coverage [CC20]. covered [BM24a]. COVID [Apo25, FBM^{+}21, HJA^{+}21, PZL^{+}23, SCC^{+}22]. COVID-19 [Apo25, FBM^{+}21, HJA^{+}21, PZL^{+}23, SCC^{+}22]. Cowlitz [CBR^{+}23]. CPUE [Det21a, LZY24b]. Crab [CMRP20, MDS^{+}20, AIJ^{+}23, AAH^{+}23, ASD^{+}22, BMSM22, BWN^{+}21, BJK24, CDAK23, CBJ24, DMNV^{+}24, FBB20, FJS^{+}25, GZLCRG25, GC21, HLG^{+}21, HSSD^{+}21, HRC23, HdLHD22, HKCW24, JMS25, JYH21, KAC^{+}23, KK21, KDF^{+}25, LYH^{+}21, aLBK^{+}21, LZW^{+}21, LDC24, LTR20a, LTR20b, MHD^{+}21, MDJP24, MFM^{+}20, MRUG^{+}23, MOTL25, MHL^{+}23b, MCMM20, MSJ21, MPEBdR23, MBZSM20, MB20, MBP20, MBD^{+}21, Mur20, Mur21, NLW^{+}22, NGDC25, NBD^{+}20, OMG^{+}23, PBM^{+}23, PDF20, PDD^{+}22, RPH20, RL24, SDdMG^{+}20, SDV^{+}22, SSG^{+}22, SV23, Szu22, TNTN20, WCGB22, WPLF20, YHC^{+}24, ZZH^{+}24, HTSJ23]. crabs [ASD^{+}22, CVM^{+}20, HHJ^{+}22, MWTH23, ML24, PFGQ20, SSFL24, SSP^{+}23]. Crack [VK21a, VK21b]. Crangon [TBH^{+}22, VMFF^{+}20]. crappie [HBMC21]. Crassostrea [OYOO21]. crayfish [DD25]. creating [LLS23]. creel [JDP22, TCL^{+}24]. crescendo [MPSM25]. crisis [Sat23]. criteria [BMA^{+}24, CTS^{+}23]. critically [PBB20, dS21]. croaker [ABT^{+}24, LLC^{+}20, SXK^{+}24, WLZ^{+}21, YLP^{+}23, ZLXL20]. crocea [LLC^{+}20, WLZ^{+}21]. cross [Hut22]. cross-regional [Hut22]. crucial [FDS^{+}23]. crumenophthalmus [WSL21]. Crustacea [PDG^{+}22]. crustacean [DMM^{+}21, KHMC23, MDC^{+}22a, SDdMG^{+}20]. crustaceans [KAC^{+}23]. cryptic [DOB^{+}24, WJN^{+}25]. ctenophore [SMC^{+}24]. cubera [MFO21, MFR^{+}22b]. cuchia [NK23]. cuckoo [BMM^{+}24, NDRR20]. cucumber [AHEV24, CPPK23, PMS^{+}20, SWH^{+}24a, XDX^{+}23]. cues [CWM^{+}23]. Culter [LCG^{+}21]. cultural [RGP^{+}23]. culture [LPRB^{+}21]. cultured [CHL^{+}20]. cultures [BHB^{+}22]. cumulative [RBH^{+}24]. Cunene [MPV^{+}24]. curema [ÁHGCVAI22, CCRGC^{+}24]. Current [AOA^{+}22, DSB^{+}21, DD25, GBO^{+}20, MBE^{+}20, HEG^{+}23b, KTFY22, Pun23, RDR^{+}23]. curve [CMD^{+}23, KHPB20, YCC22]. cusk [BÁP^{+}23]. cusk-eel [BÁP^{+}23]. Customary [RGP^{+}23]. customization [Peñ21]. cuticle [SDdMG^{+}20]. cutthroat [HQWD20]. cuttlefish [BCPH22, GCK^{+}21, GKC21, GMERCM^{+}24, KSI^{+}22, LCD^{+}23, SPC^{+}25, YLX^{+}24]. Cuvier [ALRB^{+}20, SRT^{+}20]. cyanea [CMBL21, HMY25, SKS^{+}23]. cyanopterus [MFO21, MFR^{+}22b]. cycle [CSDH^{+}23, HGC24, MPM22, OMG^{+}23]. cyclones [LNP25]. Cyclopterus [Ken21]. Cynoscion [AVACA^{+}23, SdOR^{+}23]. cyprinid [AK23, ŠBB^{+}22]. cyprinids [KC22, ÖSL^{+}23]. Cyprinus [AK23, CWM^{+}23, Lya20a, NA22]. Cyprus [MKC20].

D [CMRP20, DMM^{+}21, MCMM20, WZL^{+}22]. daily [ATA^{+}24, ECK^{+}21, HPPT24, PJSQ20, PvZ22, UFYT23]. Dakota [OGFW24]. dam [dAdSR^{+}20]. Damage [GdSPL21, BHG^{+}24a, KFI^{+}21]. damaged [BLK23]. dams [SdFZFJ21]. dangers [GPW^{+}20]. Danish [EFM25, WBBG^{+}23]. dart [MCC^{+}22, SAA^{+}22]. Dashentang [HPL^{+}24]. Data [TMH23, WKSF20, dLH23, AMHR22, AAG22, Aks24, ATAS20, AAT^{+}21, ASJ^{+}20, AVCA22, ASDW24, AGB^{+}24, BNTK23, BGG^{+}22, BDA^{+}20, BLE^{+}22, CYBW22, CPM21, CEAL21, CCR24, Cla22, Cop24, CBHS24, CCCM^{+}20, CHAY^{+}25, DHX^{+}23, DTSR22, Det21a, DMM^{+}23, DMS22, DC24, DAD^{+}22, EBGE21, FGTA24, FDB^{+}20, FCSA21, Fis25, FCMP23, GGG^{+}22, GC25, GM21, GWMC21, GBWM22, GRJW20, GML^{+}23, GAB^{+}20, HKG^{+}21, HPPT24, HL21, HEG^{+}23a, HFKS20, HCWH23, HSM^{+}24, JDP22, JM23, KM23, KMA23, LVL25, LBLF20, LWH^{+}23b, LBD24, LZY^{+}24c, MCS^{+}24, MLCMdS23, MCPJET^{+}20, MTX^{+}20, MKS^{+}22, MDJP24, MHH^{+}20, MOI23, Mon24, MPC^{+}20, MPP20, MOB^{+}22, OFF^{+}20, OUB^{+}22, ÖRS^{+}25, PVA^{+}24, PPD^{+}25, PCBL23, PJNGJ^{+}22, PGAG22, PMS^{+}20, PJOR20, PMC^{+}24, PNGGO^{+}22, PTL^{+}24, RGG22, RFMS^{+}21, RTHB25, RNP^{+}24, RRH^{+}24, SCGM^{+}21, SKW^{+}21, SSP24, SDV^{+}22, SSM^{+}23]. data [SBJ^{+}20, SBRM^{+}22, SH22b, SLW^{+}20, TMP20, TYYK21, TAK^{+}23, TDI^{+}21, TCL^{+}24, WWF^{+}20, WS24, XMCC20]. data-deficient [AGB^{+}24, FCMP23]. data-driven [AVCA22]. data-limited [CBHS24, CHAY^{+}25, DMS22, HFKS20, LZY^{+}24c, MOI23]. data-poor [CPM21, CCR24, FDB^{+}20, PMS^{+}20, SSM^{+}23, WWF^{+}20]. data-rich [Cla22, Mon24]. database [APGG22]. date [SF20, UFYT23]. day [MB23]. day-to-day [MB23]. DCN [SZS^{+}24]. DEA [SSP^{+}23]. decades [ACL^{+}20, ABK^{+}21]. decapod [SDdMG^{+}20, SB24]. Decapoda [AGNR^{+}21, GBO^{+}20, PDG^{+}22, PCGG20, SS24]. December [Ano20c, Ano21c, Ano22c, Ano23c, Ano24c]. decision [DMS22, vP20]. decision-making [DMS22]. decline [BLE^{+}22, MBP20, PNAPH24]. declines [ANB^{+}24, CKM^{+}20]. Declining [LCN^{+}20, PMC^{+}24]. Deconstructing [DHB^{+}21]. Deep [LG21, SML^{+}24, BHB24, CPL^{+}25, DMF^{+}21, DWS^{+}23, FMD^{+}24, GBB20, HdLHD22, HZZ^{+}20, IJS^{+}22, İTAD24, KFO20, PÁEMC22, PDG^{+}22, PPC^{+}21, SIM^{+}24a, SIM^{+}24b, SH22b, WWF^{+}20, YAO^{+}23]. deep-sea [KFO20]. deep-set [SH22b]. deep-water [CPL^{+}25, FMD^{+}24, IJS^{+}22, PDG^{+}22, WWF^{+}20]. deeper [CRF^{+}24]. deeply [CLD^{+}22]. deepwater [CH21b, CRS23, MOB^{+}23, SKW^{+}21]. deficient [AGB^{+}24, FCMP23]. Defining [Cad20, MBE^{+}20, MAA^{+}20, OB21, AGB^{+}24, CGBJ23, CCGR20]. deforestation [dFBPL^{+}20]. deformities [FGL22]. defy [BRN^{+}20]. degradation [BF25]. degrees [FCSA21]. dehooking [CCC^{+}21, CCC^{+}22a]. del [CRSC22]. delayed [BAC^{+}22, EMR^{+}22, LBP^{+}24]. delight [ASB^{+}24]. delineate [KHC^{+}20]. Delineating [NA22, Ten22]. Delphinus [PCC^{+}23]. delphis [PCC^{+}23]. Delta [OYOO21, ZF21, LK25, TCJ^{+}21, IOO^{+}24, NTJN21]. delta-generalized [LK25]. delta-model [TCJ^{+}21]. demand [ELM20]. demand-side [ELM20]. Demersal [AMHR22, BGM^{+}23, BCF^{+}23, CHAY^{+}25, DMF^{+}21, DWS^{+}23, EPHDB24, HKG^{+}21, HVMEO25, PTK^{+}20, RBG^{+}24, SWLH20, SOSK22, TMN^{+}21, TRWH23, WGNM24, dJDM23]. demi [BWB^{+}24]. demi-pyramid [BWB^{+}24]. demographic [DAL20, GPW^{+}20, vdHC20]. demographics [HSM^{+}24]. demonstrated [CHL^{+}20]. Demonstration [CTIC23]. dens [CHB24]. densities [DKD^{+}21, SEM^{+}23a]. Density [HGS^{+}23, MPEBdR23, TdL24, BAM^{+}24, BBČ^{+}21, CBJ24, CQA^{+}24, DKBF23, EBGE21, FBALRR^{+}22, HMC^{+}23, HLZ^{+}20, HC22c, KTFY22, LM22, Lya20b, MDL^{+}21, Mau22, SB20, WGFM21]. Density- [HGS^{+}23]. Density-dependence [TdL24, Mau22, WGFM21]. density-dependent [KTFY22, MDL^{+}21, Mau22]. Densu [OYOO21]. dentatus [WMSW22]. depend [MKH^{+}20]. dependence [CMV21, Mau22, MSS^{+}21, TdL24, WGFM21]. dependent [AGB^{+}24, DSS^{+}23, DBGV^{+}22, GML^{+}23, HCWH23, HGS^{+}23, JAN23, JSG21, KTFY22, KSS^{+}22, LVL25, LBD23, Lor22, MDL^{+}21, Mau22, RPM^{+}21]. Depensation [DSS^{+}23]. depleted [FRP22]. Depletion [Dik24, MDMS21, PPH21, SSSF25, dLH23]. deployment [UAB^{+}21]. DEPM [MSS^{+}21]. deposition [UFYT23]. depositional [TBÓ^{+}22]. Depredation [MWTH23, BTR^{+}24, CMA^{+}22, GKKL24, KLBHK23, MRE^{+}24, WOG^{+}25]. depress [HSM21]. depth [AHL20, FGCB^{+}21, LJB^{+}24, Mur20, SFC21]. derive [TMP20]. derived [ACS23, FHSC21]. deriving [GVK^{+}23]. describe [PPC^{+}23b]. description [AIM^{+}23, FCMP23]. desert [BS20b]. design [AAH^{+}23, AVCA22, BWOR23, CBA^{+}24, CFO23, DTSR22, FAK24, LRMH21, LNW20, MMML24, PPH21, SFC21, SSKS21, SDBS21, SR20, YSB^{+}21, YRTP20]. design-based [CBA^{+}24, CFO23, DTSR22]. Designing [RMD^{+}25, EBN^{+}23]. designs [BBJ22, BLFT23, CZJ^{+}24, LCMS^{+}22, MINS21, TNTN20]. desirability [CCGR20]. destination [KWW^{+}21]. destructively [PJSQ20]. detail [WHR^{+}24]. detected [DOB^{+}24, SIM^{+}24a, SIM^{+}24b]. Detecting [WPB^{+}20, DC24, SZS^{+}24]. Detection [DHH^{+}22, AFR^{+}24, LSZ^{+}23, SKY^{+}24, SJW^{+}22]. detections [RGG22]. detective [PBRT22]. Determinants [BGLP21]. determination [BRN^{+}20, SDdMG^{+}20, SLW^{+}20]. determine [CRS23, KLS20, MTX^{+}20, PvZ22]. determined [DMM^{+}23, MOI23]. Determining [BAF23, KD22, MTS^{+}21, SR20]. deterrent [LLK^{+}22, PCC^{+}23]. deterrents [LLS23]. develop [JDH22, WSL^{+}24]. Developing [CBA^{+}24, DMNV^{+}24, HMP^{+}22, MWJ^{+}24, RB22, RRH^{+}24, CGC24, GBC23a]. Development [CBD^{+}22, DNLM23, HC22a, HHD^{+}20, HdLHD22, LNMA23, PMS^{+}20, ALRB^{+}20, FFG^{+}20, FBQA20, HFMH20, HIKM21, JZQZ20, LLK^{+}22, LZCC24, MSD21, ORdlG^{+}24, SCGW24, SOSK22, TMP20, TLC^{+}22]. developments [MCGL^{+}25]. device [AUM21, BM23b, CCKL^{+}20, IKBL23, KNP^{+}20, MKH^{+}20, PCC^{+}23, QAS^{+}25, RB22, SFC21, YSB^{+}21]. Devices [BDA^{+}20, BM24b, CTM^{+}20, GGTÁVT^{+}20, HBE^{+}22, IAB20, TBF^{+}21, TFC^{+}20, ZZC^{+}21]. devil [WHR^{+}24]. dFAD [IKBL23]. DFADs [BDA^{+}20, PGAG22]. diadromous [CGT^{+}23]. diagnostic [LDM^{+}24, MUF^{+}22, PRA^{+}23]. diagnostics [CBA^{+}24, CWC^{+}21, KDBOC25, PC21]. diameter [BKM23a]. diamond [BLHS20, SHB^{+}23]. diamondback [SSV^{+}20]. diarist [LWH^{+}23b]. diary [LWH^{+}23b]. Dicentrarchus [LRW^{+}24]. dieffenbachii [CGT^{+}23]. Diel [LWX^{+}20, Žák21, OSEF22]. Diet [HUK^{+}23, HTK^{+}24, QMGRIU22, WWT^{+}25, BÁP^{+}23, CZ25, LCFJ22, LJX^{+}20, MRC24, RMRG22, RdLSdB^{+}21, RCH^{+}21]. diet-switching [HTK^{+}24]. dietary [BBC^{+}20]. diets [Lun25]. differ [SLF23]. difference [ABF^{+}21]. Differences [EWPB22, SOS^{+}21, Ash20a, Ash20b, CSRL20, DHH^{+}22, DKD^{+}21, FGCB^{+}21, HGC^{+}21, KKCP20, MDC^{+}22b, OMG^{+}23, SWLH20, SSD^{+}20a, SF20, SEM^{+}23b, WCLN20, WHCF22]. Different [LZL^{+}24, AVMBEB22, AMSC20, BSR^{+}22, CKDP^{+}20, CCC^{+}22a, FCSA21, HKCW24, JDH22, KFO20, LGD^{+}20, LTT^{+}23, LZX^{+}20, PZG^{+}20, PNRS23, SACS23, SBT^{+}20, TTL^{+}20, ZOS^{+}23]. Differentiating [RSD23]. differentiation [CPB^{+}21, FKW^{+}22, HL21, SPC^{+}23]. Digital [UVA20, AYTM21, HZZ^{+}20]. digitization [VTS^{+}22]. Dilemma [PVC^{+}22]. dimensional [SXM^{+}21, ZHC^{+}25]. dimensions [AL22]. diminish [DCL^{+}20]. dimorphic [CW22]. Direct [MFR^{+}22b, SKW^{+}21, WEH^{+}25]. direction [RWB^{+}23]. directions [MAA^{+}20]. Disaggregating [GMT24]. Disappearing [GKKL24]. discard [BKT^{+}20, BMC20, BGCCP22, BC20, EM23, EFM25, FGPVPPGG22, LMK^{+}23, LYH^{+}21, RPM^{+}21, TBH^{+}22]. discarded [BMM^{+}24, EMJ^{+}22, GDVBB^{+}20, UAB^{+}21, UAG^{+}23]. Discarding [PM23]. Discards [FOM21, AAM^{+}20, ABF^{+}21, BNL^{+}23, DML^{+}20, FOS^{+}21, MRG^{+}23, PM23, RLO^{+}21, SFPR^{+}23, TZL^{+}24, dJDM23]. disconnect [SCG^{+}24]. Discovery [RGN^{+}20]. discriminate [BTB^{+}21, DOB^{+}24]. discriminating [GFM^{+}23]. discrimination [AMdC^{+}20, BRR^{+}21, HTK^{+}24, YMYH20, dAdCdO^{+}23]. disease [FJS^{+}25]. Disko [Fre22]. dispar [ZF21]. disparate [CLB^{+}21]. Disparity [BBJ22]. Dispersal [SJB^{+}20]. Displaced [CHB24]. disposition [NMJ^{+}24]. Dissostichus [CPB^{+}21, GMPD23, ZSDZ24]. distance [GGL^{+}24]. distances [CPB^{+}21]. distant [YH21a, YH21b]. distinct [CdSLP21]. Distinguishing [BKT^{+}20, LG21, WS24]. distributed [PBD25, WST^{+}23]. Distribution [BFA^{+}21, BAW^{+}24, BHH21, LPP^{+}20, OAM^{+}21, PDC^{+}24, SSV^{+}20, TEO25, WPB22, ABK^{+}21, BNTK23, BGG^{+}22, BBČ^{+}21, BLFT23, CPL^{+}25, CKK^{+}20, CMBL21, FGSD25, GMRRG20, GWMC21, GdSPL21, HMY25, HSJ^{+}24, HLI^{+}20, JHB21, KA24, LZY24a, LR21, LCW23b, LK25, MPP20, MOB^{+}22, Mur20, NK23, NB24, PCK23, PCBL23, RCVGMI22, RL24, SB20, SWH^{+}24a, Seu22, SW24, SJ24, SFYM24, WSUN^{+}23, ZYZ^{+}23]. distributions [CG21, FAK24, HCWH23, KMA23, TYYK21]. distributor [YTHM20]. diurnal [AIW^{+}23]. divergence [AdABW^{+}22, ODM20, YFJ^{+}25]. Diverse [JvPOG25, VBL^{+}24, MCC20]. diversicolor [MOCGCC^{+}25]. Diversification [SAdC20]. diversity [APBN22, AMHR22, AVACA^{+}23, ABK^{+}21, FKS^{+}20, LRGB25, MPV^{+}24, NDRR20, OAM^{+}21, PCGG20, PNRS23, RdLSdB^{+}21, RFF^{+}22, SPW^{+}22, SWH^{+}24a, SECB21, SPM^{+}24, XWD^{+}21, YGMJ20, ZJY^{+}24]. Diving [HSM21]. d’Ivoire [AAPG21]. DNA [FMSA21, HWW22, HSL^{+}22, LNR^{+}21, PKRL21, RLQ^{+}20, RWB^{+}23, WLZ^{+}21, WMT^{+}20, WMSW22, XWD^{+}21, YLX^{+}24, ZJJ^{+}25]. Do [BCM^{+}21, CBN^{+}21, HSM21, KWMA23, KDdOM^{+}22, MFR22a, PPC23a, SLF23, Sim23, BÖN20, GLP^{+}20, KLNB^{+}24, LD25, LHH^{+}25, TLV23]. document [CC20]. Documenting [SSI^{+}23, WASS20]. Does [BHB^{+}22, MKH^{+}20, NGDC25, BKR^{+}22]. dofleini [NY23]. dogfish [HM25]. dolomieu [OGFW24]. dolphin [MSC^{+}24, PVA^{+}24, PCC^{+}23]. dolphins [BW20]. Dome [HFKS20]. Dome-shaped [HFKS20]. domestic [HSPC21, NLL^{+}25, Seu22]. domesticated [YIM^{+}20]. dominance [Lya22]. dominated [SWLH20]. Donax [dAGCR21]. door [BWR24]. dorsal [CPR^{+}24, PRRR23]. Doryteuthis [MdCG20, RdLSdB^{+}21]. Dosidicus [BCRI21, DLZ^{+}25, LXC^{+}20, QMGRIU22]. double [BLHS20, BHD^{+}23, GGG^{+}22, LCW^{+}23a, MCC^{+}22]. double-tagging [GGG^{+}22, MCC^{+}22]. down [LJB^{+}24]. down-scan [LJB^{+}24]. dragging [JHM22]. dragonfish [GBC^{+}23b]. drainage [KRH^{+}24]. drastic [NNTM^{+}20]. drawn [BAJB^{+}24]. Dredge [PHV^{+}21, KA24, MMBH23, SDBS21, TC24]. drift [JHB21, ODM20, SAHW22]. drifting [IKBL23, SFC21, TFC^{+}20, BDA^{+}20]. drive [KMA22]. driven [AVCA22, DSS^{+}23, KHMC23]. driver [CBJ24]. Drivers [KSI20b, PFdSBL25, RGP^{+}23, BAYR^{+}24, HTSJ23, JYH21, MDL^{+}21, NVB^{+}23, PM23, PHP^{+}20, PDA^{+}24, STH25, SJH^{+}23, SJ24, SFPR^{+}23, SJW^{+}22, YFJ^{+}25]. drives [HLCC22]. drum [CH21a, PHH^{+}23]. drumline [BBC^{+}25, LSM^{+}23]. drumlines [GLP^{+}20, TKB^{+}21]. dubius [SBL^{+}23]. due [MMQ21, MB25]. dumerili [AVB^{+}23]. Dungeness [RPH20]. duration [BWR24, UAB^{+}21]. during [ABT^{+}24, AUM21, Béc20, BHNP22, BKM23b, CZ25, CCC^{+}22b, EMJ^{+}22, EFM25, ECY21, ERS^{+}23, IAB20, IRJ^{+}22, MDL^{+}21, MBP20, OKKW20, PZL^{+}23, QMGRIU22, SKJJ25, TSI^{+}21, YLX^{+}24]. dusky [BAC^{+}22, SGH^{+}20, TSG25]. dussumieri [SNJ^{+}24]. dwarf [LZCC24]. dwarf-form [LZCC24]. Dynamic [HGC24, SGKAR24, BBPT^{+}22, BBPT^{+}24, LMK^{+}23, MMP^{+}24]. Dynamics [GGMMMV^{+}20, AJB20, ASO^{+}22, BPT^{+}25, CKDP^{+}20, DHCS23, FJS^{+}25, GSH22, GBB20, HCCC21, HBC^{+}22, HM25, HCK^{+}21, HKCW24, JHB21, JSG21, KHMC23, KHK^{+}20, KSI20b, aLBK^{+}21, LKSiI22, MDW^{+}21, MWJ^{+}24, MHD^{+}21, MKS^{+}22, MNS^{+}20, MOM^{+}25, MMM^{+}20, OYOO21, PPM^{+}23, PFdSBL25, RKD24, SH22a, SCCAM21, ŞGK^{+}20, SXMV^{+}21, SSP24, SCG^{+}24, SDBS21, SGKAR24, Szu22, TFC^{+}20, VTS^{+}22, WSL21, dJDM23].

Early [KKLM24, VA20, BMA^{+}20, DLZ^{+}25, ECK^{+}21, FGSD25, FNKY20, MDL^{+}21, MVLC^{+}20, OKKW20, SF20, SF22b, TLCD21, TLC^{+}22, YTH22]. East [BHB^{+}22, CTR^{+}21, NAS^{+}20, NNS^{+}22, BBM^{+}24, CLY^{+}22, MOA23, MWTH23, ODM20, YTHM20, CNE^{+}22, FHSC21, JHB21, MYKO23, NLW^{+}22, SSM^{+}23, WLZ^{+}21, YTHM20, ZZH^{+}24, BLHS20]. east-central [MOA23]. Easter [CGDTSA^{+}25, KMSJ^{+}25]. Eastern [BBPT^{+}22, KPS20, ÖÜÖG20, SFYM24, BM22, BM23b, BKM23a, BM24b, Bro25, BCOBB^{+}23, CW22, EWPB22, FJS^{+}25, GGMRC^{+}22, HMS^{+}22, KNP^{+}20, KBH22, MDW^{+}21, MWJ^{+}24, ML24, Mur20, Mur21, PAA^{+}24, PM23, PTK^{+}20, PDD^{+}22, PDA^{+}24, SXMV^{+}21, SFC21, SF22a, SJH^{+}23, SSKS21, SDV^{+}22, SC20, Szu22, TMN^{+}21, TSC^{+}20, Ten22, TFC^{+}20, TMDA22, VCPO21, XMLCMV24, ASWS^{+}21, BNTK23, CNDDAPMR21, DDCNMR21, KHK^{+}20, LCM^{+}23, LCDM^{+}24, MHH^{+}20, PGAG22, ŞGK^{+}20, SBB^{+}24, WGFM21]. eat [BR22, DDA^{+}20]. echo [RGG22]. echo-sounder [RGG22]. echosounder [BDA^{+}20, JM23, OUB^{+}22, PNGGO^{+}22]. eco [AUHK22, JK20]. eco-evolutionary [AUHK22, JK20]. Ecological [MVDH24, ARD^{+}23, BHST^{+}21, BRGB^{+}23, BHB^{+}22, FCMP23, HUK^{+}23, LHPR21, PFFdC22, SSI^{+}23, SSM^{+}23, TEO25]. ecology [Aks24, GMERCM^{+}24, IRJ^{+}22, LCM^{+}23, LPRB^{+}21, NNTM^{+}20, SHS20, SF22b, VCPO21, WHCF22]. Ecomorphological [NVRG^{+}21]. Economic [Apo25, SKST23, SRP^{+}22, dAdSR^{+}20, BDR^{+}20, BSSE21, BSA^{+}23, CEAL21, DMS22, Hut22, HBMC21, HSW25, JCL^{+}21, LCN^{+}20, MPSH25, NAS^{+}20, NK23, PHP^{+}20, PSS^{+}20, PVC^{+}22, RAE^{+}21, SGW^{+}20, SW24, ZSWK25]. economical [GTS^{+}21]. Economics [Béc20]. economy [WW21a]. ecoregion [CRCAF^{+}22, SBC^{+}22]. Ecosystem [CMIMS20, HF20, KDF^{+}25, RMNB^{+}21, AAM^{+}20, ALRA20, AAZ20, CGM^{+}22, Dik24, HPL^{+}24, HG20, KSL^{+}23, LdV22, LAG^{+}21, Mac25, MCK23, NAS^{+}20, PJNGJ^{+}22, SYZ^{+}25, TMSS^{+}23, WY20, XCB^{+}21]. Ecosystem-Based [HF20, CMIMS20, KDF^{+}25, AAZ20, HPL^{+}24, KSL^{+}23, LdV22, TMSS^{+}23]. ecosystems [ALW^{+}21, GP21, PFFdC22, TEO25]. Ecuador [SJTGAS23]. edge [BHH21, CKK^{+}20, FSS^{+}24, MOA23]. edible [ASD^{+}22]. Editorial [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25a, Ano25b, Ano25c]. EDM [SGKAR24]. eDNA [MTC^{+}22, VCG^{+}23]. eDNA-based [VCG^{+}23]. edulis [YTHM20, YTH22]. Edwards [CKD^{+}21, FBQA20]. edwardsii [LMM^{+}24, Mac25]. Eel [OCBJG20, BÁP^{+}23, BDA^{+}24, BÖN20, BLK23, CGT^{+}23, DJFU20, DKBF23, LYLC21, NK23, PRWK20, PMC^{+}24, SKJJ25]. eels [SW20, Sim23]. EEZ [AAPG21]. Effect [BWR24, EHBJ20, FG23, GKM^{+}23, JBL^{+}22, LR21, MdCG20, NLW^{+}22, WGFM21, WSWL25, AMHH21, BCS^{+}22, DC24, FWKR21, GGTÁVT^{+}20, Ham22, HEG^{+}23a, JMP^{+}21, LBD24, LTR20a, Lya20b, Lya20a, OCBJG20, PTK^{+}20, PAY23, RBM21, RS21, SJBT20, SMA^{+}24, YHC^{+}24, vdHBBR20]. effective [Bea21, BDM^{+}20, JWL^{+}24, JMS25]. Effectiveness [WOG^{+}25, BH23, KK22, KKC25, LLK^{+}22, LTE^{+}23, PGAG22, PK24]. Effects [ASDW24, AK23, dFBPL^{+}20, BM25a, CMRP20, EHE^{+}23, FG21, HPD^{+}22, IRJ^{+}22, KA24, KSI20b, LIA25, LDC24, MSW21, MCMM20, MKFF^{+}21, ORdlG^{+}24, PCM^{+}21, RKD24, RWFT25, SGD^{+}21, SdFZFJ21, TNTN20, TPD20, UAG^{+}23, AVACA^{+}23, BBPT^{+}22, BWB^{+}23, BHG^{+}24a, BKM23a, BM24a, CH21a, CKK^{+}20, CTIC23, CG21, DKD^{+}21, FNH^{+}21, GPWP20, GTS^{+}21, HSJ^{+}24, JCL^{+}21, JK20, LCB^{+}21b, LDM^{+}24, LKSiI22, MDL^{+}21, NFAL^{+}22, OOAF^{+}21, PMC^{+}24, RPL^{+}24, SCCAM21, SSRC24, SGW^{+}20, SHB^{+}23, SFJ^{+}23, Spa24, SBRM^{+}22, SM21, SBT^{+}20, TTYT24, TSPK24, TLCD21, TRS^{+}24, WGC^{+}19, WGC^{+}21, WEH^{+}25, dJDM23, vBMP^{+}23, KHS^{+}20]. Efficacy [CCC^{+}21, CHL^{+}20, DH20, BWB^{+}23, GP25, SHS20]. Efficiency [PPH21, BWR24, BLHS20, CSH^{+}21, DLKH22, DJFU20, GPWP20, IJS^{+}22, JMP^{+}21, KM23, KWE^{+}21, KA24, KKC25, KFH^{+}25, MMML24, MPSH25, MKH^{+}20, MRP^{+}23, NSQV22, NS23, PRCF22, PFGQ20, SAdC20, SSJ^{+}21, TNTN20, vORP23]. efficient [CTIC23, LdV22]. effort [AYTM21, AGB^{+}20, ASDW24, BTML20, BHNP22, BAYR^{+}24, DBDT21, DBGV^{+}22, FBB20, HEGR24, HEG^{+}23a, HLMV24, HCDB^{+}24, HCDB22, HCDF24, LRMH21, LWH^{+}23b, LLFL21, Lya20b, MMC^{+}24, MTX^{+}20, MDJP24, MOB^{+}22, PPM^{+}23, PHP^{+}20, PvZ22, PZL^{+}23, SKST23, TMP20, TDI^{+}21, TCL^{+}24, VP22, XMLCMV24]. efforts [KPK^{+}23]. Egg [KHK^{+}20, FG23, HW21, LCB^{+}21a, PSSFS24]. eggs [FG23, GBO^{+}20, KNO^{+}21]. Egyptian [SK21b]. either [SXM^{+}21]. EK60 [RGG22]. EK80 [RGG22]. Elasmobranch [GGTÁVT^{+}20, ACL^{+}20, CTM^{+}20]. elasmobranchs [BC20, DEM^{+}23, DAD^{+}22, GRG24, HS21, OPL21]. Elastomers [vBMP^{+}23]. electrical [RBM21, vORP23]. electroanaesthesia [RBS^{+}24]. electrodes [RBS^{+}24]. electrofishing [BBHF25, KHL^{+}24, PBB^{+}22]. electronic [AFR^{+}24, BKT^{+}20, BMC20, SKY^{+}24, WZX^{+}20]. ELEFAN [WZX^{+}20, WZS^{+}21]. element [LHH^{+}25]. Elemental [CNE^{+}22, BRR^{+}21, BMA^{+}20, CMTP^{+}21, CRS23, SSP^{+}22, TLC^{+}22]. elements [LPP^{+}20, TBÓ^{+}22]. Eleutheronema [SLW^{+}20]. Elevated [BKC21, JZQZ20]. elucidate [BMA^{+}20, CZ25]. Embracing [DQMV21]. Embryonic [JZQZ20, ALRB^{+}20, ANB^{+}24, ORdlG^{+}24]. emergent [FJS^{+}25, MMQ21]. emerging [GC21]. emissions [HA23]. emitted [CCKL^{+}20]. emphasis [FCSA21, LPRB^{+}21]. empirical [HHD^{+}20, SGKAR24]. enable [ECM^{+}24]. enabled [DNLM23]. enclosure [DJFU20]. encrasicolus [HLI^{+}20, VMI21]. end [DBV22, DBV23, RBM21]. endangered [ECM^{+}24, GJSW22, MWTH23, PBB20, dS21]. endeavour [LZY24b]. endemic [ZOZW22]. energy [CBD^{+}22, HGC24, HA23]. enforcement [WY20]. engagement [vdHR23]. engineering [ALRA20]. England [SXM^{+}21, EWPB22, MJC^{+}23]. English [ABF^{+}21, FMMC20, LCB^{+}21a]. Engraulis [HLI^{+}20, LIA25, MMG^{+}24, MVLC^{+}20, PAY23, DBL^{+}25, VMI21]. enhance [PK24]. enhancement [BF25, aFLpX^{+}21, LCG^{+}21, LKSiI22, LZW^{+}21, LDC24, MCS^{+}22, PBM^{+}23]. enhances [LPS^{+}25, MTC^{+}22]. Enhancing [MCS^{+}24, NZP^{+}21, MCGL^{+}25]. ensemble [AHB^{+}22, KMA23, SMA^{+}24]. ensembles [DBV22, DBV23]. Ensis [BBHF25, dAGCR21]. ensure [HLMV24]. Enteroctopus [NY23]. entrance [CSH^{+}21, TNTN20]. entrances [MTS^{+}21]. entrants [HSM21]. enumerate [WLG^{+}23]. envelope [CEAL21]. envelopment [KM23]. environment [ARD^{+}23, BHVB^{+}24, CBD^{+}22, HKG^{+}21, HSM^{+}25, HG20, LMK^{+}23, NVB^{+}23, PHP^{+}20, PSSFS24, PDD^{+}22, XDX^{+}23]. Environmental [DDCNMR21, HTSJ23, JYH21, NMJ^{+}24, PAY23, STH25, ZJJ^{+}25, AAVÁM23, AWC^{+}23, AVB^{+}23, dFBPL^{+}20, BÖN20, CMBL21, CG21, DAL20, GSH22, HSL^{+}22, KSI20b, LLFL21, LZCC24, LK25, MSW21, MPM22, MRE^{+}24, NLS21, NVB^{+}23, OCdMC24, OOAF^{+}21, PDC^{+}24, PDA^{+}24, RWB^{+}23, RRB24, SHH^{+}21, SRT^{+}20, SM21, WLZ^{+}21, WGFM21, YFJ^{+}25, YCC22]. epidemiological [GKC^{+}22]. Epinephelidae [BHST^{+}21, FMSA21]. Epinephelus [SACS23]. epipelagic [RCH^{+}21]. equation [CMD^{+}23, MSS^{+}21]. equatorial [SFC21]. Equilibrium [KSO^{+}21, DRSPTA20]. equipped [HKG^{+}21]. Erie [DAL20]. Erratum [KBPS22]. error [CHPT20, FCSA21, HW24, KA24, MUF^{+}22, NSRM22, PCJH^{+}21, WS24]. Erythrinidae [AGNS^{+}21]. Escape [RBHM24a, AIJ^{+}23, BBJ22, BBR^{+}22, BBSM24, CTM^{+}20, EMR^{+}22, GGTÁVT^{+}20, NBEI23, VS23, VRS^{+}22, YHC^{+}24, ZZH^{+}24]. escape-gap [BBJ22]. escapement [PW25, WW21b, YSB^{+}21]. escaping [BW23]. esculenta [YLX^{+}24]. Esox [ARD^{+}23, BLE^{+}22, DAR^{+}23, DSNK^{+}22, EBN^{+}23, EHE^{+}23, FG21, FDS^{+}23, HGS^{+}23, KMA22, LPG^{+}24, LRGB25, MHB^{+}23, NHE^{+}23, NRH^{+}23, PZG23, RDR^{+}23, SKBA23, SCS25, WBBG^{+}23]. Essential [PDE^{+}20, LVL25, RNKB23]. essentiality [DMS22]. Establishing [CBJ24, DSP^{+}23]. establishment [HBW21]. estimability [HK22]. estimable [Cla22]. estimate [AVMBEB22, BHNP22, CFP22, CC20, DMM^{+}21, DJFF23, GJSW22, KFO20, MSS^{+}21, PBRT22, RGN^{+}20, VP22, dJDM23]. estimated [BM23a, DLP^{+}24, GGG^{+}22, HGHH25]. Estimates [JRW^{+}21, AT20, ACS23, ASDW24, BBSM24, CMV21, CLM^{+}22, CQA^{+}24, DLKH22, DTSR22, DCR^{+}20, FFG21, FS20, HEGR24, HH20, HKCW24, JBČ^{+}22, KBH22, LWH^{+}23b, MBOCdAM24, MOB^{+}22, PDC^{+}23, PMC^{+}24, PPH21, TLAM25, TPD20, WBA23]. Estimating [AAG22, CPF20, CLY^{+}22, DCK^{+}22, GWMC21, HPPT24, HC22b, JCCAS^{+}21, JCL^{+}21, MDJP24, MBH^{+}22, RWT^{+}20, RPM^{+}21, SWH24b, Szu22, TDI^{+}21, TCL^{+}24, WSB22, WS24, AAT^{+}21, FJA^{+}20, GAB^{+}22a, HHJ^{+}22, HMC^{+}23, HHLL25, HWMVM23, KBPS21, KBPS22, LMP24, RWB^{+}23, SEM^{+}23a, SDV^{+}22, SJW^{+}22]. Estimation [AVB^{+}23, GRHHGM^{+}20, MRP^{+}23, SKW^{+}21, TYYK21, ZF21, ATA^{+}24, ALRB^{+}20, BMC20, BCOBB^{+}23, CYBW22, CMP20, CH22, DLKH22, DET21b, DAD^{+}22, ECK^{+}21, FCKG^{+}22, FOS^{+}21, FMLSP20, HIMP23, HW24, KHL^{+}24, KM23, KMA23, KK22, KA24, LCDM^{+}24, MCS^{+}24, MLCMdS23, MHL^{+}23a, MYKO23, NFAL^{+}22, PPC^{+}21, PHH^{+}23, PNGGO^{+}22, PCJH^{+}21, SIM^{+}24a, SIM^{+}24b, VK21a, VK21b, VBB20, VP23, WZS^{+}21, PVPN22]. estimations [HP23, SF20, UFYT23]. estimators [CFO23, Det23]. estuaries [AMdC^{+}20, CdSLP21]. estuarine [Bro25, BAC^{+}22, CHL^{+}20, DBDT21, LBLF20, MOTL25, OMG^{+}23, TC24]. estuarine-angler [BAC^{+}22]. estuarine-resident [CHL^{+}20]. estuary [CSDH^{+}23, PRCF22, SWLH20, XqRJ^{+}23, ZHM23, CRS23, GLA21]. Etelis [DOB^{+}24, WWF^{+}20]. ethological [LDS^{+}21]. Etrumeus [Dik24]. EU [AGB^{+}20, CCGR20, LdV22]. Eugerres [dAdCdO^{+}23]. Euphausia [ZZ22]. euphausiid [LD25]. Eurasian [VTSI^{+}24]. Europe [HLI^{+}20, LCD^{+}23, Lya20b, PRF^{+}21, TSS^{+}23]. European [AUC25, BDA^{+}24, BÖN20, DML^{+}20, EDA^{+}22, EFM25, FBB20, FSS^{+}24, FBR^{+}24, HGHH25, KZT^{+}23, LRW^{+}24, LPRB^{+}21, Lun25, Lya22, MAH^{+}22, MNPMM^{+}22, NOL23, NSM^{+}21, OCBJG20, PGD^{+}25, PRWK20, PFGQ20, RMNB^{+}21, RTHB25, SCGM^{+}21, SBC^{+}22, SDdMG^{+}20, SW20, SPC^{+}23, SSD^{+}20a, SRB^{+}25, SKJJ25, UAB^{+}21, UAG^{+}23, VCG^{+}23, WKSF20, ZOS^{+}23]. Euthynnus [BNTK23]. eutrophication [TSS^{+}23]. evaluate [ASJ^{+}20, CAYM^{+}23, DHB^{+}21, FDB^{+}20, HZZ^{+}20, LJB^{+}24, TMH23]. evaluated [vP20]. Evaluating [CFO23, CCC^{+}22a, GMK23, GP25, GSH22, HCK^{+}21, KKC25, KMC20, LAG^{+}21, MSC^{+}24, MPH21, PSSFS24, RBS^{+}24, SH22a, SJ24, SHC21, XCB^{+}21, XMLCMV24, YSB^{+}21, ZSWK25, ZOS^{+}23, ATA^{+}24, CSDH^{+}23]. Evaluation [BHD^{+}23, EVS^{+}23, GZLCRG25, HF20, HCDB22, LKSiI22, OUB^{+}22, TMN^{+}21, BDR^{+}20, BSSE21, BBPT^{+}24, BS20b, CZJ^{+}24, CH22, DHX^{+}23, DSC24, FS20, HK22, KSV^{+}22, KRH^{+}24, LDM^{+}24, PPH21, RAE^{+}21, SPC22, SFC21]. evaluations [CBHS24]. Event [DNLM23, CC20, PHH^{+}23, SKY^{+}24]. events [BSR^{+}22]. eversion [AGL^{+}24]. every [Peñ21]. Evidence [AUC25, ABF^{+}21, ÁHGCVAI22, CCRGC^{+}24, GLP^{+}20, KPWS21, Sch23, WSB24, AMSC20, BJK24, CWM^{+}23, GHAZ21, MMG^{+}24, MFR^{+}22b, NDRR20, OBB^{+}20, PBM^{+}23, RDR^{+}23, TMSS^{+}23, ZLXL20, YCC22]. evidences [VMI21]. Evolution [DEM^{+}23, YH21a, YH21b, AAVÁM23, PVA^{+}24]. evolutionary [AUHK22, JK20]. Ex [YRTP20, SJTGAS23, YLP^{+}23]. Ex-vessel [YRTP20, SJTGAS23]. examination [SSP24]. examine [GBWM22]. Examining [BPT^{+}20, CFB^{+}23, FNH^{+}21, dLH23, PUC^{+}23]. example [BDM^{+}20, BCOBB^{+}23, JWL^{+}24, LCW^{+}23a, PDD^{+}22, SDV^{+}22, WST^{+}23]. Examples [DML^{+}20]. exceptional [CW22, WCLN20]. excluder [CTM^{+}20]. excluding [BM24b]. exclusion [IAB20, QAS^{+}25]. Exclusive [SKST23]. existence [KFI^{+}21]. exogenonus [PAY23]. exoskeleton [AIM^{+}23]. exotics [PVC^{+}22]. expanding [SMLT24]. expansion [CCCM^{+}20, DPL20, Har21]. Expected [Fis25]. Experience [Mel23, BMOC22, CBN^{+}21, HSM21, SF20]. experienced [DLZ^{+}25]. experiment [AAG22, DJFU20, FGL22, HGS^{+}23, MCC^{+}22]. Experimental [GHAZ21, JMP^{+}21, FBPCC^{+}21, LM22, PRCF22, PPH21, vBMP^{+}23]. experiments [BKT^{+}20, BMC20, EMJ^{+}22, EM23, HTK^{+}24, MMQ21, MRP^{+}23, PPH21]. expert [GBWM22]. expertise [JDH22]. experts [FGTA24]. explain [HCCC21, KPUB22]. Explaining [SKBA23]. explicit [KMA23, MPP20]. Exploitation [GWGM24, AHX^{+}24, AVB^{+}23, NFdSJO25, PDJ24, PDF20, RFMS^{+}21, RPH20, RUHM20, TSPK24]. exploited [BCRI21, DQK^{+}23, FDS^{+}23, GFDN^{+}22, LVL25, MGC^{+}22, RTB^{+}21, WS20, WCC24, YHST22, vDJB^{+}23]. exploratory [KLS20]. explored [RTB^{+}21, SGW^{+}22]. Exploring [ALRB^{+}20, BBR^{+}22, DSC24, DMF^{+}21, HG20, JK20, LYLC21, MCK23, PFGQ20, SGKAR24, TSG25, VRS^{+}22, WW21a, YH21a, YH21b]. export [MCH21, RBHM24b, VKS^{+}25]. exports [Seu22]. expose [MJD^{+}21]. exposed [AGNS^{+}21, DH20, NGDC25, SSG^{+}22]. exposure [BKR^{+}22, CGB^{+}22]. extended [CSRL20]. Extending [LD25]. extension [DQK^{+}23, Fu22]. extensive [CAYM^{+}23, TDI^{+}21]. external [BBHF25, SSY20]. extractive [MOCGCC^{+}25, MPEBdR23]. extreme [BHVB^{+}24, LM22]. extremely [CCR24]. extruded [Lun25]. eye [GFDN^{+}22, LXC^{+}20, SSY20]. eyestalks [SDdMG^{+}20].

F. [PFFdC22]. face [WQGS25]. face-off [WQGS25]. facilitate [PRWK20]. facing [FBM^{+}21]. factor [MRC24]. Factors [AdlBCN25, CTM^{+}20, CFP22, DCL^{+}20, MHD^{+}21, TZL^{+}24, YTSS22, BÖN20, DAL20, GW21, HTK^{+}24, HL20, LS24, LZCC24, MPM22, MOI23, NMJ^{+}24, PNAPH24, PDC^{+}24, WGFM21, vdHBBR20]. FAD [BGL^{+}22, FFG^{+}20, PNGGO^{+}22]. FADs [TFC^{+}20]. fail [Pun23]. falciformis [SASB24, SFCG^{+}21]. Falkland [IAB20]. fall [EVS^{+}23, GW21]. false [ABBO20]. families [BHST^{+}21]. Family [MRC24, LLH^{+}25]. fangsiao [JZQZ20]. Farfante [AGNR^{+}21]. Farfantepenaeus [AGNR^{+}21, PFFdC22]. farm [WHR^{+}24]. farmed [LLC^{+}20]. farms [KDdOM^{+}22]. fast [LTT^{+}23, LIA25]. fat [ASD^{+}22, EPHDB24]. fate [UAG^{+}23]. fatty [ASD^{+}22, CZ25, PZG^{+}20, QMGRIU22, TNDM23]. fauna [CRSC22]. faunal [TKB^{+}21]. Fear [Sch23]. Feasibility [LLK^{+}22]. feasible [LDS^{+}21]. feature [JMS25]. features [KHC^{+}20, PDE^{+}20, VR20]. February [Ano20p, Ano21p, Ano22p, Ano22q, Ano23p, Ano24p, Ano25d]. fecund [EPHDB24]. Fecundity [KSI20a, DMM^{+}21, HHJ^{+}22, SSFL24]. feed [KKC25]. feedbacks [DQK^{+}23, JK20]. Feeding [GMERCM^{+}24, UPBH^{+}20, HSM^{+}25, TNDM23, BCRI21]. female [BMSM22, EWPB22, EPHDB24, FBQA20, LZCC24, MSS^{+}21, PZG^{+}20, SSFL24, SKD^{+}20, SWIRF21, TSPK24]. Fennoscandian [HKKa^{+}25]. feriata [TNTN20]. fertilized [FG23]. fewer [LD25]. fiber [WST^{+}23]. fiber-optic [WST^{+}23]. fidelity [MJD^{+}21]. field [BVR^{+}21, BKT^{+}20, DMNV^{+}24, FKS^{+}20, GGMRC^{+}22, SSM^{+}23, ZHC^{+}25]. fifty [GBWM22, SCW21]. fight [SCD^{+}22]. filamentosus [SKW^{+}21]. fill [LBLF20]. filter [Che21]. fimbria [CGC24]. Fin [AMdC^{+}20, BTB^{+}21]. financial [FGPVPPGG22]. find [FWL20]. findings [ALRB^{+}20]. fine [BSKL^{+}22, DHCS23, MWJ^{+}24, WPB^{+}20]. fine- [DHCS23]. fine-scale [BSKL^{+}22, MWJ^{+}24, WPB^{+}20]. fingerling [EVS^{+}23, ZSWK25]. fingerprint [CRS23]. Fingerprinting [KSJM^{+}20]. fingerprints [BMA^{+}20, FJJT^{+}21, SSS^{+}23]. Finland [LVP22, SHV20]. Finnish [JAN23]. FIPs [SK21a]. First [FMD^{+}24, GGTÁVT^{+}20, MSD21, NRH^{+}23, OBB^{+}20, PFC^{+}21, DHC^{+}20, KBB^{+}21, MGB24, MCS^{+}25, WSB24]. first-sale [WSB24]. First-season [NRH^{+}23]. Fish [Ash20a, BDA^{+}20, BLC^{+}21, CAGLT23, DBV23, ERS^{+}23, JBČ^{+}22, KBPS22, MOA23, MAH^{+}22, OS21a, Tho21, VK21a, WQGS25, WGW^{+}23, YH21a, AdABW^{+}22, AUHK22, APGG22, Alv21, AGNS^{+}21, APB^{+}20, ALW^{+}21, AUM21, AK23, ASJ^{+}22, AMdC^{+}20, ACP^{+}23, BWG^{+}21, BBČ^{+}21, dFBPL^{+}20, BLFT23, BTC23, BHVB^{+}24, BGBM22, BHST^{+}21, BČD^{+}21, BN21, BALBC23, BM22, BKM23a, BKM23b, BM24b, BM25a, CRSC22, CLB^{+}21, CGDTSA^{+}25, CKK^{+}20, CHPT20, CGT^{+}23, CGB^{+}22, CSH^{+}21, CDAK23, CTS^{+}23, CCC^{+}21, CCC^{+}22a, CBTH20, dSCCC^{+}22, CHL^{+}20, CHAY^{+}25, CBD^{+}22, CAH^{+}20, DCK^{+}22, DQK^{+}23, DSS^{+}23, DCS24, DET21b, Dik24, DCR^{+}20, DWS^{+}23, DHC^{+}20, DKD^{+}21, DJFF23, EBGE21, ESB^{+}24, EPHDB24, FG21, FGSD25, FRP22, FBALRR^{+}22, FWKR21, GVK^{+}23, GGTÁVT^{+}20, GKKL24, GWMC21, GFDN^{+}22, GPT^{+}21, HKG^{+}21, HBW21, HLCC22, HL21, HFKS20, HSPC21, HSL^{+}22, HMS^{+}22, HC22c]. fish [IKBL23, İTAD24, JSG21, JGU21, JCL^{+}21, JND^{+}23, KLS20, KMA23, KNP^{+}20, KDdOM^{+}22, KPWS21, KHGB25, KHPB20, KMO20, KBM23, KMC^{+}23, LVL25, LBLF20, LOFS22, LPG^{+}24, LZY24a, LTE^{+}23, LJX^{+}20, LWX^{+}20, LDS^{+}21, LYLC21, LYX^{+}21, LCG22, Lor22, Lya20b, MA20, MBOCdAM24, MRS^{+}25, MLS^{+}21, MSLMOC^{+}24, MBB^{+}23, MCH21, MKFF^{+}21, NLW^{+}22, NZP^{+}21, NVSG24, NBEI23, NFdSJO25, NVRG^{+}21, OSEF22, OMK24, OAM^{+}21, PÁEMC22, PYX^{+}20, PCK23, PKRL21, PdAMdM^{+}23, PPC^{+}21, PBB^{+}22, PDJ24, PDA^{+}24, RRSP^{+}24, RMRG22, RBHM24a, RWFT25, RPL^{+}24, RBH^{+}24, RSD23, RNKB23, RBG^{+}24, SSSF25, SHH^{+}21, SHV20, SFC21, SCGW24, SSKS21, Sch23, SWLH20, SECB21, SIM^{+}24a, SIM^{+}24b, SEM^{+}23a, ŠBB^{+}22, SGZD21, SGZH^{+}22, SdFZFJ21, SCHSC21, SCS25, SZXC25, SBT^{+}20, SAA23, SEM^{+}23b, SV23, TBF^{+}21, TFC^{+}20, TBÓ^{+}22, TPD20, TCVG20, TJS23, VR20, VA20]. fish [VCG^{+}23, VRS^{+}22, WKSF20, WGC^{+}19, WGC^{+}21, WASS20, WPB22, WSB22, WLG^{+}23, WSB24, XWD^{+}21, YAO^{+}23, YSB^{+}21, YiTM23, ZZC^{+}21, ZOZW22, ZYZ^{+}23, ZJJ^{+}25, ZOS^{+}23, dCHdMS^{+}23, dAdCdO^{+}23, dAdSR^{+}20, dSTMV20, vDJB^{+}23, vdHR23, DDA^{+}20, GVK^{+}23]. fish-aggregating [SFC21, ZZC^{+}21]. fish-attractive [APGG22]. fish-like [KMO20]. fish-markets [SCHSC21]. FishAgePredictioNet [İTAD24]. fished [EHBJ20, MJD^{+}21, WCGB22]. fisher [BMM^{+}21, BSA^{+}23, HSM21, LVL25, RBG^{+}24, SPC22, SSI^{+}23, SSM^{+}23]. fisherie [WCT^{+}20]. Fisheries [Che21, GHAZ21, HF20, MPSH25, PRF^{+}21, Pra24, RMNB^{+}21, WY20, AFR^{+}24, AAVÁM23, AMHH21, AMM^{+}22, AWC^{+}23, AAZ20, ASJ^{+}20, ABBO20, Apo25, ARD^{+}23, AAFLL^{+}25, AK23, BW20, BDR^{+}20, BMM^{+}24, Bea21, BJS^{+}22a, BJS^{+}22b, BSSE21, BSAP22, BALBC23, BKM^{+}23c, Bro24, BWOR23, BLK23, BAYR^{+}24, CJC^{+}20, CEAL21, CRL21, CCGR20, CCKL^{+}20, CTR^{+}21, CMIMS20, Cop24, CCCM^{+}20, CAYM^{+}23, CTCB22, CRF^{+}24, DRSPTA20, DPL20, DSC24, DML^{+}20, Det21a, DSP22, DEM^{+}23, DET21b, DMS22, DC24, DBGV^{+}22, DHCS23, DD25, EMR^{+}22, FTB^{+}21, FMLC^{+}22, FMD^{+}24, FGPVPPGG22, FMLSP20, FDB^{+}20, FFG^{+}20, Fis25, FBPCC^{+}21, FKS^{+}20, FHHH20, GC25, GP21, GBWM22, GML^{+}23, GAB^{+}20, HFMH20, HM25, Har21, HSM21, HSSD^{+}21, HEG^{+}23a, HLMV24, HJA^{+}21, HW24, HBMC21, IJS^{+}22, JC21, JYH21, KAB^{+}22, KSV^{+}22, KSL^{+}23, KHS^{+}20, KWC^{+}20]. fisheries [KWMA23, KFH^{+}25, KDF^{+}25, KFI^{+}21, KSI20b, LVL25, LCM^{+}23, LWH^{+}23a, LBD23, LCN^{+}20, LZY^{+}24c, LdV22, LMG^{+}24, LCHB^{+}24, MGB24, MSC^{+}24, MKS^{+}22, MPH21, MUF^{+}22, MKC20, MKS^{+}21b, MOI23, MB23, Mon24, MOTL25, MHL^{+}23b, MVDH24, MPP20, MKFF^{+}21, MBZSM20, Mun24, NOL23, NLS21, NLL^{+}25, NSQV22, NS23, NMS^{+}22, NFdSJO25, NB24, NC20, NNTM^{+}20, OPL21, ÖÜÖG20, VS23, OYOO21, ÖSL^{+}23, OWF^{+}23, PW25, PM23, Peñ21, PK24, PDM^{+}24, PBD25, PRK23, PBM^{+}23, PJP20, PDE^{+}20, PDD^{+}22, QSMG^{+}23, RAG22, RAE^{+}21, RUHM20, RNKB23, RL24, STH25, SK21b, SO25, SKST23, SBC^{+}23, SMK^{+}24, SBD^{+}22, SPC^{+}25, SW24, SSI^{+}23, SK22, SDV^{+}22, SLA24, SSM^{+}23, SRF^{+}24, SBB^{+}24, SAdC20, SFPR^{+}23, Spa24, SMKJ21, SLW^{+}20, SEM^{+}23b, TSPK24, TP24, TGG^{+}24, TYYK21, TMDA22, TZL^{+}24, VR20, VCPO21, VTS^{+}22, WW21a]. fisheries [WSL^{+}24, WXJ^{+}24, WST^{+}23, XMCC20, YTSS22, YGMJ20, YRTP20, YH21a, YH21b, YLS^{+}23, ZHM23, ZDF^{+}22, ZCFG23, dCHdMS^{+}23, dJDM23, vPDD^{+}25, Mel23]. fisheries-dependent [DBGV^{+}22, GML^{+}23, LVL25, LBD23]. fisheries-independent [BWOR23]. fisheries-related [LCHB^{+}24]. fishermen [BHST^{+}21, VCPO21]. Fishers [FBW^{+}21, AMM^{+}22, ALW^{+}21, BMOC22, DTSR22, MA20, MRS^{+}25, MFR22a, OOAF^{+}21, PvZ22, RFMS^{+}21, SMKJ21, dCHdMS^{+}23]. Fishery [AGB^{+}24, BHVB^{+}24, BBPT^{+}22, DMS22, LZY24b, AAM^{+}20, AHX^{+}24, ALRA20, ABF^{+}21, AAH^{+}23, AVCA22, APB^{+}20, ASWS^{+}21, AVACA^{+}23, ARD^{+}23, ASE21, AOA^{+}22, ASJ^{+}22, AdlBCN25, BBJ22, Béc20, BTR^{+}24, BGM^{+}23, BNL^{+}23, BaLK^{+}21, BGCCP22, BHG^{+}24b, BLHS20, BN21, BM23b, BBR^{+}22, BBSM24, Cad20, CTM^{+}20, COG22, CEAL21, CPF20, CHT20, CGC24, CZJ^{+}24, CHGC25, CDAK23, CSTR^{+}21, CRCAF^{+}22, CNDDAPMR21, CAH^{+}20, DHB^{+}21, DMNV^{+}24, DDCNMR21, DMF^{+}21, DSJG20, DBGV^{+}22, ELM20, EM23, ECM^{+}24, FWL20, FMMA20, FMMC20, Fre22, FKW^{+}22, GKC21, GP25, GDVBB^{+}20, GAB^{+}24, GPWP20, GKKL24, GC21, GBWM22, GFC^{+}22, GMPD23, GTS^{+}21, GFM^{+}23, GBO^{+}20, Ham22, HIMP23, HPS^{+}24, HdLHD22, HCWH23, HCDB22, HCDF24, IB20, JLYR24, JBL^{+}22, KM23, KWC^{+}20, KHMC23, KNP^{+}20, KBPS21, KBPS22, KLNB^{+}24, KK22, KFDE^{+}22, KPS20, KK21, KBB^{+}21]. fishery [KPUB22, LRMH21, LG21, LCMS^{+}22, LMJ^{+}23, LMM^{+}24, LLFL21, LPRB^{+}21, LCW^{+}23a, LHPR21, LFdB^{+}21, LAG^{+}21, Mac25, MINS21, MPM22, MSD21, MHL^{+}23a, MDMS21, MMP^{+}24, MMML24, MRG^{+}23, MZ24, MDJP24, MFM^{+}20, MOB^{+}23, ML24, MDC^{+}22a, MCMM20, MDC^{+}22b, MB20, MBP20, MFJ^{+}24, NLW^{+}22, NCB^{+}23, NFAL^{+}22, OOM^{+}23, OYOO21, PAA^{+}24, PMS^{+}23, PPM^{+}23, PHP^{+}20, PRA^{+}23, PKRL21, PASdCF23, PJNGJ^{+}22, PTK^{+}20, PHV^{+}21, PMS^{+}20, PBDM23, PCM^{+}21, PCC^{+}23, PBDM21, QAS^{+}25, RGG22, RBHM24a, RdLSdB^{+}21, RPH20, RTHB25, RPM^{+}21, RNKB23, RLO^{+}21, RSPE22, RGP^{+}23, RRH^{+}24, SSRC24, SXMV^{+}21, SFC21, SGW^{+}20, SJH^{+}23, SSP^{+}23, SB20, SPG^{+}21, SPC^{+}25, SAW^{+}20, SKS^{+}23, SJBT20, SHS20, SKBA23, SJ24, SGZH^{+}22, SCW21, SHC21, SCC^{+}22, SH22b, SZXC25, SGH^{+}20, SAHW22, SMLT24, SME^{+}24, TSC^{+}20, TBH^{+}22, TSC^{+}22, TMP20, TPW23, TNTN20, TTC^{+}25]. fishery [TCL^{+}24, VMFF^{+}20, WOG^{+}25, WLZ^{+}21, WPLF20, WKBMW24, YHC^{+}24, ZZH^{+}24, ZOS^{+}23, dLH23, vORP23, HPD^{+}22, KMC20, SK21a]. fishery-collected [RRH^{+}24]. fishery-dependent [HCWH23, RPM^{+}21]. Fishery-independent [AGB^{+}24, CZJ^{+}24, HdLHD22, HCWH23, TSC^{+}20, TSC^{+}22]. fishes [BPT^{+}20, BCM^{+}21, CKDP^{+}20, CSSB22, CNE^{+}22, DSP^{+}23, DNLM23, FS20, GdSPL21, KFO20, Lya22, MTE^{+}20, PRN^{+}24, SAA^{+}22, SJW^{+}22, WQGS25, WCC24, ZCM^{+}23]. fisheye [BM23b]. Fishing [GCK^{+}21, MRE^{+}24, SKY^{+}24, SBD^{+}22, SGZD20, SGZH^{+}22, WZS^{+}21, WCC24, ASB^{+}24, AMHR22, ASE21, BGL^{+}22, BTML20, BCSM20, BHNP22, BUG^{+}24, BCF^{+}23, CYBW22, COG22, CMA^{+}22, CDAK23, CBN^{+}21, CAAFH21, DGMG^{+}22, DBDT21, Dik24, EBdRPC24, FMLC^{+}22, FCML^{+}22, FM21, FBM^{+}21, FBW^{+}21, FDdCS^{+}20, GZLCRG25, GWMC21, GFC^{+}22, HMY25, HEGR24, HG20, HEG^{+}23a, HBE^{+}22, HVMEO25, Hut22, HBMC21, IOO^{+}24, JCL^{+}21, JK20, KKLM24, KWW^{+}21, KMA22, LCB^{+}21b, LNP25, LVP22, LNMA23, LLK^{+}22, LLS23, LZC^{+}21, LSZ^{+}23, LBD23, LLFL21, LYX^{+}21, LS24, LCL25, LFdB^{+}21, LTR20b, LRGB25, Lya20b, Mac22, MZ24, MDJP24, MCK23, MB23, MOTL25, MZSZVP^{+}23, MOB^{+}22, MKFF^{+}21, NCB^{+}23, NNS^{+}22, NVB^{+}23, ÖÜÖG20, ÖRS^{+}25, PVA^{+}24, PPM^{+}23, PHP^{+}20, PSS^{+}20, PFdSBL25, PGAG22, PZL^{+}23, PFGQ20, RNKB23, SFCG^{+}21, SSS^{+}23, SCD^{+}22, SBZ^{+}21, SML^{+}24, SRB^{+}25, SZS^{+}24]. fishing [SBT^{+}20, SSG^{+}22, SAHW22, TMN^{+}21, TTK^{+}25, TRWH23, TDI^{+}21, UPBH^{+}20, VKS^{+}25, VP22, WZL^{+}22, WSWL25, WHM23, WEH^{+}25, WKBMW24, XCB^{+}21, XMLCMV24, YTSS22, YCC22, YMS21, ZHC^{+}25, ZJJ^{+}25, dS21, vDJB^{+}23, vdHBBR20]. fishing-effort [HEG^{+}23a]. fishway [CTS^{+}23]. fitted [GML^{+}23]. fitting [FCSA21, PBG24]. fixation [GKM^{+}23]. fixed [BJHS^{+}23, CHM24, GRJW20, HEGR24]. fixed-shape [BJHS^{+}23]. fjord [NEBP^{+}23]. flag [SFCG^{+}21]. Flashing [SBB^{+}24]. flat [BAJB^{+}24]. Flatfish [UAB^{+}21, HML^{+}20, RBM21, RL24, SCHSC21, UVA20]. flatfishes [JMP^{+}21, TYYK21]. flathead [KKC25]. flatnose [KFO20]. flavescens [IAB20]. fleet [CGC24, CAAFH21, FGPVPPGG22, HSM21, HSM^{+}24, LCC25, MTC^{+}22, ÖRS^{+}25, PCCMOA^{+}24, PFdSBL25, PCF23, SASB24, SPC^{+}25, SRB^{+}25]. fleets [AAG22]. flesh [GKC^{+}22, RS21]. flesus [FSS^{+}24, HML^{+}20, KZT^{+}23]. flexibility [DBL^{+}25]. flexible [WGNM24]. flexigrid [IB20]. flindersi [BM22]. Flint [KSI20a]. float [KC22]. floating [HPL^{+}24, ZCM^{+}23]. floodplain [dFBPL^{+}20]. floodplains [PASdCF23]. Florida [ASJ^{+}22, BBR^{+}22, BBSM24, CFB^{+}23, GAB^{+}24, HMR24a, HBMC21, HSW25, KAC^{+}23, MOA23, MCC20, RBHM24a, WCGB22]. flounder [BAF23, Dra22, FSS^{+}24, KZT^{+}23, LKSiI22, SXM^{+}21, TLCD21, TLC^{+}22, UFYT23, WMSW22, YMYH20]. flow [ÁHGCVAI22, CPPK23]. flowing [SNJ^{+}24]. flows [Hut22]. fluctuating [FGL22]. fluctuations [LZW^{+}21]. fluid [SDBS21]. Flume [HVMEO25, SDBS21]. fluttering [LTT^{+}23]. fluviatilis [EHE^{+}23, SSD^{+}20a, VTSI^{+}24]. fly [GFC^{+}22, JHM22, OOM^{+}23]. fly-dragging [JHM22]. fly-shoot [OOM^{+}23]. flying [WHCF22]. flyshooting [VS23]. focus [GPASM22]. Focusing [AL22, DBV22, DBV23]. foils [WGNM24]. foldable [PFGQ20]. follicle [CKDP^{+}20]. following [AWVS21, CAYM^{+}23, EMR^{+}22, LYX^{+}21, Pun23, SFCG^{+}21, SZXC25]. Food [Dra22, RRSP^{+}24, DDA^{+}20, EHE^{+}23, FKS^{+}20, KTFY22, KLS20, LPS^{+}25, NRH^{+}23, SK21b, DBL^{+}25]. footgear [ASWS^{+}21]. footprint [JHM22, YCC22]. footrope [DCK^{+}22, RBM21]. Forage [RMRG22, BWG^{+}21, LBLF20, TBF^{+}21]. foragers [FBW^{+}21]. foraging [IRJ^{+}22, Žák21]. forbesii [LCB^{+}21a]. forecast [HC22c]. forecasted [MBP20]. Forecasting [OMK24, BDA^{+}24]. forecasts [DBS^{+}21, SS23]. foreland [ZOS^{+}23]. Forensic [OPL21]. forest [Aks24, GC25, LZX^{+}20, YMS21]. forgotten [LHH^{+}25]. form [LZCC24]. formalin [GKM^{+}23]. formation [PRRR23]. forms [AVMBEB22]. Forster [BÁP^{+}23]. foster [AWC^{+}23]. found [SAA^{+}22]. foundations [WHR^{+}24]. four [ABK^{+}21, BGG^{+}22, CHAY^{+}25, FBB20, IB20, JMP^{+}21, PW25, PRN^{+}24, SNJ^{+}24, SYZ^{+}25]. four-panel [IB20]. fourfinger [SLW^{+}20]. Fourier [BHB24, DOB^{+}24, LTT^{+}23, LIA25, PJSQ20]. Fourier-transform [PJSQ20]. fraction [AIW^{+}23]. fragmented [ACP^{+}23]. framework [BGCCP22, CTIC23, DSP^{+}23, DBV22, DBV23, HKG^{+}21, İTAD24, LBD23, PDD^{+}22, SM21, WSL^{+}24]. France [WSB24, ZDF^{+}22]. Fraser [ZF21]. free [APGG22, BGL^{+}22, LLS23, WKSF20]. free-swimming [BGL^{+}22, WKSF20]. freezer [PCF23]. French [BTR^{+}24]. frequencies [CPM21, LD25]. Frequency [WZX^{+}20, ASJ^{+}20, LIA25, MOB^{+}22, RGN^{+}20, SKW^{+}21, SFMA23, SV23]. fresh [dHJTCE23]. freshness [HZZ^{+}20]. freshwater [BALBC23, CH21a, CGT^{+}23, CGB^{+}22, HSL^{+}22, HGS^{+}23, HBMC21, IRJ^{+}22, LTE^{+}23, MNS^{+}20, MCC20, PNRS23, RWB^{+}23, SHV20, SPE^{+}23, SPC^{+}23, TCL^{+}21, VTS^{+}22, WLG^{+}23, ZOS^{+}23, dSTMV20]. frisii [RKN23]. front [DBV22, DBV23]. frontier [NS23]. fronts [Pra24]. frozen [DH20]. fry [FGL22, MMQ21, ZSWK25]. FT [DOB^{+}24, PJSQ20]. FT-NIR [DOB^{+}24]. FT-NIRS [PJSQ20]. Fuel [BGL^{+}22, ZHM23]. Fujian [XqRJ^{+}23]. Fukui [PFGQ20]. fulgens [PRA^{+}23, VAVQGD^{+}20]. Full [Peñ21]. Fully [SHC21]. function [JSG21]. functional [AVMBEB22]. functionality [BBG^{+}24]. functions [GWMC21, KBPS21, KBPS22]. fundamental [MTS^{+}21]. Fundamentals [Cop24]. funded [HdLHD22]. furcatus [NFC20]. furnieri [ABT^{+}24, AAR^{+}21, HCCC21]. furrow [MGB24]. future [BBG^{+}24, FBR^{+}24, GdSPL21, HEG^{+}23b, MAA^{+}20, Pun24, RWB^{+}23, SBC^{+}23, TRN^{+}23, ZCFG23]. fyke [OCBJG20, PFGQ20, PMC^{+}24, WEH^{+}25].

gadid [LADA^{+}22]. Gadiformes [KFO20]. gadoid [BLHS20, BMA^{+}20, IB20]. Gadus [BWB^{+}23, CWRR24, CSH^{+}21, EM23, HMM^{+}21, HOQ20, JBL^{+}22, MHH^{+}20, NEBP^{+}23, PNAPH24, SJB^{+}20, SBL^{+}23, SF22b, SWIRF21, UPBH^{+}20]. gains [MMP^{+}24]. Galician [GMRRG20]. gallina [MPM22, ÖA21, PHV^{+}21, dAGCR21]. gamefish [TCL^{+}21]. gammarus [HGHH25, Lun25]. gap [BBJ22]. Gaps [ZCM^{+}23, BBR^{+}22, LBLF20, RBHM24a]. gas [HA23]. Gasterosteus [MHB^{+}23]. Gastric [HMR^{+}24b]. gastropod [CTCB22, MGC^{+}22, RLQ^{+}20]. Gastropoda [BQGV^{+}24]. Gaussian [BVR^{+}21, İTAD24]. GBS [YGMJ20]. gear [AWVS21, BKM23a, BS20b, BCF^{+}23, CJC^{+}20, COG22, CDAK23, CCC^{+}22a, Dik24, GWMC21, JHM22, LCC25, LS24, MLS^{+}21, MRP^{+}23, PHH^{+}23, RNP^{+}24, ŠBB^{+}22, SGD^{+}21, SMKJ21, SH22b, SEM^{+}23b, WSWL25]. gear-set [LS24]. gears [BRvL^{+}22, CSSB22, DHH^{+}22, GAB^{+}22a, HVMEO25, MOTL25, SBT^{+}20]. gemfish [ODM20]. gene [BQGV^{+}24, CPPK23]. general [DRSPTA20, SM21]. generalized [LCL25, Lor22, LK25, MMQ21, MDMS21]. Generalizing [MSS^{+}21]. generating [BVR^{+}21]. generation [BBG^{+}24, HMP^{+}22, KSO^{+}21, PDE^{+}20]. generations [RGG22]. generative [MZ24]. generic [BM23b]. Genetic [AGNR^{+}21, CKD^{+}21, HLI^{+}20, KNS^{+}22, KSV^{+}22, KPWS21, LLC^{+}20, LZW^{+}21, NDRR20, ÖA21, WBBG^{+}23, WBD^{+}21, WMSW22, APBN22, AVACA^{+}23, ÁHGCVAI22, Bea21, BJS^{+}22a, BJS^{+}22b, CPB^{+}21, CPPK23, DSB^{+}21, DSNK^{+}22, aFLpX^{+}21, GMRRG20, GAW^{+}22, HML^{+}20, JMS25, KMSJ^{+}25, MPV^{+}24, NA22, NTJN21, ODM20, PCBL23, PBM^{+}23, PCGG20, PNRS23, RFF^{+}22, WQGS25, WWO20, YGMJ20, YFJ^{+}25, ZLXL20, ZOZW22]. Genetic-based [KSV^{+}22]. genetically [PNRS23]. genetics [MSV21, TTC^{+}25]. Genidens [AMdC^{+}20]. Genome [CPB^{+}21, CPPK23, HIF^{+}24, YGMJ20, ZJY^{+}24, LLC^{+}20, RKN23]. Genome-wide [CPB^{+}21, CPPK23, HIF^{+}24, YGMJ20, ZJY^{+}24, LLC^{+}20, RKN23]. Genomic [KRH^{+}24, PBM^{+}23, PCBL23, SPC^{+}23, ZLXL20]. Genomics [CRF^{+}24, BJK24, LvCdGSL23, MGC^{+}22, PBD25, WPB^{+}20]. genotyping [YGMJ20]. genotyping-by-sequencing [YGMJ20]. Genus [AIM^{+}23]. Genypterus [BÁP^{+}23]. geochemistry [LMT^{+}22]. geoduck [dlTGPLC23, SCW21]. Geographic [DPL20, HIF^{+}24]. geographically [CG21]. geography [LXC^{+}20]. geolocation [HKG^{+}21]. geological [ZOS^{+}23]. Geometric [TP24]. Georges [WGFM21, BAW^{+}24, EWPB22]. Georgia [CZ25, KSI20a, AHX^{+}24, GQP^{+}25, ZZ22]. Geospatial [YFJ^{+}25]. geostatistical [HCK^{+}21]. German [BKM^{+}23c, DKBF23, ÖRS^{+}25]. Germany [BHB^{+}22, FBR^{+}24, SKBA23]. get [KLS20, vdHR23]. Ghana [AOA^{+}22, MA20, OYOO21, SBD^{+}22]. giant [BCM^{+}21, DMM^{+}23, GFC^{+}22, PNRS23, WWF^{+}20, WS20, HTSJ23, NY23]. gigas [BCRI21, DLZ^{+}25, LXC^{+}20, QMGRIU22]. gilberti [BFA^{+}21]. Gill [HMC^{+}23, CLD^{+}22, FMMC20, LVPS20]. Gillnet [GLA^{+}20, KMC20, BČD^{+}21, CCKL^{+}20, FM21, MB23, OAM^{+}21, SAHW22, TYYK21, WOG^{+}25]. gillnets [BC20, CDAK23, GMK23, HFKS20, JBČ^{+}22, KBB^{+}21, KSS^{+}22, PRCF22, RMD^{+}25, SV23]. gilthead [vBMP^{+}23]. girls [KLNB^{+}24]. gizzard [XqRJ^{+}23]. glaciale [KHGB25]. glacier [KHGB25]. gladius [FCKG^{+}22, MPM^{+}23, SB20, TPW23]. glance [IOO^{+}24]. glass [BDA^{+}24, Sim23]. glauca [MCC^{+}22, MCC^{+}23, PDC^{+}24, SCD^{+}22]. global [CCCM^{+}20, FMLC^{+}22, HWMVM23, PK24, SK21a, SJTGAS23]. glove [RBS^{+}24]. Glugea [KZT^{+}23]. Gmelin [NVSG24, SCN^{+}24]. goals [HFMH20]. Gobiidae [SYZ^{+}25]. goby [BBC^{+}20, DMZH21]. goes [CRF^{+}24]. golanii [Dik24]. golden [SDV^{+}22, YLX^{+}24]. gonad [CB20, JFS21]. gonad-based [JFS21]. Gonadal [FBQA20, CB20, FSS^{+}24, LZCC24, MDS^{+}20]. gonads [PZG^{+}20]. Gone [MB20]. Good [GBC23a, KDBOC25, KBK^{+}24, LMP24, Cop24, HCDB^{+}24, HSW25, KHS^{+}20, MPSM25, Mon24, Pun23]. Goodbye [PZG23]. gorbuscha [HKKa^{+}25]. Gordon [DQK^{+}23]. governance [AAVÁM23, LdV22]. governed [KWMA23]. GPS [DNLM23]. grade [LJB^{+}24]. gradient [KKLM24, PASdCF23, VBL^{+}24]. grahami [HLC^{+}25]. Grand [WCN^{+}24, CWRR24, FGPVPPGG22]. gray [PRRR23, PBDM21, CPR^{+}24, HMY25]. Great [AJB20, WPB^{+}20, WBD^{+}21]. Greater [SBC^{+}22]. greatest [LPRB^{+}21]. Greece [AHEV24]. green [FBB20, OMG^{+}23, PRA^{+}23, PFGQ20, SOS^{+}21, VAVQGD^{+}20]. greenhouse [HA23]. Greenland [BKHA21, FJJT^{+}21, FJHT^{+}22, Fre22, GLA21, JHB21, NEBP^{+}23, SC20, TTK^{+}25]. greenling [OMK24]. grenadier [BHH21]. grenadiers [MPP22]. Grey [GKKL24, KLBHK23]. grid [BLHS20, GRJW20, KFI^{+}21, OSEF22, VMFF^{+}20]. grid-summarization [GRJW20]. grids [Bro25]. ground [BKM23a, GSS^{+}23, RBG^{+}20, SML^{+}24]. ground-gear [BKM23a]. Groundfish [LD23, AHB^{+}22, CCR24, JLYR24, LWH^{+}21, MJC^{+}23]. groundfishes [HCWH23, LCW23b, RWT^{+}20]. grounds [BBHF25, CPB^{+}21, GSS^{+}23, LCB^{+}21a, LCD^{+}23, SSS^{+}23, TRWH23, VKS^{+}25]. grouper [ASJ^{+}20, ABK^{+}21, SR21, SACS23]. groupers [FMSA21, WSUN^{+}23]. groups [GMPD23, LZL^{+}24, ÖRS^{+}25, SKBA23]. Growth [CVM^{+}20, MGB24, MABR^{+}20, WS20, AUC25, BHH21, BBHF25, BGLP21, BHG^{+}24a, BCM^{+}21, CMD^{+}23, CGM^{+}22, CSRL20, CW22, DLZ^{+}25, DMZH21, DAD^{+}22, EHBJ20, FG21, FDS^{+}23, FS20, FHE^{+}24, FNKY20, GLA21, GFDN^{+}22, HLCC22, HW24, HKCW24, JCCAS^{+}21, KTFY22, KHC^{+}20, KHMC23, KBH22, KKC25, LR21, LZL^{+}24, LCG^{+}21, MHH^{+}20, MCHA21, NRH^{+}23, OYOO21, PJSQ20, PUC^{+}23, PRRR23, PD25, RdBAT^{+}23, SSV^{+}20, SSSF25, SR21, SdOR^{+}23, SKW^{+}21, SFJ^{+}23, SL24, SF20, SBBH24, SBRM^{+}22, SWH24b, SNHM23, TTYT24, TSS^{+}23, TdL24, VA20, VDCMRF^{+}21, WCLN20, WZX^{+}20, WGFM21, WZS^{+}21, WAA^{+}22, WPGO21, dSTMV20]. growth-selective [AUC25]. grunniens [CH21a]. Guajira [VCPO21]. Guam [STH25]. Guamá [FDdCS^{+}20]. guarding [GP25]. guards [OCBJG20]. Guatemala [KMSJ^{+}25]. guide [FMLC^{+}22]. guides [FCML^{+}22]. guilds [SdFZFJ21]. Gulf [AdlBCN25, CVM^{+}20, CRS23, EBGE21, GCK^{+}21, GLA21, HSM^{+}25, KPS20, Mac25, NB24, PBPM^{+}23, SZXC25, TLAM25, WPB22, AAM^{+}20, AHB^{+}22, AVACA^{+}23, ÁHGCVAI22, BAW^{+}24, BCG^{+}25, BRGB^{+}23, BWOR23, CH21b, DB22, EBdRPC24, GGTÁVT^{+}20, GPWP20, GSS^{+}23, GGL^{+}24, GGMRC^{+}22, GBWM22, JRW^{+}21, KSI^{+}22, KWC^{+}20, KPWS21, LADA^{+}22, MPSH25, MSLMOC^{+}24, PCM^{+}21, PHH^{+}23, SR21, SSKS21, SCG^{+}24, SRT^{+}20, SNHM23, SMLT24, VKS^{+}25, WASS20, NMJ^{+}24]. gunnari [ZZ22]. Günther [NTJN21]. Guyana [WS20]. Gymnocypris [DHC^{+}20].

H. [CKD^{+}21, FBQA20, HSM21, PDG^{+}22]. haanii [BaLK^{+}21, aLBK^{+}21]. habit [Dra22]. Habitat [HLZ^{+}20, LMG^{+}24, AAR^{+}21, BPT^{+}25, BWG^{+}21, BF25, BWOR23, ESB^{+}24, FHSC21, HEG^{+}23a, KWMA23, LP23, MDC^{+}22b, MKFF^{+}21, NVB^{+}23, NMS^{+}22, NRH^{+}23, PDC^{+}24, PD25, RWT^{+}20, SIMT24, SWLH20, SYZ^{+}25, TSC^{+}22, XqRJ^{+}23, YWC^{+}21]. habitat-based [TSC^{+}22]. habitat-specific [NVB^{+}23, RWT^{+}20]. habitats [GPT^{+}21, LVL25, LCW23b, PFGQ20, RNKB23, TMN^{+}21]. habits [HSM^{+}25]. haddock [SKD^{+}20, SGZD20, WGFM21]. Haemulon [NVRG^{+}21]. Hainan [RLQ^{+}20]. Hakaluki [BMA^{+}24]. hake [LMT^{+}22, MDL^{+}21, MNPMM^{+}22, QAS^{+}25, SCGM^{+}21]. half [RPH20]. halfbeak [KBPS21, KBPS22]. halibut [BKHA21, Fre22, GLA21, Hut22, JLYR24, LBW21, LWH^{+}21, MCS^{+}22, PRK23, SHC21, TTK^{+}25, TRN^{+}23]. Haliotis [HSM21, NAV^{+}23, PRA^{+}23, RGP^{+}23, VAVQGD^{+}20]. hammerhead [GAW^{+}22]. hammerheads [BFA^{+}21]. handline [BWB^{+}23]. handling [BSR^{+}22, GFC^{+}22, LCB^{+}21b, LTE^{+}23, RPL^{+}24, TCL^{+}21]. Haor [BMA^{+}24]. happens [PTD^{+}20]. harbor [BSKL^{+}22]. harbour [CCKL^{+}20, MB23]. hard [FCKG^{+}22, HKCW24, KFO20]. hard-to-age [HKCW24]. harengus [íKMíH^{+}22, RS21, SMC^{+}24]. harmful [Béc20]. harmonized [CCCM^{+}20]. Harnessing [DGMG^{+}22]. harvest [BBPT^{+}22, BBPT^{+}24, BGCCP22, BKM^{+}23c, CFB^{+}23, CLM^{+}22, DAL20, FWL20, GPWP20, HSSD^{+}21, HTSJ23, HHD^{+}20, JvPOG25, KWMA23, KMA22, LBP^{+}24, LWH^{+}23b, LMJ^{+}23, LMM^{+}24, Lya20b, Lya20a, MRS^{+}25, NNS^{+}22, PMS^{+}20, PBDM23, PDD^{+}22, SLF23, SCW21, WH23]. harvestable [EVS^{+}23, RBHM24a]. harvested [Lya20a, SKST23]. Harvesting [MOB^{+}23, MBP20, MOM^{+}25, XSS^{+}23, dAdSR^{+}20, MBB^{+}23]. hatch [BAF23, SF20, UFYT23]. hatcheries [GQP^{+}25]. Hatchery [RFF^{+}22, BTB^{+}21, BQBW20, BJS^{+}22a, BJS^{+}22b, GQP^{+}25, ZOZW22]. hatchery-origin [BQBW20]. Hatchery-produced [RFF^{+}22]. hatchery-reared [GQP^{+}25]. haul [ABBO20, LTR20b]. Hauraki [Mac25]. Hawai’i [WSL21, SKW^{+}21, SB20]. Hawaii-based [SB20]. Hawaiian [LG21, WPLF20]. Hawai‘ [BW20]. hazard [KBPS21, KBPS22]. headline [BGM^{+}23, BM25a]. heads [SCC^{+}22]. healing [SFJ^{+}23]. health [LPP^{+}20]. Heart [MTE^{+}20]. heavily [MCK23]. hedging [PvZ22]. Hediste [MOCGCC^{+}25]. hedonic [KAC^{+}23]. height [BM25a, ZZH^{+}24]. help [KPUB22, MJD^{+}21]. hepatopancreas [HLG^{+}21]. hepatosomatic [PNAPH24]. herding [MKH^{+}20]. Hermit [ML24]. herring [CRL21, FSP22, íKMíH^{+}22, MJD^{+}21, RMRG22, RS21, SMC^{+}24, TSI^{+}21, WYM^{+}25]. Heterocarpus [SCCAM21]. heterogeneity [HSM^{+}24]. heterogeneous [PJOR20]. heteroskedastic [PAY23]. Hickory [HRH22]. Hidden [Che21]. Hierarchical [JC21, SBBH24]. High [CPB^{+}21, NTJN21, CGM^{+}22, CBD^{+}22, DJFF23, GPASM22, GTS^{+}21, HEGR24, JLYR24, JWL^{+}24, KHMC23, KHPB20, LFdB^{+}21, MOB^{+}22, MMBH23, NDRR20, OB21, PVA^{+}24, PÁEMC22, PCBL23, RRSP^{+}24, RPH20, SECB21, TBH^{+}22, ZLXL20]. high-frequency [MOB^{+}22]. high-resolution [HEGR24, PÁEMC22]. high-rise [JLYR24]. high-value [KHMC23]. Highlands [WST^{+}23]. highlight [AAFLL^{+}25, FAK24]. highlights [FTH^{+}23]. Highly [CSB^{+}23, BAYR^{+}24, DLKH22, DSB^{+}21, PYX^{+}20, PBDM23, SJ24]. hilsa [ZHM23, FM21]. Himalayas [ASB^{+}24]. hinders [BDM^{+}20]. hinge [Sch23]. hippoglossoides [BKHA21, GLA21, SC20, TTK^{+}25]. Hippoglossus [JLYR24, LBW21, PRK23, TRN^{+}23]. histological [vBMP^{+}23]. histology [CB20]. historic [RDR^{+}23]. Historical [MHH^{+}20, CCCM^{+}20, CQA^{+}24, GMT24, KTFY22, MCPJET^{+}20, SSI^{+}23]. histories [CHAY^{+}25, XqRJ^{+}23]. history [AUHK22, ACP^{+}23, BMA^{+}20, BCM^{+}21, CBHS24, HCCC21, HQWD20, LPRB^{+}21, MVLC^{+}20, NFC20, PDG^{+}22, Pun23, RdBAT^{+}23, SSFL24, SNHM23, TLV23, TEO25, VA20, VBL^{+}22, WWF^{+}20, WCT^{+}20, ZYZ^{+}23]. Hole [SM21]. Holothuria [AHEV24, CPPK23, RFF^{+}22]. holothurian [RAG22]. Holothuroidea [GBC^{+}23b]. Homarus [BAW^{+}24, EWPB22, GGL^{+}24, HGHH25, HSM^{+}24, JGG^{+}24, LPS^{+}25, LBP^{+}24, Lun25, MKS^{+}21a]. home [CHB24]. Honduras [CPF20]. Hook [HGC^{+}21, SCS25, AHL20, AK23, BW20, CCC^{+}22a, CLD^{+}22, CWM^{+}23, GPWP20, KC22, NMJ^{+}24, TCL^{+}21]. hook-size [GPWP20]. hooked [CCC^{+}22a, CLD^{+}22]. hooking [AHL20, HGC^{+}21, KC22, LPG^{+}24, SGW^{+}22, TCL^{+}21]. hooks [BCSM20, CCC^{+}21, CCC^{+}22a, FMD^{+}24, LTE^{+}23]. Hoplias [AGNS^{+}21]. Hoplostethus [BCOBB^{+}23]. horizontal [BKM23b]. horizontal-separator [BKM23b]. horneri [MKS^{+}21b]. horse [DCL^{+}20, GFDN^{+}22, MPV^{+}24]. horse-eye [GFDN^{+}22]. host [PNAPH24]. hosts [FRP22]. hot [LNR^{+}21, OB21]. hotspot [WJN^{+}25]. households [FBM^{+}21]. hovering [LLH^{+}25]. hubbsi [TNDM23]. Human [LPP^{+}20, NNTM^{+}20, AL22, CRL21, DQK^{+}23]. Human-induced [NNTM^{+}20]. human-natural [CRL21]. Humboldt [DSB^{+}21, GPASM22, GBO^{+}20]. Hungary [SSJ^{+}21]. hungry [KLS20]. Hunt [KMSJ^{+}25]. Huron [DMZH21, TPD20]. hybridisation [BJK24]. hybridization [PBM^{+}23]. hybrids [HML^{+}20]. hydraulic [MMBH23]. hydraulics [LP23]. Hydroacoustic [BLFT23, DKD^{+}21, WKSF20, ZZC^{+}21, DCR^{+}20, DQMV21, EBGE21, LNW20]. Hydroacoustics [LM22, JBČ^{+}22]. Hydrodynamic [TTL^{+}20]. hydrodynamics [BCF^{+}23]. hydroelectric [dAdSR^{+}20]. hydropower [RBH^{+}24]. hyperstability [FWL20, MDC^{+}22b].

Hypophthalmichthys [aFLpX^{+}21, SLW^{+}21]. Hyporhamphus [KBPS21, KBPS22]. Hyporthodus [SR21]. Hypoxia [CKK^{+}20].

Iberian [AMM^{+}22, MSC^{+}24, CRCAF^{+}22, GGMRC^{+}22, PIP^{+}22]. Ice [LCB^{+}21b, SCS25, AHL20, BLC^{+}22, SLF23, TTK^{+}25]. ice-angled [AHL20, BLC^{+}22]. Ice-fishing [LCB^{+}21b]. icefish [ZZ22]. icefishes [LGD^{+}20]. Iceland [JHB21, Ken21]. Icelandic [SJB^{+}20]. iconic [BAC^{+}22, DMM^{+}23, HMS^{+}22]. ICP [AMdC^{+}20, BTB^{+}21]. ICP-MS [BTB^{+}21]. ICP-MS/MC [BTB^{+}21]. Ictalurus [BKR^{+}22, NFC20]. ID [WQGS25]. Identification [dSCCC^{+}22, HLG^{+}21, LCB^{+}21a, PVA^{+}24, QMC^{+}22, AFR^{+}24, Bea21, BJS^{+}22a, BJS^{+}22b, CTR^{+}21, dSRFFN^{+}20, JGU21, MPP22, MCHA21, OPL21, RLQ^{+}20, SV23, TP24, WQGS25, WJN^{+}25, YAO^{+}23]. identified [RdLSdB^{+}21, WYM^{+}25]. identify [APGG22, BBG^{+}24, KSJM^{+}20, TSG25]. Identifying [BCG^{+}25, Fre22, ÖRS^{+}25, PIP^{+}22, QSMG^{+}23, CHPT20]. identity [PSSFS24]. if [Cla22]. ignobilis [DMM^{+}23, GFC^{+}22]. ignore [LHH^{+}25]. ignoring [MJC^{+}23]. II [MAA^{+}20, SCG^{+}24]. Illex [OBB^{+}20]. illuminated [BGM^{+}23]. Illumination [OSEF22, JLYR24, LWH^{+}21]. illustrated [SJW^{+}22]. illustrative [PDA^{+}24]. image [DMM^{+}21, HHJ^{+}22, KSJM^{+}20, MCS^{+}24, RRH^{+}24, UVA20]. image-based [DMM^{+}21, RRH^{+}24]. ImageJ [APB^{+}20]. imagery [MOB^{+}22]. images [HZZ^{+}20, İTAD24, PPC^{+}21, TRWH23]. Imaging [JGU21, HL21, KSS^{+}22, LBLF20, SEM^{+}23a]. immobilisation [RBS^{+}24]. Impact [AHL20, DLZ^{+}25, LNP25, MOTL25, ZYZ^{+}23, AAM^{+}20, BSA^{+}23, BKR^{+}22, EDA^{+}22, HLG^{+}21, LWH^{+}23a, LWH^{+}23b, LK25, MMC^{+}24, MB23, MRUG^{+}23, NVRG^{+}21, PDF20, PTD^{+}21, RAE^{+}21, RRB24, SB24, TMSS^{+}23, ZCM^{+}23]. impacted [TCVG20]. Impacts [DBGV^{+}22, NVB^{+}23, Apo25, BDR^{+}20, BCSM20, COG22, CGB^{+}22, DPL20, Dik24, FAK24, Har21, HG20, MJC^{+}23, PJNGJ^{+}22, RBH^{+}24, SHH^{+}21, SH22b, TMN^{+}21, XCB^{+}21, XMLCMV24]. impairment [AGNS^{+}21, BSR^{+}22, BKR^{+}22, LBP^{+}24, RPL^{+}24, RBHM24b]. impedance [CHPT20]. imperfect [SJW^{+}22]. imperiled [LYLC21]. Implant [vBMP^{+}23]. implantation [TRS^{+}24, ZDZR^{+}22]. Implementation [AVMBEB22, IAB20, KSL^{+}23, LYX^{+}21, QAS^{+}25]. Implementing [SPC^{+}25, SXM^{+}21, BBPT^{+}22]. Implications [ARD^{+}23, BCOBB^{+}23, CJC^{+}20, CLB^{+}21, CGSL22, CNE^{+}22, EPHDB24, FFG^{+}20, FBR^{+}24, JZQZ20, MGB24, MMC^{+}24, MPV^{+}24, NCB^{+}23, NB24, OYOO21, RBHM24b, SK22, SKS^{+}23, SXK^{+}24, TTC^{+}25, WS20, YGMJ20, AAM^{+}20, ALRB^{+}20, BHVB^{+}24, GC21, KWC^{+}20, LCDM^{+}24, LPP^{+}20, Lor22, MFR^{+}22b, NOL23, PVA^{+}24, PYX^{+}20, PBM^{+}23, RAG22, RDR^{+}23, SHS20, TLV23, dCHdMS^{+}23, KWMA23]. Importance [BCPH22, LVP22, SBD^{+}22, AAFLL^{+}25, PRF^{+}21, PSSFS24]. Importance-performance [LVP22]. important [DSP^{+}23, FMSA21, GFM^{+}23, KPWS21, MRP^{+}23, MKS^{+}21b, MTE^{+}20, NVRG^{+}21, PCK23, TLV23]. impoundment [LCN^{+}20]. improve [BM25b, DDDP21, JC21, KFH^{+}25, NFAL^{+}22, RMNB^{+}21, RRB24, SSKS21, WSL^{+}24]. Improved [BMC20, BM23a, AWC^{+}23, BaLK^{+}21, HL21]. Improvement [SK21a, SPC^{+}25]. Improvements [Det21a, Fis25, Sch23]. improves [BHB24, CMV21, CQA^{+}24, FCMP23, PMC^{+}24, SXM^{+}21, WHM23]. Improving [AAH^{+}23, BM23b, DBS^{+}21, DTSR22, LCFJ22, MINS21, MOB^{+}22, XMCC20, BBJ22, BDR^{+}20, BM22, DMM^{+}21, SDBS21, dJDM23]. impure [HSW25]. imputation [MCS^{+}24]. in-river [EMR^{+}22]. in-season [HC22c]. in-situ [DJFF23]. in-trawl [TRWH23]. in-water [CCC^{+}22a]. inadvertent [Sat23]. Incentives [GAB^{+}24]. Incidental [PMS^{+}23, BS20a, IAB20, QAS^{+}25]. inclined [OSEF22]. include [HSSD^{+}21, HPD^{+}22]. Including [PTL^{+}24, GPW^{+}20, PDF20]. Inclusion [HW24, MTX^{+}20]. income [HLCC22, RRSP^{+}24]. incomplete [Che21]. Inconnu [WBD^{+}21]. incorporate [RRH^{+}24]. incorporates [SM21]. Incorporating [BAC^{+}22, MKS^{+}22, MSS^{+}21, SPC22, WCT^{+}20, DBV22, DBV23, MMP^{+}24]. incorporation [TLCD21, TLC^{+}22]. Incorrect [MB25]. increase [KTFY22, MFR22a]. Increased [BQBW20, MBD^{+}21, NBD^{+}20, WH23, BWN^{+}21]. increases [CAAFH21, MMML24, PZL^{+}23, TdL24]. Increasing [CLM^{+}22, IJS^{+}22, Lya22, FHHH20, KLBHK23, LCN^{+}20]. increment [DHC^{+}20]. increments [ATA^{+}24, PUC^{+}23]. independence [PBG24]. independent [AGB^{+}24, BWOR23, CZJ^{+}24, HM25, HdLHD22, HCWH23, MMG^{+}24, TSC^{+}20, TSC^{+}22]. index [LBD23, LHPR21, MYKO23, PNAPH24, TCJ^{+}21]. Indexing [RBD^{+}22]. India [MRC24, SNJ^{+}24]. Indian [GAW^{+}22, GFC^{+}22, KSI^{+}22, AAFLL^{+}25, BQGV^{+}24, CKD^{+}21, CW22, GPW^{+}20, HL20, IKBL23, LK25, PVC^{+}22, PCGG20, SASB24, WWF^{+}20]. indicate [PBM^{+}23, SECB21]. indicated [FJJT^{+}21, JBČ^{+}22]. Indicating [Mac22]. Indications [SAA23, LNvD^{+}23]. indicator [KZT^{+}23, LBW21, SDC^{+}22]. indicators [DWS^{+}23, EYAO20, GVK^{+}23, MSLMOC^{+}24, ÖSL^{+}23]. indices [AAT^{+}21, BČD^{+}21, CBA^{+}24, CB20, FAK24, HdLHD22, KHK^{+}20, MTX^{+}20, MKS^{+}21a, TSC^{+}22, TMP20]. indigenous [AVCA22]. Individual [ASE21, FBW^{+}21, KPS20, BKC21, JK20, KHS^{+}20, KSJM^{+}20, MdCG20, PSSFS24, SWH24b, VDCMRF^{+}21]. individual-based [MdCG20]. individuals [LZCC24, PZG^{+}20]. Indo [FTB^{+}21, YFJ^{+}25]. Indo-West [YFJ^{+}25]. Indonesia [AIM^{+}23, ASJ^{+}20, DWS^{+}23, FMSA21, JvPOG25, KNP^{+}20, SWH^{+}24a]. Indonesian [DMF^{+}21, HHD^{+}20]. induce [KHMC23]. induced [ANB^{+}24, DQK^{+}23, FGL22, GKC^{+}22, NVB^{+}23, NNTM^{+}20]. industrial [PPM^{+}23]. Industry [HSM^{+}24, HdLHD22, JvPOG25, MSD21, MRP^{+}23, SCC^{+}22, UPBH^{+}20]. industry-based [MRP^{+}23]. industry-funded [HdLHD22]. Infection [KZT^{+}23]. infective [HKKa^{+}25]. infer [BGG^{+}22, CCCM^{+}20, HBE^{+}22]. inference [DMZH21, GRHHGM^{+}20, MB25, ZJY^{+}24]. Inferences [HQWD20, MBOCdAM24, SSKS21]. inferred [AAR^{+}21, ÁHGCVAI22, CMTP^{+}21, LXC^{+}20, MMM^{+}20, SIMT24, SCN^{+}24, SSP^{+}22, SSS^{+}23, WHCF22, YGMJ20]. Inferring [PJNGJ^{+}22]. Influence [BF25, BS20a, CLD^{+}22, CBHS24, CBR^{+}23, HCDF24, KC22, Lun25, NSRM22, NHE^{+}23, NMS^{+}22, PHP^{+}20, SK22, TCL^{+}21, YTH22, ZDZR^{+}22, dAGCR21, BKC21, CMBL21, CBN^{+}21, DHB^{+}21, HMY25, HCDB22, KWW^{+}21, LJX^{+}20, LSM^{+}23, LZW^{+}21, LP23, MPM22, MRE^{+}24, MDC^{+}22b, PDC^{+}23, SL24, SRT^{+}20, SEM^{+}23b]. influenced [NA22, vdHR23]. influences [AVB^{+}23, SKS^{+}23, SWIRF21]. influencing [CFP22, MHD^{+}21, VP23]. inform [BMM^{+}21, GC25, LCM^{+}23, LCLM23, LMG^{+}24, TP24]. information [Che21, DCS24, FCMP23, HIF^{+}24, PCK23, Ten22, vP20]. informed [PK24]. informing [MOB^{+}22]. informs [HSM^{+}24, JND^{+}23]. infrared [BHB24, DOB^{+}24, PJSQ20]. Ingestion [Mac22]. inhabiting [RDR^{+}23]. Initial [EMJ^{+}22, EM23]. injection [vBMP^{+}23]. injuries [BRvL^{+}22, LCHB^{+}24]. injury [GMK23, HGC^{+}21, KC22, LTE^{+}23, PPC^{+}23b, TCL^{+}21, UVA20]. Inland [LKSiI22, FKS^{+}20, HFKS20, MA20]. Inlet [Har21]. Inner [KPS20]. innovation [FKW^{+}22]. innovations [COG22]. innovative [CSH^{+}21, SPD^{+}24]. Input [DRSPTA20, Apo25, HW24, WW21a, WY20, XMCC20]. input-output [Apo25]. inputs [MSS^{+}21]. Inshore [SMKJ21, WOG^{+}25, ZLXL20]. insidious [PBG24]. Insight [KSI^{+}22, CPM21, LDS^{+}21, SWH^{+}24a]. Insights [BNTK23, BHVB^{+}24, COG22, FM21, GAB^{+}20, HFMH20, LPAE^{+}24, MCC^{+}22, RNKB23, SCN^{+}24, VCPO21, YTSS22, AVB^{+}23, CTR^{+}21, FMMA20, HJA^{+}21, LVL25, MCC^{+}23, SK21b, SO25, SLW^{+}21]. Instinctive [PNRS23]. instrumental [BSAP22]. insularis [GGMMMV^{+}20]. integrate [TDI^{+}21, TCL^{+}24]. integrated [BQBW20, CWC^{+}21, DBS^{+}21, DBV22, DBV23, KDBOC25, KHPB20, LRMH21, LMP24, MUF^{+}22, MPP20, PJMP22, PDE^{+}20, SMK^{+}24, SBZ^{+}21, SJW^{+}22, VBB20, VP23, WS24]. Integrating [BS20b, HM25, HKCW24, PRK23, RRB24, Fis25, SKW^{+}21]. Integration [JDH22, HSPC21]. integrative [FDB^{+}20]. intelligence [SKY^{+}24]. intensity [ZHM23]. intensive [MDC^{+}22a]. Inter [DKD^{+}21, TBF^{+}21, AAPG21, CGM^{+}22, CSRL20]. Inter-annual [TBF^{+}21, AAPG21, CGM^{+}22]. inter-oceanic [CSRL20]. Inter-vessel [DKD^{+}21]. interacting [CTR^{+}21, PDD^{+}22]. interaction [BCG^{+}25]. Interactions [AMM^{+}22, ALRA20, BBM^{+}24, CHT20, YiTM23, ABBO20, DSS^{+}23, LYLC21, MSC^{+}24, PCCMOA^{+}24, PCF23, PBDM21, RLO^{+}21, SHS20, SMA^{+}24, ZYZ^{+}23, vDJB^{+}23]. Interactive [RPL^{+}24]. Interannual [WHCF22, GBO^{+}20]. intercept [PF20]. Interconnected [vPDD^{+}25]. interdependencies [Apo25]. interfere [KLNB^{+}24]. interference [KDdOM^{+}22]. intermediate [GFM^{+}23, OCdMC24, dLH23]. Internal [BRvL^{+}22]. international [Ham22, MHH^{+}20, CSTdL25]. interplay [MABR^{+}20]. interpopulation [SSD^{+}20a]. interpretable [GC25]. interpretations [Cam23]. interpreting [MPSH25]. interspecies [HML^{+}20]. interspecific [BJK24, Mur20, PBM^{+}23, RdBAT^{+}23]. intervals [CAZN24]. intervention [CAYM^{+}23]. interventions [NA22]. interview [AMM^{+}22, DSP22, SSM^{+}23]. interview-based [DSP22]. Intra [CJC^{+}20, MKS^{+}22]. intra-annual [MKS^{+}22]. Intra-gear [CJC^{+}20]. intraperitoneal [ZDZR^{+}22]. introduced [GWGM24, YiTM23]. invaded [BN21]. invasions [KK21]. invasive [ALRA20, FBB20, GBB20, HKKa^{+}25, HSW25, NFC20, SCGW24, SMC^{+}24, SEM^{+}23b]. Inventory [CRSC22, MTE^{+}20]. inverse [Lor22]. invertebrate [NZP^{+}21, SME^{+}24]. invertebrates [MVDH24, PPH21]. investigate [BBČ^{+}21, CSRL20, CG21, HRH22, HVMEO25]. Investigating [AJB20, AIJ^{+}23, ASWS^{+}21, FOS^{+}21, GAB^{+}22b, LDS^{+}21, MUF^{+}22, SBL^{+}23, TBÓ^{+}22, WWO20, WSF22, SDdMG^{+}20]. Investigation [BGBM22, PCGG20, GKC^{+}22, HML^{+}20]. investigations [MFJ^{+}24]. investment [ANB^{+}24, KAB^{+}22]. investments [CCGR20]. involving [BMC20, BHB^{+}22]. Ionian [SPM^{+}24]. Iowa [GW21]. IPMs [SMK^{+}24]. Iran [TAA^{+}20]. iris [NAV^{+}23, RGP^{+}23]. Irish [BMM^{+}24, EHBJ20, LRW^{+}24, MHL^{+}23b, OOM^{+}23]. Island [CRSC22, CGDTSA^{+}25, CMBL21, RLQ^{+}20, YTH22, LWX^{+}20, WSL21, GLP^{+}20]. Islands [ASO^{+}22, IAB20, LG21, MTE^{+}20, NAV^{+}23, RGN^{+}20, SDV^{+}22, SBT^{+}20, KDF^{+}25, SC20]. Islas [SSI^{+}23]. isn’t [vdHR23]. isolated [CPPK23]. isolated-genetic [CPPK23]. isotope [ACP^{+}23, BRGB^{+}23, BBHF25, CZ25, SIMT24, WHCF22, WWT^{+}25, YLX^{+}24]. isotopes [AAFLL^{+}25, HTK^{+}24, LCFJ22, LXC^{+}20, SPW^{+}22]. isotopic [BÁP^{+}23, CMTP^{+}21, FHSC21]. Issue [Mel23]. Issues [APGG22, ECM^{+}24, LFdB^{+}21, MPSM25, PDC^{+}24]. Italian [DSP22]. Italy [CGSL22, ZDF^{+}22]. items [Mac22]. iterative [TMH23]. IUU [SBZ^{+}21].

JABBA [WCT^{+}20]. JABBA-Select [WCT^{+}20]. jack [GFDN^{+}22, HIF^{+}24, MYKO23]. January [Ash20a, VK21a, Ano20q, Ano21q, Ano22r, Ano23q, Ano24q, Ano25e]. Japan [ASO^{+}22, AIW^{+}23, FHE^{+}24, KM23, MYKO23, TMN^{+}21, TSI^{+}21, YTHM20, YTH22, YHST22, YMS21]. Japanese [FHE^{+}24, FNKY20, HIF^{+}24, KTFY22, KSI20b, LKSiI22, LIA25, MYKO23, PYX^{+}20, TTYT24, UFYT23, WHM23, YGMJ20]. japonica [YGMJ20]. japonicus [BM24b, BAC^{+}22, HIF^{+}24, HMS^{+}22, LIA25, MYKO23, RTB^{+}21, SSY20, WYM^{+}25]. Jasus [LMM^{+}24, Mac25, MCHA21, vZvdHCA25]. Java [ASJ^{+}20]. jaw [CCC^{+}21, CCC^{+}22a]. jaw-hooked [CCC^{+}22a]. Jersey [BSA^{+}23]. joint [SJ24]. Jonah [GC21, HKCW24]. Jorge [CVM^{+}20]. José [TLAM25]. July [Ano20r, Ano21r, Ano22s, Ano23r, Ano24r]. jumbo [BCRI21, LXC^{+}20, QMGRIU22]. June [OS21a, Ano20s, Ano21s, Ano22t, Ano23s, Ano24s]. Jurisdictional [ECM^{+}24]. just [KLNB^{+}24]. Juvenile [GLA21, BTB^{+}21, BM24b, CHB24, CVM^{+}20, DNLM23, FGSD25, FFG^{+}20, FAK24, JHB21, LVL25, OKKW20, PJSQ20, PGAG22, PFFdC22, RKD24, RGN^{+}20, RTB^{+}21, SJB^{+}20, SSKS21, SWLH20, SFYM24, SNHM23, SLW^{+}20, TBH^{+}22, UFYT23, WGC^{+}19, WGC^{+}21, YMYH20, ZDZR^{+}22, ZZC^{+}21]. juveniles [ABT^{+}24, LR21, SIMT24, SF20]. JZB [XCB^{+}21].

kōura [Mac25]. Katsuwonus [ATA^{+}24, AUM21, Ash20a, Ash20b, dSCCC^{+}22, ZZC^{+}21]. Kenya [FBM^{+}21]. Key [PBD25, MBE^{+}20, MAA^{+}20, OMG^{+}23, SJTGAS23, ZF21]. keys [BVR^{+}21]. keystone [RGP^{+}23]. khudree [BSR^{+}22]. Kiddi [CKD^{+}21]. kill [JSKM20]. killer [ABBO20, CTR^{+}21]. kin [Fis25, PTL^{+}24]. king [ASB^{+}24, AAH^{+}23, CVM^{+}20, EDA^{+}22, HRC23, KK21, LYH^{+}21, LDC24, LTR20a, LTR20b, MFM^{+}20, MRUG^{+}23, MSJ21, RL24, SDV^{+}22]. kisutch [LCHB^{+}24]. kite [TTK^{+}25]. km [PCBL23]. knifejaw [CW22]. knowledge [BMOC22, BHST^{+}21, FMLC^{+}22, FCMP23, GBWM22, LVL25, MBE^{+}20, SSI^{+}23, SSM^{+}23, VCPO21]. kob [BAC^{+}22]. Kona [WPLF20]. Korea [Seu22]. krill [KHGB25, MZ24, ZZ22]. Kudoa [GKC^{+}22]. kuruma [SSY20]. Kutum [RKN23]. Kuznets [YCC22]. Kyushu [FHE^{+}24, YTH22].

L. [AK23, BLE^{+}22, CMD^{+}23, EHE^{+}23, EHBJ20, FG21, LCB^{+}21a, LCD^{+}23, LR21, MPM^{+}23, ORdlG^{+}24, RDR^{+}23, SJB^{+}20, SF22b, SSJ^{+}21, WKSF20, vBMP^{+}23]. LA-ICP-MS [AMdC^{+}20]. labelling [ZDF^{+}22]. Labeo [SNJ^{+}24]. Laboratory [MFJ^{+}24, AAVÁM23, GRG24]. Labrador [MBZSM20, MB20, MBP20]. labrax [LRW^{+}24]. labrid [CGDTSA^{+}25]. Lack [PBG24, CPB^{+}21, MHD^{+}21]. lacking [LPAE^{+}24]. lacustris [LR21]. laevigata [HSM21]. lagoon [ALW^{+}21, ESB^{+}24, SPC^{+}25, dAdCdO^{+}23, FGSD25]. lagoons [DAR^{+}23, RDR^{+}23]. Lagrangian [KPUB22, Pra24]. Lake [BH23, DH20, WBD^{+}21, ZJJ^{+}25, AJB20, AGB^{+}20, GBB20, HSL^{+}22, HGS^{+}23, KSV^{+}22, LPG^{+}24, MMC^{+}24, NRH^{+}23, SHV20, SPE^{+}23, SBBH24, TAA^{+}20, TSS^{+}23, VTSI^{+}24, WPB^{+}20, BQBW20, DAL20, DMZH21, FFG21, FBM^{+}21, HJMS20, LCFJ22, MCH21, NCB^{+}23, NAS^{+}20, NNS^{+}22, NC20, NNTM^{+}20, OGFW24, PvZ22, SPW^{+}22, SBBH24, SSJ^{+}21, TPD20]. lake-run [TAA^{+}20]. lakes [CNE^{+}22, GW21, LOFS22, LCG^{+}21, NMS^{+}22, PRWK20, RGG22, VTSI^{+}24, VCG^{+}23, AJB20, EVS^{+}23, WPB^{+}20]. lalandii [MCHA21, vZvdHCA25]. lampreys [AJB20]. lance [SBL^{+}23]. land [NVRG^{+}21, PASdCF23]. land-based [NVRG^{+}21]. landed [OPL21]. landing [YHST22]. Landings [RLO^{+}21, ACL^{+}20, ABF^{+}21, ECM^{+}24, HMR24a, Ken21, KK22, PÁEMC22, PAA^{+}24, PKRL21, SCGM^{+}21, Žák21]. landlocked [PRWK20]. Landscape [CRSC22, MDC^{+}22b, TDI^{+}21]. Lane [CPF20, SFMA23, QMC^{+}22]. Lanka [ABK^{+}21, PKRL21, PDJ24, RKD24]. lanternfish [KHGB25]. Lao [RBH^{+}24]. Lar [TAA^{+}20]. Large [BMSM22, EYAO20, NFdSJO25, SKD^{+}20, Bea21, BBC^{+}25, BCM^{+}21, GBB20, JM23, KSV^{+}22, KFI^{+}21, LLC^{+}20, LP23, Lya20a, MMC^{+}24, MCC^{+}22, SAA^{+}22, WLZ^{+}21, WLG^{+}23, ZCM^{+}23, ZSWK25]. large-scale [Bea21, BBC^{+}25, KFI^{+}21, Lya20a]. Largemouth [LCB^{+}21b, CBN^{+}21, MSW21, RBS^{+}24]. larger [PZG^{+}20]. largest [AGB^{+}20, DDA^{+}20, SJH^{+}23]. Larimichthys [LLC^{+}20, WLZ^{+}21, YLP^{+}23, ZLXL20]. Larmichthys [SXK^{+}24]. larseni [CZ25]. larvae [AUC25, BAW^{+}24, LR21, DBL^{+}25, SF20, SMC^{+}24, TTYT24]. Larval [JHB21, LPS^{+}25, ANB^{+}24, FG21, FSP22, FGSD25, Lun25, UFYT23, VKS^{+}25]. lasting [CWM^{+}23]. lastridge [IB20, SHB^{+}23]. late [LM22]. Latitudinal [WCLN20]. latus [GFDN^{+}22]. lavaretus [WKSF20]. law [DSP^{+}23]. Lawrence [CRS23, GLA21, PCM^{+}21, SMLT24]. layers [BBČ^{+}21]. LB [HFKS20]. LB-SPR [HFKS20]. LBPA [CPM21]. LBSPR [CBHS24]. LC [Lun25]. LC-PUFA [Lun25]. lead [BAJB^{+}24, BS20a, CLB^{+}21]. lead-ahead [BS20a]. Leader [SSRC24, SGD^{+}21]. leading [EYAO20]. leads [BKR^{+}22]. leaps [GTS^{+}21]. learn [Pun23]. learned [LMJ^{+}23, Mel23, SMK^{+}24, TDJ^{+}21]. Learning [PNGGO^{+}22, BDA^{+}20, BHB24, BGBM22, CPL^{+}25, CWM^{+}23, GC25, HZZ^{+}20, JMS25, MCS^{+}24, NY23, OMK24, OUB^{+}22, PÁEMC22, PPC^{+}21, SIM^{+}24a, SIM^{+}24b, SPD^{+}24, SML^{+}24, SRB^{+}25, YAO^{+}23]. learning-based [MCS^{+}24, SML^{+}24]. lebranche [APBN22]. Led [SSG^{+}22, BWB^{+}23, SSG^{+}22]. left [HLC^{+}25, MAH^{+}22]. legal [ECM^{+}24, KSI^{+}22]. leidyi [SMC^{+}24]. LEngth [WZX^{+}20, ASJ^{+}20, ASJ^{+}22, FDS^{+}23, GBC^{+}23b, HFKS20, MRC24, SIM^{+}24a, SIM^{+}24b, SDV^{+}22, AVMBEB22, BVR^{+}21, CPM21, CSB^{+}23, CSRL20, CRCAF^{+}22, CHAY^{+}25, DSP^{+}23, DMF^{+}21, FSP22, Fu22, HW24, JFS21, LMP24, Lor22, MMP^{+}24, PÁEMC22, PTD^{+}21, RGN^{+}20, SKW^{+}21, SFMA23, SLW^{+}20, VP22, WKSF20, WSB22, CBHS24]. length- [JFS21]. length-at-age [CSB^{+}23, LMP24, SFMA23]. Length-Based [HFKS20, ASJ^{+}22, FDS^{+}23, SDV^{+}22, CPM21, CRCAF^{+}22, CHAY^{+}25, DMF^{+}21, Fu22, CBHS24]. length-frequencies [CPM21]. Length-frequency [ASJ^{+}20, SKW^{+}21, SFMA23]. length-inverse [Lor22]. Length-weight [MRC24]. length-within-age [MMP^{+}24]. lens [LXC^{+}20]. Lepidonotothen [CZ25]. Less [BČD^{+}21, LD25]. Lessons [LMJ^{+}23, PDM^{+}24, SMK^{+}24, TDJ^{+}21, BAJB^{+}24, Mel23]. Let [HEGR24, Tho21, Tho19]. lethostigma [BAF23]. leucas [GLP^{+}20]. leucichthys [WBD^{+}21]. Leucoraja [NDRR20]. leucospilota [CPPK23]. level [HEGR24, KM23, MMC^{+}24]. levels [GMK23, SWIRF21, TSPK24]. lewini [BFA^{+}21, GAW^{+}22]. licences [MFR22a]. Life [HQWD20, VBL^{+}22, AUHK22, ACP^{+}23, BMA^{+}20, BCM^{+}21, CSDH^{+}23, CBHS24, CHAY^{+}25, DLZ^{+}25, ECK^{+}21, GQP^{+}25, HCCC21, HGC24, IRJ^{+}22, LPRB^{+}21, MDL^{+}21, MVLC^{+}20, NFC20, OKKW20, PDG^{+}22, RdBAT^{+}23, DBL^{+}25, SF22b, SNHM23, TLV23, TLCD21, TLC^{+}22, TEO25, VA20, WWF^{+}20, WCT^{+}20, XqRJ^{+}23, ZYZ^{+}23]. life-history [BCM^{+}21, NFC20, PDG^{+}22, RdBAT^{+}23, TLV23]. lifespan [WZS^{+}21].

Lifetime [FJHT^{+}22]. light [AMHH21, LZC^{+}21, LDS^{+}21, NCB^{+}23, SSG^{+}22, TLAM25, ZZC^{+}21, ZHC^{+}25]. lights [EDA^{+}22, SSG^{+}22, ZHC^{+}25]. lightweight [TC24]. like [KMO20]. likelihood [SKW^{+}21]. likelihoods [CHGC25, FCSA21]. Limanda [Ten22]. limbatus [BRGB^{+}23]. limit [CBHS24, ELM20, Mau22, SHB^{+}23]. limitation [SSFL24]. limitations [Cop24, DH20, FHHH20, WHM23]. Limited [CTCB22, CBHS24, CHAY^{+}25, DMS22, HFKS20, LZY^{+}24c, LZX^{+}20, MOI23, MDC^{+}22a, RTHB25, SH22b, WKSF20, ZOS^{+}23]. limiting [PFGQ20]. Limits [NC20, BCS^{+}22, WH23]. line [BW20, FMMA20, KNP^{+}20, MBH^{+}22]. linear [HK22, LIA25, WKSF20]. linkage [WSF22, ZZ22]. linkages [RTB^{+}21]. linked [AWVS21]. Linking [OCdMC24, TMH23]. links [SM21]. Linnaeus [dSCCC^{+}22, MPEBdR23, OMG^{+}23, SCGM^{+}21, VMFF^{+}20, WBBG^{+}23, WPLF20]. lion [Mac22]. Lionfish [HSW25]. lions [QAS^{+}25, BCG^{+}25]. literature [FGL22, SB24, WY20]. Lithodes [CVM^{+}20, DMM^{+}21, HRC23, LTR20a, MFM^{+}20, MRUG^{+}23, SDV^{+}22]. Lithodidae [DMM^{+}21]. Little [MTC^{+}22, BNTK23, MB23]. live [BUG^{+}24, GC21, LTR20a, MSW21, MSJ21, RBHM24b]. live-bait [BUG^{+}24]. live-well [MSW21]. lived [BHB24, CFB^{+}23, HP23]. livelihoods [KMC^{+}23, SBD^{+}22]. liver [PNAPH24]. Living [AAVÁM23]. liza [APBN22]. Lobster [BHG^{+}24a, CSTdL25, AIM^{+}23, AMSC20, ANB^{+}24, BAW^{+}24, BHG^{+}24b, BF25, BBR^{+}22, BBSM24, CQA^{+}24, GAB^{+}24, GJGW23, HGHH25, HSM^{+}24, HMR^{+}24b, HMR24a, LPS^{+}25, LBP^{+}24, LMJ^{+}23, LMM^{+}24, Lun25, Mac25, MKS^{+}21a, ML24, MHL^{+}23b, Pun24, RBHM24a, RRH^{+}24, RRB24, SCG^{+}24, SRF^{+}24, Spa24, TSPK24, TdL24, YFJ^{+}25, dLH23, dLHER24, MMML24]. lobsters [CHB24, EWPB22, FBQA20, GAB^{+}24, GGL^{+}24, JGG^{+}24, KKC25, MCHA21, PDM^{+}24, PBD25, RBHM24a]. Local [BMOC22, HBW21, LP23, AAFLL^{+}25, BHST^{+}21, FFG^{+}20, JDH22, LOFS22, PPC23a, SSI^{+}23, SSM^{+}23]. location [KC22, LPG^{+}24, SGW^{+}22]. loci [BQGV^{+}24]. locomotor [BLC^{+}22]. logbook [Det21a, RFMS^{+}21, TYYK21]. logbooks [SBJ^{+}20]. loggerhead [SOS^{+}21]. logistic [HK22, JSG21]. loliginid [GMRRG20]. Loligo [LCB^{+}21a]. Long [BLE^{+}22, Cam23, GKC^{+}22, GGL^{+}24, HCCC21, HPS^{+}24, PW25, SJH^{+}23, SW20, VAVQGD^{+}20, BWN^{+}21, BTML20, BHB24, BBPT^{+}24, BBSM24, CPB^{+}21, CHL^{+}20, FWKR21, HP23, LTR20b, MBH^{+}22, MBD^{+}21, NBD^{+}20, PDM^{+}24, PMC^{+}24]. long-haul [LTR20b]. long-line [MBH^{+}22]. long-lived [BHB24, HP23]. long-range [BTML20]. Long-run [HPS^{+}24]. Long-term [BLE^{+}22, Cam23, GKC^{+}22, HCCC21, PW25, SJH^{+}23, SW20, VAVQGD^{+}20, BBPT^{+}24, BBSM24, CHL^{+}20, PDM^{+}24, PMC^{+}24]. longevity [BKHA21, CSRL20, CW22, HC22a, WCLN20]. longevity-based [HC22a]. longfin [CGT^{+}23]. longinaris [SS24]. longirostris [PDG^{+}22]. longitudinal [PASdCF23]. Longline [NMJ^{+}24, AMHH21, ABBO20, BTR^{+}24, CHT20, CTR^{+}21, FMD^{+}24, LCMS^{+}22, MRE^{+}24, PMS^{+}23, PPM^{+}23, SSRC24, SXMV^{+}21, SFCG^{+}21, SB20, SGD^{+}21, SGH^{+}20, TSC^{+}20, TTK^{+}25, WSWL25, XMLCMV24]. longliners [NFdSJO25]. longnose [GM21]. look [FWL20]. looking [JM23]. Lophius [CKM^{+}20]. loss [MLCMdS23]. losses [Béc20, DCK^{+}22, MRS^{+}25]. lottery [OWF^{+}23]. Low [BMSM22, BCS^{+}22, APBN22, ÁHGCVAI22, CW22, EDA^{+}22, HZZ^{+}20, KHPB20, SDBS21, TLAM25]. low-cost [HZZ^{+}20, TLAM25]. lower [KTFY22, KSI20a, Sim23, ZHM23, LvCdGSL23, PRN^{+}24]. LPUE [MDJP24]. Lucas [PDG^{+}22]. lucioperca [SSD^{+}20b, SSJ^{+}21, TSS^{+}23]. lucius [ARD^{+}23, BLE^{+}22, DAR^{+}23, DSNK^{+}22, EBN^{+}23, EHE^{+}23, FG21, FDS^{+}23, HGS^{+}23, KMA22, LPG^{+}24, LRGB25, MHB^{+}23, NHE^{+}23, NRH^{+}23, PZG23, RDR^{+}23, SKBA23, SCS25, WBBG^{+}23]. luminescent [BWN^{+}21, MBD^{+}21, NBD^{+}20]. luminescent-netting [BWN^{+}21, MBD^{+}21, NBD^{+}20]. lumpfish [Ken21]. lumps [Ken21]. lumpus [Ken21]. Lunella [KBH22]. lupus [GKM^{+}23]. lure [CBN^{+}21, GKC21, KLS20, LNMA23, TCL^{+}21]. luridus [Dik24]. Lutjanidae [BHST^{+}21]. Lutjanus [AVMBEB22, BPT^{+}25, CPF20, FGCB^{+}21, GAB^{+}22a, MFO21, MFR^{+}22b, NVRG^{+}21, PJSQ20, QMC^{+}22, SFMA23, SACS23, WCLN20, WSB22]. LWRs [DSP^{+}23].

M. [MPP22]. MA [BMM^{+}21]. MacCall [Det21a]. Machine [BDA^{+}20, PNGGO^{+}22, BGBM22, CPL^{+}25, GC25, MCS^{+}24, NY23, OMK24, OUB^{+}22, SPD^{+}24, SRB^{+}25, YAO^{+}23]. mackerel [CMTP^{+}21, DCL^{+}20, FHE^{+}24, GKC^{+}22, HGC24, HIF^{+}24, KNO^{+}21, KPS20, KPWS21, MPV^{+}24, MMM^{+}20, MYKO23, PYX^{+}20, WYM^{+}25, ZZ22]. Macro [OMG^{+}23]. macroalgae [MKS^{+}21b]. Macrobrachium [PNRS23]. macrocephalus [NTJN21]. Macrodon [AdABW^{+}22]. Macrognathus [TTC^{+}25]. macroinvertebrates [DCR^{+}20]. macrostructures [SdOR^{+}23]. Macrourus [BHH21, MPP22]. made [UPBH^{+}20]. maenas [FBB20, MOM^{+}25, OMG^{+}23, PFGQ20]. Mafia [CMBL21]. Magallanes [HRC23]. magellenicus [HC22b]. magister [RPH20]. magnet [GRG24]. magnitude [MKC20]. mahseer [ASB^{+}24]. Main [VR20, AAPG21, LG21, MBE^{+}20]. Maine [BAW^{+}24, GGL^{+}24, KWC^{+}20, NMJ^{+}24, SCG^{+}24]. maintaining [RBHM24a]. Major [LMT^{+}22, GMPD23, HSPC21, KSI20b, SSS^{+}23]. Makaira [WWO20]. makes [KHS^{+}20]. Making [GRJW20, LD25, DMS22, SPC22]. mako [OB21]. ’Malabar [SNJ^{+}24]. malabaricus [AGNS^{+}21]. Malacca [WY20]. Malawi [MCH21]. Malay [FTB^{+}21]. Malaysia [NTJN21, NLL^{+}25, WY20]. Male [BGK22, HSSD^{+}21, MCHA21, vZvdHCA25]. male-only [HSSD^{+}21]. males [BMSM22]. Mallotus [ASGG21, FJJT^{+}21, FJHT^{+}22, SBJ^{+}20, SGKAR24]. manage [DMS22, SPC^{+}25]. managed [NMS^{+}22]. Management [BBPT^{+}24, HF20, RMNB^{+}21, AAM^{+}20, AAVÁM23, AAZ20, AVCA22, ARD^{+}23, ASJ^{+}22, BMM^{+}21, BDM^{+}20, BaLK^{+}21, BGCCP22, BMJ^{+}24, BHG^{+}24b, CJC^{+}20, CGSL22, Che21, CFB^{+}23, CMIMS20, Cop24, CAYM^{+}23, DSC24, DCS24, DHB^{+}21, EBN^{+}23, ELM20, FTB^{+}21, FTH^{+}23, FBR^{+}24, Ham22, JZQZ20, JC21, KSV^{+}22, KSL^{+}23, KWC^{+}20, KWMA23, KDF^{+}25, KMC^{+}23, LCM^{+}23, LCLM23, LMJ^{+}23, LMG^{+}24, LAG^{+}21, Mac25, MGB24, MPM22, MJC^{+}23, MDC^{+}22a, MZSZVP^{+}23, MPV^{+}24, MFR^{+}22b, MKFF^{+}21, MBZSM20, NB24, NC20, ÖÜÖG20, PRA^{+}23, PFFdC22, PK24, PRF^{+}21, PDM^{+}24, PBD25, PRK23, PBM^{+}23, PTD^{+}20, PBDM21, PTL^{+}24, RAG22, RUHM20, RDR^{+}23, SK21b, SSP^{+}23, SK22, SSM^{+}23, SKBA23, SRF^{+}24, Spa24, TP24, TLV23, VR20, WHM23, WJN^{+}25, WXJ^{+}24, YTSS22, YGMJ20, YH21a, YH21b, ZOS^{+}23, CSTdL25]. managers [KWMA23]. Managing [Har21, HdLHD22, PW25]. mangrove [ASD^{+}22]. mangroves [MPEBdR23]. Mannar [KSI^{+}22]. manned [BWG^{+}21]. Manta [SJTGAS23]. many [Har21]. Maori [WCLN20]. Mapping [LVL25, vPDD^{+}25, ACP^{+}23, FSP22]. maps [Peñ21]. March [Ano20t, Ano21t, Ano22u, Ano23t, Ano24t, Ano25f]. Marías [SSI^{+}23]. Marine [DML^{+}20, FFG^{+}20, FJJT^{+}21, GGMMMV^{+}20, Mac25, PCCMOA^{+}24, RAE^{+}21, Spa24, AMM^{+}22, ALRA20, BMOC22, BCG^{+}25, CMV21, CTCB22, DSC24, DSP^{+}23, DSP22, DHCS23, FKW^{+}22, HIF^{+}24, JGU21, LVL25, LWH^{+}23a, LJX^{+}20, LWX^{+}20, LCL25, LCHB^{+}24, MCS^{+}22, MRS^{+}25, MGC^{+}22, PFFdC22, PK24, PDJ24, RLQ^{+}20, RMD^{+}25, SWLH20, SSP24, SYZ^{+}25, SLA24, SB24, TKB^{+}21, TEO25, TdL24, XSS^{+}23, BSA^{+}23, SHS20]. marine-dominated [SWLH20]. mark [Det23, DAD^{+}22, Fis25, GW21, HGHH25, SHS20]. mark-recapture [Det23, Fis25, GW21, PTL^{+}24, SHS20]. marker [dSCCC^{+}22, GAW^{+}22]. markers [AIW^{+}23, CKD^{+}21, PCGG20, RKN23]. Market [HSPC21, SCGW24, DDA^{+}20, MZSZVP^{+}23, RBHM24b, SK21b, SECB21, SK22, SJTGAS23, HLCC22]. markets [HBW21, HSPC21, KMSJ^{+}25, PPC23a, PKRL21, SCHSC21, WSB24]. marking [CDAK23, DH20, LR21, SV23]. Markov [Che21]. marks [BBHF25, SW20]. marlin [WWO20]. Marphysa [MOCGCC^{+}25]. Maryland [MWJ^{+}24, DWLT21, MDW^{+}21]. mass [CB20, JBČ^{+}22, LZW^{+}21, PHH^{+}23]. mass-based [CB20]. Massachusetts [WPGO21]. masses [LCB^{+}21a]. Mastacembelidae [TTC^{+}25]. Matching [KAB^{+}22]. mate [PNRS23]. material [LPAE^{+}24, SSRC24]. materials [CDAK23, TTL^{+}20]. maternal [ANB^{+}24, MDL^{+}21, PSSFS24]. mathematical [CHGC25]. Matías [AdlBCN25]. matrix [ZHC^{+}25]. matter [BMSM22]. matters [ATAS20, PZG^{+}20]. Maturation [JFS21, CCC^{+}22b, FSS^{+}24, OMG^{+}23, WPGO21, YTH22]. maturity [AVMBEB22, BHG^{+}24b, CB20, EWPB22, GKC21, GRHHGM^{+}20, KHMC23, LCM^{+}23, MGB24, MDS^{+}20, MDC^{+}22a, Mur21, SWH24b]. Mauritanian [DEM^{+}23]. Maurolicus [KHGB25]. mawsoni [CPB^{+}21, GMPD23, ZSDZ24]. Maximising [CHPT20]. Maximizing [KKC25, TJS23]. maximum [BKHA21, BCOBB^{+}23]. maximus [BGK22, EDA^{+}22]. may [ANB^{+}24, BAJB^{+}24, CLB^{+}21, GdSPL21, ODM20, Sch23, Ano20u, Ano21u, Ano22v, Ano23u, Ano24u]. MC [BTB^{+}21]. mean [RGN^{+}20]. measure [FMD^{+}24, HEGR24]. measurement [PFC^{+}21, PDC^{+}23]. measurements [BUG^{+}24, DJFF23, LIA25, RWFT25]. Measures [KFI^{+}21, AYTM21, GVK^{+}23, HLMV24, SZXC25]. Measuring [CEAL21, JHM22, KWE^{+}21, Ken21, MPSH25, APB^{+}20, CHPT20]. meat [KMSJ^{+}25]. mechanical [WSWL25, vORP23]. mechanics [VRS^{+}22]. mechanism [LZCC24]. mechanisms [HGS^{+}23, RMNB^{+}21]. mechanistic [JDP22]. medaka [XSS^{+}23]. media [AWC^{+}23]. medicine [NLL^{+}25]. medicine-based [NLL^{+}25]. mediocris [HRH22]. Mediterranean [AMSC20, BMOC22, PDG^{+}22, PDC^{+}24, SPG^{+}21, SBT^{+}20, VMI21, BNL^{+}23, CGSL22, DRSPTA20, Dik24, GP25, GDVBB^{+}20, LVPS20, MDMS21, MPM^{+}23, MABR^{+}20, MNPMM^{+}22, NSM^{+}21, NFdSJO25, ÖÜÖG20, PFC^{+}21, PDC^{+}23, PGD^{+}25, PAA^{+}24, PM23, PTK^{+}20, PHV^{+}21, RAG22, RCVGMI22, RBG^{+}20, SCGM^{+}21, ŞGK^{+}20, SPM^{+}24, SBB^{+}24, TZL^{+}24, dJDM23]. mediterraneus [PFC^{+}21, RCVGMI22]. meets [LFdB^{+}21, Sat23]. mega [SB24]. mega-decapod [SB24]. megafauna [AMM^{+}22, PCCMOA^{+}24]. Megalops [ECK^{+}21]. Meghna [ZHM23]. Mekong [KMC^{+}23, NTJN21, PRN^{+}24, RBH^{+}24, TTC^{+}25, VBL^{+}22, VBL^{+}24]. Melanogrammus [SKD^{+}20, WGFM21]. melanostictus [WHM23]. melanostromus [BBC^{+}20]. melastigma [XSS^{+}23]. memorable [JSKM20]. memorable-sized [JSKM20]. menhaden [LNW20]. mentella [CRS23, HP23]. menu [BCRI21]. Merlangius [BRvL^{+}22, VS23, ŞGK^{+}20]. merlangus [BRvL^{+}22, VS23, ŞGK^{+}20]. Merluccius [KNS^{+}22, LMT^{+}22, MNPMM^{+}22, SCGM^{+}21, TNDM23]. mesh [AIJ^{+}23, BW23, BLHS20, BM25b, CAH^{+}20, FWKR21, GTS^{+}21, LTT^{+}23, MOB^{+}23, VS23, SHB^{+}23, TTL^{+}20, VRS^{+}22]. meshes [BJHS^{+}23, BM24a]. mesocosm [MMQ21]. mesopelagic [KHGB25, SSSF25, WQGS25]. Mesoscale [FBALRR^{+}22]. messy [Aks24]. meta [BCOBB^{+}23, DSP^{+}23, OFF^{+}20, RDR^{+}23, RPM^{+}21]. meta-analysis [DSP^{+}23, OFF^{+}20, RPM^{+}21]. meta-analytic [BCOBB^{+}23]. meta-population [RDR^{+}23]. metabarcoding [HSL^{+}22, MTC^{+}22, RdLSdB^{+}21, YLX^{+}24]. Metabolomics [NAV^{+}23]. metal [TTK^{+}25]. metals [AMdC^{+}20]. metamorphosis [BAF23]. Metanephrops [MTS^{+}21]. metapopulation [CMV21, MPC^{+}20]. method [APB^{+}20, AMdC^{+}20, CSH^{+}21, CLD^{+}22, CH22, CB20, DMS22, DJFF23, EDA^{+}22, FG23, GRJW20, GJGW23, HBMC21, MPM22, OWF^{+}23, RGN^{+}20, SSY20, SIM^{+}24a, SIM^{+}24b, TMSS^{+}23, VK21a, VK21b, WYM^{+}25]. methodologies [DC24, LCW23b, SPD^{+}24]. Methodology [MCS^{+}25, HIKM21, TCL^{+}24]. Methods [MMBH23, SS23, BDA^{+}24, BRN^{+}20, Cop24, CBR^{+}23, CRCAF^{+}22, DHX^{+}23, DSJG20, ECK^{+}21, FJA^{+}20, GFM^{+}23, HMY25, HSL^{+}22, HKCW24, KHK^{+}20, LGD^{+}20, LTT^{+}23, LTR20b, MHL^{+}23a, MOI23, NDRR20, OMK24, PTD^{+}21, RRH^{+}24, SKW^{+}21, SSP24, SDBS21, SGZD20, SACS23, WSWL25, WASS20]. métiers [GMK23]. metrics [JFS21, LYH^{+}21, MTS^{+}21, PBDM21]. Mexican [ÁHGCVAI22, dlTGPLC23, SFCG^{+}21]. Mexico [ÁHGCVAI22, BRGB^{+}23, EBGE21, EBdRPC24, GRHHGM^{+}20, GPWP20, GGMMMV^{+}20, PBPM^{+}23, PHH^{+}23, SR21, SSKS21, SSI^{+}23, VKS^{+}25, WASS20, WPB22]. Michigan [DMZH21]. micro [Lun25, SdOR^{+}23]. microchemistry [FJHT^{+}22, HQWD20, LBW21, MNPMM^{+}22, VBL^{+}22, XqRJ^{+}23]. microlepis [KFO20]. Micromesistius [BMA^{+}20, MABR^{+}20]. microPIT [WGC^{+}21, WGC^{+}19]. Micropogonias [ABT^{+}24, AAR^{+}21, HCCC21]. Micropterus [KSI20a, OGFW24]. microsatellite [BQGV^{+}24, CKD^{+}21, HWW22, MPC^{+}20, PCGG20, SLW^{+}21, WMT^{+}20, WMSW22]. microscopic [OMG^{+}23]. microstructure [AUC25, BWB^{+}24, BMA^{+}20, DHC^{+}20, SF22b, SLW^{+}20]. Mid [SXM^{+}21, BLFT23, FMMA20, CG21]. Mid-Atlantic [CG21]. mid-north [FMMA20]. mid-water [BLFT23]. midwestern [ZSWK25, WW21b]. migrating [HL21, LCHB^{+}24, YTHM20]. migration [EMR^{+}22, JAN23, KNS^{+}22, KPWS21, MZZ^{+}21, PPC^{+}23b, PRWK20, SBJ^{+}20, VBL^{+}24, vZvdHCA25]. migrations [CLB^{+}21, CRS23, RBH^{+}24, TSI^{+}21]. migratory [HWW22, KPS20, PYX^{+}20, TTC^{+}25, XCB^{+}21, YTH22]. miles [Dik24]. mill [HMR^{+}24b]. million [MB20]. Milne [CKD^{+}21, FBQA20]. Min [XqRJ^{+}23]. mini [PK24]. mini-review [PK24]. Miniaturized [DNLM23]. minimise [RBG^{+}24]. Minimising [MRG^{+}23]. minimize [vBMP^{+}23]. minimizing [GP25]. minimum [CBHS24, KSI^{+}22, Lya20a]. minor [GKM^{+}23]. mis [TMH23]. mis-specification [TMH23]. misalignment [AJB20]. misclassification [MOI23]. Misidentification [APGG22, KNO^{+}21]. mislabeling [HOQ20]. mislabelling [DDA^{+}20]. mismatch [BAM^{+}24]. misreported [DHX^{+}23]. missing [ÖÜÖG20, PAY23]. misspecification [SH22a]. misuse [MB25]. Mitchill [SPD^{+}24]. mitigate [IAB20, SMLT24]. Mitigating [BC20, QAS^{+}25]. Mitigation [KFI^{+}21, FMD^{+}24, OB21, PCC^{+}23, SACS23]. Mixed [KWC^{+}20, TAA^{+}20, ARD^{+}23, BJS^{+}22a, BJS^{+}22b, BGM^{+}23, GMT24, HPPT24, KPS20, KPWS21, KFH^{+}25, MPH21, PPD^{+}25, SJ24, SBRM^{+}22, WXJ^{+}24, dJDM23]. mixed-effects [SBRM^{+}22]. mixed-origin [HPPT24]. mixed-stock [BJS^{+}22a, BJS^{+}22b, KPS20, KPWS21]. mm [ZDZR^{+}22]. Mnemiopsis [SMC^{+}24]. mobile [FG23, GAB^{+}22a, LLK^{+}22]. Model [AT20, Che21, DET21b, FAK24, SM21, TAK^{+}23, ATAS20, CBA^{+}24, CYBW22, CAZN24, CWRR24, CWC^{+}21, CSDH^{+}23, CAAFH21, CFP22, DQK^{+}23, DBS^{+}21, DMM^{+}21, DWLT21, DMZH21, DBV22, DBV23, EYAO20, Fu22, GRHHGM^{+}20, HC22b, HGC24, HLZ^{+}20, HKCW24, HPD^{+}22, HK22, KNO^{+}21, KDBOC25, KHPB20, LVL25, LBD24, LKSiI22, LAG^{+}21, LK25, MWJ^{+}24, MMQ21, MDMS21, MKS^{+}21a, MZ24, MVMdS^{+}24, OS21a, OS21b, OCdMC24, OOAF^{+}21, PPD^{+}25, PBG24, PJOR20, PTD^{+}20, RRB24, SBBH24, SXM^{+}21, SEM^{+}23b, TCJ^{+}21, VBB20, VP23, WCN^{+}24, XCB^{+}21, YIM^{+}20, dJDM23, dLH23]. Model-based [AT20, DET21b, FAK24, MKS^{+}21a, PPD^{+}25, PTD^{+}20, TCJ^{+}21]. modeled [BPT^{+}20]. Modeling [CC22, CRL21, CPL^{+}25, CFP21, DSP^{+}23, DMZH21, GPT^{+}21, HCDF24, MCC20, Bro24, JDP22, LZY24a, MCS^{+}22, MLCMdS23, MdCG20, MTX^{+}20, PGD^{+}25, SH22a, SYZ^{+}25, Szu22]. modeling-based [SYZ^{+}25]. modelled [BCOBB^{+}23]. Modelling [CQA^{+}24, NLS21, SFPR^{+}23, SGKAR24, BAM^{+}24, Fre22, HCDB^{+}24, HCK^{+}21, KMA23, MVMdS^{+}24, MPP20, PJNGJ^{+}22, PDA^{+}24, SMA^{+}24]. models [AVMBEB22, Aks24, AAT^{+}21, BP20, BCM^{+}21, CMP20, CC22, CGT^{+}23, CGC24, CFP21, FCML^{+}22, GPW^{+}20, GJSW22, GML^{+}23, HIMP23, HW24, JC21, KSL^{+}23, KBK^{+}24, LDM^{+}24, LCL25, Lor22, LZX^{+}20, MTX^{+}20, Mau22, MMP^{+}24, MVMdS^{+}24, MPP20, Mun24, NAS^{+}20, OUB^{+}22, PC21, PNGGO^{+}22, PCJH^{+}21, SMK^{+}24, SJ24, SFYM24, SFPR^{+}23, SBRM^{+}22, TDJ^{+}21, TMP20, TMH23, TAK^{+}23, TDI^{+}21, VDCMRF^{+}21, WAA^{+}22, WCT^{+}20, XMCC20]. modern [Mon24, OUB^{+}22]. modest [HW21]. modification [AAH^{+}23, FOM21, VK21a, VK21b]. modifications [KSS^{+}22]. modified [TC24]. modifier [DB22]. modify [BCF^{+}23, OSEF22]. Modifying [FMMC20]. Modular [MOB^{+}23]. modulate [HGS^{+}23]. modulated [LIA25]. Molecular [dSRFFN^{+}20, GKC^{+}22, HLG^{+}21]. molitrix [aFLpX^{+}21, SLW^{+}21]. Mollusca [RdLSdB^{+}21]. Molting [VDCMRF^{+}21, SSY20]. monacanthid [GFM^{+}23]. mongolicus [LCG^{+}21]. monitor [BTML20, SCC^{+}22, VCG^{+}23]. Monitoring [MBOCdAM24, SSKS21, AFR^{+}24, AYTM21, BSSE21, BHNP22, CHL^{+}20, DMM^{+}21, GRJW20, GLP^{+}20, LZC^{+}21, LvCdGSL23, MTC^{+}22, MPSH25, PÁEMC22, PDM^{+}24, RTHB25, RMD^{+}25, SKY^{+}24, SJBT20, SRF^{+}24, SOSK22, YRTP20, dS21]. monkfish [CKM^{+}20]. monodon [FBQA20, GBO^{+}20]. monofilament [SME^{+}24]. Monomia [BaLK^{+}21, aLBK^{+}21]. monophasic [CMD^{+}23]. Monopterus [NK23]. monotone [CYBW22]. monotuberculatus [GBC^{+}23b]. montagui [BAM^{+}24, TLV23]. Monte [Alv21, dAdSR^{+}20]. monthly [HCDF24]. months [CWM^{+}23]. moratorium [MB20, PGAG22]. Moreton [LMG^{+}24]. morhua [BWB^{+}23, CWRR24, CSH^{+}21, EM23, HMM^{+}21, HOQ20, JBL^{+}22, MHH^{+}20, NEBP^{+}23, PNAPH24, SJB^{+}20, SBL^{+}23, SF22b, SWIRF21, UPBH^{+}20]. Moridae [KFO20]. morio [SACS23]. Morlet [LTT^{+}23]. Morocco [LMT^{+}22]. Morone [GSH22, WMT^{+}20]. morphological [BPT^{+}20]. Morphology [HML^{+}20, LGD^{+}20]. morphometric [MDS^{+}20]. Morphometrics [TP24]. morphometry [DOB^{+}24, MOCGCC^{+}25, TLV23]. mortalities [BM23a, IAB20, TBH^{+}22]. Mortality [SFJ^{+}23, AHB^{+}22, AUC25, AAT^{+}21, AK23, BQBW20, BCS^{+}22, BKT^{+}20, BMC20, BGLP21, BGCCP22, BKC21, BHG^{+}24a, BC20, BCOBB^{+}23, BBSM24, CWRR24, CC22, CPF20, Cla22, CH22, CW22, CLY^{+}22, CHAY^{+}25, CFP22, DWLT21, DB22, FDS^{+}23, HMM^{+}21, HC22a, HIMP23, HC22b, dlTGPLC23, HW21, HHLL25, HP23, HWMVM23, KC22, KBPS21, KBPS22, KFDE^{+}22, KHPB20, LBP^{+}24, LBD23, LCG22, Lor22, LCHB^{+}24, MHL^{+}23a, MBH^{+}22, PVPN22, PC21, PBRT22, PGRD21, PHH^{+}23, PDF20, PSS^{+}21, PCJH^{+}21, RKD24, RBD^{+}22, RBHM24b, RPM^{+}21, RSPE22, RBG^{+}24, SH22a, SR21, SDV^{+}22, SXM^{+}21, SGH^{+}20, SKJJ25, TdL24, VA20, VP23, WCGB22, WS20, WW21b, WCN^{+}24, WPLF20, vDJB^{+}23, JSKM20]. mortem [RS21]. most [GMT24, GRJW20, SBC^{+}23]. motivations [BR22, vdHR23]. motives [GAB^{+}22b]. mottled [HTK^{+}24]. mountain [YMS21]. Movement [BNTK23, CMRP20, DMM^{+}23, HMS^{+}22, PDC^{+}24, ALRA20, AGB^{+}24, NGDC25, ÖRS^{+}25, OCdMC24, ZSDZ24]. Movements [FFG^{+}20, AWVS21, BMJ^{+}24, BHD^{+}23, CTR^{+}21, HKG^{+}21, SC20]. Moving [BAM^{+}24]. Mozambique [NLS21]. mPER [DNLM23]. MS [AMdC^{+}20, BTB^{+}21, LCW^{+}23a]. MS-based [LCW^{+}23a]. MS/MC [BTB^{+}21]. MSY [ZF21]. MSY-based [ZF21]. Mud [HTSJ23, GGMRC^{+}22, KRH^{+}24]. muelleri [KHGB25]. Mugil [APBN22, ÁHGCVAI22, CCRGC^{+}24]. Mugilidae [CCRGC^{+}24]. Muller [KHGB25]. mullet [ÁHGCVAI22, CCRGC^{+}24, SCC^{+}22]. mullets [RdBAT^{+}23]. mulloway [BM24b, HMS^{+}22, RTB^{+}21]. Mullowney [BWN^{+}21]. multi [BGCCP22, DMZH21, GRHHGM^{+}20, GAW^{+}22, HL20, İTAD24, JC21, LCC25, MDMS21, NCB^{+}23, QMC^{+}22, RGN^{+}20, TMSS^{+}23, WWT^{+}25, YIM^{+}20]. multi-annual [MDMS21]. multi-frequency [RGN^{+}20]. multi-gear [LCC25]. multi-marker [GAW^{+}22]. multi-method [TMSS^{+}23]. multi-model [DMZH21, GRHHGM^{+}20]. multi-proxy [QMC^{+}22]. multi-region [HL20]. multi-sector [BGCCP22]. multi-species [JC21, NCB^{+}23]. multi-stage [İTAD24]. multi-tissue [WWT^{+}25]. multi-variate [YIM^{+}20]. multidecadal [YWC^{+}21]. multidirectional [MPC^{+}20]. multidisciplinary [GAB^{+}22a]. multifaceted [OGFW24]. multifleet [NAS^{+}20]. multigear [FMLSP20]. Multiple [MSV21, BRGB^{+}23, BSR^{+}22, BHD^{+}23, CPM21, LCFJ22, LDM^{+}24, MJD^{+}21, PDD^{+}22, SXK^{+}24, TSC^{+}22, TDI^{+}21, TCL^{+}24]. Multiple-scale [MSV21]. multipliers [SW24]. multiscalar [CRL21]. Multispecies [PDF20, ASJ^{+}22, Det21a, FMLSP20, HCDF24, NAS^{+}20, PBRT22, QSMG^{+}23, SMA^{+}24]. multitask [PPC^{+}21]. multivariate [BCG^{+}25, QSMG^{+}23, SRB^{+}25]. Munididae [GBO^{+}20]. Muskellunge [PW25, BCS^{+}22, WW21b]. Mya [LCM^{+}23]. Myanmar [CTS^{+}23]. Myliobatis [GRHHGM^{+}20]. myriaster [MZZ^{+}21]. myxosporean [GKC^{+}22]. Myxozoa [GKC^{+}22].

N [FRP22]. naevus [NDRR20]. namaycush [DH20]. Nansei [ASO^{+}22]. narrow [BM22]. natal [HRH22, SPW^{+}22]. National [FFG^{+}20, TAA^{+}20, HEGR24, WW21a]. native [BN21, DD25, Lya22, YIM^{+}20, YiTM23]. Natural [HIMP23, HP23, LCG22, PC21, AAT^{+}21, BQBW20, BKT^{+}20, BCOBB^{+}23, CC22, CRL21, Cla22, CH22, dSCCC^{+}22, CW22, CLY^{+}22, CFP22, DWLT21, DB22, FDS^{+}23, HC22a, HPPT24, HC22b, HHLL25, HWMVM23, KHPB20, LR21, Lor22, MHL^{+}23a, NTJN21, PVPN22, PBRT22, PGRD21, PBM^{+}23, PCJH^{+}21, SDV^{+}22, SXM^{+}21, SSJ^{+}21, TdL24, VP23, WPB22]. natural- [BQBW20]. natural-origin [HPPT24]. naturalized [ECY21]. naturally [HGS^{+}23]. navigational [LCW^{+}23a]. NE-Atlantic [OBB^{+}20]. near [BHB24, DOB^{+}24, PJSQ20, SSI^{+}23, TRWH23]. near-infrared [DOB^{+}24]. Nearly [RPH20]. nearshore [DLP^{+}24, GBWM22, GPT^{+}21, GLP^{+}20, JDH22, LCW23b, VKS^{+}25]. need [BMJ^{+}24, MTX^{+}20, WKBMW24]. needed [CC20]. needs [BBG^{+}24]. neglected [Mac22]. Neiden [HKKa^{+}25]. Nelusetta [GFM^{+}23]. Nematistius [JCCAS^{+}21]. nematodes [DCL^{+}20, PNAPH24]. neodymium [GRG24]. Neogobius [BBC^{+}20]. neon [WHCF22]. Neotropical [AGNS^{+}21, KDdOM^{+}22, LFdB^{+}21]. Nephrops [AMSC20, GDVBB^{+}20, IJS^{+}22]. nerka [EMR^{+}22, PPC^{+}23b]. nests [SSJ^{+}21]. net [CSH^{+}21, EM23, FOM21, FMMC20, GP25, GKKL24, HMC^{+}23, HEGR24, JMP^{+}21, KLNB^{+}24, KFI^{+}21, LLK^{+}22, LS24, LVPS20, MSC^{+}24, OCBJG20, PMC^{+}24, SME^{+}24, WEH^{+}25]. net-pen-based [CSH^{+}21]. Netherlands [vdHC20]. NetLights [SBB^{+}24]. nets [CHM24, EMJ^{+}22, FSP22, GCK^{+}21, GMK23, GP25, Har21, PFGQ20, RMD^{+}25]. netters [PVA^{+}24]. netting [BWN^{+}21, MBD^{+}21, NBD^{+}20]. network [CGT^{+}23, FFG^{+}20, İTAD24, MZ24, WZL^{+}22, MAH^{+}22]. network-based [MZ24]. networks [BCG^{+}25, JvPOG25, LZY24b, LZY24a, QSMG^{+}23]. neural [LZY24b, LZY24a, QSMG^{+}23, WZL^{+}22]. neutral [SPC^{+}23]. Newfoundland [BMSM22, CWRR24, MBZSM20, MB20, MBP20, RMRG22]. newly [DOB^{+}24]. next [BBG^{+}24, HMP^{+}22, KSO^{+}21, PDE^{+}20]. next-generation [PDE^{+}20]. Nguyen [MBD^{+}21]. Niche [ALW^{+}21, KDdOM^{+}22, LCFJ22, SWIRF21, YLX^{+}24]. niches [WWT^{+}25]. Nigeria [IOO^{+}24]. nightly [LZC^{+}21]. Nighttime [LSZ^{+}23]. nigricans [WWO20]. nigritus [SR21]. Nile [PvZ22]. Ningaloo [AGL^{+}24]. niphonius [FHE^{+}24, PYX^{+}20]. NIR [DOB^{+}24]. NIRS [PJSQ20]. nitrogen [HTK^{+}24]. No [AVACA^{+}23, BKM23a, CWM^{+}23, KMC^{+}23, NDRR20, SCGM^{+}21]. no-take [KMC^{+}23, SCGM^{+}21]. non [BSAP22, BC20, CG21, FOM21, HK22, LOFS22, LSM^{+}23, Lya22, MOB^{+}23, NLL^{+}25, PJSQ20, STH25, SIM^{+}24a, SIM^{+}24b, TBF^{+}21, TTC^{+}25]. non-commercial [FOM21, STH25]. non-compliance [BSAP22]. non-destructively [PJSQ20]. non-linear [HK22]. non-local [LOFS22]. non-mesh [MOB^{+}23]. non-migratory [TTC^{+}25]. non-native [Lya22]. non-occluded [SIM^{+}24a, SIM^{+}24b]. non-resident [LOFS22]. non-stationary [CG21]. non-target [BC20, LSM^{+}23, NLL^{+}25]. non-traditional [TBF^{+}21]. nonparametric [CYBW22]. nonparametric-monotone [CYBW22]. Nordmann [RKN23]. Nordmøre [Bro25]. Nordmøre-grids [Bro25]. Norte [CRSC22]. North [SBC^{+}22, Sch23, CAGLT23, FMMA20, GGMRC^{+}22, MMG^{+}24, MPM^{+}23, SPM^{+}24, TMN^{+}21, TMDA22, WCLN20, ABF^{+}21, ABBO20, BFA^{+}21, BLHS20, CBJ24, GLA^{+}20, GTS^{+}21, HJMS20, HLI^{+}20, KTFY22, KNO^{+}21, KPUB22, LCD^{+}23, MMM^{+}20, OBB^{+}20, PRK23, PDF20, ŞGK^{+}20, SMA^{+}24, TBH^{+}22, vDJB^{+}23]. North-east [BLHS20]. north-eastern [GGMRC^{+}22, TMN^{+}21, TMDA22]. north-western [CAGLT23, MPM^{+}23, SPM^{+}24, WCLN20, HLI^{+}20]. northcentral [PBPM^{+}23]. northeast [CRF^{+}24, MHL^{+}23b, SWLH20, GM21, GMT24, GKC^{+}22, HUK^{+}23, HP23, KHC^{+}20, íKMíH^{+}22, MPC^{+}20, NDRR20, PMS^{+}23, PPM^{+}23, SKD^{+}20, TRN^{+}23]. Northeastern [MFO21, BHH21, CdSLP21, KC22, SFCG^{+}21, WPB22]. Northern [DSNK^{+}22, FNH^{+}21, LZY24b, AHL20, ARD^{+}23, BMM^{+}24, BHB24, BLE^{+}22, CMD^{+}23, DCK^{+}22, EBN^{+}23, FG21, FG23, FRP22, FHE^{+}24, GPWP20, HPS^{+}24, HGS^{+}23, JAN23, LOFS22, LPG^{+}24, LRGB25, MHB^{+}23, NB24, NDRR20, PIP^{+}22, PAY23, PHH^{+}23, PZG23, RDR^{+}23, RSPE22, SCGM^{+}21, SBL^{+}23, SCS25, Ten22, TSI^{+}21, WJN^{+}25, dAdSR^{+}20, BLC^{+}22, EVS^{+}23, GPASM22, HPD^{+}22, KCB^{+}24, MSV21, RBD^{+}22, SGZD21, VMI21]. Northwest [LvCdGSL23, WBD^{+}21, CKM^{+}20, HBC^{+}22, HM25, LZL^{+}24, RCH^{+}21, SdFZFJ21, TTK^{+}25, GC21, HGC24, PBDM21, SCN^{+}24]. northwestern [DSP^{+}23, GRHHGM^{+}20, SLW^{+}20, HCK^{+}21, HCDB22, SCD^{+}22]. norvegicus [AMSC20, GDVBB^{+}20]. Norway [AMSC20, KK21, SGZD21, SGZH^{+}22, SV23]. Norwegian [AAH^{+}23, Apo25, BHH21, HA23, MB23, NBF20, OWF^{+}23, SV23]. notching [HGHH25]. notialis [AGNR^{+}21]. notothen [CZ25]. Nototheniidae [BRR^{+}21]. Nototodarus [LPP^{+}20]. Novel [PHH^{+}23, SSY20, SLA24, BDA^{+}24, CBTH20, EDA^{+}22, FBW^{+}21, GJGW23, HGHH25, LLK^{+}22, LBD23, MOB^{+}23, MBB^{+}23, OWF^{+}23, PRA^{+}23, RB22, SFMA23, SSP24, SCC^{+}22, TSC^{+}22, TCL^{+}24, dLH23]. November [Ano20v, Ano21v, Ano22w, Ano23v, Ano24v]. NSW [ML24]. nucleotide [ÁHGCVAI22, LLC^{+}20, RKN23]. Nui [CGDTSA^{+}25]. number [MSW21, MSS^{+}21, NSRM22]. numbers [MMP^{+}24, PÁEMC22]. Nunavut [LvCdGSL23]. nurseries [RTB^{+}21]. nursery [AAFLL^{+}25, ESB^{+}24, GSS^{+}23, LBW21, MJD^{+}21, SSS^{+}23]. nutritional [ABT^{+}24, CNE^{+}22, Lun25, NK23]. NW [OOAF^{+}21, BNL^{+}23, GDVBB^{+}20, GMRRG20, MABR^{+}20, SCGM^{+}21]. nybelini [TEO25]. nylon [SCCAM21].

O [PD25, RWFT25]. Oahe [OGFW24, FFG21]. O’ahu [WSL21]. obesus [LK25, SXMV^{+}21, WAA^{+}22]. Object [HL21, XDX^{+}23]. objectives [NAS^{+}20]. ObsCovgTools [CC20]. obscurus [SGH^{+}20]. observation [CSH^{+}21, FCSA21, LSZ^{+}23, SSM^{+}23]. Observations [BW23, BMM^{+}21, GKC^{+}22, PGD^{+}25]. observed [SAHW22, WSB22]. observer [CC20, DC24, MDJP24]. Observing [SGH^{+}20]. obtain [TRWH23]. obtained [SLW^{+}20, ZHC^{+}25]. occluded [SIM^{+}24a, SIM^{+}24b]. occurrence [LPG^{+}24]. Ocean [Ash20a, CBD^{+}22, MTC^{+}22, PDF20, WJN^{+}25, ABT^{+}24, AAFLL^{+}25, Ash20b, AAR^{+}21, BNTK23, BFA^{+}21, CHT20, CVM^{+}20, dSCCC^{+}22, CW22, CNDDAPMR21, DDCNMR21, FCKG^{+}22, GGG^{+}22, GPW^{+}20, GM21, GMT24, GFC^{+}22, HL20, HCK^{+}21, HCDB22, KSI^{+}22, íKMíH^{+}22, LCDM^{+}24, LZL^{+}24, LK25, MCC^{+}23, MBE^{+}20, MAA^{+}20, MBH^{+}22, NFdSJO25, PVPN22, PGAG22, PRK23, QMGRIU22, SASB24, SSRC24, SXMV^{+}21, SFCG^{+}21, SF22a, SSP^{+}22, SAA^{+}22, SDC^{+}22, SML^{+}24, TRN^{+}23, TFC^{+}20, VP23, WAA^{+}22, WWO20, XMLCMV24, dAdCdO^{+}23]. oceanic [CSRL20, JWL^{+}24, NSQV22, Mel23]. Oceanographic [KHC^{+}20, KNP^{+}20]. oceanography [PNGGO^{+}22]. Oceans [GAW^{+}22, IKBL23, WWF^{+}20]. ocellatus [PHH^{+}23]. October [WGC^{+}21, Ano20w, Ano21w, Ano22x, Ano23w, Ano24w]. octopods [FHSC21]. Octopus [SRT^{+}20, MSD21, NY23, PRF^{+}21, SKS^{+}23, WKBMW24, ALRB^{+}20, CMBL21, GGMMMV^{+}20, HMY25, SKS^{+}23, SRT^{+}20]. Odontesthes [DSB^{+}21]. odontocetes [KSS^{+}22]. off [AYTM21, ASO^{+}22, BMSM22, CQA^{+}24, FHE^{+}24, KBH22, LCD^{+}23, MSC^{+}24, MWTH23, MVLC^{+}20, QMGRIU22, RCH^{+}21, SCCAM21, SLW^{+}20, TMN^{+}21, WQGS25, YTH22]. offer [PPC23a]. officinalis [LCD^{+}23, SPC^{+}25]. offs [BGCCP22, HLMV24, KDBOC25, NAS^{+}20, PBDM23, SHC21]. Offshore [WHR^{+}24, ASWS^{+}21, BTML20, DEM^{+}23, EWPB22, KWW^{+}21, VKS^{+}25]. ogives [MBOCdAM24]. Ohio [PW25]. ok [ABT^{+}24]. Oklahoma [JCL^{+}21]. old [EPHDB24]. olivaceus [TLCD21, TLC^{+}22, UFYT23]. Olo [AAVÁM23]. Ommastrephes [WHCF22]. onboard [FOS^{+}21]. Oncorhynchus [EMR^{+}22, HKKa^{+}25, IRJ^{+}22, LCHB^{+}24, PPC^{+}23b, YSB^{+}21]. one [CWM^{+}23]. one-trial [CWM^{+}23]. ongrowing [GAB^{+}24]. online [GVK^{+}23, HJA^{+}21]. only [DHX^{+}23, FJA^{+}20, HSSD^{+}21]. onset [Mau22]. Ontario [HJA^{+}21, LCFJ22]. Ontogenetic [CLB^{+}21, FHSC21, LXC^{+}20, SSD^{+}20a, AAR^{+}21, CZ25, WHCF22, ZSDZ24]. ontogeny [LPS^{+}25, YLX^{+}24]. oocytes [GKM^{+}23]. open [BBČ^{+}21, LZC^{+}21, SLF23]. opening [Bro25, YHC^{+}24]. openings [AIJ^{+}23]. operated [TTK^{+}25, TJS23]. operating [LAG^{+}21, MWJ^{+}24]. operation [SZS^{+}24, dAdSR^{+}20]. operational [PÁEMC22]. operations [MB23, QSMG^{+}23]. opercularis [OOAF^{+}21]. opilio [AIJ^{+}23, BMSM22, BWN^{+}21, HLG^{+}21, MBD^{+}21, Mur20, Mur21, NBD^{+}20]. Oplegnathus [CW22, FCMP23]. opportunistic [Fis25]. opportunities [PPC23a]. optic [JRW^{+}21, WST^{+}23]. optical [MOB^{+}22, SEM^{+}23a]. Optimal [ZHC^{+}25, Che21, WZX^{+}20, ZZH^{+}24]. Optimising [Ken21, PBB^{+}22]. optimize [AIJ^{+}23, vBMP^{+}23]. Optimizing [HIKM21, JM23, LDC24, MCS^{+}22]. options [ASJ^{+}22, GBB20, PFGQ20]. orange [BCOBB^{+}23]. Ordering [WSL^{+}24]. Ordway [GZLCRG25]. Oregon [HCWH23]. organic [RWFT25]. organizations [HFMH20]. organosomatic [SWIRF21]. orientalis [ASO^{+}22, AIW^{+}23, SAW^{+}20]. orientation [BS20a, CHB24, SPE^{+}23]. orientations [BKM23b]. origin [AAFLL^{+}25, BTB^{+}21, BQBW20, BKC21, HPPT24, KWC^{+}20, RSD23, SPC^{+}23]. origins [CRS23, LZL^{+}24, LBW21, MJD^{+}21, SPW^{+}22]. ornamental [FDdCS^{+}20, MCH21, PDJ24, dAdSR^{+}20]. ornatus [YFJ^{+}25]. Oryzias [XSS^{+}23]. Oscillation [CBJ24]. oscillations [AUHK22]. OSMOSE [XCB^{+}21]. OSMOSE-JZB [XCB^{+}21]. ossicles [HMR^{+}24b]. Osteichthyes [SBT^{+}20]. Otaria [IAB20]. Other [Mun24, dFBPL^{+}20, BM24b, SKBA23]. othonopterus [AVACA^{+}23]. Otolith [AAFLL^{+}25, BRR^{+}21, CGDTSA^{+}25, CdSLP21, DOB^{+}24, DHC^{+}20, MPP22, SNJ^{+}24, SDC^{+}22, SSD^{+}20b, TLC^{+}22, WJN^{+}25, XqRJ^{+}23, ZSDZ24, AUC25, ACS23, AAR^{+}21, BMA^{+}20, Cam23, CPR^{+}24, dSCCC^{+}22, CMTP^{+}21, CRS23, FG21, FGL22, FJHT^{+}22, GFM^{+}23, HLC^{+}25, HQWD20, HRH22, İTAD24, KD22, LGD^{+}20, LMT^{+}22, LHH^{+}25, LBW21, MJD^{+}21, MCGL^{+}25, MNPMM^{+}22, MMM^{+}20, MZZ^{+}21, NSM^{+}21, NVSG24, PYX^{+}20, PPC^{+}21, PD25, QMC^{+}22, RdBAT^{+}23, RTB^{+}21, SIMT24, SBC^{+}22, SSSF25, SCN^{+}24, SSP^{+}22, SSS^{+}23, SFMA23, SXK^{+}24, SSD^{+}20a, SF22b, SLW^{+}20, TTYT24, TSG25, UFYT23, VK21a, VK21b, VBL^{+}22, WQGS25]. otolith-based [GFM^{+}23, SXK^{+}24]. otoliths [ATA^{+}24, BHB24, CHL^{+}20, CSRL20, DOB^{+}24, FJJT^{+}21, GQP^{+}25, JCCAS^{+}21, PJSQ20, PRRR23, RWFT25, SdOR^{+}23, SPW^{+}22, SNHM23, TLCD21, ZOS^{+}23, dAdCdO^{+}23]. otter [CAH^{+}20, OCBJG20, OFF^{+}20, WGNM24]. oualaniensis [LZCC24]. our [SSP24]. outcome [HEG^{+}23a, HGS^{+}23]. outcomes [BKM^{+}23c, CBN^{+}21, FDS^{+}23]. outer [FGSD25]. outliers [XMCC20]. output [Apo25, WW21a, WHM23]. over-exploited [WS20]. overcapitalization [DRSPTA20]. Overexploitation [BAJB^{+}24, DWS^{+}23, Fre22, SECB21, WXJ^{+}24]. Overfishing [SHV20, PdAMdM^{+}23]. overlap [ALW^{+}21, KA24, NOL23, SBL^{+}23]. Overlapping [WBA23]. overlook [BCM^{+}21]. overlooked [OYOO21]. overthinking [BHG^{+}24b]. overview [BF25, LVPS20, SRF^{+}24, WPLF20]. overwintering [Lya20b]. ovigerous [JGG^{+}24]. ovoviviparous [HP23]. ovulatory [CKDP^{+}20]. oxygen [BBHF25, SIMT24, SSSF25]. oyster [DWLT21, MDW^{+}21, MWJ^{+}24, OYOO21]. Ozark [WST^{+}23].

pāua [RGP^{+}23]. pace [SCW21]. Pacific [Ash20a, CNDDAPMR21, DDCNMR21, GM21, GMT24, HCK^{+}21, HCDB22, LCDM^{+}24, OS21a, PRK23, QMGRIU22, SML^{+}24, TRN^{+}23, AVMBEB22, ABBO20, Ash20b, ASO^{+}22, AIW^{+}23, ÁHGCVAI22, CGDTSA^{+}25, CPPK23, CMIMS20, DLP^{+}24, FCKG^{+}22, FHHH20, FAK24, GZLCRG25, GAW^{+}22, GHAZ21, HGC24, HCK^{+}21, HCDB22, HLZ^{+}20, Hut22, JLYR24, KTFY22, KNO^{+}21, KHC^{+}20, KFO20, KPUB22, LZL^{+}24, LBW21, LWH^{+}21, MCS^{+}24, MBE^{+}20, MAA^{+}20, MBH^{+}22, NY23, OS21b, PVPN22, PRK23, PDF20, PDA^{+}24, RNP^{+}24, SXMV^{+}21, SFC21, SFCG^{+}21, SF22a, SWLH20, SAW^{+}20, SFYM24, SHC21, TMN^{+}21, TRN^{+}23, TSI^{+}21, VP23, WWF^{+}20, WYM^{+}25, WWO20, XMLCMV24, YFJ^{+}25, YSB^{+}21, YWC^{+}21]. package [PDE^{+}20, TDJ^{+}21]. packages [DDDP21]. packaging [LTR20b]. Paddlefish [WLG^{+}23, LJB^{+}24]. Pagellus [CRF^{+}24]. pagurus [CBJ24, MDS^{+}20, MDJP24]. painted [CZ25]. pair [PCC^{+}23]. paired [CYBW22, CSSB22, MFR22a, MRP^{+}23, NSRM22]. paired-age [NSRM22]. paired-tow [CYBW22]. PAL [CCKL^{+}20]. Palau [FFG^{+}20, MKFF^{+}21]. pallasii [WYM^{+}25]. pan [OAB^{+}23]. pan-Baltic [OAB^{+}23]. Panay [NB24]. Pandalopsis [ZF21]. Pandalus [BAM^{+}24, IJS^{+}22]. pandemic [KKLM24, MLCMdS23]. pandemic-related [MLCMdS23]. panel [BS20a, BKM23b, CAH^{+}20, IB20, VS23, PRA^{+}23]. pangasiid [VBL^{+}22]. panmixia [HIF^{+}24]. Panopea [dlTGPLC23]. pantropical [BW20]. Panulirus [AIM^{+}23, BBR^{+}22, BBSM24, HMR^{+}24b, HMR24a, RBHM24a, RBHM24b, YFJ^{+}25]. Pará [FDdCS^{+}20]. paradigm [GSH22, Lor22]. paralarvae [GMRRG20]. Paralichthys [BAF23, MCS^{+}22, TLCD21, TLC^{+}22, UFYT23, WMSW22]. Paralithodes [AAH^{+}23, LYH^{+}21, MSJ21]. Parameter [VBB20, CRCAF^{+}22, FS20]. Parameterizations [BP20]. parameterizing [CHGC25]. parameters [BBHF25, CH21b, DBV22, DBV23, GBO^{+}20, dlTGPLC23, HK22, MCHA21, RCVGMI22, SKW^{+}21, SRT^{+}20, WZS^{+}21, ZF21]. parametric [MLCMdS23]. Parapenaeopsis [CKD^{+}21]. Parapenaeus [PDG^{+}22]. parasite [CLB^{+}21, GKC^{+}22, MMG^{+}24]. Parasites [MSLMOC^{+}24, BRR^{+}21, FRP22]. Parastichopus [RAG22]. Parentage [Bea21, BJS^{+}22b, BJS^{+}22a]. Parentage-based [Bea21, BJS^{+}22b, BJS^{+}22a]. Park [Mac25, TAA^{+}20]. Parker [MSS^{+}21]. part [ATAS20, DKBF23, NDRR20, CWRR24, WCN^{+}24]. Participation [vdHC20, LWH^{+}23b, NC20]. Partitioning [DHCS23, LMG^{+}24, YLX^{+}24]. partners [PNRS23]. partnerships [vPDD^{+}25]. parts [FCKG^{+}22]. Pass [MMML24, DJFU20]. Passage [LvCdGSL23, BH23, OSEF22]. passive [BHNP22, MLS^{+}21]. passively [TJS23]. Past [Pun24, WH23]. Patagonia [AdlBCN25, HRC23, MSV21]. Patagonian [MFM^{+}20, AHX^{+}24, BRR^{+}21, CTCB22, MMG^{+}24, PCF23, TNDM23]. patagonica [PCF23]. Patagonotothen [BRR^{+}21]. patchy [LPS^{+}25]. path [SBJ^{+}20]. pathways [PDC^{+}24]. pattern [LZW^{+}21, MPC^{+}20, Žák21]. patterning [DAL20]. Patterns [DWLT21, ASGG21, ABF^{+}21, BNTK23, BHST^{+}21, BBC^{+}20, CGM^{+}22, CAGLT23, DSB^{+}21, DMM^{+}23, DBGV^{+}22, EBGE21, FJHT^{+}22, FSS^{+}24, GMPD23, dlTGPLC23, HMS^{+}22, KKLM24, LCDM^{+}24, LZL^{+}24, LNvD^{+}23, NNS^{+}22, PAA^{+}24, PFFdC22, SSSF25, SGW^{+}22, YWC^{+}21]. paulensis [PFFdC22]. pay [BKM^{+}23c, TCVG20]. PCR [HSL^{+}22]. PDR [RBH^{+}24]. Pearls [KLNB^{+}24]. pearlside [KHGB25]. Pecten [EDA^{+}22]. pectoralis [JCCAS^{+}21]. pelagic [AMHH21, AOA^{+}22, BAYR^{+}24, CHT20, CNE^{+}22, CAAFH21, DCR^{+}20, GPASM22, GdSPL21, IRJ^{+}22, MCC^{+}22, MRE^{+}24, MSLMOC^{+}24, NVSG24, NFdSJO25, OWF^{+}23, PMS^{+}23, PPM^{+}23, RLO^{+}21, SHH^{+}21, SSRC24, SAA^{+}22, SGH^{+}20, SAA23, TPD20, WLG^{+}23, YSB^{+}21, ABBO20]. pelagicus [SSG^{+}22]. pelamis [ATA^{+}24, AUM21, Ash20a, Ash20b, dSCCC^{+}22, ZZC^{+}21]. pen [CSH^{+}21]. penaeid [BS20a, BM23b, BM24a, BM24b, CTM^{+}20, SJBT20]. penaeid-trawl [BM23b, CTM^{+}20]. Penaeidae [AGNR^{+}21, SS24]. Penaeus [AGNR^{+}21, RKD24, SSY20]. Peninsula [GGMRC^{+}22]. Peninsular [NTJN21]. Pentaceros [CSRL20]. people [AOA^{+}22]. peppery [MGB24]. Perca [EHE^{+}23, SSD^{+}20a, VTSI^{+}24]. perceptions [BR22, vdHR23]. Perch [GWGM24, EHE^{+}23, PvZ22, SPC^{+}23, SSD^{+}20a, VTSI^{+}24, VCG^{+}23]. Perciformes [AVACA^{+}23]. Pérez [AGNR^{+}21]. perfect [ODM20]. Performance [CZJ^{+}24, DHX^{+}23, FJA^{+}20, LZC^{+}21, MMP^{+}24, PBDM21, SFC21, TLAM25, ASWS^{+}21, ANB^{+}24, BM23b, BKM23a, BM24b, BCF^{+}23, CFB^{+}23, ECY21, Fu22, JC21, KC22, LVP22, LBD24, Lun25, MCC^{+}22, MHL^{+}23a, PTD^{+}21, SK21a, TTL^{+}20, VBB20, WGC^{+}19, WGC^{+}21]. performances [LZX^{+}20]. period [KTFY22]. Periodic [SC20]. permeability [MMQ21]. permits [RFMS^{+}21]. Persian [BTB^{+}21]. persistence [BCSM20]. Persistent [FWL20, PIP^{+}22, SBBH24]. personal [Har21]. Perspective [Tho19, Tho21, AOA^{+}22, BSA^{+}23, Bro24, LMJ^{+}23, MCS^{+}25, MAH^{+}22, PVC^{+}22, PDE^{+}20, Pun23]. Perspectives [BMM^{+}21, LYLC21, MBZSM20, ZDF^{+}22, ZCM^{+}23, ZCFG23]. Peru [MRG^{+}23, AVMBEB22]. Peruvian [CCR24, PCCMOA^{+}24]. Pervasive [CKM^{+}20]. Peterson [Fu22]. petroleum [EBGE21]. phase [HLCC22]. phasic [WKSF20]. phenological [HIF^{+}24]. Phenomenological [JDP22]. Phenotypic [NFC20, SDC^{+}22]. Philippines [CRSC22, LNP25, MTE^{+}20, NB24]. philopatry [CCRGC^{+}24, HRH22]. Phocoena [CCKL^{+}20]. phospholipid [Lun25]. photo [CTR^{+}21, JSKM20]. photo-identification [CTR^{+}21]. photo-releases [JSKM20]. Phragmites [NHE^{+}23]. phylogeographic [BQGV^{+}24, ZJY^{+}24]. Phylogeography [ŞGK^{+}20]. physical [CEAL21, GMK23, HRC23]. Physiological [CH21a, MSJ21, SBT^{+}20, AGNS^{+}21, SCD^{+}22]. physiology [BKR^{+}22, KFDE^{+}22]. pica [BQGV^{+}24]. pick [FCKG^{+}22]. picturatus [MPC^{+}20]. pike [AHL20, ARD^{+}23, BLE^{+}22, DAR^{+}23, DSNK^{+}22, EBN^{+}23, EHE^{+}23, FG21, FGL22, FG23, FDS^{+}23, FNH^{+}21, FTH^{+}23, HGS^{+}23, KMA22, KSJM^{+}20, LPG^{+}24, LRGB25, MHB^{+}23, NHE^{+}23, NRH^{+}23, OAB^{+}23, PZG23, RDR^{+}23, SKBA23, SCS25, WBBG^{+}23, BLC^{+}22]. pikeperch [LM22, SSD^{+}20b, SSJ^{+}21, TSS^{+}23]. pikes [KSJM^{+}20]. pilchardus [LMT^{+}22, NSM^{+}21]. pilot [HHD^{+}20, KBB^{+}21]. pinger [PCC^{+}23]. Pingers [MB23, BSKL^{+}22]. pink [AGNR^{+}21, BÁP^{+}23, HKKa^{+}25, PFFdC22, VAVQGD^{+}20]. pink-shrimps [PFFdC22]. pinniped [PSS^{+}21]. pioneering [MAH^{+}22]. Pisces [ÁHGCVAI22, CCRGC^{+}24]. piscivore [HGS^{+}23]. piscivorous [BN21]. PIT [BS20b, PRN^{+}24, RBH^{+}24, ZDZR^{+}22]. pitfall [PBG24]. placement [TGG^{+}24]. Placopecten [HC22b]. plaice [EMJ^{+}22, EFM25, OFF^{+}20, UAB^{+}21, UAG^{+}23]. plan [AAVÁM23]. plana [MGB24]. planktivorous [SHV20]. plankton [VTS^{+}22]. planktonic [LPS^{+}25]. planning [AVCA22, NFdSJO25, SLA24]. plans [MPH21]. plants [RBH^{+}24]. Plastic [KLNB^{+}24]. plasticity [NFC20]. Plata [ACP^{+}23]. plate [TTK^{+}25]. platessa [EMJ^{+}22, EFM25, HML^{+}20, UAB^{+}21, UAG^{+}23]. platform [GAB^{+}20, WASS20, WSUN^{+}23]. platforms [AWC^{+}23, BBG^{+}24, EBGE21]. Platichthys [FSS^{+}24, HML^{+}20, KZT^{+}23]. player [SJTGAS23]. Playing [PBRT22, YTHM20]. Plectropomus [SK22]. pleii [MdCG20]. plenty [AUHK22]. Pleuroncodes [FBQA20, GBO^{+}20]. Pleuronectes [EMJ^{+}22, EFM25, HML^{+}20, UAB^{+}21, UAG^{+}23]. plumbeus [PCBL23]. plumierii [NVRG^{+}21]. Poey [PJSQ20]. point [DTSR22, KSO^{+}21, ÖÜÖG20, VP22]. pointhead [YMYH20]. points [AT20, DQK^{+}23, LWH^{+}23a, LZY^{+}24c, SMA^{+}24, ZF21]. Poland [KC22]. pole [FMMA20, KNP^{+}20]. pole-and-line [FMMA20, KNP^{+}20]. poleward [Mur20]. poli [AHEV24]. Policy [RMNB^{+}21, BDR^{+}20, FGPVPPGG22, SPC22, YH21a, YH21b]. Polish [PZG23, WBBG^{+}23]. pollock [DB22, EYAO20, YMYH20]. polyactis [SXK^{+}24, YLP^{+}23, ZLXL20]. polychaetes [MOCGCC^{+}25]. Polydactylus [BCM^{+}21]. polymorphism [RKN23]. polymorphisms [ÁHGCVAI22, LLC^{+}20]. polynemid [BCM^{+}21]. Polynesia [BTR^{+}24]. Polynesian [AVCA22]. polynya [ZSDZ24]. pompanos [RFMS^{+}21]. ponds [GWGM24]. pontoon [LLH^{+}25, RSPE22]. pontoon-trap [RSPE22]. pooled [GWMC21]. Poor [RWB^{+}23, CPM21, CCR24, DWS^{+}23, FDB^{+}20, GSH22, PMS^{+}20, SSM^{+}23, WWF^{+}20]. poorly [GFDN^{+}22]. pop [NGDC25, NEBP^{+}23, OOM^{+}23]. pop-up [NGDC25, NEBP^{+}23, OOM^{+}23]. Population [APBN22, AAR^{+}21, BJK24, CMTP^{+}21, FTB^{+}21, GGMMMV^{+}20, HSM^{+}25, MCS^{+}25, MGC^{+}22, MPV^{+}24, MMM^{+}20, NSM^{+}21, PYX^{+}20, PCBL23, RKN23, SBC^{+}22, SXK^{+}24, TP24, TTC^{+}25, ZLXL20, AHX^{+}24, AUHK22, ASDW24, BQBW20, BÖN20, CPM21, DBS^{+}21, Det23, ECY21, FBPCC^{+}21, HCCC21, HSM^{+}24, HLI^{+}20, JSG21, JK20, KHMC23, íKMíH^{+}22, KHPB20, KZT^{+}23, LMT^{+}22, LKSiI22, LLC^{+}20, Lya20b, MJD^{+}21, MDW^{+}21, MWJ^{+}24, MSV21, MHD^{+}21, MKS^{+}22, MSS^{+}21, MOM^{+}25, MSLMOC^{+}24, MPP20, OOAF^{+}21, ÖA21, RKD24, RBD^{+}22, RDR^{+}23, ŞGK^{+}20, SHV20, SMK^{+}24, SNJ^{+}24, SL24, Spa24, SGKAR24, Szu22, TBF^{+}21, TMH23, WBBG^{+}23, WS20, WCN^{+}24, WPB^{+}20, WMT^{+}20, WMSW22, YGMJ20, ZF21, vZvdHCA25]. populations [ÁHGCVAI22, BGK22, BLE^{+}22, BF25, CPPK23, DD25, dlTGPLC23, LBLF20, LvCdGSL23, LLC^{+}20, LZW^{+}21, LCG22, Lor22, MKFF^{+}21, NTJN21, NZP^{+}21, OAB^{+}23, PDA^{+}24, PCGG20, RTB^{+}21, SS24, VKS^{+}25, WBBG^{+}23, ZLXL20, ZOZW22, vDJB^{+}23]. porbeagle [BMJ^{+}24]. porpoise [CCKL^{+}20, MB23, CCKL^{+}20]. porpoises [BSKL^{+}22]. port [LCDM^{+}24]. port-sampling [LCDM^{+}24]. portfolio [ASE21, LdV22, WXJ^{+}24]. ports [CSTR^{+}21]. Portugal [AAVÁM23, MGB24, RMD^{+}25]. Portuguese [ACL^{+}20, PMS^{+}23, PPM^{+}23]. portunid [BBJ22, HHJ^{+}22, SJH^{+}23]. Portunus [BJK24, CBTH20, JYH21, MHD^{+}21, SSG^{+}22, YHC^{+}24]. posed [dAdSR^{+}20]. position [LSM^{+}23]. positional [KA24]. positioning [LS24].

Possible [Har21, GSH22, GBB20, IOO^{+}24, KZT^{+}23, MCHA21, RWFT25]. Post [BLC^{+}22, KFDE^{+}22, MRS^{+}25, SFCG^{+}21, SAW^{+}20, SFJ^{+}23, SACS23, TPW23, AWVS21, AK23, CKDP^{+}20, CHM24, CHAY^{+}25, GC25, HUK^{+}23, LCB^{+}21b, LBP^{+}24, LCHB^{+}24, MSW21, MBH^{+}22, PSS^{+}21, RBG^{+}24, SGH^{+}20, TBÓ^{+}22]. post-capture [CHM24]. post-depositional [TBÓ^{+}22]. Post-harvest [MRS^{+}25, LBP^{+}24]. post-ovulatory [CKDP^{+}20]. Post-release [BLC^{+}22, KFDE^{+}22, SFCG^{+}21, SAW^{+}20, SFJ^{+}23, SACS23, AWVS21, AK23, CHAY^{+}25, LCB^{+}21b, LCHB^{+}24, MBH^{+}22, PSS^{+}21, RBG^{+}24, SGH^{+}20]. post-smolts [HUK^{+}23]. post-stratification [GC25]. postlarval [GGL^{+}24, HMR24a]. pot [AIJ^{+}23, AAH^{+}23, CGC24, CSH^{+}21, HGHH25, MTS^{+}21, MMML24, MHL^{+}23b, YHC^{+}24, ZZH^{+}24]. pot-entrance [CSH^{+}21]. Potential [CBHS24, DLKH22, LLS23, LWH^{+}23b, PVA^{+}24, RTB^{+}21, AIJ^{+}23, BGK22, BBC^{+}25, BBR^{+}22, CSDH^{+}23, CBJ24, CRS23, DKD^{+}21, EM23, ESB^{+}24, EPHDB24, FNH^{+}21, FBPCC^{+}21, KSI20a, LZL^{+}24, LRW^{+}24, LPRB^{+}21, LBW21, MFM^{+}20, MPEBdR23, NK23, OGFW24, PSSFS24, SSFL24, SECB21, SHB^{+}23, Spa24, SSG^{+}22, TSPK24]. potentially [MKS^{+}21b]. pots [BWN^{+}21, MBD^{+}21, NBD^{+}20, SV23, ZZH^{+}24]. potting [EDA^{+}22]. poutassou [MABR^{+}20]. power [CAAFH21, DGMG^{+}22, DSP^{+}23]. powerful [APB^{+}20]. Poyang [ZJJ^{+}25]. prabahari [KSI^{+}22]. practicable [Cop24]. practical [CH22, DH20]. practice [Cla22, PJMP22]. Practices [LMP24, BR22, CGBJ23, Cop24, GBC23a, HEG^{+}23b, HCDB^{+}24, KDBOC25, KBK^{+}24, LCB^{+}21b, LG21, MPSM25, Mon24, PM23, Pun23]. Pragmatic [Bro24, GBC23a]. Prawn [HPD^{+}22, MPSH25, GGTÁVT^{+}20, HPS^{+}24, LZY24b, PHP^{+}20, PNRS23, SSY20]. prawns [HPD^{+}22]. Pre [LTR20a, RS21]. Pre-cooling [LTR20a]. precautionary [SMA^{+}24, WH23]. precautions [KDBOC25]. Precision [FSP22, ASDW24, Det23, Fis25, SFMA23, TLAM25, MBB^{+}23]. Predation [DNLM23, SMC^{+}24, AHB^{+}22, DB22, JK20, SCCAM21, WEH^{+}25]. Predator [BLK23, LRGB25, AJB20, DSS^{+}23, GPASM22, MHB^{+}23, RMRG22]. predator-prey [GPASM22]. predators [PRCF22]. predict [LYH^{+}21, MPM22, PPC^{+}23b]. predictability [SJ24]. Predictable [RFMS^{+}21]. predicted [KWW^{+}21, MBOCdAM24]. Predicting [LBP^{+}24, LCC25, LRW^{+}24, LADA^{+}22, NY23, RL24, GSH22]. prediction [BHB24, İTAD24, LCL25, SRT^{+}20, SH22b, SML^{+}24]. predictions [Aks24, LCW^{+}23a]. predictive [LZX^{+}20]. predictor [RBHM24b, WCGB22]. predictors [KC22, MCHA21, PSSFS24]. predicts [KPK^{+}23, PJSQ20]. Preface [HMP^{+}22]. Prefecture [KM23]. preference [BKM^{+}23c, JCL^{+}21]. preferences [DDCNMR21, LVP22, NLS21, RSD23, SPE^{+}23, vdHBBR20]. Preferential [AVB^{+}23]. Preferred [vBMP^{+}23]. Preliminary [HJA^{+}21, PDC^{+}23, SMKJ21]. premiums [HBW21]. preparation [OUB^{+}22]. Preparing [SBC^{+}23]. presence [AMSC20, CTR^{+}21, DCR^{+}20, MLCMdS23, NGDC25, PRCF22, PNAPH24]. present [Pun24]. preserve [SDdMG^{+}20]. prespawn [BKC21]. prespawning [BQBW20]. pressing [SBC^{+}23]. pressure [LRGB25, MOCGCC^{+}25, NNS^{+}22, RFMS^{+}21, WZS^{+}21, WHM23]. Prevalence [BCSM20, GKC21, SCHSC21]. prevent [ODM20]. Prey [BÁP^{+}23, AJB20, BAW^{+}24, GPASM22, LPG^{+}24, MHB^{+}23, NZP^{+}21, PRCF22]. Price [dHJTCE23, GAB^{+}24, DQK^{+}23, HBW21, KAC^{+}23, NFAL^{+}22, SGZD20, YRTP20]. Prices [SK21b, LD23, LCN^{+}20]. primary [dAGCR21]. primers [HSL^{+}22]. Principles [HHLL25, Cop24]. Prionace [MCC^{+}22, MCC^{+}23, PDC^{+}24, SCD^{+}22]. prior [BCOBB^{+}23, HC22a]. prioritization [Sat23]. Pristiophorus [BRN^{+}20]. Pristipomoides [SKW^{+}21]. private [CWM^{+}23]. probabilistic [OWF^{+}23]. probability [JND^{+}23, ŠBB^{+}22, WH23, WSF22]. problem [DCS24, Mac22]. problems [IOO^{+}24]. procedures [BSR^{+}22, TRS^{+}24]. process [FCSA21, İTAD24, LWH^{+}23b, LBD24, MUF^{+}22, Sat23, TMH23]. processes [LDM^{+}24, MDL^{+}21, MSV21, MdCG20, SM21, Szu22]. processing [HA23]. Procypris [ZOZW22]. produced [BKHA21, RFF^{+}22]. producers [PPC23a]. producing [Aks24]. product [LSZ^{+}23, ZDF^{+}22]. production [BP20, CKDP^{+}20, CAAFH21, GPW^{+}20, HLCC22, HK22, JC21, KBK^{+}24, KHK^{+}20, LKSiI22, WCT^{+}20]. productive [CGM^{+}22, DSB^{+}21]. Productivity [CCR24, NSQV22, BTC23, BBPT^{+}24, FM21, HPS^{+}24, NS23, PJNGJ^{+}22, PDJ24, SAdC20, dAGCR21]. products [HOQ20, RSD23, UPBH^{+}20]. profile [LTT^{+}23, QMGRIU22, SDBS21, vdHC20]. profiles [MZZ^{+}21, PZG^{+}20, PJOR20]. profit [PDF20]. profound [SAA23]. program [BBC^{+}25, LSM^{+}23]. programme [SPC^{+}25]. programming [JMS25]. programs [HMS^{+}22]. progress [Ano22q, SBC^{+}23]. prohibited [ECM^{+}24]. project [MAH^{+}22]. Projects [SK21a]. promote [KDdOM^{+}22]. proof [CHGC25, WGNM24]. properties [ALRA20, CHGC25, LPAE^{+}24, TAK^{+}23, VTSI^{+}24]. prospects [YH21a, YH21b]. protandrous [BCM^{+}21]. Protected [BSA^{+}23, CRSC22, GGMMMV^{+}20, RAE^{+}21, SHS20, AL22, CTCB22, DSP22, KMC20, NOL23, NMS^{+}22, RLO^{+}21, Spa24, TdL24, WEH^{+}25]. Protection [RGP^{+}23, BC20, LSM^{+}23, SLW^{+}21, SZXC25, TCVG20]. protections [TSPK24]. protocols [CPR^{+}24]. prototype [SFC21]. provenance [KD22]. provide [BRR^{+}21, CPM21, CTR^{+}21, KHK^{+}20]. provides [BJK24, FCMP23, SMLT24, Ten22]. providing [NZP^{+}21, PFFdC22]. Province [AIM^{+}23, XqRJ^{+}23, DDA^{+}20]. provincial [GLA^{+}20, HMC^{+}23]. provision [FKS^{+}20]. provisioning [ESB^{+}24]. proxies [Alv21, HH20]. proxy [LZY^{+}24c, QMC^{+}22]. PSATs [NEBP^{+}23]. pseudo [CPM21]. pseudo-cohort [CPM21]. Pterois [Dik24]. public [HSW25, SPE^{+}23]. Puck [FGSD25]. Puerto [ECM^{+}24]. puerulus [JvPOG25]. PUFA [Lun25]. pulchra [HTK^{+}24]. Pulling [RMNB^{+}21]. pulse [BRvL^{+}22]. pulse-trawl [BRvL^{+}22]. punctatum [FTB^{+}21]. punctatus [BKR^{+}22]. purse [BGL^{+}22, CEAL21, CNDDAPMR21, DDCNMR21, LCLM23, LCDM^{+}24, NLS21, PCCMOA^{+}24, RS21, RBG^{+}20, RLO^{+}21, SASB24, SFC21, TFC^{+}20]. purse-seine [CNDDAPMR21, DDCNMR21, LCLM23, LCDM^{+}24, PCCMOA^{+}24, RLO^{+}21, SFC21]. pushing [RMNB^{+}21]. Pyhäjärvi [SHV20]. pyramid [BWB^{+}24].

quadrifilis [BCM^{+}21]. Quality [SGZD20, BBC^{+}20, DCS24, JBL^{+}22, JM23, JCL^{+}21, LBD24, MBB^{+}23, RS21, SGZH^{+}22, VTSI^{+}24]. quantification [DHH^{+}22]. quantify [BKM23b, DNLM23, DBDT21]. Quantifying [NLL^{+}25, NBEI23, OOAF^{+}21, SGW^{+}20, SB20, BCPH22, GJGW23, MJD^{+}21, PJP20]. quantitative [BČD^{+}21, HRH22, VCG^{+}23]. quantities [PM23]. quantity [JM23]. quarter [SBJ^{+}20]. queen [OOAF^{+}21]. Queensland [CLY^{+}22, MFR22a, SSP^{+}23]. Questioning [TGG^{+}24]. questionnaire [YTSS22]. questions [SBC^{+}23]. quickly [RB22]. quiet [ERS^{+}23]. quota [LD23, LMJ^{+}23, WZL^{+}22]. quotas [ASE21].

R [KK22]. r4ss [TDJ^{+}21]. rāhui [AVCA22]. rabaudi [ZOZW22]. radio [SV23]. radiocarbon [BKHA21, CSB^{+}23]. Rainbow [RPL^{+}24, ECY21]. rainfall [dAGCR21]. raised [SME^{+}24]. Raja [TLV23]. Rajidae [ABF^{+}21, CHM24, SHS20]. Rajiformes [FMLSP20]. Ramadan [PCGG20]. RAMP [WCGB22]. Ramsar [DBDT21]. ramsayi [BRR^{+}21]. Random [Aks24, GC25, GRJW20, LDM^{+}24, LZX^{+}20, SM21]. range [BTML20, BNL^{+}23, BRN^{+}20, ÖA21, WKSF20]. ranges [CdSLP21, CLM^{+}22, SB24]. Ranina [MWTH23, WPLF20]. Rapa [CGDTSA^{+}25]. RAPFISH [BMA^{+}24]. rapid [BAC^{+}22, HZZ^{+}20]. rapidly [CGC24, PJSQ20]. rare [CC20, MMBH23]. Rastrelliger [KPS20, KPWS21]. rate [BSR^{+}22, CLY^{+}22, FNKY20, GKC21, HMY25, HC22a, HHLL25, LZY24a, LCHB^{+}24, MDC^{+}22b, Sim23, SF20, vdHC20]. rates [AAG22, AAT^{+}21, BMC20, BBC^{+}20, BN21, CAYM^{+}23, DHH^{+}22, GGG^{+}22, HSM21, HRC23, HTSJ23, HW21, HTK^{+}24, JYH21, KBPS21, KBPS22, KLBHK23, LOFS22, LWH^{+}23b, Lya20b, Lya20a, MCC^{+}22, NLW^{+}22, PPM^{+}23, PVPN22, PASdCF23, PBRT22, PDF20, SLF23, SEM^{+}23b, YIM^{+}20, YSB^{+}21]. Ratio [CBHS24, KSI20a]. ratios [ZOS^{+}23]. ray [BTB^{+}21, BMM^{+}24, dSRFFN^{+}20, GRHHGM^{+}20, KMSJ^{+}25, TLV23]. rays [CSTR^{+}21, NDRR20]. razor [BBHF25]. Re [ATA^{+}24, BBHF25, PUC^{+}23, HPD^{+}22, SSP24]. Re- [PUC^{+}23]. Re-assessing [BBHF25]. Re-evaluating [ATA^{+}24]. re-examination [SSP24]. re-specifying [HPD^{+}22]. reached [Mau22]. readers [SF20]. reading [AYTM21, KD22]. real [RB22]. real-time [RB22]. reared [BTB^{+}21, GQP^{+}25]. rearing [JAN23]. rebuilding [BDM^{+}20, MPH21, OS21a, OS21b, ODM20]. recapture [BNTK23, Det23, DMM^{+}23, DAD^{+}22, Fis25, GW21, HPPT24, HGHH25, LCG^{+}21, Lya20b, Lya20a, PTL^{+}24, SKW^{+}21, SHS20, VBB20]. recapture-conditioned [VBB20]. recaptured [TBÓ^{+}22]. recaptures [CLY^{+}22]. receiver [KWE^{+}21]. Recognition [WZL^{+}22, KSJM^{+}20, XDX^{+}23]. recommendation [KSI^{+}22]. Recommendations [TCJ^{+}21, EBN^{+}23, GBC23a, PK24]. recommended [BBJ22]. reconstruct [GM21, HKG^{+}21]. reconstructing [SNHM23]. Reconstruction [dlTGPLC23]. record [HMR^{+}24b, SDdMG^{+}20, WSF22]. Recorder [DNLM23]. records [SSM^{+}23]. Recovery [vP20, CSDH^{+}23, KFDE^{+}22, MHD^{+}21, SHV20, SCW21, WS20]. Recreational [AGB^{+}20, LVP22, LCHB^{+}24, MKC20, RBG^{+}24, SCD^{+}22, dS21, AYTM21, AMHR22, AWC^{+}23, ALRA20, ARD^{+}23, AK23, BLE^{+}22, BGCCP22, BSAP22, BHNP22, BALBC23, BKM^{+}23c, BAYR^{+}24, CLM^{+}22, CAYM^{+}23, DGMG^{+}22, DBDT21, DTSR22, DET21b, DAL20, FMLC^{+}22, FCML^{+}22, FWL20, GPWP20, GFC^{+}22, GAB^{+}20, HJA^{+}21, HMS^{+}22, HCDF24, JDP22, KAB^{+}22, KC22, KKLM24, KWMA23, KFDE^{+}22, LRMH21, LPG^{+}24, LTE^{+}23, LWH^{+}23a, LHPR21, LJB^{+}24, LRGB25, MFR22a, MOB^{+}22, MDC^{+}22b, NMS^{+}22, PZL^{+}23, RFMS^{+}21, RTHB25, SBC^{+}23, SGW^{+}20, Sch23, SAW^{+}20, SLA24, SR20, SRF^{+}24, SL24, TC24, TPW23, TCL^{+}24, VP22, YTSS22, ZCFG23, vdHBBR20, vdHC20]. recreational-grade [LJB^{+}24]. recreationally [BCPH22, CGB^{+}22]. recruit [HW21, SCG^{+}24]. recruiting [HGS^{+}23]. Recruitment [dlTGPLC23, BDA^{+}24, Bro24, DSS^{+}23, DBS^{+}21, EYAO20, FNKY20, GSH22, HW21, JHB21, KSI20b, LWH^{+}23b, MDL^{+}21, MdCG20, Mau22, MNS^{+}20, OMK24, PZL^{+}23, SCCAM21, SSKS21, SSP24, vDJB^{+}23]. recruits [AGS20, AAFLL^{+}25]. Red [OAM^{+}21, VKS^{+}25, AVMBEB22, AAH^{+}23, BPT^{+}25, BHNP22, CPL^{+}25, FBQA20, FGCB^{+}21, GPWP20, GAB^{+}22a, KK21, LR21, LYH^{+}21, LDC24, MSJ21, PJSQ20, PHH^{+}23, RL24, SH22a, SW20, SACS23, SBT^{+}20, WSB22, AAZ20, SECB21]. Redfish [SMLT24, BW23, CRS23, HP23]. redleg [PHP^{+}20]. reduce [AL22, CCKL^{+}20, DML^{+}20, DQMV21, FMMC20, JLYR24, KBB^{+}21, KSS^{+}22, LWH^{+}21, MB23, RBHM24a, SBB^{+}24]. reduced [AIJ^{+}23, BČD^{+}21, LPG^{+}24, PZL^{+}23, UAB^{+}21, XMLCMV24]. reduces [KNO^{+}21]. Reducing [BDR^{+}20, BM25b, VS23, SME^{+}24, FS20, LTE^{+}23, ZZH^{+}24, dJDM23]. reduction [BGCCP22, BM23b, BM24b, FOM21, GBB20, HLMV24, JBČ^{+}22, RB22, YSB^{+}21, YLS^{+}23, ABT^{+}24]. reed [NHE^{+}23]. reedi [SCCAM21]. Reef [AGL^{+}24, GGMMMV^{+}20, ASJ^{+}22, BTML20, BTFL22, BLFT23, BPT^{+}20, CSSB22, HPL^{+}24, LWX^{+}20, MTE^{+}20, MKFF^{+}21, NZP^{+}21, PDJ24, RBHM24a, STH25, SSKS21, WCC24, WSUN^{+}23]. reef-associated [BPT^{+}20]. reefs [BBM^{+}24, BHNP22, FGCB^{+}21, MOA23, SEM^{+}23a, VR20, WHR^{+}24, WASS20, WPB22]. reel [KWC^{+}20]. Reexamining [CWM^{+}23]. reference [AT20, KSO^{+}21, LZY^{+}24c, SMA^{+}24, ZF21]. refine [CTS^{+}23]. Refining [BM25c, BBSM24]. reflected [FGL22]. Reflecting [LMJ^{+}23]. reflectivity [KSS^{+}22]. reflects [SSSF25, VA20]. Reflex [AGNS^{+}21, RBHM24b, BSR^{+}22, BKR^{+}22, KC22, LBP^{+}24, RPL^{+}24, WCGB22]. refugia [NZP^{+}21]. regalis [RAG22]. Regan [BRR^{+}21]. regia [DSB^{+}21]. regime [HUK^{+}23]. region [BTB^{+}21, CCC^{+}21, FTB^{+}21, GMPD23, HMC^{+}23, HL20, KNS^{+}22, MPP20, SPC^{+}23, SdFZFJ21]. Regional [HFMH20, EHBJ20, GPT^{+}21, Hut22, KHS^{+}20, SRP^{+}22]. regions [FDdCS^{+}20, SSM^{+}23]. regression [Aks24, CYBW22, CG21, İTAD24]. regular [IB20]. regulated [YTSS22]. regulation [GHAZ21, RBD^{+}22, WY20]. regulations [BKM^{+}23c, CFB^{+}23, GPWP20, KWMA23, KMA22, LWH^{+}23a, SRF^{+}24]. regulatory [SGW^{+}20]. Reinhardtius [BKHA21, GLA21, SC20, TTK^{+}25]. rejected [PTD^{+}20]. related [Béc20, EFM25, GMK23, LCHB^{+}24, MLCMdS23, YCC22]. relation [FTB^{+}21, JAN23, NNS^{+}22, PDC^{+}24, DBL^{+}25, TSS^{+}23]. relations [HLCC22]. Relationship [aFLpX^{+}21, BPT^{+}20, FG21, GBC^{+}23b, MRC24, TTYT24, WKSF20]. Relationships [BALBC23, HMR24a, PASdCF23, FNKY20, GPASM22, HW21, SWH^{+}24a, SAA^{+}22]. Relative [BM24b, IB20, BGG^{+}22, BFA^{+}21, BH23, CMV21, FRP22, HCDF24, JMP^{+}21, PMC^{+}24, TMP20]. Release [CGSL22, RSPE22, AWVS21, AGNS^{+}21, AK23, BNTK23, BCS^{+}22, BLC^{+}22, BR22, BSR^{+}22, BALBC23, BAC^{+}22, CH21a, CAYM^{+}23, CHAY^{+}25, CAH^{+}20, GFC^{+}22, HEG^{+}23b, JAN23, KFDE^{+}22, LCB^{+}21b, LPG^{+}24, LCG^{+}21, LDC24, LCHB^{+}24, MSW21, MBH^{+}22, PSS^{+}21, RB22, RBG^{+}24, SFCG^{+}21, SAW^{+}20, SGW^{+}22, SFJ^{+}23, SACS23, SGH^{+}20, TPW23, YTSS22]. released [CBR^{+}23]. releases [JSKM20]. relevance [BHB^{+}22, FDS^{+}23, GAB^{+}22b]. relevant [KSL^{+}23, MDL^{+}21]. Reliable [DAD^{+}22, DCS24, MBOCdAM24]. remains [BTFL22]. Remote [KMO20, KBM23, CSSB22, LVP22]. Removal [MB20, BCPH22, CCC^{+}21, CCC^{+}22a, CLD^{+}22, DJFU20, GC21]. removals [ALW^{+}21, GM21]. rent [Ham22]. rent-seeking [Ham22]. repeat [PPC^{+}23b, Pun23]. replaced [vORP23]. replacement [Bea21]. replacing [Tho19, Tho21]. replicate [ELM20]. replicated [HGS^{+}23, JC21]. replication [PF20]. reported [HSM^{+}24]. reporting [AAG22, FFG21, vdHBBR20]. reports [WSF22]. representation [AGS20]. representativeness [FOS^{+}21]. reproducibility [Sat23]. reproducible [CTIC23]. Reproduction [CW22, FCMP23, FHE^{+}24, MGB24, MBOCdAM24, MABR^{+}20, Ten22]. Reproductive [AHEV24, ASO^{+}22, ECY21, GGMMMV^{+}20, KCB^{+}24, LCM^{+}23, aLBK^{+}21, MFM^{+}20, MPM^{+}23, OMG^{+}23, PBPM^{+}23, RAG22, Ash20a, Ash20b, BGK22, CH21b, EPHDB24, FNH^{+}21, FBPCC^{+}21, GBO^{+}20, HSSD^{+}21, HIF^{+}24, KSI^{+}22, MPM22, MCC^{+}23, MSS^{+}21, MRUG^{+}23, PZL^{+}23, PDC^{+}24, SF22a]. reproductive-phenological [HIF^{+}24]. Republic [GFC^{+}22]. requests [YTSS22]. requires [CHPT20]. requiring [TRS^{+}24]. Res [Ash20a, DBV23, KBPS22, OS21a, SIM^{+}24a, Tho21, VK21a, WGC^{+}21, YH21a]. Research [DSJG20, ZCFG23, BMJ^{+}24, CBD^{+}22, DHH^{+}22, HEG^{+}23b, KK21, MCGL^{+}25, PRK23, SO25, SBC^{+}23, TLAM25, vBMP^{+}23, Mel23]. reserve [LWX^{+}20]. reserves [BMSM22]. reservoir [KDdOM^{+}22, LJB^{+}24, PW25, ZSWK25]. reservoirs [PW25, PSS^{+}20, PRWK20, WW21b, WLG^{+}23]. residency [CSDH^{+}23, FJHT^{+}22, FGCB^{+}21]. resident [CHL^{+}20, LOFS22]. residual [BCSM20]. residuals [TAK^{+}23]. resilience [BALBC23, HCCC21, LPS^{+}25]. Resolution [JGU21, Dik24, HEGR24, HEG^{+}23a, LCFJ22, PVA^{+}24, PÁEMC22]. resource [Cop24, GZLCRG25, GBO^{+}20, HdLHD22, JZQZ20, KAB^{+}22, KA24, SHC21]. resources [BMOC22, CRCAF^{+}22, DSJG20, Ham22, KK21, KSI20b, MCH21, SZXC25, TTC^{+}25, TCVG20, WLZ^{+}21, dCHdMS^{+}23]. respect [VTSI^{+}24]. responding [SS23]. Response [BWN^{+}21, TRN^{+}23, AGNS^{+}21, CCC^{+}22b, GAB^{+}22a, GRG24, GFC^{+}22, LMM^{+}24, LZCC24, SSG^{+}22, TLC^{+}22, UPBH^{+}20, WGW^{+}23]. responses [BPT^{+}20, CMA^{+}22, LRGB25, MSJ21, RNP^{+}24]. restaurants [SCHSC21]. restocked [BÖN20, Sim23]. restocking [CHL^{+}20]. restoration [LDS^{+}21]. restricted [DTSR22, SR20]. restriction [Lya20a]. restrictions [HJA^{+}21, SCC^{+}22]. result [HW21, SCC^{+}22]. resulting [PZL^{+}23]. Results [AMM^{+}22, PRN^{+}24, TSC^{+}20, EYAO20, Mel23]. retention [FFG21, HGC^{+}21, MdCG20, MCC^{+}22, NMJ^{+}24, SW20, vBMP^{+}23]. retentive [SSY20]. retinal [SSG^{+}22]. retraction [AGL^{+}24]. retrieval [AHL20]. return [AAG22, HC22c]. returns [CSTdL25]. Réunion [GLP^{+}20]. reuse [BBC^{+}25]. reveal [BLFT23, CPPK23, CSRL20, LM22, LZL^{+}24, NVRG^{+}21, NDRR20, SPW^{+}22]. revealed [BÁP^{+}23, BLE^{+}22, FJHT^{+}22, FBW^{+}21, HMS^{+}22, JCL^{+}21, LMT^{+}22, PKRL21, SBC^{+}22, SAA^{+}22, VBL^{+}22, WWT^{+}25, YLX^{+}24]. Revealing [SCG^{+}24, MZZ^{+}21]. reveals [BQGV^{+}24, CPB^{+}21, DJFF23, HIF^{+}24, KHPB20, MGC^{+}22, NAV^{+}23, PW25, PMC^{+}24, SNJ^{+}24, TSI^{+}21, XqRJ^{+}23, ZSDZ24, ZJJ^{+}25]. revenue [ASE21, MRG^{+}23]. reversal [MHB^{+}23]. Review [DC24, VR20, DD25, FGL22, GPASM22, LPRB^{+}21, MHL^{+}23a, NCB^{+}23, NK23, PVC^{+}22, PK24, PJP20, Sat23, Spa24, SF22b, SB24, WY20, YAO^{+}23, dSTMV20]. revisit [ABK^{+}21]. revisited [SAA23, WKSF20]. Revisiting [AIW^{+}23, BTFL22]. Rexea [ODM20]. rhadinum [SLW^{+}20]. rheophilic [Lya20b]. Rhizoprionodon [KCB^{+}24]. Rhomboplites [CSB^{+}23]. rhombus [SSV^{+}20]. Ria [MGB24, OOAF^{+}21]. Rican [OPL21]. Rice [FKS^{+}20]. rich [Cla22, Mon24]. richardsoni [CSRL20]. Rico [ECM^{+}24]. Right [GM21, HLC^{+}25]. Rights [GHAZ21]. ring [PCM^{+}21, UFYT23]. ringens [MVLC^{+}20, PAY23]. riparian [NMS^{+}22]. Rise [LPP^{+}20, JLYR24]. rising [GLA21]. Risk [TMDA22, ASE21, ASJ^{+}22, DCL^{+}20, aFLpX^{+}21, JDH22, MVDH24, PBDM23, PDJ24, WXJ^{+}24]. risk-catch-cost [PBDM23]. risks [SH22a]. Risso [MABR^{+}20]. River [AAVÁM23, BKC21, CBR^{+}23, KSI20a, Sch23, ZF21, BN21, EMR^{+}22, HL21, JAN23, LVP22, LFdB^{+}21, LP23, PRN^{+}24, SdFZFJ21, aFLpX^{+}21, FDdCS^{+}20, KMC^{+}23, LYX^{+}21, PASdCF23, RBH^{+}24, SLW^{+}21, TTC^{+}25, VBL^{+}24, XqRJ^{+}23, ZHM23, ZOZW22]. riverine [SJW^{+}22]. Rivers [HKKa^{+}25, ACP^{+}23, JM23, JCL^{+}21, LCN^{+}20, LHPR21, Lya22, SNJ^{+}24, WLG^{+}23]. rivulatus [Dik24, WCLN20]. robust [CYBW22, DCR^{+}20, SBZ^{+}21]. robustness [SSP24, XMCC20]. Rock [BHG^{+}24a, MMML24, BHG^{+}24b, ML24, PDM^{+}24, Pun24, dLHER24]. rockfish [BHB24, GBWM22, JRW^{+}21, KHE^{+}22, TSG25]. rockfishes [CH21b]. rod [KWC^{+}20]. rodgersii [BWB^{+}24]. Role [DAL20, LPAE^{+}24, CBJ24, FCML^{+}22, MCK23, MHB^{+}23, PBD25, PPC^{+}21, RBD^{+}22, WW21a, YTHM20, YSB^{+}21]. roller [ASWS^{+}21]. room [dLHER24]. roosterfish [JCCAS^{+}21]. rope [KFI^{+}21]. ropes [IB20, SHB^{+}23]. rose [PDG^{+}22]. rosenbergii [PNRS23]. Ross [GMPD23, MPP20]. rostrata [LYLC21]. rotundicauda [HSM^{+}25]. Roughhead [BHH21]. roughy [BCOBB^{+}23]. round [BAJB^{+}24, BBC^{+}20, CBTH20, DMZH21]. route [SBJ^{+}20, YTH22]. roving [DBDT21]. Roweothuria [AHEV24]. rubra [HSM21]. ruby [WWF^{+}20]. Rügen [FDS^{+}23]. rule [BBPT^{+}22, WH23]. Rules [MJD^{+}21, BBPT^{+}24, HSSD^{+}21]. run [HPS^{+}24, TAA^{+}20]. rural [KKLM24]. Russia [DD25]. Rutilus [RKN23].

S [LR21, MFM^{+}20, SW20, TNDM23]. S. [BFA^{+}21, Dik24]. sablefish [CGC24, GSS^{+}23, KHC^{+}20, SNHM23]. sac [SMC^{+}24]. safety [Sch23]. sagax [DLP^{+}24]. Sagmariasus [ML24]. saira [HLZ^{+}20, LZL^{+}24]. salar [HUK^{+}23]. sale [WSB24, dCHdMS^{+}23]. salinity [CdSLP21, TLCD21, VBL^{+}24]. Salminus [ACP^{+}23]. Salmo [HUK^{+}23, JAN23, LR21, TAA^{+}20, ZDZR^{+}22]. salmon [ACS23, AGB^{+}20, BQBW20, Bea21, BJS^{+}22a, BJS^{+}22b, BKC21, CSDH^{+}23, CCC^{+}22b, CBR^{+}23, DBS^{+}21, EMR^{+}22, FHHH20, FAK24, GQP^{+}25, HPPT24, HUK^{+}23, HKKa^{+}25, IRJ^{+}22, KD22, KFI^{+}21, LVP22, LP23, LCHB^{+}24, MOB^{+}22, OS21a, OS21b, OKKW20, PPC^{+}23b, RGN^{+}20, RSPE22, SS23, SFYM24, TBÓ^{+}22, WBA23, YSB^{+}21]. salmonid [KHL^{+}24, LCFJ22, TJS23, YTSS22]. Salvelinus [DH20, JSKM20, LvCdGSL23]. SAM [ATAS20, CAZN24]. same [BTFL22, MKC20]. Sample [EYAO20, DH20, HW24, LJB^{+}24, LZX^{+}20, NSRM22, WSF22]. Sample-Based [EYAO20]. sampled [HFKS20]. samplers [MTC^{+}22]. Sampling [MAA^{+}20, MMBH23, ASDW24, BČD^{+}21, BAC^{+}22, CZJ^{+}24, DBGV^{+}22, FOS^{+}21, FSP22, GAB^{+}22a, GRJW20, KHL^{+}24, Ken21, KK22, LCMS^{+}22, LCDM^{+}24, LCW23b, OWF^{+}23, PF20, PTD^{+}21, RPM^{+}21, ŠBB^{+}22, TBF^{+}21]. SAMs [SMK^{+}24]. San [AdlBCN25, CVM^{+}20, TLAM25]. Sanctuary [FFG^{+}20]. sand [SBL^{+}23]. sandbar [PCBL23]. sandeel [MDMS21]. Sander [EVS^{+}23, GLA^{+}20, GW21, HMC^{+}23, OGFW24, PSSFS24, SSD^{+}20b, SPD^{+}24, SSJ^{+}21]. sandfish [RFF^{+}22]. sanguinea [MOCGCC^{+}25]. sanpaulensis [RdLSdB^{+}21]. santolla [CVM^{+}20, DMM^{+}21, HRC23, LTR20a, MFM^{+}20, MRUG^{+}23]. Santos [MPEBdR23]. sapidissima [BHD^{+}23]. sapidus [DMNV^{+}24, WCGB22]. Sarda [ORdlG^{+}24, SKST23]. Sardina [LMT^{+}22, NSM^{+}21]. sardine [BGLP21, CMIMS20, DLP^{+}24, FNKY20, KTFY22, LMT^{+}22, MVLC^{+}20, NSM^{+}21, OCdMC24, RdLSdB^{+}21, SIMT24, SSP^{+}22, SSS^{+}23, WHM23]. sardinella [BAJB^{+}24, SSP^{+}22, SSS^{+}23]. Sardinia [SPG^{+}21]. Sardinops [DLP^{+}24, WHM23]. Sargassum [MKS^{+}21b]. satellite [FDdCS^{+}20, LZC^{+}21, NGDC25, NEBP^{+}23, OOM^{+}23]. satisfaction [BHB^{+}22, GAB^{+}22b, MFR22a, YTSS22, vdHR23]. saucer [CLY^{+}22]. Saudi [AAM^{+}20]. saury [HCK^{+}21, HCDB22, HLZ^{+}20, KPUB22, LZL^{+}24]. sawfish [PBB20]. sawsharks [BRN^{+}20]. saxatilis [GSH22, WMT^{+}20]. sBCB [VK21a, VK21b]. scabra [RFF^{+}22]. scad [WSL21]. scale [ACS23, APB^{+}20, BAM^{+}24, Bea21, BBC^{+}25, BSKL^{+}22, BALBC23, COG22, CPF20, CGT^{+}23, CDAK23, CTCB22, DPL20, DSP22, DSNK^{+}22, DMS22, DHCS23, ELM20, ECM^{+}24, FBM^{+}21, FBW^{+}21, GFM^{+}23, HA23, IOO^{+}24, KM23, KHS^{+}20, KFI^{+}21, LLFL21, Lya20a, Mac25, MWJ^{+}24, MSV21, NB24, NFAL^{+}22, PAA^{+}24, PM23, PSS^{+}20, PJNGJ^{+}22, QAS^{+}25, RFMS^{+}21, RUHM20, SCG^{+}24, SSM^{+}23, SBB^{+}24, TCJ^{+}21, WPB^{+}20, WKBMW24, YRTP20, YLS^{+}23]. Scalefish [AMHR22, BBPT^{+}22]. scales [BHD^{+}23, JDH22, KHS^{+}20, TBÓ^{+}22, WBA23]. Scaling [HL20, DAR^{+}23]. Scallop [EDA^{+}22, SDBS21, CFB^{+}23, MSV21, OOAF^{+}21, PCM^{+}21, PCF23, TC24, TLAM25]. scalloped [BFA^{+}21, GAW^{+}22]. scallops [CLY^{+}22, HC22b]. scampi [MTS^{+}21]. scan [LJB^{+}24, PBB20, WLG^{+}23]. scattering [BPT^{+}20]. scenarios [CCR24, GdSPL21, Har21, TRN^{+}23]. Schaefer [DQK^{+}23]. schemes [PTD^{+}21, WXJ^{+}24]. school [APGG22, BGL^{+}22, BM22]. schooling [VMI21]. Sciaenidae [AVACA^{+}23, HCCC21]. Sciaenops [PHH^{+}23]. Science [PK24, BMM^{+}21, Cop24, CRF^{+}24, GAB^{+}20, HMS^{+}22, PJNGJ^{+}22, PRK23, Sat23, SGW^{+}22, WSL^{+}24]. science-based [Cop24]. Science-informed [PK24]. scientific [GWMC21, HEG^{+}23b, JDH22, LVL25]. scientists [SCC^{+}22]. Scomber [CMTP^{+}21, MMM^{+}20, NVSG24, PFC^{+}21, SCN^{+}24, WYM^{+}25]. Scomberomorus [FHE^{+}24, PYX^{+}20, SDC^{+}22]. scombrus [MMM^{+}20]. Scope [CCC^{+}22b]. Scophthalmus [BGK22]. Scorpaena [SBT^{+}20]. Scorpaenidae [SBT^{+}20]. scorpion [SBT^{+}20]. Scottish [BBHF25, MDS^{+}20, MDJP24, SHH^{+}21, SMKJ21]. Scrabbling [PGRD21]. script [KK22]. Scrobicularia [MGB24]. scrofa [SBT^{+}20]. sculpin [TEO25]. scup [CG21]. Scylla [HTSJ23]. Sea [ASJ^{+}20, AMSC20, DSJG20, MPM^{+}23, OBB^{+}20, ÖÜÖG20, PDG^{+}22, SKBA23, SZXC25, VMI21, YLS^{+}23, AJB20, AUHK22, AHEV24, BWB^{+}24, BR22, CJC^{+}20, CPPK23, CG21, DDA^{+}20, HC22b, HBE^{+}22, HdLHD22, KFO20, Mac22, OKKW20, OWF^{+}23, PMS^{+}23, PMS^{+}20, QAS^{+}25, SPC^{+}25, SWH^{+}24a, SGW^{+}22, SFJ^{+}23, TTK^{+}25, VCPO21, XDX^{+}23, ASGG21, AAZ20, ABF^{+}21, AIJ^{+}23, ALW^{+}21, ARD^{+}23, AIW^{+}23, BGK22, BTB^{+}21, BQGV^{+}24, BDR^{+}20, BMOC22, BHH21, BLE^{+}22, BGM^{+}23, BLK23, CAZN24, CMD^{+}23, CPL^{+}25, CDAK23, DRSPTA20, DAR^{+}23, DSNK^{+}22, Dik24, DKBF23, EBN^{+}23, EYAO20, EHE^{+}23, EHBJ20, EM23, EFM25, FHSC21, FJS^{+}25, FGSD25, FDS^{+}23, FTH^{+}23, GCK^{+}21, GMPD23, GTS^{+}21, HSJ^{+}24, HML^{+}20, HSSD^{+}21, HG20, IB20, JBL^{+}22, JFS21, KLBHK23, KMA22, KDF^{+}25, LCB^{+}21a, LWH^{+}23a, LZC^{+}21, LSZ^{+}23, LKSiI22, LZCC24, LVPS20, LNvD^{+}23, MHL^{+}23b, MPP20, MYKO23]. Sea [Mur20, Mur21, NLW^{+}22, NVB^{+}23, NFdSJO25, OAM^{+}21, ÖSL^{+}23, PFC^{+}21, PDC^{+}23, PGD^{+}25, PM23, PDC^{+}24, PDD^{+}22, PDA^{+}24, RKN23, RAG22, RDR^{+}23, RSPE22, RL24, SSV^{+}20, ŞGK^{+}20, SBC^{+}22, SECB21, SPM^{+}24, SC20, SPC^{+}23, SFYM24, SXK^{+}24, SMA^{+}24, SMC^{+}24, SV23, Szu22, TAA^{+}20, TBH^{+}22, Ten22, TRWH23, WLZ^{+}21, WBBG^{+}23, XWD^{+}21, YTHM20, ZZC^{+}21, ZZH^{+}24, ZSDZ24, vDJB^{+}23]. Sea-gear [Dik24]. seabed [HEG^{+}23a, MKH^{+}20, TRWH23, TLAM25, XDX^{+}23]. Seabird [PCF23, OKKW20]. seabream [CRF^{+}24, vBMP^{+}23]. seafloor [BDR^{+}20]. Seafood [MBB^{+}23, Béc20, PPC23a, SK21b, Seu22]. seagrass [FWKR21]. seahorse [NLL^{+}25]. seal [GKKL24, IAB20, KLBHK23, LLK^{+}22, LLS23, PBDM21, WOG^{+}25]. seal-commercial [PBDM21]. seal-free [LLS23]. seals [KFI^{+}21]. seam [JMP^{+}21]. seamounts [APGG22, GSS^{+}23]. seas [GdSPL21, JWL^{+}24, LRW^{+}24, BMM^{+}24, MZZ^{+}21]. Seascape [CRSC22, SKS^{+}23]. season [BHNP22, EMJ^{+}22, HC22c, NRH^{+}23, PZL^{+}23]. Seasonal [AAPG21, BBC^{+}20, GDVBB^{+}20, GBO^{+}20, MOCGCC^{+}25, TTYT24, TYYK21, vZvdHCA25, EHE^{+}23, FAK24, HLI^{+}20, JWL^{+}24, KNS^{+}22, LJX^{+}20, LWX^{+}20, NLS21, OYOO21, SSM^{+}23, ASD^{+}22]. seasonality [HCDF24]. seasonally [KHPB20]. seasonally-varying [KHPB20]. seasons [ECY21]. seaward [PRWK20]. seawater [TLC^{+}22, UAB^{+}21, YLP^{+}23]. seaweed [ALRA20]. Sebastes [BW23, CRS23, HP23, TSG25]. secondary [AdABW^{+}22, BM23b]. sectioned [CSRL20]. sector [Apo25, BGCCP22, WW21a, MFM^{+}20]. sectoral [Apo25]. sectors [SHC21]. security [CNE^{+}22, SK21b]. sedentary [FBALRR^{+}22, PPH21]. sediment [LCW23b, SOSK22]. see [vdHR23]. seeding [AAG22]. seeking [Ham22]. sees [SEM^{+}23a]. SEEZ [SKST23]. segmentation [İTAD24, SRB^{+}25]. segregation [CNDDAPMR21]. seine [BGL^{+}22, BM23a, BM25c, CEAL21, CNDDAPMR21, DDCNMR21, EFM25, LCLM23, LCDM^{+}24, NLS21, OOM^{+}23, PCCMOA^{+}24, RS21, RLO^{+}21, SASB24, SFC21, TFC^{+}20]. seiner [RBG^{+}20]. Seismic [CMRP20, dLHER24, HLG^{+}21, MCMM20]. Selar [WSL21]. select [JRW^{+}21, WCT^{+}20]. Selecting [LYH^{+}21, WZX^{+}20]. Selection [dCHdMS^{+}23, AAH^{+}23, BJHS^{+}23, BBJ22, BM23a, BM25c, CBA^{+}24, EMR^{+}22, LP23, MCC20, OFF^{+}20, PF20, PD25, RBM21, RB22, SIMT24]. selective [AUC25, BGLP21, CTCB22, FBPCC^{+}21, LRGB25, XSS^{+}23]. Selectivity [ZZH^{+}24, AK23, BLHS20, BM22, BM25b, CSSB22, CFP21, DHB^{+}21, FMMC20, FS20, FHHH20, GWMC21, HGHH25, HFKS20, IB20, KHGB25, LVPS20, MINS21, MOB^{+}23, MB25, MOTL25, PHV^{+}21, PCM^{+}21, PHH^{+}23, PJMP22, SHB^{+}23, TBF^{+}21, WCT^{+}20, YHC^{+}24, dJDM23, vORP23]. self [GHAZ21]. self-regulation [GHAZ21]. selincuoensis [DHC^{+}20]. semi [ACP^{+}23, MGC^{+}22, YIM^{+}20]. semi-continuous [MGC^{+}22]. semi-fragmented [ACP^{+}23]. semi-native [YIM^{+}20]. semisulcatus [RKD24]. Senegalese [SKST23]. sensing [WST^{+}23]. sensitivity [CRCAF^{+}22, TCJ^{+}21]. separation [BGM^{+}23, LMT^{+}22]. separator [BKM23b]. Sepia [GMERCM^{+}24, KSI^{+}22, LCD^{+}23, SPC^{+}25, YLX^{+}24]. Sepiidae [GMERCM^{+}24]. September [Ano20x, Ano21x, Ano22y, Ano23x, Ano24x]. sequencing [YGMJ20]. Sergipe [NFAL^{+}22]. series [BCG^{+}25, GJSW22, LD25, MPSM25, SSSF25]. Seriola [AVB^{+}23, BGG^{+}22, WWT^{+}25]. Serranidae [SK22]. serrata [HTSJ23]. service [LVP22]. services [ESB^{+}24, FKS^{+}20]. SESSF [BBPT^{+}24]. set [EM23, HSL^{+}22, KLNB^{+}24, KFI^{+}21, LCLM23, LS24, MSC^{+}24, SH22b]. set-net [EM23]. Setipinna [HSJ^{+}24]. setnet [SOS^{+}21]. Seto [LKSiI22]. sets [OB21]. Setting [MHH^{+}20, WH23]. settings [PBB^{+}22]. settlement [HMR24a, SCG^{+}24]. settling [GSS^{+}23]. seven [CWM^{+}23, FGTA24]. several [RWT^{+}20]. severity [PPC^{+}23b]. Sex [CBTH20, CHGC25, MCS^{+}22, SPD^{+}24, WCLN20]. sex-composition [CHGC25]. Sex-specific [CBTH20, MCS^{+}22, WCLN20]. sexual [ASD^{+}22, CB20, GRHHGM^{+}20]. sexually [CW22]. Seychelles [GFC^{+}22]. Shad [HRH22, BHD^{+}23, XqRJ^{+}23]. Shag [NOL23]. shake [dLHER24]. shallow [LBLF20, LMK^{+}23, PFFdC22, PBB^{+}22, SFC21]. shallowest [BBČ^{+}21]. shape [BJHS^{+}23, CGDTSA^{+}25, CdSLP21, HLC^{+}25, JvPOG25, MSV21, MPP22, MNPMM^{+}22, MMM^{+}20, NSM^{+}21, NVSG24, SBC^{+}22, SCN^{+}24, SNJ^{+}24, SSD^{+}20a, SSD^{+}20b, TSG25, WQGS25, WJN^{+}25, dAdCdO^{+}23]. shaped [HFKS20, YLS^{+}23]. share [SAdC20]. shared [BAJB^{+}24, CPB^{+}21, SYZ^{+}25]. Shark [BBPT^{+}22, TKB^{+}21, AWVS21, AGL^{+}24, BTR^{+}24, CMA^{+}22, ECM^{+}24, FTB^{+}21, FMMA20, GPW^{+}20, HEG^{+}23b, KMSJ^{+}25, KFDE^{+}22, LCMS^{+}22, LNR^{+}21, MCC^{+}22, MCC^{+}23, PKRL21, PCBL23, SBD^{+}22, SSI^{+}23, TRS^{+}24, YHST22, dS21]. shark-tuna [FMMA20]. Sharks [KMSJ^{+}25, LNR^{+}21, BCSM20, BRGB^{+}23, BMJ^{+}24, CSTR^{+}21, DDCNMR21, FMD^{+}24, GLP^{+}20, KCB^{+}24, KFDE^{+}22, LSM^{+}23, MCPJET^{+}20, MCC^{+}22, MRC24, PDC^{+}24, SASB24, SFCG^{+}21, SCD^{+}22, SGD^{+}21, SGH^{+}20, TSC^{+}20, BBM^{+}24]. Shedding [MCC^{+}22, GGG^{+}22]. shelf [ABK^{+}21, BPT^{+}25, BHH21, CZ25, CBD^{+}22, DCK^{+}22, HCWH23, MFO21, MOA23, MCK23, Ten22, BRR^{+}21]. shelf-edge [MOA23]. shell [BQGV^{+}24, BBHF25, LCM^{+}23, MGB24, MB20, ZJY^{+}24]. shift [HUK^{+}23, JHB21]. shifts [Cam23, KHMC23, LXC^{+}20, Mur20]. ships [ERS^{+}23]. shoal [KSI20a]. shoot [OOM^{+}23]. shore [HJMS20]. Short [CGB^{+}22, EFM25, GFC^{+}22, HMM^{+}21, MVLC^{+}20, TRS^{+}24, AK23, BDA^{+}24, BKR^{+}22, CFB^{+}23, DBS^{+}21, DMS22, EMJ^{+}22, EM23, IB20, KPS20, KPWS21, LCB^{+}21b, NVB^{+}23, PVA^{+}24, PCC^{+}23]. short-beaked [PVA^{+}24, PCC^{+}23]. short-lived [CFB^{+}23]. Short-term [CGB^{+}22, EFM25, GFC^{+}22, HMM^{+}21, MVLC^{+}20, TRS^{+}24, BDA^{+}24, DBS^{+}21, DMS22, EMJ^{+}22, EM23, LCB^{+}21b, NVB^{+}23]. short-time [AK23]. shortened [LTE^{+}23]. shortening [SHB^{+}23]. Shortfin [OB21]. Should [HSSD^{+}21]. show [CHB24, RFF^{+}22, SBJ^{+}20]. shrimp [AAM^{+}20, ASWS^{+}21, AGNR^{+}21, BAM^{+}24, CKD^{+}21, FDB^{+}20, GTS^{+}21, IJS^{+}22, LLFL21, MRG^{+}23, NFAL^{+}22, ÖRS^{+}25, PDG^{+}22, PBB^{+}22, SCCAM21, SS24, TBH^{+}22, VMFF^{+}20, ZF21]. shrimps [CPL^{+}25, PFFdC22]. shrinks [WCC24]. siamensis [TTC^{+}25]. Siargao [CRSC22]. Siberian [TEO25]. sibling [LLH^{+}25]. side [ELM20, JM23, LJB^{+}24, PBB20, WLG^{+}23]. side-looking [JM23]. side-scan [LJB^{+}24, PBB20, WLG^{+}23]. sides [YTHM20]. sidestripe [ZF21]. Siganus [Dik24]. sigmoidal [AVMBEB22]. signals [CCKL^{+}20, LIA25]. signatures [CMTP^{+}21, MMM^{+}20, SSP^{+}22]. significance [TEO25]. significantly [CCKL^{+}20]. siliqua [dAGCR21, BBHF25]. silky [SASB24, SFCG^{+}21]. Sillago [BM22, BM23a, BM25b, BM25c, YGMJ20]. silver [BSAP22, aFLpX^{+}21, SLW^{+}21]. silverside [CdSLP21, DSB^{+}21]. similarly [KWMA23]. simple [Det23, FOM21, MINS21]. simplify [Tho19, Tho21]. SIMS [RWFT25]. SIMS-based [RWFT25]. simulate [MZ24]. simulated [BSR^{+}22, EMR^{+}22]. Simulating [GPWP20, NAS^{+}20]. simulation [BGCCP22, DBGV^{+}22, HCDB22, LNW20, SOS^{+}21, WSB22]. Simulations [IKBL23]. since [HG20, MB20, PVA^{+}24]. Singapore [CSTR^{+}21, LNR^{+}21]. Single [LCMS^{+}22, ÁHGCVAI22, LLC^{+}20, MPP20, MGC^{+}22, RKN23]. single-area [MPP20]. Single-cluster [LCMS^{+}22]. sinking [AUM21]. site [FSP22, JAN23, MJD^{+}21]. sites [ASGG21, KNP^{+}20, MFO21, vBMP^{+}23]. situ [BMC20, DJFF23, PGD^{+}25, RGG22, RRB24, YLP^{+}23]. situations [CPM21]. six [AIM^{+}23, DSP^{+}23]. Size [BLHS20, KLS20, KHGB25, Lor22, MDS^{+}20, MOB^{+}23, PCM^{+}21, SXMV^{+}21, XSS^{+}23, AIJ^{+}23, AAH^{+}23, APB^{+}20, ANB^{+}24, AK23, BNTK23, BJHS^{+}23, BHVB^{+}24, BHH21, BKC21, BHG^{+}24b, BM23a, BM25b, BM25c, CC22, CKM^{+}20, CSSB22, CBHS24, CNDDAPMR21, CFP21, CFP22, DHB^{+}21, DAR^{+}23, EBGE21, ELM20, EWPB22, FG21, FWKR21, GRHHGM^{+}20, GPWP20, GKM^{+}23, GTS^{+}21, HMY25, HC22b, HGS^{+}23, HW24, HKCW24, IB20, JCL^{+}21, KAB^{+}22, KPK^{+}23, KSI^{+}22, LCM^{+}23, LCMS^{+}22, LCG^{+}21, LIA25, LCG22, LZX^{+}20, LVPS20, LCHB^{+}24, Lya20a, MINS21, MGB24, MSS^{+}21, MB25, MRUG^{+}23, MOTL25, MDC^{+}22a, Mur21, NSRM22, NAV^{+}23, OFF^{+}20, PZG^{+}20, PCBL23, SKST23, SK22, SHB^{+}23, Spa24, SBRM^{+}22, TTYT24, TBF^{+}21, TSPK24, TSI^{+}21, TPD20, VMFF^{+}20, WZX^{+}20, WPGO21, WCC24, dJDM23]. Size- [Lor22]. size-at-age [SBRM^{+}22]. Size-at-maturity [MDS^{+}20, LCM^{+}23, Mur21]. size-based [HC22b]. size-dependent [HGS^{+}23]. size-selection [OFF^{+}20]. Size-selective [XSS^{+}23]. size-selectivity [MB25]. Size-specific [SXMV^{+}21, NAV^{+}23, TSI^{+}21]. size-spectrum [dJDM23]. size-structured [CC22, CFP21, CFP22]. sized [JSKM20]. sizes [LTT^{+}23, TTL^{+}20]. sizing [KMO20, KBM23]. skate [GM21, HTK^{+}24]. skates [GMT24, TLV23]. skeletal [FGL22]. Skill [KSL^{+}23]. skipjack [ATA^{+}24, AUM21, Ash20a, Ash20b, HHD^{+}20, VP23, RCH^{+}21]. skipped [SKD^{+}20]. Slave [WBD^{+}21]. sloanii [LPP^{+}20]. slope [KRH^{+}24]. slots [BGCCP22]. Small [DPL20, FBM^{+}21, GTS^{+}21, PM23, XDX^{+}23, APB^{+}20, AOA^{+}22, COG22, CPF20, CDAK23, CNE^{+}22, CTCB22, DSP22, DSNK^{+}22, DMS22, ELM20, ECM^{+}24, FG23, FMMA20, FBW^{+}21, GdSPL21, HA23, IOO^{+}24, KM23, KSI^{+}22, KHS^{+}20, KCB^{+}24, KBB^{+}21, LTE^{+}23, LLFL21, LFdB^{+}21, MSLMOC^{+}24, NMS^{+}22, NB24, NFAL^{+}22, PAA^{+}24, PSS^{+}20, PJNGJ^{+}22, PBDM23, QAS^{+}25, RFMS^{+}21, RUHM20, RLO^{+}21, SHV20, SSM^{+}23, SBB^{+}24, SXK^{+}24, VTSI^{+}24, WGC^{+}19, WGC^{+}21, WKBMW24, YRTP20, YLS^{+}23, ZLXL20]. small-bodied [KCB^{+}24, LTE^{+}23, WGC^{+}19, WGC^{+}21]. Small-scale [FBM^{+}21, PM23, APB^{+}20, COG22, CPF20, CDAK23, DPL20, DSP22, DSNK^{+}22, DMS22, ELM20, ECM^{+}24, FBW^{+}21, HA23, IOO^{+}24, KM23, KHS^{+}20, NB24, NFAL^{+}22, PAA^{+}24, PJNGJ^{+}22, QAS^{+}25, RFMS^{+}21, RUHM20, SSM^{+}23, SBB^{+}24, WKBMW24, YRTP20, YLS^{+}23]. smaller [BHVB^{+}24, PZG^{+}20]. Smallmouth [OGFW24, CLD^{+}22]. smalltooth [PBB20]. SMART [BBC^{+}25, GLP^{+}20, LSM^{+}23, TKB^{+}21]. smartphone [SIM^{+}24a, SIM^{+}24b]. smartphones [JSKM20]. smolts [HPPT24, HUK^{+}23]. snagging [HVMEO25]. snail [AdlBCN25, KBH22]. snapper [AVMBEB22, ASJ^{+}20, BPT^{+}25, BHNP22, CPF20, CSB^{+}23, FGCB^{+}21, GPWP20, GAB^{+}22a, MFO21, MFR^{+}22b, PJSQ20, PJNGJ^{+}22, QMC^{+}22, SKW^{+}21, SFMA23, SACS23, TP24, VKS^{+}25, WCLN20, WWF^{+}20, WSB22]. SNE [BMM^{+}21]. SNE/MA [BMM^{+}21]. snorkel [SJW^{+}22]. Snow [CMRP20, HLG^{+}21, AIJ^{+}23, BMSM22, BWN^{+}21, FJS^{+}25, MCMM20, MBZSM20, MB20, MBP20, MBD^{+}21, Mur20, Mur21, NBD^{+}20, Szu22]. SNP [CPPK23, HIF^{+}24, PRA^{+}23, ZJY^{+}24]. SNP-based [ZJY^{+}24]. SNPs [BQGV^{+}24]. SO-YOLOv5 [XDX^{+}23]. soak [AAH^{+}23, BWN^{+}21, MBD^{+}21, NBD^{+}20, YHC^{+}24]. soaking [NLW^{+}22]. social [AWC^{+}23, ARD^{+}23, BHB^{+}22, CWM^{+}23, JvPOG25, RAE^{+}21, vdHBBR20]. social-ecological [ARD^{+}23, BHB^{+}22]. Socio [BSSE21, BSA^{+}23, NAS^{+}20, PSS^{+}20, PVC^{+}22]. Socio-economic [BSSE21, BSA^{+}23, NAS^{+}20, PSS^{+}20, PVC^{+}22]. socioeconomic [PRF^{+}21]. sockeye [BQBW20, EMR^{+}22, PPC^{+}23b, vP20, ACS23]. soft [GKC^{+}22, LCM^{+}23, LNMA23, LCW23b, PFGQ20, SPD^{+}24]. soft-bottom [PFGQ20]. soft-sediment [LCW23b]. soft-shell [LCM^{+}23]. software [BBG^{+}24, HIKM21, HMP^{+}22, TDJ^{+}21, VTS^{+}22]. solandri [ODM20]. sold [DCL^{+}20]. sole [CMD^{+}23, FMMC20, PIP^{+}22, Ten22, vORP23, FGPVPPGG22]. Solea [CMD^{+}23]. solida [dAGCR21]. solutions [IOO^{+}24, SMLT24, VR20].

Somatic [TSS^{+}23, FG21, TTYT24, WGFM21]. some [BGBM22]. Sonar [JGU21, LBLF20, GAB^{+}22a, HL21, KSS^{+}22, LJB^{+}24, PBB20, WLG^{+}23]. sonars [SEM^{+}23a]. Sonic [TSI^{+}21]. SORT [SZS^{+}24]. sorting [BLHS20, VMFF^{+}20]. Sound [SEM^{+}23a, ZCM^{+}23, LNvD^{+}23]. sounder [RGG22]. soup [LNR^{+}21, WSL^{+}24]. source [dS21]. Sources [Sat23, ASDW24, CHPT20, PPH21, DBL^{+}25, TDI^{+}21, TCL^{+}24]. South [ACP^{+}23, OGFW24, SZXC25, BM22, BM23b, BKM23a, BM24b, Bro25, BCRI21, CW22, GGMRC^{+}22, KBH22, MFM^{+}20, SJH^{+}23, AHX^{+}24, AdABW^{+}22, BSAP22, CZ25, CTR^{+}21, FCMP23, LZC^{+}21, LZCC24, Mac22, MCC^{+}23, MPSH25, MFM^{+}20, OCdMC24, QMC^{+}22, SS24, XWD^{+}21, ZZC^{+}21, ZZ22, ZCFG23, dSTMV20]. south-central [BCRI21]. south-eastern [BM22, BM23b, BKM23a, BM24b, Bro25, CW22, KBH22, SJH^{+}23]. south-western [GGMRC^{+}22]. Southeast [KMC20, BGG^{+}22, DSC24, HTSJ23, BPT^{+}20, KFDE^{+}22, QMGRIU22]. southeastern [BPT^{+}25, BM24a, BM25a, CGDTSA^{+}25, EYAO20, MdCG20, MFR^{+}22b, SSS^{+}23]. Southern [BMM^{+}24, BBPT^{+}22, BHG^{+}24a, LTR20a, LTR20b, MTE^{+}20, ALW^{+}21, ARD^{+}23, ASO^{+}22, ASJ^{+}22, BQGV^{+}24, BAF23, BaLK^{+}21, BHG^{+}24b, BRR^{+}21, BMA^{+}20, CWRR24, CVM^{+}20, CSRL20, DML^{+}20, DAR^{+}23, DMM^{+}23, DKBF23, EBN^{+}23, ECY21, FGSD25, FDS^{+}23, FSS^{+}24, FBR^{+}24, HRC23, JvPOG25, Mac22, MFM^{+}20, MRC24, MRUG^{+}23, NLW^{+}22, PCM^{+}21, PZG23, RKN23, RdLSdB^{+}21, RWT^{+}20, RDR^{+}23, SKBA23, SXK^{+}24, SMLT24, WBBG^{+}23, dS21, AMM^{+}22, BTB^{+}21, EWPB22, HSM^{+}25, KHE^{+}22, MSC^{+}24, MMML24, PDG^{+}22, SXM^{+}21, WCN^{+}24, WPGO21]. Southwest [IAB20, FCKG^{+}22, ABT^{+}24, dSCCC^{+}22, MOCGCC^{+}25, SSP^{+}22, SW24, dAdCdO^{+}23]. Southwestern [BHST^{+}21, CVM^{+}20, BÁP^{+}23, SDC^{+}22, AAR^{+}21]. sp [WWF^{+}20]. space [AT20, BP20, CAZN24, CWRR24, DAR^{+}23, HK22, LDM^{+}24, LBD23, LBD24, LKSiI22, PC21, SXM^{+}21, SM21, YIM^{+}20, ZHC^{+}25]. Spain [GMRRG20, dHJTCE23, HOQ20, MOCGCC^{+}25, OOAF^{+}21]. Spanish [DRSPTA20, DCL^{+}20, FGPVPPGG22, FHE^{+}24, PYX^{+}20, RRSP^{+}24, RLO^{+}21]. spanner [MWTH23, SSP^{+}23]. sparse [WS24]. sparse-spike [WS24]. Sparus [vBMP^{+}23]. Spatial [AMHR22, ASGG21, Ash20a, Ash20b, BHST^{+}21, BAYR^{+}24, CMP20, CGM^{+}22, CNDDAPMR21, DAL20, FSS^{+}24, FKW^{+}22, GMRRG20, GBC23a, KKCP20, LCD^{+}23, LZY24a, LJX^{+}20, MDW^{+}21, MDC^{+}22a, MYKO23, Mur20, NOL23, NFAL^{+}22, RCVGMI22, SSP^{+}23, Seu22, SW24, TNDM23, WPGO21, WSUN^{+}23, AVCA22, BVR^{+}21, BBG^{+}24, BHD^{+}23, Cad20, CGBJ23, CPL^{+}25, CMV21, CG21, DML^{+}20, DBGV^{+}22, DHCS23, HLMV24, HCDB22, IRJ^{+}22, JDH22, KSO^{+}21, KA24, LNW20, LK25, Mac25, MWJ^{+}24, MPP20, Mur21, NFdSJO25, OOAF^{+}21, PAA^{+}24, PCK23, PFdSBL25, SLA24, SBL^{+}23, SHS20, SR20, TSG25, TSI^{+}21, ZJJ^{+}25]. Spatial-temporal [FKW^{+}22, LZY24a]. spatially [CZJ^{+}24, GBC23a, JC21, MPP20, PDF20, TDI^{+}21]. spatially-explicit [MPP20]. spatially-replicated [JC21]. Spatio [DSB^{+}21, GGMRC^{+}22, HBC^{+}22, KNO^{+}21, MLCMdS23, MRG^{+}23, MPC^{+}20, CGT^{+}23, DBDT21, DTSR22, EYAO20, GWMC21, GML^{+}23, GMPD23, MTX^{+}20, NB24, PPM^{+}23, SPM^{+}24, SFPR^{+}23, TMP20, ZYZ^{+}23]. Spatio-temporal [DSB^{+}21, GGMRC^{+}22, HBC^{+}22, KNO^{+}21, MLCMdS23, MRG^{+}23, MPC^{+}20, CGT^{+}23, DBDT21, DTSR22, EYAO20, GWMC21, GML^{+}23, GMPD23, MTX^{+}20, NB24, PPM^{+}23, SPM^{+}24, SFPR^{+}23, TMP20, ZYZ^{+}23]. Spatiotemporal [BPT^{+}25, BAM^{+}24, CRS23, HRC23, NNS^{+}22, SF22a, SMLT24, BCPH22, BAF23, CBA^{+}24, DBS^{+}21, FTH^{+}23, GBB20, HM25, HCK^{+}21, SXMV^{+}21, SFYM24, WCN^{+}24]. spatiotemporally [KMA23]. spawner [KSI20b, SCG^{+}24]. Spawning [CBHS24, MFO21, SSFL24, ASGG21, AIW^{+}23, BGK22, BHST^{+}21, CPB^{+}21, CKDP^{+}20, EHE^{+}23, ECY21, FSP22, FSS^{+}24, GSS^{+}23, JAN23, KNO^{+}21, KSI20a, KHK^{+}20, LCB^{+}21a, LCD^{+}23, LCM^{+}23, LRW^{+}24, MSS^{+}21, MPM^{+}23, MFR^{+}22b, OBB^{+}20, PZL^{+}23, PPC^{+}23b, RFMS^{+}21, RBG^{+}20, SKD^{+}20, SSJ^{+}21, TSPK24, TSI^{+}21, VCG^{+}23, VAVQGD^{+}20, WSUN^{+}23]. spearfisher [BLC^{+}21, FCMP23]. spearfishing [HSW25, SRP^{+}22]. Special [Mel23, MPSM25, NOL23]. specialised [BAYR^{+}24]. Specialist [FCMP23, FBR^{+}24]. Species [MDC^{+}22b, PKRL21, SPM^{+}24, WXJ^{+}24, AdABW^{+}22, AIM^{+}23, AL22, ASD^{+}22, BGG^{+}22, BGM^{+}23, BKM23b, CMV21, CGT^{+}23, CGB^{+}22, CHL^{+}20, CHAY^{+}25, DOB^{+}24, Dik24, DQMV21, ECK^{+}21, ECM^{+}24, FOM21, FM21, dSRFFN^{+}20, FCMP23, FKS^{+}20, FAK24, GGG^{+}22, GM21, GMT24, GFDN^{+}22, GMPD23, HIF^{+}24, HP23, HZZ^{+}20, HMS^{+}22, HCDF24, JDH22, JC21, JRW^{+}21, KNO^{+}21, KHE^{+}22, KFH^{+}25, KMC20, LPAE^{+}24, LG21, LYH^{+}21, LJX^{+}20, LPRB^{+}21, LNR^{+}21, LMG^{+}24, LZX^{+}20, Lya20b, MINS21, MRP^{+}23, MPP22, MCC20, NLW^{+}22, NCB^{+}23, NVRG^{+}21, OPL21, OKKW20, OAM^{+}21, PCK23, PM23, PPD^{+}25, PPM^{+}23, PRN^{+}24, PBM^{+}23, RLQ^{+}20, RLO^{+}21, SSSF25, SSRC24, SKY^{+}24, SCGW24, SWH^{+}24a, SECB21, SYZ^{+}25, SDdMG^{+}20, SJ24, SAA23, TFC^{+}20, TRS^{+}24, WQGS25, WWT^{+}25, WJN^{+}25, XCB^{+}21, YHST22, YAO^{+}23, YiTM23, YWC^{+}21, ZYZ^{+}23, ZOS^{+}23]. species-specific [BKM23b, FAK24, HCDF24, KFH^{+}25, LG21, YiTM23]. specific [BKM23b, CBTH20, DEM^{+}23, DJFU20, FAK24, HCDF24, KFH^{+}25, LG21, LYH^{+}21, MCS^{+}22, NAV^{+}23, NVB^{+}23, NRH^{+}23, RWT^{+}20, SXMV^{+}21, TSI^{+}21, WCLN20, YiTM23]. specification [ATAS20, TMH23]. specifying [HPD^{+}22, PJMP22]. spectra [LIA25]. spectroscopy [BHB24, DOB^{+}24, PJSQ20]. spectrum [dJDM23]. Spencer [MPSH25]. sperm [BMSM22, SSFL24]. spheres [KLNB^{+}24]. Sphyrna [BFA^{+}21, GAW^{+}22]. spider [SDdMG^{+}20]. spike [WS24]. spillover [SCGM^{+}21]. spine [AMdC^{+}20, CPR^{+}24]. spined [GBB20, MHB^{+}23]. spines [PRRR23]. Spiny [SRF^{+}24, BF25, BBR^{+}22, BBSM24, CHB24, GAB^{+}24, HM25, HMR^{+}24b, HMR24a, LMJ^{+}23, LMM^{+}24, Mac25, MCHA21, PBD25, RBHM24a, TdL24, YFJ^{+}25, dLH23]. Spisula [dAGCR21]. split [BTB^{+}21]. sport [ASB^{+}24]. spotted [BW20, GKM^{+}23, KNO^{+}21, TLV23]. spp [BW23, BM23a, BM25b, BM25c, dlTGPLC23, YSB^{+}21]. SPR [HFKS20]. sprat [PGD^{+}25, SBC^{+}22]. Sprattus [PDC^{+}23, SBC^{+}22]. spread [BWR24, JMP^{+}21]. Spreading [WGNM24]. spring [Sim23]. Squalus [HM25, YHST22]. square [BJHS^{+}23, CAH^{+}20, VS23]. square-meshes [BJHS^{+}23]. squat [CQA^{+}24, FBQA20]. squid [Bro25, BCRI21, HZZ^{+}20, KM23, LCB^{+}21a, LPP^{+}20, LXC^{+}20, MdCG20, QMGRIU22, SSV^{+}20, WHCF22, YTHM20, YTH22, YWC^{+}21, ZHC^{+}25]. squids [BCPH22, JWL^{+}24]. Sr [ZOS^{+}23, ZOS^{+}23]. Sr/ [ZOS^{+}23]. Sr/Ca [ZOS^{+}23]. Sri [ABK^{+}21, PKRL21, PDJ24, RKD24]. St [CRS23, GLA21, PCM^{+}21, SMLT24]. stable [AAFLL^{+}25, BÁP^{+}23, BRGB^{+}23, CZ25, DSNK^{+}22, HTK^{+}24, LCFJ22, LXC^{+}20, SIMT24, WHCF22, WWT^{+}25, YLX^{+}24]. stacked [SJ24]. stage [İTAD24, IRJ^{+}22, KK22, Mun24, OMG^{+}23, WHCF22]. stages [BMA^{+}20, CGT^{+}23, ECK^{+}21, GQP^{+}25, MDL^{+}21, SF22b, TLCD21, TLC^{+}22, ZYZ^{+}23]. stained [CHL^{+}20]. stakeholders [EBN^{+}23, HFMH20]. standard [JBČ^{+}22, RGN^{+}20, SFC21, TRS^{+}24]. Standardising [PMC^{+}24]. standardization [DKD^{+}21, DBGV^{+}22, HCDB22, LZY24b, LZY24a, TCJ^{+}21, XMLCMV24]. Standardized [SKST23, VK21a, VK21b]. standardizing [MKS^{+}21a]. standards [GLA^{+}20, HMC^{+}23]. Startle [WOG^{+}25]. starvation [CWRR24, RBD^{+}22, WCN^{+}24]. State [IOO^{+}24, NFAL^{+}22, AT20, BP20, CAZN24, CWRR24, HK22, LDM^{+}24, LBD23, LBD24, LKSiI22, MFO21, PC21, SXM^{+}21, SM21, YIM^{+}20]. state-space [AT20, BP20, CAZN24, CWRR24, HK22, LDM^{+}24, LBD23, LBD24, LKSiI22, PC21, SXM^{+}21, SM21, YIM^{+}20]. stated [BKM^{+}23c]. States [BGG^{+}22, BPT^{+}25, BPT^{+}20, DSC24]. static [BBPT^{+}24, Fre22, PVA^{+}24, SMKJ21]. stationary [BGG^{+}22, CG21, RWT^{+}20]. stations [GRJW20]. statistical [BCG^{+}25, CHGC25, LGD^{+}20, Mon24, PTL^{+}24, SSKS21, VP23]. statistics [SRB^{+}25, Tho19, Tho21]. status [CPM21, CBHS24, CRCAF^{+}22, DMF^{+}21, DWS^{+}23, DD25, FJA^{+}20, GKC21, HEG^{+}23a, LWH^{+}23a, MOI23, NA22, PKRL21, RGP^{+}23, SK21a, SCD^{+}22, WZL^{+}22]. stays [VP22]. steelhead [CBR^{+}23]. Steep [HW21]. Steindachner [RCVGMI22]. Stenodus [WBD^{+}21]. stenolepis [JLYR24, LBW21, PRK23, TRN^{+}23]. Stenotomus [CG21]. stephani [KZT^{+}23]. Stephens [Det21a]. steps [GTS^{+}21]. stereo [BWG^{+}21, CSSB22, RWT^{+}20, WSB22]. stereo-baited [CSSB22]. stereo-camera [BWG^{+}21]. stereo-video [WSB22]. Sthenoteuthis [LZCC24]. Stichopus [GBC^{+}23b]. stickleback [MHB^{+}23]. sticklebacks [GBB20]. still [RWT^{+}20]. Stimulating [CAH^{+}20]. stimulation [RBM21, vORP23]. Stizostedion [TSS^{+}23]. stochastic [Che21]. Stock [CRCAF^{+}22, HWW22, LCG^{+}21, LNvD^{+}23, Mau22, MPSM25, MNS^{+}20, Pun24, SSP^{+}22, AT20, AAT^{+}21, Alv21, AGS20, AGB^{+}24, AMdC^{+}20, Bea21, BJS^{+}22a, BJS^{+}22b, BAW^{+}24, BBG^{+}24, BBPT^{+}24, BDM^{+}20, BRR^{+}21, Bro24, CAZN24, CWRR24, Cad20, CMP20, CGBJ23, CWC^{+}21, CTIC23, CGC24, CHGC25, Cla22, Cop24, dSCCC^{+}22, CBHS24, DDDP21, DMF^{+}21, DBV22, DBV23, EPHDB24, aFLpX^{+}21, FCSA21, Fis25, FDS^{+}23, FJA^{+}20, FHHH20, GVK^{+}23, GBC23a, GML^{+}23, GBB20, GFM^{+}23, HCCC21, HIMP23, HC22b, HIKM21, HL20, HMP^{+}22, HWMVM23, HCDB^{+}24, HW24, JC21, JBČ^{+}22, KTFY22, KNO^{+}21, KDBOC25, KWC^{+}20, KPS20, KPWS21, KHK^{+}20, KHPB20, LMP24, LWH^{+}23a, LDM^{+}24, LBD23, LBD24, LKSiI22, LZW^{+}21, LDC24, MCS^{+}22, MJD^{+}21, MMG^{+}24, MTX^{+}20, MHL^{+}23a, MJC^{+}23, MPH21, MUF^{+}22, MVMdS^{+}24, MOI23, Mon24, MBE^{+}20, MAA^{+}20, MZSZVP^{+}23, MNPMM^{+}22, NVRG^{+}21]. stock [NDRR20, OBB^{+}20, ÖSL^{+}23, ODM20, PC21, PBM^{+}23, PJP20, PJMP22, PDE^{+}20, PTD^{+}20, PCJH^{+}21, PTD^{+}21, Pun23, PTL^{+}24, PCGG20, RRH^{+}24, SH22a, Sat23, SCN^{+}24, SMK^{+}24, SDC^{+}22, SBZ^{+}21, TAA^{+}20, TDJ^{+}21, TNDM23, TSG25, Tho19, Tho21, TAK^{+}23, WH23, WWO20, WS24, XMCC20, vPDD^{+}25, vdHR23]. stock-increase [KTFY22]. Stock-recruitment [Mau22, MNS^{+}20]. stock-selectivity [FHHH20]. stocked [EVS^{+}23, GW21, JAN23, PW25, SSD^{+}20b]. Stockholm [HBW21, SAA23]. stocking [HGS^{+}23, KK22, KWMA23, Lya20b, SPE^{+}23, ZSWK25]. stocks [BTC23, BAJB^{+}24, BCRI21, CAZN24, CAAFH21, DQK^{+}23, DWS^{+}23, DHCS23, GQP^{+}25, GM21, HFKS20, HWMVM23, HLI^{+}20, KPS20, LADA^{+}22, MMBH23, PDG^{+}22, PdAMdM^{+}23, PDF20, PDD^{+}22, Pun24, QMC^{+}22, RNKB23, SCW21]. Stomach [AGL^{+}24, BÁP^{+}23, WWT^{+}25]. stomachs [SAA^{+}22]. stomias [Dra22]. stone [FJA^{+}20, KAC^{+}23]. storage [DH20, HKG^{+}21, LTR20b]. storm [ODM20]. Strain [JAN23]. Strain-dependent [JAN23]. Strait [BaLK^{+}21, GQP^{+}25, NY23]. Straits [ŞGK^{+}20, WY20]. Strangomera [MVLC^{+}20]. Strategies [WSL^{+}24, BDR^{+}20, BGL^{+}22, HHD^{+}20, KKC25, LDC24, LAG^{+}21, MCS^{+}24, MA20, MBP20, OUB^{+}22, PSS^{+}20, PvZ22, PBDM21, PTL^{+}24, RMNB^{+}21, RBG^{+}24, VBL^{+}22, VBL^{+}24]. Strategy [HF20, BBPT^{+}24, DSC24, KK22, LMJ^{+}23, LMM^{+}24, PMS^{+}20, PBDM23, PDD^{+}22]. stratification [CKK^{+}20, GC25, DBL^{+}25]. stratified [HKG^{+}21]. stream [BTB^{+}21, BKC21, CGT^{+}23, LR21, YTSS22]. streamlining [WQGS25]. streams [JCL^{+}21, LHPR21, Lya22, PBB^{+}22]. strength [BUG^{+}24, DJFF23, LPAE^{+}24, LADA^{+}22, LIA25, PFC^{+}21, PDC^{+}23, SGD^{+}21, WYM^{+}25, WKSF20, YLP^{+}23]. Stress [FGL22, AGNS^{+}21, BALBC23, CCC^{+}22b, GMK23, KFDE^{+}22, RS21]. stressed [LADA^{+}22]. stresses [HSJ^{+}24]. stressor [FJS^{+}25]. stressors [NVRG^{+}21]. striata [CG21]. strictly [YTSS22]. striking [CPB^{+}21]. striped [BAM^{+}24, GSH22, HWW22, KSI^{+}22, MPM22, ÖA21, PHV^{+}21, WMT^{+}20]. strong [CHB24, Mau22]. Stronger [KTFY22]. Strontium [ACP^{+}23, SPW^{+}22, TLCD21]. Structural [FGCB^{+}21]. Structure [CMP20, GGMMMV^{+}20, AHX^{+}24, APBN22, AGB^{+}24, AGNR^{+}21, AAR^{+}21, ÁHGCVAI22, BHH21, BJK24, Cad20, CGBJ23, CKM^{+}20, CGC24, CPPK23, CMTP^{+}21, CSRL20, DSNK^{+}22, FTB^{+}21, GMRRG20, GFM^{+}23, HSM^{+}25, HQWD20, dlTGPLC23, HLI^{+}20, íKMíH^{+}22, LVP22, LLC^{+}20, MJD^{+}21, MVMdS^{+}24, MNPMM^{+}22, MMM^{+}20, NTJN21, NSM^{+}21, NMS^{+}22, NVRG^{+}21, NDRR20, ÖA21, PBM^{+}23, PCGG20, RKN23, SBC^{+}22, SCN^{+}24, SSP^{+}22, SNJ^{+}24, SDC^{+}22, Spa24, TSG25, TBF^{+}21, TSPK24, WBBG^{+}23, WASS20, WPB^{+}20, WWO20, WMT^{+}20, WMSW22, WCC24, YGMJ20, ZOZW22]. Structured [LZY24b, CC22, CMV21, CFP21, CFP22, GBC23a, MZZ^{+}21, Mun24, PJMP22]. structures [KFO20, MLCMdS23, MBE^{+}20, MAA^{+}20]. structuring [LNvD^{+}23]. Struggling [BRN^{+}20]. studied [GFDN^{+}22]. studies [CAZN24, HHD^{+}20, LHH^{+}25, LPRB^{+}21, MB25, NS23, SF22b, ZZC^{+}21, dSTMV20]. Study [ZOZW22, AVMBEB22, ACP^{+}23, BBPT^{+}24, BHB^{+}22, BSAP22, BKM^{+}23c, BCM^{+}21, CFB^{+}23, CHL^{+}20, DBGV^{+}22, EBN^{+}23, EBdRPC24, FCKG^{+}22, FDdCS^{+}20, GCK^{+}21, GPW^{+}20, GM21, GP21, GPT^{+}21, HSSD^{+}21, HA23, HCDB22, Hut22, HSW25, KC22, KFO20, LZY24b, LNW20, LKSiI22, MdCG20, MPP20, MBZSM20, OMK24, PIP^{+}22, PHH^{+}23, PSS^{+}21, RKD24, RRSP^{+}24, SSP^{+}23, SFMA23, SPC^{+}25, SKBA23, SMA^{+}24, SMKJ21, SH22b, Szu22, TMN^{+}21, TRWH23, TLV23, VTS^{+}22, WCN^{+}24, WPB^{+}20, WSB22, ZCM^{+}23]. stunts [TdL24]. Sturgeon [BH23, WPB^{+}20]. sturgeons [BTB^{+}21]. style [SCS25]. stylifera [CKD^{+}21]. Sub [AGS20, LNvD^{+}23, vZvdHCA25]. Sub-annual [AGS20]. sub-population [vZvdHCA25]. sub-stock [LNvD^{+}23]. subadult [CHB24]. subarctic [BAM^{+}24]. subcontinent [PVC^{+}22]. subject [TSPK24]. sublegal [ZZH^{+}24]. submerged [NHE^{+}23]. submersible [BWG^{+}21]. subpopulational [dAdCdO^{+}23]. subpopulations [AMSC20]. subsampling [SJBT20]. subscribed [SHC21]. subsequent [LR21, RS21]. subsidies [PFFdC22, VKS^{+}25]. subsistence [LFdB^{+}21]. substantial [CPPK23, FGL22]. substantially [SEM^{+}23b]. Substitution [SCHSC21]. subtidal [BWG^{+}21]. Subtle [SEM^{+}23b, PBG24]. subtropical [Ash20a, Ash20b, ECK^{+}21, FDB^{+}20, VR20]. subtropical-tropical [ECK^{+}21]. success [JHB21, MRUG^{+}23, PZL^{+}23, PPC^{+}23b]. suckleyi [YHST22]. Sudanese [OAM^{+}21]. suggest [BALBC23, MPC^{+}20, PCBL23, WKSF20]. suggests [BKHA21, DWS^{+}23]. Suitability [PRN^{+}24, HLZ^{+}20, RBS^{+}24, dLH23]. suitable [HMC^{+}23, RBH^{+}24]. summarization [GRJW20]. summary [HCDB^{+}24]. summer [EFM25, LM22, MLS^{+}21, RMRG22, WMSW22]. sunfish [KRH^{+}24]. sunlight [DH20]. sunset [KHE^{+}22]. superba [ZZ22]. superciliosus [AWVS21]. Superior [HJMS20, SBBH24]. supermarkets [DCL^{+}20]. supervised [SPD^{+}24]. supplementary [DML^{+}20]. Supplemented [DBDT21]. supply [LPS^{+}25]. support [BMJ^{+}24, Cop24, CBD^{+}22, LZY^{+}24c, PDM^{+}24, PBD25, TDJ^{+}21, WKBMW24]. Supporting [GML^{+}23, MSD21]. suppressing [SOSK22]. sure [DDA^{+}20]. surface [AUM21, CB20, ERS^{+}23, LD25, WSUN^{+}23]. surface-based [CB20]. surface-type [AUM21]. surgical [TRS^{+}24]. Surigao [CRSC22]. surplus [BP20, CAAFH21, GPW^{+}20, JC21, KBK^{+}24, WCT^{+}20]. surprising [TCJ^{+}21]. Survey [GVK^{+}23, HKKa^{+}25, HH20, LFdB^{+}21, AMM^{+}22, BWOR23, CYBW22, DCK^{+}22, DBDT21, DTSR22, DKD^{+}21, FGTA24, FAK24, GWMC21, HLG^{+}21, HdLHD22, HJA^{+}21, LVL25, LRMH21, LNW20, MRP^{+}23, OWF^{+}23, PBB20, PMC^{+}24, RMD^{+}25, SSKS21, SR20, SGD^{+}21, TSC^{+}20, TLAM25, TCL^{+}24, VP22, WASS20, WSL21, YTSS22, NMJ^{+}24]. Survey-based [HH20, LVL25]. surveying [MCMM20, MMBH23]. surveys [BLFT23, CZJ^{+}24, DGMG^{+}22, DLP^{+}24, ERS^{+}23, HM25, JM23, JRW^{+}21, LBLF20, OCBJG20, OAM^{+}21, PF20, SJW^{+}22, TC24, TGG^{+}24, TSC^{+}22, TMH23, dLHER24]. survival [CHM24, CBR^{+}23, EVS^{+}23, EMJ^{+}22, EM23, EFM25, GKC21, GDVBB^{+}20, GW21, GWGM24, LYH^{+}21, MSW21, Mau22, OOM^{+}23, SFCG^{+}21, SBD^{+}22, Sim23, SACS23, SXM^{+}21, SBT^{+}20, TPW23, TJS23, UAB^{+}21, YIM^{+}20, dLHER24]. survivors [UAB^{+}21]. Survivorship [BMM^{+}24, AWVS21, OOM^{+}23, SAW^{+}20]. susceptibility [CCR24, FM21, PJNGJ^{+}22, PDJ24]. Sustainability [EBdRPC24, BDR^{+}20, BMA^{+}24, FDdCS^{+}20, HLMV24, LLFL21, PK24, WKBMW24, dCHdMS^{+}23]. Sustainable [MZSZVP^{+}23, AAM^{+}20, AVB^{+}23, DMNV^{+}24, HFMH20, IOO^{+}24, KSV^{+}22, LS24, LZY^{+}24c, MGB24, MSD21, MOM^{+}25, Mun24, NFdSJO25, RPH20, RUHM20, SKS^{+}23, TP24, VTSI^{+}24]. sustaining [AdlBCN25, KMC^{+}23, PNRS23]. SW [HCCC21, SHV20, SFMA23]. swamp [NK23]. Swedish [MVDH24, ÖSL^{+}23, vdHBBR20]. swept [BWR24]. swimbladder [LPAE^{+}24]. swimmer [BJK24, JYH21, MHD^{+}21, PBM^{+}23]. swimmers [GGL^{+}24]. swimming [BGL^{+}22, CTS^{+}23, GZLCRG25, aLBK^{+}21, LZW^{+}21, LP23, SSG^{+}22, TNTN20, WKSF20, WGC^{+}19, WGC^{+}21, YHC^{+}24, ZZH^{+}24]. switching [HTK^{+}24]. swordfish [CHT20, FCKG^{+}22, MPM^{+}23, SB20, SAHW22, TPW23]. swordtip [YTHM20, YTH22]. sympatric [FHSC21, RdBAT^{+}23, WWT^{+}25, YiTM23]. synagris [CPF20, NVRG^{+}21, QMC^{+}22, SFMA23]. Synbranchiformes [TTC^{+}25]. synthesis [ARD^{+}23, TMH23]. Synthesizing [BBG^{+}24, RDR^{+}23]. Synthetic [CCKL^{+}20, DMNV^{+}24]. System [GGMMMV^{+}20, AFR^{+}24, AAZ20, FG23, GPASM22, HHJ^{+}22, HZZ^{+}20, KSV^{+}22, SOSK22, TLAM25, DSB^{+}21, GBO^{+}20, MOB^{+}23, ŞGK^{+}20]. systematic [LCMS^{+}22, PVC^{+}22, SB24, YAO^{+}23]. systems [BH23, CRL21, LJB^{+}24, OMG^{+}23, RBH^{+}24, dAdCdO^{+}23].

T0 [BM24a]. T45 [BM24a]. T90 [BW23, BM24a, JBL^{+}22]. T90-codend [JBL^{+}22]. tackle [CBR^{+}23]. tactics [SGZH^{+}22, vP20]. Tag [GGG^{+}22, AAG22, BNTK23, CLY^{+}22, DMM^{+}23, FFG21, NGDC25, RNP^{+}24, SSY20, SKW^{+}21, SBZ^{+}21, VP23, vBMP^{+}23]. tag-based [SBZ^{+}21]. tag-integrated [VP23]. tag-recapture [DMM^{+}23, SKW^{+}21]. tag-recaptures [CLY^{+}22]. tag-reporting [AAG22]. Tag-shedding [GGG^{+}22]. tagged [AGL^{+}24, TBÓ^{+}22]. Tagging [Mel23, RBG^{+}20, Bea21, BJS^{+}22a, BJS^{+}22b, BKT^{+}20, BMC20, BHD^{+}23, DAD^{+}22, GGG^{+}22, HMM^{+}21, HMS^{+}22, KHPB20, MCC^{+}22, MHH^{+}20, PW25, PRN^{+}24, RBH^{+}24, SSY20, SPG^{+}21, SDV^{+}22, SHS20, TSI^{+}21, VBB20]. tags [CLB^{+}21, HKG^{+}21, MMG^{+}24, MCC^{+}22, NEBP^{+}23, OOM^{+}23, SAA^{+}22, SW20, SV23, WGC^{+}19, WGC^{+}21]. Taiwan [BaLK^{+}21, SLW^{+}20]. take [KMC^{+}23, SCGM^{+}21]. Taking [JWL^{+}24, NMJ^{+}24, MRG^{+}23]. tale [GMT24, KHE^{+}22]. tales [UAB^{+}21]. Tana [HKKa^{+}25]. tank [HVMEO25, SDBS21, YLP^{+}23]. Tanner [FJS^{+}25, HSSD^{+}21, Mur20, Mur21, NGDC25]. Tanzania [CMBL21, HMY25]. Tarek [SPW^{+}22]. Target [BUG^{+}24, BC20, DJFF23, GM21, KLNB^{+}24, KFH^{+}25, LPAE^{+}24, LSM^{+}23, LIA25, MCC20, NLL^{+}25, PFC^{+}21, PDC^{+}23, PPM^{+}23, Peñ21, SSRC24, WYM^{+}25, WKSF20, YLP^{+}23]. Targeted [WOG^{+}25, BS20a, GM21, OOAF^{+}21]. targeting [BM22, KHGB25, LG21, ÖSL^{+}23, PCF23, RLO^{+}21, TSC^{+}20]. targets [KMO20, NLW^{+}22, SZS^{+}24]. tarichi [SPW^{+}22]. tarpon [ECK^{+}21]. Tasmanian [BHG^{+}24b]. taty [HSJ^{+}24]. Tawi [MTE^{+}20]. Tawi-Tawi [MTE^{+}20]. taxa [CFB^{+}23]. taxonomy [XWD^{+}21]. Tchang [ZOZW22]. teams [FGTA24]. technical [KM23, NSQV22]. techniques [BHD^{+}23, GW21, PGD^{+}25]. technological [DGMG^{+}22, NSQV22]. Technology [WOG^{+}25, BS20b, FBW^{+}21, MBB^{+}23, SBL^{+}23]. TEDs [CTM^{+}20]. Tegulidae [BQGV^{+}24]. Tehuelche [MSV21, TLAM25]. tehuelchus [MSV21, TLAM25]. telemetry [CBD^{+}22, JND^{+}23, KWE^{+}21]. teleost [EPHDB24, RWB^{+}23]. Teleostei [HCCC21, KFO20]. tell [UAB^{+}21]. temperate [Ash20a, Ash20b, BLFT23, CW22, PRCF22, TPW23, XCB^{+}21]. Temperature [MDL^{+}21, TLCD21, BKC21, DLZ^{+}25, EHBJ20, FG21, GLA21, ORdlG^{+}24, RPL^{+}24, TTYT24, TSS^{+}23, WST^{+}23, dAGCR21]. temperatures [JZQZ20]. Temporal [ABK^{+}21, CH21b, GKC^{+}22, Mur21, SWLH20, ASGG21, Ash20a, Ash20b, BHST^{+}21, CMBL21, CGT^{+}23, DBDT21, DTSR22, DSB^{+}21, EYAO20, FJHT^{+}22, FSS^{+}24, FKW^{+}22, GGMRC^{+}22, GWMC21, GML^{+}23, GMPD23, HBC^{+}22, HLMV24, HCDF24, KNO^{+}21, KKCP20, LCD^{+}23, LZY24a, MLCMdS23, MTX^{+}20, MRG^{+}23, MDC^{+}22a, MPC^{+}20, MOB^{+}22, NB24, NFAL^{+}22, OAB^{+}23, PPM^{+}23, RCVGMI22, SSP^{+}23, SPM^{+}24, SR20, SFPR^{+}23, TMP20, ZYZ^{+}23, ZJJ^{+}25]. Temporally [DSNK^{+}22]. temporary [WKBMW24]. Teno [LVP22]. term [BDA^{+}24, BLE^{+}22, BBPT^{+}24, BKR^{+}22, BBSM24, Cam23, CGB^{+}22, CHL^{+}20, DBS^{+}21, DMS22, EMJ^{+}22, EM23, EFM25, GKC^{+}22, GFC^{+}22, HMM^{+}21, HCCC21, LCB^{+}21b, MVLC^{+}20, NVB^{+}23, PW25, PDM^{+}24, PMC^{+}24, SJH^{+}23, SW20, TRS^{+}24, VAVQGD^{+}20]. terminal [CBR^{+}23]. Territorial [GHAZ21]. Territories [WBD^{+}21]. Territory [KCB^{+}24]. Test [SOSK22, OYOO21]. tested [ERS^{+}23]. Testing [BDA^{+}24, QAS^{+}25, SSJ^{+}21, VMFF^{+}20, DMNV^{+}24, HLG^{+}21, RB22, SDBS21]. tether [WYM^{+}25]. tethered [KBM23]. Texas [BAF23]. Thailand [CDAK23, HSM^{+}25, KPS20, KPWS21, SPC22]. theft [HBE^{+}22]. their [ALW^{+}21, BTB^{+}21, BBJ22, CCGR20, CCKL^{+}20, DD25, FCML^{+}22, FRP22, KSS^{+}22, LCB^{+}21b, LR21, MBOCdAM24, MHL^{+}23a, MPSM25, MABR^{+}20, OCBJG20, PVA^{+}24, RNKB23, DBL^{+}25, Spa24, SCC^{+}22, YTSS22, Žák21]. them [Pun23]. Theory [CMP20, HIMP23, Cla22]. therapeutic [NK23]. there [AUHK22, FDdCS^{+}20, SHB^{+}23]. Thermaikos [GCK^{+}21]. Thermal [EMR^{+}22, JHB21]. thermoregulation [JGG^{+}24]. thicknesses [TTL^{+}20]. Thin [CSRL20, FWKR21]. Thin-sectioned [CSRL20]. think [KWMA23]. thornback [TLV23]. Those [Pun23, NVB^{+}23]. threadfin [BCM^{+}21, SLW^{+}20]. threatened [APBN22, BC20, HS21, RWB^{+}23, SR21]. threatens [BQBW20]. threats [VR20, dAdSR^{+}20]. three [ACL^{+}20, CGB^{+}22, CDAK23, CNE^{+}22, ECY21, GBB20, LOFS22, LCFJ22, Lya20b, MHB^{+}23, OAM^{+}21, RdBAT^{+}23, RUHM20, TYYK21, ZHC^{+}25]. three-dimensional [ZHC^{+}25]. three-spined [GBB20, MHB^{+}23]. thresher [AWVS21, MRC24]. thresholds [RBS^{+}24]. thrissa [XqRJ^{+}23]. through-the-gill [CLD^{+}22]. throughout [CKM^{+}20]. Thunnus [AAFLL^{+}25, ASO^{+}22, AIW^{+}23, BUG^{+}24, FFG^{+}20, HBC^{+}22, LK25, NBF20, ORdlG^{+}24, PBPM^{+}23, PZG^{+}20, SXMV^{+}21, SF22a, SAW^{+}20, SBRM^{+}22, SWH24b, WAA^{+}22, ZZC^{+}21]. thynnus [BUG^{+}24, HBC^{+}22, NBF20, ORdlG^{+}24, SBRM^{+}22, SWH24b]. thyrsites [GKC^{+}22]. Thysanoteuthidae [SSV^{+}20]. Thysanoteuthis [SSV^{+}20]. Tibetan [DHC^{+}20]. tickler [BRvL^{+}22]. tickler-chain [BRvL^{+}22]. tide [SH22a]. tiger [ASB^{+}24, AGL^{+}24, BCSM20]. tilt [WKSF20]. Time [CWRR24, DB22, FNKY20, SSSF25, AHL20, AAH^{+}23, AK23, BWN^{+}21, BCG^{+}25, CC22, CFP21, FG23, GJSW22, JSG21, LCM^{+}23, LTE^{+}23, LD25, MAH^{+}22, NLW^{+}22, OS21a, OS21b, RB22, SM21, Szu22, YHC^{+}24]. time- [SM21]. time-dependent [JSG21]. time-variation [Szu22]. Time-varying [DB22, FNKY20, CC22, CFP21]. Times [PVPN22, MBD^{+}21, NBD^{+}20, SCD^{+}22]. timing [BAF23, LDC24, SBJ^{+}20]. tip [AHL20, SCS25]. tip-ups [AHL20]. tipping [DQK^{+}23]. tissue [HTK^{+}24, WWT^{+}25]. tissues [BRGB^{+}23, LCFJ22, LPP^{+}20]. Todos [MPEBdR23]. Toki [TLAM25]. tolerance [SSSF25]. tonnes [MB20]. tool [BMA^{+}24, DML^{+}20, DMM^{+}21, DMS22, JGU21, LvCdGSL23, MVMdS^{+}24, MPP22, RTHB25, WSL21, WJN^{+}25, ZOS^{+}23, dAdCdO^{+}23]. tools [BRR^{+}21, CCC^{+}21, DOB^{+}24, LDM^{+}24, SLA24, WSL^{+}24]. Toothed [BTR^{+}24, MRE^{+}24]. toothfish [AHX^{+}24, CPB^{+}21, GMPD23, MPP20, ZSDZ24]. top [BQGV^{+}24, BS20a, Bro25]. top- [Bro25]. top-panel [BS20a]. Tor [BSR^{+}22]. torquata [KBH22]. tougher [LLH^{+}25]. tourism [LVP22]. tournament [MSW21]. tow [BWR24, CYBW22, KA24]. towed [HVMEO25]. toxotes [GZLCRG25]. trace [BRR^{+}21, LPP^{+}20, TBÓ^{+}22]. traceability [RSD23]. Trachurus [DCL^{+}20, HIF^{+}24, MPC^{+}20, MPV^{+}24, MYKO23, PFC^{+}21, RCVGMI22]. track [BHD^{+}23, JSG21]. tracking [FBW^{+}21, HBE^{+}22]. Trade [AYTM21, HS21, HLMV24, BGCCP22, KDBOC25, MCH21, NAS^{+}20, NLL^{+}25, PBDM23, PBD25, SHC21]. Trade-off [AYTM21]. Trade-offs [HLMV24, BGCCP22, KDBOC25, NAS^{+}20, PBDM23, SHC21]. traded [dSRFFN^{+}20, LNR^{+}21]. tradeoffs [DSC24, VA20]. traditional [AVCA22, BMC20, BRN^{+}20, CBTH20, FG23, GZLCRG25, HPL^{+}24, NLL^{+}25, SFC21, TBF^{+}21, VCPO21]. trained [PNGGO^{+}22]. trait [FBPCC^{+}21]. trait-selective [FBPCC^{+}21]. traits [AUHK22, Ash20a, Ash20b, BLC^{+}21, LZX^{+}20, MDS^{+}20, MPM^{+}23, MVLC^{+}20, MOM^{+}25, PDG^{+}22, PSSFS24, RdBAT^{+}23, DBL^{+}25, TLV23, TEO25, XSS^{+}23]. trajectories [IKBL23, PVA^{+}24, SWH24b]. trammel [EMJ^{+}22, FMMC20, GCK^{+}21, GMK23, GP25, LVPS20, RMD^{+}25, SME^{+}24]. TRAMORE [SSY20]. trans [SSY20, YWC^{+}21]. trans-molting [SSY20]. Trans-Pacific [YWC^{+}21]. Transboundary [BMJ^{+}24, CMIMS20, KPWS21]. transcriptome [HLG^{+}21]. transform [BHB24, DOB^{+}24, LTT^{+}23, LIA25, PJSQ20]. transient [LTR20b, SBBH24]. transmission [dHJTCE23]. transmitter [TRS^{+}24]. transport [GGL^{+}24, LTR20a, MSJ21, PRWK20, SKJJ25]. Transportation [FG23, LTR20b]. trap [BBJ22, BBSM24, GKC21, GJGW23, LLK^{+}22, LLH^{+}25, MKS^{+}21a, MDJP24, NLW^{+}22, OAM^{+}21, PRWK20, RSPE22, SPG^{+}21, SKJJ25, SSG^{+}22, TNTN20, VR20, YLS^{+}23]. trap-net [LLK^{+}22]. traps [CBTH20, GCK^{+}21, MMQ21, PFGQ20, RBHM24a, TJS23]. trauma [EFM25]. travel [HBMC21, LCHB^{+}24, PJOR20]. trawl [ASE21, BMM^{+}24, BWR24, BGM^{+}23, BNL^{+}23, BRvL^{+}22, BLHS20, BS20a, BM23b, BKM23a, BKM23b, CYBW22, CTM^{+}20, CAH^{+}20, DCK^{+}22, DML^{+}20, GDVBB^{+}20, HEG^{+}23a, IB20, IJS^{+}22, JLYR24, JBL^{+}22, JMP^{+}21, KFH^{+}25, LWH^{+}21, MINS21, MRG^{+}23, NBEI23, OFF^{+}20, PTK^{+}20, PCF23, PCC^{+}23, QAS^{+}25, RBM21, RNP^{+}24, RL24, SJBT20, SBL^{+}23, SJ24, SFPR^{+}23, TRWH23, TGG^{+}24, TYYK21, TZL^{+}24, VMFF^{+}20, WGNM24, YSB^{+}21, vORP23]. trawlers [LCW^{+}23a, MTC^{+}22, UAB^{+}21, UAG^{+}23]. Trawling [TMSS^{+}23, AAM^{+}20, GGMRC^{+}22, IAB20, LM22, TPD20]. trawls [BM22, BM24a, BM24b, Bro25, BM25a, BM25b, GGTÁVT^{+}20, KHGB25, RB22, SOSK22, TTL^{+}20]. treatments [HCDB22]. trecae [MPV^{+}24]. tree [HF20]. Trends [AHX^{+}24, WCN^{+}24, ACL^{+}20, BAF23, CAYM^{+}23, DKBF23, GKC^{+}22, GGMRC^{+}22, MUF^{+}22, MYKO23, OAB^{+}23, OOAF^{+}21, SS24, YHST22, ZCFG23]. trevally [DMM^{+}23, GFC^{+}22]. triad [CTIC23]. trial [CWM^{+}23]. Trials [SV23, HVMEO25, KBB^{+}21]. Triangle [MTE^{+}20]. tributaries [LBLF20]. tributary [BS20b]. Triggerfish [CPR^{+}24, PRRR23]. triggering [RB22]. Triglops [TEO25]. Trinity [RMRG22]. trips [FOS^{+}21, HBMC21, HCDF24, SL24]. trituberculatus [YHC^{+}24]. trolling [AGB^{+}20]. Trophic [SAA^{+}22, SWIRF21, YLX^{+}24, ZZ22, FHSC21, GPASM22, KDdOM^{+}22, LCFJ22, LXC^{+}20, DBL^{+}25, WHCF22, WWT^{+}25]. Trophodynamics [QMGRIU22]. tropical [AAPG21, AAG22, BDA^{+}20, BGL^{+}22, CTM^{+}20, CHT20, ECK^{+}21, GGG^{+}22, GFDN^{+}22, HHD^{+}20, HL20, KHS^{+}20, LNP25, LCLM23, LCN^{+}20, LLFL21, MdCG20, MBH^{+}22, NLS21, OYOO21, PSS^{+}20, PVPN22, PDM^{+}24, PRN^{+}24, PBB^{+}22, QSMG^{+}23, RdBAT^{+}23, SSV^{+}20, SASB24, SKW^{+}21, SJTGAS23, TFC^{+}20, VR20, BHST^{+}21]. Trout [JSKM20, RPL^{+}24, AJB20, AGB^{+}20, BR22, CGSL22, DH20, ECY21, HQWD20, JAN23, SGW^{+}22, SFJ^{+}23, SBBH24, TAA^{+}20, YiTM23, ZDZR^{+}22]. truncation [MRUG^{+}23]. trunculus [dAGCR21]. truth [WS24]. trutta [LR21, JAN23, ZDZR^{+}22]. TS [RGG22]. tshawytscha [IRJ^{+}22]. Tsugaru [NY23]. tulipa [OYOO21]. TUN [PNGGO^{+}22]. TUN-AI [PNGGO^{+}22]. Tuna [Mel23, PNGGO^{+}22, AAG22, ATA^{+}24, AGS20, AAFLL^{+}25, AUM21, Ash20a, Ash20b, ASO^{+}22, AIW^{+}23, BDA^{+}20, BW20, BGL^{+}22, BUG^{+}24, CHT20, CCCM^{+}20, CNDDAPMR21, DDCNMR21, FFG^{+}20, FMMA20, GGG^{+}22, HBC^{+}22, HHD^{+}20, HWMVM23, KWC^{+}20, KNP^{+}20, LCLM23, LCDM^{+}24, LK25, MCS^{+}24, MLCMdS23, NLS21, NSQV22, NBF20, ORdlG^{+}24, PBPM^{+}23, PZG^{+}20, RBG^{+}20, SASB24, SXMV^{+}21, SFC21, SF22a, SPG^{+}21, SAW^{+}20, SJTGAS23, SBRM^{+}22, SWH24b, SML^{+}24, TFC^{+}20, WSWL25, WAA^{+}22, XMLCMV24]. tuna/swordfish [CHT20]. Tunas [RCH^{+}21, AAPG21, APGG22, HL20, MBE^{+}20, MAA^{+}20, MBH^{+}22, PVPN22, PGAG22]. tuning [Tho19, Tho21]. tunny [BNTK23]. turban [KBH22]. turbot [BGK22, SC20]. turbulent [NVB^{+}23]. TURF [GHAZ21]. Turkey [ÖÜÖG20, SPW^{+}22]. Turkish [ŞGK^{+}20]. turnover [HSM21, HTK^{+}24]. turtle [CJC^{+}20, CTM^{+}20]. turtles [PMS^{+}23, SOS^{+}21, VCPO21]. Twenty [MCGL^{+}25]. twine [TTL^{+}20]. two [AdABW^{+}22, BLFT23, BM24b, CHGC25, CdSLP21, CAAFH21, FGCB^{+}21, GMT24, GW21, KHE^{+}22, KCB^{+}24, KK22, LPAE^{+}24, LZL^{+}24, MINS21, MFR22a, Mun24, NVRG^{+}21, PDG^{+}22, PBM^{+}23, RGG22, SDdMG^{+}20, SXM^{+}21, TSPK24, WASS20, YiTM23, YWC^{+}21]. two-dimensional [SXM^{+}21]. two-stage [KK22]. two-stage-structured [Mun24]. type [AHL20, AUM21, AK23, KC22, LCLM23, PCM^{+}21, SGD^{+}21, TCL^{+}21]. types [FBB20, GP25, KKCP20, LRGB25]. typology [SPE^{+}23]. Tyrrhenian [PDG^{+}22].

U.S. [Dra22, EBGE21, JLYR24, KKLM24, KWC^{+}20, KFDE^{+}22, KMC20, LWH^{+}21, PBPM^{+}23, SS23, SGH^{+}20]. U.S.A [MWJ^{+}24]. U.S.A. [LHPR21]. ucá [MPEBdR23]. Ucides [ASD^{+}22, MPEBdR23]. Uganda [NCB^{+}23]. UK [BCPH22]. Unaccounted [WPLF20]. unbiased [KK22]. uncertain [SBC^{+}23]. uncertainties [MBE^{+}20]. uncertainty [Aks24, BDM^{+}20, CH22, DQMV21, DBV22, DBV23, KMA22, MBOCdAM24, MZSZVP^{+}23, PPH21, PJP20, PSS^{+}21, PDA^{+}24, SS23, SJW^{+}22, WH23]. uncoupling [FG21]. Uncovering [SPC^{+}23]. uncrewed [ERS^{+}23, LD25]. undatum [CB20, EHBJ20]. underestimated [FJS^{+}25]. Underpinning [BMA^{+}24]. underscores [FDS^{+}23]. undersized [CAH^{+}20, VS23]. understand [AHB^{+}22, DMS22]. Understanding [CMA^{+}22, EBGE21, GMPD23, HSJ^{+}24, PSS^{+}20, PJOR20, HEG^{+}23b, KWE^{+}21, PCK23, PGRD21, WSL21]. Underwater [SPG^{+}21, ASWS^{+}21, CSSB22, FBALRR^{+}22, LMK^{+}23, WGW^{+}23]. Undetected [SASB24]. undisputed [ASB^{+}24]. Unexpected [RdLSdB^{+}21, TBH^{+}22]. Union [RMNB^{+}21]. unit [DBGV^{+}22, FBB20, HCDB^{+}24, HCDB22, MTX^{+}20, SKST23, TMP20, XMLCMV24]. United [BPT^{+}20, BGG^{+}22, BPT^{+}25, DSC24]. units [KSV^{+}22]. universal [HSL^{+}22]. Unlocking [CTIC23]. unobserved [SAHW22]. unregulated [BS20b]. unreported [CFO23]. unsupervised [SPD^{+}24]. untrawlable [BWOR23]. unvegetated [ESB^{+}24]. unwanted [FMMC20]. update [SBC^{+}23]. Updated [GAW^{+}22, WAA^{+}22]. Upgrading [CH22]. upper [ALRB^{+}20, ZOZW22]. ups [AHL20, SCS25]. upstream [RBH^{+}24]. upwelling [OOAF^{+}21]. urban [KKLM24]. urban-to-rural [KKLM24]. urchin [BWB^{+}24]. Uroteuthis [YTHM20, YTH22]. USA [HJMS20, HMR24a, SL24, ASJ^{+}22, CG21, EWPB22, KSI20a, MCC20]. usage [VTS^{+}22]. Use [BWG^{+}21, BHNP22, CG21, GHAZ21, HLC^{+}25, LWH^{+}21, MCPJET^{+}20, AAR^{+}21, BPT^{+}25, BBR^{+}22, CLB^{+}21, DAR^{+}23, GMK23, GRJW20, Har21, HA23, JLYR24, KAB^{+}22, KWW^{+}21, KSJM^{+}20, KFI^{+}21, LJB^{+}24, MVMdS^{+}24, OCBJG20, PDC^{+}24, PTD^{+}20, SMKJ21, TSG25, ZHM23, dAdCdO^{+}23, dLH23]. used [AWC^{+}23, CSRL20, LCC25, MOTL25]. useful [BRR^{+}21, Mun24]. user [SKBA23]. users [GAB^{+}20]. uses [XqRJ^{+}23]. Using [BGG^{+}22, BW20, BTB^{+}21, BKM23b, BCF^{+}23, CZ25, CSH^{+}21, CTS^{+}23, CCCM^{+}20, CAYM^{+}23, DMM^{+}21, FDB^{+}20, GJSW22, GW21, HBE^{+}22, KFO20, KFH^{+}25, KBB^{+}21, LJB^{+}24, MZ24, PBB20, PPC^{+}23b, PRWK20, PBRT22, PD25, SSM^{+}23, SF22b, SCC^{+}22, WST^{+}23, AFR^{+}24, AAG22, AAT^{+}21, APB^{+}20, ATA^{+}24, ASWS^{+}21, BCG^{+}25, BMA^{+}24, BRGB^{+}23, BBHF25, BWB^{+}24, BCOBB^{+}23, BRN^{+}20, BBSM24, CPM21, CEAL21, CWC^{+}21, CTIC23, CKD^{+}21, CSSB22, CBHS24, DSC24, DBS^{+}21, DMF^{+}21, DMM^{+}23, DWLT21, EBGE21, FCKG^{+}22, FM21, FBALRR^{+}22, Fre22, GM21, GAW^{+}22, GPT^{+}21, HPL^{+}24, HC22b, HWW22, HL21, HRH22, HCK^{+}21, HSL^{+}22, HW24, JDP22, JMS25, KM23, KMA23, KBPS21, KBPS22, KK22, KSS^{+}22, KMO20, KBM23, LMT^{+}22, LCFJ22, LMP24, LBP^{+}24, LMK^{+}23, LD25, LBD23, LKSiI22, LIA25, LAG^{+}21, MLCMdS23, MMQ21, MDJP24, Mon24, MOB^{+}22, MMBH23]. using [NY23, NAS^{+}20, OMK24, OOM^{+}23, VS23, OCdMC24, OOAF^{+}21, PÁEMC22, PRCF22, PHH^{+}23, RKN23, RdBAT^{+}23, RGG22, RNP^{+}24, RWT^{+}20, RBG^{+}20, RPM^{+}21, RMD^{+}25, RTB^{+}21, SCGM^{+}21, SKY^{+}24, SKW^{+}21, SSP^{+}23, SSI^{+}23, SIM^{+}24a, SIM^{+}24b, SDV^{+}22, SDBS21, SBJ^{+}20, SJ24, SFYM24, SFPR^{+}23, SPD^{+}24, SMA^{+}24, SACS23, SGKAR24, SLW^{+}20, SME^{+}24, TTK^{+}25, TP24, TBÓ^{+}22, TDI^{+}21, TCL^{+}24, WLZ^{+}21, WZS^{+}21, WYM^{+}25, WGNM24, WMT^{+}20, WMSW22, XCB^{+}21, XWD^{+}21, YIM^{+}20, YAO^{+}23, ZHC^{+}25]. Utah [SL24]. utilisation [PAA^{+}24]. Utility [TC24, BBJ22, CHPT20, GFM^{+}23, LBW21, RWB^{+}23, SDdMG^{+}20, SNHM23]. utilization [CEAL21, DSJG20, KM23, SLW^{+}21]. Utilizing [BMM^{+}21, GC25, SHC21].

v [HGHH25]. v-notching [HGHH25]. Valenciennes [SNJ^{+}24]. validated [CSB^{+}23]. validates [CPR^{+}24]. Validating [BM22]. Validation [ACS23, HKG^{+}21, NEBP^{+}23, PRRR23, UFYT23, BBHF25, BKHA21, DSP^{+}23, GFDN^{+}22, TAK^{+}23]. Validity [ECK^{+}21, KHL^{+}24]. Valley [LP23]. valuable [KK21]. value [AOA^{+}22, CAGLT23, DSP22, dHJTCE23, HBMC21, JCL^{+}21, KHMC23, LFdB^{+}21, RRSP^{+}24, SO25, SRP^{+}22, Seu22, SK22, SL24, vP20]. values [PAY23]. Vänern [AGB^{+}20]. variabilis [TSG25]. variability [AAPG21, CGDTSA^{+}25, CH21b, CRS23, Dra22, DJFF23, HSJ^{+}24, HW24, HLI^{+}20, KPUB22, LCD^{+}23, LBD24, MJD^{+}21, MKS^{+}22, MRE^{+}24, Mur21, NVB^{+}23, PAY23, SF22a, SJH^{+}23, SWIRF21, TTYT24, TBF^{+}21, TZL^{+}24, WZX^{+}20, WBA23, ZOS^{+}23]. variable [CSB^{+}23, DLKH22, GSH22, PW25, PBDM23, PAY23, VP22]. variables [dFBPL^{+}20, CEAL21, CMBL21, KHS^{+}20, OCdMC24]. variance [DLP^{+}24]. variate [YIM^{+}20]. Variation [FTH^{+}23, GQP^{+}25, ABK^{+}21, ASD^{+}22, BPT^{+}20, CJC^{+}20, DSB^{+}21, DHCS23, FHSC21, FJJT^{+}21, FBW^{+}21, GDVBB^{+}20, GAB^{+}24, HTSJ23, LM22, LWX^{+}20, LLC^{+}20, MDC^{+}22a, NNS^{+}22, OSEF22, RdBAT^{+}23, SKD^{+}20, SDC^{+}22, SBBH24, Szu22, TSS^{+}23, TSI^{+}21, WPGO21, ZJJ^{+}25]. Variations [PFFdC22, SZXC25, WWF^{+}20, CZ25, MVLC^{+}20, NAV^{+}23, PSS^{+}20, PTK^{+}20, SSD^{+}20b, TNDM23]. various [GP25, XqRJ^{+}23]. vary [BHB^{+}22, PDF20]. Varying [RWFT25, CWRR24, CC22, CHAY^{+}25, CFP21, DB22, FNKY20, KHPB20, SM21, TSPK24]. VAST [CBA^{+}24, EYAO20]. Vateritic [GQP^{+}25]. VBD [LSZ^{+}23]. vector [CBA^{+}24]. vegetated [ESB^{+}24]. vegetation [NHE^{+}23]. vehicle [LD25]. vendace [KSV^{+}22, SHV20]. vent [ZZH^{+}24]. ventilation [BSR^{+}22]. ventless [MKS^{+}21a]. vents [ZZH^{+}24]. venus [MPM22, ÖA21, PHV^{+}21, PUC^{+}23]. Veracruz [GGMMMV^{+}20]. Vermilion [KHE^{+}22, CSB^{+}23]. verreauxi [ML24]. verrucosa [PUC^{+}23]. version [CH22]. versus [BGL^{+}22, BN21, CMD^{+}23, MLS^{+}21, PFGQ20, SFC21, SGZD20]. vertebrae [PRRR23]. Vertical [SIMT24, BKM23b, FBALRR^{+}22]. vessel [BUG^{+}24, BM23a, DHH^{+}22, DKD^{+}21, HBE^{+}22, KM23, LSZ^{+}23, ÖRS^{+}25, SJTGAS23, SZS^{+}24, TPD20, YRTP20]. vessel-level [KM23]. vessels [DHH^{+}22, ERS^{+}23, FBW^{+}21, LZC^{+}21, LCL25, RNP^{+}24, SHH^{+}21, SFCG^{+}21, WZL^{+}22]. via [BJS^{+}22a, CSB^{+}23, SM21]. Viability [WLG^{+}23, DMS22]. viable [Bea21]. Victoria [FBM^{+}21, NAS^{+}20, NNS^{+}22, NC20, NNTM^{+}20, PvZ22]. video [ASWS^{+}21, BGG^{+}22, BBSM24, CSSB22, LMK^{+}23, SBL^{+}23, TSC^{+}22, TLAM25, WSB22]. Vietnamese [CEAL21, NSQV22]. viewpoint [Mau22]. VIIRS [LSZ^{+}23]. villosus [ASGG21, FJJT^{+}21, FJHT^{+}22, SBJ^{+}20, SGKAR24]. Visible [vBMP^{+}23, KBB^{+}21]. vision [SKY^{+}24, SLA24]. visitor [KWW^{+}21]. visual [CB20, WASS20]. Visualizing [Peñ21, TDJ^{+}21]. vitality [LYH^{+}21]. vitreus [EVS^{+}23, GLA^{+}20, GW21, HMC^{+}23, OGFW24, PSSFS24, SPD^{+}24]. VMS [LZC^{+}21, MDJP24, SCGM^{+}21]. vocabulary [WSL^{+}24]. volatility [dHJTCE23]. volcano [GGMRC^{+}22]. Volta [MA20]. volumetric [DMM^{+}21]. voluntary [WSF22]. vs [BM24a, LCB^{+}21a, Mur20, TBF^{+}21, WKSF20]. vulgaris [ALRB^{+}20, LCB^{+}21a, SRT^{+}20]. vulnerabilities [PSS^{+}20]. Vulnerability [MCH21, PdAMdM^{+}23, CBTH20, FM21, MCPJET^{+}20, PDJ24]. vulnerable [BALBC23].

W [SPG^{+}21]. Walleye [FFG21, GLA^{+}20, OGFW24, SPD^{+}24, ZSWK25, DAL20, DB22, EYAO20, EVS^{+}23, GW21, HMC^{+}23, PSSFS24, YMYH20]. warm [BCS^{+}22, KWW^{+}21]. warm-water [BCS^{+}22]. warming [GGL^{+}24, WJN^{+}25]. Warsaw [SR21]. was [MAH^{+}22]. washer [PCM^{+}21]. Washington [BQBW20, CBR^{+}23]. waste [RRSP^{+}24]. Water [MMC^{+}24, ASB^{+}24, BBČ^{+}21, BCS^{+}22, BLFT23, BMM^{+}21, CPL^{+}25, CKK^{+}20, CCC^{+}22a, DLZ^{+}25, FMD^{+}24, IJS^{+}22, PDG^{+}22, RPL^{+}24, SLF23, SPC^{+}23, VTSI^{+}24, WWF^{+}20, YH21a, YH21b]. Water-level [MMC^{+}24]. Waterbody [KPK^{+}23, KAB^{+}22]. waters [AAM^{+}20, AMM^{+}22, CW22, DML^{+}20, DLP^{+}24, GGL^{+}24, GLP^{+}20, HBC^{+}22, HM25, HHD^{+}20, KWW^{+}21, MSC^{+}24, MMG^{+}24, MDS^{+}20, NSM^{+}21, NBF20, PIP^{+}22, SJB^{+}20, TPW23, ZLXL20]. waved [CB20]. wavelet [LTT^{+}23]. waves [KWW^{+}21]. way [BBČ^{+}21, JWL^{+}24]. ways [KWW^{+}21]. weak [AdABW^{+}22]. weakening [KZT^{+}23]. Weakfish [SdOR^{+}23]. weather [SL24]. Web [RFMS^{+}21]. Web-based [RFMS^{+}21]. webs [EHE^{+}23]. weekdays [KWW^{+}21]. weight [CTIC23, CLM^{+}22, DSP^{+}23, GBC^{+}23b, GKM^{+}23, Lya20a, MRC24, SR20]. weight-at-age [CTIC23]. weighted [CG21, HLZ^{+}20]. weighting [TMH23]. weights [DET21b]. Welfare [Ham22, CBN^{+}21, RS21, SCS25, ZDZR^{+}22]. well [LCDM^{+}24, MSW21]. wells [SASB24]. West [BTC23, JLYR24, LWH^{+}21, SS23, Dra22, MVDH24, SNJ^{+}24, VAVQGD^{+}20, YTHM20, BQGV^{+}24, BHB^{+}22, CAAFH21, FJHT^{+}22, MPSH25, YFJ^{+}25]. west-coast [MVDH24]. west-flowing [SNJ^{+}24]. Western [AGL^{+}24, GFC^{+}22, SBT^{+}20, AGS20, AAFLL^{+}25, Ash20a, Ash20b, BFA^{+}21, CHT20, CAGLT23, CCKL^{+}20, EHE^{+}23, FHE^{+}24, GGMRC^{+}22, GMERCM^{+}24, KTFY22, KNO^{+}21, KPUB22, LWH^{+}23a, LNvD^{+}23, MPM^{+}23, PVPN22, QMC^{+}22, RNKB23, SPM^{+}24, SPC^{+}23, VP23, WCLN20, YTH22, dLHER24, BMOC22, BJK24, CSTdL25, CPPK23, HdLHD22, JYH21, LRMH21, PDC^{+}24, RCVGMI22, SBD^{+}22, SFYM24, SML^{+}24, HLI^{+}20]. westslope [HQWD20]. wetland [DBDT21]. Whale [FMMA20, ABBO20, BTR^{+}24, CTR^{+}21]. whales [MRE^{+}24]. WHAM [SM21]. whelk [CB20, EHBJ20, MFJ^{+}24, WPGO21]. Which [MOI23, KWW^{+}21]. while [KLBHK23, MRG^{+}23, RBHM24a]. white [ÁHGCVAI22, CCRGC^{+}24, BBM^{+}24]. whitefish [HA23, VCG^{+}23, WKSF20]. Whitemouth [ABT^{+}24]. whiting [BRvL^{+}22, BM22, BM23a, BM25b, BM25c, BMA^{+}20, MABR^{+}20, VS23, ŞGK^{+}20, YGMJ20]. whitsoni [MPP22]. who [Pun23]. whole [HGS^{+}23]. whole-lake [HGS^{+}23]. wide [BNL^{+}23, CPB^{+}21, CPPK23, HIF^{+}24, LLC^{+}20, RKN23, YGMJ20, ZJY^{+}24]. widely [PBD25]. widely-distributed [PBD25]. widespread [MB25]. Wild [FKS^{+}20, DHC^{+}20, GAB^{+}24, KDdOM^{+}22, LLC^{+}20, LZW^{+}21, NA22, RSPE22, SK22, TBÓ^{+}22, ZOZW22]. wild-caught [GAB^{+}24]. will [FWL20]. Willingness [BKM^{+}23c]. win [RNKB23]. wind [WHR^{+}24, dAGCR21]. window [LIA25]. windowpane [BMM^{+}21]. Winds [KWW^{+}21]. wing [JMP^{+}21]. winter [ABT^{+}24, CZ25, EMJ^{+}22, KHL^{+}24, MLS^{+}21, QMGRIU22, Sim23, TCL^{+}24]. wire [Bea21, RBHM24a, SW20]. Wisconsin [EVS^{+}23, FWL20, LOFS22, SLF23]. Within [LCDM^{+}24, EHBJ20, FFG^{+}20, FCSA21, LG21, MMP^{+}24, NRH^{+}23, PBM^{+}23, RMNB^{+}21, SL24, WS24]. Within-well [LCDM^{+}24]. without [WGNM24]. wolf [AGNS^{+}21]. wolffish [GKM^{+}23]. Women [NC20]. Woods [SM21]. woodwardi [CW22]. workflow [CBA^{+}24, Mon24]. workflows [HIKM21]. working [WZL^{+}22]. Workshop [CSTdL25, MPSM25]. world [aLBK^{+}21]. wound [SFJ^{+}23].

Xiphias [FCKG^{+}22, MPM^{+}23, SB20, TPW23].

Yangtze [aFLpX^{+}21, LCG^{+}21, SLW^{+}21, ZOZW22]. year [AGS20, BAF23, CGM^{+}22, CTIC23, LADA^{+}22, LCW23b, PD25, ZJJ^{+}25]. yearly [MYKO23]. years [AHX^{+}24, BTFL22, GBWM22, HPPT24, LMJ^{+}23, MCGL^{+}25, MCS^{+}25, MAH^{+}22, SCW21, YHST22]. Yellow [DKBF23, GWGM24, CQA^{+}24, DJFU20, LLC^{+}20, SXK^{+}24, WLZ^{+}21, YLP^{+}23, ZLXL20, HSJ^{+}24, SXK^{+}24, YLS^{+}23]. yellowfin [AAFLL^{+}25, BW20, FFG^{+}20, HHD^{+}20, HWMVM23, PBPM^{+}23, PZG^{+}20, SF22a, SML^{+}24, Ten22]. yellowtail [SXM^{+}21]. yield [BÖN20, HLZ^{+}20, LCW^{+}23a, Lya22, PDF20, SHC21]. yield-density [HLZ^{+}20]. yields [LZY^{+}24c, VTSI^{+}24]. Ylistrum [CLY^{+}22]. yolk [SMC^{+}24]. yolk-sac [SMC^{+}24]. YOLOv5 [XDX^{+}23]. YOLOv7 [SZS^{+}24]. YOLOv7-DCN-SORT [SZS^{+}24]. young [AGS20, BAF23, CGM^{+}22, DHC^{+}20, LCW23b, PD25]. young-of-the-year [AGS20, LCW23b, PD25]. young-of-year [BAF23, CGM^{+}22]. younger [LLH^{+}25]. YOY [NRH^{+}23].

Zanzibar [HMY25]. Zealand [LPP^{+}20, Mac25, MTS^{+}21, MOB^{+}23, NAV^{+}23, RGP^{+}23, SRF^{+}24]. Zhongjieshan [LWX^{+}20]. zonation [KHC^{+}20]. Zone [SKST23, PRRR23, SCGM^{+}21]. zones [KMC^{+}23, MFR22a]. zoning [BHG^{+}24b]. zooplankton [BAW^{+}24, KTFY22, SCG^{+}24, VTS^{+}22]. Zygochlamys [PCF23].

References

Artetxe-Arrate:2025:OSI

[AAFLL^{+}25] Iraide Artetxe-Arrate, Igaratza Fraile, Patricia Lastra-Luque, Jessica Farley, Naomi Clear, Umair Shahid, Shoaib Abdul Razzaque, Mohamed Ahusan, Annie Vidot, Denham Parker, Francis Marsac, Hilario Murua, Gorka Merino, and Iker Zudaire. Otolith stable isotopes highlight the importance of local nursery areas as the origin of recruits to yellowfin tuna (Thunnus albacares) fisheries in the western Indian Ocean. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003059.

Akia:2022:ETR

[AAG22] S. Akia, M. Amandé, and D. Gaertner. Estimating tagreporting rates for Atlantic tropical tuna fleets using coincidental tag return and tag seeding experiment data. Fisheries Research, 253(??):Article 106372, September 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001497.

Anders:2023:ISS

[AAH^{+}23] Neil Anders, Kenneth Arnesen, Anette Hustad, Terje Jørgensen, Svein Løkkeborg, Sten Siikavuopio, Tina Thesslund, and Anne Christine Utne-Palm. Improving size selection in the Norwegian red king crab (Paralithodes camtschaticus) fishery through modification to pot design and soak time. Fisheries Research, 261(??):Article 106641, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000346.

Abdulqader:2020:BDS

[AAM^{+}20] Ebrahim A. A. Abdulqader, Pulikkodan Abdurahiman, Lamjed Mansour, Abdel Halim Harrath, Mohammad A. Qurban, and Lotfi Rabaoui. Bycatch and discards of shrimp trawling in the Saudi waters of the Arabian Gulf: ecosystem impact assessment and implications for a sustainable fishery management. Fisheries Research, 229(??):Article 105596, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301132.

Akia:2021:SIA

[AAPG21] S. Akia, M. Amandé, P. Pascual, and D. Gaertner. Seasonal and inter-annual variability in abundance of the main tropical tunas in the EEZ of Côte d’Ivoire (2000–2019). Fisheries Research, 243(??):Article 106053, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001818.

Avigliano:2021:PSO

[AAR^{+}21] Esteban Avigliano, Nadia M. Alves, M. Rita Rico, Claudio O. Ruarte, Luciana D’Atri, Ana Méndez, Jorge Pisonero, Alejandra V. Volpedo, and Claudia Borstelmann. Population structure and ontogenetic habitat use of Micropogonias furnieri in the Southwestern Atlantic Ocean inferred by otolith chemistry. Fisheries Research, 240(??):Article 105953, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621000813.

Aldrin:2021:CEN

[AAT^{+}21] M. Aldrin, F. L. Aanes, I. F. Tvete, S. Aanes, and S. Subbey. Caveats with estimating natural mortality rates in stock assessment models using age aggregated catch data and abundance indices. Fisheries Research, 243(??):Article 106071, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001995.

Acuna-Alonso:2023:AEF

[AAVÁM23] Carolina Acuña-Alonso, Simone Varandas, Xana Álvarez, and António Martinho. Analysis of the evolution of a fisheries management plan based on environmental governance: Living laboratory in the Olo River, Portugal. Fisheries Research, 260(??):Article 106595, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003721.

Alsolami:2020:EBF

[AAZ20] Lafi S. Alsolami, Mohammed Abdelaty, and Chang-Ik Zhang. An ecosystem-based fisheries assessment approach and management system for the Red Sea. Fisheries Research, 227(??):Article 105551, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300680.

Anderson:2020:IAA

[ABBO20] David Anderson, Robin W. Baird, Amanda L. Bradford, and Erin M. Oleson. Is it all about the haul? Pelagic false killer whale interactions with longline fisheries in the central North Pacific. Fisheries Research, 230(??):Article 105665, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030182X.

Amelot:2021:EDL

[ABF^{+}21] Morgane Amelot, Jurgen Batsleer, Eric Foucher, Raphaël Girardin, Paul Marchal, Jan Jaap Poos, and Klaas Sys. Evidence of difference in landings and discards patterns in the English Channel and North Sea Rajidae complex fishery. Fisheries Research, 242(??):Article 106028, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001569.

Athukoorala:2021:TVG

[ABK^{+}21] A. A. S. H. Athukoorala, Ram C. Bhujel, Jens-Otto Krakstad, Takuji W. Tsusaka, and Takashi Fritz Matsuishi. Temporal variation of grouper diversity and distribution on the continental shelf of Sri Lanka: a revisit after four decades. Fisheries Research, 242(??):Article 106017, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001454.

Alves:2024:WCM

[ABT^{+}24] Nadia M. Alves, Mara S. Braverman, Brenda Temperoni, Julieta S. Rodríguez, and Marina V. Diaz. Whitemouth croaker (Micropogonias furnieri) are you ok? Reduction in nutritional condition of juveniles during winter in the Southwest Atlantic Ocean. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000134.

Alves:2020:ATP

[ACL^{+}20] Luís M. F. Alves, João P. S. Correia, Marco F. L. Lemos, Sara C. Novais, and Henrique Cabral. Assessment of trends in the Portuguese elasmobranch commercial landings over three decades (1986–2017). Fisheries Research, 230(??):Article 105648, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030165X.

Avigliano:2023:SIM

[ACP^{+}23] Esteban Avigliano, Ming-Tsung Chung, Marc Pouilly, Kuo-Fang Huang, Miguel Casalinuovo, Jael Dominino, Natalia Silva, Sebastian Sánchez, Juan F. Facetti, and Alejandra V. Volpedo. Strontium isotope mapping and its application to study the fish life history (Salminus brasiliensis) in semi-fragmented rivers (La Plata Basin, South America). Fisheries Research, 265(??):??, September 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001340.

Anderson:2023:VAE

[ACS23] Austin J. Anderson, Andrew M. Claiborne, and Wade Smith. Validation of age estimates for Chum and Sockeye salmon derived from otolith and scale analysis. Fisheries Research, 259(??):Article 106556, March 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003332.

Aderne:2022:ADS

[AdABW^{+}22] Alexandre Falcão Aderne, Jamille de Araújo Bitencourt, Luciana Almeida Watanabe, Horacio Schneider, Paulo Roberto A. de Mello Affonso, and Iracilda Sampaio. Allopatric divergence and secondary contact of two weak fish species (Macrodon ancylodon and Macrodon atricauda) from the South Atlantic. Fisheries Research, 245 (??):Article 106126, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362100254X.

Avaca:2025:FSS

[AdlBCN25] María Soledad Avaca, Paula de la Barra, Ana Cinti, and Maite Narvarte. Factors sustaining the snail artisanal fishery in the San Matías Gulf, Patagonia, Argentina. Fisheries Research, 282(??):??, February 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000190.

Fang:2021:RBG

[aFLpX^{+}21] Di an Fang, Yu-Ting Luo, Dong po Xu, Xi wen Yang, and Xiao hao Wang. Relationship between genetic risk and stock enhancement of the silver carp (Hypophthalmichthys molitrix) in the Yangtze River. Fisheries Research, 235(??):Article 105829, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303465.

Acharya:2024:TAA

[AFR^{+}24] Debaditya Acharya, Moshiur Farazi, Vivien Rolland, Lars Petersson, Uwe Rosebrock, Daniel Smith, Jessica Ford, Dadong Wang, Geoffrey N. Tuck, L. Richard Little, and Chris Wilcox. Towards automatic anomaly detection in fisheries using electronic monitoring and automatic identification system. Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000031.

Andersson:2020:RTE

[AGB^{+}20] A. Andersson, L. A. Greenberg, E. Bergman, Z. Su, M. Andersson, and J. J. Piccolo. Recreational trolling effort and catch of Atlantic salmon and brown trout in Vänern, the EU’s largest lake. Fisheries Research, 227(??):Article 105548, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300655.

Arostegui:2024:FID

[AGB^{+}24] M. C. Arostegui, P. Gaube, M. Bowman, K. Nakamaru, and C. D. Braun. Fishery-independent and -dependent movement data aid in defining the stock structure of a data-deficient billfish. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003168.

Andrzejaczek:2024:SER

[AGL^{+}24] Samantha Andrzejaczek, Adrian Gleiss, Karissa Lear, Frazer McGregor, Taylor Chapple, and Mark Meekan. Stomach eversion and retraction by a tagged tiger shark at Ningaloo Reef, Western Australia. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002680.

Atencia-Galindo:2021:GSP

[AGNR^{+}21] Mayra A. Atencia-Galindo, Juan C. Narvaéz, Argiro Ramírez, Jorge Paramo, and Juan C. Aguire-Pabón. Genetic structure of the pink shrimp Penaeus (Farfantepenaeus) notialis (Pérez–Farfante, 1967) (Decapoda: Penaeidae) in the Colombian Caribbean. Fisheries Research, 243(??):Article 106052, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001806.

Andrade:2021:RIP

[AGNS^{+}21] Laura Simões Andrade, Domingos Garrone-Neto, Manuela Alves Nobre Sales, Luciana Rodrigues de Souza-Bastos, Ursulla Pereira Souza, and Helen Sadauskas-Henrique. Reflex impairment and physiological stress response in the neotropical wolf fish Hoplias malabaricus (Characiformes, Erythrinidae) exposed to catch-and-release angling. Fisheries Research, 239(??):Article 105940, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000680.

Arai:2020:SAC

[AGS20] Kohma Arai, John E. Graves, and David H. Secor. Subannual cohort representation among young-of-the-year recruits of the western stock of Atlantic bluefin tuna. Fisheries Research, 225(??):Article 105476, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303315.

Adams:2022:EAU

[AHB^{+}22] Grant D. Adams, Kirstin K. Holsman, Steven J. Barbeaux, Martin W. Dorn, James N. Ianelli, Ingrid Spies, Ian J. Stewart, and André E. Punt. An ensemble approach to understand predation mortality for groundfish in the Gulf of Alaska. Fisheries Research, 251(??):Article 106303, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000807.

Apostologamvrou:2024:RBC

[AHEV24] Chrysoula Apostologamvrou, Marianthi Hatziioannou, Athanasios Exadactylos, and Dimitris Vafidis. Reproductive biology of the commercial sea cucumber Holothuria (Roweothuria) poli, in the Central Aegean Sea, Greece. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783624002212.

Avila-Herrera:2022:EFL

[ÁHGCVAI22] Concepción Lizette Ávila-Herrera, Luis Manuel Guevara-Chumacero, José Antonio Velázquez-Aragón, and Ana Laura Ibáñez. Evidence of flow and low genetic structure between populations of the Gulf of Mexico and the Mexican Pacific of white mullet, Mugil curema (Pisces) inferred from single nucleotide polymorphisms. Fisheries Research, 256(??):Article 106484, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002612.

Althoff:2020:IRT

[AHL20] Andrew L. Althoff, Caleb T. Hasler, and Michael J. Louison. Impact of retrieval time and hook type on hooking depth in ice-angled northern pike caught on tip-ups. Fisheries Research, 225(??):Article 105502, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300199.

Abreu:2024:TPS

[AHX^{+}24] José Abreu, Philip R. Hollyman, José C. Xavier, Connor C. G. Bamford, Richard A. Phillips, and Martin A. Collins. Trends in population structure of Patagonian toothfish over 25 years of fishery exploitation at South Georgia. Fisheries Research, 279(??):??, November 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001863.

Anders:2023:IPE

[AIJ^{+}23] Neil Anders, Ólafur Arnar Ingólfsson, Terje Jørgensen, Svein Løkkeborg, and Odd-Børre Humborstad. Investigating the potential of escape openings and reduced mesh size to optimize snow crab (Chionoecetes opilio) pot catches in the Barents Sea. Fisheries Research, 258(??):Article 106517, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002946.

Akmal:2023:CDE

[AIM^{+}23] Yusrizal Akmal, Irfannur Irfannur, Muliari Muliari, Agung Setia Batubara, Muchammad Yunus, Hani Plumeriastuti, and Yeni Dhamayanti. A comprehensive description of the exoskeleton of six lobster species (genus Panulirus) in Aceh Province, Indonesia. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001248.

Ashida:2023:RSF

[AIW^{+}23] Hiroshi Ashida, Yukio Ishihara, Shuuyo Watanabe, Seiji Ohshimo, and Yosuke Tanaka. Revisiting the spawning fraction of Pacific bluefin tuna (Thunnus orientalis) caught in the Sea of Japan based on diurnal changes in spawning markers. Fisheries Research, 266(??):??, October 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001819.

Adams:2020:IAM

[AJB20] Jean V. Adams, Michael L. Jones, and James R. Bence. Investigating apparent misalignment of predator–prey dynamics: Great Lakes lake trout and sea lampreys. Fisheries Research, 232(??):Article 105734, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302514.

Atessahin:2023:EHS

[AK23] Tuncay Ateşşahin and Umar Khan. Effects of hook size and bait type on the size selectivity and short-time post-release mortality of a cyprinid fish (Cyprinus carpio L.) in recreational fisheries. Fisheries Research, 261(??):Article 106640, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000334.

Akselrud:2024:RFR

[Aks24] Caitlin I. Allen Akselrud. Random forest regression models in ecology: Accounting for messy biological data and producing predictions with uncertainty. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400225X.

Ayers:2022:FHD

[AL22] Adam L. Ayers and Kirsten Leong. Focusing on the human dimensions to reduce protected species bycatch. Fisheries Research, 254(??):Article 106432, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002090.

Lin:2021:RDS

[aLBK^{+}21] Bai an Lin, Robert Boenish, Jacob P. Kritzer, Yan Jiang, Song lin Wang, and Min Liu. Reproductive dynamics of a swimming crab (Monomia haanii) in the world’s crab basket. Fisheries Research, 236(??):Article 105828, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303453.

Alos:2020:IBA

[ALRA20] Josep Alós, Arancha Lana, Josep Ramis, and Robert Arlinghaus. Interactions between angler movement behaviour and an invasive seaweed with ecosystem engineering properties in a marine recreational fishery. Fisheries Research, 230(??):Article 105624, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301417.

Armelloni:2020:EED

[ALRB^{+}20] Enrico Nicola Armelloni, María Jesús Lago-Rouco, Aurora Bartolomé, Beatriz C. Felipe, Eduardo Almansa, and Catalina Perales-Raya. Exploring the embryonic development of upper beak in Octopus vulgaris Cuvier, 1797: New findings and implications for age estimation. Fisheries Research, 221(??):Article 105375, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302309.

Alvarez:2021:CPF

[Alv21] Antonio M. Alvarez. Comparison of proxies for fish stock. A Monte Carlo analysis. Fisheries Research, 238(??):Article 105901, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621000291.

Arlinghaus:2021:NOA

[ALW^{+}21] Robert Arlinghaus, Jorrit Lucas, Marc Simon Weltersbach, Dieter Kömle, Helmut M. Winkler, Carsten Riepe, Carsten Kühn, and Harry V. Strehlow. Niche overlap among anglers, fishers and cormorants and their removals of fish biomass: a case from brackish lagoon ecosystems in the southern Baltic Sea. Fisheries Research, 238(??):Article 105894, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000229.

Avigliano:2020:FSM

[AMdC^{+}20] Esteban Avigliano, Nathan Miller, Barbara Maichak de Carvalho, Sofía Córdoba Gironde, Andrea Tombari, and Alejandra V. Volpedo. Fin spine metals by LA-ICP-MS as a method for fish stock discrimination of Genidens barbus in anthropized estuaries. Fisheries Research, 230 (??):Article 105625, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301429.

Afonso:2021:ELA

[AMHH21] André S. Afonso, Bruno Mourato, Humberto Hazin, and Fábio H. V. Hazin. The effect of light attractor color in pelagic longline fisheries. Fisheries Research, 235 (??):Article 105822, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303398.

Aidoo:2022:SCD

[AMHR22] Eric N. Aidoo, Ute Mueller, Glenn A. Hyndes, and Karina L. Ryan. Spatial characterisation of demersal scalefish diversity based on recreational fishing data. Fisheries Research, 254(??):Article 106403, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001801.

Alexandre:2022:IBA

[AMM^{+}22] Sofia Alexandre, Ana Marçalo, Tiago A. Marques, Alexandra Pires, Mafalda Rangel, Adriana Ressurreição, Pedro Monteiro, Karim Erzini, and Jorge M. S. Gonçalves. Interactions between air-breathing marine megafauna and artisanal fisheries in Southern Iberian Atlantic waters: Results from an interview survey to fishers. Fisheries Research, 254(??):Article 106430, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002077.

Angelini:2020:BEP

[AMSC20] Silvia Angelini, Michela Martinelli, Alberto Santojanni, and Sabrina Colella. Biological evidence of the presence of different subpopulations of Norway lobster (Nephrops norvegicus) in the Adriatic Sea (Central Mediterranean Sea). Fisheries Research, 221(??):Article 105365, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302206.

Ascher:2024:CID

[ANB^{+}24] Alexander Ascher, Maura Niemisto, Donaven Baughman, Grace Andrews, Curtis Morris, Emily Patrick, Richard A. Wahle, and David M. Fields. Climate induced declines in maternal size may come at a cost to embryonic investment and larval performance in the American lobster. Fisheries Research, 276(??):??, August 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001231.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Fisheries Research, 224(??):??, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Fisheries Research, 228(??):??, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. Fisheries Research, 232(??):??, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:EBa

[Ano20d] Anonymous. Editorial Board. Fisheries Research, 221 (??):Article 105422, January 2020. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302772.

Anonymous:2020:EBb

[Ano20e] Anonymous. Editorial Board. Fisheries Research, 222 (??):Article 105448, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619303030.

Anonymous:2020:EBc

[Ano20f] Anonymous. Editorial Board. Fisheries Research, 223 (??):Article 105488, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300059.

Anonymous:2020:EBd

[Ano20g] Anonymous. Editorial Board. Fisheries Research, 224(??): Article 105513, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620300308.

Anonymous:2020:EBe

[Ano20h] Anonymous. Editorial Board. Fisheries Research, 225(??): Article 105544, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620300618.

Anonymous:2020:EBf

[Ano20i] Anonymous. Editorial Board. Fisheries Research, 226(??): Article 105581, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620300989.

Anonymous:2020:EBg

[Ano20j] Anonymous. Editorial Board. Fisheries Research, 227(??): Article 105602, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620301193.

Anonymous:2020:EBh

[Ano20k] Anonymous. Editorial Board. Fisheries Research, 228 (??):Article 105629, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301466.

Anonymous:2020:EBi

[Ano20l] Anonymous. Editorial Board. Fisheries Research, 229 (??):Article 105657, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301740.

Anonymous:2020:EBj

[Ano20m] Anonymous. Editorial Board. Fisheries Research, 230 (??):Article 105697, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620302149.

Anonymous:2020:EBk

[Ano20n] Anonymous. Editorial Board. Fisheries Research, 231 (??):Article 105741, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620302587.

Anonymous:2020:EBl

[Ano20o] Anonymous. Editorial Board. Fisheries Research, 232 (??):Article 105775, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620302927.

Anonymous:2020:F

[Ano20p] Anonymous. February 2020. Fisheries Research, 222(??):??, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Ja

[Ano20q] Anonymous. January 2020. Fisheries Research, 221(??):??, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Jc

[Ano20r] Anonymous. July 2020. Fisheries Research, 227(??):??, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Jb

[Ano20s] Anonymous. June 2020. Fisheries Research, 226(??):??, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Ma

[Ano20t] Anonymous. March 2020. Fisheries Research, 223(??):??, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:Mb

[Ano20u] Anonymous. May 2020. Fisheries Research, 225(??):??, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:N

[Ano20v] Anonymous. November 2020. Fisheries Research, 231(??):??, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:O

[Ano20w] Anonymous. October 2020. Fisheries Research, 230(??):??, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2020:S

[Ano20x] Anonymous. September 2020. Fisheries Research, 229(??):??, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Fisheries Research, 236(??):??, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Fisheries Research, 240(??):??, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Fisheries Research, 244(??):??, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Fisheries Research, 233 (??):Article 105791, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303088.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Fisheries Research, 234 (??):Article 105835, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303520.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Fisheries Research, 235 (??):Article 105861, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303787.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Fisheries Research, 236(??): Article 105889, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783621000175.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Fisheries Research, 237(??): Article 105911, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783621000394.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Fisheries Research, 238(??): Article 105931, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578362100059X.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Fisheries Research, 239(??): Article 105979, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783621001077.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Fisheries Research, 240 (??):Article 106010, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621001387.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Fisheries Research, 241 (??):Article 106025, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621001533.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Fisheries Research, 242 (??):Article 106080, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621002083.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. Fisheries Research, 243 (??):Article 106112, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362100240X.

Anonymous:2021:EBl

[Ano21o] Anonymous. Editorial Board. Fisheries Research, 244 (??):Article 106144, December 2021. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621002721.

Anonymous:2021:F

[Ano21p] Anonymous. February 2021. Fisheries Research, 234(??):??, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. January 2021. Fisheries Research, 233(??):??, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Jc

[Ano21r] Anonymous. July 2021. Fisheries Research, 239(??):??, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Fisheries Research, 238(??):??, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021. Fisheries Research, 235(??):??, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021. Fisheries Research, 237(??):??, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:N

[Ano21v] Anonymous. November 2021. Fisheries Research, 243(??):??, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021. Fisheries Research, 242(??):??, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2021:S

[Ano21x] Anonymous. September 2021. Fisheries Research, 241(??):??, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Fisheries Research, 248(??):??, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Fisheries Research, 252(??):??, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Fisheries Research, 256(??):??, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Fisheries Research, 245 (??):Article 106161, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621002897.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Fisheries Research, 246 (??):Article 106185, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0165783621003131.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Fisheries Research, 247 (??):Article 106215, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362100343X.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Fisheries Research, 248(??): Article 106244, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622000212.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Fisheries Research, 249(??): Article 106282, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622000595.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Fisheries Research, 250(??): Article 106312, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622000893.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Fisheries Research, 251(??): Article 106347, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622001242.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Fisheries Research, 252 (??):Article 106376, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622001539.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Fisheries Research, 253 (??):Article 106399, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362200176X.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Fisheries Research, 254 (??):Article 106446, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622002235.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Fisheries Research, 255 (??):Article 106471, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362200248X.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Fisheries Research, 256 (??):Article 106506, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622002831.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Fisheries Research, 246(??):??, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:PF

[Ano22q] Anonymous. In progress (February 2022). Fisheries Research, 246(??):??, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Ja

[Ano22r] Anonymous. January 2022. Fisheries Research, 245(??):??, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Jc

[Ano22s] Anonymous. July 2022. Fisheries Research, 251(??):??, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Jb

[Ano22t] Anonymous. June 2022. Fisheries Research, 250(??):??, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Ma

[Ano22u] Anonymous. March 2022. Fisheries Research, 247(??):??, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:Mb

[Ano22v] Anonymous. May 2022. Fisheries Research, 249(??):??, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:N

[Ano22w] Anonymous. November 2022. Fisheries Research, 255(??):??, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:O

[Ano22x] Anonymous. October 2022. Fisheries Research, 254(??):??, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2022:S

[Ano22y] Anonymous. September 2022. Fisheries Research, 253(??):??, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Fisheries Research, 260(??):??, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Fisheries Research, 257 (??):Article 106528, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622003058.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Fisheries Research, 258 (??):Article 106570, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622003472.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Fisheries Research, 259 (??):Article 106600, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622003770.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Fisheries Research, 260(??): Article 106630, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783623000231.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Fisheries Research, 261(??): Article 106656, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783623000498.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Fisheries Research, 262(??): Article 106700, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783623000930.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Fisheries Research, 263(??): ??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001200.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Fisheries Research, 264(??): ??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001455.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001625.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Fisheries Research, 266(??): ??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001984.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002308.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362300262X.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Fisheries Research, 258(??):??, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Fisheries Research, 257(??):??, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Fisheries Research, 262(??):??, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Fisheries Research, 259(??):??, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Fisheries Research, 261(??):??, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:EBa

[Ano24d] Anonymous. Editorial Board. Fisheries Research, 269(??): ??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002813.

Anonymous:2024:EBb

[Ano24e] Anonymous. Editorial Board. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362300303X.

Anonymous:2024:EBc

[Ano24f] Anonymous. Editorial Board. Fisheries Research, 271(??): ??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000079.

Anonymous:2024:EBd

[Ano24g] Anonymous. Editorial Board. Fisheries Research, 272(??): ??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000298.

Anonymous:2024:EBe

[Ano24h] Anonymous. Editorial Board. Fisheries Research, 273(??): ??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000511.

Anonymous:2024:EBf

[Ano24i] Anonymous. Editorial Board. Fisheries Research, 274(??): ??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400081X.

Anonymous:2024:EBg

[Ano24j] Anonymous. Editorial Board. Fisheries Research, 275(??): ??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001140.

Anonymous:2024:EBh

[Ano24k] Anonymous. Editorial Board. Fisheries Research, 276(??): ??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001401.

Anonymous:2024:EBi

[Ano24l] Anonymous. Editorial Board. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001541.

Anonymous:2024:EBj

[Ano24m] Anonymous. Editorial Board. Fisheries Research, 278(??): ??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001930.

Anonymous:2024:EBk

[Ano24n] Anonymous. Editorial Board. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002339.

Anonymous:2024:EBl

[Ano24o] Anonymous. Editorial Board. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002480.

Anonymous:2024:F

[Ano24p] Anonymous. February 2024. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Ja

[Ano24q] Anonymous. January 2024. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Jc

[Ano24r] Anonymous. July 2024. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Jb

[Ano24s] Anonymous. June 2024. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Ma

[Ano24t] Anonymous. March 2024. Fisheries Research, 271(??):??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:Mb

[Ano24u] Anonymous. May 2024. Fisheries Research, 273(??):??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:N

[Ano24v] Anonymous. November 2024. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:O

[Ano24w] Anonymous. October 2024. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2024:S

[Ano24x] Anonymous. September 2024. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2025:EBa

[Ano25a] Anonymous. Editorial Board. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000153.

Anonymous:2025:EBb

[Ano25b] Anonymous. Editorial Board. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000268.

Anonymous:2025:EBc

[Ano25c] Anonymous. Editorial Board. Fisheries Research, 283(??): ??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000517.

Anonymous:2025:F

[Ano25d] Anonymous. February 2025. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2025:Ja

[Ano25e] Anonymous. January 2025. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2025:Ma

[Ano25f] Anonymous. March 2025. Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Asiedu:2022:PFP

[AOA^{+}22] Berchie Asiedu, Paulina Okpei, Samuel K. K. Amponsh, Pierre Failler, Bara Deme, and Rashid Ussif Sumaila. The people’s fishery in perspective: Current analysis of the small pelagic fishery value chain of Ghana. Fisheries Research, 254(??):Article 106426, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200203X.

Andrialovanirina:2020:PMM

[APB^{+}20] Nicolas Andrialovanirina, Dominique Ponton, Faustinato Behivoke, Jamal Mahafina, and Marc Léopold. A powerful method for measuring fish size of small-scale fishery catches using ImageJ. Fisheries Research, 223(??):Article 105425, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619302802.

Aguirre-Pabon:2022:PSL

[APBN22] Juan C. Aguirre-Pabon, Gilberto Orozco Berdugo, and Juan C. Narváez. Population structure and low genetic diversity in the threatened lebranche Mugil liza in the Colombian Caribbean. Fisheries Research, 256(??):Article 106485, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002624.

Akia:2022:MFS

[APGG22] Sosthène Akia, Ilan Perez, Loreleï Guéry, and Daniel Gaertner. Misidentification of free school tunas in the AOTTP database: Issues to identify fish-attractive seamounts. Fisheries Research, 251(??):Article 106324, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001011.

Aponte:2025:EIS

[Apo25] Fabian Rocha Aponte. Economic interdependencies and sectoral impacts: an input-output analysis of the Norwegian fisheries and aquaculture sector amidst COVID-19. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003114.

Arlinghaus:2023:SCN

[ARD^{+}23] Robert Arlinghaus, Timo Rittweg, Félicie Dhellemmes, Dieter Koemle, Rob van Gemert, Hendrik Schubert, Dominique Niessner, Sören Möller, Jan Droll, René Friedland, Wolf-Christian Lewin, Malte Dorow, Linda Westphal, Elias Ehrlich, Harry V. Strehlow, Marc Simon Weltersbach, Phillip Roser, Marlon Braun, Fritz Feldhege, and Helmut Winkler. A synthesis of a coastal northern pike (Esox lucius) fishery and its social-ecological environment in the southern Baltic Sea: Implications for the management of mixed commercialrecreational fisheries. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000565.

Abass:2024:MTW

[ASB^{+}24] Zaheer Abass, Tasaduq Hussain Shah, Farooz Ahmed Bhat, Karankumar Ramteke, Aadil Hussain Magloo, Ishfaq Hamid, Rinkesh Nemichand Wanjari, and Iyyappan Somasundharam. The mahseer: the tiger of water — an angler’s delight in the Himalayas and the undisputed king of sport fishing. Fisheries Research, 279(??):??, November 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400211X.

Azevedo:2022:WMC

[ASD^{+}22] Laíse S. Azevedo, Darlan J. B. Simith, Kelly G. Duarte, Marcus E. B. Fernandes, and Evaldo M. Silva. When are mangrove crabs fat? Seasonal and sexual variation in the fatty acid composition of edible crab species Ucides cordatus from the Brazilian Amazon coast. Fisheries Research, 251(??):Article 106330, July 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001072.

Annear:2024:ESE

[ASDW24] Andrew R. Annear, Martin A. Simonson, Philip M. Dixon, and Michael J. Weber. Effects of sampling effort and data sources on precision of common carp and bigmouth buffalo population estimates. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002078.

Ashrafi:2021:IQR

[ASE21] Tannaz Alizadeh Ashrafi, Shaheen Syed, and Arne Eide. Individual quotas and revenue risk of fishing portfolio in the trawl fishery. Fisheries Research, 241(??):Article 105990, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001181.

Alrabeei:2021:STP

[ASGG21] Salah Alrabeei, Sam Subbey, Sofie Gundersen, and Harald Gjøsæter. Spatial and temporal patterns of capelin (Mallotus villosus) spawning sites in the Barents Sea. Fisheries Research, 244(??):Article 106117, December 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002459.

Ashida:2020:CST

[Ash20a] Hiroshi Ashida. Corrigendum to “Spatial and temporal differences in the reproductive traits of skipjack tuna Katsuwonus pelamis between the subtropical and temperate western Pacific Ocean” [Fish. Res. 221 (January) (2020) 105352]. Fisheries Research, 227(??):Article 105566, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300837. See [Ash20b].

Ashida:2020:STD

[Ash20b] Hiroshi Ashida. Spatial and temporal differences in the reproductive traits of skipjack tuna Katsuwonus pelamis between the subtropical and temperate western Pacific Ocean. Fisheries Research, 221(??):Article 105352, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302073. See corrigen-

dum [Ash20a].

Amorim:2020:LFD

[ASJ^{+}20] Patrícia Amorim, Pedro Sousa, Ernesto Jardim, Manuela Azevedo, and Gui M. Menezes. Length-frequency data approaches to evaluate snapper and grouper fisheries in the Java Sea, Indonesia. Fisheries Research, 229(??):Article 105576, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362030093X.

Ault:2022:LBR

[ASJ^{+}22] Jerald S. Ault, Steven G. Smith, Matthew W. Johnson, Laura Jay W. Grove, James A. Bohnsack, Gerard T. Di-Nardo, Caroline McLaughlin, Nelson M. Ehrhardt, Vanessa McDonough, Michael P. Seki, Steven L. Miller, Jiangang Luo, Jeremiah Blondeau, Michael P. Crosby, Glenn Simpson, Mark E. Monaco, Clayton G. Pollock, Michael W. Feeley, and Alejandro Acosta. Length-based risk analysis of management options for the southern Florida, USA multispecies coral reef fish fishery. Fisheries Research, 249(??):Article 106210, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003386.

Ashida:2022:RDP

[ASO^{+}22] Hiroshi Ashida, Tamaki Shimose, Yumi Okochi, Yosuke Tanaka, and Sho Tanaka. Reproductive dynamics of Pacific bluefin tuna (Thunnus orientalis) off the Nansei Islands, southern Japan. Fisheries Research, 249(??):Article 106256, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000339.

Araya-Schmidt:2021:IPR

[ASWS^{+}21] T. Araya-Schmidt, P. D. Winger, M. R. Santos, K. Moret, H. DeLouche, G. Legge, and S. M. Bayse. Investigating the performance of a roller footgear in the offshore shrimp fishery of Eastern Canada using underwater video. Fisheries Research, 240(??):Article 105968, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000965.

Albertsen:2020:MBE

[AT20] Christoffer Moesgaard Albertsen and Vanessa Trijoulet. Model-based estimates of reference points in an agebased state-space stock assessment model. Fisheries Research, 230(??):Article 105618, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301351.

Aoki:2024:REA

[ATA^{+}24] Yoshinori Aoki, Fumiya Tanaka, Akiko Aoki, Shinpei Ohashi, and Hidetada Kiyofuji. Re-evaluating age estimation using daily increments on otoliths of the skipjack tuna Katsuwonus pelamis. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002485.

Aldrin:2020:SDM

[ATAS20] M. Aldrin, I. F. Tvete, S. Aanes, and S. Subbey. The specification of the data model part in the SAM model matters. Fisheries Research, 229(??):Article 105585, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301028.

Aldanondo:2025:EGS

[AUC25] Naroa Aldanondo, Agurtzane Urtizberea, and Unai Cotano. Evidence of growth-selective mortality of European anchovy larvae in the Bay of Biscay based on otolith microstructure analysis. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003230.

Ahti:2022:TPF

[AUHK22] Pauliina A. Ahti, Silva Uusi-Heikkilä, and Anna Kuparinen. Are there plenty of fish in the sea? How life history traits affect the eco-evolutionary consequences of population oscillations. Fisheries Research, 254(??):Article 106409, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001862.

Asai:2021:BST

[AUM21] Saki Asai, Keiichi Uchida, and Yoshinori Miyamoto. Behavior of skipjack tuna (Katsuwonus pelamis) during sinking of a surface-type fish aggregating device. Fisheries Research, 239(??):Article 105925, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000539.

Arce-Valdes:2023:NEF

[AVACA^{+}23] Luis Rodrigo Arce-Valdés, Alicia Abadía-Cardoso, Maria Clara Arteaga, Laura Victoria Peñaranda-Gonzalez, Gorgonio Ruiz-Campos, and Luis Manuel Enríquez-Paredes. No effects of fishery collapse on the genetic diversity of the Gulf of California Corvina, Cynoscion othonopterus (Perciformes: Sciaenidae). Fisheries Research, 261(??):Article 106608, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000012.

Aoufi:2023:PAA

[AVB^{+}23] Sofya Aoufi, Maria Valls, Oona C. Bienentreu, Ines R. Pereira, Yansong Huang, Baptiste Mourre, and Miguel Cabanellas-Reboredo. Preferential aggregation areas of Seriola dumerili: Estimation of environmental influences and insights towards sustainable exploitation. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623001868.

Andre:2022:BCB

[AVCA22] Laure Vaitiare André, Simon Van Wynsberge, Mireille Chinain, and Serge Andréfouët. Benefits of collaboration between indigenous fishery management and data-driven spatial planning approaches: the case of a Polynesian traditional design (rāhui). Fisheries Research, 256(??):Article 106475, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002521.

Aguirre-Villasenor:2022:ISM

[AVMBEB22] Hugo Aguirre-Villaseñor, Enrique Morales-Bojórquez, and Elaine Espino-Barr. Implementation of sigmoidal models with different functional forms to estimate length at 50% maturity: a case study of the Pacific red snapper Lutjanus peru. Fisheries Research, 248(??):Article 106204, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003325.

Allison:2023:CSM

[AWC^{+}23] C. Allison, A. C. Winkler, A.-R. Childs, C. Muller, and W. M. Potts. Can social media platforms be used to foster improved environmental behaviour in recreational fisheries? Fisheries Research, 258(??):Article 106544, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003216.

Aalbers:2021:BTS

[AWVS21] Scott A. Aalbers, Michael Wang, Charles Villafana, and Chugey A. Sepulveda. Bigeye thresher shark Alopias superciliosus movements and post-release survivorship following capture on linked buoy gear. Fisheries Research, 236(??):Article 105857, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362030374X.

Afrifa-Yamoah:2021:TAB

[AYTM21] Ebenezer Afrifa-Yamoah, Stephen M. Taylor, and Ute Mueller. Trade-off assessments between reading cost and accuracy measures for digital camera monitoring of recreational boating effort. Fisheries Research, 233(??):Article 105757, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302745.

Butler:2022:IEA

[BAC^{+}22] Edward C. Butler, Nicky K. Arkert, Amber-Robyn Childs, Brett A. Pringle, Michael R. Skeeles, Ryan M. Foster, Matthew W. Farthing, Alexander C. Winkler, and Warren M.

Potts. Incorporating estuarine-angler behaviour and delayed blood sampling into the rapid assessment of catch-and release angling on the iconic dusky kob Argyrosomus japonicus. Fisheries Research, 253(??):Article 106364, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001412.

Beeken:2023:DST

[BAF23] Nicolette S. Beeken, Joel Anderson, and Mark Fisher. Determining spatiotemporal trends in hatch and metamorphosis timing of young-of-year southern flounder Paralichthys lethostigma in Texas bays. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001157.

Braham:2024:ORS

[BAJB^{+}24] Cheikh-Baye Braham, Mohamed Ahmed-Jeyid, Jilali Bensbai, Fambye Ngoum, Ad Corten, and Jo Gascoigne. Overexploitation of round sardinella may lead to the collapse of flat sardinella: What lessons can be drawn for shared stocks. Fisheries Research, 269(??):??, January 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002667.

Britton:2023:RSC

[BALBC23] J. R. Britton, D. Andreou, M. Lopez-Bejar, and A. Carbajal. Relationships of scale cortisol content suggest stress resilience in freshwater fish vulnerable to catch-and-release angling in recreational fisheries. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001698.

Boenish:2021:BAT

[BaLK^{+}21] Robert Boenish, Bai an Lin, Jacob P. Kritzer, Michael J. Wilberg, Chang chun Shen, Yan Jiang, and Min Liu. A bioeconomic approach towards improved fishery management of Monomia haanii in the southern Taiwan Strait, China. Fisheries Research, 240(??):Article 105969, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000977.

Baker:2024:MSM

[BAM^{+}24] Krista D. Baker, Sean C. Anderson, Darrell R. J. Mullowney, Wojciech Walkusz, and Katherine R. Skanes. Moving away from a scale mismatch: Spatiotemporal modelling of striped shrimp (Pandalus montagui) density in Canada’s subarctic. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002916.

Belleggia:2023:PCD

[BÁP^{+}23] Mauro Belleggia, Cecilia Diamela Álvarez, Emiliano Pisani, Mariana Descalzo, and Eliana Zuazquita. Prey contribution to the diet of pink cusk-eel Genypterus blacodes (Forster, 1801) revealed by stomach content and stable isotopic analyses in the southwestern Atlantic. Fisheries Research, 262(??):Article 106660, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300053X.

Benfer:2024:DLL

[BAW^{+}24] Caroline Benfer, Eric Annis, Jesica Waller, Joshua T. Carloni, Kathleen Reardon, LeAnn Whitney, Rachel Lasley-Rasher, and Heidi Henninger. Distribution of lobster larvae, Homarus americanus, and zooplankton prey in the Gulf of Maine and Georges Bank stock area. Fisheries Research, 278(??):??, October 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001851.

Burchert:2024:SCD

[BAYR^{+}24] Shannon J. Burchert, Ebenezer Afrifa-Yamoah, Karina L. Ryan, Ute Mueller, and Glenn A. Hyndes. Spatial characterisation and drivers of catch and effort in highly specialised recreational pelagic fisheries. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001358.

Brauer:2020:SPR

[BBC^{+}20] Meike Brauer, Jane W. Behrens, Mads Christoffersen, Grethe Hyldig, Charlotte Jacobsen, Katla H. Björnsdottir, and Mikael van Deurs. Seasonal patterns in round goby (Neogobius melanostromus) catch rates, catch composition, and dietary quality. Fisheries Research, 222(??):Article 105412, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578361930267X.

Baran:2021:NWI

[BBČ^{+}21] Roman Baran, Petr Blabolil, Martin Čech, Vladislav Draštík, Jaroslava Frouzová, Michaela Holubová, Tomáš Jůza, Ievgen Koliada, Milan Muška, Jiří Peterka, Marie Prchalová, Milan Říha, Zuzana Sajdlová, Marek Šmejkal, Michal Tušer, Lukáš Vejřík, and Jan Kubečka. New way to investigate fish density and distribution in the shallowest layers of the open water. Fisheries Research, 238(??):Article 105907, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000357.

Bell:2025:APB

[BBC^{+}25] Craig Bell, Paul A. Butcher, Keith Cox, Stephen Morris, Curtis Champion, and Brendan Kelaher. Assessing the potential of bait reuse in a large-scale SMART drumline program. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002406.

Berger:2024:SSF

[BBG^{+}24] Aaron M. Berger, Caren Barceló, Daniel R. Goethel, Simon D. Hoyle, Patrick D. Lynch, Jeremy McKenzie, Alistair Dunn, André E. Punt, Richard D. Methot, John Hampton, Clay E. Porch, Richard McGarvey, James T. Thorson, Z. Teresa A’mar, Jonathan J. Deroba, Bjarki Tór Elvarsson, Steven J. Holmes, Daniel Howell, Brian J. Langseth, Craig Marsh, Mark N. Maunder, Sophie Mormede, and Scott Rasmussen. Synthesizing the spatial functionality of contemporary stock assessment software to identify future needs for next generation assessment platforms.

Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000729.

Blackman:2025:RAG

[BBHF25] Chloe Blackman, Lynda Blackadder, Elizabeth M. Harper, and Clive Fox. Re-assessing the growth parameters for the razor clam, Ensis siliqua, from Scottish electrofishing grounds using external shell marks and oxygen isotope validation. Fisheries Research, 283(??):??, March 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000311.

Barnes:2022:DAR

[BBJ22] Thomas C. Barnes, Matt K. Broadhurst, and Daniel D. Johnson. Disparity among recommended and adopted escape-gap designs and their utility for improving selection in an Australian portunid trap fishery. Fisheries Research, 248(??):Article 106219, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003477.

Becker:2024:IBW

[BBM^{+}24] Alistair Becker, Paul A. Butcher, Stephen Morris, Curtis Champion, Victor M. Peddemors, Michael B. Lowry, and Matthew D. Taylor. Interactions between White Sharks (Carcharodon carcharias) and artificial reefs along the east coast of Australia. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000250.

Bessell-Browne:2022:EIL

[BBPT^{+}22] Pia Bessell-Browne, Andre E. Punt, Geoffrey N. Tuck, Jemery Day, Neil Klaer, and Andrew Penney. The effects of implementing a ‘dynamic B_{0}’ harvest control rule in Australia’s southern and eastern scalefish and shark fishery. Fisheries Research, 252(??):Article 106306, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000832.

Bessell-Browne:2024:MSE

[BBPT^{+}24] Pia Bessell-Browne, André E. Punt, Geoffrey N. Tuck, Paul Burch, and Andrew Penney. Management strategy evaluation of static and dynamic harvest control rules under longterm changes in stock productivity: a case study from the SESSF. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000365.

Butler:2022:EPU

[BBR^{+}22] Casey B. Butler, Jack Butler, Gabrielle F. Renchen, Emily Hutchinson, and Thomas R. Matthews. Exploring the potential use of escape gaps in the Florida spiny lobster, Panulirus argus, fishery. Fisheries Research, 254(??):Article 106421, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622001989.

Butler:2024:RSL

[BBSM24] Casey B. Butler, Jack Butler, William C. Sharp, and Thomas R. Matthews. Refining spiny lobster (Panulirus argus) escape and mortality estimates in Florida’s lobster trap fishery using long-term video. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002030.

Broadhurst:2020:MDM

[BC20] Matt K. Broadhurst and Brian R. Cullis. Mitigating the discard mortality of non-target, threatened elasmobranchs in bather-protection gillnets. Fisheries Research, 222(??):Article 105435, February 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302905.

Blabolil:2021:LMB

[BČD^{+}21] Petr Blabolil, Martin Čech, Vladislav Draštík, Michaela Holubová, Luboš Kočvara, Jan Kubečka, Milan Muška, Marie Prchalová, Milan Říha, Zuzana Sajdlová, Marek Šmejkal, Michal Tušer, Mojmír Vašek, Lukáš Vejřík, Ivana Vejříková, Jiří Peterka, and Tomáš Jůza. Less is more — basic quantitative indices for fish can be achieved with reduced gillnet sampling. Fisheries Research, 240(??):Article 105983, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001119.

Burgaard:2023:UHM

[BCF^{+}23] Karen B. Burgaard, Stefan Carstensen, David R. Fuhrman, Camille Saurel, and Finbarr G. O’Neill. Using hydrodynamics to modify fishing performance of a demersal fishing gear. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002242.

Bensebaini:2025:ISI

[BCG^{+}25] Cyria Meriem Bensebaini, Grégoire Certain, Sophie Gourguet, Olivier Thébaud, Tarek Hattab, Norbert Billet, Angélique Jadaud, and Jean-Marc Fromentin. Identifying statistical interaction networks in marine communities using multivariate time series analysis: an application in the Gulf of Lions. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002418.

Butler:2021:DCA

[BCM^{+}21] Edward C. Butler, Amber-Robyn Childs, Marianne V. Milner, Matthew W. Farthing, Murray I. Duncan, Alexander C. Winkler, and Warren M. Potts. Do contemporary agegrowth models overlook life-history complexities in protandrous fishes? A case study on the large protandrous polynemid, the giant African threadfin Polydactylus quadrifilis. Fisheries Research, 233(??):Article 105770, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302873.

Burch:2023:IMM

[BCOBB^{+}23] Paul Burch, Sandra Curin-Osorio, Pia Bessell-Browne, Geoffrey N. Tuck, Malcolm Haddon, Nils C. Krueck, and André E. Punt. Implications of the maximum modelled age on the estimation of natural mortality when using a meta-analytic prior: the example of eastern Australian orange roughy (Hoplostethus atlanticus). Fisheries Research, 258(??):Article 106534, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003113.

Barrett:2022:IQS

[BCPH22] C. J. Barrett, A. Cook, J. K. Pinnegar, and K. Hyder. Importance of quantifying spatiotemporal biomass removal of recreationally caught UK squids and cuttlefish. Fisheries Research, 252(??):Article 106332, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001096.

Bruno:2021:WMF

[BCRI21] Claudia Bruno, Claudio F. Cornejo, Rodrigo Riera, and Christian M. Ibáñez. What is on the menu? Feeding, consumption and cannibalism in exploited stocks of the jumbo squid Dosidicus gigas in south-central Chile. Fisheries Research, 233(??):Article 105722, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302393.

Bauerlien:2022:LCL

[BCS^{+}22] Cory J. Bauerlien, Derek P. Crane, Scott Smith, George Palmer, Tyler Young, Daniel B. Goetz, Jeff Hansbarger, and Kyle Hartman. Low catchability limits the effect of warmwater catch-and-release mortality on muskellunge. Fisheries Research, 254(??):Article 106434, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002119.

Begue:2020:PPI

[BCSM20] Michel Bègue, Eric Clua, Gilles Siu, and Carl Meyer. Prevalence, persistence and impacts of residual fishing hooks on tiger sharks. Fisheries Research, 224(??):Article 105462, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619303170.

Baidai:2020:MLC

[BDA^{+}20] Y. Baidai, L. Dagorn, M. J. Amande, D. Gaertner, and M. Capello. Machine learning for characterizing tropical tuna aggregations under Drifting Fish Aggregating Devices (DFADs) from commercial echosounder buoys data. Fisheries Research, 229(??):Article 105613, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301302.

Benezech:2024:TNM

[BDA^{+}24] Mathilde Bénézech, Hilaire Drouineau, Anthony Acou, Agnès Bardonnet, Clarisse Boulenger, Patrick Lambert, Etienne Prévost, and Laurent Beaulaton. Testing novel methods for short-term forecasting of European glass eel recruitment. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003089.

Beveren:2020:EHC

[BDM^{+}20] Elisabeth Van Beveren, Daniel E. Duplisea, Julie R. Marentette, Andrew Smith, and Martin Castonguay. An example of how catch uncertainty hinders effective stock management and rebuilding. Fisheries Research, 224(??):Article 105473, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619303285.

Bastardie:2020:RFI

[BDR^{+}20] Francois Bastardie, Jules Danto, Marie-Christine Rufener, Daniel van Denderen, Ole R. Eigaard, Grete E. Dinesen, and J. Rasmus Nielsen. Reducing fisheries impacts on the seafloor: a bio-economic evaluation of policy strategies for improving sustainability in the Baltic Sea. Fisheries Research, 230(??):Article 105681, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301983.

Beacham:2021:PBT

[Bea21] Terry D. Beacham. Parentage-based tagging combined with genetic stock identification is a cost-effective and viable replacement for coded-wire tagging in large-scale assessments of Canadian salmon fisheries. Fisheries Research, 239(??):Article 105920, July 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000485.

Bechard:2020:ELF

[Béc20] Andrew Béchard. Economics losses to fishery and seafood related businesses during harmful algal blooms. Fisheries Research, 230(??):Article 105678, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301958.

Briones-Fourzan:2025:IHE

[BF25] Patricia Briones-Fourzán. Influence of habitat enhancement and habitat degradation on spiny lobster populations in the Caribbean: an overview. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002571.

Barker:2021:DRA

[BFA^{+}21] Amanda M. Barker, Bryan S. Frazier, Douglas H. Adams, Christine N. Bedore, Carolyn N. Belcher, William B. Driggers, Ashley S. Galloway, James Gelsleichter, R. Dean Grubbs, Eric A. Reyier, and David S. Portnoy. Distribution and relative abundance of scalloped (Sphyrna lewini) and Carolina (S. gilberti) hammerheads in the western North Atlantic Ocean. Fisheries Research, 242(??):Article 106039, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001673.

Benzer:2022:ISM

[BGBM22] Semra Benzer, Farid Hassanbaki Garabaghi, Recep Benzer, and Homay Danaei Mehr. Investigation of some machine learning algorithms in fish age classification. Fisheries Research, 245(??):Article 106151, January 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002794.

Bohaboy:2022:SFA

[BGCCP22] Erin C. Bohaboy, Daniel R. Goethel, Shannon L. Cass-Calay, and William F. Patterson. A simulation framework to assess management trade-offs associated with recreational harvest slots, discard mortality reduction, and bycatch accountability in a multi-sector fishery. Fisheries Research, 250(??):Article 106268, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000455.

Bacheler:2022:USV

[BGG^{+}22] Nathan M. Bacheler, Kevan C. Gregalis, Zachary D. Gillum, Erin P. Pickett, Christina M. Schobernd, Zebulon H. Schobernd, and Bradford Z. Teer. Using stationary video data to infer relative abundance and distribution of four Seriola species along the southeast United States Atlantic coast. Fisheries Research, 249(??):Article 106238, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000157.

Baiandina:2022:MRP

[BGK22] Iuliia Baiandina, Vitaly Giragosov, and Antonina Khanaychenko. Male reproductive potential in the Black Sea turbot (Scophthalmus maximus) spawning populations. Fisheries Research, 253(??):Article 106367, September 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001448.

Basurko:2022:FCF

[BGL^{+}22] Oihane C. Basurko, Gorka Gabiña, Jon Lopez, Igor Granado, Hilario Murua, Jose A. Fernandes, Iñigo Krug, Jon Ruiz, and Zigor Uriondo. Fuel consumption of free-swimming school versus FAD strategies in tropical tuna purse seine fishing. Fisheries Research, 245(??):Article 106139, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002678.

Boens:2021:DGS

[BGLP21] Andy Boëns, Patrick Grellier, Christophe Lebigre, and Pierre Petitgas. Determinants of growth and selective mortality in anchovy and sardine in the Bay of Biscay. Fisheries Research, 239(??):Article 105947, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000758.

Birch:2023:HIH

[BGM^{+}23] Samantha F. Birch, Stephen D. Gregory, David L. Maxwell, Marieke Desender, and Thomas L. Catchpole. How an illuminated headline affects catches and species separation in a Celtic Sea mixed demersal trawl fishery. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002254.

Bruch:2023:CRE

[BH23] Ronald M. Bruch and Tim J. Haxton. Cost and relative effectiveness of lake sturgeon passage systems in the US and Canada. Fisheries Research, 257(??):Article 106510, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002879.

Birdsong:2022:DRC

[BHB^{+}22] Max Birdsong, Len M. Hunt, Ben Beardmore, Malte Dorow, Thilo Pagel, and Robert Arlinghaus. Does the relevance of catch for angler satisfaction vary with social-ecological context? A study involving angler cultures from West and East Germany. Fisheries Research, 254(??):Article 106414, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001916.

Benson:2024:FTN

[BHB24] Irina M. Benson, Thomas E. Helser, and Beverly K. Barnett. Fourier transform near infrared spectroscopy of otoliths coupled with deep learning improves age prediction for longlived northern rockfish. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001802.

Bunch:2023:EDA

[BHD^{+}23] Aaron J. Bunch, James P. Henne, Dennis R. DeVries, Russell A. Wright, David L. Smith, and Troy M. Farmer. Evaluation of double acoustic tagging techniques to track American shad Alosa sapidissima movements at multiple spatial scales. Fisheries Research, 261(??):Article 106636, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000292.

Bradshaw:2024:ADE

[BHG^{+}24a] Stephen Bradshaw, Klaas Hartmann, Caleb Gardner, Katherine A. Cresswell, and Denham Parker. Appendage damage effects on southern rock lobster growth and mortality. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002170.

Bradshaw:2024:WOM

[BHG^{+}24b] Stephen Bradshaw, Klaas Hartmann, Caleb Gardner, Katie Cresswell, and Denham Parker. Are we overthinking management zoning for the Tasmanian southern rock lobster fishery? A size of maturity approach. Fisheries Research, 274(??): ??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000614.

Bergstad:2021:RGM

[BHH21] Odd Aksel Bergstad, Hege Øverbø Hansen, and Alf Harbitz. Roughhead grenadier (Macrourus berglax) on the shelf edge of the northeastern Norwegian Sea, 1997–2020: Distribution, abundance, size and age structure, growth. Fisheries Research, 240(??):Article 105957, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000850.

Boyle:2022:UPA

[BHNP22] Kelly S. Boyle, Crystal L. Hightower, T. Reid Nelson, and Sean P. Powers. Use of passive acoustic monitoring to estimate fishing effort on artificial reefs in Alabama during the recreational red snapper fishing season. Fisheries Research, 249(??):Article 106262, May 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200039X.

Bezerra:2021:STP

[BHST^{+}21] Inajara Marques Bezerra, Mauricio Hostim-Silva, Jessyca Luana Silva Teixeira, Carlos Werner Hackradt, Fabiana C. Félix-Hackradt, and Alexandre Schiavetti. Spatial and temporal patterns of spawning aggregations of fish from the Epinephelidae and Lutjanidae families: an analysis by the local ecological knowledge of fishermen in the Tropical Southwestern Atlantic. Fisheries Research, 239(??):Article 105937, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000655.

Ben-Hasan:2024:FIS

[BHVB^{+}24] Abdulrahman Ben-Hasan, Arezoo Vahabnezhad, John A. Burt, Tariq Alrushaid, and Carl J. Walters. Fishery implications of smaller asymptotic body size: Insights from fish in an extreme environment. Fisheries Research, 271(??):??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623003119.

Bak-Jensen:2023:CSM

[BJHS^{+}23] Zita Bak-Jensen, Bent Herrmann, Juan Santos, Valentina Melli, Daniel Stepputtis, and Jordan P. Feekings. The capability of square-meshes and fixed-shape meshes to control codend size selection. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000978.

Briggs:2024:PGP

[BJK24] Jeremy Briggs, Danielle Johnston, and W. Jason Kennington. Population genomics provides evidence of interspecific hybridisation and population structure in the blue-swimmer crab (Portunus armatus) along the Western Australian coastline. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002862.

Beacham:2022:AMS

[BJS^{+}22a] Terry D. Beacham, Kim Jonsen, Ben J. G. Sutherland, Cheryl Lynch, and Eric B. Rondeau. Assessment of mixed-stock fisheries and hatchery broodstocks for Coho salmon in British Columbia, Canada via parentage-based tagging and genetic stock identification. Fisheries Research, 245(??):Article 106136, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002642.

Beacham:2022:PBT

[BJS^{+}22b] Terry D. Beacham, Kim Jonsen, Ben J. G. Sutherland, Brock Ramshaw, and Eric B. Rondeau. Parentage-based tagging and genetic stock identification applied to assessment of mixed-stock fisheries and hatchery broodstocks for Chinook salmon in British Columbia, Canada. Fisheries Research, 253(??):Article 106369, September 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001461.

Bowerman:2021:EST

[BKC21] Tracy E. Bowerman, Matthew L. Keefer, and Christopher C. Caudill. Elevated stream temperature, origin, and individual size influence Chinook salmon prespawn mortality across the Columbia River Basin. Fisheries Research, 237(??):Article 105874, May 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000023.

Brogan:2021:BPR

[BKHA21] John D. Brogan, Craig R. Kastelle, Thomas E. Helser, and Delsa M. Anderl. Bomb-produced radiocarbon age validation of Greenland halibut (Reinhardtius hippoglossoides) suggests a new maximum longevity. Fisheries Research, 241(??):Article 106000, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001284.

Broadhurst:2023:NEG

[BKM23a] Matt K. Broadhurst, Ian A. Knuckey, and Russell B. Millar. No effects of ground-gear diameter on the performance of a south-eastern Australian fish trawl. Fisheries Research, 263 (??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000887.

Broadhurst:2023:UHS

[BKM23b] Matt K. Broadhurst, Ian A. Knuckey, and Russell B. Millar. Using a horizontal-separator panel in an Australian fish trawl to quantify species-specific changes in vertical orientations during capture. Fisheries Research, 261(??):Article 106618, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000115.

Bronnmann:2023:WPH

[BKM^{+}23c] Julia Bronnmann, Dieter Koemle, Jürgen Meyerhoff, Marc Simon Weltersbach, Harry V. Strehlow, and Robert Arlinghaus. Willingness to pay for harvest regulations and catch outcomes in recreational fisheries: a stated preference study of German cod anglers. Fisheries Research, 259(??):Article 106536, March 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003137.

Brand:2022:AEL

[BKR^{+}22] Nathan R. Brand, Emma M. Knoebel, Riley M. Ross, Robb D. VanPutte, and Michael J. Louison. Air exposure leads to short term reflex impairment, but does not impact blood physiology, in angled channel catfish Ictalurus punctatus. Fisheries Research, 256(??):Article 106490, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002673.

Benoit:2020:DDM

[BKT^{+}20] Hugues P. Benoît, Jeff Kneebone, Sean R. Tracey, Diego Bernal, Klaas Hartmann, and Walt Golet. Distinguishing discard mortality from natural mortality in field experiments based on electronic tagging. Fisheries Research, 230(??):Article 105642, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301594.

Barbosa:2021:FST

[BLC^{+}21] Moysés C. Barbosa, Osmar J. Luiz, Cesar A. M. M. Cordeiro, Vinicius J. Giglio, and Carlos E. L. Ferreira. Fish and spearfisher traits contributing to catch composition. Fisheries Research, 241(??):Article 105988, September 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001168.

Bieber:2022:PRL

[BLC^{+}22] John F. Bieber, Luc LaRochelle, Steven J. Cooke, Cory D. Suski, and Michael J. Louison. Post-release locomotor activity of ice-angled Northern Pike. Fisheries Research, 256(??):Article 106481, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002582.

Bergstrom:2022:LTD

[BLE^{+}22] Ulf Bergström, Stefan Larsson, Mårten Erlandsson, Maria Ovegård, Henrik Ragnarsson Stabo, Örjan Östman, and Göran Sundblad. Long-term decline in northern pike (Esox lucius L.) populations in the Baltic Sea revealed by recreational angling data. Fisheries Research, 251(??): Article 106307, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622000844.

Becker:2023:HSR

[BLFT23] Alistair Becker, Michael B. Lowry, Ashley M. Fowler, and Matthew D. Taylor. Hydroacoustic surveys reveal the distribution of mid-water fish around two artificial reef designs in temperate Australia. Fisheries Research, 257(??):Article 106509, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002867.

Brinkhof:2020:SSC

[BLHS20] Jesse Brinkhof, Roger B. Larsen, Bent Herrmann, and Manu Sistiaga. Size selectivity and catch efficiency of bottom trawl with a double sorting grid and diamond mesh codend in the North-east Atlantic gadoid fishery. Fisheries Research, 231(??):Article 105647, November 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301648.

Bryhn:2023:PDE

[BLK23] Andreas C. Bryhn, Sven Gunnar Lunneryd, and Sara Königson. Predator damaged eel caught in coastal fisheries of the Baltic Sea. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002205.

Broadhurst:2022:VNC

[BM22] Matt K. Broadhurst and Russell B. Millar. Validating a narrow codend cover and improving selectivity in south-eastern Australian fish trawls targeting eastern school whiting, Sillago flindersi. Fisheries Research, 251(??):Article 106302, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000790.

Broadhurst:2023:ISS

[BM23a] Matt K. Broadhurst and Russell B. Millar. Improved size selection and estimated at-vessel mortalities for an Australian whiting (Sillago spp.) boat seine. Fisheries Research, 266(??): ??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001716.

Broadhurst:2023:IPG

[BM23b] Matt K. Broadhurst and Russell B. Millar. Improving the performance of a generic fisheye as a secondary bycatch reduction device in a south-eastern Australian penaeid-trawl fishery. Fisheries Research, 259(??):Article 106562, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003393.

Broadhurst:2024:ATV

[BM24a] Matt K. Broadhurst and Russell B. Millar. Assessing T45 vs T0 and T90 meshes and covered codend effects in southeastern Australian penaeid trawls. Fisheries Research, 270(??):??, February 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002758.

Broadhurst:2024:RPT

[BM24b] Matt K. Broadhurst and Russell B. Millar. Relative performance of two bycatch reduction devices for excluding juvenile mulloway, Argyrosomus japonicus and other fish from south-eastern Australian penaeid trawls. Fisheries Research, 271(??):??, March 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002989.

Broadhurst:2025:EHH

[BM25a] Matt K. Broadhurst and Russell B. Millar. Effects of headline height on catches in southeastern Australian fish trawls. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002832.

Broadhurst:2025:RCM

[BM25b] Matt K. Broadhurst and Russell B. Millar. Reducing codend mesh size and changing configurations to improve selectivity in Australian whiting (Sillago spp.) trawls. Fisheries Research, 282(??):??, February 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000104.

Broadhurst:2025:RSS

[BM25c] Matt K. Broadhurst and Russell B. Millar. Refining size selection in an Australian whiting (Sillago spp.) boat seine. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003151.

Busbridge:2020:COM

[BMA^{+}20] Thomas A. J. Busbridge, C. Tara Marshall, Alexander I. Arkhipkin, Zhanna Shcherbich, Andy L. Marriott, and Paul Brickle. Can otolith microstructure and elemental fingerprints elucidate the early life history stages of the gadoid southern blue whiting (Micromesistius australis australis)? Fisheries Research, 228(??):Article 105572, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300898.

BinAziz:2024:UCS

[BMA^{+}24] Md. Saifullah Bin Aziz, Md. Mostafizur Rahman Mondol, Md. Mehedi Alam, Mohammad Mahfujul Haque, and Seikh Razibul Islam. Underpinning the criteria for the sustainability assessment of Hakaluki Haor using the RAPFISH tool. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001449.

Benoit:2020:IED

[BMC20] Hugues P. Benoît, Marie Morfin, and Connor W. Capizzano. Improved estimation of discard mortality rates with in situ experiments involving electronic and traditional tagging. Fisheries Research, 221(??):Article 105398, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361930253X.

Bortoluzzi:2024:TMP

[BMJ^{+}24] Jenny R. Bortoluzzi, Grace E. McNicholas, Andrew L. Jackson, C. Antonia Klöcker, Keno Ferter, Claudia Junge, Otte Bjelland, Adam Barnett, Austin J. Gallagher, Neil Hammerschlag, William K. Roche, and Nicholas L. Payne. Transboundary movements of porbeagle sharks support need for continued cooperative research and management approaches. Fisheries Research, 275(??):??, July 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000717.

Bell:2021:PWU

[BMM^{+}21] Richard J. Bell, M. Conor McManus, Jason McNamee, James Gartland, Ben Galuardi, and Chris McGuire. Perspectives from the water: Utilizing fisher’s observations to inform SNE/MA windowpane science and management. Fisheries Research, 243(??):Article 106090, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002186.

Baulier:2024:SDC

[BMM^{+}24] Loïc Baulier, Matthew McHugh, Cóilín Minto, Fabien Morandeau, Shane Murphy, Quiterie Sourget, Martin Oliver, and Marie Morfin. Survivorship of discarded cuckoo ray in bottom trawl fisheries in the northern Bay of Biscay, Southern Celtic and Irish Seas. Fisheries Research, 273(??):??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000353.

Bastari:2022:LFE

[BMOC22] Azzurra Bastari, Yesmina Mascarell, Miquel Ortega, and Marta Coll. Local fishers experience can contribute to a better knowledge of marine resources in the Western Mediterranean Sea. Fisheries Research, 248(??):Article 106222, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003507.

Baker:2022:LMM

[BMSM22] Krista D. Baker, Darrell R. J. Mullowney, and Bernard Sainte-Marie. Large males matter: Low sperm reserves in female snow crab (Chionoecetes opilio) off Newfoundland, Canada. Fisheries Research, 253(??):Article 106385, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362200162X.

Britton:2021:CAC

[BN21] J. Robert Britton and Emma T. Nolan. Comparative angler catch rates of native versus alien piscivorous fish in an invaded river fishery. Fisheries Research, 240(??):Article 105970, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000989.

Blanco:2023:CDA

[BNL^{+}23] Marta Blanco, David Nos, Antoni Lombarte, Laura Recasens, Joan B. Company, and Eve Galimany. Characterization of discards along a wide bathymetric range from a trawl fishery in the NW Mediterranean. Fisheries Research, 258(??):Article 106552, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003290.

Ba:2023:MPS

[BNTK23] Kamarel Ba, Fambaye Ngom, Salimata Tall, and Justin Kantoussan. Movement patterns and size distribution of little tunny (Euthynnus alletteratus) in the Eastern Central Atlantic Ocean: Insights from conventional tag release– recapture data. Fisheries Research, 266(??):??, October 2023.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001546.

Bernotas:2020:HDE

[BÖN20] Priit Bernotas, Burak Öğlü, and Peeter Nõges. How do environmental factors affect the yield of European eel (Anguilla anguilla) in a restocked population? Fisheries Research, 230(??):Article 105649, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301661.

Best:2020:PBS

[BP20] John K. Best and André E. Punt. Parameterizations for Bayesian state-space surplus production models. Fisheries Research, 222(??):Article 105411, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302668.

Boswell:2020:ERB

[BPT^{+}20] Kevin M. Boswell, Geir Pedersen, J. Christopher Taylor, Savannah LaBua, and William F. Patterson. Examining the relationship between morphological variation and modeled broadband scattering responses of reef-associated fishes from the Southeast United States. Fisheries Research, 228(??):Article 105590, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301077.

Bacheler:2025:SDH

[BPT^{+}25] Nathan M. Bacheler, William F. Patterson III, Joseph H. Tarnecki, Kyle W. Shertzer, Jeffrey A. Buckel, Nathan J. Hostetter, Krishna Pacifici, Viviane Zulian, and Walter J. Bubley. Spatiotemporal dynamics and habitat use of red snapper (Lutjanus campechanus) on the southeastern United States Atlantic continental shelf. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002649.

Barnett:2020:IPM

[BQBW20] Heidy K. Barnett, Thomas P. Quinn, Mary Bhuthimethee, and James R. Winton. Increased prespawning mortality threatens an integrated natural- and hatchery-origin sockeye salmon population in the Lake Washington Basin. Fisheries Research, 227(??):Article 105527, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300448.

Barandica:2024:CBM

[BQGV^{+}24] Juan Carlos Narváez Barandica, Julian F. Quintero-Galvis, Ana Caruso Vargas, Germán Blanco Cervantes, Juan Carlos Aguirre Pabón, Lyda Castro García, Ricardo Betancur-R., and Arturo Acero P. Congruence between microsatellite loci, SNPs, and COI gene reveals a phylogeographic break in the southern Caribbean Sea: the case of West Indian top shell Cittarium pica (Gastropoda: Tegulidae). Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002503.

Blyth:2022:EEC

[BR22] Samuel Blyth and Patrik Rönnbäck. To eat or not to eat, coastal sea trout anglers’ motivations and perceptions of best practices for catch and release. Fisheries Research, 254(??):Article 106412, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001898.

Binstock:2023:AEC

[BRGB^{+}23] Addie L. Binstock, Travis M. Richards, Kesley Gibson-Banks, J. Marcus Drymon, R. J. David Wells, and John A. Mohan. Assessing ecological connectivity of blacktip sharks (Carcharhinus limbatus) in the Gulf of Mexico using stable isotope analysis of multiple tissues. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623002424.

Burke:2020:SAC

[BRN^{+}20] Patrick J. Burke, Vincent Raoult, Lisa J. Natanson, Timothy D. Murphy, Victor Peddemors, and Jane E. Williamson. Struggling with age: Common sawsharks (Pristiophorus cirratus) defy age determination using a range of traditional methods. Fisheries Research, 231(??):Article 105706, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030223X.

Brooks:2024:PAM

[Bro24] Elizabeth N. Brooks. Pragmatic approaches to modeling recruitment in fisheries stock assessment: a perspective. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002898.

Broadhurst:2025:CTB

[Bro25] Matt K. Broadhurst. Comparing top- and bottom-opening Nordmøre-grids in south-eastern Australian estuarine squid trawls. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003102.

Brickle:2021:OTE

[BRR^{+}21] Paul Brickle, Haseeb S. Randhawa, Malcolm R. Reid, Brendon Lee, Zhanna Shcherbich, and Alexander I. Arkhipkin. Otolith trace elemental analyses and parasites provide useful tools for the stock discrimination of Patagonotothen ramsayi (Regan, 1913) (Nototheniidae) on the southern Patagonian Shelf. Fisheries Research, 244(??):Article 106129, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002575.

Boute:2022:IIW

[BRvL^{+}22] Pim G. Boute, Adriaan D. Rijnsdorp, Johan L. van Leeuwen, W. Sarina M. Versteeg, Remco P. M. Pieters, and Martin J. Lankheet. Internal injuries in whiting (Merlangius merlangus) caught by tickler-chain and pulse-trawl gears. Fisheries Research, 253(??):Article 106351, September 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200128X.

Broadhurst:2020:ITP

[BS20a] Matt K. Broadhurst and David J. Sterling. Influence of top-panel orientation (lead-ahead) on targeted and incidental catches in a penaeid trawl. Fisheries Research, 228(??):Article 105562, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300795.

Bunch:2020:IPT

[BS20b] Aaron J. Bunch and William T. Stewart. Integrating PIT technology into gear evaluation in an unregulated desert tributary. Fisheries Research, 221(??):Article 105366, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302218.

Blampied:2023:SEI

[BSA^{+}23] Samantha R. Blampied, Emma V. Sheehan, Martin J. Attrill, Francis C. T. Binney, and Sian E. Rees. The socioeconomic impact of Marine Protected Areas in Jersey: a fishers’ perspective. Fisheries Research, 259(??):Article 106555, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003320.

Bova:2022:IAL

[BSAP22] Christopher S. Bova, Jed Stephens, Shankar Aswani, and Warren M. Potts. Is the instrumental approach a ‘silver bullet’ for addressing non-compliance in recreational fisheries: a South African case study. Fisheries Research, 255(??):Article 106439, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002168.

Brennecke:2022:FSB

[BSKL^{+}22] Dennis Brennecke, Ursula Siebert, Lotte Kindt-Larsen, Henrik Skov Midtiby, Henrik Dyrberg Egemose, Sara Torres Ortiz, Katrin Knickmeier, and Magnus Wahlberg. The finescale behavior of harbor porpoises towards pingers. Fisheries Research, 255(??):Article 106437, November 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002144.

Bower:2022:CSM

[BSR^{+}22] Shannon D. Bower, Petra Szekeres, Rajeev Raghavan, Andy J. Danylchuk, and Steven J. Cooke. Consequences of simulated multiple catch-and-release events and different handling procedures on reflex impairment, ventilation rate, and body condition in Tor khudree. Fisheries Research, 246(??):Article 106175, February 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0165783621003039.

Bennett:2021:SEM

[BSSE21] Nathan J. Bennett, Anna Schuhbauer, Daniel Skerritt, and Naazia Ebrahim. Socio-economic monitoring and evaluation in fisheries. Fisheries Research, 239(??):Article 105934, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362100062X.

Bakhshalizadeh:2021:UFR

[BTB^{+}21] Shima Bakhshalizadeh, Anastassiya Tchaikovsky, Ali Bani, Thomas Prohaska, and Andreas Zitek. Using fin ray chemistry to discriminate hatchery reared juvenile age-0 Persian sturgeons by their origin in the Southern Caspian Sea region using split stream ICP-MS/MC ICP-MS. Fisheries Research, 243(??):Article 106093, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002216.

Bell:2023:CPU

[BTC23] Richard J. Bell, Adrien Tableau, and Jeremy S. Collie. Changes in the productivity of US West Coast fish stocks.

Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001054.

Becker:2022:RAR

[BTFL22] Alistair Becker, Matthew Taylor, Heath Folpp, and Michael Lowry. Revisiting an artificial reef after 10 years: What has changed and what remains the same? Fisheries Research, 249(??):Article 106261, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000388.

Becker:2020:ALR

[BTML20] Alistair Becker, Matthew Taylor, James McLeod, and Michael Lowry. Application of a long-range camera to monitor fishing effort on an offshore artificial reef. Fisheries Research, 228(??):Article 105589, August 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301065.

Biquet:2024:TWS

[BTR^{+}24] Juliette Aminian Biquet, Paul Tixier, Gaëtan Richard, Marie Soehnlen, Thibaut Thellier, Pamela Carzon, Eric Clua, and Christophe Guinet. Toothed whale and shark depredation and bycatch in the longline fishery of French Polynesia. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003211.

Boyra:2024:TSM

[BUG^{+}24] Guillermo Boyra, Jon Uranga, Nicolas Goñi, Iann Godard, and Bea Sobradillo. Target strength measurements of Atlantic bluefin tuna (Thunnus thynnus) from a live-bait fishing vessel in the Bay of Biscay. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002145.

Babyn:2021:GFA

[BVR^{+}21] Jonathan Babyn, Divya Varkey, Paul Regular, Danny Ings, and Joanna Mills Flemming. A Gaussian field approach to generating spatial age length keys. Fisheries Research, 240(??):Article 105956, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000849.

Baird:2020:UDC

[BW20] Robin W. Baird and Daniel L. Webster. Using dolphins to catch tuna: Assessment of associations between pantropical spotted dolphins and yellowfin tuna hook and line fisheries in Hawai‘ i. Fisheries Research, 230(??):Article 105652, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301697.

Bayse:2023:ORS

[BW23] S. M. Bayse and P. D. Winger. Observations of redfish (Sebastes spp.) escaping a T90 mesh codend. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002400.

Blackmore:2023:ELH

[BWB^{+}23] Robert J. Blackmore, Paul D. Winger, Pierre-Paul Bitton, Shannon Bayse, Kira Whittaker, and William A. Montevecchi. The effects of LED handline attachments on Atlantic cod (Gadus morhua) catch efficacy and bycatch. Fisheries Research, 258(??):Article 106543, February 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003204.

Blount:2024:ASU

[BWB^{+}24] Craig Blount, Duncan G. Worthington, Maria Byrne, Rowan C. Chick, and Neil L. Andrew. Aging of the sea urchin Centrostephanus rodgersii using demi-pyramid microstructure. Fisheries Research, 270(??):??, February 2024.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002928.

Baker:2021:UMS

[BWG^{+}21] Matthew R. Baker, Kresimir Williams, H. G. Greene, Casey Greufe, Heather Lopes, John Aschoff, and Rick Towler. Use of manned submersible and autonomous stereo-camera array to assess forage fish and associated subtidal habitat. Fisheries Research, 243(??):Article 106067, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001958.

Bayse:2021:RCI

[BWN^{+}21] Shannon M. Bayse, Paul D. Winger, Khanh Q. Nguyen, Meghan Donovan, Rioghnach Steiner, and Scott M. Grant. Response to comments on ’Increased catches of snow crab (Chionoecetes opilio) with luminescent-netting pots at long soak times’ by Mullowney et al. 2021. Fisheries Research, 239(??):Article 105923, July 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000515.

Bryan:2023:DCB

[BWOR23] David R. Bryan, Kresimir Williams, Cecilia A. O’Leary, and Christopher N. Rooper. The design of a camera-based fisheries-independent survey for untrawlable habitat in the Gulf of Alaska. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001789.

Berg:2024:ETD

[BWR24] Casper W. Berg, Kai Wieland, and Anna Rindorf. Effect of tow duration, door spread, and swept area on the catch efficiency of a bottom trawl. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001723.

Cook:2021:FBS

[CAAFH21] Robin Cook, Emmanuel Acheampong, Joseph Aggrey-Fynn, and Michael Heath. A fleet based surplus production model that accounts for increases in fishing power with application to two West African pelagic stocks. Fisheries Research, 243(??):Article 106048, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001764.

Cadrin:2020:DSS

[Cad20] Steven X. Cadrin. Defining spatial structure for fishery stock assessment. Fisheries Research, 221(??):Article 105397, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302528.

Chea:2023:FCP

[CAGLT23] Ratha Chea, Dewan Ahsan, Iria García-Lorenzo, and Louise Teh. Fish consumption patterns and value chain analysis in north-western Cambodia. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362300070X.

Cuende:2020:SRU

[CAH^{+}20] Elsa Cuende, Luis Arregi, Bent Herrmann, Manu Sistiaga, and Iñigo Onandia. Stimulating release of undersized fish through a square mesh panel in the Basque otter trawl fishery. Fisheries Research, 224(??):Article 105431, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302863.

Campana:2023:LTS

[Cam23] Steven E. Campana. Long-term shifts in otolith age interpretations. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000747.

Crisafulli:2023:UIA

[CAYM^{+}23] Brett Crisafulli, Ebenezer Afrifa-Yamoah, Ute Mueller, Karina Ryan, David Fairclough, and Johnny Lo. Using intervention analysis to evaluate the trends in release rates of recreational fisheries following extensive management changes. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002114.

Cadigan:2024:SSS

[CAZN24] Noel G. Cadigan, Christoffer Moesgaard Albertsen, Nan Zheng, and Anders Nielsen. Are state-space stock assessment model confidence intervals accurate? Case studies with SAM and Barents Sea stocks. Fisheries Research, 272(??): ??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000146.

Couillard:2020:CVM

[CB20] Catherine M. Couillard and Sylvie Brulotte. Comparison of a visual method, mass-based and surface-based gonadal indices and gonad histology to assess sexual maturity in the waved whelk, Buccinum undatum. Fisheries Research, 224(??):Article 105468, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303236.

Cacciapaglia:2024:DWD

[CBA^{+}24] Christopher Cacciapaglia, Elizabeth N. Brooks, Charles F. Adams, Christopher M. Legault, Charles T. Perretti, and Deborah Hart. Developing workflow and diagnostics for model selection of a vector autoregressive spatiotemporal (VAST) model in comparison to design-based indices. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000730.

Cudney:2022:DCS

[CBD^{+}22] Jennifer L. Cudney, Charles W. Bangley, Andrea Dell’Apa, Eric Diaddorio, and Roger A. Rulifson. Development of a continental shelf acoustic telemetry array to support behavioral research of fish in a high energy ocean environment. Fisheries Research, 247(??):Article 106177, March 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003052.

Coscino:2024:ILH

[CBHS24] Connor L. Coscino, Lyall Bellquist, William J. Harford, and Brice X. Semmens. Influence of life history characteristics on data-limited stock status assertions and minimum size limit evaluations using Length-Based Spawning Potential Ratio (LBSPR). Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001000.

Coleman:2024:ERN

[CBJ24] Matthew T. Coleman, Isobel S. M. Bloor, and Stuart R. Jenkins. Establishing the role of the North Atlantic Oscillation as a potential driver of brown crab Cancer pagurus density. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400184X.

Clarke:2021:DAE

[CBN^{+}21] Shannon H. Clarke, Jacob W. Brownscombe, Liane Nowell, Aaron J. Zolderdo, Andy J. Danylchuk, and Steven J. Cooke. Do angler experience and fishing lure characteristics influence welfare outcomes for largemouth bass? Fisheries Research, 233(??):Article 105756, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302733.

Courter:2023:IAM

[CBR^{+}23] Ian I. Courter, Thomas Buehrens, Mark Roes, Tara E. Blackman, Benjamin Briscoe, and Sean Gibbs. Influence of angling methods and terminal tackle on survival of salmon and steelhead caught and released in the Cowlitz River, Washington. Fisheries Research, 268(??):??, December 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002412.

Cooling:2020:SSV

[CBTH20] Kris Cooling, Matt K. Broadhurst, Daniel J. Tolhurst, and Brian Hughes. Sex-specific vulnerability of Portunus armatus to capture in round traps with traditional and novel fish baits. Fisheries Research, 226(??):Article 105518, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300357.

Curtis:2020:POA

[CC20] K. Alexandra Curtis and James V. Carretta. ObsCovgTools: Assessing observer coverage needed to document and estimate rare event bycatch. Fisheries Research, 225(??): Article 105493, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620300102.

Cao:2022:MTV

[CC22] Jie Cao and Yong Chen. Modeling time-varying natural mortality in size-structured assessment models. Fisheries Research, 250(??):Article 106290, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000674.

Cooke:2021:EDT

[CCC^{+}21] Steven J. Cooke, Cameron J. A. Cooke, Joshua T. H. Cooke, Benjamin W. C. Cooke, Andy J. Danylchuk, and Jacob W. Brownscombe. Efficacy of dehooking tools for the removal of hooks from the jaw region of angled fish. Fisheries Research, 240(??):Article 105965, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100093X.

Cooke:2022:EDH

[CCC^{+}22a] Steven J. Cooke, Benjamin W. C. Cooke, Joshua T. H. Cooke, Cameron J. A. Cooke, Luc LaRochelle, Andy J. Danylchuk, Sascha Clark Danylchuk, and Robert J. Lennox. Evaluating different hook removal gear for in-water dehooking of jaw-hooked fish captured with barbed or barbless hooks. Fisheries Research, 248(??):Article 106201, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003295.

Couch:2022:SCS

[CCC^{+}22b] Claire E. Couch, Michael E. Colvin, Rob L. Chitwood, James T. Peterson, and Carl B. Schreck. Scope of the cortisol stress response in Chinook salmon during maturation. Fisheries Research, 254(??):Article 106416, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200193X.

Coulter:2020:UHH

[CCCM^{+}20] Angie Coulter, Tim Cashion, Andrés M. Cisneros-Montemayor, Sarah Popov, Gordon Tsui, Frédéric Le Manach, Laurenne Schiller, Maria Lourdes D. Palomares, Dirk Zeller, and Daniel Pauly. Using harmonized historical catch data to infer the expansion of global tuna fisheries. Fisheries Research, 221(??):Article 105379, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302346.

Carvalho:2020:CIE

[CCGR20] Natacha Carvalho, John Casey, Jordi Guillen, and Philip Rodgers. Characterising investments in EU fisheries and defining their desirability. Fisheries Research, 221 (??):Article 105396, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302516.

Chladek:2020:SHP

[CCKL^{+}20] Jérôme Chladek, Boris Culik, Lotte Kindt-Larsen, Christoffer Moesgaard Albertsen, and Christian von Dorrien. Synthetic harbour porpoise (Phocoena phocoena) communication signals emitted by acoustic alerting device (Porpoise ALert, PAL) significantly reduce their bycatch in western Baltic gillnet fisheries. Fisheries Research, 232(??):Article 105732, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302496.

Castagnino:2024:PSA

[CCR24] Fabio Castagnino, Matías Caillaux, and Elmer Ramos. Productivity susceptibility analysis in extremely data-poor scenarios: the case of Peruvian coastal groundfish. Fisheries Research, 276(??):??, August 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001085.

Colin:2024:EPW

[CCRGC^{+}24] Angélica Colín, Susette Castañeda-Rico, Luis M. Guevara-Chumacero, Eloisa Pacheco-Almanzar, and Ana L. Ibáñez. Evidence of philopatry in white mullet Mugil curema (Pisces: Mugilidae). Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400239X.

Chumchuen:2023:ATM

[CDAK23] Watcharapong Chumchuen, Sahaphat Duerasor, Chalermchat Arunrojprapai, and Chakhrit Keesaphong. Application of three materials of fishing gear marking for crab and fish gillnets in small-scale fishery in the Andaman Sea coast of Thailand. Fisheries Research, 262(??):Article 106668, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000619.

Clark:2021:OSA

[CdSLP21] Fernando José König Clark, Caroline Stefani da Silva Lima, and André Luiz Machado Pessanha. Otolith shape analysis of the Brazilian silverside in two northeastern Brazilian estuaries with distinct salinity ranges. Fisheries Research, 243(??):Article 106094, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002228.

Cao:2021:MCU

[CEAL21] Nga Thi Hong Cao, Arne Eide, Claire W. Armstrong, and Long Kim Le. Measuring capacity utilization in fisheries using physical or economic variables: a data envelope analysis of a Vietnamese purse seine fishery. Fisheries Research, 243(??):Article 106087, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002150.

Chong:2023:EPA

[CFB^{+}23] Lisa Chong, Nicholas Fisch, John Scott Borsum, Jennifer Granneman, Diana Perry, Gabrielle Love, Brittany Hall-Scharf, Robert Botta, Kai Lorenzen, Edward Camp, and Zachary Siders. Examining the performance of alternative harvest regulations for short-lived taxa: a case study of Florida Bay scallop management. Fisheries Research, 263 (??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000760.

Clegg:2023:EAB

[CFO23] Thomas L. Clegg, Edvin Fuglebakk, and Kotaro Ono. Evaluating assumptions behind design-based estimators for unreported catches. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000796.

Cronin-Fine:2021:MTV

[CFP21] Lee Cronin-Fine and André E. Punt. Modeling time-varying selectivity in size-structured assessment models. Fisheries Research, 239(??):Article 105927, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000552.

Cronin-Fine:2022:FIS

[CFP22] Lee Cronin-Fine and André E. Punt. Factors influencing size-structured models’ ability to estimate natural mortality. Fisheries Research, 250(??):Article 106292, June 2022.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000698.

Cullen:2021:UGW

[CG21] Daniel W. Cullen and Vincent Guida. Use of geographically weighted regression to investigate spatial non-stationary environmental effects on the distributions of black sea bass (Centropristis striata) and scup (Stenotomus chrysops) in the Mid-Atlantic Bight, USA. Fisheries Research, 234(??):Article 105795, February 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030312X.

Chhor:2022:STB

[CGB^{+}22] Auston D. Chhor, Daniel M. Glassman, Jacob W. Brownscombe, Alexandria T. Trahan, Andy J. Danylchuk, and Steven J. Cooke. Short-term behavioural impacts of airexposure in three species of recreationally angled freshwater fish. Fisheries Research, 253(??):Article 106342, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001199.

Cadrin:2023:BPD

[CGBJ23] Steven X. Cadrin, Daniel R. Goethel, Aaron Berger, and Ernesto Jardim. Best practices for defining spatial boundaries and spatial structure in stock assessment. Fisheries Research, 262(??):Article 106650, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000437.

Cheng:2024:ACF

[CGC24] Matthew L. H. Cheng, Daniel R. Goethel, and Curry J. Cunningham. Addressing complex fleet structure in fishery stock assessment models: Accounting for a rapidly developing pot fishery for Alaska sablefish (Anoplopoma fimbria). Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003107.

Castro-Garcia:2025:OSV

[CGDTSA^{+}25] Andrés Castro-García, Erwan Delrieu-Trottin, Pablo Saenz-Agudelo, Cristian Rapu-Edmunds, Guido Plaza, Federico Márquez, and Mauricio F. Landaeta. Otolith shape variability of labrid fish from Rapa Nui (Easter Island), southeastern Pacific. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002972.

Cerna:2022:SIA

[CGM^{+}22] Francisco Cerna, Mackarena Gómez, Guillermo Moyano, Guido Plaza, and Beatriz Morales-Nin. Spatial and interannual changes in the growth patterns of young-of-year anchovy in a high productive ecosystem. Fisheries Research, 249(??):Article 106236, May 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000133.

Carosi:2022:CRA

[CGSL22] A. Carosi, L. Ghetti, A. Soresina, and M. Lorenzoni. Catch and release angling: Implications for the management and conservation of the Mediterranean trout in central Italy. Fisheries Research, 250(??):Article 106285, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000625.

Charsley:2023:CSS

[CGT^{+}23] Anthony R. Charsley, Arnaud Grüss, James T. Thorson, Merrill B. Rudd, Shannan K. Crow, Bruno David, Erica K. Williams, and Simon D. Hoyle. Catchment-scale stream network spatio-temporal models, applied to the freshwater stages of a diadromous fish species, longfin eel (Anguilla dieffenbachii). Fisheries Research, 259(??):Article 106583, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003605.

Card:2021:PEC

[CH21a] Jamie T. Card and Caleb T. Hasler. Physiological effects of catch-and-release angling on freshwater drum (Aplodinotus grunniens). Fisheries Research, 237(??):Article 105881, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000096.

Conrath:2021:TVR

[CH21b] Christina L. Conrath and Peter-John F. Hulson. Temporal variability in the reproductive parameters of deepwater rockfishes in the Gulf of Alaska. Fisheries Research, 237(??):Article 105876, May 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000047.

Cope:2022:UVA

[CH22] Jason M. Cope and Owen S. Hamel. Upgrading from M version 0.2: an application-based method for practical estimation, evaluation and uncertainty characterization of natural mortality. Fisheries Research, 256(??):Article 106493, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002703.

Crisafulli:2025:APR

[CHAY^{+}25] Brett Crisafulli, Sybrand Alexander Hesp, Ebenezer Afrifa-Yamoah, Ute A. Mueller, Karina L. Ryan, David V. Fairclough, Norman G. Hall, and Johnny Lo. Accounting for post-release mortality in data-limited, length-based assessments for four demersal fish species with varying life histories. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003023.

Childress:2024:DJS

[CHB24] Michael J. Childress, Coral Holt, and Rodney D. Bertelsen. Displaced juvenile and subadult Caribbean spiny lobsters show strong orientation toward home dens. Fisheries Research, 279(??):??, November 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001966.

Chen:2021:FMU

[Che21] Xi Chen. Fisheries management under incomplete information by optimal stochastic control and hidden Markov model filter. Fisheries Research, 243(??):Article 106047, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001752.

Cheng:2025:MPC

[CHGC25] Matthew LH. Cheng, Peter-John F. Hulson, Daniel R. Goethel, and Curry J. Cunningham. A mathematical proof comparing the statistical properties between two common approaches for parameterizing sex-composition likelihoods in fishery stock assessments. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002959.

Cottingham:2020:ERE

[CHL^{+}20] Alan Cottingham, Norman G. Hall, Neil R. Loneragan, Gregory I. Jenkins, and Ian C. Potter. Efficacy of restocking an estuarine-resident species demonstrated by long-term monitoring of cultured fish with alizarin complexone-stained otoliths. a case study. Fisheries Research, 227(??):Article 105556, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300734.

Clarke:2024:APC

[CHM24] Leo J. Clarke, Phoebe N. Holding, and Ian D. McCarthy. An assessment of post-capture condition and survival of Rajidae caught in fixed nets. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400105X.

Champion:2020:MUB

[CHPT20] Curtis Champion, Alistair J. Hobday, Gretta T. Pecl, and Sean R. Tracey. Maximising the utility of bioelectrical impedance analysis for measuring fish condition requires identifying and controlling for sources of error. Fisheries Research, 229(??):Article 105575, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300928.

Charles:2020:IBC

[CHT20] William Dantas Charles, Humberto Hazin, and Paulo Travassos. Interactions between cetaceans and the tuna/swordfish pelagic longline fishery in the tropical western Atlantic Ocean. Fisheries Research, 226(??):Article 105530, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300473.

Cambie:2020:IGV

[CJC^{+}20] Giulia Cambiè, Imed Jribi, Irene Cambera, Giulia Vagnoli, Daniela Freggi, and Paolo Casale. Intra-gear variation in sea turtle bycatch: Implications for fisheries management. Fisheries Research, 221(??):Article 105405, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302607.

Chowdhury:2021:GCK

[CKD^{+}21] Labrechai Mog Chowdhury, A. Kathirvelpandian, P. R. Divya, V. S. Basheer, Chelath Mohitha, A. Pavan-Kumar, and Gopal Krishna. Genetic characterization of kiddi shrimp, Parapenaeopsis stylifera (H. Milne Edwards, 1837) along the Indian coast using microsatellite markers. Fisheries Research, 244(??):Article 106128, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002563.

Charitonidou:2020:CPO

[CKDP^{+}20] Katerina Charitonidou, Olav Sigurd Kjesbu, Rosario Dominguez-Petit, Dolores Garabana, Maria Albisua Korta, Maria Santos, Cindy J. G. van Damme, Anders Thorsen, and Kostas Ganias. Contrasting post-ovulatory follicle production in fishes with different spawning dynamics. Fisheries Research, 231(??):Article 105710, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302277.

Chamberlin:2020:HAE

[CKK^{+}20] Derek W. Chamberlin, Carey T. Knight, Richard T. Kraus, Ann Marie Gorman, Wenzhao Xu, and Paris D. Collingsworth. Hypoxia augments edge effects of water column stratification on fish distribution. Fisheries Research, 231(??):Article 105684, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302010.

Charbonneau:2020:PDM

[CKM^{+}20] Julie A. Charbonneau, David M. Keith, M. Aaron Mac-Neil, Jessica A. Sameoto, and Jeffrey A. Hutchings. Pervasive declines in monkfish (Lophius americanus) size structure throughout the northwest Atlantic. Fisheries Research, 230(??):Article 105633, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301508.

Clark:2022:WNM

[Cla22] William G. Clark. Why natural mortality is estimable, in theory if not in practice, in a data-rich stock assessment. Fisheries Research, 248(??):Article 106203, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003313.

Canel:2021:OAF

[CLB^{+}21] Delfina Canel, Eugenia Levy, Paola E. Braicovich, Manuel Haimovici, and Juan T. Timi. Ontogenetic asynchrony in fish migrations may lead to disparate parasite assemblages: Implications for its use as biological tags. Fisheries Research, 239(??):Article 105941, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000692.

Cooke:2022:IHB

[CLD^{+}22] Steven J. Cooke, Luc LaRochelle, Andy J. Danylchuk, Sascha Clark Danylchuk, and Lucas P. Griffin. Influence of hook barbs on the “through-the-gill” hook removal method for deeply hooked Smallmouth Bass. Fisheries Research, 251(??):Article 106322, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000996.

Crisafulli:2022:ICE

[CLM^{+}22] Brett Crisafulli, Johnny Lo, Ute Mueller, Karina Ryan, and David Fairclough. Increasing confidence in estimates of average weight and recreational harvest ranges. Fisheries Research, 248(??):Article 106208, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003362.

Courtney:2022:ENM

[CLY^{+}22] Anthony J. Courtney, George M. Leigh, Wen-Hsi Yang, Matthew J. Campbell, and Mark F. McLennan. Estimating the natural mortality rate of saucer scallops (Ylistrum balloti) on the Queensland east coast from tag-recaptures. Fisheries Research, 250(??):Article 106273, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000509.

Casselberry:2022:WFB

[CMA^{+}22] Grace A. Casselberry, Ezra M. Markowitz, Kelly Alves, Joseph Dello Russo, Gregory B. Skomal, and Andy J. Danylchuk. When fishing bites: Understanding angler responses to shark depredation. Fisheries Research, 246 (??):Article 106174, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0165783621003027.

Chande:2021:IEV

[CMBL21] Muhaji A. Chande, Yunus D. Mgaya, Lusato B. Benno, and Samwel M. Limbu. The influence of environmental variables on the abundance and temporal distribution of Octopus cyanea around Mafia Island, Tanzania. Fisheries Research, 241(??):Article 105991, September 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001193.

Carbonara:2023:BVM

[CMD^{+}23] Pierluigi Carbonara, Francesco Masnadi, Fortunata Donato, Laura Sabatini, Giulio Pellini, Massimiliano Cardinale, and Giuseppe Scarcella. Biphasic versus monophasic growth curve equation, an application to common sole (Solea solea, L.) in the northern and central Adriatic Sea. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000875.

Cisneros-Montemayor:2020:EBM

[CMIMS20] Andrés M. Cisneros-Montemayor, Gakushi Ishimura, Gordon R. Munro, and U. Rashid Sumaila. Ecosystem-based management can contribute to cooperation in transboundary fisheries: the case of Pacific sardine. Fisheries Research, 221(??):Article 105401, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302565.

Cadrin:2020:SST

[CMP20] Steven X. Cadrin, Mark N. Maunder, and André E. Punt. Spatial structure: Theory, estimation and application in stock assessment models. Fisheries Research, 229(??):Article 105608, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301259.

Cote:2020:ESS

[CMRP20] D. Cote, C. J. Morris, P. M. Regular, and M. G. Piersiak. Effects of 2D seismic on snow crab movement behavior. Fisheries Research, 230(??):Article 105661, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301788.

Correia:2021:PSC

[CMTP^{+}21] A. T. Correia, A. Moura, R. Triay-Portella, P. T. Santos, E. Pinto, A. A. Almeida, A. N. Sial, and A. A. Muniz. Population structure of the chub mackerel (Scomber colias) in the NE Atlantic inferred from otolith elemental and isotopic signatures. Fisheries Research, 234(??):Article 105785, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303027.

Cavieres:2021:ASD

[CMV21] Joaquin Cavieres, Cole C. Monnahan, and Aki Vehtari. Accounting for spatial dependence improves relative abundance estimates in a benthic marine species structured as a metapopulation. Fisheries Research, 240(??):Article 105960, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621000886.

Crespo-Neto:2021:SSS

[CNDDAPMR21] Osman Crespo-Neto, Eric Díaz-Delgado, Tatiana A. Acosta-Pachón, and Raúl O. Martínez-Rincón. Spatial segregation by size of billfishes bycaught by the tuna purseseine fishery in the Eastern Pacific Ocean. Fisheries Research, 241(??):Article 106001, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001296.

Clarke:2022:ECS

[CNE^{+}22] Shelby B. Clarke, William A. Nesbitt, Jackson Efitre, Margaret Masette, and Lauren J. Chapman. Elemental composition of small pelagic fishes in three East African lakes: Implications for nutritional security. Fisheries Research, 256(??):Article 106479, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002569.

Canete:2022:AIF

[COG22] Felipe Torres Cañete, Rodrigo Oyanedel, and Stefan Gelcich. Adoption and impacts of fishing gear innovations: Insights from a small-scale fishery in Chile. Fisheries Research, 248(??):Article 106200, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003283.

Cope:2024:GPP

[Cop24] Jason M. Cope. The good practices of practicable alchemy in the stock assessment continuum: Fundamentals and principles of analytical methods to support science-based fisheries management under data and resource limitations. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002527.

Ceballos:2021:GWA

[CPB^{+}21] Santiago G. Ceballos, Chiara Papetti, Massimiliano Babbucci, Daniel A. Fernández, Luca Schiavon, and C.- H. Christina Cheng. Genome-wide analysis reveals striking lack of genetic differentiation over long distances for the Antarctic toothfish Dissostichus mawsoni: High genetic connectivity or shared spawning grounds? Fisheries Research, 243(??):Article 106074, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002022.

Castillo:2020:EMA

[CPF20] Liliana Sierra Castillo, Michaela Pawluk, and Masami Fujiwara. Estimating mortality for the assessment of a smallscale fishery: Lane snapper (Lutjanus synagris) in Honduras. Fisheries Research, 231(??):Article 105709, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302265.

Catucci:2025:MSD

[CPL^{+}25] Elena Catucci, Diego Panzeri, Simone Libralato, Gianpiero Cossarini, Germana Garofalo, Irida Maina, Stefanos Kavadas, Federico Quattrocchi, Giulia Cipriano, Roberto Carlucci, Sergio Vitale, Chryssi Mytilineou, Fabio Fiorentino, and Tommaso Russo. Modeling the spatial distribution and abundance of deep-water red shrimps in the 4Mediterranean Sea: a machine learning approach. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003217.

Canales:2021:CLB

[CPM21] Cristian M. Canales, André E. Punt, and Mauricio Mardones. Can a length-based pseudo-cohort analysis (LBPA) using multiple catch length-frequencies provide insight into population status in data-poor situations? Fisheries Research, 234(??):Article 105810, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303271.

Chieu:2023:GWS

[CPPK23] Hoang Dinh Chieu, H. K. A. Premachandra, Daniel Powell, and Wayne Knibb. Genome-wide SNP analyses reveal a substantial gene flow and isolated-genetic structure of sea cucumber Holothuria leucospilota populations in Western Central Pacific. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300111X.

Chamberlin:2024:BVG

[CPR^{+}24] Derek W. Chamberlin, Jennifer C. Potts, Walter D. Rogers, Zachary A. Siders, and William F. Patterson. Bomb ^{14}C validates Gray triggerfish (Balistes capriscus) dorsal spine and otolith ageing protocols. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001875.

Cusba:2024:MDH

[CQA^{+}24] José Cusba, Dante Queirolo, Mauricio Ahumada, Pedro Apablaza, Jorge Paramo, and Rodrigo Wiff. Modelling density on historical aggregation areas improves biomass estimates in yellow squat lobster off Chile. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000900.

Cousido-Rocha:2022:ALB

[CRCAF^{+}22] Marta Cousido-Rocha, Santiago Cerviño, Alexandre Alonso-Fernández, Juan Gil, Isabel González Herraiz, Margarita María Rincón, Fernando Ramos, Cristina Rodríguez-Cabello, Paz Sampedro, Yolanda Vila, and Maria Grazia Pennino. Applying length-based assessment methods to fishery resources in the Bay of Biscay and Iberian Coast ecoregion: Stock status and parameter sensitivity. Fisheries Research, 248(??):Article 106197, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003258.

Cunha:2024:GGD

[CRF^{+}24] Regina L. Cunha, Joana I. Robalo, Sara M. Francisco, Inês Farias, Rita Castilho, and Ivone Figueiredo. Genomics goes deeper in fisheries science: the case of the blackspot seabream (Pagellus bogaraveo) in the northeast Atlantic. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002849.

Carlson:2021:MAH

[CRL21] Andrew K. Carlson, Daniel I. Rubenstein, and Simon A. Levin. Modeling Atlantic herring fisheries as multiscalar human-natural systems. Fisheries Research, 236(??):Article 105855, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783620303726.

Coussau:2023:SVO

[CRS23] Lola Coussau, Dominique Robert, and Pascal Sirois. Spatiotemporal variability in otolith elemental fingerprint and the potential to determine deepwater redfish (Sebastes mentella) origins and migrations in the Estuary and Gulf of St. Lawrence, Canada. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001327.

Calagui:2022:IFF

[CRSC22] Laurence B. Calagui, Jashin J. Rosal, Romell A. Seronay, and Shirlamaine Irina M. Calagui. Inventory of fish fauna in Siargao Island Protected Landscape and Seascape, Surigao del Norte, Philippines. Fisheries Research, 251(??): Article 106325, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783622001023.

Chamberlin:2023:HVL

[CSB^{+}23] Derek W. Chamberlin, Zachary A. Siders, Beverly K. Barnett, Robert N. M. Ahrens, and William F. Patterson. Highly variable length-at-age in vermilion snapper (Rhomboplites aurorubens) validated via Bayesian analysis of bomb radiocarbon. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300125X.

Chen:2023:LCM

[CSDH^{+}23] Emily K. Chen, Nicholas A. Som, John D. Deibner-Hanson, David G. Anderson, and Mark J. Henderson. A life cycle model for evaluating estuary residency and recovery potential in Chinook salmon. Fisheries Research, 257(??):Article 106511, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002880.

Chladek:2021:UIN

[CSH^{+}21] Jérôme Chladek, Daniel Stepputtis, Andreas Hermann, Isabella M. F. Kratzer, Peter Ljungberg, Paco Rodriguez-Tress, Juan Santos, and Jon C. Svendsen. Using an innovative net-pen-based observation method to assess and compare fish pot-entrance catch efficiency for Atlantic cod (Gadus morhua). Fisheries Research, 236(??):Article 105851, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303684.

Coulson:2020:TSO

[CSRL20] P. G. Coulson, R. Shotton, S. Robertson, and J. H. Lee. Thin-sectioned otoliths reveal extended longevity of southern boarfish (Pentaceros richardsoni) and are used to investigate inter-oceanic differences in length and age structure and growth. Fisheries Research, 231(??):Article 105691, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302083.

Christiansen:2022:ASS

[CSSB22] Heather M. Christiansen, Justin J. Solomon, Theodore S. Switzer, and Russell B. Brodie. Assessing the size selectivity of capture gears for reef fishes using paired stereo-baited remote underwater video. Fisheries Research, 249(??):Article 106234, May 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200011X.

Caputi:2025:ICW

[CSTdL25] Nick Caputi, Nic Sofoulis, Matt Taylor, and Simon de Lestang. The International Conference and Workshop on Lobster Biology and Management returns to Western Australia! Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000293.

Clark-Shen:2021:SRS

[CSTR^{+}21] Naomi Clark-Shen, Kathy Xu Tingting, Madhu Rao, Shannon Cosentino-Roush, Rajkumar Sandrasegeren, Anya R. Gajanur, Demian D. Chapman, Esther Lee Xin Ying, Kathryn I. Flowers, Kevin A. Feldheim, B. Mabel Manjaji-Matsumoto, and Sirius Ng Zheng Hui. The sharks and rays at Singapore’s fishery ports. Fisheries Research, 235(??):Article 105805, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303222.

Cumplido:2022:LGA

[CTCB22] Mariano Cumplido, Gastón Trobbiani, Alvar Carranza, and Gregorio Bigatti. Limited gastropod abundances call for selective, small scale artisanal fisheries in a Patagonian marine protected area. Fisheries Research, 250(??):Article 106291, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000686.

Cheng:2023:UTA

[CTIC23] Matthew L. H. Cheng, James T. Thorson, James N. Ianelli, and Curry J. Cunningham. Unlocking the triad of age, year, and cohort effects for stock assessment: Demonstration of a computationally efficient and reproducible framework using weight-at-age. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001480.

Campbell:2020:FAE

[CTM^{+}20] Matthew J. Campbell, Mark L. Tonks, Margaret Miller, David T. Brewer, Anthony J. Courtney, and Colin A. Simpfendorfer. Factors affecting elasmobranch escape from turtle excluder devices (TEDs) in a tropical penaeid-trawl fishery. Fisheries Research, 224(??):Article 105456, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361930311X.

Cieslak:2021:API

[CTR^{+}21] Matthew Cieslak, Paul Tixier, Gaétan Richard, Mark Hindell, John P. Y. Arnould, and Mary-Anne Lea. Acoustics and photo-identification provide new insights on killer whale presence and movements when interacting with longline fisheries in South East Australia. Fisheries Research, 233(??):Article 105748, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302654.

Conallin:2023:UFS

[CTS^{+}23] John Conallin, Nyi Nyi Tun, Aye Myint Swe, Lee J. Baumgartner, Zau Lunn, Martin Mallen-Cooper, Tim Marsden, Nathan Ning, Wayne Robinson, Lalantha Senevirathna, and Peter Thew. Using fish swimming ability to refine criteria for fishway construction in Myanmar. Fisheries Research, 262(??):Article 106680, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000735.

Colombo:2020:GJS

[CVM^{+}20] Julia Colombo, Julio Vinuesa, Bernardo Marqués, Tomas Isola, and Martín Varisco. Growth of juvenile southern king crabs (Lithodes santolla) in San Jorge Gulf, Southwestern Atlantic Ocean. Fisheries Research, 226(??):Article 105519, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300369.

Coulson:2022:RSD

[CW22] Peter G. Coulson and Corey B. Wakefield. Reproduction, sexually dimorphic growth, exceptional longevity and low natural mortality of the knifejaw, Oplegnathus woodwardi, from temperate waters in the south-eastern Indian Ocean. Fisheries Research, 256(??):Article 106466, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002430.

Carvalho:2021:CUM

[CWC^{+}21] Felipe Carvalho, Henning Winker, Dean Courtney, Maia Kapur, Laurence Kell, Massimiliano Cardinale, Michael Schirripa, Toshihide Kitakado, Dawit Yemane, Kevin R. Piner, Mark N. Maunder, Ian Taylor, Chantel R. Wetzel, Kathryn Doering, Kelli F. Johnson, and Richard D. Methot, Jr. A cookbook for using model diagnostics in integrated stock assessments. Fisheries Research, 240 (??):Article 105959, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621000874.

Czapla:2023:ROT

[CWM^{+}23] Philipp Czapla, Magnus Lovén Wallerius, Christopher T. Monk, Steven J. Cooke, and Robert Arlinghaus. Reexamining one-trial learning in common carp (Cyprinus carpio) through private and social cues: No evidence for hook avoidance lasting more than seven months. Fisheries Research, 259(??):Article 106573, March 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003502.

Cadigan:2024:TVS

[CWRR24] Noel G. Cadigan, S. J. W. W. M. M. P. Weerasekera, Paul M. Regular, and Rick M. Rideout. Time varying M with starvation mortality in a state-space stock assessment model: Part 2: Atlantic cod (Gadus morhua) on the southern Grand Bank of Newfoundland. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002388.

Cadigan:2022:NMR

[CYBW22] Noel Cadigan, Yihao Yin, Hugues P. Benoît, and Stephen J. Walsh. A nonparametric-monotone regression model and robust estimation for paired-tow bottom-trawl survey comparative fishing data. Fisheries Research, 254(??):Article 106422, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622001990.

Cao:2025:UFA

[CZ25] Zhi Cao and Guoping Zhu. Using fatty acids and stable isotope analyses to elucidate ontogenetic variations in diet of painted notothen Lepidonotothen larseni during winter at the South Georgia shelf, Antarctic. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002820.

Cheng:2024:PES

[CZJ^{+}24] Wen Cheng, Chongliang Zhang, Yupeng Ji, Ying Xue, Yiping Ren, and Binduo Xu. Performance evaluation of spatially balanced sampling designs in fishery-independent surveys. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002722.

Dureuil:2022:RGE

[DAD^{+}22] Manuel Dureuil, William H. Aeberhard, Michael Dowd, Sebastián A. Pardo, Frederick G. Whoriskey, and Boris Worm. Reliable growth estimation from mark–recapture tagging data in elasmobranchs. Fisheries Research, 256 (??):Article 106488, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362200265X.

deAlmeida:2023:USC

[dAdCdO^{+}23] Paulo Roberto Camponez de Almeida, Marcus Rodrigues da Costa, Raiane Soares Cirino de Oliveira, Agostinho Almeida, Rui Azevedo, Cassiano Monteiro-Neto, and Alberto Teodorico Correia. The use of the shape and chemistry of fish otoliths as a subpopulational discrimination tool for Eugerres brasilianus in lagoon systems in the Southwest Atlantic Ocean. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001881.

deAraujo:2020:EAT

[dAdSR^{+}20] Janayna Galvão de Araújo, Marcos Antônio Souza dos Santos, Fabrício Khoury Rebello, Gregor Prang, Morgana Carvalho de Almeida, and Victoria Judith Isaac. Economic analysis of the threats posed to the harvesting of ornamental fish by the operation of the Belo Monte hydroelectric dam in northern Brazil. Fisheries Research, 225(??):Article 105483, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619303388.

deAlmeida:2021:IWR

[dAGCR21] Joana M. Bento de Almeida, Miguel B. Gaspar, Margarida Castro, and Marta M. Rufino. Influence of wind, rainfall, temperature, and primary productivity, on the biomass of the bivalves Spisula solida, Donax trunculus, Chamelea gallina and Ensis siliqua. Fisheries Research, 242(??):Article 106044, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001727.

Dippold:2020:SPW

[DAL20] David A. Dippold, Grant D. Adams, and Stuart A. Ludsin. Spatial patterning of walleye recreational harvest in Lake Erie: Role of demographic and environmental factors. Fisheries Research, 230(??):Article 105676, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301934.

Dhellemmes:2023:BSS

[DAR^{+}23] F. Dhellemmes, E. Aspillaga, T. Rittweg, J. Alós, P. Möller, and R. Arlinghaus. Body size scaling of space use in coastal pike (Esox lucius) in brackish lagoons of the southern Baltic Sea. Fisheries Research, 260(??):Article 106560, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200337X.

Dorn:2022:TVP

[DB22] Martin W. Dorn and Cheryl L. Barnes. Time-varying predation as a modifier of constant natural mortality for Gulf of Alaska walleye pollock. Fisheries Research, 254 (??):Article 106391, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622001680.

Desfosses:2021:SRS

[DBDT21] Cameron J. Desfosses, Stuart J. Blight, Ainslie M. Denham, and Stephen M. Taylor. Supplemented roving survey to quantify spatio-temporal recreational fishing effort in an estuarine Ramsar wetland. Fisheries Research, 242(??):Article 106042, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001703.

Ducharme-Barth:2022:IFD

[DBGV^{+}22] Nicholas D. Ducharme-Barth, Arnaud Grüss, Matthew T. Vincent, Hidetada Kiyofuji, Yoshinori Aoki, Graham Pilling, John Hampton, and James T. Thorson. Impacts of fisheriesdependent spatial sampling patterns on catch-per-unit-effort standardization: a simulation study and fishery application. Fisheries Research, 246(??):Article 106169, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002976.

Souto:2025:CST

[DBL^{+}25] Marina Do Souto, Daniel Roberto Brown, Ezequiel Leonarduzzi, Ricardo Ismael Silva, Ana Martínez, Georgina Cepeda, Gustavo Javier Macchi, David Edgardo Galván, and Marina Vera Diaz. Comfort in stratification and trophic flexibility: Argentine anchovy, Engraulis anchoita, larvae life traits in relation to their food sources. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002790.

DeFilippo:2021:IST

[DBS^{+}21] Lukas B. DeFilippo, Thomas W. Buehrens, Mark Scheuerell, Neala W. Kendall, and Daniel E. Schindler. Improving short-term recruitment forecasts for Coho salmon using a spatiotemporal integrated population model. Fisheries Research, 242(??):Article 106014, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001429.

Ducharme-Barth:2022:FFE

[DBV22] Nicholas D. Ducharme-Barth and Matthew T. Vincent. Focusing on the front end: a framework for incorporating uncertainty in biological parameters in model ensembles of integrated stock assessments. Fisheries Research, 255(??):Article 106452, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783622002296. See corrigendum [DBV23].

Ducharme-Barth:2023:CFF

[DBV23] Nicholas D. Ducharme-Barth and Matthew T. Vincent. Corrigendum to “Focusing on the front end: a framework for incorporating uncertainty in biological parameters in model ensembles of integrated stock assessments” [Fish. Res. 255 (2022) 106452]. Fisheries Research, 260(??):Article 106593, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003708. See [DBV22].

Duarte:2024:RMD

[DC24] Debra Duarte and Steven X. Cadrin. Review of methodologies for detecting an observer effect in commercial fisheries data. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400064X.

daCunha:2023:SFR

[dCHdMS^{+}23] Suelen Maria Beeck da Cunha, Dannieli Firme Herbst, Luis C. P. de Macedo-Soares, Marta Jussara Cremer, and Natalia Hanazaki. Selection of fish resources for consumption and sale by artisanal fishers and implications to fisheries sustainability. Fisheries Research, 261(??):Article 106615, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000085.

Danby:2022:EAF

[DCK^{+}22] R. E. Danby, E. D. Clarke, R. J. Kynoch, D. G. Reid, and P. G. Fernandes. Estimating and accounting for fish losses under the footrope of a survey trawl: the case of northern shelf anglerfish. Fisheries Research, 255(??):Article 106431, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002089.

Debenedetti:2020:ANH

[DCL^{+}20] Ángela L. Debenedetti, Francisco Codes, Susana Laza, Sandra Hernández, Elena Madrid, María Trelis, and Màrius V. Fuentes. Ascaridoid nematodes in horse mackerel, Trachurus trachurus, sold in Spanish supermarkets — factors able to diminish consumer risk. Fisheries Research, 230(??):Article 105669, October 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301867.

Dillon:2020:TMR

[DCR^{+}20] Rebecca A. Dillon, Joseph D. Conroy, Lars G. Rudstam, Peter F. Craigmile, Doran M. Mason, and Stuart A. Ludsin. Towards more robust hydroacoustic estimates of fish abundance in the presence of pelagic macroinvertebrates. Fisheries Research, 230(??):Article 105667, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301843.

Davison:2024:PAA

[DCS24] Micah Davison, Timothy Copeland, and Dennis Scarnecchia. A problem for the ages: Achieving reliable fish age information through quality management. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783624001656.

Dvoretsky:2025:CSN

[DD25] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Current status of native crayfish populations in Russia: a brief review of their biology and fisheries. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400290X.

Delpiani:2020:WSW

[DDA^{+}20] G. Delpiani, S. M. Delpiani, M. Y. Deli Antoni, M. Covatti Ale, L. Fischer, L. O. Lucifora, and J. M. Díaz de Astarloa. Are we sure we eat what we buy? Fish mislabelling in Buenos Aires province, the largest sea food market in Argentina. Fisheries Research, 221(??):Article 105373, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302280.

Diaz-Delgado:2021:EPS

[DDCNMR21] Eric Díaz-Delgado, Osman Crespo-Neto, and Raúl O. Martínez-Rincón. Environmental preferences of sharks bycaught by the tuna purse-seine fishery in the Eastern Pacific Ocean. Fisheries Research, 243(??):Article 106076, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002046.

Dichmont:2021:CSA

[DDDP21] Catherine M. Dichmont, Roy A. Deng, Natalie Dowling, and André E. Punt. Collating stock assessment packages to improve stock assessments. Fisheries Research, 236(??):Article 105844, April 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303611.

Dia:2023:ECS

[DEM^{+}23] Mamadou Dia, Yeslem El Vally, Beyah Meissa, Inejih Cheikh Abdallahi, Mika Diop, Mohamed Moustapha Bouzouma, Ousmane Sarr, Ely Beibou, Cheikh Baye Braham, and Rima W. Jabado. Evolution of catches and specific composition of elasmobranchs in Mauritanian artisanal, coastal and offshore fisheries. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002035.

Dettloff:2021:ISM

[Det21a] Kyle Dettloff. Improvements to the Stephens–MacCall approach for calculating CPUE from multispecies fisheries logbook data. Fisheries Research, 242(??):Article 106038, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001661.

Dick:2021:MBE

[DET21b] E. J. Dick, Jason Edwards, and Tien-Shui Tsou. Modelbased estimation of average fish weights from recreational fisheries. Fisheries Research, 241(??):Article 106002, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001302.

Dettloff:2023:ABP

[Det23] Kyle Dettloff. Assessment of bias and precision among simple closed population mark-recapture estimators. Fisheries Research, 265(??):??, September 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001492.

Barros:2020:EDO

[dFBPL^{+}20] Daniela de França Barros, Miguel Petrere, Vincent Lecours, Davi Butturi-Gomes, Leandro Castello, and Victoria Judith Isaac. Effects of deforestation and other environmental variables on floodplain fish catch in the Amazon. Fisheries Research, 230(??):Article 105643, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301600.

Dainys:2022:ACH

[DGMG^{+}22] Justas Dainys, Harry Gorfine, Fernando Mateos-González, Christian Skov, Robertas Urbanavičius, and Asta Audzijonyte. Angling counts: Harnessing the power of technological advances for recreational fishing surveys. Fisheries Research, 254(??):Article 106410, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001874.

Davis:2020:EPL

[DH20] Robert P. Davis and Dale C. Honeyfield. Efficacy and practical limitations of calcein as a marking agent in lake trout (Salvelinus namaycush) exposed to sunlight and frozen sample storage. Fisheries Research, 232(??):Article 105736, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302538.

Dean:2021:DSS

[DHB^{+}21] Micah J. Dean, William S. Hoffman, Nicholas C. Buchan, Steven X. Cadrin, and Jonathan H. Grabowski. Deconstructing size selectivity to evaluate the influence of fishery management. Fisheries Research, 234(??):Article 105782, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030299X.

Ding:2020:OMA

[DHC^{+}20] Chengzhi Ding, Dekui He, Yifeng Chen, Yintao Jia, and Juan Tao. Otolith microstructure analysis based on wild young fish and its application in confirming the first annual increment in Tibetan Gymnocypris selincuoensis. Fisheries Research, 221(??):Article 105386, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302413.

Duskey:2023:PSD

[DHCS23] E. Duskey, D. R. Hart, J.-H. Chang, and P. J. Sullivan. Partitioning spatial dynamics in abundance of marine fisheries stocks between fine- and broad-scale variation: a Bayesian approach. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002096.

Delargy:2022:DQD

[DHH^{+}22] Adam J. Delargy, Natalie Hold, Charlotte Heney, Ruth P. Cann, Kiran Bhandari, Charlotte N. Colvin, Alec B. M. Moore, Harriet Lincoln, Ian D. McCarthy, and Jan G. Hiddink. Detection and quantification of differences in catch rates among research vessel gears and commercial vessels. Fisheries Research, 254(??):Article 106371, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001485.

Garcia-del-Hoyo:2023:PVT

[dHJTCE23] Juan José García del Hoyo, Ramón Jiménez-Toribio, and David Castilla-Espino. Price volatility transmission in the value chain of fresh anchovies in Spain. Fisheries Research, 260(??):Article 106606, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003836.

Dai:2023:PEC

[DHX^{+}23] Libin Dai, Cameron T. Hodgdon, Luoliang Xu, Jiaqi Wang, Siquan Tian, and Yong Chen. Performance evaluation of catch-only methods when catch data are misreported. Fisheries Research, 258(??):Article 106520, February 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002971.

Dikou:2024:DFA

[Dik24] Angela Dikou. Depletion fishing of the alien fish species Siganus luridus, S. rivulatus, Pterois miles, and Etrumeus golanii in the Mediterranean Sea-gear, ecosystem impacts, and resolution. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001590.

deJuan:2023:MSS

[dJDM23] Silvia de Juan, Gustav Delius, and Francesc Maynou. A model of size-spectrum dynamics to estimate the effects of improving fisheries selectivity and reducing discards in Mediterranean mixed demersal fisheries. Fisheries Research, 266(??): ??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001571.

Dunning:2023:NSM

[DJFF23] James Dunning, Teunis Jansen, Alan J. Fenwick, and Paul G. Fernandes. A new in-situ method to estimate fish target strength reveals high variability in broadband measurements. Fisheries Research, 261(??):Article 106611, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000048.

Dorow:2020:APR

[DJFU20] Malte Dorow, Jakob Jünger, Jens Frankowski, and Claus Ubl. Application of a 3-pass removal experiment to assess the yellow eel specific capture efficiency of a 1-ha enclosure. Fisheries Research, 221(??):Article 105409, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302644.

Dorow:2023:YEA

[DKBF23] Malte Dorow, Laura Kullmann, Melanie Buck, and Jens Frankowski. Yellow eel (Anguilla anguilla) density trends along the German part of the southern Baltic Sea between 2009 and 2020. Fisheries Research, 257(??):Article 106497, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002740.

DuFour:2021:HSS

[DKD^{+}21] Mark R. DuFour, Patrick M. Kocovsky, John Deller, Paul W. Simonin, and Lars G. Rudstam. Hydroacoustic survey standardization: Inter-vessel differences in fish densities and potential effects of vessel avoidance. Fisheries Research, 239(??):Article 105948, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100076X.

deLestang:2023:DIA

[dLH23] Simon de Lestang and Jason How. Data intermediate assessment: Examining the suitability of a novel depletion model for use in a spiny lobster fishery. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300245X.

deLestang:2024:BSR

[dLHER24] Simon de Lestang, Jason How, Christine Erbe, and Kelvin Rushworth. Boom, shake the room: Seismic surveys affect behaviour and survival of western rock lobster. Fisheries Research, 277(??):??, September 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362400136X.

Delargy:2022:PHV

[DLKH22] Adam J. Delargy, Gwladys I. Lambert, Michel J. Kaiser, and Jan G. Hiddink. Potential highly variable catch efficiency estimates complicate estimation of abundance. Fisheries Research, 245(??):Article 106138, January 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002666.

Dorval:2024:CBV

[DLP^{+}24] Emmanis Dorval, Kirk Lynn, Dianna Porzio, Trung Nguyen, and Katie Grady. Computing bias and variance for Pacific sardine (Sardinops sagax) biomass estimated from aerial surveys in California nearshore waters. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000638.

Hidalgo-de-la-Toba:2023:RPM

[dlTGPLC23] José Angel Hidalgo de-la Toba, Sergio Scarry González-Peláez, and Daniel B. Lluch-Cota. Recruitment patterns in Mexican geoduck (Panopea spp.) populations: Reconstruction from age structure and mortality parameters. Fisheries Research, 257(??):Article 106512, January 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002892.

Dan:2025:IWT

[DLZ^{+}25] Yating Dan, Bilin Liu, Leilei Zou, Jie Lu, and Linwei Song. Impact of water temperature experienced in early life of Dosidicus gigas on its adult growth. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624003242.

Dimarchopoulou:2021:ESI

[DMF^{+}21] Donna Dimarchopoulou, Peter J. Mous, Edwison Firmana, Elle Wibisono, Gianpaolo Coro, and Austin T. Humphries. Exploring the status of the Indonesian deep demersal fishery using length-based stock assessments. Fisheries Research, 243(??):Article 106089, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002174.

Despoti:2020:MSC

[DML^{+}20] Smaragda Despoti, Giacomo Milisenda, Alessandro Ligas, Luis Bentes, Francesc Maynou, Sergio Vitale, Germana Garofalo, Mario Sbrana, Karim Erzini, George Tserpes, Konstantinos Tsagarakis, Irida Maina, Maria-Myrto Pyrounaki, Nadia Papadopoulou, Athanassios Machias, Francesco Colloca, Fabio Fiorentino, Konstantinos I. Stergiou, and Marianna Giannoulaki. Marine spatial closures as a supplementary tool to reduce discards in bottom trawl fisheries: Examples from southern European waters. Fisheries Research, 232(??):Article 105714, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302319.

Diaz:2021:UIB

[DMM^{+}21] Manuel Díaz, Carlos Molinet, Thamara Matamala, Ricardo Diaz, Daniela Uribe, Patricio A. Diaz, Kurt Paschke, and Paulina Gebauer. Using a 3D image-based volumetric model to estimate fecundity in Lithodes santolla: a tool for improving Lithodidae crustacean monitoring. Fisheries Research, 234(??):Article 105803, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303209.

Dixon:2023:MPI

[DMM^{+}23] R. B. Dixon, T. S. Murray, B. Q. Mann, P. D. Cowley, and G. L. Jordaan. Movement patterns of the iconic giant trevally Caranx ignobilis from southern Africa, determined using tagrecapture data. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000863.

Derby:2024:DFT

[DMNV^{+}24] Charles D. Derby, Kevin M. Mesape, Hanh Ngo-Vu, Lisa C. Gentit, and Bryan A. Fluech. Developing and field testing sustainable synthetic baits for the blue crab (Callinectes sapidus) fishery. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001061.

Dorta:2022:FES

[DMS22] Carmelo Dorta and Pablo Martín-Sosa. Fishery essentiality: a short-term decision-making method based on economic viability as a tool to understand and manage data-limited smallscale fisheries. Fisheries Research, 246(??):Article 106171, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016578362100299X.

Duan:2021:MRG

[DMZH21] Youjian Duan, Charles P. Madenjian, Yingming Zhao, and Bin Huo. Modeling round goby growth in Lake Michigan and Lake Huron with multi-model inference. Fisheries Research, 236(??):Article 105842, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303593.

Demetras:2023:DCG

[DNLM23] Nicholas J. Demetras, T. Reid Nelson, Brendan M. Lehman, and Cyril J. Michel. Development of a castable, GPS enabled, Miniaturized Predation Event Recorder (mPER) to quantify predation of juvenile fishes. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000991.

Dahl:2024:OMF

[DOB^{+}24] Kristen Dahl, Joseph O’Malley, Beverly Barnett, Bill Kline, and Joseph Widdrington. Otolith morphometry and Fourier transform near-infrared (FT-NIR) spectroscopy as tools to discriminate archived otoliths of newly detected cryptic species, Etelis carbunculus and Etelis boweni. Fisheries Research, 272(??):??, April 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300320X.

Damasio:2020:SCB

[DPL20] Ludmila M. A. Damasio, Maria Grazia Peninno, and Priscila F. M. Lopes. Small changes, big impacts: Geographic expansion in small-scale fisheries. Fisheries Research, 226(??):Article 105533, June 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300503.

Dao:2023:CPF

[DQK^{+}23] Thang Dao, Martin Quaas, Dieter Koemle, Elias Ehrlich, and Robert Arlinghaus. Can price feedbacks cause human behavior-induced tipping points in exploited fish stocks? An extension of the bioeconomic Gordon–Schaefer model. Fisheries Research, 259(??):Article 106550, March 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003277.

DuFour:2021:EUR

[DQMV21] Mark R. DuFour, Song S. Qian, Christine M. Mayer, and Christopher S. Vandergoot. Embracing uncertainty to reduce bias in hydroacoustic species apportionment. Fisheries Research, 233(??):Article 105750, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302678.

Draper:2022:FHV

[Dra22] Douglas L. Draper. Food habit variability of arrowtooth flounder (Atheresthes stomias) along the U.S. west coast. Fisheries Research, 248(??):Article 106205, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003337.

Da-Rocha:2020:ICO

[DRSPTA20] José-María Da-Rocha, Jaume Sempere, Raúl Prellezo, and Luís Taboada-Antelo. Input controls and overcapitalization: a general equilibrium analysis of the Spanish Mediterranean Sea fisheries. Fisheries Research, 228(??):Article 105559, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030076X.

dosSantos:2021:RFS

[dS21] Paulo Roberto Santos dos Santos. Recreational fishing as a source for the monitoring of a critically endangered shark in southern Brazil. Fisheries Research, 241(??):Article 106006, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362100134X.

Deville:2021:STP

[DSB^{+}21] Diego Deville, Gustavo Sanchez, Sergio P. Barahona, Carmen Yamashiro, Daniel Oré-Chávez, Roger Quiroz Bazán, and Tetsuya Umino. Spatio-temporal patterns of genetic variation of the silverside Odontesthes regia in the highly productive Humboldt Current System. Fisheries Research, 244(??):Article 106127, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002551.

Damiano:2024:ETS

[DSC24] Matthew D. Damiano, Kyle W. Shertzer, and Jie Cao. Exploring tradeoffs in southeast United States marine fisheries management using management strategy evaluation. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000924.

Correa:2022:IFS

[dSCCC^{+}22] Gabriel Marcel de Souza Corrêa, Juliano Lauser Coletto, Jorge Pablo Castello, Nathaniel R. Miller, Rafael de Almeida Tubino, Cassiano Monteiro-Neto, and Marcus Rodrigues da Costa. Identification of fish stock based on otolith as a natural marker: the case of Katsuwonus pelamis (Linnaeus, 1758) in the Southwest Atlantic Ocean. Fisheries Research, 255(??):Article 106436, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002132.

Ding:2020:RUC

[DSJG20] Qi Ding, Xiujuan Shan, Xianshi Jin, and Harry Gorfine. Research on utilization conflicts of fishery resources and catch allocation methods in the Bohai Sea, China. Fisheries Research, 225(??):Article 105477, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303327.

Diaz-Suarez:2022:TSS

[DSNK^{+}22] Alfonso Diaz-Suarez, Kristina Noreikiene, Veljo Kisand, Oksana Burimski, Roland Svirgsden, Mehis Rohtla, Mikhail Ozerov, Riho Gross, Markus Vetemaa, and Anti Vasemägi. Temporally stable small-scale genetic structure of northern pike (Esox lucius) in the coastal Baltic Sea. Fisheries Research, 254(??):Article 106402, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001795.

DiCintio:2022:ASS

[DSP22] Antonio Di Cintio, Claudia Scianna, and Giulia Prato. Analysis of small-scale fisheries value chain: an interviewbased approach in Italian marine protected areas. Fisheries Research, 252(??):Article 106358, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001357.

Dash:2023:MFE

[DSP^{+}23] Gyanaranjan Dash, Swatipriyanka Sen, Rajesh Kumar Pradhan, Shubhadeep Ghosh, Jose Josileen, and Jayaraman Jayasankar. Modeling framework for establishing the power law between length and weight of fishes and a meta-analysis for validation of LWRs for six commercially important marine fishes from the northwestern Bay of Bengal. Fisheries Research, 257(??):Article 106496, January 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002739.

Filho:2020:MIR

[dSRFFN^{+}20] Luis Fernando da Silva Rodrigues Filho, Leonardo Manir Feitosa, Jorge Luiz Silva Nunes, Ana Rita Onodera Palmeira, Ana Paula Barbosa Martins, Tommaso Giarrizzo, Luís Fernando Carvalho-Costa, Iann Leonardo Pinheiro Monteiro, Romário Gemaque, Fernanda Gomes, Rosália Furtado C. Souza, Iracilda Sampaio, and João Bráullio de Luna Sales. Molecular identification of ray species traded along the Brazilian Amazon coast. Fisheries Research, 223(??):Article 105407, March 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302620.

Dassow:2023:DFR

[DSS^{+}23] Colin Dassow, Greg Sass, Stephanie Shaw, Zachary Feiner, Chelsey Nieman, and Stuart Jones. Depensation in fish recruitment driven by context-dependent interactions with another predator. Fisheries Research, 262(??):Article 106675, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000681.

deSantana:2020:RAG

[dSTMV20] Herick Soares de Santana, Claudenice Dei Tos, and Carolina Viviana Minte-Vera. A review on the age and growth studies of freshwater fish in South America. Fisheries Research, 222(??):Article 105410, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302656.

Desfosses:2022:IDB

[DTSR22] Cameron J. Desfosses, Alissa C. Tate, Claire B. Smallwood, and Karina L. Ryan. Improving design-based estimates of biological data collected from a restricted spatiotemporal access point survey of recreational fishers. Fisheries Research, 256(??):Article 106486, December 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002636.

Doering:2021:PON

[DWLT21] Kathryn L. Doering, Michael J. Wilberg, Dong Liang, and Mitchell Tarnowski. Patterns in oyster natural mortality in Chesapeake Bay, Maryland using a Bayesian model. Fisheries Research, 236(??):Article 105838, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303556.

Dimarchopoulou:2023:CCB

[DWS^{+}23] Donna Dimarchopoulou, Elle Wibisono, Steven Saul, Paul Carvalho, Angga Nugraha, Peter J. Mous, and Austin T. Humphries. Combining catch-based indicators suggests overexploitation and poor status of Indonesia’s deep demersal fish stocks. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002473.

EspinozaGuzman:2024:SFC

[EBdRPC24] Marco Antonio Espinoza Guzmán, Julliana W. Barretto, Maria del Rosario Pineda López, and Carlos Cruz Cruz. Sustainability of fishing cooperatives in the Gulf of Mexico: a case study. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001693.

Egerton:2021:UPF

[EBGE21] Jack P. Egerton, Derek G. Bolser, Arnaud Grüss, and Brad E. Erisman. Understanding patterns of fish backscatter, size and density around petroleum platforms of the U.S. Gulf of Mexico using hydroacoustic data. Fisheries Research, 233(??):Article 105752, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302691.

Ehrlich:2023:CDM

[EBN^{+}23] Elias Ehrlich, Alistair J. Bath, Dominique Niessner, Dieter Koemle, Eva-Maria Cyrus, and Robert Arlinghaus. Codesigning management recommendations with stakeholders: a case study about northern pike (Esox lucius) in the southern Baltic Sea. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000802.

Elmo:2021:VDA

[ECK^{+}21] Garrett M. Elmo, Derek P. Crane, Matthew E. Kimball, Kyle L. Williams, and Philip W. Stevens. Validity of daily and annual age estimation and back-calculation methods for early life stages of a subtropical-tropical species, the tarpon (Megalops atlanticus). Fisheries Research, 243(??):Article 106057, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001855.

Espinoza:2024:CSC

[ECM^{+}24] Raimundo Espinoza, Demian Chapman, Jack Morris, Sara N. Schoen, Valerie Hagan, Robert Hueter, Martin Soto, Rodolfo Abrams, and Paola Sotomayor. Characteristics and species composition of a small-scale shark fishery in Puerto Rico: Jurisdictional issues enable legal landings of prohibited and endangered species. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003296.

Estay:2021:RPA

[ECY21] Francisco Javier Estay, Nelson Colihueque, and Miguel Yáñez. Reproductive performance assessed during three spawning seasons in a naturalized rainbow trout population from southern Chile. Fisheries Research, 244 (??):Article 106107, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621002356.

Enever:2022:SPL

[EDA^{+}22] Robert Enever, Philip D. Doherty, Jon Ashworth, Mark Duffy, Pete Kibel, Melanie Parker, Bryce D. Stewart, and Brendan J. Godley. Scallop potting with lights: a novel, low impact method for catching European king scallop (Pecten maximus). Fisheries Research, 252(??):Article 106334, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001114.

Ern:2025:STD

[EFM25] Rasmus Ern, Andreas Faber, and Niels Madsen. Short-term discard survival and catch-related trauma in European plaice (Pleuronectes platessa) caught in the Baltic Sea by Danish seine during summer. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002686.

Emmerson:2020:ETG

[EHBJ20] J. A. Emmerson, P. R. Hollyman, I. S. M. Bloor, and S. R. Jenkins. Effect of temperature on the growth of the commercially fished common whelk (Buccinum undatum, L.): a regional analysis within the Irish Sea. Fisheries Research, 223(??):Article 105437, March 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302929.

Eklof:2023:ESS

[EHE^{+}23] Johan S. Eklöf, Joakim P. Hansen, Britas Klemens Eriksson, Örjan Östman, Åsa N. Austin, Casey Yanos, Ronny Fredriksson, Ulf Bergström, and Henrik C. Andersson. Effects of seasonal spawning closures on pike (Esox lucius L.) and perch (Perca fluviatilis L.) catches and coastal food webs in the western Baltic Sea. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362300067X.

Ekawaty:2020:CDS

[ELM20] Rani Ekawaty, John Lynham, and Peter Mous. Can demandside management replicate a size limit in a small-scale fishery? Fisheries Research, 223(??):Article 105436, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302917.

Ern:2023:IEA

[EM23] Rasmus Ern and Niels Madsen. Initial experiments to assess potential short-term discard survival of cod (Gadus morhua) in the Baltic Sea set-net fishery. Fisheries Research, 266(??):

??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001674.

Ern:2022:IEA

[EMJ^{+}22] Rasmus Ern, Katrine Molbo, Trine H. Jensen, Sergey V. Kucheryavskiy, Peter R. Møller, and Niels Madsen. Initial experiments to assess short-term survival of discarded plaice (Pleuronectes platessa) caught in trammel nets during winter season. Fisheries Research, 251(??):Article 106308, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000856.

Elmer:2022:TSD

[EMR^{+}22] Laura K. Elmer, David L. Moulton, Andrea J. Reid, Anthony P. Farrell, David A. Patterson, Brian Hendriks, Steven J. Cooke, and Scott G. Hinch. Thermal selection and delayed migration by adult sockeye salmon (Oncorhynchus nerka) following escape from simulated in-river fisheries capture. Fisheries Research, 251(??):Article 106321, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000984.

Evans-Powell:2024:IBO

[EPHDB24] Rubie T. Evans-Powell, Sybrand A. Hesp, Ainslie Denham, and Lynnath E. Beckley. Implications of big, old, fat, fecund, female fish (BOFFFFs) for the reproductive potential of a demersal teleost stock. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623003272.

Evans:2023:FAS

[ERS^{+}23] Thomas M. Evans, Lars G. Rudstam, Suresh A. Sethi, David M. Warner, S. Dale Hanson, Benjamin Turschak, Steven A. Farha, Andrew R. Barnard, Daniel L. Yule, Mark R. DuFour, Timothy P. O’Brien, Kevin N. McDonnell, James M. Watkins, Scott R. Koproski, Susan E. Wells, Patricia M. Dieter, Erik Kocher, James J. Roberts, Steven A. Senczyszyn, and Peter C. Esselman. Fish avoidance of ships during acoustic surveys tested with quiet uncrewed surface vessels. Fisheries Research, 267(??):??, November 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002102.

Erzini:2024:PFP

[ESB^{+}24] Karim Erzini, Zineb Sadat, Luís Bentes, Rui Coelho, Pedro G. Lino, Pedro Monteiro, Frederico Oliveira, Joaquim Ribeiro, and Jorge M. S. Gonçalves. The potential fish provisioning services of vegetated and unvegetated habitat in a lagoon nursery. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001796.

Elwer:2023:ESC

[EVS^{+}23] Brenden M. Elwer, Justin A. VanDeHey, Stephanie L. Shaw, Logan W. Sikora, Joseph T. Mrnak, and Greg G. Sass. Evaluation of survival and cost to harvestable age of stocked fall fingerling walleye (Sander vitreus) in Northern Wisconsin Lakes. Fisheries Research, 266(??):??, October 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001510.

Ellertson:2022:DSM

[EWPB22] Aubrey A. Ellertson, Jesica D. Waller, Tracy L. Pugh, and N. David Bethoney. Differences in the size at maturity of female American lobsters (Homarus americanus) from offshore Southern New England and eastern Georges Bank, USA. Fisheries Research, 250(??):Article 106276, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000534.

Eisner:2020:LCL

[EYAO20] Lisa B. Eisner, Ellen M. Yasumiishi, Alexander G. Andrews, and Cecilia A. O’Leary. Large copepods as leading indicators of walleye pollock recruitment in the southeastern Bering Sea: Sample-based and spatio-temporal model (VAST) results. Fisheries Research, 232(??):Article 105720, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030237X.

Freshwater:2024:MBI

[FAK24] Cameron Freshwater, Sean C. Anderson, and Jackie King. Model-based indices of juvenile Pacific salmon abundance highlight species-specific seasonal distributions and impacts of changes to survey design. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001279.

Follana-Berna:2022:MAS

[FBALRR^{+}22] Guillermo Follana-Berná, Pablo Arechavala-Lopez, Eduardo Ramirez-Romero, Elka Koleva, Amalia Grau, and Miquel Palmer. Mesoscale assessment of sedentary coastal fish density using vertical underwater cameras. Fisheries Research, 253(??):Article 106362, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001394.

Favaro:2020:CCP

[FBB20] Brett Favaro, Mary Alliston Butt, and Jonathan A. Bergshoeff. Comparison of catch per unit effort of invasive European green crab (Carcinus maenas) across four bait types. Fisheries Research, 225(??):Article 105484, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300011.

Fiorella:2021:SSF

[FBM^{+}21] Kathryn J. Fiorella, Elizabeth R. Bageant, Lia Mojica, Julia A. Obuya, Jane Ochieng, Pamela Olela, Patrick Wanguche Otuo, Horace Owiti Onyango, Christopher Mulanda Aura, and Harriet Okronipa. Small-scale fishing households facing COVID-19: the case of Lake Victoria, Kenya. Fisheries Research, 237(??):Article 105856, May 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303738.

Follana-Berna:2021:CTS

[FBPCC^{+}21] Guillermo Follana-Berná, Miquel Palmer, Andrea Campos-Candela, Josep Alós, Andrés Ospina-Alvarez, Amalia Grau, Susan Lowerre-Barbieri, and Pablo Arechavala-Lopez. Consequences of trait-selective fisheries on population reproductive potential: an experimental approach. Fisheries Research, 239(??):Article 105939, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000679.

Flores:2020:GDF

[FBQA20] Andrés Flores, Donald I. Brown, Dante Queirolo, and Mauricio Ahumada. Gonadal development of female red squat lobsters (Pleuroncodes monodon, H. Milne Edwards, 1837). Fisheries Research, 225(??):Article 105508, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300254.

Fromherz:2024:CSE

[FBR^{+}24] Matthias Fromherz, Jan Baer, Samuel Roch, Juergen Geist, and Alexander Brinker. Characterization of specialist European catfish anglers in southern Germany: Implications for future management. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362400208X.

Frawley:2021:FFI

[FBW^{+}21] Timothy H. Frawley, Hannah E. Blondin, Timothy D. White, Rachel R. Carlson, Brianna Villalon, and Larry B. Crowder. Fishers as foragers: Individual variation among smallscale fishing vessels as revealed by novel tracking technology. Fisheries Research, 238(??):Article 105896, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000242.

Farley:2022:BPA

[FCKG^{+}22] Jessica H. Farley, Naomi P. Clear, Kyne Krusic-Golub, J. Paige Eveson, and Jock W. Young. A bone to pick with age estimation using hard parts: a case study of swordfish, Xiphias gladius, in the southwest Pacific Ocean. Fisheries Research, 254(??):Article 106413, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001904.

Farthing:2022:RFG

[FCML^{+}22] M. W. Farthing, A. R. Childs, J. B. Mann-Lang, C. S. Bova, and W. M. Potts. Are recreational fishing guides role models for their clients? Fisheries Research, 254(??):Article 106408, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622001850.

Foster:2023:SSE

[FCMP23] R. M. Foster, A. R. Childs, B. Q. Mann, and W. M. Potts. Specialist spearfishers’ ecological knowledge provides accurate information that improves the description of the reproduction of a data-deficient species, Oplegnathus conwayi in South Africa. Fisheries Research, 257(??):Article 106513, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002909.

Fisch:2021:ALF

[FCSA21] Nicholas Fisch, Ed Camp, Kyle Shertzer, and Robert Ahrens. Assessing likelihoods for fitting composition data within stock assessments, with emphasis on different degrees of process and observation error. Fisheries Research, 243(??):Article 106069, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001971.

Filho:2020:UIA

[FDB^{+}20] Jorge L. Rodrigues Filho, Marina Dolbeth, Jurandir J. Bernardes, Jr., Igor Ogashawara, and Joaquim O. Branco. Using an integrative approach to evaluate shrimp bycatch from subtropical data-poor fisheries. Fisheries Research, 230(??):Article 105587, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301041.

Fujimoto:2020:TSS

[FDdCS^{+}20] Rodrigo Yudi Fujimoto, Henrique Malta Dias, Natalino da Costa Sousa, Marcia Valéria Silva do Couto, Rudã Fernandes Brandão Santos, Peterson Emmanuel Guimarães Paixão, Fernanda Dos Santos Cunha, Fabrício Menezes Ramos, Keid Nolan Silva Sousa, and Francisco Carlos Holanda. Is there sustainability for “satellite” ornamental fishing regions? A case study of Guamá River basin — Pará — Brasil. Fisheries Research, 221(??):Article 105354, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302097.

Fitzgerald:2023:LBA

[FDS^{+}23] Colm J. Fitzgerald, Jan S. Droll, Samuel Shephard, Christopher T. Monk, Timo Rittweg, and Robert Arlinghaus. Length-based assessment of an exploited coastal pike (Esox lucius) stock (Rügen, southern Baltic Sea) underscores the crucial relevance of growth and natural mortality for assessment outcomes. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000607.

Filous:2020:MJY

[FFG^{+}20] Alexander Filous, Alan M. Friedlander, Lucas Griffin, Robert J. Lennox, Andy J. Danylchuk, Geory Mereb, and Yimnang Golbuu. Movements of juvenile yellowfin tuna (Thunnus albacares) within the coastal FAD network adjacent to the Palau National Marine Sanctuary: Implications for local fisheries development. Fisheries Research, 230(??):Article 105688, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302058.

Felts:2021:ART

[FFG21] Eli A. Felts, Mark J. Fincel, and Brian D. S. Graeb. Angler reporting and tag retention estimates for walleye in Lake Oahe. Fisheries Research, 243(??):Article 106096, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002241.

Fey:2021:ETS

[FG21] Dariusz P. Fey and Martyna Greszkiewicz. Effects of temperature on somatic growth, otolith growth, and uncoupling in the otolith to fish size relationship of larval northern pike, Esox lucius L. Fisheries Research, 236(??):Article 105843, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030360X.

Fey:2023:TFN

[FG23] Dariusz P. Fey and Martyna Greszkiewicz. Transportation of fertilized northern pike eggs in a small mobile egg transportation system: Effect of time and comparison to a traditional method. Fisheries Research, 260(??):Article 106592, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003691.

Froehlich:2021:SDA

[FGCB^{+}21] Catheline Y. M. Froehlich, Andres Garcia, Carlos E. Cintra-Buenrostro, David W. Hicks, and Richard J. Kline. Structural differences alter residency and depth activity of red snapper (Lutjanus campechanus) at two artificial reefs. Fisheries Research, 242(??):Article 106043, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001715.

Fey:2022:SIS

[FGL22] Dariusz P. Fey, Martyna Greszkiewicz, and Adam M. Lejk. Stress induced by substantial skeletal deformities in pike fry is not reflected in otolith fluctuating asymmetry: an experiment and literature review. Fisheries Research, 254(??):Article 106387, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001643.

Fernandez-Gonzalez:2022:BCF

[FGPVPPGG22] Raquel Fernández-González, Raisa Pérez-Vas, Marcos Pérez-Pérez, and María Dolores Garza-Gil. Brexit, common fisheries policy and discard ban: a financial analysis of the Spanish fleet in the Grand Sole. Fisheries Research, 249(??):Article 106264, May 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000418.

Fey:2025:ADL

[FGSD25] Dariusz P. Fey, Martyna Greszkiewicz, Lena Szymanek, and Małgorzata Dembek. Abundance and distribution of larval and early juvenile fish in Puck Lagoon and outer Puck Bay in the southern Baltic Sea. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002996.

Fall:2024:CAS

[FGTA24] Johanna Fall, Harald Gjøsæter, Ingunn Fride Tvete, and Magne Aldrin. Classification of acoustic survey data: a comparison between seven teams of experts. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000699.

Fujinami:2024:AGR

[FHE^{+}24] Yuki Fujinami, Yuko Hiraoka, Ryousei Ebisu, Nobuyo Matsushima, Takeshi Itou, Kenya Matsui, Hiroyuki Kurota, and Toshiyuki Tanabe. Age, growth and reproduction of Japanese Spanish mackerel (Scomberomorus niphonius) in the coastal area off northern and western Kyushu, Japan. Fisheries Research, 275(??):??, July 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000754.

Freshwater:2020:BLI

[FHHH20] Cameron Freshwater, Kendra R. Holt, Ann-Marie Huang, and Carrie A. Holt. Benefits and limitations of increasing the stock-selectivity of Pacific salmon fisheries. Fisheries Research, 226(??):Article 105509, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300266.

Fang:2021:OVT

[FHSC21] Zhou Fang, Peiwu Han, Siwenjia Shi, and Xinjun Chen. Ontogenetic variation of trophic habitat for sympatric benthic octopods in East China Sea derived from isotopic analysis on beaks. Fisheries Research, 238(??):Article 105902, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000308.

Fisch:2025:EIP

[Fis25] Nicholas Fisch. Expected improvements in precision when integrating opportunistic close-kin mark-recapture data into fisheries stock assessments. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002868.

Free:2020:BSP

[FJA^{+}20] Christopher M. Free, Olaf P. Jensen, Sean C. Anderson, Nicolas L. Gutierrez, Kristin M. Kleisner, Catherine Longo, Cóilín Minto, Giacomo Chato Osio, and Jessica C. Walsh. Blood from a stone: Performance of catchonly methods in estimating stock biomass status. Fisheries Research, 223(??):Article 105452, March 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303078.

Fink-Jensen:2022:LRC

[FJHT^{+}22] Peter Fink-Jensen, Karin Hüssy, Tonny Bernt Thomsen, Simon Hansen Serre, Jens Søndergaard, and Teunis Jansen. Lifetime residency of capelin (Mallotus villosus) in West Greenland revealed by temporal patterns in otolith microchemistry. Fisheries Research, 247(??):Article 106172, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003003.

Fink-Jensen:2021:MCV

[FJJT^{+}21] Peter Fink-Jensen, Teunis Jansen, Tonny Bernt Thomsen, Simon Hansen Serre, and Karin Hüssy. Marine chemistry variation along Greenland’s coastline indicated by chemical fingerprints in capelin (Mallotus villosus) otoliths. Fisheries Research, 236(??):Article 105839, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303568.

Fedewa:2025:BCD

[FJS^{+}25] Erin J. Fedewa, Pamela C. Jensen, Hamish J. Small, Michael A. Litzow, Michael J. Malick, Leah S. Zacher, W. Christopher Long, and Stan Kotwicki. Bitter crab disease dynamics in eastern Bering Sea Tanner and snow crab: an underestimated and emergent stressor. Fisheries Research, 283(??):??, March 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362500044X.

Freed:2020:RFF

[FKS^{+}20] Sarah Freed, Yumiko Kura, Vichet Sean, Samonn Mith, Philippa Cohen, Miratori Kim, Somony Thay, and Savry Chhy. Rice field fisheries: Wild aquatic species diversity, food provision services and contribution to inland fisheries. Fisheries Research, 229(??):Article 105615, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301326.

Fu:2022:STD

[FKW^{+}22] Xiu-Mei Fu, Hong-Li Ku, Wan-Yu Wu, Li-Xia Wang, Han-Xue Chen, Chun-Yu Lin, and Ying Liu. Spatialtemporal differentiation and convergence analysis of marine fishery innovation ability in China. Fisheries Research, 254(??):Article 106393, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001709.

Faruque:2021:AVB

[FM21] Hasan Faruque and Hiroyuki Matsuda. Assessing the vulnerability of bycatch species from Hilsa gillnet fishing using productivity susceptibility analysis: Insights from Bangladesh. Fisheries Research, 234(??):Article 105808, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303258.

Fauconnet:2024:FAC

[FMD^{+}24] Laurence Fauconnet, Telmo Morato, Diya Das, Diana Catarino, Jorge Fontes, Eva Giacomello, and Pedro Afonso. First assessment of circle hooks as bycatch mitigation measure for deep-water sharks on longline fisheries. Fisheries Research, 270(??):??, February 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002709.

Farthing:2022:AFG

[FMLC^{+}22] M. W. Farthing, J. Mann-Lang, A. R. Childs, C. S. Bova, S. D. Bower, A. C. Pinder, K. Ferter, A. C. Winkler, E. C. Butler, J. W. Brownscombe, A. J. Danylchuk, and W. M. Potts. Assessment of fishing guide knowledge, attitudes, and behaviours in global recreational fisheries. Fisheries Research, 255(??):Article 106453, November 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002302.

Figueiredo:2020:BER

[FMLSP20] Ivone Figueiredo, Catarina Maia, Neide Lagarto, and Bárbara Serra-Pereira. Bycatch estimation of Rajiformes in multispecies and multigear fisheries. Fisheries Research, 232(??):Article 105727, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302447.

Fontes:2020:WST

[FMMA20] Jorge Fontes, Niall McGinty, Miguel Machete, and Pedro Afonso. Whale shark-tuna associations, insights from a small pole-and-line fishery from the mid-north Atlantic. Fisheries Research, 229(??):Article 105598, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301156.

Ford:2020:MSR

[FMMC20] J. Ford, D. Maxwell, E. W. Muiruri, and T. Catchpole. Modifying selectivity to reduce unwanted catches in an English trammel net and gill net common sole fishery. Fisheries Research, 227(??):Article 105531, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300485.

Fadli:2021:DBC

[FMSA21] Nur Fadli, Zainal A. Muchlisin, and Mohd N. Siti-Azizah. DNA barcoding of commercially important groupers (Epinephelidae) in Aceh, Indonesia. Fisheries Research, 234(??):Article 105796, February 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303131.

Flink:2021:EEA

[FNH^{+}21] Henrik Flink, Oscar Nordahl, Marcus Hall, Anton Rarysson, Kristofer Bergström, Per Larsson, Erik Petersson, Juha Merilä, and Petter Tibblin. Examining the effects of authentic C&R on the reproductive potential of northern pike. Fisheries Research, 243(??):Article 106068, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100196X.

Furuichi:2020:TVR

[FNKY20] Sho Furuichi, Yohei Niino, Yasuhiro Kamimura, and Ryuji Yukami. Time-varying relationships between early growth rate and recruitment in Japanese sardine. Fisheries Research, 232(??):Article 105723, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030240X.

Farriols:2021:DRN

[FOM21] M. Teresa Farriols, Francesc Ordines, and Enric Massutí. Discards reduction of non-commercial benthic species from a simple net modification. Fisheries Research, 241 (??):Article 105985, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621001132.

Fernandes:2021:IRO

[FOS^{+}21] Ana Cláudia Fernandes, Melinda Oroszlányová, Cristina Silva, Manuela Azevedo, and Rui Coelho. Investigating the representativeness of onboard sampling trips and estimation of discards based on clustering. Fisheries Research, 234(??):Article 105778, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302952.

Fredenslund:2022:IOC

[Fre22] Theresa Fredenslund. Identifying overexploitation in the coastal Greenland halibut fishery in the Disko Bay using static bioeconomic modelling. Fisheries Research, 254 (??):Article 106417, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622001941.

Finn:2022:CPD

[FRP22] Kaegan J. Finn, Karling N. Roberts, and Mark S. Poesch. Cestode parasites are depleted in ^{15}N relative to their fish hosts in northern Alberta, Canada. Fisheries Research, 248(??):Article 106193, April 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003210.

Frater:2020:CEA

[FS20] Paul N. Frater and Gunnar Stefansson. Comparison and evaluation of approaches aimed at correcting or reducing selectivity bias in growth parameter estimates for fishes. Fisheries Research, 225(??):Article 105464, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303194.

Fey:2022:PLH

[FSP22] Dariusz P. Fey and Agnieszka Szkudlarek-Pawełczyk. Precision of larval herring abundance and body length assessments when sampling with BONGO nets for spawning site mapping. Fisheries Research, 246(??):Article 106165, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002939.

Freitas:2024:STP

[FSS^{+}24] Vânia Freitas, Diana Santos, Diogo M. Silva, Jacinto Cunha, Sabrina M. Rodrigues, Vanessa Neves, Eduardo Rocha, Filipe Martinho, and Sandra Ramos. Spatial and temporal patterns of gonadal maturation and spawning in European flounder Platichthys flesus at its southern continental edge. Fisheries Research, 269(??):??, January 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002576.

Fahmi:2021:PSB

[FTB^{+}21] Fahmi, I. R. Tibbetts, M. B. Bennett, A. Ali, T. Krajangdara, and C. L. Dudgeon. Population structure of the brownbanded bamboo shark, Chiloscyllium punctatum and its relation to fisheries management in the Indo–Malay region. Fisheries Research, 240(??):Article 105972, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001004.

Flink:2023:VAB

[FTH^{+}23] Henrik Flink, Petter Tibblin, Marcus Hall, Gustav Hellström, and Oscar Nordahl. Variation among bays in spatiotemporal aggregation of Baltic Sea pike highlights management complexity. Fisheries Research, 259(??):Article 106579, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003563.

Fu:2022:LBE

[Fu22] Dan Fu. A length-based extension to the Brownie– Peterson model and its performance. Fisheries Research, 249(??):Article 106248, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362200025X.

French:2021:MSE

[FWKR21] Ben French, Shaun Wilson, Alan Kendrick, and Michael Rule. The mesh size effect: counting long thin fish in seagrass. Fisheries Research, 242(??):Article 106019, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001478.

Feiner:2020:WLY

[FWL20] Zachary S. Feiner, Max H. Wolter, and Alexander W. Latzka. “I will look for you, I will find you, and I will [harvest] you”: Persistent hyperstability in Wisconsin’s recreational fishery. Fisheries Research, 230(??):Article 105679, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030196X.

Gundelund:2020:IUC

[GAB^{+}20] Casper Gundelund, Robert Arlinghaus, Henrik Baktoft, Kieran Hyder, Paul Venturelli, and Christian Skov. Insights into the users of a citizen science platform for collecting recreational fisheries data. Fisheries Research, 229(??):Article 105597, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301144.

Garner:2022:MAE

[GAB^{+}22a] Steven B. Garner, Robert Ahrens, Kevin M. Boswell, Matthew D. Campbell, Daniel Correa, Joseph H. Tarnecki, and William F. Patterson. A multidisciplinary approach to estimating red snapper, Lutjanus campechanus, behavioral response to mobile camera and sonar sampling gears. Fisheries Research, 246(??):Article 106155, February 2022.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002836.

Gundelund:2022:IAS

[GAB^{+}22b] Casper Gundelund, Robert Arlinghaus, Max Birdsong, Hugo Flávio, and Christian Skov. Investigating angler satisfaction: the relevance of catch, motives and contextual conditions. Fisheries Research, 250(??):Article 106294, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000716.

Garlock:2024:PVC

[GAB^{+}24] Taryn Garlock, Frank Asche, Casey B. Butler, Thomas R. Matthews, and Erica Ross. Price variation in the Caribbean spiny lobster fishery: Incentives for ongrowing wild-caught lobsters in Florida. Fisheries Research, 273(??):??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000249.

Green:2022:UCA

[GAW^{+}22] M. E. Green, S. A. Appleyard, W. T. White, S. R. Tracey, M. R. Heupel, and J. R. Ovenden. Updated connectivity assessment for the scalloped hammerhead (Sphyrna lewini) in Pacific and Indian Oceans using a multi-marker genetic approach. Fisheries Research, 251(??):Article 106305, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000820.

Gugele:2020:SDI

[GBB20] S. M. Gugele, J. Baer, and A. Brinker. The spatiotemporal dynamics of invasive three-spined sticklebacks in a large, deep lake and possible options for stock reduction. Fisheries Research, 232(??):Article 105746, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302630.

Goethel:2023:SAG

[GBC23a] Daniel R. Goethel, Aaron M. Berger, and Steven X. Cadrin. Spatial awareness: Good practices and pragmatic recommendations for developing spatially structured stock assessments. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000966.

Gray:2023:LWR

[GBC^{+}23b] Benjamin C. T. Gray, Maria Byrne, Matthew Clements, Shawna A. Foo, and Steven W. Purcell. Length–weight relationship for the dragonfish, Stichopus cf. monotuberculatus (Holothuroidea). Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002448.

Guzman:2020:SIC

[GBO^{+}20] Fabián Guzmán, Miguel Bascur, Luis Olavarria, Sergio Mora, Rodrigo Riera, and Ángel Urzúa. Seasonal and interannual changes in reproductive parameters and eggs biochemical composition of the fishery resource Pleuroncodes monodon (Decapoda: Munididae) from the Humboldt Current System. Fisheries Research, 221(??):Article 105404, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302590.

Gordon:2022:BEK

[GBWM22] Jesse Y. Gordon, Anne H. Beaudreau, Benjamin C. Williams, and Scott C. Meyer. Bridging expert knowledge and fishery data to examine changes in nearshore rockfish fisheries in the Gulf of Alaska over fifty years. Fisheries Research, 252(??):Article 106333, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001102.

Goldstein:2021:AIL

[GC21] Jason S. Goldstein and Joshua T. Carloni. Assessing the implications of live claw removal on Jonah crab (Cancer borealis), an emerging fishery in the Northwest Atlantic. Fisheries Research, 243(??):Article 106046, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001740.

Gasper:2025:URF

[GC25] Jason Gasper and Jennifer Cahalan. Utilizing the random forest algorithm and interpretable machine learning to inform post-stratification of commercial fisheries data. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003175.

Ganias:2021:FCT

[GCK^{+}21] Konstantinos Ganias, Georgios Christidis, Irini-Fotini Kompogianni, Xeni Simeonidou, Eleni Voultsiadou, and Chryssanthi Antoniadou. Fishing for cuttlefish with traps and trammel nets: a comparative study in Thermaikos Gulf, Aegean Sea. Fisheries Research, 234(??):Article 105783, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303003.

Guerra:2021:DBH

[GdSPL21] Thiago Pereira Guerra, Josiene Maria Falcão Fraga dos Santos, Maria Grazia Pennino, and Priscila Fabiana Macedo Lopes. Damage or benefit? How future scenarios of climate change may affect the distribution of small pelagic fishes in the coastal seas of the Americas. Fisheries Research, 234(??):Article 105815, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303325.

Garcia-De-Vinuesa:2020:SVS

[GDVBB^{+}20] Alfredo García-De-Vinuesa, Mike Breen, Hugues P. Benoît, Francesc Maynou, and Montserrat Demestre. Seasonal variation in the survival of discarded Nephrops norvegicus in a NW Mediterranean bottom-trawl fishery. Fisheries Research, 230(??):Article 105671, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301880.

Griffin:2022:STR

[GFC^{+}22] Lucas P. Griffin, Gail Fordham, George Curd, Christopher Narty, Pierre-André Adam, Jacob W. Brownsombe, Steven J. Cooke, and Andy J. Danylchuk. Short-term response of giant trevally (Caranx ignobilis) to capture and handling in a catch-and-release fly fishing recreational fishery, Republic of the Seychelles, Western Indian Ocean. Fisheries Research, 252(??):Article 106337, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200114X.

Gonzalez:2022:AVG

[GFDN^{+}22] Júlio Guazzelli Gonzalez, Thierry Frédou, Paulo José Duarte-Neto, Cécile Petit, Maylis Labonne, Rosângela Paula Lessa, and Audrey M. Darnaude. Age validation and growth in an exploited but poorly studied tropical fish species, the horse-eye jack (Caranx latus). Fisheries Research, 253(??):Article 106368, September 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200145X.

Gunton:2023:UOB

[GFM^{+}23] Holly Gunton, Ashley M. Fowler, Marcus E. Miller, David J. Booth, and John Stewart. The utility of otolith-based methods for discriminating stock structure of the fishery important monacanthid, Nelusetta ayraud, over an intermediate scale. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300108X.

Gaertner:2022:TSR

[GGG^{+}22] D. Gaertner, L. Guéry, N. Goñi, J. Amande, P. Pascual Alayon, F. N’Gom, J. Pereira, E. Addi, L. Ailloud, and D. Beare. Tag-shedding rates for tropical tuna species in the Atlantic Ocean estimated from double-tagging data. Fisheries Research, 248(??):Article 106211, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003398.

Goldstein:2024:LDS

[GGL^{+}24] Jason S. Goldstein, Benjamin C. Gutzler, Abigail Lemmon, Joshua T. Carloni, Marlies Betka, and Steven H. Jury. Long distance swimmers in warming waters: Active transport in postlarval American lobsters (Homarus americanus) in the context of climate change in the Gulf of Maine. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001760.

Gonzalez-Gomez:2020:RDP

[GGMMMV^{+}20] Roberto González-Gómez, César Meiners-Mandujano, Piedad S. Morillo-Velarde, Lourdes Jiménez-Badillo, and Unai Markaida. Reproductive dynamics and population structure of Octopus insularis from the Veracruz Reef System Marine Protected Area, Mexico. Fisheries Research, 221(??):Article 105385, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619302401.

Gonzalez-Garcia:2022:STT

[GGMRC^{+}22] E. González-García, Á. Mateo-Ramírez, M. P. Maroto Castaño, G. Bruque, C. Farias, N. López-González, A. Punzón, and J. L. Rueda. Spatio-temporal trends of the bottom trawling activity in a mud volcano field of the north-eastern Gulf of Cádiz (south-western Iberian Peninsula). Fisheries Research, 254(??):Article 106420, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001977.

Garces-Garcia:2020:EBP

[GGTÁVT^{+}20] Karla C. Garcés-García, Javier Tovar-Ávila, Bibiana Vargas-Trejo, Darío A. Chávez-Arrenquín, Terence I. Walker, and Robert W. Day. Elasmobranch bycatch by prawn trawls in the Gulf of California: First comprehensive analysis and the effect of fish escape devices. Fisheries Research, 230(??):Article 105639, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301569.

Guerrero:2021:TUR

[GHAZ21] Daniel Guerrero, Atakelty Hailu, Jose Santiago Arroyo, and Luis Alonso Zapata. Territorial Use Rights for Fisheries (TURF) and self-regulation of behaviour: Experimental evidence from the Colombian Pacific coast. Fisheries Research, 240(??):Article 105946, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000746.

Gutzler:2023:NMQ

[GJGW23] Benjamin C. Gutzler, Steven H. Jury, Jason S. Goldstein, and Winsor H. Watson. A novel method for quantifying the attractiveness of lobster trap baits. Fisheries Research, 267(??):??, November 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002084.

Good:2022:UBT

[GJSW22] Thomas P. Good, Jason E. Jannot, Kayleigh A. Somers, and Eric J. Ward. Using Bayesian time series models to estimate bycatch of an endangered albatross. Fisheries Research, 256(??):Article 106492, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002697.

Ganias:2021:HLP

[GKC21] Konstantinos Ganias, Irini-Fotini Kompogianni, and Georgios Christidis. How lure prevalence, survival rate and maturity status affect catchability in cuttlefish trap fishery. Fisheries Research, 238(??):Article 105919, June 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000473.

Giulietti:2022:LTI

[GKC^{+}22] Lucilla Giulietti, Egil Karlsbakk, Paolo Cipriani, Miguel Bao, Julia E. Storesund, Nachiket P. Marathe, and Arne Levsen. Long-term investigation of the ‘soft flesh’ condition in Northeast Atlantic mackerel induced by the myxosporean parasite Kudoa thyrsites (Cnidaria, Myxozoa): Temporal trends and new molecular epidemiological observations. Fisheries Research, 248(??):Article 106221, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003490.

Glemarec:2024:DFG

[GKKL24] Gildas Glemarec, Anne-Mette Kroner, and Lotte Kindt-Larsen. Disappearing fish: Grey seal depredation in a Baltic net fishery. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001346.

Gunnarsson:2023:EFF

[GKM^{+}23] Á. Gunnarsson, J. Kennedy, Á. Magnússon, B. Bárdarson, and B. Elvarsson. Effect of formalin fixation on size and weight of Atlantic wolffish (Anarhichas lupus) and spotted wolffish (Anarhichas minor) oocytes. Fisheries Research, 258(??):Article 106515, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002922.

Giacomini:2020:GCW

[GLA^{+}20] Henrique Corrêa Giacomini, Nigel Lester, Pete Addison, Steve Sandstrom, Daniel Nadeau, Cindy Chu, and Derrick de Kerckhove. Gillnet catchability of walleye (Sander vitreus): comparison of North American and provincial standards. Fisheries Research, 224(??):Article 105433, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302887.

Ghinter:2021:JGH

[GLA21] Leopold Ghinter, Yvan Lambert, and Céline Audet. Juvenile Greenland halibut (Reinhardtius hippoglossoides) growth in the context of rising temperature in the Estuary and Gulf of St. Lawrence. Fisheries Research, 233(??):Article 105766, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302836.

Guyomard:2020:SDR

[GLP^{+}20] D. Guyomard, K. A. Lee, C. Perry, S. Jaquemet, and G. Cliff. SMART drumlines at Réunion Island do not attract bull sharks Carcharhinus leucas into nearshore waters: Evidence from acoustic monitoring. Fisheries Research, 225(??):Article 105480, May 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303352.

Gertseva:2021:RTU

[GM21] Vladlena V. Gertseva and Sean E. Matson. Right on target: using data from targeted stocks to reconstruct removals of bycatch species, a case study of longnose skate from Northeast Pacific Ocean. Fisheries Research, 236(??):Article 105841, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303581.

Guerra-Marrero:2024:FEA

[GMERCM^{+}24] Airam Guerra-Marrero, Ana Espino-Ruano, Lorena Couce-Montero, David Jiménez-Alvarado, and José J. Castro. Feeding ecology of the African cuttlefish Sepia bertheloti (Cephalopoda: Sepiidae) in western Africa. Fisheries Research, 269(??):??, January 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002692.

Ganias:2023:ELC

[GMK23] Konstantinos Ganias, Glykeria Malioufa, and Martha Kaloyanni. Evaluating the levels of capture-related stress and physical injury in métiers that use gillnets and trammel nets. Fisheries Research, 267(??):??, November 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002072.

Gruss:2023:SSA

[GML^{+}23] Arnaud Grüss, Jeremy R. McKenzie, Martin Lindegren, Richard Bian, Simon D. Hoyle, and Jennifer A. Devine. Supporting a stock assessment with spatio-temporal models fitted to fisheries-dependent data. Fisheries Research, 262(??):Article 106649, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623000425.

Gruss:2023:UST

[GMPD23] Arnaud Grüss, Bradley R. Moore, Matthew H. Pinkerton, and Jennifer A. Devine. Understanding the spatio-temporal abundance patterns of the major bycatch species groups in the Ross Sea region Antarctic toothfish (Dissostichus mawsoni) fishery. Fisheries Research, 262(??):Article 106647, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000401.

Garcia-Mayoral:2020:SDG

[GMRRG20] Elsa García-Mayoral, Álvaro Roura, Andrea Ramilo, and Ángel F. González. Spatial distribution and genetic structure of loliginid paralarvae along the Galician coast (NW Spain). Fisheries Research, 222(??):Article 105406, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302619.

Gertseva:2024:TTS

[GMT24] Vladlena Gertseva, Sean E. Matson, and Ian G. Taylor. A tale of two species: Disaggregating mixed historical catches of two most common skates in the Northeast Pacific Ocean. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000675.

Gonzalez:2021:ASC

[GP21] Ángel F. González and Graham J. Pierce. Advances in the study of cephalopod fisheries and ecosystems. Fisheries Research, 242(??):Article 105975, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100103X.

Ganias:2025:EEV

[GP25] Kostas Ganias and Sofia Papandreou. Evaluating the efficacy of various types of guarding nets in minimizing bycatch in a Mediterranean trammel net fishery. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003084.

Gonzalez-Pestana:2022:RHT

[GPASM22] Adriana Gonzalez-Pestana, Joanna Alfaro-Shigueto, and Jeffrey C. Mangel. A review of high trophic predatorprey relationships in the pelagic Northern Humboldt system, with a focus on anchovetas. Fisheries Research, 253(??):Article 106386, September 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001631.

Gruss:2021:MNF

[GPT^{+}21] Arnaud Grüss, Jodi L. Pirtle, James T. Thorson, Mandy R. Lindeberg, A. Darcie Neff, Steve G. Lewis, and Timothy E. Essington. Modeling nearshore fish habitats using Alaska as a regional case study. Fisheries Research, 238(??):Article 105905, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000333.

Geng:2020:DID

[GPW^{+}20] Zhe Geng, André E. Punt, Yang Wang, Jiangfeng Zhu, and Xiaojie Dai. On the dangers of including demographic analysis in Bayesian surplus production models: a case study for Indian Ocean blue shark. Fisheries Research, 230(??):Article 105636, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301533.

Garner:2020:SEH

[GPWP20] Steven B. Garner, William F. Patterson, John F. Walter, and Clay E. Porch. Simulating effects of hook-size regulations on recreational harvest efficiency in the northern Gulf of Mexico red snapper fishery. Fisheries Research, 228(??):Article 105561, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300783.

Gaffney:2025:VOH

[GQP^{+}25] Leigh P. Gaffney, Micah Quindazzi, Emma Polard, Chloe Kraemer, Lydia N. Walton, Zoe A. Molder, Wesley L. Greentree, Will Duguid, Nick Bohlender, and Francis Juanes. Vateritic otoliths in hatchery-reared strait of Georgia coho salmon: Variation among stocks, hatcheries, and life stages. Fisheries Research, 283(??):??, March 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000335.

Grew:2024:BRB

[GRG24] M. Grew, V. Raoult, and T. F. Gaston. Behavioural response of benthic elasmobranchs to a neodymium magnet under controlled laboratory conditions. Fisheries Research, 271(??): ??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623003193.

Garcia-Rodriguez:2020:ESS

[GRHHGM^{+}20] Alain García-Rodríguez, Agustín Hernández-Herrera, Felipe Galván-Magaña, Bertha Patricia Ceballos-Vázquez, Tania Pelamatti, and Javier Tovar-Ávila. Estimation of the size at sexual maturity of the bat ray (Myliobatis californica) in northwestern Mexico through a multi-model inference. Fisheries Research, 231(??):Article 105712, November 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302290.

Gruss:2020:MMA

[GRJW20] Arnaud Grüss, Kenneth A. Rose, Dubravko Justić, and Lixia Wang. Making the most of available monitoring data: a grid-summarization method to allow for the combined use of monitoring data collected at random and fixed sampling stations. Fisheries Research, 229(??):Article 105623, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301405.

Gross:2022:EPN

[GSH22] Julie M. Gross, Philip Sadler, and John M. Hoenig. Evaluating a possible new paradigm for recruitment dynamics: predicting poor recruitment for striped bass (Morone saxatilis) from an environmental variable. Fisheries Research, 252(??):Article 106329, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001060.

Gibson:2023:CSG

[GSS^{+}23] G. A. Gibson, W. T. Stockhausen, K. Shotwell, A. L. Deary, J. L. Pirtle, K. O. Coyle, and A. J. Hermann. Can seamounts in the Gulf of Alaska be a spawning ground for sablefish settling in coastal nursery grounds? Fisheries Research, 261(??):Article 106625, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000188.

Gunther:2021:SSH

[GTS^{+}21] Claudia Günther, Axel Temming, Juan Santos, Jörg Berkenhagen, Daniel Stepputtis, Sebastian Schultz, Thomas Neudecker, Gerd Kraus, Eckhard Bethke, and Marc Hufnagl. Small steps high leaps: Bio-economical effects of changing codend mesh size in the North Sea brown shrimp fishery. Fisheries Research, 234(??):Article 105797, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303143.

G:2023:FCS

[GVK^{+}23] Mini K. G., Sathianandan T. V., Somy Kuriakose, Sindhu K. Augustine, Manu V. K., Manjeesh R., Sijo Paul, Jayasankar J., Eldho Varghese, and Gopalakrishnan A. Fish Catch Survey and Analysis — an online application for deriving measures and indicators for fish stock assessment. Fisheries Research, 267(??):??, November 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300214X.

Grausgruber:2021:UMR

[GW21] Emily E. Grausgruber and Michael J. Weber. Using markrecapture techniques to assess factors affecting survival of fall stocked walleye Sander vitreus in two Iowa lakes. Fisheries Research, 242(??):Article 106029, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001570.

Grausgruber:2024:ESY

[GWGM24] Stephen J. Grausgruber, Michael J. Weber, Emily E. Grausgruber, and Joseph E. Morris. Exploitation and survival of yellow perch introduced into community ponds. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000766.

Gonzalez:2021:EST

[GWMC21] Guillermo Martin Gonzalez, Rodrigo Wiff, C. Tara Marshall, and Thomas Cornulier. Estimating spatio-temporal distribution of fish and gear selectivity functions from pooled scientific survey and commercial fishing data. Fisheries Research, 243(??):Article 106054, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100182X.

Gallego-Zerrato:2025:ESC

[GZLCRG25] Juan José Gallego-Zerrato, Brenda Natalia Londoño, Diego Fernando Córdoba-Rojas, and Alan Giraldo. Evaluation of the swimming crab (Callinectes toxotes Ordway, 1863) as an alternative resource to traditional fishing in the Colombian Pacific. Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000402.

Hoyli:2023:SEU

[HA23] Randulf Høyli and Karl Gunnar Aarsæther. A study of energy use and associated greenhouse gas emissions in Norwegian small-scale processing of whitefish. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002357.

Hamaguchi:2022:WER

[Ham22] Yoshihiro Hamaguchi. Welfare effect of rent-seeking activities under international management of fishery resources. Fisheries Research, 246(??):Article 106170, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002988.

Harrison:2021:MMN

[Har21] Hannah L. Harrison. Managing many nets: Possible scenarios and impacts for the expansion of Cook Inlet personal use fisheries. Fisheries Research, 236(??):Article 105811, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303283.

Hansell:2022:STD

[HBC^{+}22] Alexander C. Hansell, Sarah L. Becker, Steven X. Cadrin, Matthew Lauretta, John F. Walter III, and Lisa A. Kerr. Spatio-temporal dynamics of bluefin tuna (Thunnus thynnus) in US waters of the northwest Atlantic. Fisheries Research, 255(??):Article 106460, November 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002375.

Hoenner:2022:UAT

[HBE^{+}22] X. Hoenner, E. Barlian, T. Ernawati, B. D. Hardesty, D. D. Kembaren, P. J. Mous, L. Sadiyah, F. Satria, and C. Wilcox. Using anti-theft tracking devices to infer fishing vessel activity at sea. Fisheries Research, 249(??):Article 106230, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000078.

Hwang:2021:EVF

[HBMC21] Joonghyun Hwang, Xiang Bi, Nia Morales, and Edward V. Camp. The economic value of freshwater fisheries in Florida: an application of the travel cost method for black crappie fishing trips. Fisheries Research, 233(??):Article 105754, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030271X.

Hammarlund:2021:LMP

[HBW21] Cecilia Hammarlund, Johan Blomquist, and Staffan Waldo. Local markets and price premiums — the case of the establishment of the Stockholm fish auction. Fisheries Research, 236(??):Article 105853, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303702.

Hamel:2022:DCA

[HC22a] Owen S. Hamel and Jason M. Cope. Development and considerations for application of a longevity-based prior for the natural mortality rate. Fisheries Research, 256 (??):Article 106477, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622002545.

Hart:2022:ENM

[HC22b] Deborah R. Hart and Jui-Han Chang. Estimating natural mortality for Atlantic sea scallops (Placopecten magellenicus) using a size-based stock assessment model. Fisheries Research, 254(??):Article 106423, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002004.

Hyun:2022:NSF

[HC22c] Saang-Yoon Hyun and Curry J. Cunningham. A new inseason forecast density of anadromous fish return abundance. Fisheries Research, 256(??):Article 106467, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002442.

Haimovici:2021:LTC

[HCCC21] Manuel Haimovici, Leticia Maria Cavole, Jason M. Cope, and Luís Gustavo Cardoso. Long-term changes in population dynamics and life history contribute to explain the resilience of a stock of Micropogonias furnieri (Sciaenidae, Teleostei) in the SW Atlantic. Fisheries Research, 237(??):Article 105878, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000060.

Hsu:2022:EIS

[HCDB22] Jhen Hsu, Yi-Jay Chang, and Nicholas D. Ducharme-Barth. Evaluation of the influence of spatial treatments on catch-per-unit-effort standardization: a fishery application and simulation study of Pacific saury in the Northwestern Pacific Ocean. Fisheries Research, 255(??):Article 106440, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362200217X.

Hoyle:2024:CPU

[HCDB^{+}24] Simon D. Hoyle, Robert A. Campbell, Nicholas D. Ducharme-Barth, Arnaud Grüss, Bradley R. Moore, James T. Thorson, Laura Tremblay-Boyer, Henning Winker, Shijie Zhou, and Mark N. Maunder. Catch per unit effort modelling for stock assessment: a summary of good practices. Fisheries Research, 269(??):??, January 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002539.

Hyman:2024:MEM

[HCDF24] A. Challen Hyman, David Chagaris, Michael Drexler, and Thomas K. Frazer. Modeling effort in a multispecies recreational fishery; influence of species-specific temporal closures, relative abundance, and seasonality on monthly anglertrips. Fisheries Research, 279(??):??, November 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002005.

Hsu:2021:ESD

[HCK^{+}21] Jhen Hsu, Yi-Jay Chang, Toshihide Kitakado, Mikihiko Kai, Bai Li, Midori Hashimoto, Chih hao Hsieh, Vladimir Kulik, and Kyum Joon Park. Evaluating the spatiotemporal dynamics of Pacific saury in the Northwestern Pacific Ocean by using a geostatistical modelling approach. Fisheries Research, 235(??):Article 105821, March 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303386.

Howard:2023:CFI

[HCWH23] Rebecca A. Howard, Lorenzo Ciannelli, W. Waldo Wakefield, and Melissa A. Haltuch. Comparing fishery-independent and fishery-dependent data for analysis of the distributions of Oregon shelf groundfishes. Fisheries Research, 258(??):Article 106553, February 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003307.

How:2022:DIF

[HdLHD22] Jason How, Simon de Lestang, Benjamin Hebiton, and Ainslie Denham. Development of an industry-funded fisheryindependent survey and associated indices for managing a deep sea crab resource in Western Australia. Fisheries Research, 251(??):Article 106296, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200073X.

Hiddink:2023:EHF

[HEG^{+}23a] Jan Geert Hiddink, Lowri Evans, Fiona Gilmour, Gonçalo Lourenço, Shaun McLennan, Elise Quinn, and Jennifer Shepperson. The effect of habitat and fishing-effort data resolution on the outcome of seabed status assessment in bottom trawl fisheries. Fisheries Research, 259(??):Article 106578, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003551.

Horton:2023:BPC

[HEG^{+}23b] Thomas W. Horton, Owen Exeter, Francesco Garzon, Cat Gordon, Lucy A. Hawkes, Ali Hood, David Righton, Joana F. Silva, and Matthew J. Witt. Best practices for catch-and-release shark angling: current scientific understanding and future research. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001534.

Henriques:2024:LMI

[HEGR24] Nuno Sales Henriques, Karim Erzini, Jorge M. S. Gonçalves, and Tommaso Russo. Let’s measure it: an approach of highresolution estimates of bottom fixed net fishing effort at national level. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001826.

Hart:2020:ATA

[HF20] Amanda R. Hart and Gavin Fay. Applying tree analysis to assess combinations of ecosystem-based fisheries management actions in management strategy evaluation. Fisheries Research, 225(??):Article 105466, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303212.

Hommik:2020:DSS

[HFKS20] Kristiina Hommik, Colm J. Fitzgerald, Fiona Kelly, and Samuel Shephard. Dome-shaped selectivity in LB-SPR: Length-based assessment of data-limited inland fish stocks sampled with gillnets. Fisheries Research, 229 (??):Article 105574, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300916.

Haas:2020:RFO

[HFMH20] Bianca Haas, Aysha Fleming, Jeffrey McGee, and Marcus Haward. Regional fisheries organizations and sustainable development goals 13 and 14: Insights from stakeholders.

Fisheries Research, 226(??):Article 105529, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300461.

Hernvann:2020:EIF

[HG20] Pierre-Yves Hernvann and Didier Gascuel. Exploring the impacts of fishing and environment on the Celtic Sea ecosystem since 1950. Fisheries Research, 225(??):Article 105472, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619303273.

Hlina:2021:HRH

[HGC^{+}21] Benjamin L. Hlina, Daniel M. Glassman, Auston D. Chhor, Brooke S. Etherington, Chris K. Elvidge, Benjamin K. Diggles, and Steven J. Cooke. Hook retention but not hooking injury is associated with behavioral differences in bluegill. Fisheries Research, 242(??):Article 106034, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001624.

He:2024:DEB

[HGC24] Weiwei He, Wenjiang Guan, and Ruixing Cao. Dynamic energy budget model for the complete life cycle of chub mackerel in the Northwest Pacific. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002953.

Heney:2025:CPS

[HGHH25] Charlotte Heney, Julie M. Gross, John M. Hoenig, and Natalie Hold. Commercial pot selectivity for European lobster, Homarus gammarus, estimated through a novel v-notching mark–recapture approach. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000207.

Huhn:2023:DSD

[HGS^{+}23] Daniel Hühn, Daniel C. Gwinn, Stephanie L. Shaw, Josep Alós, Micheal S. Allen, Thilo Pagel, Christian Skov, and Robert Arlinghaus. Density- and size-dependent mechanisms modulate the outcome of stocking in a naturally recruiting freshwater piscivore (northern pike, Esox lucius): a replicated whole-lake experiment. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001923.

Horbowy:2020:SBE

[HH20] Jan Horbowy and Kristiina Hommik. Survey-based estimates of F_{msy} and its proxies. Fisheries Research, 229(??):Article 105607, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301247.

Hoshino:2020:DPE

[HHD^{+}20] Eriko Hoshino, Richard Hillary, Campbell Davies, Fayakun Satria, Lilis Sadiyah, Tri Ernawati, and Craig Proctor. Development of pilot empirical harvest strategies for tropical tuna in Indonesian archipelagic waters: Case studies of skipjack and yellowfin tuna. Fisheries Research, 227(??):Article 105539, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300564.

Hanamseth:2022:AIA

[HHJ^{+}22] Roshan Hanamseth, Daniel E. Hewitt, Daniel D. Johnson, Iain M. Suthers, and Matthew D. Taylor. An automated image analysis system for estimating fecundity in portunid crabs. Fisheries Research, 245(??):Article 106140, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362100268X.

Hoenig:2025:PEN

[HHLL25] John M. Hoenig, William S. Hearn, George M. Leigh, and Robert J. Latour. Principles for estimating natural mortality rate. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002595.

Hirao:2024:GWS

[HIF^{+}24] Akira S. Hirao, Junichi Imoto, Atushi Fujiwara, Chikako Watanabe, Mari Yoda, Aiko Matsuura, and Tetsuya Akita. Genome-wide SNP analysis coupled with geographic and reproductive-phenological information reveals panmixia in a classical marine species, the Japanese jack mackerel (Trachurus japonicus). Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002108.

Hayashi:2021:OSA

[HIKM21] Akira Hayashi, Momoko Ichinokawa, Junji Kinoshita, and Akihiro Manabe. Optimizing stock assessment workflows by applying software development methodology. Fisheries Research, 244(??):Article 106108, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002368.

Hamel:2023:NMT

[HIMP23] Owen S. Hamel, James N. Ianelli, Mark N. Maunder, and André E. Punt. Natural mortality: Theory, estimation and application in fishery stock assessment models. Fisheries Research, 261(??):Article 106638, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000310.

Howarth:2021:CRR

[HJA^{+}21] A. Howarth, A. L. Jeanson, A. E. I. Abrams, C. Beaudoin, I. Mistry, A. Berberi, N. Young, V. M. Nguyen, S. J. Landsman, A. N. Kadykalo, A. J. Danylchuk, and S. J. Cooke. COVID-19 restrictions and recreational fisheries in Ontario, Canada: Preliminary insights from an online angler survey. Fisheries Research, 240(??):Article 105961, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000898.

Hestetune:2020:CCA

[HJMS20] Adam Hestetune, Paul M. Jakus, Christopher Monz, and Jordan W. Smith. Climate change and angling behavior on the North shore of Lake Superior (USA). Fisheries Research, 231(??):Article 105717, November 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302344.

Hyun:2022:EEP

[HK22] Saang-Yoon Hyun and Kyuhan Kim. An evaluation of estimability of parameters in the state-space nonlinear logistic production model. Fisheries Research, 245 (??):Article 106135, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621002630.

Huntsberger:2024:IDA

[HKCW24] Carl. J. Huntsberger, Raouf Kilada, Yong Chen, and Richard A. Wahle. Integrating different aging methods to model the dynamics of hard-to-age crab growth: Age at size estimates for the Jonah crab (Cancer borealis). Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001255.

Haase:2021:VAG

[HKG^{+}21] Stefanie Haase, Uwe Krumme, Ulf Gräwe, Camrin D. Braun, and Axel Temming. Validation approaches of a geolocation framework to reconstruct movements of demersal fish equipped with data storage tags in a stratified environment. Fisheries Research, 237(??):Article 105884, May 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000126.

Holopainen:2025:SIA

[HKKa^{+}25] Riikka Holopainen, Tuija Kantala, Tiina Korkea-aho, Petra Heikkinen, Antti Oksanen, Panu Orell, and Satu Viljamaa-Dirks. Survey of the invasive alien pink salmon (Oncorhynchus gorbuscha) for infective agents in the Fennoscandian rivers Tana and Neiden. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002510.

Hoyle:2020:SFM

[HL20] Simon D. Hoyle and Adam D. Langley. Scaling factors for multi-region stock assessments, with an application to Indian Ocean tropical tunas. Fisheries Research, 228(??):Article 105586, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030103X.

Helminen:2021:OBD

[HL21] Jani Helminen and Tommi Linnansaari. Object and behavior differentiation for improved automated counts of migrating river fish using imaging sonar data. Fisheries Research, 237(??):Article 105883, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000114.

He:2025:ULR

[HLC^{+}25] Zhongtang He, Mingju Luo, Weijie Cui, Yang Zhou, and Tao He. Use left or right otolith for shape analysis in chizothorax grahami? Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000323.

Han:2022:WDC

[HLCC22] Kiuk Han, Kyounghee Leem, Young Rok Choi, and Keunsuk Chung. What drives a country’s fish consumption? Market growth phase and the causal relations among fish consumption, production and income growth. Fisheries Research, 254(??):Article 106435, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002120.

Hall:2021:SCC

[HLG^{+}21] Jennifer R. Hall, Sarah J. Lehnert, Emmanuel Gonzalez, Surendra Kumar, Jacqueline M. Hanlon, Corey J. Morris, and Matthew L. Rise. Snow crab (Chionoecetes opilio) hepatopancreas transcriptome: Identification and testing of candidate molecular biomarkers of seismic survey impact. Fisheries Research, 234(??):Article 105794, February 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303118.

Huret:2020:GPS

[HLI^{+}20] Martin Huret, Christophe Lebigre, Mikel Iriondo, Iratxe Montes, and Andone Estonba. Genetic population structure of anchovy (Engraulis encrasicolus) in North-western Europe and variability in the seasonal distribution of the stocks. Fisheries Research, 229(??):Article 105619, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301363.

Hopkins:2024:TOB

[HLMV24] Stephanie C. Hopkins, Sigrid Lehuta, Stephanie Mahevas, and Sandrine Vaz. Trade-offs between spatial temporal closures and effort reduction measures to ensure fisheries sustainability. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000626.

Hua:2020:HSP

[HLZ^{+}20] Chuanxiang Hua, Fei Li, Qingcheng Zhu, Guoping Zhu, and Lingwen Meng. Habitat suitability of Pacific saury (Cololabis saira) based on a yield-density model and weighted analysis. Fisheries Research, 221(??):Article 105408, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302632.

Hansell:2025:IFI

[HM25] Alexander C. Hansell and M. Conor McManus. Integrating fisheries independent surveys to account for the spatiotemporal dynamics of spiny dogfish (Squalus acanthias) in US waters of the northwest Atlantic. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002376.

Hansen:2023:GNC

[HMC^{+}23] Adam G. Hansen, Michael W. Miller, Erik T. Cristan, Collin J. Farrell, Paul Winkle, Mandi M. Brandt, Kyle D. Battige, and Jesse M. Lepak. Gill net catchability of walleye (Sander vitreus): Are provincial standards suitable for estimating adult density outside the region? Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001935.

He:2020:MGI

[HML^{+}20] Song He, Jarle Mork, William B. Larsen, Peter R. Møller, and Michael L. Berumen. Morphology and genetic investigation of flatfish interspecies hybrids (Pleuronectes platessa × Platichthys flesus) from the Baltic Sea. Fisheries Research, 225(??):Article 105498, May 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300151.

Haase:2021:STT

[HMM^{+}21] Stefanie Haase, Kate McQueen, Monica Mion, Magnus Andersson, Annelie Hilvarsson, Hans Jakob Olesen, Anders Svenson, Michele Casini, Karin Hüssy, Krzysztof Radtke, and Uwe Krumme. Short-term tagging mortality of Baltic cod (Gadus morhua). Fisheries Research, 234(??):Article 105804, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303210.

Hoyle:2022:PDN

[HMP^{+}22] Simon D. Hoyle, Mark N. Maunder, André E. Punt, Pamela M. Mace, Jennifer A. Devine, and Z. Teresa A’mar. Preface: Developing the next generation of stock assessment software. Fisheries Research, 246(??):Article 106176, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0165783621003040.

Hutchinson:2024:RBP

[HMR24a] Emily Hutchinson, Thomas R. Matthews, and Gabrielle F. Renchen. Relationships between postlarval settlement and commercial landings of Caribbean spiny lobster (Panulirus argus) in Florida (USA). Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002017.

Hutchinson:2024:GMO

[HMR^{+}24b] Emily Hutchinson, Thomas R. Matthews, Erica Ross, Samantha Hagedorn, and Mark J. Butler. Gastric mill ossicles record chronological age in the Caribbean spiny lobster (Panulirus argus). Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001474.

Hughes:2022:MPI

[HMS^{+}22] Julian M. Hughes, Nicholas M. Meadows, John Stewart, David J. Booth, and Ashley M. Fowler. Movement patterns of an iconic recreational fish species, mulloway (Argyrosomus japonicus), revealed by cooperative citizen-science tagging programs in coastal eastern Australia. Fisheries Research, 247(??):Article 106179, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003076.

Hamad:2025:IFM

[HMY25] Amini I. Hamad, Christopher A. Muhando, and Batuli M. Yahya. The influence of fishing methods on catch size, catch rate, and size distribution of Octopus cyanea (Gray, 1849) in Zanzibar, Tanzania. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002777.

Helgoe:2020:CAM

[HOQ20] Joshua Helgoe, Kenneth J. Oswald, and Joseph M. Quattro. A comprehensive analysis of the mislabeling of Atlantic cod (Gadus morhua) products in Spain. Fisheries Research, 222(??):Article 105400, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302553.

Hoffle:2023:NME

[HP23] Hannes Höffle and Benjamin Planque. Natural mortality estimations for beaked redfish (Sebastes mentella) — a longlived ovoviviparous species of the Northeast Arctic. Fisheries Research, 260(??):Article 106581, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003587.

Hutton:2022:ERS

[HPD^{+}22] Trevor Hutton, Sean Pascoe, Roy A. Deng, André E. Punt, and Shijie Zhou. Effects of re-specifying the Northern Prawn Fishery bioeconomic model to include banana prawns. Fisheries Research, 247(??):Article 106190, March 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003180.

Han:2024:CFT

[HPL^{+}24] Chenglong Han, Xiaoke Pang, Wanni Liu, Wenjing Fu, Biao Guo, Toshihisa Kinoshita, Debin Zheng, Osamu Yamashita, Jianfeng Feng, and Xueqiang Lu. Comparison of floating and traditional artificial reef areas using an ecosystembased approach in the dashentang area of Bohai Bay, China. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001498.

Hance:2024:BBE

[HPPT24] Dalton J. Hance, John M. Plumb, Russell W. Perry, and Kenneth F. Tiffan. Back from the brink: Estimating daily and annual abundance of natural-origin salmon smolts from 30-years of mixed-origin capture-recapture data. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001620.

Hoshino:2024:LRP

[HPS^{+}24] Eriko Hoshino, Sean Pascoe, Peggy Schrobback, Stephanie McWhinnie, and Robert Curtotti. Long-run productivity changes in the Australian northern prawn fishery. Fisheries Research, 273(??):??, May 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362400033X.

Heckel:2020:LHS

[HQWD20] John W. Heckel, Michael C. Quist, Carson J. Watkins, and Andrew M. Dux. Life history structure of westslope cutthroat trout: Inferences from otolith microchemistry. Fisheries Research, 222(??):Article 105416, February 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302711.

Hernandez-Rodriguez:2023:SPB

[HRC23] Ruth E. Hernández-Rodríguez and Luis A. Cubillos. Spatiotemporal physical barrier analysis of southern king crab (Lithodes santolla) catch rates in Magallanes, Chilean Patagonia (2014–2020). Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002138.

Hill:2022:QAI

[HRH22] Christopher Ryan Hill, Roger A. Rulifson, and Norman M. Halden. A quantitative approach to investigate natal philopatry in hickory shad (Alosa mediocris) using otolith chemistry. Fisheries Research, 249(??):Article 106232, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000091.

Haque:2021:TTE

[HS21] Alifa Bintha Haque and Julia L. Y. Spaet. Trade in threatened elasmobranchs in the Bay of Bengal, Bangladesh. Fisheries Research, 243(??):Article 106059, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001879.

Han:2024:UEC

[HSJ^{+}24] Qingpeng Han, Xiujuan Shan, Xianshi Jin, Harry Gorfine, Yue Jin, Qiang Wu, and Yongqiang Shi. Understanding the effects of climate and anthropogenic stresses on distribution variability of Setipinna taty in the Yellow Sea. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001012.

Hu:2022:CFC

[HSL^{+}22] Wenjing Hu, Chaoqun Su, Qigen Liu, Youjia Kong, Shaopeng Hua, and Zhongjun Hu. Comparison of fish communities using environmental DNA metabarcoding and capture methods in a freshwater lake: a new set of universal PCR primers. Fisheries Research, 253(??):Article 106365, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001424.

Heldt:2021:DFE

[HSM21] Katherine Heldt, Ben Stobart, and Stephen Mayfield. Diving into fisher experience: Do new entrants and fleet turnover depress catch rates in abalone (Haliotis laevigata and H. rubra) fisheries. Fisheries Research, 238(??):Article 105906, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000345.

Huntsberger:2024:IRB

[HSM^{+}24] C. J. Huntsberger, B. Shank, M. C. McManus, A. Ellertson, and N. D. Bethoney. Industry reported biological data informs population demographics and commercial fleet heterogeneity for American lobster (Homarus americanus). Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400016X.

Hajisamae:2025:PSF

[HSM^{+}25] Sukree Hajisamae, Kay Khine Soe, Sofiyudin Maae, Saweit Chaymongkol, and Akbar John. Population structure and feeding habits of Carcinoscorpius rotundicauda in the bay environment, Southern Gulf of Thailand. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002807.

Hoshino:2021:MIB

[HSPC21] Eriko Hoshino, Peggy Schrobback, Sean Pascoe, and Robert Curtotti. Market integration between the major domestic fish markets in Australia. Fisheries Research, 243(??):Article 106085, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002137.

Heller-Shipley:2021:SHC

[HSSD^{+}21] Madison A. Heller-Shipley, William T. Stockhausen, Benjamin J. Daly, André E. Punt, and Scott E. Goodman. Should harvest control rules for male-only fisheries include reproductive buffers? A Bering Sea Tanner crab (Chionoecetes bairdi) case study. Fisheries Research, 243(??):Article 106049, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001776.

Hwang:2025:EBS

[HSW25] Julian J. Hwang, Jacquelyn Strager, and Matthew Walker. The economic benefit of spearfishing as an impure public good: a case study of invasive Lionfish in Florida. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002583.

Horimoto:2024:DTD

[HTK^{+}24] Takanori Horimoto, Toshiki Tanii, Tomohiro Kuwae, Kenta Watanabe, and Motohiro Ito. Diet–tissue discrimination factors and turnover rates of carbon and nitrogen stable isotopes in the mottled skate Beringraja pulchra based on dietswitching experiments. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000705.

Hewitt:2023:EDV

[HTSJ23] Daniel E. Hewitt, Matthew D. Taylor, Iain M. Suthers, and Daniel D. Johnson. Environmental drivers of variation in southeast Australian Giant Mud Crab (Scylla serrata) harvest rates. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002436.

Hellenbrecht:2023:DAA

[HUK^{+}23] Lea Marie Hellenbrecht, Kjell Rong Utne, Ørjan Karlsen, Kevin Alan Glover, and Vidar Wennevik. Diet analysis of Atlantic salmon (Salmo salar) post-smolts after the ecological regime shift in the Northeast Atlantic. Fisheries Research, 262(??):Article 106672, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000656.

Hutniczak:2022:ACR

[Hut22] Barbara Hutniczak. Assessing cross-regional flows of economic benefits: a case study of Pacific halibut commercial fishing in Alaska. Fisheries Research, 255(??):Article 106449, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622002260.

Huda:2025:FTT

[HVMEO25] Nurul Huda, Tiago Veiga-Malta, Ole Eigaard, and Finbarr G. O’Neill. Flume tank trials to investigate the snagging of towed demersal fishing gears on boulders. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003011.

Hilborn:2021:SRR

[HW21] Ray Hilborn and Carl J. Walters. Steep recruitment relationships result from modest changes in egg to recruit mortality rates. Fisheries Research, 237(??):Article 105872, May 2021.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303891.

Hulson:2024:IAE

[HW24] Peter-John F. Hulson and Benjamin C. Williams. Inclusion of ageing error and growth variability using a bootstrap estimation of age composition and conditional age-at-length input sample size for fisheries stock assessment models. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002874.

Hoyle:2023:AEN

[HWMVM23] Simon D. Hoyle, Ashley J. Williams, Carolina V. Minte-Vera, and Mark N. Maunder. Approaches for estimating natural mortality in tuna stock assessments: Application to global yellowfin tuna stocks. Fisheries Research, 257(??):Article 106498, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002752.

Hasegawa:2022:SCA

[HWW22] Eva H. Hasegawa, John Waldman, and Isaac Wirgin. Stock composition of Atlantic coastal migratory striped bass using microsatellite DNA analysis. Fisheries Research, 254(??):Article 106384, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001618.

Hu:2020:RLC

[HZZ^{+}20] Jun Hu, Chengquan Zhou, Dandan Zhao, Linnan Zhang, Guijun Yang, and Wenxuan Chen. A rapid, low-cost deep learning system to classify squid species and evaluate freshness based on digital images. Fisheries Research, 221(??):Article 105376, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302310.

Iriarte:2020:IED

[IAB20] Verónica Iriarte, Alexander Arkhipkin, and Denise Blake. Implementation of exclusion devices to mitigate seal (Arctocephalus australis, Otaria flavescens) incidental mortalities during bottom-trawling in the Falkland Islands (Southwest Atlantic). Fisheries Research, 227(??):Article 105537, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300540.

Ingolfsson:2020:RSS

[IB20] Ólafur Arnar Ingólfsson and Jesse Brinkhof. Relative size selectivity of a four-panel codend with short lastridge ropes compared to a flexigrid with a regular codend in the Barents Sea gadoid trawl fishery. Fisheries Research, 232(??):Article 105724, December 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302411.

Ingolfsson:2022:ICE

[IJS^{+}22] Ólafur Arnar Ingólfsson, Terje Jørgensen, Manu Sistiaga, Bent Herrmann, and Liz Kvalvik. Increasing catch efficiency for Nephrops in deep-water shrimp (Pandalus borealis) trawl fisheries. Fisheries Research, 254(??):Article 106394, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001710.

Imzilen:2023:SDF

[IKBL23] Taha Imzilen, David M. Kaplan, Nicolas Barrier, and Christophe Lett. Simulations of drifting fish aggregating device (dFAD) trajectories in the Atlantic and Indian Oceans. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001042.

Kongsstovu:2022:AHC

[íKMíH^{+}22] Sunnvør í Kongsstovu, Svein-Ole Mikalsen, Eydna í Homrum, Jan Arge Jacobsen, Thomas D. Als, Hannes Gislason, Paul Flicek, Einar Eg Nielsen, and Hans Atli Dahl. Atlantic herring (Clupea harengus) population structure in the Northeast Atlantic Ocean. Fisheries Research, 249(??):Article 106231, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362200008X.

Irabor:2024:SSF

[IOO^{+}24] Arnold Ebuka Irabor, Oghenebrorhie Obakanurhe, Augustine Onyemaechi Ozor, Hardin Aaron Jn pierre, Oghenefejiro Adagha, Jovita Oghenenyerhovwo Sanubi, Augustine Ikechukwu Chukwurah, Ifie Idolo, Oster Francis Nwachi, Jerimoth Kesena Ekelemu, Florence Nkeonyeasua Olele, and Samuel Azubuike Zelibe. Is small-scale fishing sustainable in Delta State, Nigeria? A glance into the problems and possible solutions. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000456.

Ivanova:2022:ELS

[IRJ^{+}22] Silviya V. Ivanova, Graham Raby, Timothy B. Johnson, Sarah M. Larocque, and Aaron T. Fisk. Effects of life stage on the spatial ecology of Chinook salmon (Oncorhynchus tshawytscha) during pelagic freshwater foraging. Fisheries Research, 254(??):Article 106395, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001722.

Isguzar:2024:FMS

[İTAD24] Seda İşgüzar, Muammer Türkoğlu, Tuncay Ateşşahin, and Ömerhan Dürrani. FishAgePredictioNet: a multi-stage fish age prediction framework based on segmentation, deep convolution network, and Gaussian process regression with otolith images. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003090.

Janhunen:2023:SDS

[JAN23] Matti Janhunen, Nico Alioravainen, and Teuvo Niva. Straindependent spawning migration of stocked brown trout (Salmo trutta) to a northern Finnish river, in relation to rearing site and release area. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001121.

Juza:2022:FSM

[JBČ^{+}22] Tomáš Jůza, Petr Blabolil, Martin Čech, Vladislav Draštík, Jaroslava Frouzová, Zuzana Sajdlová, Michaela Holubová, Luboš Kočvara, Tomáš Kolařík, Karlos R. Moraes, Milan Muška, Allan T. Souza, Mojmír Vašek, Milan Říha, Michal Tušer, Marek Šmejkal, Jiří Peterka, Marie Prchalová, and Jan Kubečka. Fish stock mass reduction is indicated in standard abundance and biomass estimates from gillnets and hydroacoustics. Fisheries Research, 253(??):Article 106389, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001667.

Jensen:2022:ETC

[JBL^{+}22] Tonje K. Jensen, Jesse Brinkhof, Stein-Kato Lindberg, Torbjørn Tobiassen, Karsten Heia, Stein Harris Olsen, Roger B. Larsen, and Margrethe Esaiassen. Effect of the T90-codend on the catch quality of cod (Gadus morhua) compared to the conventional codend configuration in the Barents Sea bottom trawl fishery. Fisheries Research, 250(??):Article 106277, June 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000546.

Johnson:2021:HSP

[JC21] Samuel D. N. Johnson and Sean P. Cox. Hierarchical surplus production stock assessment models improve management performance in multi-species, spatially-replicated fisheries. Fisheries Research, 238(??):Article 105885, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000138.

Jakes-Cota:2021:EAG

[JCCAS^{+}21] Ulianov Jakes-Cota, Rafael Chavéz-Arellano, Chugey Sepulveda, Scott Aalbers, and Sofía Ortega-García. Estimating age and growth of roosterfish (Nematistius pectoralis) from otoliths. Fisheries Research, 240(??):Article 105958, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000862.

Joshi:2021:EEF

[JCL^{+}21] Omkar Joshi, Binod P. Chapagain, James M. Long, Betsey York, and Andrew T. Taylor. Estimating the effects of fish quality and size on the economic value of fishing in Oklahoma streams and rivers: a revealed preference and contingent behavior approach. Fisheries Research, 244(??):Article 106116, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002447.

Jara:2022:ISL

[JDH22] Andrea Jara, Matthew D. Damiano, and Selina S. Heppell. Integration of scientific and local expertise to develop risk assessments for nearshore species at different spatial scales. Fisheries Research, 245(??):Article 106153, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002812.

Jensen:2022:PMM

[JDP22] Alexander J. Jensen, Steven J. Dundas, and James T. Peterson. Phenomenological and mechanistic modeling of recreational angling behavior using creel data. Fisheries Research, 249(??):Article 106235, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000121.

Jourdain:2021:MBS

[JFS21] N. O. A. S. Jourdain, E. Fuglebakk, and S. Subbey. Maturation in the Barents Sea capelin — contrasting length- and gonad-based metrics. Fisheries Research, 237(??):Article 105880, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621000084.

Jury:2024:BTO

[JGG^{+}24] Steven H. Jury, Benjamin C. Gutzler, Jason S. Goldstein, Joshua T. Carloni, and Winsor H. Watson. Behavioral thermoregulation of ovigerous American lobsters (Homarus americanus). Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001322.

Jones:2021:ARI

[JGU21] Robyn E. Jones, Ross A. Griffin, and Richard K. F. Unsworth. Adaptive Resolution Imaging Sonar (ARIS) as a tool for marine fish identification. Fisheries Research, 243(??):Article 106092, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002204.

Jansen:2021:LDD

[JHB21] Teunis Jansen, Flemming Thorbjørn Hansen, and Birkir Bardarson. Larval drift dynamics, thermal conditions and the shift in juvenile capelin distribution and recruitment success around Iceland and East Greenland. Fisheries Research, 236(??):Article 105845, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303623.

Jung:2022:MFF

[JHM22] Jung-Mo Jung, Miyuki Hirose, and Yoshiki Matsushita. Measuring the footprint of fly-dragging gear. Fisheries Research, 255(??):Article 106465, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002429.

Jusufovski:2020:EIP

[JK20] Dunja Jusufovski and Anna Kuparinen. Exploring individual and population eco-evolutionary feedbacks under the coupled effects of fishing and predation. Fisheries Research, 231(??):Article 105713, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302307.

Jackson:2024:UAI

[JLYR24] Derek N. Jackson, Mark J. M. Lomeli, Noëlle Yochum, and David B. Rudders. The use of artificial illumination to reduce Pacific halibut (Hippoglossus stenolepis) bycatch for a high-rise bottom trawl in the U.S. West Coast groundfish fishery. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001784.

Johnson:2023:ODQ

[JM23] Garrett R. Johnson and Benjamin J. Marcek. Optimizing data quantity and quality for side-looking echosounder surveys in large rivers. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001066.

Jones:2021:EAE

[JMP^{+}21] Andrew W. Jones, Timothy J. Miller, Philip J. Politis, David E. Richardson, Anna M. Mercer, Michael V. Pol, and Christopher D. Roebuck. Experimental assessment of the effect of net wing spread on relative catch efficiency of four flatfishes by a four seam bottom trawl. Fisheries Research, 244(??):Article 106106, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002344.

Jin:2025:EFL

[JMS25] Yiheng Jin, Lingcheng Meng, and Tao Shi. An effective feature learning approach using genetic programming for crab age classification. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002613.

Jubb:2023:ATI

[JND^{+}23] W. M. Jubb, R. A. A. Noble, J. R. Dodd, A. D. Nunn, A. J. Lothian, A. J. Albright, D. H. Bubb, M. C. Lucas, and J. D. Bolland. Acoustic telemetry informs conditional capture probability of an anadromous fish. Fisheries Research, 264(??):??, August 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001303.

Jones:2021:EAC

[JRW^{+}21] Darin T. Jones, Christopher N. Rooper, Christopher D. Wilson, Paul D. Spencer, Dana H. Hanselman, and Rachel E. Wilborn. Estimates of availability and catchability for select rockfish species based on acoustic-optic surveys in the Gulf of Alaska. Fisheries Research, 236(??):Article 105848, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303659.

Jokar:2021:LFT

[JSG21] Mahmood Jokar, Sam Subbey, and Harald Gjøsæter. A logistic function to track time-dependent fish population dynamics. Fisheries Research, 236(??):Article 105840, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362030357X.

Joubert:2020:CSK

[JSKM20] B. A. Joubert, M. G. Sullivan, B. C. Kissinger, and A. T. Meinke. Can smartphones kill trout? Mortality of memorable-sized bull trout (Salvelinus confluentus) after photo-releases. Fisheries Research, 223(??):Article 105458, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619303133.

Jayanti:2025:DSB

[JvPOG25] Anes Dwi Jayanti, Ingrid van Putten, Emily Ogier, and Caleb Gardner. Diverse social and business networks shape the puerulus harvest industry along southern Indonesia. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003278.

Jiang:2024:SCE

[JWL^{+}24] Mingfeng Jiang, Jintao Wang, Gang Li, Bilin Liu, and Xinjun Chen. Is seasonal closure an effective way to conserve oceanic squids — taking Chinese autonomic seasonal closure on the high seas as an example. Fisheries Research, 271(??): ??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623003077.

Johnston:2021:EDC

[JYH21] Danielle J. Johnston, Daniel E. Yeoh, and David C. Harris. Environmental drivers of commercial blue swimmer crab (Portunus armatus) catch rates in Western Australian fisheries. Fisheries Research, 235(??):Article 105827, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303441.

Jiang:2020:EDA

[JZQZ20] Dianhang Jiang, Xiaodong Zheng, Yaosen Qian, and Qingqi Zhang. Embryonic development of Amphioctopus fangsiao under elevated temperatures: Implications for resource management and conservation. Fisheries Research, 225(??):Article 105479, May 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303340.

Kittlein:2024:ERS

[KA24] Marcelo J. Kittlein and Juan Alberti. Effects of resource spatial distribution, tow overlap, and positional error in the estimation of dredge efficiency. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002029.

Kaemingk:2022:MRU

[KAB^{+}22] Mark A. Kaemingk, Robert Arlinghaus, Max H. Birdsong, Christopher J. Chizinski, Roman Lyach, Kyle L. Wilson, and Kevin L. Pope. Matching of resource use and investment according to waterbody size in recreational fisheries. Fisheries Research, 254(??):Article 106388, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001655.

Kehoe:2023:CCH

[KAC^{+}23] Liam Kehoe, Frank Asche, Claire Crowley, Ryan Gandy, and David Chagaris. Costly crustaceans: a hedonic price analysis of the Florida stone crab. Fisheries Research, 258(??):Article 106541, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003186.

Kratzer:2021:UAV

[KBB^{+}21] Isabella Maria Friederike Kratzer, Mollie Elizabeth Brooks, Sabri Bilgin, Süleyman Özdemir, Lotte Kindt-Larsen, Finn Larsen, and Daniel Stepputtis. Using acoustically visible gillnets to reduce bycatch of a small cetacean: first pilot trials in a commercial fishery. Fisheries Research, 243(??):Article 106088, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002162.

Kienzle:2022:BET

[KBH22] Marco Kienzle, Matt K. Broadhurst, and Gary Hamer. Bayesian estimates of turban snail (Lunella torquata) growth off south-eastern Australia. Fisheries Research, 248(??):Article 106218, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621003465.

Kokkalis:2024:GPS

[KBK^{+}24] A. Kokkalis, C. W. Berg, M. S. Kapur, H. Winker, N. S. Jacobsen, M. H. Taylor, M. Ichinokawa, M. Miyagawa, W. Medeiros-Leal, J. R. Nielsen, and T. K. Mildenberger. Good practices for surplus production models. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000742.

Kubilius:2023:RAS

[KBM23] Rokas Kubilius, Benoît Bergès, and Gavin J. Macaulay. Remote acoustic sizing of tethered fish using broadband acoustics. Fisheries Research, 260(??):Article 106585, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003629.

Kienzle:2021:BAE

[KBPS21] Marco Kienzle, Matt K. Broadhurst, Alexander Pletzer, and John Stewart. A Bayesian approach to estimating mortality rates using hazard functions: Application to an Australian halfbeak, Hyporhamphus australis fishery. Fisheries Research, 243(??):Article 106066, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001946. See erratum [KBPS22].

Kienzle:2022:EBA

[KBPS22] Marco Kienzle, Matt K. Broadhurst, Alexander Pletzer, and John Stewart. Erratum to “A Bayesian approach to estimating mortality rates using hazard functions: Application to an Australian halfbeak, Hyporhamphus australis fishery” [Fish. Res. 243 (2021) 10606]. Fisheries Research, 245(??):Article 106137, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002654. See [KBPS21].

Kapusta:2022:IHT

[KC22] Andrzej Kapusta and Tomasz K. Czarkowski. Influence of hook type on performance, hooking location, injury, and reflex action mortality predictors in float recreational angling for cyprinids: a case study in northeastern Poland. Fisheries Research, 254(??):Article 106390, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001679.

Kirke:2024:RBT

[KCB^{+}24] A. K. Kirke, D. A. Crook, S. C. Banks, O. J. Luiz, T. M. Saunders, A. J. King., and G. J. Johnson. Reproductive biology of two small-bodied sharks, Carcharhinus coatesi and Rhizoprionodon acutus, in the Northern Territory, Australia. Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000109.

Kemp:2022:DSP

[KD22] Chandler E. Kemp and Susan K. Doherty. Determining salmon provenance with automated otolith reading. Fisheries Research, 250(??):Article 106295, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000728.

Kapur:2025:GPT

[KDBOC25] Maia S. Kapur, Nicholas Ducharme-Barth, Megumi Oshima, and Felipe Carvalho. Good practices, trade-offs, and precautions for model diagnostics in integrated stock assessments. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002704.

Kliemann:2022:DCF

[KDdOM^{+}22] Bruna Caroline Kotz Kliemann, Rosilene Luciana Delariva, Letícia de Oliveira Manoel, Amanda Pereira dos Santos Silva, Rosicleire Veríssimo-Silveira, and Igor Paiva Ramos. Do cage fish farms promote interference in the trophic niche of wild fish in neotropical reservoir? Fisheries Research, 248(??):Article 106198, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100326X.

Kruse:2025:EBF

[KDF^{+}25] Gordon H. Kruse, Benjamin J. Daly, Erin J. Fedewa, Diana L. Stram, and Cody S. Szuwalski. Ecosystem-based fisheries management of crab fisheries in the Bering Sea and Aleutian Islands. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400300X.

Kennedy:2021:MLO

[Ken21] James Kennedy. Measuring lumps: Optimising sampling of the commercial landings of lumpfish (Cyclopterus lumpus) in Iceland. Fisheries Research, 244(??):Article 106130, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002587.

Knotek:2022:PRM

[KFDE^{+}22] R. J. Knotek, B. S. Frazier, T. S. Daly-Engel, C. F. White, S. N. Barry, E. J. Cave, and N. M. Whitney. Post-release mortality, recovery, and stress physiology of blacknose sharks, Carcharhinus acronotus, in the Southeast U.S. recreational shark fishery. Fisheries Research, 254(??):Article 106406, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001837.

Krag:2025:USS

[KFH^{+}25] Ludvig A. Krag, Rikke P. Frandsen, Bent Herrmann, Grete E. Dinesen, and Junita D. Karlsen. Using species-specific behavior to improve catch efficiency of target species in mixed trawl fisheries. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002960.

Kuramoto:2021:MCE

[KFI^{+}21] Yosuke Kuramoto, Yasuzumi Fujimori, Ryohei Ito, Yumi Kobayashi, and Yasunori Sakurai. Measures for co-existence between seals and coastal large-scale salmon set net fisheries: Mitigation of catch damage by the use of rope grid. Fisheries Research, 242(??):Article 106041, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001697.

Korostelev:2020:UDH

[KFO20] N. B. Korostelev, P. H. Frey, and A. M. Orlov. Using different hard structures to estimate the age of deep-sea fishes: a case study of the Pacific flatnose, Antimora microlepis (Moridae, Gadiformes, Teleostei). Fisheries Research, 232(??):Article 105731, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302484.

Kapur:2020:OFD

[KHC^{+}20] M. Kapur, M. Haltuch, B. Connors, L. Rogers, A. Berger, E. Koontz, J. Cope, K. Echave, K. Fenske, D. Hanselman, and A. E. Punt. Oceanographic features delineate growth zonation in Northeast Pacific sablefish. Fisheries Research, 222(??):Article 105414, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302693.

Keller:2022:TTS

[KHE^{+}22] Aimee A. Keller, John H. Harms, Anna Elz, John R. Wallace, Jim A. Benante, and Aaron Chappell. A tale of two species: Vermilion and sunset rockfish in the Southern California Bight. Fisheries Research, 250(??):Article 106275, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000522.

Kostak:2025:SSM

[KHGB25] Enis N. Kostak, Bent Herrmann, Eduardo Grimaldo, and Jesse Brinkhof. Size selectivity of Muller’s pearlside (Maurolicus muelleri), glacier lanternfish (Benthosema glaciale) and krill in trawls targeting the mesopelagic fish. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002935.

Koster:2020:EPM

[KHK^{+}20] Friedrich W. Köster, Bastian Huwer, Gerd Kraus, Rabea Diekmann, Margit Eero, Andrei Makarchouk, Serra Örey, Jan Dierking, Piotr Margonski, Jens Peter Herrmann, Jonna Tomkiewicz, Daniel Oesterwind, Paul Kotterba, Holger Haslob, Rüdiger Voss, and Thorsten B. H. Reusch. Egg production methods applied to Eastern Baltic cod provide indices of spawning stock dynamics. Fisheries Research, 227(??):Article 105553, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300709.

Kambestad:2024:VWS

[KHL^{+}24] Marius Kambestad, Bjart Are Hellen, Robert J. Lennox, Gaute Velle, and Harald Sægrov. Validity of winter sampling for estimation of salmonid abundance by electrofishing. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003223.

Khalsa:2023:CDS

[KHMC23] Noah S. Khalsa, Cameron T. Hodgdon, Mackenzie D. Mazur, and Yong Chen. Climate-driven shifts in growth and maturity induce changes to the population and fishery dynamics of a high-value crustacean. Fisheries Research, 259(??):Article 106574, March 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003514.

Krause:2020:ITC

[KHPB20] Jacob R. Krause, Joseph E. Hightower, Stephen J. Poland, and Jeffrey A. Buckel. An integrated tagging and catchcurve model reveals high and seasonally-varying natural mortality for a fish population at low stock biomass. Fisheries Research, 232(??):Article 105725, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302423.

Keppeler:2020:WMG

[KHS^{+}20] Friedrich Wofgang Keppeler, Gustavo Hallwass, Franciele Santos, Luís Henrique Tomazzoni da Silva, and Renato Azevedo Matias Silvano. What makes a good catch? Effects of variables from individual to regional scales on tropical small-scale fisheries. Fisheries Research, 229 (??):Article 105571, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300886.

Kourantidou:2021:ARR

[KK21] Melina Kourantidou and Brooks A. Kaiser. Allocation of research resources for commercially valuable invasions: Norway’s red king crab fishery. Fisheries Research, 237(??):Article 105871, May 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030388X.

Kitada:2022:SUE

[KK22] Shuichi Kitada and Hirohisa Kishino. An R script for unbiased estimation of stocking effectiveness using a twostage sampling strategy based on fishery landings. Fisheries Research, 250(??):Article 106266, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000431.

Klahan:2025:MGE

[KKC25] Rungkan Klahan, Kritsana Krajabthong, and Krit Chaiwong. Maximizing growth, efficiency, and cost-effectiveness: Evaluating feed strategies for flathead lobsters. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003138.

Kane:2020:STB

[KKCP20] Derek S. Kane, Mark A. Kaemingk, Christopher J. Chizinski, and Kevin L. Pope. Spatial and temporal behavioral differences between angler-access types. Fisheries Research, 224(??):Article 105463, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303182.

Kaz:2024:EPR

[KKLM24] Anna L. Kaz, Michael D. Kaller, Abigail J. Lynch, and Stephen R. Midway. Early pandemic recreational fishing patterns across the urban-to-rural gradient in the U.S. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000857.

Kindt-Larsen:2023:AGS

[KLBHK23] Lotte Kindt-Larsen, Casper Willestofte Berg, Maria Hedgärde, and Sara Königson. Avoiding grey seal depredation in the Baltic Sea while increasing catch rates of cod. Fisheries Research, 261(??):Article 106609, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000024.

Kindt-Larsen:2024:PJG

[KLNB^{+}24] Lotte Kindt-Larsen, Thomas Noack, Mollie Elizabeth Brooks, Anne-Mette Kroner, and Gildas Glemarec. Pearls are not just for girls: Plastic spheres do not interfere with target catches in a set net fishery. Fisheries Research, 276(??): ??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000961.

Keiling:2020:BHF

[KLS20] Toniann D. Keiling, Michael J. Louison, and Cory D. Suski. Big, hungry fish get the lure: Size and food availability determine capture over boldness and exploratory behaviors. Fisheries Research, 227(??):Article 105554, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300710.

Kamiyama:2023:ADE

[KM23] Ryutaro Kamiyama and Tsutom Miyata. Applicability of data envelopment analysis using vessel-level data to the estimation of technical efficiency and capacity utilization of the small-scale squid angling fishery in Aomori Prefecture, Japan. Fisheries Research, 262(??):Article 106664, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000577.

Koemle:2022:HCU

[KMA22] Dieter Koemle, Jürgen Meyerhoff, and Robert Arlinghaus. How catch uncertainty and harvest regulations drive anglers’ choice for pike (Esox lucius) fishing in the Baltic Sea. Fisheries Research, 256(??):Article 106480, December 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002570.

Kelly:2023:EMA

[KMA23] Cian Kelly, Finn Are Michelsen, and Morten Omholt Alver. An ensemble modelling approach for spatiotemporally explicit estimation of fish distributions using data assimilation. Fisheries Research, 261(??):Article 106624, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000176.

Kroetz:2020:EPS

[KMC20] Andrea M. Kroetz, Alyssa N. Mathers, and John K. Carlson. Evaluating protected species bycatch in the U.S. Southeast Gillnet Fishery. Fisheries Research, 228(??):Article 105573, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620300904.

Kura:2023:CSL

[KMC^{+}23] Yumiko Kura, Kosal Mam, Seila Chea, Dyna Eam, Kaitlin Almack, and Hiroe Ishihara. Conservation for sustaining livelihoods: Adaptive co-management of fish no-take zones in the Mekong River. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001376.

Kubilius:2020:RSF

[KMO20] Rokas Kubilius, Gavin J. Macaulay, and Egil Ona. Remote sizing of fish-like targets using broadband acoustics. Fisheries Research, 228(??):Article 105568, August 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300850.

Kasana:2025:HES

[KMSJ^{+}25] Devanshi Kasana, Hector Daniel Martinez, Julio Sánchez-Jiménez, Elisa M. Areano-Barillas, Kevin A. Feldheim, and Demian D. Chapman. Hunt for the Easter sharks: a genetic analysis of shark and ray meat markets in Guatemala. Fisheries Research, 283(??):??, March 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000372.

Kanamori:2021:STM

[KNO^{+}21] Yuki Kanamori, Shota Nishijima, Hiroshi Okamura, Ryuji Yukami, Mikio Watai, and Akinori Takasuka. Spatiotemporal model reduces species misidentification bias of spawning eggs in stock assessment of spotted mackerel in the western North Pacific. Fisheries Research, 236(??):Article 105825, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620303428.

Khan:2020:OCF

[KNP^{+}20] Alexander M. A. Khan, Anta M. Nasution, Noir P. Purba, Achmad Rizal, Zahidah, Herman Hamdani, Lantun P. Dewanti, Junianto, Isni Nurruhwati, Asep Sahidin, Dedi Supriyadi, Heti Herawati, Izza M. Apriliani, Monika Ridwan, Tim S. Gray, Mingguo Jiang, Hazmi Arief, Aileen C. Mill, and Nicholas V. C. Polunin. Oceanographic characteristics at fish aggregating device sites for tuna poleand-line fishery in eastern Indonesia. Fisheries Research, 225(??):Article 105471, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619303261.

Kapula:2022:GAS

[KNS^{+}22] VK Kapula, H. O. N. Ndjaula, M. Schulze, D. Durholtz, D. Japp, L. Singh, C. A. Matthee, S. von der Heyden, and R. Henriques. Genetic assessment of seasonal alongshore migration in Merluccius capensis in the Benguela region. Fisheries Research, 250(??):Article 106293, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000704.

Kane:2023:WSP

[KPK^{+}23] Derek S. Kane, Kevin L. Pope, Keith D. Koupal, Mark A. Pegg, Christopher J. Chizinski, and Mark A. Kaemingk. Waterbody size predicts bank- and boat-angler efforts. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001947.

Kongseng:2020:IAM

[KPS20] Sirithorn Kongseng, Ratanavaree Phoonsawat, and Akarapong Swatdipong. Individual assignment and mixed-stock analysis of short mackerel (Rastrelliger brachysoma) in the Inner and Eastern Gulf of Thailand: Contrast migratory behavior among the fishery stocks. Fisheries Research, 221(??):Article 105372, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302279.

Kulik:2022:LCW

[KPUB22] Vladimir V. Kulik, Sergey V. Prants, Michael Yu. Uleysky, and Maxim V. Budyansky. Lagrangian characteristics in the western North Pacific help to explain variability in Pacific saury fishery. Fisheries Research, 252(??):Article 106361, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622001382.

Kongseng:2021:GMS

[KPWS21] Sirithorn Kongseng, Ratanavaree Phoonsawat, Worawit Wanchana, and Akarapong Swatdipong. Genetic mixedstock analysis of short mackerel, Rastrelliger brachysoma, catches in the Gulf of Thailand: Evidence of transboundary migration of the commercially important fish. Fisheries Research, 235(??):Article 105823, March 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303404.

Knott:2024:GEM

[KRH^{+}24] Stephen Knott, Richard Raesly, Daniel Hocking, Jay Stauffer, and Amy Welsh. Genomic evaluation of mud sunfish in the Atlantic slope drainage. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001073.

Kern:2020:FSP

[KSI20a] Aaron I. Kern, Steven M. Sammons, and Travis R. Ingram. Fecundity and spawning potential ratio of shoal bass Micropterus cataractae in the lower Flint River, Georgia, USA. Fisheries Research, 231(??):Article 105687, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302046.

Kurota:2020:DRD

[KSI20b] Hiroyuki Kurota, Cody S. Szuwalski, and Momoko Ichinokawa. Drivers of recruitment dynamics in Japanese major fisheries resources: Effects of environmental conditions and spawner abundance. Fisheries Research, 221 (??):Article 105353, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302085.

Kavitha:2022:IRB

[KSI^{+}22] Mookaiah Kavitha, Geetha Sasikumar, Jagadis Iyadurai, Ranjith Lakshmanan, and Jasmin Felix. Insight on the reproductive biology of small striped cuttlefish, Sepia prabahari in Gulf of Mannar, Indian Ocean and recommendation for a minimum legal size. Fisheries Research, 248(??):Article 106227, April 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000042.

Kristensen:2020:FPU

[KSJM^{+}20] Emil Kristensen, Kaj Sand-Jensen, Kenneth Thorø Martinsen, Mikkel Madsen-Østerbye, and Theis Kragh. Fingerprinting pike: the use of image recognition to identify individual pikes. Fisheries Research, 229(??):Article 105622, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301399.

Kempf:2023:SAM

[KSL^{+}23] Alexander Kempf, Michael A. Spence, Sigrid Lehuta, Vanessa Trijoulet, Valerio Bartolino, Maria Ching Villanueva, and Sarah K. Gaichas. Skill assessment of models relevant for the implementation of ecosystem-based fisheries management. Fisheries Research, 268(??):??, December 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002382.

Kapur:2021:ERP

[KSO^{+}21] M. S. Kapur, M. C. Siple, M. Olmos, K. M. Privitera-Johnson, G. Adams, J. Best, C. Castillo-Jordán, L. Cronin-Fine, A. M. Havron, Q. Lee, R. D. Methot, Jr., and André E. Punt. Equilibrium reference point calculations for the next generation of spatial assessments. Fisheries Research, 244(??):Article 106132, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002605.

Kratzer:2022:ADA

[KSS^{+}22] Isabella Maria Friederike Kratzer, Daniel Stepputtis, Juan Santos, Frauke Lütkefedder, Arne Stoltenberg, Lea Hartkens, Matthias Schaber, Lotte Kindt-Larsen, and Finn Larsen. Angle-dependent acoustic reflectivity of gillnets and their modifications to reduce bycatch of odontocetes using sonar imaging. Fisheries Research, 250(??):Article 106278, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000558.

Karjalainen:2022:GBE

[KSV^{+}22] Juha Karjalainen, Rosanna Sjövik, Tuula Väänänen, Tiina Sävilammi, Lotta-Riina Sundberg, Silva Uusi-Heikkilä, and Timo J. Marjomäki. Genetic-based evaluation of management units for sustainable vendace (Coregonus albula) fisheries in a large lake system. Fisheries Research, 246(??):Article 106173, February 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0165783621003015.

Kamimura:2022:SDD

[KTFY22] Yasuhiro Kamimura, Kazuaki Tadokoro, Sho Furuichi, and Ryuji Yukami. Stronger density-dependent growth of Japanese sardine with lower food availability: Comparison of growth and zooplankton biomass between a historical and current stock-increase period in the western North Pacific. Fisheries Research, 255(??):Article 106461, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002387.

Kerr:2020:MSO

[KWC^{+}20] Lisa A. Kerr, Zachary T. Whitener, Steven X. Cadrin, Molly R. Morse, David H. Secor, and Walter Golet. Mixed stock origin of Atlantic bluefin tuna in the U.S. rod and reel fishery (Gulf of Maine) and implications for fisheries management. Fisheries Research, 224(??):Article 105461, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619303169.

Kendall:2021:MUR

[KWE^{+}21] M. S. Kendall, B. L. Williams, R. D. Ellis, K. E. Flaherty-Walia, A. B. Collins, and K. W. Roberson. Measuring and understanding receiver efficiency in your acoustic telemetry array. Fisheries Research, 234(??):Article 105802, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303192.

Klefoth:2023:DAM

[KWMA23] Thomas Klefoth, Nicola Wegener, Jürgen Meyerhoff, and Robert Arlinghaus. Do anglers and managers think similarly about stocking, habitat management and harvest regulations? Implications for the management of community-governed recreational fisheries. Fisheries Research, 260(??):Article 106589, April 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003666.

Kendall:2021:WWW

[KWW^{+}21] Matthew S. Kendall, Bethany L. Williams, Arliss J. Winship, Mark Carson, Karen Grissom, Timothy J. Rowell, Jenni Stanley, and Kimberly W. Roberson. Winds, waves, warm waters, weekdays, and which ways boats are counted influence predicted visitor use at an offshore fishing destination. Fisheries Research, 237(??):Article 105879, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000072.

Kucinski:2023:IEF

[KZT^{+}23] Marcin Kuciński, Ilona Złoch, Paulina Trzeciak, Anna Kycko, Katarzyna Nadolna-Ałtyn, and Katarzyna Mierzejewska. Infection of the European flounder (Platichthys flesus) with Glugea stephani, a possible new indicator of the weakening of the Baltic population. Fisheries Research, 260(??):Article 106590, April 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003678.

Litzow:2022:PYC

[LADA^{+}22] Michael A. Litzow, Alisa A. Abookire, Janet T. Duffy-Anderson, Benjamin J. Laurel, Michael J. Malick, and Lauren A. Rogers. Predicting year class strength for climate-stressed gadid stocks in the Gulf of Alaska. Fisheries Research, 249(??):Article 106250, May 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000273.

Lucey:2021:EFM

[LAG^{+}21] Sean M. Lucey, Kerim Y. Aydin, Sarah K. Gaichas, Steven X. Cadrin, Gavin Fay, Michael J. Fogarty, and André Punt. Evaluating fishery management strategies using an ecosystem model as an operating model. Fisheries Research, 234(??):Article 105780, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302976.

Liljestrand:2023:ANS

[LBD23] Emily M. Liljestrand, James R. Bence, and Jonathan J. Deroba. Applying a novel state-space stock assessment framework using a fisheries-dependent index of fishing mortality. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001005.

Liljestrand:2024:EPV

[LBD24] Emily M. Liljestrand, James R. Bence, and Jonathan J. Deroba. The effect of process variability and data quality on performance of a state-space stock assessment model. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000870.

Lankowicz:2020:SIS

[LBLF20] Katelynn M. Lankowicz, Hongsheng Bi, Dong Liang, and Chunlei Fan. Sonar imaging surveys fill data gaps in forage fish populations in shallow estuarine tributaries. Fisheries Research, 226(??):Article 105520, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300370.

Leeman:2024:PDP

[LBP^{+}24] Cassandra Leeman, Deborah A. Bouchard, Eric Payne, Richard A. Wahle, and Damian C. Brady. Predicting delayed post-harvest mortality in American lobster (Homarus americanus) using reflex impairment. Fisheries Research, 272(??): ??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000122.

Loher:2021:PUO

[LBW21] Timothy Loher, Gretchen Elizabeth Bath, and Stephen Wischniowski. The potential utility of otolith microchemistry as an indicator of nursery origins in Pacific halibut (Hippoglossus stenolepis). Fisheries Research, 243(??):Article 106072, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002009.

Laptikhovsky:2021:IBE

[LCB^{+}21a] Vladimir Laptikhovsky, Gavan Cooke, Christopher Barrett, Sophie Lozach, Eleanor MacLeod, Daniel Oesterwind, Edel Sheerin, Michael Petroni, Leigh Barnwall, Jean-Paul Robin, Louise Allcock, and Anne Marie Power. Identification of benthic egg masses and spawning grounds in commercial squid in the English Channel and Celtic Sea: Loligo vulgaris vs L. forbesii. Fisheries Research, 241(??):Article 106004, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001326.

LaRochelle:2021:IFH

[LCB^{+}21b] L. LaRochelle, A. D. Chhor, J. W. Brownscombe, A. J. Zolderdo, A. J. Danylchuk, and S. J. Cooke. Ice-fishing handling practices and their effects on the short-term post-release behaviour of largemouth bass. Fisheries Research, 243(??):Article 106084, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002125.

Leitao:2025:PGU

[LCC25] P. Leitão, A. Campos, and M. Castro. Predicting gear used in a multi-gear coastal fleet. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002637.

Laptikhovsky:2023:STV

[LCD^{+}23] Vladimir Laptikhovsky, Gavan Cooke, Christian Drerup, Angus Jackson, Eleanor MacLeod, and Jean-Paul Robin. Spatial and temporal variability of common cuttlefish, Sepia officinalis, L. spawning grounds off North Europe. Fisheries Research, 263(??):??, July 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000814.

Lennert-Cody:2024:WWP

[LCDM^{+}24] Cleridy E. Lennert-Cody, Cristina De La Cadena, Marti McCracken, Luis Chompoy, Nickolas W. Vogel, Mark N. Maunder, Brad A. Wiley, Ernesto Altamirano Nieto, and Alexandre Aires da Silva. Within-well patterns in bigeye tuna catch composition and implications for purse-seine port-sampling and catch estimation for the Eastern Pacific Ocean. Fisheries Research, 277(??):??, September 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001437.

Larocque:2022:ITN

[LCFJ22] Sarah M. Larocque, Scott F. Colborne, Aaron T. Fisk, and Timothy B. Johnson. Improving trophic niche and diet resolution of the salmonid community of Lake Ontario using three stable isotopes and multiple tissues. Fisheries Research, 255(??):Article 106455, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002326.

Lin:2021:SEC

[LCG^{+}21] Mingli Lin, Sibao Chen, Rodolphe E. Gozlan, Peng Ren, Wei Li, Shaowen Ye, Qidong Wang, Tanglin Zhang, Jiashou Liu, Songhai Li, and Zhongjie Li. Stock enhancement of Culter mongolicus: Assessment of growth, recapture and release size in the Yangtze lakes. Fisheries Research, 234(??):Article 105809, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030326X.

Lorenzen:2022:NMB

[LCG22] Kai Lorenzen, Edward V. Camp, and Taryn M. Garlock. Natural mortality and body size in fish populations. Fisheries Research, 252(??):Article 106327, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001047.

Lunzmann-Cooke:2024:RFR

[LCHB^{+}24] Emma L. Lunzmann-Cooke, Scott G. Hinch, Arthur L. Bass, Stephen D. Johnston, Brian J. Hendriks, Aswea D. Porter, Steven J. Cooke, and David W. Welch. Recreational fisheriesrelated injuries and body size affect travel rate and postrelease mortality in marine migrating Coho salmon (Oncorhynchus kisutch). Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001267.

Liu:2025:APF

[LCL25] Shuang Liu, Xiu-Jie Cui, and Chao Lyu. Analysis and prediction of fishing capacity of marine fishing vessels based on generalized additive models. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000050.

Lennert-Cody:2023:APS

[LCLM23] Cleridy E. Lennert-Cody, Jon Lopez, and Mark N. Maunder. An automatic purse-seine set type classification algorithm to inform tropical tuna management. Fisheries Research, 262(??):Article 106644, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000371.

Ledoux:2023:RES

[LCM^{+}23] Tamara Ledoux, Jeff C. Clements, Michelle Maillet, Daniel Gallant, Rémi Sonier, and Gilles Miron. Reproductive ecology of the soft-shell clam (Mya arenaria) in Eastern New Brunswick, Canada: Assessing size-at-maturity and spawning time to inform fisheries management. Fisheries Research, 265(??):??, September 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001522.

Lennert-Cody:2022:SCS

[LCMS^{+}22] Cleridy E. Lennert-Cody, Marti McCracken, Salvador Siu, Ricardo Oliveros-Ramos, Mark N. Maunder, Alexandre Aires da Silva, José Miguel Carvajal-Rodríguez, Jean D. Opsomer, and Pedro de Barros. Single-cluster systematic sampling designs for shark catch size composition in a Central American longline fishery. Fisheries Research, 251(??):Article 106320, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000972.

Lima:2020:DFI

[LCN^{+}20] Maria Alice Leite Lima, Adriana Rosa Carvalho, Marcus Alexandre Nunes, Ronaldo Angelini, and Carolina Rodrigues da Costa Doria. Declining fisheries and increasing prices: the economic cost of tropical rivers impoundment. Fisheries Research, 221(??):Article 105399, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302541.

Liu:2023:EFY

[LCW^{+}23a] Qi Liu, Yunxia Chen, Jing Wang, Haihong Miao, and Yingbin Wang. An example of fishery yield predictions from MSbased navigational characteristics applied to double trawlers in China. Fisheries Research, 261(??):Article 106614, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000073.

Lockhart:2023:CAS

[LCW23b] Katlyn Lockhart, Lorenzo Ciannelli, and W. Waldo Wakefield. A comparative analysis of sampling methodologies for assessing abundance and distribution of youngof-the-year groundfishes in nearshore soft-sediment habitats. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001273.

Lee:2023:GQP

[LD23] Min-Yang Lee and Chad Demarest. Groundfish quota prices. Fisheries Research, 260(??):Article 106605, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003824.

Levine:2025:MDL

[LD25] Mike Levine and Alex De Robertis. Making do with less: Extending an acoustic-based time series of euphausiid abundance using an uncrewed surface vehicle with fewer frequencies. Fisheries Research, 282(??):??, February 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000074.

Long:2024:ORS

[LDC24] William Christopher Long, Benjamin J. Daly, and Peter A. Cummiskey. Optimizing release strategies for red king crab stock enhancement: Effects of release timing. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000390.

Li:2024:ECS

[LDM^{+}24] Chengxue Li, Jonathan J. Deroba, Timothy J. Miller, Christopher M. Legault, and Charles T. Perretti. An evaluation of common stock assessment diagnostic tools for choosing among state-space models with multiple random effects processes. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000328.

Lin:2021:IFL

[LDS^{+}21] Chenyu Lin, Huichao Dai, Xiaotao Shi, Zhiqun Daniel Deng, Jingqiao Mao, Jia Luo, Wenqin Huang, Jiawei Xu, Ning Zhang, and Shuangke Sun. Investigating feasible light configurations for fish restoration: an ethological insight. Fisheries Research, 234(??):Article 105807, February 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303246.

Lopetegui:2022:EPA

[LdV22] Itsaso Lopetegui and Ikerne del Valle. An efficient portfolio approach towards ecosystem-based fisheries governance in EU. Fisheries Research, 254(??):Article 106427, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002041.

Longin:2021:WSF

[LFdB^{+}21] G. Longin, G. Fontenelle, L. Bonneau de Beaufort, C. Delord, S. Launey, R. Rinaldo, G. Lassalle, P.-Y. Le Bail, and J. M. Roussel. When subsistence fishing meets conservation issues: Survey of a small fishery in a neotropical river with high biodiversity value. Fisheries Research, 241(??):Article 105995, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001235.

Langseth:2021:DSS

[LG21] Brian J. Langseth and Sarah Glover. Distinguishing species-specific targeting practices within the main Hawaiian Islands Deep 7 bottomfish fishery. Fisheries Research, 244(??):Article 106105, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002332.

LaMesa:2020:CAO

[LGD^{+}20] Mario La Mesa, Stefano Guicciardi, Fortunata Donato, Emilio Riginella, Luca Schiavon, and Chiara Papetti. Comparative analysis of otolith morphology in icefishes (Channichthyidae) applying different statistical classification methods. Fisheries Research, 230(??):Article 105668, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301855.

Limburg:2025:FEW

[LHH^{+}25] Karin E. Limburg, Yvette Heimbrand, Karin Hüssy, Martina Blass, Jay B. Thomas, Katja Mäkinen, and Tomas Næraa. The forgotten element: Why do we ignore calcium in otolith studies? Fisheries Research, 283(??):??, March 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000347.

Lomnicky:2021:CBR

[LHPR21] Gregg A. Lomnicky, Robert M. Hughes, David V. Peck, and Paul L. Ringold. Correspondence between a recreational fishery index and ecological condition for U.S.A. streams and rivers. Fisheries Research, 233(??):Article 105749, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620302666.

Liu:2025:EFF

[LIA25] Jing Liu, Tomohito Imaizumi, and Kazuo Amakasu. Effects of fast Fourier transform window size on the target strength spectra measurements of Japanese anchovy (Engraulis japonicus) using linear frequency modulated signals. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000049.

Long:2024:USC

[LJB^{+}24] James M. Long, Patrick Joyce, Lindsey A. Bruckerhoff, Robert C. Lonsinger, and Wyatt Wolfenkoehler. Using downscan capabilities from recreational-grade side-scan sonar systems to sample paddlefish and evaluate depth use in a reservoir. Fisheries Research, 269(??):??, January 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002655.

Li:2020:SAI

[LJX^{+}20] Min Li, Yan Jiao, Binduo Xu, Chongliang Zhang, Ying Xue, and Yiping Ren. Spatial analyses of the influence of autocorrelation on seasonal diet composition of a marine fish species. Fisheries Research, 228(??):Article 105563, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300801.

Lurkpranee:2025:ADG

[LK25] Supatcha Lurkpranee and Toshihide Kitakado. Application of a delta-generalized additive model to assess the impact of environmental changes on the spatial distribution of bigeye tuna (Thunnus obesus) in the Indian Ocean. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003205.

Lin:2022:EES

[LKSiI22] Zhen Lin, Toshihide Kitakado, Naoki Suzuki, and Shin ichi Ito. Evaluation of the effects of stock enhancement on population dynamics using a state-space production model: a case study of Japanese flounder in the Seto Inland Sea. Fisheries Research, 251(??):Article 106299, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000765.

Liu:2020:GVP

[LLC^{+}20] Qiaohong Liu, Hungdu Lin, Jia Chen, Junkai Ma, Ruiqi Liu, and Shaoxiong Ding. Genetic variation and population genetic structure of the large yellow croaker (Larimichthys crocea) based on genome-wide single nucleotide polymorphisms in farmed and wild populations. Fisheries Research, 232(??):Article 105718, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302356.

Lira:2021:HFE

[LLFL21] Alex Souza Lira, Flávia Lucena-Frédou, and François Le Loc’h. How the fishing effort control and environmental changes affect the sustainability of a tropical shrimp small scale fishery. Fisheries Research, 235(??):Article 105824, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303416.

Ljungberg:2025:HPT

[LLH^{+}25] Peter Ljungberg, Sven-Gunnar Lunneryd, Lars Hillström, Glenn Fridh, and Mikael Lundin. The hovering pontoon trap: the tougher, younger sibling in the pontoon trap family. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002789.

Lehtonen:2022:FES

[LLK^{+}22] Esa Lehtonen, Roope Lehmonen, Joel Kostensalo, Mika Kurkilahti, and Petri Suuronen. Feasibility and effectiveness of seal deterrent in coastal trap-net fishing — development of a novel mobile deterrent. Fisheries Research, 252(??):Article 106328, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001059.

Lehtonen:2023:PCS

[LLS23] Esa Lehtonen, Roope Lehmonen, and Petri Suuronen. Potential of creating seal-free fishing areas with seal deterrents. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001297.

Lappalainen:2022:HCE

[LM22] Jyrki Lappalainen and Tommi Malinen. Hydroacoustics and concurrent experimental trawling reveal extreme annual variation in the density of 0+ pikeperch in late summer. Fisheries Research, 251(??):Article 106316, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000935.

Louw:2024:HPM

[LMG^{+}24] Nora R. Louw, Matthew N. McMillan, Naomi M. Gardiner, James Daniell, and Eric M. Roberts. Habitat partitioning in Moreton Bay bug species to inform fisheries management. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000201.

Linnane:2023:RYQ

[LMJ^{+}23] Adrian Linnane, Richard McGarvey, Annabel Jones, Nathan Kimber, John E. Feenstra, and Janet M. Matthews. Reflecting on 25 years of quota management in a spiny lobster fishery: Lessons learned from a harvest strategy perspective. Fisheries Research, 257(??):Article 106516, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002934.

Lejeune:2023:ADC

[LMK^{+}23] Benjamin Lejeune, Anna Marcout, Dorothée Kopp, Fabien Morandeau, Sonia Mehault, and Maud Aline Mouchet. Assessing discard consumption dynamic in shallow coastal environment using underwater video. Fisheries Research, 260(??):Article 106587, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622003642.

Linnane:2024:ACC

[LMM^{+}24] Adrian Linnane, Richard McGarvey, Janet M. Matthews, John E. Feenstra, Annabel Jones, and Nathan Kimber. Adapting to climate change in a spiny lobster (Jasus edwardsii) fishery: a harvest strategy response. Fisheries Research, 276(??):??, August 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001188.

Lee:2024:GPE

[LMP24] HuiHua Lee, Mark N. Maunder, and Kevin R. Piner. Good practices for estimating and using length-at-age in integrated stock assessments. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362300276X.

Labonne:2022:MPS

[LMT^{+}22] Maylis Labonne, Hicham Masski, Sophia Talba, Imane Tai, Khalid Manchih, Rachid Chfiri, and Raymond Lae. Major population’s separation area for sardine (Sardina pilchardus) and hake (Merluccius merluccius) revealed using otolith geochemistry on the Atlantic coast of Morocco. Fisheries Research, 254(??):Article 106415, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001928.

Legault:2023:DBS

[LNMA23] Ryan Legault, Emad Naseri, Elias Madadian, and Ali Ahmadi. Development of a biodegradable soft bait fishing lure. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001315.

Lacheheb:2025:ITC

[LNP25] Miloud Lacheheb, Ilan Noy, and Madhavi Pundit. Impact of tropical cyclones on fishing activity in the Philippines. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002856.

Liu:2021:SHS

[LNR^{+}21] Celine J. N. Liu, Sean Neo, Nathalia M. Rengifo, Ian French, Sarah Chiang, Mathias Ooi, Jie Min Heng, Nathaniel Soon, Jing Ying Yeo, Haaken Z. Bungum, Kurumi Ota, Arina A. Koul, Yan Hong Poh, and Benjamin J. Wainwright. Sharks in hot soup: DNA barcoding of shark species traded in Singapore. Fisheries Research, 241(??):Article 105994, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001223.

Lundgreen:2023:SCP

[LNvD^{+}23] Regitze B. C. Lundgreen, Anders Nielsen, Mikael van Deurs, Hans Jakob Olesen, Monica Mion, Stefanie Haase, Michele Casini, Uwe Krumme, and Karin Hüssy. Stock connectivity patterns and indications of sub-stock component structuring of cod in the Sound in the western Baltic Sea. Fisheries Research, 261(??):Article 106617, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000103.

Liang:2020:SSA

[LNW20] Dong Liang, Geneviève M. Nesslage, and Michael J. Wilberg. A spatial simulation approach to hydroacoustic survey design: a case study for Atlantic menhaden. Fisheries Research, 222(??):Article 105402, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302577.

Lant:2022:CFC

[LOFS22] Michael J. Lant, Derek H. Ogle, Zachary S. Feiner, and Greg G. Sass. A comparison of fish catch rates among local, non-local, and non-resident anglers of three northern Wisconsin lakes. Fisheries Research, 250(??):Article 106286, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000637.

Lorenzen:2022:SAD

[Lor22] Kai Lorenzen. Size- and age-dependent natural mortality in fish populations: Biology, models, implications, and a generalized length-inverse mortality paradigm. Fisheries Research, 255(??):Article 106454, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002314.

Luis:2023:LHI

[LP23] Sean M. Luis and Gregory B. Pasternack. Local hydraulics influence habitat selection and swimming behavior in adult California Central Valley Chinook salmon at a large river confluence. Fisheries Research, 261(??):Article 106634, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000279.

Ladino:2024:RMP

[LPAE^{+}24] A. Ladino, I. Pérez-Arjona, V. Espinosa, M. Chillarón, V. Vidal, L. M. Godinho, G. Moreno, and G. Boyra. Role of material properties in acoustical target strength: Insights from two species lacking a swimbladder. Fisheries Research, 270(??):??, February 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002886.

Larsen:2024:AHL

[LPG^{+}24] Martin H. Larsen, O. Jonas Palder, Casper Gundelund, Nicolas Azana Schnedler-Meyer, Henrik D. Ravn, and Christian Skov. Anatomical hooking location and bleeding occurrence in northern pike (Esox lucius) caught in recreational catchand-release angling in a lake with reduced prey fish availability. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002990.

Lischka:2020:DTE

[LPP^{+}20] Alexandra Lischka, Chris J. Pook, Jennifer L. Pannell, Heather E. Braid, Sally Gaw, and Kathrin S. R. Bolstad. Distribution of trace elements in the tissues of arrow squid (Nototodarus sloanii) from the Chatham Rise, New Zealand: Human health implications. Fisheries Research, 221(??):Article 105383, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302383.

Lishchenko:2021:RRS

[LPRB^{+}21] F. Lishchenko, C. Perales-Raya, C. Barrett, D. Oesterwind, A. M. Power, A. Larivain, V. Laptikhovsky, A. Karatza, N. Badouvas, A. Lishchenko, and G. J. Pierce. A review of recent studies on the life history and ecology of European cephalopods with emphasis on species with the greatest commercial fishery and culture potential. Fisheries Research, 236(??):Article 105847, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303647.

Layland:2025:LOE

[LPS^{+}25] Evelyn M. Layland, Emily Patrick, Molly Spencer, Rachel Lasley-Rasher, David M. Fields, and Richard A. Wahle. Larval ontogeny enhances resilience to a patchy planktonic food supply in the American lobster (Homarus americanus). Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002431.

Lejk:2021:EMS

[LR21] Adam M. Lejk and Grzegorz Radtke. Effect of marking Salmo trutta lacustris L. larvae with alizarin red S on their subsequent growth, condition, and distribution as juveniles in a natural stream. Fisheries Research, 234(??):Article 105786, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303039.

Lucas:2025:PRR

[LRGB25] Jorrit Lucas, Albert Ros, Juergen Geist, and Alexander Brinker. Predator responses in recreational fishing: Assessing selective pressure of bait types on behavioral diversity in northern pike (Esox lucius). Fisheries Research, 283(??): ??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000360.

Lai:2021:CEC

[LRMH21] Eva K. M. Lai, Karina L. Ryan, Ute Mueller, and Glenn A. Hyndes. Corroborating effort and catch from an integrated survey design for a boat-based recreational fishery in Western Australia. Fisheries Research, 236(??):Article 105865, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303829.

Lincoln:2024:PPS

[LRW^{+}24] H. Lincoln, P. E. Robins, S-B. Wilmes, S. Pérez-Mayol, A. Moore, S. Simpson, A. Goward-Brown, C. Heney, S. Malham, B. Morales-Nin, N. Hold, and I. D. McCarthy. Predicting potential spawning areas of European bass, Dicentrarchus labrax, in the Irish and Celtic seas. Fisheries Research, 270(??):??, February 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002771.

Liu:2024:APF

[LS24] Jen-Ming Liu and Po-Yuk So. Analysis of the positioning factors affecting the sustainable fishing gear-set net. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002588.

Lipscombe:2023:IBP

[LSM^{+}23] Rebecca S. Lipscombe, Anna Scott, Stephen Morris, Victor M. Peddemors, Amy F. Smoothey, and Paul A. Butcher. The influence of bait position on the catch of target and non-target sharks in a SMART drumline bather protection program. Fisheries Research, 257(??):Article 106501, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002788.

Li:2023:NFV

[LSZ^{+}23] Yiping Li, Lun Song, Sufang Zhao, Dongyang Zhao, Yingchao Wu, Guangran You, Zhongren Kong, Xiaohui Xi, and Zhe Yu. Nighttime fishing vessel observation in Bohai Sea based on VIIRS fishing vessel detection product (VBD). Fisheries Research, 258(??):Article 106539, February 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003162.

Lepine:2023:EBS

[LTE^{+}23] Tara M. Lepine, William M. Twardek, Brooke Etherington, Mike Dusevic, Luc LaRochelle, Andy J. Danylchuk, and Steven J. Cooke. The effectiveness of bite-shortened hooks for reducing handling time and injury of small-bodied freshwater fish captured by recreational anglers. Fisheries Research, 259(??):Article 106558, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003356.

Lorenzo:2020:PCE

[LTR20a] Rodrigo A. Lorenzo, Federico Tapella, and M. Carolina Romero. Pre-cooling effect on live transport of the southern king crab, Lithodes santolla. Fisheries Research, 227(??):Article 105552, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300692.

Lorenzo:2020:TMS

[LTR20b] Rodrigo A. Lorenzo, Federico Tapella, and M. Carolina Romero. Transportation methods for southern king crab: From fishing to transient storage and long-haul packaging. Fisheries Research, 223(??):Article 105441, March 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302966.

Liu:2023:PFC

[LTT^{+}23] Wei Liu, Hao Tang, Nyatchouba Nsangue Bruno Thierry, Jian Zhang, Feng Zhang, Meixi Zhu, Qiuyang Sun, Liuxiong Xu, and Fuxiang Hu. The profile and fluttering characteristics of a codend with different mesh sizes and catch by fast Fourier transform and Morlet wavelet methods. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001078.

Lund:2025:IPL

[Lun25] Ivar Lund. Influence of phospholipid and LC-PUFA content in extruded micro diets on European lobster larval (Homarus gammarus) performance and nutritional composition. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003199.

Li:2023:NGT

[LvCdGSL23] Peiwen Li, Peter van Coeverden de Groot, Zhengxin Sun, and Stephen C. Lougheed. A new genomics tool for monitoring Arctic char (Salvelinus alpinus) populations in the Lower Northwest Passage, Nunavut. Fisheries Research, 258(??):Article 106523, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003009.

Lahellec:2025:MEJ

[LVL25] Gabriel Lahellec, Youen Vermard, and Olivier Le Pape. Mapping essential juvenile habitats of exploited marine fish: Complementary insights from a scientific surveybased model, fishers’ knowledge and fisheries-dependent data. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002819.

Lankia:2022:IPA

[LVP22] Tuija Lankia, Riikka Venesjärvi, and Eija Pouta. Importanceperformance analysis of the fishing tourism service structure: Recreational anglers’ preferences on the remote salmon river of Teno in Finland. Fisheries Research, 254(??):Article 106425, October 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002028.

Lucchetti:2020:OGN

[LVPS20] A. Lucchetti, M. Virgili, A. Petetta, and P. Sartor. An overview of gill net and trammel net size selectivity in the Mediterranean Sea. Fisheries Research, 230 (??):Article 105677, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301946.

Lomeli:2021:UAI

[LWH^{+}21] Mark J. M. Lomeli, W. Waldo Wakefield, Bent Herrmann, Claude L. Dykstra, Anna Simeon, Dana M. Rudy, and Josep V. Planas. Use of artificial illumination to reduce Pacific halibut bycatch in a U.S. West Coast groundfish bottom trawl. Fisheries Research, 233(??):Article 105737, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362030254X.

Lewin:2023:CPM

[LWH^{+}23a] Wolf-Christian Lewin, Marc Simon Weltersbach, Kevin Haase, Robert Arlinghaus, and Harry V. Strehlow. Change points in marine recreational fisheries — the impact of stock status and fisheries regulations: a case from the western Baltic Sea. Fisheries Research, 258(??):Article 106548, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003253.

Lewin:2023:PBA

[LWH^{+}23b] Wolf-Christian Lewin, Marc Simon Weltersbach, Kevin Haase, Carsten Riepe, and Harry V. Strehlow. Potential biases in angler diary data: the impact of the diarist recruitment process on participation rates, catch, harvest, and effort estimates. Fisheries Research, 258(??):Article 106551, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003289.

Liang:2020:DSV

[LWX^{+}20] Jun Liang, Weiding Wang, Hanxiang Xu, Yongdong Zhou, Kaida Xu, Hongliang Zhang, and Kankai Lu. Diel and seasonal variation in fish communities in the zhongjieshan marine island reef reserve. Fisheries Research, 227(??):Article 105549, July 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300667.

Liu:2020:OST

[LXC^{+}20] Bi Lin Liu, Wei Xu, Xin Jun Chen, Meng Yao Huan, and Na Liu. Ontogenetic shifts in trophic geography of jumbo squid, Dosidicus gigas, inferred from stable isotopes in eye lens. Fisheries Research, 226(??):Article 105507, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300242.

Lyach:2020:ELS

[Lya20a] Roman Lyach. The effect of a large-scale angling restriction in minimum angling size on harvest rates, recapture rates, and average body weight of harvested common carps Cyprinus carpio. Fisheries Research, 223(??):Article 105438, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619302930.

Lyach:2020:EFE

[Lya20b] Roman Lyach. The effect of fishing effort, fish stocking, and population density of overwintering cormorants on the harvest and recapture rates of three rheophilic fish species in central Europe. Fisheries Research, 223(??):Article 105440, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619302954.

Lyach:2022:IDN

[Lya22] Roman Lyach. Increasing dominance of non-native fishes in the yield of central European streams and rivers. Fisheries Research, 254(??):Article 106433, October 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002107.

Lescher:2021:SSS

[LYH^{+}21] Cory Lescher, Noëlle Yochum, Brad Harris, Nathan Wolf, and John Gauvin. Selecting species specific vitality metrics to predict red king crab (Paralithodes camtschaticus) discard survival. Fisheries Research, 240(??):Article 105964, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621000928.

Litt:2021:AIA

[LYLC21] M. Aline Litt, Nathan Young, Nicolas W. R. Lapointe, and Steven J. Cooke. Angler interactions with American eel (Anguilla rostrata): Exploring perspectives and behaviors toward an imperiled fish. Fisheries Research, 234(??):Article 105781, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302988.

Liu:2021:CFA

[LYX^{+}21] Fei Liu, Fandong Yu, Zhijun Xia, Qiang Qin, Chunsen Xu, Jianwei Wang, and Huanzhang Liu. Changes in fish assemblages following the implementation of a complete fishing closure in the Chishui River. Fisheries Research, 243(??):Article 106099, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002277.

Li:2021:PVN

[LZC^{+}21] Jiajun Li, Peng Zhang, Yancong Cai, Qingling Zhang, Kui Zhang, Zhiyou Jing, Qiaer Wu, Yongsong Qiu, Shengwei Ma, and Zuozhi Chen. Performance of VMS and nightly satellite in monitoring light fishing vessels in the open South China Sea. Fisheries Research, 243(??):Article 106100, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002289.

Lu:2024:MGD

[LZCC24] Huajie Lu, Maolin Zhao, Ziyue Chen, and Xinjun Chen. The mechanism of gonadal development in response to environmental factors in dwarf-form individuals of female Sthenoteuthis oualaniensis in the South China Sea. Fisheries Research, 272(??):??, April 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000171.

Li:2024:DGP

[LZL^{+}24] Wenjia Li, Chi Zhang, Yang Liu, Shigang Liu, Hao Tian, Chang Cao, Yoshiro Watanabe, and Yongjun Tian. Different growth patterns reveal the potential origins of two Pacific saury (Cololabis saira) groups in the northwest Pacific Ocean. Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003260.

Liu:2021:GPF

[LZW^{+}21] Binwei Liu, Xi Zhang, Ziwei Wang, Weiyuan Li, Qi Zhang, Qi Liu, Wenlei Liu, Lei Zhang, Ying Liu, and Chenqi Wang. Genetic pattern fluctuations in wild swimming crab populations, under the influence of continuous mass stock enhancement. Fisheries Research, 243(??):Article 106075, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002034.

Luan:2020:PPR

[LZX^{+}20] Jing Luan, Chongliang Zhang, Binduo Xu, Ying Xue, and Yiping Ren. The predictive performances of random forest models with limited sample size and different species traits. Fisheries Research, 227(??):Article 105534, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300515.

Lei:2024:STN

[LZY24a] Yeming Lei, Shijie Zhou, and Nan Ye. Spatial-temporal neural networks for catch rate standardization and fish distribution modeling. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001619.

Lei:2024:SNN

[LZY24b] Yeming Lei, Shijie Zhou, and Nan Ye. Structured neural networks for CPUE standardization: a case study of the blue endeavour prawn in Australia’s northern prawn fishery. Fisheries Research, 279(??):??, November 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002042.

Liu:2024:BPR

[LZY^{+}24c] Yiwen Liu, Chongliang Zhang, Ji Yupeng, Binduo Xu, Ying Xue, and Yiping Ren. Bridging proxy reference points to sustainable yields to support data-limited fisheries. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002965.

Mabe:2020:CCA

[MA20] Franklin Nantui Mabe and Amos Asase. Climate change adaptation strategies and fish catchability: the case of inland artisanal fishers along the Volta Basin in Ghana. Fisheries Research, 230(??):Article 105675, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301922.

Moore:2020:DSSb

[MAA^{+}20] Bradley R. Moore, Tim Adams, Valerie Allain, Johann D. Bell, Mark Bigler, Don Bromhead, Sangaa Clark, Campbell Davies, Karen Evans, Ueta Faasili, Jessica Farley, Mark Fitchett, Peter M. Grewe, John Hampton, John Hyde, Bruno Leroy, Antony Lewis, Anne Lorrain, Jed I. Macdonald, Amandine D. Marie, Carolina Minte-Vera, Janice Natasha, Simon Nicol, Pablo Obregon, Thomas Peatman, Carlo Pecoraro, N. Bradley Phillip, Graham M. Pilling, Ciro Rico, Caroline Sanchez, Robert Scott, Joe Scutt Phillips, Brian Stockwell, Laura Tremblay-Boyer, Thomas Usu, Ashley J. Williams, and Neville Smith. Defining the stock structures of key commercial tunas in the Pacific Ocean II: Sampling considerations and future directions. Fisheries Research, 230(??):Article 105524, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300412.

Mir-Arguimbau:2020:GRT

[MABR^{+}20] Joan Mir-Arguimbau, Marc Balcells, Nuria Raventós, Paloma Martín, and Ana Sabatés. Growth, reproduction and their interplay in blue whiting (Micromesistius poutassou, Risso, 1827) from the NW Mediterranean. Fisheries Research, 227(??):Article 105540, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300576.

Machado:2022:IFI

[Mac22] Rodrigo Machado. Ingestion of fishing items by South American sea lion in southern Brazil: Indicating a neglected problem. Fisheries Research, 255(??):Article 106459, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002363.

MacDiarmid:2025:WAS

[Mac25] Alison MacDiarmid. What is an appropriate spatial scale for ecosystem based fishery management of kōura, spiny lobster Jasus edwardsii, in the Hauraki Gulf Marine Park, Aotearoa New Zealand? Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003254.

Moksness:2022:WLW

[MAH^{+}22] Erlend Moksness, Magnus Appelberg, Cornelius Hammer, Beatriz Morales Nin, and Peter J. Wright. What is left and what was achieved? A time perspective of a pioneering project 20 years after the European Fish Ageing Network. Fisheries Research, 252(??):Article 106340, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001175.

Maunder:2022:SRM

[Mau22] Mark N. Maunder. Stock-recruitment models from the viewpoint of density-dependent survival and the onset of strong density-dependence when a carrying capacity limit is reached. Fisheries Research, 249(??):Article 106249, May 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000261.

Mullowney:2020:GSR

[MB20] Darrell R. J. Mullowney and Krista D. Baker. Gone to shell: Removal of a million tonnes of snow crab since cod moratorium in the Newfoundland and Labrador fishery. Fisheries Research, 230(??):Article 105680, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301971.

Moan:2023:PRH

[MB23] André Moan and Arne Bjørge. Pingers reduce harbour porpoise bycatch in Norwegian gillnet fisheries, with little impact on day-to-day fishing operations. Fisheries Research, 259(??):Article 106564, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003411.

Millar:2025:IIS

[MB25] Russell B. Millar and Matt K. Broadhurst. Incorrect inference from size-selectivity studies due to widespread misuse of bootstrap confidence bands. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002893.

Moran:2023:CBQ

[MBB^{+}23] Damian Moran, Suzy E. Black, Erin Bell, Peter Bell, Benita Chambers, Sharon Ford, Jason Hamill, Greg Knox, Annalise Runarsson, Igor Ruza, Sebastian Horn, Louis Olsen, Jacqui Day, Samuel Thomas, Dave Woods, and Gerard Janssen. Catching better quality fish with novel codend technology: Precision Seafood Harvesting. Fisheries Research, 260(??):Article 106604, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003812.

Mullowney:2021:CIC

[MBD^{+}21] Darrell Mullowney, Krista Baker, Ben Davis, Katherine Skanes, Julia Pantin, William Coffey, Elizabeth Coughlan, Sanaollah Zabihi-Seissan, and Derek Osborne. Comments on ‘Increased catches of snow crab (Chionoecetes opilio) with luminescent-netting pots at long soak times by Nguyen et al. 2020’. Fisheries Research, 239(??):Article 105935, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000631.

Moore:2020:DSSa

[MBE^{+}20] Bradley R. Moore, Johann D. Bell, Karen Evans, Jessica Farley, Peter M. Grewe, John Hampton, Amandine D. Marie, Carolina Minte-Vera, Simon Nicol, Graham M. Pilling, Joe Scutt Phillips, Laura Tremblay-Boyer, Ashley J. Williams, and Neville Smith. Defining the stock structures of key commercial tunas in the Pacific Ocean I: Current knowledge and main uncertainties. Fisheries Research, 230(??):Article 105525, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300424.

Muir:2022:EPR

[MBH^{+}22] Jeff A. Muir, Richard J. Barker, Melanie R. Hutchinson, Bruno M. Leroy, Simon J. Nicol, and Joe Scutt Phillips. Estimating post-release mortality of long-line caught tropical tunas in the Pacific Ocean. Fisheries Research, 249(??):Article 106194, May 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003222.

Mainguy:2024:MRF

[MBOCdAM24] Julien Mainguy, Martin Bélanger, Geneviève Ouellet-Cauchon, and Rafael de Andrade Moral. Monitoring reproduction in fish: Assessing the adequacy of ogives and the predicted uncertainty of their L_{50} estimates for more reliable biological inferences. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002564.

Mullowney:2020:HSD

[MBP20] Darrell R. J. Mullowney, Krista D. Baker, and Eric J. Pedersen. Harvesting strategies during a forecasted decline in the Newfoundland and Labrador snow crab fishery. Fisheries Research, 232(??):Article 105707, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302241.

Mullowney:2020:BPC

[MBZSM20] Darrell R. J. Mullowney, Krista D. Baker, Sana Zabihi-Seissan, and Corey Morris. Biological perspectives on complexities of fisheries co-management: a case study of Newfoundland and Labrador snow crab. Fisheries Research, 232(??):Article 105728, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302459.

Morales:2020:MTS

[MCC20] Nia Morales, Ed Camp, and Matteo Cleary. Modeling target species selection behavior among diverse freshwater anglers in Florida, USA. Fisheries Research, 231(??):Article 105693, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302101.

Mas:2022:SRR

[MCC^{+}22] Federico Mas, Enric Cortés, Rui Coelho, Omar Defeo, Rodrigo Forselledo, Sebastián Jiménez, Philip Miller, and Andrés Domingo. Shedding rates and retention performance of conventional dart tags in large pelagic sharks: Insights from a double-tagging experiment on blue shark (Prionace glauca). Fisheries Research, 255(??):Article 106462, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002399.

Mas:2023:NIR

[MCC^{+}23] F. Mas, E. Cortés, R. Coelho, O. Defeo, R. Forselledo, and A. Domingo. New insights into the reproductive biology of the blue shark (Prionace glauca) in the South Atlantic Ocean. Fisheries Research, 262(??):Article 106643, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300036X.

Manjabacas:2025:TYA

[MCGL^{+}25] A. Manjabacas, Ò. Chic, E. García-Ladona, V. M. Tuset, J. R. Morros, E. Sayrol, P. Martí-Puig, J. Forest, J. Vasconcelos, and A. Lombarte. Twenty years of AFORO:

New developments and connections enhancing otolith research. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003060.

Msukwa:2021:VAL

[MCH21] Amulike V. Msukwa, Ian G. Cowx, and Jonathan P. Harvey. Vulnerability assessment of Lake Malawi’s ornamental fish resources to export ornamental trade. Fisheries Research, 238(??):Article 105869, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303866.

Munian:2021:ABB

[MCHA21] André Munian, Andrew Cockcroft, Louwrens C. Hoffman, and Lutz Auerswald. Analysis of biological and biochemical parameters of adult male spiny lobsters Jasus lalandii for identification of possible growth predictors. Fisheries Research, 243(??):Article 106061, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001892.

Michailidis:2023:ERF

[MCK23] Nikolas Michailidis, Niki Chartosia, and Stelios Katsanevakis. Exploring the role of fishing in a heavily bioinvaded shelf ecosystem. Fisheries Research, 259(??):Article 106554, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003319.

Morris:2020:ESS

[MCMM20] Corey J. Morris, David Cote, S. Bruce Martin, and Darrell Mullowney. Effects of 3D seismic surveying on snow crab fishery. Fisheries Research, 232(??):Article 105719, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620302368.

Martinez-Candelas:2020:UHD

[MCPJET^{+}20] I. A. Martínez-Candelas, J. C. Pérez-Jiménez, A. Espinoza-Tenorio, L. McClenachan, and I. Méndez-Loeza. Use of historical data to assess changes in the vulnerability of sharks. Fisheries Research, 226(??):Article 105526, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300436.

MacNamara:2022:OMS

[MCS^{+}22] Ruairi MacNamara, Edward Camp, Michael Shane, Kai Lorenzen, and Mark Drawbridge. Optimizing marine stock enhancement through modeling: a sex-specific application with California halibut Paralichthys californicus. Fisheries Research, 252(??):Article 106341, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001187.

Ma:2024:EML

[MCS^{+}24] Tsung-Hsiang Ma, Yi-Jay Chang, Jen-Chieh Shiao, Chien-Bang Jin, and Yan-Fu Kuo. Enhancing machine learningbased age estimation for Pacific bluefin tuna: an approach with data imputation and image augmentation strategies. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000560.

Maunder:2025:CAP

[MCS^{+}25] Mark N. Maunder, Paul R. Crone, Brice X. Semmens, Juan L. Valero, Lynn Waterhouse, Richard D. Methot, and André E. Punt. The Center for the Advancement of Population Assessment Methodology (CAPAM): a perspective on the first 10 years. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002261.

Moore:2022:STV

[MDC^{+}22a] Alec B. M. Moore, Adam J. Delargy, Ruth P. Cann, Charlotte Heney, Lewis Le Vay, Harriet Lincoln, Ian D. McCarthy, and Natalie Hold. Spatial and temporal variation of size at maturity in an intensive crustacean fishery with limited management. Fisheries Research, 255(??):Article 106450, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002272.

Mosley:2022:SDH

[MDC^{+}22b] Camille L. Mosley, Colin J. Dassow, John Caffarelli, Alexander J. Ross, Greg G. Sass, Stephanie L. Shaw, Christopher T. Solomon, and Stuart E. Jones. Species differences, but not habitat, influence catch rate hyperstability across a recreational fishery landscape. Fisheries Research, 255(??):Article 106438, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002156.

Martins:2020:ERP

[MdCG20] Rodrigo Silvestre Martins, Ricardo de Camargo, and Maria A. Gasalla. Effect of retention processes on the recruitment of tropical arrow squid (Doryteuthis pleii): an individual-based modeling case study in southeastern Brazil. Fisheries Research, 224(??):Article 105455, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303108.

Mesquita:2024:EFE

[MDJP24] Carlos Mesquita, Helen Dobby, Catherine S. Jones, and Graham J. Pierce. Estimating fishing effort and LPUE for the Scottish brown crab (Cancer pagurus) trap fishery using VMS and observer data. Fisheries Research, 274(??): ??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000389.

Macchi:2021:TME

[MDL^{+}21] Gustavo J. Macchi, Marina V. Diaz, Ezequiel Leonarduzzi, Martín Ehrlich, Laura Machinandiarena, Mariana Cadaveira, and María Inés Militelli. Temperature, maternal effects and density-dependent processes during early life stages of Argentine hake as relevant recruitment drivers. Fisheries Research, 238(??):Article 105898, June 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000266.

Maynou:2021:AMA

[MDMS21] F. Maynou, M. Demestre, P. Martín, and P. Sánchez. Application of a multi-annual generalized depletion model to the Mediterranean sandeel fishery in Catalonia. Fisheries Research, 234(??):Article 105814, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303313.

Mesquita:2020:SMB

[MDS^{+}20] Carlos Mesquita, Helen Dobby, Stephanie Sweeting, Catherine S. Jones, and Graham J. Pierce. Size-at-maturity of brown crab (Cancer pagurus) in Scottish waters based on gonadal and morphometric traits. Fisheries Research, 229(??):Article 105610, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301272.

Mace:2021:SPD

[MDW^{+}21] Marvin M. Mace, Kathryn L. Doering, Michael J. Wilberg, Amy Larimer, Frank Marenghi, Alexei Sharov, and Mitchell Tarnowski. Spatial population dynamics of eastern oyster in the Chesapeake Bay, Maryland. Fisheries Research, 237(??):Article 105854, May 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303714.

Melvin:2023:SIF

[Mel23] Gary D. Melvin. Special issue Fisheries Research — experience, results and lessons learned from an Oceanic Tuna Tagging Campaign. Fisheries Research, 260(??):Article 106588, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003654.

Munley:2024:LIA

[MFJ^{+}24] Mary Kate Munley, Elizabeth A. Fairchild, Steven H. Jury, Winsor H. Watson, and Shelley A. Edmundson. Laboratory investigations into alternative baits for the channeled whelk (Busycotypus canaliculatus) fishery. Fisheries Research, 271(??):??, March 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623003132.

Militelli:2020:RPS

[MFM^{+}20] María I. Militelli, Carla Firpo, A. Cecilia Mauna, Karina A. Rodrigues, and Gustavo J. Macchi. Reproductive potential of southern king crab (Lithodes santolla) in South Patagonian Sector (south 48^{o}S), a new fishery area. Fisheries Research, 229(??):Article 105595, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301120.

Malafaia:2021:SAS

[MFO21] Priscilla Nogueira Malafaia, Aline Rocha França, and George Olavo. Spawning aggregation sites of the cubera snapper, Lutjanus cyanopterus, on the continental shelf of Bahia state, Northeastern Brazil. Fisheries Research, 242(??):Article 106037, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100165X.

Marine:2022:DBC

[MFR22a] Sabiha Sultana Marine, Nicole Flint, and John Rolfe. Do buyouts of commercial licences increase satisfaction of recreational fishers? A paired comparison of two Queensland zones. Fisheries Research, 254(??):Article 106428, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002053.

Motta:2022:DES

[MFR^{+}22b] Fabio S. Motta, Matheus O. Freitas, Fernanda A. Rolim, Vinícius Abilhoa, and Guilherme H. Pereira Filho. Direct evidence of a spawning aggregation of cubera snapper (Lutjanus cyanopterus) in southeastern Brazil and its management implications. Fisheries Research, 252(??):Article 106339, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001163.

Maia:2024:GRS

[MGB24] Francisco Maia, Miguel B. Gaspar, and Carlos M. Barroso. Growth, reproduction, and size at first maturity of the peppery furrow shell Scrobicularia plana in the Ria de Aveiro, Portugal: Implications for sustainable fisheries management. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000420.

Morrissey:2022:PGR

[MGC^{+}22] Declan Morrissey, Jake Goodall, Rita Castilho, Tom C. Cameron, and Michelle L. Taylor. Population genomics reveals a single semi-continuous population of a commercially exploited marine gastropod. Fisheries Research, 254(??):Article 106418, October 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001953.

Munnich:2023:RCC

[MHB^{+}23] Jasper Münnich, Frederike Hoppmann, Hanna Berggren, Oscar Nordahl, and Petter Tibblin. The role of chemical communication in the predator–prey role reversal of northern pike (Esox lucius) and three-spined stickleback (Gasterosteus aculeatus). Fisheries Research, 258(??):Article 106537, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003149.

Marks:2021:FID

[MHD^{+}21] Rachel Marks, Sybrand A. Hesp, Ainslie Denham, Neil R. Loneragan, Danielle Johnston, and Norman Hall. Factors influencing the dynamics of a collapsed blue swimmer crab (Portunus armatus) population and its lack of recovery. Fisheries Research, 242(??):Article 106035, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001636.

Mion:2020:HGE

[MHH^{+}20] Monica Mion, Annelie Hilvarsson, Karin Hüssy, Uwe Krumme, Maria Krüger-Johnsen, Kate McQueen, Esha Mohamed, Roman Motyka, Alessandro Orio, Maris Plikshs, Krzysztof Radtke, and Michele Casini. Historical growth of Eastern Baltic cod (Gadus morhua): Setting a baseline with international tagging data. Fisheries Research, 223(??):Article 105442, March 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302978.

Maunder:2023:REM

[MHL^{+}23a] Mark N. Maunder, Owen S. Hamel, Hui-Hua Lee, Kevin R. Piner, Jason M. Cope, André E. Punt, James N. Ianelli, Claudio Castillo-Jordán, Maia S. Kapur, and Richard D. Methot. A review of estimation methods for natural mortality and their performance in the context of fishery stock assessment. Fisheries Research, 257(??):Article 106489, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002661.

Moore:2023:BNA

[MHL^{+}23b] Alec B. M. Moore, Charlotte Heney, Harriet Lincoln, Charlotte Colvin, Hadley Newell, Rebecca Turner, Ian D. Mc-Carthy, and Natalie Hold. Bycatch in northeast Atlantic lobster and crab pot fisheries (Irish Sea, Celtic Sea and Bristol Channel). Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001388.

Madsen:2021:ISS

[MINS21] Niels Madsen, Ólafur A. Ingólfsson, Hans Nilsson, and Petri Suuronen. Improving species and size selectivity in the Baltic cod trawl fishery with two simple codend designs. Fisheries Research, 236(??):Article 105846, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303635.

Mazur:2023:CIC

[MJC^{+}23] Mackenzie D. Mazur, Jerelle Jesse, Steven X. Cadrin, Samuel B. Truesdell, and Lisa Kerr. Consequences of ignoring climate impacts on New England groundfish stock assessment and management. Fisheries Research, 262(??):Article 106652, June 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000450.

Macdonald:2021:RQO

[MJD^{+}21] Jed I. Macdonald, Ingibjörg G. Jónsdóttir, Russell N. Drysdale, Roman Witt, Torsteinn Sigurdsson, Gudmundur J. Óskarsson, Zsófia Cságoly, and Gudrún Marteinsdóttir. Rules for quantifying otolith chemical variability help expose nursery population structure, site fidelity and multiple origins in a fished herring stock. Fisheries Research, 242(??):Article 106040, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001685.

Michailidis:2020:RFC

[MKC20] Nikolas Michailidis, Stelios Katsanevakis, and Niki Chartosia. Recreational fisheries can be of the same magnitude as commercial fisheries: the case of Cyprus. Fisheries Research, 231(??):Article 105711, November 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302289.

Muller-Karanassos:2021:EHF

[MKFF^{+}21] Christina Muller-Karanassos, Alex Filous, Alan M. Friedlander, Javier Cuetos-Bueno, Marine Gouezo, Steven J. Lindfield, Victor Nestor, Lincy Lee Marino, Geory Mereb, Dawnette Olsudong, and Yimnang Golbuu. Effects of habitat, fishing, and fisheries management on reef fish populations in Palau. Fisheries Research, 241(??):Article 105996, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001247.

Melli:2020:DEC

[MKH^{+}20] Valentina Melli, Ludvig A. Krag, Bent Herrmann, Junita D. Karlsen, and Jordan P. Feekings. Does the efficiency of a counter-herding device depend on seabed contact? Fisheries Research, 230(??):Article 105686, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302034.

McManus:2021:MBA

[MKS^{+}21a] M. Conor McManus, Jeff Kipp, Burton Shank, Kathleen Reardon, Tracy L. Pugh, Josh Carloni, and Kim McKown. A model-based approach to standardizing American lobster (Homarus americanus) ventless trap abundance indices. Fisheries Research, 238(??):Article 105899, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000278.

Minami:2021:ACP

[MKS^{+}21b] Kenji Minami, Chihomi Kita, Hokuto Shirakawa, Yohei Kawauchi, Huamei Shao, Makoto Tomiyasu, Yuka Iwahara, Hideo Takahara, Takashi Kitagawa, and Kazushi Miyashita. Acoustic characteristics of a potentially important macroalgae, Sargassum horneri, for coastal fisheries. Fisheries Research, 240(??):Article 105955, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000837.

McDonald:2022:IIA

[MKS^{+}22] Raphaël R. McDonald, David M. Keith, Jessica A. Sameoto, Jeffrey A. Hutchings, and Joanna Mills Flemming. Incorporating intra-annual variability in fisheries abundance data to better capture population dynamics. Fisheries Research, 246(??):Article 106152, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002800.

Miller:2024:HCA

[ML24] Marcus E. Miller and Geoffrey W. Liggins. Hermit crabs associated with catches from the eastern rock lobster (Sagmariasus verreauxi) fishery along the coast of NSW, Australia. Fisheries Research, 280(??):??, December 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002182.

Majumdar:2023:STM

[MLCMdS23] Anandamayee Majumdar, Cleridy E. Lennert-Cody, Mark N. Maunder, and Alexandre Aires da Silva. Spatio-temporal modeling for estimation of bigeye tuna catch in the presence of pandemic-related data loss using parametric adjacency structures. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002060.

Mehdi:2021:CPA

[MLS^{+}21] Hossein Mehdi, Samantha C. Lau, Caitlyn Synyshyn, Matthew G. Salena, Markelle E. Morphet, Jonathan Hamilton, Melissa N. Muzzatti, Erin S. McCallum, Jonathan D. Midwood, and Sigal Balshine. A comparison of passive and active gear in fish community assessments in summer versus winter. Fisheries Research, 242(??):Article 106016, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001442.

Munroe:2023:SHB

[MMBH23] Daphne Munroe, Jason Morson, Sarah Borsetti, and Daniel Hennen. Sampling high biomass but rare benthic animals: Methods for surveying commercial clam stocks using a hydraulic dredge. Fisheries Research, 258(??):Article 106538, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003150.

Maldonado:2024:WLC

[MMC^{+}24] Matthew L. Maldonado, Taufique H. Mahmood, David P. Coulter, Alison A. Coulter, Steve R. Chipps, Maddy K.

Siller, Michaela L. Neal, Ayon Saha, and Mark A. Kaemingk. Water-level changes impact angler effort in a large lake: Implications for climate change. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002200.

Marcotegui:2024:ISE

[MMG^{+}24] P. Marcotegui, M. J. Merlo, M. P. Gutiérrez, C. C. Buratti, M. Parietti, and J. T. Timi. An independent stock of Engraulis anchoita in north Patagonian waters? Parasite tags as evidence. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002121.

Moura:2020:PSD

[MMM^{+}20] A. Moura, A. A. Muniz, E. Mullis, J. M. Wilson, R. P. Vieira, A. A. Almeida, E. Pinto, G. J. A. Brummer, P. V. Gaever, J. M. S. Gonçalves, and A. T. Correia. Population structure and dynamics of the Atlantic mackerel (Scomber scombrus) in the North Atlantic inferred from otolith chemical and shape signatures. Fisheries Research, 230(??):Article 105621, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620301387.

McLeay:2024:PBA

[MMML24] Lachlan McLeay, Kevin Mark, Richard McGarvey, and Adrian Linnane. Pass the batten! Alternative pot design increases catch efficiency in a Southern Rock Lobster fishery. Fisheries Research, 280(??):??, December 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002285.

McGarvey:2024:PGI

[MMP^{+}24] Richard McGarvey, Richard D. Methot, André E. Punt, Janet M. Matthews, Ian G. Taylor, John E. Feenstra, and Kathryn Doering. Performance gains from incorporating dynamic numbers by length-within-age in fishery assessment models. Fisheries Research, 276(??):??, August 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001036.

Marsden:2021:GMC

[MMQ21] J. Ellen Marsden and Benjamin Marcy-Quay. A generalized model for correcting bias due to permeability using emergent fry traps in mesocosm experiments. Fisheries Research, 233(??):Article 105769, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302861.

Morales-Nin:2022:EHM

[MNPMM^{+}22] Beatriz Morales-Nin, Sílvia Pérez-Mayol, Kirsteen MacKenzie, Ignacio A. Catalán, Miquel Palmer, Thibault Kersaudy, and Kélig Mahé. European hake (Merluccius merluccius) stock structure in the Mediterranean as assessed by otolith shape and microchemistry. Fisheries Research, 254(??):Article 106419, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001965.

Miranda:2020:SRD

[MNS^{+}20] L. E. Miranda, D. M. Norris, V. R. Starnes, N. M. Faucheux, and T. Holman. Stock-recruitment dynamics of a freshwater clupeid. Fisheries Research, 221(??):Article 105378, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302334.

McCallister:2023:FCA

[MOA23] Michael P. McCallister, James Oppenborn, and Matthew J. Ajemian. Fish community assessment of shelf-edge artificial reefs along east-central Florida. Fisheries Research, 259(??):Article 106561, March 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003381.

Morrow:2022:IEE

[MOB^{+}22] Benjamin D. Morrow, Patrick D. O’Hara, Natalie C. Ban, Tunai P. Marques, Molly D. Fraser, Norma S. Serra-Sogas, and Christopher E. Bone. Improving effort estimates and informing temporal distribution of recreational salmon fishing in British Columbia, Canada using high-frequency optical imagery data. Fisheries Research, 249(??):Article 106251, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000285.

Millar:2023:SSN

[MOB^{+}23] Russell B. Millar, Richard L. O’Driscoll, Suzy Black, Gerard Janssen, Jason Hamill, David Woods, and Damian Moran. Size selectivity of a novel non-mesh codend (the Modular Harvesting System) in a New Zealand deepwater fishery. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300098X.

Medina-Ortega:2025:SME

[MOCGCC^{+}25] Leticia Medina-Ortega, Jairo Castro-Gutiérrez, Remedios Cabrera-Castro, Anyell Caderno, and Francisco Hortas. Seasonal morphometry and extractive pressure on polychaetes Hediste diversicolor and Marphysa sanguinea from Southwest Spain. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003163.

Miyagawa:2023:WFB

[MOI23] Mitsuyo Miyagawa, Hiroshi Okamura, and Momoko Ichinokawa. Which fisheries and biological factors affect the misclassification of stock status determined by data-limited methods? Fisheries Research, 257(??):Article 106491, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002685.

Monteiro:2025:BTP

[MOM^{+}25] João N. Monteiro, Andreia Ovelheiro, Francisco Maia, Maria Alexandra Teodósio, and Francisco Leitão. Biological traits and population dynamics for sustainable harvesting of Carcinus maenas. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624003072.

Monnahan:2024:TGP

[Mon24] Cole C. Monnahan. Toward good practices for Bayesian data-rich fisheries stock assessments using a modern statistical workflow. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000882.

Monteiro:2025:ISS

[MOTL25] João N. Monteiro, Andreia Ovelheiro, Maria Alexandra Teodósio, and Francisco Leitão. Impact and size selectivity of fishing gears used in estuarine crab fisheries. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000219.

Moreira:2020:STM

[MPC^{+}20] Cláudia Moreira, Pablo Presa, Alberto Teodorico Correia, Paulo Vaz-Pires, and Elsa Froufe. Spatio-temporal microsatellite data suggest a multidirectional connectivity pattern in the Trachurus picturatus metapopulation from the Northeast Atlantic. Fisheries Research, 225(??):Article 105499, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300163.

Mota:2023:DEP

[MPEBdR23] Thaís Arrais Mota, Marcelo Antonio Amaro Pinheiro, Norma Suely Evangelista-Barreto, and Sérgio Schwarz da Rocha. Density and extractive potential of “uçá”- crab, Ucides cordatus (Linnaeus, 1763), in mangroves of the “Todos os Santos” Bay, Bahia, Brazil. Fisheries Research, 265(??):??, September 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001261.

McQuaw:2021:EAR

[MPH21] Kristin McQuaw, André E. Punt, and Ray Hilborn. Evaluating alternative rebuilding plans for mixed stock fisheries. Fisheries Research, 240(??):Article 105984, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001120.

Martinez-Pita:2022:MPR

[MPM22] Inés Martínez-Pita and Oscar Moreno. A method to predict the reproductive cycle of the striped venus clam Chamelea gallina based on the influence of environmental factors: Application in its fishery management. Fisheries Research, 245(??):Article 106133, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002617.

Millot:2023:RTS

[MPM^{+}23] Rémi Millot, François Poisson, David Macías, Sámar Saber, Antoine Aiello, and Eric Dominique Henri Durieux. Reproductive traits and spawning activity of swordfish Xiphias gladius L. in the north-western Mediterranean Sea (Corsica). Fisheries Research, 267(??):??, November 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002047.

Mormede:2020:CSD

[MPP20] Sophie Mormede, Steven J. Parker, and Matthew H. Pinkerton. Comparing spatial distribution modelling of fisheries data with single-area or spatially-explicit integrated population models, a case study of toothfish in the Ross Sea region. Fisheries Research, 221(??):Article 105381, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361930236X.

Moore:2022:OST

[MPP22] Bradley R. Moore, Steven J. Parker, and Matthew H. Pinkerton. Otolith shape as a tool for species identification of the grenadiers Macrourus caml and M. whitsoni. Fisheries Research, 253(??):Article 106370, September 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001473.

McWhinnie:2025:MIM

[MPSH25] Stephanie McWhinnie, Sean Pascoe, Peggy Schrobback, and Eriko Hoshino. Measuring, interpreting and monitoring economic efficiency in South Australia’s Spencer Gulf and West Coast prawn fisheries. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002558.

Maunder:2025:SAG

[MPSM25] Mark N. Maunder, Andre E. Punt, Rishi Sharma, and Richard D. Methot. Stock assessment good practices: the crescendo of CAPAM’s workshop series and their consequent special issues. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002753.

Mota:2024:PGD

[MPV^{+}24] Kenneth Gabriel Mota, Caio Augusto Perazza, Filomena Vaz Velho, Maria de Lourdes Sardinha, Letícia Morais, Jussara Oliveira Vaini, Eric Hallerman, and Alexandre Wagner Silva Hilsdorf. Population genetic diversity of Cunene horse mackerel Trachurus trecae on the Angolan coast: Implications for management and conservation. Fisheries Research, 277(??):??, September 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001334.

Mohanraj:2024:LWR

[MRC24] T. Mohanraj, T. Jebarani Rajathy, and S. R. T. Sherly Cross. Length-weight relationship, condition factor, and diet analysis of thresher sharks (family: Alopiidae) along the southern coast of India. Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001310.

Monaghan:2024:FBE

[MRE^{+}24] Erin Monaghan, Phillip Ravanello, David Ellis, Jessica A. Bolin, David Schoeman, and Kylie L. Scales. Fishing behaviour and environmental variability influence depredation of pelagic longline catch by toothed whales. Fisheries Research, 273(??):??, May 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000237.

Mendo:2023:MDW

[MRG^{+}23] T. Mendo, J. M. Ransijn, I. Gomez, R. Gozzer-Wuest, I. Paradinas, M. James, and J. Mendo. Minimising discards while taking revenue into account: Spatio-temporal assessment of catches in an artisanal shrimp trawl fishery in Peru. Fisheries Research, 261(??):Article 106623, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000164.

Miller:2023:ESE

[MRP^{+}23] Timothy J. Miller, David E. Richardson, Philip J. Politis, Christopher D. Roebuck, John P. Manderson, Michael H. Martin, and Andrew W. Jones. Estimation of survey efficiency and biomass for commercially important species from industry-based paired gear experiments. Fisheries Research, 259(??):Article 106565, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003423.

Mandal:2025:PHL

[MRS^{+}25] Anup Kumar Mandal, Md. Mamun Or Rashid, Md. Sujahangir Kabir Sarkar, Badiuzzaman, Md. Sazedul Hoque, and Afjal Hossain. Post-harvest losses of marine fish at fishers in Bangladesh. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002625.

Molinet:2023:IST

[MRUG^{+}23] Carlos Molinet, Rubén H. Roa-Ureta, Paulina Gebauer, Manuel Díaz, Patricio A. Díaz, Thamara Matamala, Katherine Espinoza, Jorge Henríquez, Daniela Uribe, Oscar de Lázaro, Andrés Olguín, Kurt Paschke, José Valenzuela, and Yohnatan Jaramillo. The impact of size truncation on reproductive success in the southern king crab (Lithodes santolla). Fisheries Research, 258(??):Article 106522, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002995.

Marcalo:2024:EDI

[MSC^{+}24] Ana Marçalo, Vighnesh Samel, Flávia Carvalho, Magda Frade, Karim Erzini, and Jorge MS Gonçalves. Evaluating dolphin interactions with bottom-set net fisheries off Southern Iberian Atlantic waters. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001644.

Martino:2021:SSD

[MSD21] Jasmin C. Martino, Michael Steer, and Zoë A. Doubleday. Supporting the sustainable development of Australia’s octopus industry: First assessment of an artisanal fishery. Fisheries Research, 241(??):Article 105999, September 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001272.

Mota:2021:PRL

[MSJ21] Vasco C. Mota, Sten I. Siikavuopio, and Philip James. Physiological responses to live air transport of red king crab (Paralithodes camtschaticus). Fisheries Research, 237(??):Article 105882, May 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621000102.

Morales-Serna:2024:PIP

[MSLMOC^{+}24] Francisco Neptalí Morales-Serna, Dania López-Moreno, Juan M. Osuna-Cabanillas, Emigdio Marín-Enríquez, Juan R. F. Vallarta-Zárate, and Felipe Amezcua. Parasites as indicators of population connectivity of a small pelagic fish in the Gulf of California. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400122X.

McGarvey:2021:GPE

[MSS^{+}21] Richard McGarvey, Mike A. Steer, Jonathan J. Smart, Damian J. Matthews, and Janet M. Matthews. Generalizing the Parker equation of DEPM: Incorporating the size dependence of population number and reproductive inputs to estimate spawning biomass and female population by size. Fisheries Research, 242(??):Article 105992, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362100120X.

Mamet:2021:MSP

[MSV21] Leandro Nicolás Getino Mamet, Gaspar Soria, and Adrián Munguía Vega. Multiple-scale processes shape the population genetics of Tehuelche scallop, Aequipecten tehuelchus, in Northern Patagonia. Fisheries Research, 240(??):Article 105971, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621000990.

Maahs:2021:ENL

[MSW21] Brandon Maahs, Andrea Sylvia, and Michael J. Weber. Effects of number of largemouth bass per live-well compared to environmental conditions on post release tournament survival. Fisheries Research, 243(??):Article 106101, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002290.

Maiello:2022:LSB

[MTC^{+}22] Giulia Maiello, Lorenzo Talarico, Paolo Carpentieri, Flavio De Angelis, Simone Franceschini, Lynsey R. Harper, Erika F. Neave, Olga Rickards, Alice Sbrana, Peter Shum, Virginia Veltre, Stefano Mariani, and Tommaso Russo. Little samplers, big fleet: eDNA metabarcoding from commercial trawlers enhances ocean monitoring. Fisheries Research, 249(??):Article 106259, May 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000364.

Muallil:2020:ICI

[MTE^{+}20] Richard N. Muallil, Ahalnida M. Tambihasan, Marylyn J. Enojario, Yunadzmal N. Ong, and Cleto L. Nañola. Inventory of commercially important coral reef fishes in Tawi-Tawi Islands, Southern Philippines: the Heart of the Coral Triangle. Fisheries Research, 230(??):Article 105640, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301570.

Major:2021:DFM

[MTS^{+}21] Robert N. Major, David I. Taylor, Michael Scott, John Radford, Stephen Connor, Geoffrey Connor, and Shaun Ogilvie. Determining the fundamental metrics of pot entrances for New Zealand scampi, Metanephrops challengeri. Fisheries Research, 233(??):Article 105760, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302770.

Maunder:2020:NST

[MTX^{+}20] Mark N. Maunder, James T. Thorson, Haikun Xu, Ricardo Oliveros-Ramos, Simon D. Hoyle, Laura Tremblay-Boyer, Hui Hua Lee, Mikihiko Kai, Shui-Kai Chang, Toshihide Kitakado, Christoffer M. Albertsen, Carolina V. Minte-Vera, Cleridy E. Lennert-Cody, Alexandre M. Aires da Silva, and Kevin R. Piner. The need for spatio-temporal modeling to determine catch-per-unit effort based indices of abundance and associated composition data for inclusion in stock assessment models. Fisheries Research, 229(??):Article 105594, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620301119.

Merino:2022:ITP

[MUF^{+}22] Gorka Merino, Agurtzane Urtizberea, Dan Fu, Henning Winker, Massimiliano Cardinale, Matthew V. Lauretta, Hilario Murua, Toshihide Kitakado, Haritz Arrizabalaga, Robert Scott, Graham Pilling, Carolina Minte-Vera, Haikun Xu, Ane Laborda, Maite Erauskin-Extramiana, and Josu Santiago. Investigating trends in process error as a diagnostic for integrated fisheries stock assessments. Fisheries Research, 256(??):Article 106478, December 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002557.

Munyandorero:2024:OUC

[Mun24] Joseph Munyandorero. Other useful considerations on sustainable benchmarks associated with two-stage-structured models of fisheries. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000687.

Murphy:2020:CCI

[Mur20] James T. Murphy. Climate change, interspecific competition, and poleward vs. depth distribution shifts: Spatial analyses of the eastern Bering Sea snow and Tanner crab (Chionoecetes opilio and C. bairdi). Fisheries Research, 223(??):Article 105417, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302723.

Murphy:2021:TSV

[Mur21] James T. Murphy. Temporal and spatial variability in sizeat-maturity for the eastern Bering Sea snow and Tanner crab (Chionoecetes opilio and C. bairdi). Fisheries Research, 234(??):Article 105761, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302782.

Morgan:2024:ERA

[MVDH24] Linnéa Morgan, Daniel Valentinsson, Thomas G. Dahlgren, and Sara Hornborg. Ecological risk assessment of invertebrates caught in Swedish west-coast fisheries. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000468.

Molina-Valdivia:2020:STV

[MVLC^{+}20] Victor Molina-Valdivia, Mauricio F. Landaeta, Manuel I. Castillo, Darly Alarcón, and Guido Plaza. Short-term variations in the early life history traits of common sardine Strangomera bentincki and anchoveta Engraulis ringens off central Chile. Fisheries Research, 224(??):Article 105460, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303157.

Minte-Vera:2024:UCM

[MVMdS^{+}24] Carolina V. Minte-Vera, Mark N. Maunder, Alexandre Aires da Silva, Haikun Xu, Juan L. Valero, Steven L. H. Teo, Patrício Barría, and Nicholas D. Ducharme-Barth. The use of conceptual models to structure stock assessments: a tool for collaboration and for “modelling what to model”. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001991.

Mace:2024:DFS

[MWJ^{+}24] Marvin M. Mace, Michael J. Wilberg, Jerelle Jesse, Elizabeth North, Rasika Gawde, Malcolm E. Scully, and Lisa Wainger. Developing a fine-scale spatial operating model of eastern oyster population dynamics in Chesapeake Bay, Maryland, U.S.A. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002091.

Milburn:2023:DSC

[MWTH23] J. R. Milburn, S. M. Williams, K. A. Townsend, and B. J. Holmes. Depredation of spanner crabs (Ranina ranina) by endangered batoids off the east coast of Australia. Fisheries Research, 261(??):Article 106619, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000127.

Muko:2023:SEY

[MYKO23] Soyoka Muko, Mari Yoda, Hiroyuki Kurota, and Seiji Ohshimo. Spatial estimation and yearly trends in abundanceindex of Japanese jack mackerel (Trachurus japonicus) in the East China Sea and Sea of Japan. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623001741.

Meng:2024:UGA

[MZ24] Fanyi Meng and Guoping Zhu. Using a generative adversarial network-based model to simulate fishing behavior in Antarctic krill fishery. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001292.

Mora-Zamacona:2023:SMN

[MZSZVP^{+}23] Pablo Mora-Zamacona, César A. Salinas-Zavala, Raúl R. Villanueva-Poot, Enrique Morales-Bojórquez, Fernando I. González Laxe, and Kristin N. Marshall. Sustainable management of a new fishing stock: Bioeconomic approach under biological and market uncertainty. Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001583.

Mu:2021:ASO

[MZZ^{+}21] Xiuxia Mu, Chi Zhang, Chongliang Zhang, Jian Yang, and Yiping Ren. Age-structured otolith chemistry profiles revealing the migration of Conger myriaster in China Seas. Fisheries Research, 239(??):Article 105938, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000667.

Naz:2022:DGS

[NA22] Sumra Naz and Khalid Abbas. Delineating the genetic status of wild Cyprinus carpio as influenced by anthropogenic interventions. Fisheries Research, 251(??):Article 106300, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000777.

Natugonza:2020:STO

[NAS^{+}20] Vianny Natugonza, Cameron Ainsworth, Erla Sturludóttir, Laban Musinguzi, Richard Ogutu-Ohwayo, Tumi Tomasson, Chrisphine Nyamweya, and Gunnar Stefansson. Simulating trade-offs between socio-economic and conservation objectives for Lake Victoria (East Africa) using multispecies, multifleet ecosystem models. Fisheries Research, 229(??):Article 105593, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301107.

Nguyen:2023:MAR

[NAV^{+}23] Thao Van Nguyen, Andrea C. Alfaro, Leonie Venter, Jessica A. Ericson, Norman L. C. Ragg, Tom McCowan, and Craig Mundy. Metabolomics approach reveals sizespecific variations of blackfoot abalone (Haliotis iris) in Chatham Islands, New Zealand. Fisheries Research, 262(??):Article 106645, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623000383.

Nochete:2024:STD

[NB24] Charmane B. Nochete and Rex B. Baleña. The spatiotemporal distribution of small-scale fisheries along the northern Panay Gulf, Philippines: Implications for management. Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003284.

Nguyen:2020:ICS

[NBD^{+}20] Khanh Q. Nguyen, Shannon M. Bayse, Meghan Donovan, Paul D. Winger, Svein Løkkeborg, and Odd-Børre Humborstad. Increased catches of snow crab (Chionoecetes opilio) with luminescent-netting pots at long soak times. Fisheries Research, 230(??):Article 105685, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302022.

Nguyen:2023:QFE

[NBEI23] Vang Y. Nguyen, Shannon M. Bayse, Haraldur Arnar Einarsson, and Ólafur Arnar Ingólfsson. Quantifying fish escape under a bottom trawl. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001893.

Nottestad:2020:CAB

[NBF20] Leif Nøttestad, Erling Boge, and Keno Ferter. The comeback of Atlantic bluefin tuna (Thunnus thynnus) to Norwegian waters. Fisheries Research, 231(??):Article 105689, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362030206X.

Nunan:2020:WFC

[NC20] Fiona Nunan and Dražen Cepić. Women and fisheries comanagement: Limits to participation on Lake Victoria. Fisheries Research, 224(??):Article 105454, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303091.

Nakiyende:2023:RLF

[NCB^{+}23] Herbert Nakiyende, Lauren Chapman, Anthony Basooma, Dismas Mbabazi, Robinson Odong, Everest Nduwayesu, Samuel Bassa, Bairon Mugeni, Winnie Nkalubo, Alex Mulowoza, Richard Sande Mangeni, Anthony Taabu-Munyaho, and Jackson Efitre. A review of light fishing on Lake Albert, Uganda: Implications for a multi-species artisanal fishery. Fisheries Research, 258(??):Article 106535, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003125.

Nykanen:2020:GMR

[NDRR20] Milaja Nykänen, Eileen Dillane, David Reid, and Emer Rogan. Genetic methods reveal high diversity and no evidence of stock structure among cuckoo rays (Leucoraja naevus) in the northern part of Northeast Atlantic. Fisheries Research, 232(??):Article 105715, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302320.

Nielsen:2023:VPS

[NEBP^{+}23] Julius Nielsen, Daniel Estévez-Barcia, Søren Post, Helle Torp Christensen, Anja Retzel, Lorenz Meire, Frank Rigét, John Frederik Strøm, Audun Rikardsen, and Rasmus Hedeholm. Validation of pop-up satellite archival tags (PSATs) on Atlantic cod (Gadus morhua) in a Greenland fjord. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001753.

Noleto-Filho:2022:STE

[NFAL^{+}22] Eurico Mesquita Noleto-Filho, Ronaldo Angelini, Maria Alice Leite Lima, Sebastián Villasante, Mario J. F. Thomé-Souza, and Adriana Rosa Carvalho. Spatial and temporal effects improve Bayesian price estimation for the smallscale shrimp fishery in Sergipe State, Brazil. Fisheries Research, 247(??):Article 106189, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003179.

Nepal:2020:PPL

[NFC20] Vaskar Nepal, Mary C. Fabrizio, and William J. Connelly. Phenotypic plasticity in life-history characteristics of invasive blue catfish, Ictalurus furcatus. Fisheries Research, 230(??):Article 105650, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301673.

Nobrega:2025:LPF

[NFdSJO25] Marcelo Francisco Nóbrega, Luis Henrique França, Alexandre Ricardo dos Santos Junior, and Jorge Eduardo Lins Oliveira. Large pelagic fish exploitation by longliners in the Atlantic Ocean and Mediterranean Sea: a contribution to spatial planning and sustainable fisheries. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362400242X.

Nault:2025:DPP

[NGDC25] Andrew J. Nault, William B. Gaeuman, Benjamin J. Daly, and Aspen E. Coyle. Does the presence of a pop-up satellite archival tag affect movement of Tanner crab (Chionoecetes bairdi) in an exposed Alaskan bay? Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002923.

Niemi:2023:IRB

[NHE^{+}23] Niklas Niemi, Joakim P. Hansen, Johan S. Eklöf, Britas Klemens Eriksson, Henrik C. Andersson, Ulf Bergström, and Örjan Östman. Influence of reed beds (Phragmites australis) and submerged vegetation on pike (Esox lucius). Fisheries Research, 261(??):Article 106621, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000140.

Neog:2023:DEA

[NK23] Panchi Rani Neog and Bolin Kumar Konwar. The distribution, economic aspects, nutritional, and therapeutic potential of swamp eel Monopterus cuchia: a review. Fisheries Research, 261(??):Article 106635, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000280.

Ng:2025:QNT

[NLL^{+}25] Reana May Yen Ng, Adam Chee Ooi Lim, Chin Nurhiqwanalina Binti Henry Chin Siew Lee, Norazliana Binti Abdul Majib, Muhammad Ali Syed Hussein, and Amy Yee-Hui Then. Quantifying non-target seahorse fisheries and domestic traditional medicine-based trade in Malaysia. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002522.

Nataniel:2021:MSE

[NLS21] Anildo Nataniel, Jon Lopez, and Maria Soto. Modelling seasonal environmental preferences of tropical tuna purse seine fisheries in the Mozambique Channel. Fisheries Research, 243(??):Article 106073, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002010.

Naimullah:2022:EST

[NLW^{+}22] Muhamad Naimullah, Wei-Yu Lee, Yan-Lun Wu, Yi-Kai Chen, Yu-Cih Huang, Cheng-Hsin Liao, and Kuo-Wei Lan. Effect of soaking time on targets and bycatch species catch rates in fish and crab trap fishery in the southern East China Sea. Fisheries Research, 250(??):Article 106258, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000352.

Nelson:2024:TBE

[NMJ^{+}24] Lindsey Noel Nelson, W. David McElroy, Andrew W. Jones, George A. Maynard, Chris Tholke, and Anna J. M. Mercer. Taking the bait: Environmental factors affecting bait retention and hook disposition in the Gulf of Maine Bottom Longline Survey. Fisheries Research, 273(??):??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000213.

Nikolaus:2022:IPR

[NMS^{+}22] Robert Nikolaus, Sven Matern, Malwina Schafft, Andreas Maday, Christian Wolter, Thomas Klefoth, and Robert Arlinghaus. Influence of protected riparian areas on habitat structure and biodiversity in and at small lakes managed by recreational fisheries. Fisheries Research, 256 (??):Article 106476, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622002533.

Natugonza:2022:SVF

[NNS^{+}22] Vianny Natugonza, Chrispine Nyamweya, Erla Sturludóttir, Laban Musinguzi, Richard Ogutu-Ohwayo, Sam Bassa, Enock Mlaponi, Tumi Tomasson, and Gunnar Stefansson. Spatiotemporal variation in fishing patterns and fishing pressure in Lake Victoria (East Africa) in relation to balanced harvest. Fisheries Research, 252(??):Article 106355, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001321.

Nyamweya:2020:CDC

[NNTM^{+}20] Chrispine Sangara Nyamweya, Vianny Natugonza, Anthony Taabu-Munyaho, Christopher Mulanda Aura, James Murithi Njiru, Collins Ongore, Richard Mangeni-Sande, Benedicto Boniphace Kashindye, Cyprian Ogombe Odoli, Zachary Ogari, and Robert Kayanda. A century of drastic change: Human-induced changes of Lake Victoria fisheries and ecology. Fisheries Research, 230(??):Article 105564, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300813.

Nascimento:2023:SOB

[NOL23] Tânia Nascimento, Nuno Oliveira, and António Luís. Spatial overlap between the European Shag and commercial fisheries in a special protected area: implications for conservation. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000826.

Nilsson:2023:FSG

[NRH^{+}23] P. A. Nilsson, L. Ranåker, K. Hulthén, V. Nilsson-Örtman, C. Brönmark, and J. Brodersen. First-season growth and food of YOY pike (Esox lucius) are habitat specific within a lake. Fisheries Research, 259(??):Article 106563, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362200340X.

Nguyen:2023:AFA

[NS23] Quang Van Nguyen and Kok Fong See. Application of the frontier approach in capture fisheries efficiency and productivity studies: a bibliometric analysis. Fisheries Research, 263(??):??, July 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000693.

Neves:2021:PSE

[NSM^{+}21] João Neves, Alexandra Almeida Silva, Ana Moreno, Ana Veríssimo, António Múrias Santos, and Susana Garrido. Population structure of the European sardine Sardina pilchardus from Atlantic and Mediterranean waters based on otolith shape analysis. Fisheries Research, 243(??):Article 106050, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001788.

Nguyen:2022:PTE

[NSQV22] Thanh Viet Nguyen, Michel Simioni, Cao Le Quyen, and Hreidar Pór Valtýsson. Productivity, technical efficiency, and technological change in Vietnamese oceanic tuna fisheries. Fisheries Research, 248(??):Article 106202, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003301.

Nesslage:2022:ISS

[NSRM22] Geneviève Nesslage, Amy M. Schueller, Amanda R. Rezek, and Raymond M. Mroch. Influence of sample size and number of age classes on characterization of ageing error in pairedage comparisons. Fisheries Research, 249(??):Article 106255, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000327.

Nazia:2021:HGS

[NTJN21] Abdul Kadar Nazia, Bui Minh Tam, Jamsari Amirul Firdaus Jamaluddin, and Siti Azizah Mohd Nor. High genetic structure between natural populations of bighead catfish Clarias macrocephalus (Günther, 1864) from the Mekong Delta and Peninsular Malaysia. Fisheries Research, 241(??):Article 105993, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001211.

Nielsen:2023:IHS

[NVB^{+}23] J. Rasmus Nielsen, Berthe M. J. Vastenhoud, Sieme Bossier, Flemming Møhlenberg, Asbjørn Christensen, Rabea Diekman, Grete E. Dinesen, Ole R. Eigaard, Mayya Gogina, Michael L. Zettler, Alexander Darr, and Francois Bastardie. Impacts of habitat-specific benthic fishing compared to those of short-term induced variability by environmental drivers in a turbulent Baltic Sea environment. Fisheries Research, 257(??):Article 106514, January 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002910.

Nunez-Vallecillo:2021:EAR

[NVRG^{+}21] Mayra Nuñez-Vallecillo, Antonella Rivera, Konrad Górski, Antonio Brante, and Hugo A. Benítez. Ecomorphological analyses reveal impact of land-based stressors on stock structure of two commercially important fish species (Lutjanus synagris and Haemulon plumierii) in the Caribbean. Fisheries Research, 234(??):Article 105812, February 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303295.

Neves:2024:AAO

[NVSG24] João Neves, Ana Veríssimo, António Múrias Santos, and Susana Garrido. Age affects otolith shape in a coastal pelagic fish (Scomber colias Gmelin, 1789). Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623002746.

Nagano:2023:PCG

[NY23] Kosuke Nagano and Orio Yamamura. Predicting catch of Giant Pacific octopus Enteroctopus dofleini in the Tsugaru Strait using a machine learning approach. Fisheries Research, 261(??):Article 106622, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000152.

Neely:2021:EAR

[NZP^{+}21] Karen L. Neely, Tracy A. Ziegler, Margaret Peloso, Mark Hooper, Chesson O’Briant, Maria Wise, and Daniel Rittschof. Enhancing artificial reef fish populations by providing invertebrate prey refugia. Fisheries Research, 241(??):Article 106003, September 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001314.

Ozturk:2021:GPS

[ÖA21] Rafet Çağrı Öztürk and İlhan Altınok. Genetic population structure of the striped venus clam Chamelea gallina across its range. Fisheries Research, 234(??):Article 105758, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620302757.

Olsson:2023:PBA

[OAB^{+}23] Jens Olsson, Matilda L. Andersson, Ulf Bergström, Robert Arlinghaus, Asta Audzijonyte, Soren Berg, Laura Briekmane, Justas Dainys, Henrik Dalby Ravn, Jan Droll, Łukasz Dziemian, Dariusz P. Fey, Rob van Gemert, Martyna Greszkiewicz, Adam Grochowski, Egle Jakubavičiūtė, Linas Lozys, Adam M. Lejk, Noora Mustamäki, Rahmat Naddafi, Mikko Olin, Lauri Saks, Christian Skov, Szymon Smoliński, Roland Svirgsden, Joni Tiainen, and Örjan Östman. A pan-Baltic assessment of temporal trends in coastal pike populations. Fisheries Research, 260(??):Article 106594, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200371X.

Olsen:2021:DDF

[OAM^{+}21] Erik Olsen, Bjørn Erik Axelsen, Even Moland, Anne Christine Utne-Palm, Elamin Mohammed Elamin, Motassim Ali Mukhtar, Adel Mohamed Saleh, Sheikheldin Mohamed Elamin, Mohamed Abdelhameed Iragi, and Said Gumaa Fadul Gumaa. Distribution and diversity of fish species along the Sudanese Red Sea coast based on three combined trap and gillnet surveys. Fisheries Research, 242(??):Article 106032, October 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001600.

OFarrell:2021:SMH

[OB21] Halie B. O’Farrell and Elizabeth A. Babcock. Shortfin mako hot sets — defining high bycatch conditions as a basis for bycatch mitigation. Fisheries Research, 244(??):Article 106123, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002514.

Oesterwind:2020:FEN

[OBB^{+}20] Daniel Oesterwind, Bianca T. C. Bobowski, Anika Brunsch, Vladimir Laptikhovsky, Ralf van Hal, Anne F. Sell, and Graham J. Pierce. First evidence of a new spawning stock of Illex coindetii in the North Sea (NE-Atlantic). Fisheries Research, 221(??):Article 105384, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302395.

OLeary:2020:UOG

[OCBJG20] Ciara O’Leary, Robert Cruikshanks, Gustavo Becerra-Jurado, and Patrick Gargan. The use of otter guards in fyke net surveys and their effect on catches of European eel Anguilla anguilla. Fisheries Research, 228(??):Article 105570, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620300874.

Ortega-Cisneros:2024:LMS

[OCdMC24] Kelly Ortega-Cisneros, Carryn L. de Moor, and Kevern Cochrane. Linking the movement of South African sardine and anchovy to environmental variables using a model of intermediate complexity. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000651.

Ovenden:2020:PSG

[ODM20] Jennifer Ovenden, Danielle Davenport, and Anthony Moore. A perfect storm of genetic drift and divergence may prevent the rebuilding of the gemfish (Rexea solandri) stock on the east Australian coast. Fisheries Research, 230 (??):Article 105645, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301624.

ONeill:2020:MAP

[OFF^{+}20] F. G. O’Neill, R. J. Fryer, R. P. Frandsen, B. Herrmann, N. Madsen, and B. Mieske. A meta-analysis of plaice size-selection data in otter trawl codends. Fisheries Research, 227(??):Article 105558, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300758.

Olivencia:2024:MAA

[OGFW24] Kyle Olivencia, Emily E. Grausgruber, Mark J. Fincel, and Michael J. Weber. A multifaceted approach for assessing potential competition between smallmouth bass Micropterus dolomieu and walleye Sander vitreus in Lake Oahe, South Dakota. Fisheries Research, 276(??):??, August 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001243.

Okado:2020:CJC

[OKKW20] Jumpei Okado, Yosuke Koshino, Hideaki Kudo, and Yutaka Watanuki. Consumption of juvenile chum salmon by a seabird species during early sea life. Fisheries Research, 222(??):Article 105415, February 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578361930270X.

Ovelheiro:2023:MMM

[OMG^{+}23] Andreia Ovelheiro, João Monteiro, Patrícia Gonçalves, Marco António Campinho, Francisco Maia, Maria A. Teodósio, and Francisco Leitão. Macro and microscopic maturation stage key of green crab (Carcinus maenas, Linnaeus 1758): Reproductive cycle and differences among estuarine systems. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002217.

Okamura:2024:FFR

[OMK24] Hiroshi Okamura, Shoko Morita, and Hiroshi Kuroda. Forecasting fish recruitment using machine learning methods: a case study of arabesque greenling. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783624001607.

Outeiro:2021:QAT

[OOAF^{+}21] Luis Outeiro, Jaime Otero, Alexandre Alonso-Fernández, Rafael Bañón, and Juliano Palacios-Abrantes. Quantifying abundance trends and environmental effects on a population of queen scallop Aequipecten opercularis targeted by artisanal fishers in a coastal upwelling area (Ría de Arousa, NW Spain) using a Bayesian spatial model. Fisheries Research, 240(??):Article 105963, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000916.

Oliver:2023:ACS

[OOM^{+}23] Martin Oliver, Ross O’Neill, Matthew McHugh, Daragh Browne, Shane Murphy, Cóilín Minto, and Ronán Cosgrove. Assessment of cod survival in the Irish fly-shoot seine fishery using survivorship pop-up satellite archival tags. Fisheries Research, 266(??):??, October 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001765.

OBryhim:2021:FSI

[OPL21] Jason R. O’Bryhim, E. C. M. Parsons, and Stacey L. Lance. Forensic species identification of elasmobranchs landed in Costa Rican artisanal fisheries. Fisheries Research, 233(??):Article 105755, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302721.

Ortega:2024:ETE

[ORdlG^{+}24] Aurelio Ortega, Patricia Reglero, Fernando de la Gándara, Gabriel Mourente, and Edurne Blanco. Effects of temperature on embryonic development of Atlantic bluefin tuna (Thunnus thynnus, L. 1758) and Atlantic bonito (Sarda sarda, Bloch 1793). Fisheries Research, 277(??):??, September 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001309.

Orey:2025:IFB

[ÖRS^{+}25] Serra Örey, Jennifer Rehren, Torsten Schulze, Oscar Puebla, and Rabea Diekmann. Identifying fishing behavior groups from vessel movement data: Application to the German brown shrimp fleet. Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783625000220.

OFarrell:2021:CRT

[OS21a] Michael R. O’Farrell and William H. Satterthwaite. Corrigendum to “A rebuilding time model for Pacific salmon” [Fish. Res. 238 (June) 2021 105900]. Fisheries Research, 240(??):Article 105986, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001144. See [OS21b].

OFarrell:2021:RTM

[OS21b] Michael R. O’Farrell and William H. Satterthwaite. A rebuilding time model for Pacific salmon. Fisheries Research, 238(??):Article 105900, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016578362100028X. See corrigendum [OS21a].

ONeill:2022:IDV

[OSEF22] F. G. O’Neill, K. Summerbell, A. Edridge, and R. J. Fryer. Illumination and diel variation modify fish passage through an inclined grid. Fisheries Research, 250(??):Article 106297, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000741.

Ostman:2023:CBS

[ÖSL^{+}23] Örjan Östman, Göran Sundblad, Peter Ljungberg, Sandra Levin, Martina Blass, Marju Kaljuste, Iris Dahlin, Rebecka Svensson, and Jens Olsson. Catches, bycatches and stock indicators of fisheries targeting cyprinids along the Swedish Baltic Sea coast. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002229.

Ordonez:2022:EED

[OUB^{+}22] Alba Ordoñez, Ingrid Utseth, Olav Brautaset, Rolf Korneliussen, and Nils Olav Handegard. Evaluation of echosounder data preparation strategies for modern machine learning models. Fisheries Research, 254(??):Article 106411, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622001886.

Ondes:2020:CFA

[ÖÜÖG20] Fikret Öndes, Vahdet Ünal, Hakan Öndes, and Ana Gordoa. Charter fishing in the Aegean Sea (Turkey), Eastern Mediterranean: the missing point of fisheries management. Fisheries Research, 224(??):Article 105457, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303121.

Otteraa:2023:NPS

[OWF^{+}23] Håkon Otterå, Jens A. Wathne, Edvin Fuglebakk, Aril Slotte, Bjørn Vidar Svendsen, and Jon Helge Vølstad. A novel probabilistic survey method for at sea sampling in pelagic fisheries — the Norwegian catch sampling lottery. Fisheries Research, 260(??):Article 106584, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003617.

Osei:2021:IOS

[OYOO21] Isaac Kofi Osei, Kobina Yankson, Edward Adzesiwor Obodai, and Isaac Okyere. Implications of overlooked seasonal growth dynamics in tropical fisheries assessment: a test case of an oyster (Crassostrea tulipa) fishery in the densu delta, Ghana. Fisheries Research, 244(??):Article 106118, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002460.

Palmer:2024:BPS

[PAA^{+}24] Josie L. Palmer, Carina Armstrong, Hasan D. Akbora, Damla Beton, Çiğdem Çağlar, Brendan J. Godley, Kristian Metcalfe, Meryem Özkan, Robin T. E. Snape, and Annette C. Broderick. Behavioural patterns, spatial utilisation and landings composition of a small-scale fishery in the eastern Mediterranean. Fisheries Research, 269(??):??, January 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002540.

Palmer:2022:AOH

[PÁEMC22] Miquel Palmer, Amaya Álvarez-Ellacuría, Vicenç Moltó, and Ignacio A. Catalán. Automatic, operational, high-resolution monitoring of fish length and catch numbers from landings using deep learning. Fisheries Research, 246(??):Article 106166, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002940.

Pereira:2023:RBF

[PASdCF23] Diego V. Pereira, Caroline C. Arantes, Keid Nolan S. Sousa, and Carlos Edwar de C. Freitas. Relationships between fishery catch rates and land cover along a longitudinal gradient in floodplains of the Amazon River. Fisheries Research, 258(??):Article 106521, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002983.

Plaza:2023:EEV

[PAY23] Francisco Plaza, Héctor Araya, and Eleuterio Yáñez. Environmental effect on the variability of anchovy (Engraulis ringens) in northern Chile: Autoregressive conditional heteroskedastic approach with exogenonus variable and missing values. Fisheries Research, 260(??):Article 106607, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003848.

Papastamatiou:2020:USS

[PBB20] Yannis P. Papastamatiou, Chris Britton, and George H. Burgess. Using side-scan sonar to survey critically endangered smalltooth sawfish. Fisheries Research, 228 (??):Article 105577, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620300941.

Pottier:2022:OES

[PBB^{+}22] Gaétan Pottier, Nicolas Bargier, Yoann Bennevault, Régis Vigouroux, Didier Azam, Frédéric Marchand, Marie Nevoux, and Jean-Marc Roussel. Optimising electrofishing settings for shrimp and fish in shallow tropical streams. Fisheries Research, 256(??):Article 106457, December 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200234X.

Plaganyi:2025:KRF

[PBD25] Éva E. Plagányi, Laura K. Blamey, and Roy Aijun Deng. Key role of fisheries genomics to support conservation, fisheries management and trade of widely-distributed spiny lobsters. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000037.

Punt:2021:PMA

[PBDM21] André E. Punt, John R. Brandon, Douglas P. DeMaster, and Paula T. Moreno. Performance metrics for alternative management strategies for gray seal-commercial fishery interactions in the Northwest Atlantic. Fisheries Research, 243(??):Article 106060, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001880.

Plaganyi:2023:ARC

[PBDM23] Éva E. Plagányi, Laura K. Blamey, Roy Aijun Deng, and Margaret Miller. Accounting for risk-catch-cost trade-offs in a harvest strategy for a small, highly variable fishery. Fisheries Research, 258(??):Article 106518, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002958.

Petrere:2024:LII

[PBG24] Miguel Petrere and Davi Butturi-Gomes. Lack of independence: an insidious and subtle pitfall in model fitting. Fisheries Research, 270(??):??, February 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002941.

Premachandra:2023:GAI

[PBM^{+}23] HKA Premachandra, Alistair Becker, Kerry Millard, Danielle Johnston, Sankar Subramanian, Manoharan Kumar, Matthew Taylor, and Wayne Knibb. Genomic analyses indicate two blue swimmer crab species in Australia, evidence for natural interspecific hybridization and genetic structure within species with implications for fisheries management and stock enhancement. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001509.

Pacicco:2023:RBY

[PBPM^{+}23] Ashley E. Pacicco, Nancy J. Brown-Peterson, Debra J. Murie, Robert J. Allman, Derke Snodgrass, and James S.

Franks. Reproductive biology of yellowfin tuna (Thunnus albacares) in the northcentral U.S. Gulf of Mexico. Fisheries Research, 261(??):Article 106620, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000139.

Plaganyi:2022:PDU

[PBRT22] Éva E. Plagányi, Laura K. Blamey, Jacob G. D. Rogers, and Vivitskaia J. D. Tulloch. Playing the detective: Using multispecies approaches to estimate natural mortality rates. Fisheries Research, 249(??):Article 106229, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000066.

Perreault:2021:NMD

[PC21] Andrea M. J. Perreault and Noel G. Cadigan. Natural mortality diagnostics for state-space stock assessment models. Fisheries Research, 243(??):Article 106062, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001909.

Pember:2023:PGS

[PCBL23] Brenton M. Pember, Jennifer A. Chaplin, Matias Braccini, and Neil R. Loneragan. Population genomic and size distribution data suggest high genetic connectivity in the sandbar shark (Carcharhinus plumbeus) along a 2700 km coastline. Fisheries Research, 266(??):??, October 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001728.

Puente:2023:BSB

[PCC^{+}23] Esteban Puente, Leire Citores, Elsa Cuende, Iñigo Krug, and Mikel Basterretxea. Bycatch of short-beaked common dolphin (Delphinus delphis) in the pair bottom trawl fishery of the Bay of Biscay and its mitigation with an active acoustic deterrent device (pinger). Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002126.

Pena-Cutimbo:2024:MMI

[PCCMOA^{+}24] Nelly Peña-Cutimbo, Cristel Cordero-Maldonado, Clara Ortiz-Alvarez, Joanna Alfaro-Shigueto, and Jeffrey C. Mangel. Marine megafauna interactions with the Peruvian artisanal purse-seine fleet. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002710.

Pon:2023:SIB

[PCF23] Juan Pablo Seco Pon, Sofía Copello, and Marco Favero. Seabird interactions and bycatch in the Argentine freezer trawl fleet targeting Patagonian scallop (Zygochlamys patagonica). Fisheries Research, 262(??):Article 106661, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000541.

Purushothaman:2020:IGD

[PCGG20] Paramasivam Purushothaman, Rekha Devi Chakraborty, Maheswarudu Gidda, and Kuberan Ganesan. Investigation of genetic diversity and stock structure of Aristeus alcocki Ramadan, 1938 (Decapoda: Aristeidae) populations in the Indian coast with microsatellite markers. Fisheries Research, 227(??):Article 105550, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300679.

Punt:2021:CEN

[PCJH^{+}21] André E. Punt, Claudio Castillo-Jordán, Owen S. Hamel, Jason M. Cope, Mark N. Maunder, and James N. Ianelli. Consequences of error in natural mortality and its estimation in stock assessment models. Fisheries Research, 233(??):Article 105759, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620302769.

Pan:2023:BIC

[PCK23] Ruo-Yu Pan, Ke-Yang Chang, and Ting-Chun Kuo. Bycatch information complements the understanding of spatial distribution for commercially-important fish species.

Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001662.

Poirier:2021:SSS

[PCM^{+}21] Luke A. Poirier, Jeff C. Clements, Russell B. Millar, Rémi Sonier, and Monique Niles. Size selectivity of the scallop fishery in the southern Gulf of St. Lawrence: Effects of ring size and washer type. Fisheries Research, 243(??):Article 106103, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621002319.

Power:2025:UOD

[PD25] M. Power and J. B. Dempson. Using otolith δ^{18}O to assess habitat selection and growth in young-of-the-year Arctic charr. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003096.

Punt:2024:CUW

[PDA^{+}24] André E. Punt, Michael G. Dalton, Grant D. Adams, Steven J. Barbeaux, Wei Cheng, Albert J. Hermann, Kirstin K. Holsman, Peter-John F. Hulson, Thomas P. Hurst, and Alberto Rovellini. Capturing uncertainty when modelling environmental drivers of fish populations, with an illustrative application to Pacific cod in the eastern Bering Sea. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000158.

Pereira:2023:VOF

[PdAMdM^{+}23] Diego V. Pereira, Marcos de A. Mereles, Octávio F. de Matos, Giulia Cristina dos S. Lopes, Katrine G. da Conceição, and Carlos Edwar de C. Freitas. Vulnerability to overfishing of fish stocks in the Amazon Basin. Fisheries Research, 265(??):??, September 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001339.

Palermino:2023:PTS

[PDC^{+}23] Antonio Palermino, Andrea De Felice, Giovanni Canduci, Ilaria Biagiotti, Ilaria Costantini, Michele Centurelli, and Iole Leonori. Preliminary target strength measurement of Sprattus sprattus and its influence on biomass estimates in the Adriatic Sea (Mediterranean Sea). Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001704.

Poisson:2024:MPH

[PDC^{+}24] François Poisson, Hervé Demarcq, Sylvain Coudray, Jens Bohn, Juan Antonio Camiñas, Jean-Marc Groul, and David March. Movement pathways and habitat use of blue sharks (Prionace glauca) in the Western Mediterranean Sea: Distribution in relation to environmental factors, reproductive biology, and conservation issues. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362300293X.

Punt:2022:FAH

[PDD^{+}22] André E. Punt, Michael G. Dalton, Benjamin Daly, Tyler Jackson, W. Christopher Long, William T. Stockhausen, Cody Szuwalski, and Jie Zheng. A framework for assessing harvest strategy choice when considering multiple interacting fisheries and a changing environment: the example of eastern Bering Sea crab stocks. Fisheries Research, 252(??):Article 106338, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001151.

Punt:2020:EFN

[PDE^{+}20] André E. Punt, Alistair Dunn, Bjarki Tór Elvarsson, John Hampton, Simon D. Hoyle, Mark N. Maunder, Richard D. Methot, Jr., and Anders Nielsen. Essential features of the next-generation integrated fisheries stock assessment package: a perspective. Fisheries Research, 229(??):Article 105617, September 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362030134X.

Punt:2020:MYP

[PDF20] André E. Punt, Michael G. Dalton, and Robert J. Foy. Multispecies yield and profit when exploitation rates vary spatially including the impact on mortality of ocean acidification on North Pacific crab stocks. Fisheries Research, 225(??):Article 105481, May 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303364.

Perdichizzi:2022:CLH

[PDG^{+}22] Anna Perdichizzi, Claudio D’Iglio, Daniela Giordano, Adriana Profeta, Sergio Ragonese, and Paola Rinelli. Comparing life-history traits in two contiguous stocks of the deep-water rose shrimp Parapenaeus longirostris (H. Lucas, 1846) (Crustacea: Decapoda) in the Southern Tyrrhenian Sea (Central Mediterranean Sea). Fisheries Research, 248(??):Article 106206, April 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003349.

Priyashadi:2024:AEV

[PDJ24] M. S. V. H. Priyashadi, K. H. M. Ashoka Deepananda, and Asanka Jayasinghe. Assessing exploitation vulnerability risk of marine ornamental reef fish in Sri Lanka: a productivity susceptibility analysis. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002066.

Plaganyi:2024:LLT

[PDM^{+}24] Éva Plagányi, Leo Dutra, Nicole Murphy, Steven Edgar, Kinam Salee, Roy Aijun Deng, Laura K. Blamey, Denham Parker, and Stephanie Brodie. Lessons from long-term monitoring of tropical rock lobsters to support fisheries management. Fisheries Research, 275(??):??, July 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000948.

Pena:2021:FCC

[Peñ21] Marian Peña. Full customization of color maps for fisheries acoustics: Visualizing every target. Fisheries Research, 240(??):Article 105949, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000771.

Papacostas:2020:RAC

[PF20] Katherine J. Papacostas and John Foster. A replication approach to controlled selection for catch sampling intercept surveys. Fisheries Research, 229(??):Article 105609, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301260.

Palermino:2021:FTS

[PFC^{+}21] Antonio Palermino, Andrea De Felice, Giovanni Canduci, Ilaria Biagiotti, Ilaria Costantini, Sara Malavolti, and Iole Leonori. First target strength measurement of Trachurus mediterraneus and Scomber colias in the Mediterranean Sea. Fisheries Research, 240(??):Article 105973, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001016.

Pereira:2025:DCS

[PFdSBL25] Samantha Aquino Pereira, Nídia Noemi Fabré, Vandick da Silva Batista, and Sophie Lanco. Drivers of change in the spatial dynamics of the Central Amazon artisanal fishing fleet. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002844.

Perroca:2022:VPS

[PFFdC22] Júlia Fernandes Perroca, Jorge Luiz Rodrigues Filho, Adilson Fransozo, and Rogerio Caetano da Costa. Variations in pink-shrimps Farfantepenaeus brasiliensis and F. paulensis juvenile abundance: clarifying ecological patterns and providing subsidies to management in shallow marine ecosystems. Fisheries Research, 256(??):Article 106482, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002594.

Poirier:2020:FFT

[PFGQ20] Luke A. Poirier, Paula Tummon Flynn, Hannah Gehrels, and Pedro A. Quijón. Fukui foldable traps versus fyke nets as options for fishing European green crabs (Carcinus maenas) in soft-bottom habitats: Exploring efficiency and limiting bycatch. Fisheries Research, 230(??):Article 105637, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301545.

Perez:2022:AED

[PGAG22] Ilan Perez, Loreleï Guéry, Matthieu Authier, and Daniel Gaertner. Assessing the effectiveness of dFADs fishing moratorium in the Eastern Atlantic Ocean for conservation of juvenile tunas from AOTTP data. Fisheries Research, 253(??):Article 106360, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001370.

Palermino:2025:CES

[PGD^{+}25] Antonio Palermino, Sven Gastauer, Andrea De Felice, Giovanni Canduci, Ilaria Biagiotti, and Iole Leonori. Characterization of European sprat acoustic backscatter through modeling techniques: a comparison with in situ observations in the Mediterranean Sea. Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000359.

Pope:2021:SAU

[PGRD21] J. G. Pope, H. Gislason, J. C. Rice, and N. Daan. Scrabbling around for understanding of natural mortality. Fisheries Research, 240(??):Article 105952, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000801.

Powers:2023:NEG

[PHH^{+}23] Sean P. Powers, Crystal L. Hightower, John M. Hoenig, Jeffrey D. Plumlee, T. Reid Nelson, and J. Marcus Drymon. Novel estimation of gear selectivity using a concurrent mass mortality event: a case study using red drum (Sciaenops ocellatus) in the northern Gulf of Mexico. Fisheries Research, 263 (??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000838.

Pascoe:2020:IEE

[PHP^{+}20] Sean Pascoe, Trevor Hutton, Éva Plagányi, Roy Aijun Deng, Margaret Miller, Chris Moeseneder, and Stephen Eves. Influence of environment and economic drivers on fishing effort in Australia’s redleg banana prawn fishery. Fisheries Research, 227(??):Article 105555, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300722.

Petetta:2021:DSM

[PHV^{+}21] Andrea Petetta, Bent Herrmann, Massimo Virgili, Giada Bargione, Claudio Vasapollo, and Alessandro Lucchetti. Dredge selectivity in a Mediterranean striped venus clam (Chamelea gallina) fishery. Fisheries Research, 238(??):Article 105895, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000230.

Pennino:2022:IPB

[PIP^{+}22] Maria Grazia Pennino, Francisco Izquierdo, Iosu Paradinas, Marta Cousido, Francisco Velasco, and Santiago Cerviño. Identifying persistent biomass areas: the case study of the common sole in the northern Iberian waters. Fisheries Research, 248(??):Article 106196, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003246.

Privitera-Johnson:2022:TBP

[PJMP22] Kristin M. Privitera-Johnson, Richard D. Methot, and André E. Punt. Towards best practice for specifying selectivity in age-structured integrated stock assessments. Fisheries Research, 249(??):Article 106247, May 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000248.

Perez-Jimenez:2022:IEI

[PJNGJ^{+}22] Juan C. Pérez-Jiménez, Adrian Núñez, Mónica González-Jaramillo, Manuel Mendoza-Carranza, Jaime Acosta-Cetina, Alesa Flores-Guzmán, and Lorena Rocha-Tejeda. Inferring ecosystem impacts of a small-scale snapper fishery through citizen science data, productivity and susceptibility analysis, and ecosystem modelling. Fisheries Research, 250(??):Article 106269, June 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000467.

Pokki:2020:UAP

[PJOR20] H. Pokki, J. B. Jacobsen, S. B. Olsen, and A. Romakkaniemi. Understanding angler profiles in cases of heterogeneous count data — a travel cost model. Fisheries Research, 221(??):Article 105377, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302322.

Privitera-Johnson:2020:RAQ

[PJP20] Kristin M. Privitera-Johnson and André E. Punt. A review of approaches to quantifying uncertainty in fisheries stock assessments. Fisheries Research, 226(??):Article 105503, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300205.

Passerotti:2020:FTN

[PJSQ20] Michelle S. Passerotti, Christian M. Jones, Christopher E. Swanson, and Joseph M. Quattro. Fourier-transform near infrared spectroscopy (FT-NIRS) rapidly and non-destructively predicts daily age and growth in otoliths of juvenile red snapper Lutjanus campechanus (Poey, 1860). Fisheries Research, 223(??):Article 105439, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302942.

Petza:2024:SIR

[PK24] Dimitra Petza and Stelios Katsanevakis. Science-informed recommendations to enhance the effectiveness of areabased fisheries management for fisheries sustainability and marine conservation: a global mini-review. Fisheries Research, 272(??):??, April 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000110.

Peiris:2021:SCC

[PKRL21] M. Aravinda Kishan Peiris, Terney Pradeep Kumara, R. R. M. K. P. Ranatunga, and Shang-Yin Vanson Liu. Species composition and conservation status of shark from fishery landings and fish markets in Sri Lanka revealed by DNA barcoding. Fisheries Research, 242(??):Article 106045, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001739.

Papageorgiou:2023:SSF

[PM23] Marios Papageorgiou and Dimitrios K. Moutopoulos. Smallscale fisheries discards in the eastern Mediterranean Sea: Discarding species, quantities, practices and drivers. Fisheries Research, 267(??):??, November 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001911.

Poole:2024:SLT

[PMC^{+}24] Russell Poole, Cóilín Minto, Joseph Cooney, Alan Drumm, Pat Hughes, Michael Murphy, Pat Nixon, Ger Rogan, David Sweeney, and Ciara O’Leary. Standardising long-term eel (Anguilla anguilla) fyke net survey data reveals covariate effects and improves estimates of declining relative abundance. Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400002X.

Plaganyi:2020:DDP

[PMS^{+}20] Éva E. Plagányi, Nicole Murphy, Timothy Skewes, Leo X. C. Dutra, Natalie Dowling, and Mibu Fischer. Development of a data-poor harvest strategy for a sea cucumber fishery. Fisheries Research, 230(??):Article 105635, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301521.

Parra:2023:ICS

[PMS^{+}23] Hugo Parra, Miguel Machete, Marco Santos, Karen A. Bjorndal, and Frederic Vandeperre. Incidental capture of sea turtles in the Northeast Atlantic Portuguese pelagic longline fishery. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000668.

Podolska:2024:PNL

[PNAPH24] Magdalena Podolska, Katarzyna Nadolna-Ałtyn, Joanna Pawlak, and Jan Horbowy. The presence of nematodes in the liver of Baltic cod, Gadus morhua, is associated with a decline in condition factors and hepatosomatic index of the host. Fisheries Research, 273(??):??, May 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000225.

Precioso:2022:TAT

[PNGGO^{+}22] Daniel Precioso, Manuel Navarro-García, Kathryn Gavira-O’Neill, Alberto Torres-Barrán, David Gordo, Víctor Gallego, and David Gómez-Ullate. TUN-AI: Tuna biomass estimation with machine learning models trained on oceanography and echosounder FAD data. Fisheries Research, 250(??):Article 106263, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622000406.

Puttasan:2023:IMC

[PNRS23] Mayuree Puttasan, Mingkwan Nipitwattanaphon, Wikrom Rungsin, and Akarapong Swatdipong. Instinctive mate choice with genetically different partners assists sustaining genetic diversity in giant freshwater prawn (Macrobrachium rosenbergii). Fisheries Research, 262(??):Article 106648, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000413.

Politikos:2021:AFA

[PPC^{+}21] Dimitris V. Politikos, Georgios Petasis, Archontia Chatzispyrou, Chryssi Mytilineou, and Aikaterini Anastasopoulou. Automating fish age estimation combining otolith images and deep learning: the role of multitask learning. Fisheries Research, 242(??):Article 106033, October 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001612.

Pascoe:2023:DLM

[PPC23a] Sean Pascoe, Samantha Paredes, and Louisa Coglan. Do “local” markets offer new opportunities to Australian seafood producers? Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300084X.

Philipp:2023:URI

[PPC^{+}23b] Madison A. Philipp, Kathryn S. Peiman, Glenn T. Crossin, Scott G. Hinch, David A. Patterson, Chris K. Elvidge, and Steven J. Cooke. Using repeat injury assessments in adult sockeye salmon (Oncorhynchus nerka) to predict spawning success and describe severity of migration conditions. Fisheries Research, 266(??):??, October 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300190X.

Parker:2025:ACS

[PPD^{+}25] D. Parker, A. E. Punt, R. A. Deng, W. N. Venables, C. M. Dichmont, A. G. Donovan, R. A. Kenyon, T. van der Velde, M. Miller, and T. Hutton. Allocating catches to species in mixed species data – a model-based approach. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003047.

Poussard:2021:EED

[PPH21] Leanne M. Poussard, Eric N. Powell, and Daniel R. Hennen. Efficiency estimates from depletion experiments for sedentary invertebrates: evaluation of sources of uncertainty in experimental design. Fisheries Research, 234(??):Article 105806, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303234.

Parra:2023:PIP

[PPM^{+}23] Hugo Parra, Christopher K. Pham, Miguel Machete, Marco Santos, Karen A. Bjorndal, and Frederic Vandeperre. The Portuguese industrial pelagic longline fishery in the Northeast Atlantic: Catch composition, spatio-temporal dynamics of fishing effort, and target species catch rates. Fisheries Research, 264(??):??, August 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001236.

Paulina:2023:NDS

[PRA^{+}23] Mejía-Ruíz Paulina, Perez-Enriquez Ricardo, Severin Andrew, Martínez-Rincón Raúl Octavio, Garza John Carlos, Mares-Mayagoitia Jorge Alberto, and Valenzuela-Quiñonez Fausto. A novel diagnostic SNP panel for fishery management of green abalone (Haliotis fulgens). Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783623001790.

Prants:2024:FLF

[Pra24] S. V. Prants. Fisheries at Lagrangian fronts. Fisheries Research, 279(??):??, November 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001899.

Plumlee:2022:ACE

[PRCF22] Jeffrey D. Plumlee, Grace Roskar, J. Kevin Craig, and F. Joel Fodrie. Assessing the catch efficiency of predators in the presence of prey using experimental gillnets in a temperate estuary. Fisheries Research, 253(??):Article 106383, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001606.

Pita:2021:FCO

[PRF^{+}21] Cristina Pita, Katina Roumbedakis, Teresa Fonseca, Fábio L. Matos, João Pereira, Sebastián Villasante, Pablo Pita, José Maria Bellido, Angel F. Gonzalez, Manuel García-Tasende, Evgenia Lefkaditou, Aggeliki Adamidou, Danila Cuccu, Paola Belcari, Ana Moreno, and Graham J. Pierce. Fisheries for common octopus in Europe: socioeconomic importance and management. Fisheries Research, 235(??):Article 105820, March 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303374.

Planas:2023:IBR

[PRK23] Josep V. Planas, Christopher N. Rooper, and Gordon H. Kruse. Integrating biological research, fisheries science and management of Pacific halibut (Hippoglossus stenolepis) across the North Pacific Ocean. Fisheries Research, 259(??):Article 106559, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622003368.

Poomchaivej:2024:STR

[PRN^{+}24] Thanasak Poomchaivej, Wayne Robinson, Nathan Ning, Lee J. Baumgartner, and Xiaodi Huang. Suitability of tropical river fishes for PIT tagging: Results for four Lower Mekong species. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003235.

Potts:2023:VAG

[PRRR23] Jennifer C. Potts, Walter D. Rogers, Troy C. Rezek, and Amanda R. Rezek. Validation of annual growth zone formation in gray triggerfish Balistes capriscus dorsal spines, vertebrae, and otoliths. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002023.

Piper:2020:UTT

[PRWK20] Adam T. Piper, Paula J. Rosewarne, Rosalind M. Wright, and Paul S. Kemp. Using ‘trap and transport’ to facilitate seaward migration of landlocked European eel (Anguilla anguilla) from lakes and reservoirs. Fisheries Research, 228(??):Article 105567, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300849.

Paul:2020:UVS

[PSS^{+}20] Thankam Theresa Paul, Shyam S. Salim, Manoharan S., U. K. Sarkar, and B. K. Das. Understanding variations in socio-economic vulnerabilities and the strategies adopted by small scale fishing communities of tropical reservoirs. Fisheries Research, 226(??):Article 105523, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300400.

Punt:2021:APB

[PSS^{+}21] André E. Punt, Maritza Sepúlveda, Margaret C. Siple, Jeffrey Moore, Tessa B. Francis, Philip S. Hammond, Dennis Heinemann, Kristy J. Long, Doris Oliva, Randall R. Reeves, Gudjón Már Sigurdsson, Gísli Víkingsson, Paul R. Wade, Rob Williams, and Alexandre N. Zerbini. Assessing pinniped bycatch mortality with uncertainty in abundance and post-release mortality: a case study from Chile. Fisheries Research, 235(??):Article 105816, March 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303337.

Preul-Stimetz:2024:EPI

[PSSFS24] Taylor N. Preul-Stimetz, Stephanie L. Shaw, Zachary S. Feiner, and Greg G. Sass. Evaluating the potential importance of individual identity, maternal traits, and environment as predictors of egg characteristics in walleye Sander vitreus. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001711.

Punt:2020:WMB

[PTD^{+}20] André E. Punt, Geoffrey N. Tuck, Jemery Day, Cristian M. Canales, Jason M. Cope, Carryn L. de Moor, José A. A. De Oliveira, Mark Dickey-Collas, Bjarki T. Elvarsson, Melissa A. Haltuch, Owen S. Hamel, Allan C. Hicks, Christopher M. Legault, Patrick D. Lynch, and Michael J. Wilberg. When are model-based stock assessments rejected for use in management and what happens then? Fisheries Research, 224(??):Article 105465, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303200.

Punt:2021:IAA

[PTD^{+}21] André E. Punt, Geoffrey N. Tuck, Jemery Day, Paul Burch, Robin B. Thomson, and Pia Bessell-Browne. The impact of alternative age–length sampling schemes on the performance of stock assessment methods. Fisheries Research, 238(??):Article 105904, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000321.

Peristeraki:2020:EBT

[PTK^{+}20] Panagiota Peristeraki, George Tserpes, Stefanos Kavadas, Argyris Kallianiotis, and Kostantinos I. Stergiou. The effect of bottom trawl fishery on biomass variations of demersal chondrichthyes in the eastern Mediterranean. Fisheries Research, 221(??):Article 105367, January 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361930222X.

Punt:2024:ICK

[PTL^{+}24] André E. Punt, Robin Thomson, L. Richard Little, Pia Bessell-Browne, Paul Burch, and Mark Bravington. Including close-kin mark-recapture data in statistical catchat-age stock assessments and management strategies. Fisheries Research, 276(??):??, August 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001218.

Peharda:2023:REA

[PUC^{+}23] Melita Peharda, Hana Uvanović, Maxi Castrillejo, Daria Ezgeta-Balić, and Nedo Vrgoč. (Re-)Examining age and growth increments of the clam Venus verrucosa. Fisheries Research, 267(??):??, November 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002059.

Punt:2023:WFL

[Pun23] André E. Punt. Those who fail to learn from history are condemned to repeat it: a perspective on current stock assessment good practices and the consequences of not following them. Fisheries Research, 261(??):Article 106642, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000358.

Punt:2024:SAR

[Pun24] André E. Punt. Stock assessment of rock lobster stocks: Past, present and future. Fisheries Research, 274(??):??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000602.

Paille:2024:ISN

[PVA^{+}24] Jade Paillé, Corentin Vignard, Matthieu Authier, Emeric Bidenbach, Camille Deslias, Stéphanie Tachoires, and Hélène Peltier. Identification of static netters fishing trajectories with high resolution data and their evolution in the Bay of Biscay since 2015: Potential implications for short-beaked common dolphin bycatch. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001838.

Paul:2022:DBE

[PVC^{+}22] Thankam Theresa Paul, G. D. Vandana, Albin Albert C., Arun Pandit, B. K. Das, and Shyam S. Salim. Dilemma on the ban on exotics in Indian subcontinent: a systematic review in a socio-economic perspective. Fisheries Research, 246(??):Article 106150, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002782.

Peatman:2022:TCN

[PVPN22] Tom Peatman, Matthew T. Vincent, Joe Scutt Phillips, and Simon Nicol. Times are changing, but has natural mortality? Estimation of mortality rates for tropical tunas in the western and central Pacific Ocean. Fisheries Research, 256(??):Article 106463, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002405.

Peter:2022:BHS

[PvZ22] Happy K. Peter and Paul A. M. van Zwieten. Bethedging strategies determine daily choices in effort allocation for Nile perch fishers of Lake Victoria. Fisheries Research, 253(??):Article 106363, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001400.

Page:2025:LTT

[PW25] Kevin S. Page and Curtis P. Wagner. Long-term tagging reveals variable escapement of stocked muskellunge among four Ohio reservoirs: Consequences for managing reservoir fisheries. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002947.

Pan:2020:PCH

[PYX^{+}20] Xindong Pan, Zhenjiang Ye, Binduo Xu, Tao Jiang, Jian Yang, and Yongjun Tian. Population connectivity in a highly migratory fish, Japanese Spanish mackerel (Scomberomorus niphonius), along the Chinese coast, implications from otolith chemistry. Fisheries Research, 231(??):Article 105690, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302071.

Pecoraro:2020:WSM

[PZG^{+}20] C. Pecoraro, I. Zudaire, G. Galimberti, M. Romeo, H. Murua, C. Fruciano, C. Scherer, F. Tinti, N. C. Diaha, N. Bodin, and E. Chassot. When size matters: the gonads of larger female yellowfin tuna (Thunnus albacares) have different fatty acid profiles compared to smaller individuals. Fisheries Research, 232(??):Article 105726, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302435.

Psuty:2023:GNP

[PZG23] Iwona Psuty, Radosław Zaporowski, and Władysław Gaweł. Goodbye to northern pike (Esox lucius) in the Polish southern Baltic? Fisheries Research, 258(??):Article 106549, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003265.

Philipp:2023:CRR

[PZL^{+}23] David P. Philipp, Aaron Zolderdo, Michael J. Lawrence, Julie E. Claussen, Liane Nowell, Peter Holder, and Steven J. Cooke. COVID-19 reduced recreational fishing effort during the black bass spawning season, resulting in increases in black bass reproductive success and annual recruitment. Fisheries Research, 259(??):Article 106580, March 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003575.

Queirolo:2025:TIE

[QAS^{+}25] Dante Queirolo, Mauricio Ahumada, Marcelo A. San Martín, Catalina Román, Tomás Araya-Schmidt, Pedro Apablaza, and Claudio Bernal. Testing and implementation of an exclusion device for mitigating the incidental capture of sea lions in the Chilean small-scale trawl hake fishery. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002698.

QuintelaDosSantos:2022:IWG

[QMC^{+}22] Jessica Borba Quintela Dos Santos, Cristiano Mazur Chiessi, Stefano Crivellari, Jonas Eloi de Vasconcelos Filho, Natan Silva Pereira, Matheus Oliveira Freitas, and Beatrice Padovani Ferreira. Identification of western South Atlantic stocks of the Lane snapper (Lutjanus synagris) from an otolith based multi-proxy approach. Fisheries Research, 253(??):Article 106356, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001333.

Quispe-Machaca:2022:TJS

[QMGRIU22] Marco Quispe-Machaca, Fabián A. Guzmán-Rivas, Christian M. Ibáñez, and Ángel Urzúa. Trophodynamics of the jumbo squid Dosidicus gigas during winter in the Southeast Pacific Ocean off the coast of Chile: Diet analyses and fatty acid profile. Fisheries Research, 245(??):Article 106154, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002824.

Quijano:2023:IFO

[QSMG^{+}23] Daniel Quijano, Silvia Salas, Carmen Monroy-García, Michel Dreyfus-León, and Edgar Torrres-Irineo. Identifying fisheries operations in tropical multispecies fisheries: a comparative analysis of multivariate approaches and neural networks. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000851.

Rees:2021:ESE

[RAE^{+}21] Siân E. Rees, Matthew Ashley, Louisa Evans, Stephen Mangi, Emma V. Sheehan, Tom Mullier, Adam Rees, and Martin J. Attrill. An evaluation of the social and economic impact of a marine protected area on commercial fisheries. Fisheries Research, 235(??):Article 105819, March 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303362.

Ramon:2022:RBH

[RAG22] M. Ramón, M. J. Amor, and E. Galimany. Reproductive biology of the holothurian Parastichopus regalis in the Mediterranean Sea and its implications for fisheries management. Fisheries Research, 247(??):Article 106191, March 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003192.

Rose:2022:DTN

[RB22] Craig S. Rose and David Barbee. Developing and testing a novel active-selection (ActSel) bycatch reduction device to quickly alternate trawls between capture and release configurations with real-time triggering. Fisheries Research, 254(??):Article 106380, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001576.

Regular:2022:ISM

[RBD^{+}22] Paul M. Regular, Alejandro D. Buren, Karen S. Dwyer, Noel G. Cadigan, Robert S. Gregory, Mariano Koen-Alonso, Rick M. Rideout, Gregory J. Robertson, Matthew D. Robertson, Garry B. Stenson, Laura J. Wheeland, and Fan Zhang. Indexing starvation mortality to assess its role in the population regulation of Northern cod. Fisheries Research, 247(??):Article 106180, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003088.

Rouyer:2020:TAB

[RBG^{+}20] Tristan Rouyer, Sylvain Bonhommeau, Nicolas Giordano, François Giordano, Saviour Ellul, Giovanni Ellul, Simeon Deguara, Bertrand Wendling, Serge Bernard, and Vincent Kerzerho. Tagging Atlantic bluefin tuna from a Mediterranean spawning ground using a purse seiner. Fisheries Research, 226(??):Article 105522, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300394.

Rynvis:2024:RFA

[RBG^{+}24] Liv Rynvis, Sam Bock, Matt Gillett, Shane Walters, David V. Fairclough, and Matt Navarro. Recreational fishers’ awareness and adoption of strategies to minimise post-release mortality in demersal fish. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001206.

Robinson:2024:PTS

[RBH^{+}24] Wayne Robinson, Lee J. Baumgartner, Khampheng Homsombath, Nathan Ning, Khamla Phommachanh, Thonglom Phommavong, Thanasak Poomchaivej, Karl Pomorin, Dulce Simmanivong, Douangkham Singhanouvong, and Phousone Vorasane. PIT tagging systems are suitable for assessing cumulative impacts of Mekong River hydropower plants on (upstream) fish migrations in Lao PDR. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000596.

Renchen:2024:EGW

[RBHM24a] Gabrielle F. Renchen, Casey B. Butler, Emily Hutchinson, and Thomas R. Matthews. Escape gaps in wire lobster traps reduce bycatch of coral reef fish while maintaining catch of harvestable lobsters in Florida’s Caribbean spiny lobster (Panulirus argus), fishery. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002977.

Ross:2024:RIP

[RBHM24b] Erica P. Ross, Casey B. Butler, Carly M. Hannah, and Thomas R. Matthews. Reflex impairment as a predictor of mortality in Panulirus argus: Implications for the live export market. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001814.

Rijnsdorp:2021:EES

[RBM21] A. D. Rijnsdorp, J. Batsleer, and P. Molenaar. The effect of electrical stimulation on the footrope and codend selection of a flatfish bottom trawl. Fisheries Research, 243(??):Article 106104, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002320.

Reid:2024:EIT

[RBS^{+}24] Connor H. Reid, Albana I. Berberi, Kara M. Scott, Sam J. Woods, Jonathan D. Midwood, and Steven J. Cooke. Evaluating immobilisation thresholds and suitability of conductive glove electrodes for largemouth bass electroanaesthesia. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623003247.

Romero:2021:TNA

[RCH^{+}21] Joana Romero, Paulo Catry, Margarida Hermida, Verónica Neves, Bárbara Cavaleiro, Lídia Gouveia, and José Pedro Granadeiro. Tunas off northwest Africa: the epipelagic diet of the Bigeye and Skipjack tunas. Fisheries Research, 238(??):Article 105914, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000424.

Rodriguez-Castaneda:2022:STA

[RCVGMI22] José Carlos Rodríguez-Castañeda, Ana Ventero, Manuela Gertrudis García-Márquez, and Magdalena Iglesias. Spatial and temporal analysis (2009–2020) of the biological parameters, abundance and distribution of Trachurus mediterraneus (Steindachner, 1868) in the Western Mediterranean. Fisheries Research, 256(??):Article 106483, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002600.

Rangely:2023:AIV

[RdBAT^{+}23] Jordana Rangely, Matheus S. F. de Barros, Mônica D. Albuquerque-Tenório, Reginaldo Medeiros, Richard James Ladle, and Nidia Noemi Fabré. Assessing interspecific variation in life-history traits of three sympatric tropical mullets using age, growth and otolith allometry. Fisheries Research, 260(??):Article 106577, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200354X.

Ribas:2021:UDD

[RdLSdB^{+}21] Talita Fernanda Augusto Ribas, João Bráullio de Luna Sales, Hugo de Boer, Jarl Andreas Anmarkrud, Renato Renison Moreira Oliveira, Marcele Laux, Fabricio dos Anjos Santa Rosa, Guilherme Corrêa Oliveira, Felippe A. Postuma, Maria A. Gasalla, and Jonathan Stuart Ready. Unexpected diversity in the diet of Doryteuthis sanpaulensis (Brakoniecki, 1984) (Mollusca: Cephalopoda) from the southern Brazilian sardine fishery identified by metabarcoding. Fisheries Research, 239(??):Article 105936, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000643.

Roser:2023:SHC

[RDR^{+}23] Phillip Roser, Félicie Dhellemmes, Timo Rittweg, Sören Möller, Helmut Winkler, Olga Lukyanova, Dominique Niessner, Jörg Schütt, Carsten Kühn, Stefan Dennenmoser, Arne W. Nolte, Johannes Radinger, Dieter Koemle, and Robert Arlinghaus. Synthesizing historic and current evidence for anadromy in a northern pike (Esox lucius L.) meta-population inhabiting brackish lagoons of the southern Baltic Sea, with implications for management. Fisheries Research, 263(??):??, July 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000632.

Riquet:2022:HPS

[RFF^{+}22] Florentine Riquet, Cécile Fauvelot, Pauline Fey, Daphné Grulois, and Marc Leopold. Hatchery-produced sandfish (Holothuria scabra) show altered genetic diversity in New Caledonia. Fisheries Research, 252(??):Article 106343, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001205.

Reis-Filho:2021:WBL

[RFMS^{+}21] José Amorim Reis-Filho, Ricardo J. Miranda, Claudio L. S. Sampaio, José Anchieta C. C. Nunes, and Antoine O. H. C. Leduc. Web-based and logbook catch data of permits and pompanos by small-scale and recreational fishers: Predictable spawning aggregation and exploitation pressure. Fisheries Research, 243(??):Article 106064, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001922.

Rautureau:2022:STD

[RGG22] Clément Rautureau, Chloé Goulon, and Jean Guillard. In situ TS detections using two generations of echo-sounder, EK60 and EK80: the continuity of fishery acoustic data in lakes. Fisheries Research, 249(??):Article 106237, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000145.

Rousseau:2020:MFA

[RGN^{+}20] Shani Rousseau, Stéphane Gauthier, Chrys Neville, Stewart Johnson, and Marc Trudel. A multi-frequency acoustic method to estimate mean standard length of juvenile salmon in the Discovery Islands, British Columbia. Fisheries Research, 227(??):Article 105536, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300539.

Ryder:2023:DFS

[RGP^{+}23] Finn J. Ryder, Gaya Gnanalingam, Daniel W. Pritchard, William J. Rayment, Nigel J. Scott, and Christopher D. Hepburn. Drivers of fishery status for the cultural keystone pāua (Haliotis iris) in Customary fishery Protection Areas in Aotearoa New Zealand. Fisheries Research, 261(??):Article 106613, May 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623000061.

Ragavan:2024:EJM

[RKD24] N. Ragavan, S. Kuganathan, and D. C. T. Dissanayake. Effects of juvenile mortality on population dynamics of Penaeus semisulcatus: a case study from Sri Lanka. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000559.

Rahbar:2023:PSC

[RKN23] Mahdieh Rahbar, Majidreza Khoshkholgh, and Sajad Nazari. Population structure of Caspian Kutum (Rutilus frisii, Nordmann, 1840) in the southern coast of Caspian Sea using genome-wide single nucleotide polymorphism markers. Fisheries Research, 257(??):Article 106499, January 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002764.

Ryznar:2024:PDR

[RL24] Emily R. Ryznar and Michael A. Litzow. Predicting the distribution of red king crab bycatch in Bering Sea flatfish trawl fisheries. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002224.

Ruiz:2021:SPS

[RLO^{+}21] Jon Ruiz, Maite Louzao, Iñaki Oyarzabal, Luis Arregi, Estanis Mugerza, and Andres Uriarte. The Spanish purse-seine fishery targeting small pelagic species in the Bay of Biscay:

Landings, discards and interactions with protected species. Fisheries Research, 239(??):Article 105951, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000795.

Ran:2020:DBI

[RLQ^{+}20] Ke Ran, Qi Li, Lu Qi, Weidong Li, and Lingfeng Kong. DNA barcoding for identification of marine gastropod species from Hainan Island, China. Fisheries Research, 225(??):Article 105504, May 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300217.

Rufino:2025:DCM

[RMD^{+}25] Marta M. Rufino, Catarina Maia, David Dinis, Inês Farias, Teresa Moura, Pedro Gomes, Ivania Quaresma, Rogélia Martins, Ivone Figueiredo, Bárbara Serra-Pereira, Miguel B. Gaspar, and Ana Moreno. Designing a coastal monitoring marine biodiversity survey, using trammel nets and gillnets in Portugal. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002662.

Ramirez-Monsalve:2021:PMP

[RMNB^{+}21] P. Ramírez-Monsalve, K. N. Nielsen, M. Ballesteros, T. S. Kirkfeldt, M. Dickey-Collas, A. Delaney, T. J. Hegland, J. Raakjær, and P. Degnbol. Pulling mechanisms and pushing strategies: How to improve ecosystem approach fisheries management advice within the European Union’s common fisheries policy. Fisheries Research, 233(??):Article 105751, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362030268X.

Randall:2022:FFP

[RMRG22] Jessica R. Randall, Hannah M. Murphy, Dominique Robert, and Maxime Geoffroy. Forage fish as a predator: summer and autumn diet of Atlantic herring in Trinity Bay, Newfoundland. Fisheries Research, 252(??):Article 106331, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622001084.

Rufener:2023:CCE

[RNKB23] Marie-Christine Rufener, J. Rasmus Nielsen, Kasper Kristensen, and Francois Bastardie. Closing certain essential fish habitats to fishing could be a win–win for fish stocks and their fisheries — insights from the western Baltic cod fishery. Fisheries Research, 268(??):??, December 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002461.

Rohan:2024:CBR

[RNP^{+}24] Sean K. Rohan, Julie K. Nielsen, Bianca K. Prohaska, Alex De Robertis, Steve G. Lewis, and Susanne F. McDermott. Characterizing behavioral responses of Pacific cod to bottom trawl vessels and gear using archival tag accelerometer data. Fisheries Research, 275(??):??, July 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000973.

Richerson:2020:NHC

[RPH20] Kate Richerson, André E. Punt, and Daniel S. Holland. Nearly a half century of high but sustainable exploitation in the Dungeness crab (Cancer magister) fishery. Fisheries Research, 226(??):Article 105528, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362030045X.

Robichaud:2024:IEF

[RPL^{+}24] Jessica A. Robichaud, Morgan L. Piczak, Luc LaRochelle, Jessica L. Reid, Auston D. Chhor, Peter E. Holder, Liane B. Nowell, Jacob W. Brownscombe, Andrew J. Danylchuk, and Steven J. Cooke. Interactive effects of fish handling and water temperature on reflex impairment of angled rainbow trout. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000572.

Rudershausen:2021:EDM

[RPM^{+}21] P. J. Rudershausen, S. J. Poland, J. H. Merrell, C. B. Pelletier, C. S. Mikles, and J. A. Buckel. Estimating discard mortality using meta-analysis and fishery-dependent sampling. Fisheries Research, 240(??):Article 105962, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000904.

Rzeszowski:2024:ISE

[RRB24] Everett J. Rzeszowski, Kathleen M. Reardon, and Damian C. Brady. Integrating in situ environmental covariates in an American lobster catch model to improve impact assessment. Fisheries Research, 280(??):??, December 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002273.

Rzeszowski:2024:BCD

[RRH^{+}24] Everett J. Rzeszowski, Kathleen M. Reardon, Heidi Henninger, Joshua T. Carloni, and Damian C. Brady. Building confidence: Developing image-based methods to incorporate fishery-collected data in the American lobster stock assessment. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400119X.

Ramirez-Rodriguez:2024:FWH

[RRSP^{+}24] Saray Ramírez-Rodríguez, Amèlia Sarroca, Santiago Pelosso, David Fernández-Guerrero, and Lourdes Reig Puig. Food waste in high income countries: Spanish fish value chain as a case study. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400198X.

Roth:2021:PMC

[RS21] Bjørn Roth and Torstein Skåra. Pre mortem capturing stress of Atlantic herring (Clupea harengus) in purse seine and subsequent effect on welfare and flesh quality. Fisheries Research, 244(??):Article 106124, December 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002526.

Rodriguez-Salvador:2023:DFP

[RSD23] Beatriz Rodriguez-Salvador and Domingo Calvo Dopico. Differentiating fish products: Consumers’ preferences for origin and traceability. Fisheries Research, 262(??):Article 106682, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000759.

Ruokonen:2022:RMW

[RSPE22] Timo J. Ruokonen, Petri Suuronen, Henni Pulkkinen, and Jaakko Erkinaro. Release mortality of wild Atlantic salmon in coastal pontoon-trap fishery in the northern Baltic Sea. Fisheries Research, 252(??):Article 106336, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001138.

Russell:2021:PLB

[RTB^{+}21] Angela Russell, Matthew D. Taylor, Thomas C. Barnes, Daniel D. Johnson, and Bronwyn M. Gillanders. Potential linkages between juvenile nurseries and exploited populations of mulloway (Argyrosomus japonicus), explored using otolith chemistry. Fisheries Research, 243(??):Article 106063, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783621001910.

Roberts:2025:ACD

[RTHB25] C. Gutmann Roberts, A. S. Tarkan, M. E. Hanley, and J. R. Britton. Angler catch data as a monitoring tool for European barbel Barbus barbus in a data limited recreational fishery. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002881.

Roa-Ureta:2020:ASE

[RUHM20] Ruben H. Roa-Ureta, Jorge Henríquez, and Carlos Molinet. Achieving sustainable exploitation through comanagement in three Chilean small-scale fisheries. Fisheries Research, 230(??):Article 105674, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301910.

Rourke:2023:PUE

[RWB^{+}23] Meaghan L. Rourke, Jackson Wilkes Walburn, Matt K. Broadhurst, Ashley M. Fowler, Julian M. Hughes, D. Stewart Fielder, Joseph D. DiBattista, and Elise M. Furlan. Poor utility of environmental DNA for estimating the biomass of a threatened freshwater teleost; but clear direction for future candidate assessments. Fisheries Research, 258(??):Article 106545, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003228.

Rittweg:2025:VOC

[RWFT25] Timo D. Rittweg, Michael Wiedenbeck, Jan Fietzke, and Clive Trueman. Varying organic content in fish otoliths: Effects on SIMS-based δ^{18}O measurements and possible corrections. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624003035.

Rooper:2020:EHS

[RWT^{+}20] Christopher N. Rooper, Kresimir Williams, Richard H. Towler, Rachel Wilborn, and Pam Goddard. Estimating habitat-specific abundance and behavior of several groundfishes using stationary stereo still cameras in the southern California Bight. Fisheries Research, 224(??):Article 105443, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578361930298X.

Silva:2022:TRR

[SAA^{+}22] Guelson Batista Silva, Lisa Elma Ailloud, Justin Monin Amandé, Rafael Ferreira Muniz, Fabio Hissa Vieira Hazin, and Doug Beare. Trophic relationships revealed by dart tags found in the stomachs of large pelagic fishes in the Atlantic Ocean. Fisheries Research, 248(??):Article 106224, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000017.

Svedang:2023:BPF

[SAA23] Henrik Svedäng, Gustaf Almqvist, and Thomas Axenrot. A Baltic pelagic fish community revisited: Indications of profound changes in species composition in the Stockholm Archipelago. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300173X.

Stallings:2023:PRS

[SACS23] Christopher D. Stallings, Oscar Ayala, Tiffanie A. Cross, and Beverly Sauls. Post-release survival of red snapper (Lutjanus campechanus) and red grouper (Epinephelus morio) using different barotrauma mitigation methods. Fisheries Research, 264(??):??, August 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001108.

Solis:2020:DEP

[SAdC20] Daniel Solís, Juan J. Agar, and Julio del Corral. Diversification, efficiency and productivity in catch share fisheries. Fisheries Research, 226(??):Article 105532, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300497.

Suter:2022:COU

[SAHW22] Jenny M. Suter, Robert T. Ames, Brett Holycross, and Jordan T. Watson. Comparing observed and unobserved fishing characteristics in the drift gillnet fishery for swordfish. Fisheries Research, 256(??):Article 106456, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002338.

SanJuan:2024:USS

[SASB24] A. Pérez San Juan, M. L. Ramos Alonso, V. Sierra, and J. C. Báez. Undetected silky sharks (Carcharhinus falciformis) in the wells of the tropical tuna purse seine fleet in the Indian Ocean. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001735.

Satterthwaite:2023:RCM

[Sat23] William H. Satterthwaite. The reproducibility crisis meets stock assessment science: Sources of inadvertent bias in the stock assessment prioritization and review process. Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300156X.

Sepulveda:2020:PRS

[SAW^{+}20] C. A. Sepulveda, S. A. Aalbers, Michael Wang, Jeff Kneebone, and Diego Bernal. Post-release survivorship of Pacific bluefin tuna (Thunnus orientalis) captured in the California recreational fishery. Fisheries Research, 223(??):Article 105413, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302681.

Sculley:2020:QDS

[SB20] Michelle L. Sculley and Jon Brodziak. Quantifying the distribution of swordfish (Xiphias gladius) density in the Hawaii-based longline fishery. Fisheries Research, 230 (??):Article 105638, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301557.

Strang:2024:ICC

[SB24] Cerys I. Strang and Thijs Bosker. The impact of climate change on marine mega-decapod ranges: a systematic literature review. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002297.

Smejkal:2022:BAC

[ŠBB^{+}22] Marek Šmejkal, Daniel Bartoň, Marek Brabec, Zuzana Sajdlová, Allan T. Souza, Karlos R. Moraes, Petr Blabolil, Lukáš Vejřík, and Jan Kubečka. Behaviour affects capture probability by active sampling gear in a cyprinid fish. Fisheries Research, 249(??):Article 106267, May 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000443.

Snape:2024:FNR

[SBB^{+}24] Robin T. E. Snape, Damla Beton, Annette C. Broderick, Lucy C. M. Omeyer, and Brendan J. Godley. Flashing Net-Lights reduce bycatch in small-scale fisheries of the Eastern Mediterranean. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003120.

Stebbins:2024:HMP

[SBBH24] Elizabeth Stebbins, James R. Bence, Travis O. Brenden, and Michael J. Hansen. A hierarchical model of persistent and transient growth variation applied to Lake Superior lake trout. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001450.

Saltalamacchia:2022:PSE

[SBC^{+}22] Francesco Saltalamacchia, Florian Berg, Michele Casini, Julie Coad Davies, and Valerio Bartolino. Population structure of European sprat (Sprattus sprattus) in the Greater North Sea ecoregion revealed by otolith shape analysis. Fisheries Research, 245(??):Article 106131, January 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621002599.

Sbragaglia:2023:PRF

[SBC^{+}23] Valerio Sbragaglia, Jacob W. Brownscombe, Steven J. Cooke, Anthonie D. Buijse, Robert Arlinghaus, and Warren M.

Potts. Preparing recreational fisheries for the uncertain future: an update of progress towards answering the 100 most pressing research questions. Fisheries Research, 263(??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000553.

Seidu:2022:FSI

[SBD^{+}22] Issah Seidu, Lawrence K. Brobbey, Emmanuel Danquah, Samuel K. Oppong, David van Beuningen, Moro Seidu, and Nicholas K. Dulvy. Fishing for survival: Importance of shark fisheries for the livelihoods of coastal communities in Western Ghana. Fisheries Research, 246(??):Article 106157, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362100285X.

Singh:2020:WLS

[SBJ^{+}20] Warsha Singh, Birkir Bárdarson, Sigurdur T. Jónsson, Bjarki Elvarsson, and Christophe Pampoulie. When logbooks show the path: Analyzing the route and timing of capelin (Mallotus villosus) migration over a quarter century using catch data. Fisheries Research, 230(??):Article 105653, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301703.

Silva:2023:ISO

[SBL^{+}23] Tammy L. Silva, Tabitha Breault, Travis M. Lowery, Nicholas M. Calabrese, Kevin D. E. Stokesbury, David N. Wiley, and Gavin Fay. Investigating spatial overlap between northern sand lance (Ammodytes dubius) and Atlantic cod (Gadus morhua) with video trawl technology. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300139X.

Stewart:2022:AME

[SBRM^{+}22] Nathan D. Stewart, Dheeraj S. Busawon, Enrique Rodriguez-Marin, Matt Siskey, and Alex R. Hanke. Applying mixedeffects growth models to back-calculated size-at-age data for Atlantic bluefin tuna (Thunnus thynnus). Fisheries Research, 250(??):Article 106260, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000376.

Sureda:2020:PSE

[SBT^{+}20] Antoni Sureda, Carlos Barceló, Silvia Tejada, Inmaculada Montero, Elisa Langley, and Antonio Box. Physiological and survival effects of capture of red scorpion fish Scorpaena scrofa (Osteichthyes: Scorpaenidae) by different fishing gears in the Balearic Islands (Western Mediterranean). Fisheries Research, 229(??):Article 105616, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301338.

Stacy:2021:TBI

[SBZ^{+}21] Brett Stacy, Paul Burch, Philippe E. Ziegler, Katherine A. Cresswell, Klaas Hartmann, and Richard M. Hillary. Are tag-based integrated stock assessments robust to IUU fishing? Fisheries Research, 243(??):Article 106098, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002265.

Siwicke:2020:PMG

[SC20] Kevin A. Siwicke and Karson Coutré. Periodic movements of Greenland turbot Reinhardtius hippoglossoides in the eastern Bering Sea and Aleutian Islands. Fisheries Research, 229(??):Article 105612, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301296.

Stewart:2022:UTH

[SCC^{+}22] John Stewart, James R. Craig, Chantelle Clain, Anne-Marie Hegarty, Nicholas M. Meadows, Antony Gould, and Caitlin Young. Using their heads — a novel, collaborative approach between industry and scientists to monitor a commercial mullet fishery as a result of COVID-19 restrictions. Fisheries Research, 250(??):Article 106272, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000492.

Sagua:2021:RDN

[SCCAM21] Camila Sagua, Luis A. Cubillos, Cristian M. Canales, and Rubén Alarcón-Muñoz. The recruitment dynamics of the nylon shrimp Heterocarpus reedi, effects of climate and predation off Chile. Fisheries Research, 235(??):Article 105817, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620303349.

Shea:2022:RFF

[SCD^{+}22] Brendan D. Shea, Sydney K. Coulter, Kelly E. Dooling, Hana L. Isihara, Jessica C. Roth, Elliot Sudal, Donald J. Donovan, Lisa A. Hoopes, Alistair D. M. Dove, Steven J. Cooke, and Austin J. Gallagher. Recreational fishing fight times are not correlated with physiological status of blue sharks (Prionace glauca) in the Northwestern Atlantic. Fisheries Research, 248(??):Article 106220, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003489.

Shank:2024:BSR

[SCG^{+}24] Burton Shank, Joshua T. Carloni, Paul Geoghegan, David M. Fields, Andrew G. Goode, Harvey J. Walsh, and Richard A. Wahle. Bridging the spawner–recruit disconnect II: Revealing basin-scale correlations between zooplankton and lobster settlement dynamics in the Gulf of Maine. Fisheries Research, 278(??):??, October 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001462.

Sala-Coromina:2021:EHM

[SCGM^{+}21] Joan Sala-Coromina, Jose Antonio García, Paloma Martín, Ulla Fernandez-Arcaya, and Laura Recasens. European hake (Merluccius merluccius, Linnaeus 1758) spillover analysis using VMS and landings data in a no-take zone in the northern Catalan coast (NW Mediterranean). Fisheries Research, 237(??):Article 105870, May 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303878.

Scheld:2024:MDI

[SCGW24] Andrew M. Scheld, W. Reid Calhoun, Caela B. Gilsinan, and Shelby B. White. Market development for an invasive fish species: Blue catfish in the Chesapeake Bay, US. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001632.

Schuhmann:2023:BRC

[Sch23] Peter W. Schuhmann. Benefits from recreational catch improvements may hinge on fish consumption safety: Evidence from the Cape Fear River, North Carolina. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002266.

Souza:2021:FMR

[SCHSC21] Daniela Santos Souza, Weidy Rozendo Clemente, Frederico Henning, and Antonio Mateo Solé-Cava. From fishmarkets to restaurants: Substitution prevalence along the flatfish commercialization chain in Brazil. Fisheries Research, 243(??):Article 106095, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100223X.

Sbiba:2024:ISS

[SCN^{+}24] Salah Eddine Sbiba, Alberto Teodorico Correia, Nikolaos Nikolioudakis, Hocein Bazairi, and Malika Chlaida. Insights into the stock structure of Scomber colias Gmelin, 1789 along the Northwest African coast inferred from otolith shape analysis. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000195.

Stewart:2025:IAN

[SCS25] Julee Stewart, Shayna Cossette, and Christopher M. Somers. Ice angling for northern pike (Esox lucius) with tip ups: Hook style affects angler catch and fish welfare. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002911.

Stevick:2021:PHR

[SCW21] Bethany C. Stevick, Henry S. Carson, and Ocean Working. The pace of harvest and recovery in geoduck clam stocks fifty years into the fishery. Fisheries Research, 242(??):Article 106018, October 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001466.

Siemann:2021:SDD

[SDBS21] Liese A. Siemann, Farrell H. Davis, Tor A. Bendiksen, and Ronald J. Smolowitz. Scallop dredge design using computational fluid dynamics and flume tank testing and the application of both methods to improving a low profile dredge. Fisheries Research, 241(??):Article 105998, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001260.

Soeth:2022:OPV

[SDC^{+}22] Marcelo Soeth, Felippe Alexandre Daros, Alberto Teodorico Correia, Nidia Noemi Fabré, Reginaldo Medeiros, Caroline Vieira Feitosa, Oscar de Sousa Duarte, Tiago Moraes Lenz, and Henry Louis Spach. Otolith phenotypic variation as an indicator of stock structure of Scomberomorus brasiliensis from the southwestern Atlantic Ocean. Fisheries Research, 252(??):Article 106357, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001345.

Sheridan:2020:CCC

[SDdMG^{+}20] Michael Sheridan, Juana Durán, Maria del Mar Gil, Elena Pastor, and Ian O’Connor. Can crustacean cuticle preserve a record of chronological age? Investigating the utility of decapod crustacean eyestalks for age determination of two European spider crab species. Fisheries Research, 224(??):Article 105467, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303224.

Sousa:2021:ERD

[SdFZFJ21] Raniere Garcez Costa Sousa, Helen Cristina Parazzi de Freitas, Diego Maia Zacardi, and Charles Hanry Faria-Junior. Effects of river dams on the fish guilds in the northwest region of the Brazilian Amazon. Fisheries Research, 243(??):Article 106091, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002198.

Santana:2023:AGA

[SdOR^{+}23] Francisco Marcante Santana, Cícero Diogo Lins de Oliveira, Amanda Graziele Araújo Resende, Antônio Carlos Leal de Castro, Zafira da Silva Almeida, and Rosângela Lessa. Age and growth of the Acoupa Weakfish (Cynoscion acoupa) of the Brazilian Amazonian Coast, through micro and macrostructures in otoliths. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002199.

Siddeek:2022:LBA

[SDV^{+}22] M. S. M. Siddeek, B. Daly, V. Vanek, J. Zheng, and C. Siddon. Length-based approaches to estimating natural mortality using tagging and fisheries data: the example of the eastern Aleutian Islands, Alaska golden king crab (Lithodes aequispinus). Fisheries Research, 251(??):Article 106304, July 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000819.

Shellem:2021:RSF

[SECB21] Claire T. Shellem, Joanne I. Ellis, Darren J. Coker, and Michael L. Berumen. Red Sea fish market assessments indicate high species diversity and potential overexploitation.

Fisheries Research, 239(??):Article 105922, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000503.

Sibley:2023:SSM

[SEM^{+}23a] Edward C. P. Sibley, Travis S. Elsdon, Michael J. Marnane, Alethea S. Madgett, Euan S. Harvey, Thomas Cornulier, Damon Driessen, and Paul G. Fernandes. Sound sees more: a comparison of imaging sonars and optical cameras for estimating fish densities at artificial reefs. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001133.

Synyshyn:2023:SFG

[SEM^{+}23b] Caitlyn Synyshyn, Lucas Eckert, Adrienne McLean, Megan Cyr, and Sigal Balshine. Subtle fisheries gear model differences substantially influence catch rates of an invasive fish. Fisheries Research, 258(??):Article 106524, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003010.

Seung:2022:SDV

[Seu22] Chang K. Seung. Spatial distribution of the value added from seafood exports: a domestic value chain analysis for Korea. Fisheries Research, 247(??):Article 106181, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362100309X.

Spich:2020:CDA

[SF20] Katarzyna Spich and Dariusz P. Fey. Consequences of differences among readers in age estimations of Baltic cod larvae and early juveniles for growth rate and hatch date analysis. is more experience always better? Fisheries Research, 225(??):Article 105500, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300175.

Schaefer:2022:SVR

[SF22a] Kurt M. Schaefer and Daniel W. Fuller. Spatiotemporal variability in the reproductive biology of yellowfin tuna (Thunnus albacares) in the eastern Pacific Ocean. Fisheries Research, 248(??):Article 106225, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000029.

Spich:2022:UOM

[SF22b] K. Spich and D. P. Fey. Using otolith microstructure analysis in studies on the ecology of the early life stages of cod, Gadus morhua L.: a review. Fisheries Research, 250(??):Article 106265, June 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200042X.

Schaefer:2021:PES

[SFC21] Kurt M. Schaefer, Daniel W. Fuller, and Milani Chaloupka. Performance evaluation of a shallow prototype versus a standard depth traditional design drifting fish-aggregating device in the equatorial eastern Pacific tuna purse-seine fishery. Fisheries Research, 233(??):Article 105763, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302800.

Schaefer:2021:PRS

[SFCG^{+}21] Kurt Schaefer, Daniel Fuller, José Leonardo Castillo-Geniz, Carlos Javier Godinez-Padilla, Michel Dreyfus, and Alexandre Aires da Silva. Post-release survival of silky sharks (Carcharhinus falciformis) following capture by Mexican flag longline fishing vessels in the northeastern Pacific Ocean. Fisheries Research, 234(??):Article 105779, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302964.

Skov:2023:PRE

[SFJ^{+}23] Christian Skov, Keno Ferter, Niels Jepsen, Lars-Flemming Pedersen, Wolf-Christian Lewin, Casper Gundelund, and Marc Simon Weltersbach. Post-release effects of catch and release angling for sea trout: Mortality, growth and wound healing. Fisheries Research, 261(??):Article 106637, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000309.

Schwamborn:2023:CAP

[SFMA23] Ralf Schwamborn, Matheus Oliveira Freitas, Rodrigo Leão Moura, and Alexandre Aschenbrenner. Comparing the accuracy and precision of novel bootstrapped length-frequency and length-at-age (otolith) analyses, with a case study of lane snapper (Lutjanus synagris) in the SW Atlantic. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001285.

Soto:2023:MDT

[SFPR^{+}23] M. Soto, L. Fernández-Peralta, J. Rey, I. Czerwisnki, R. García-Cancela, M. Llope, J. Cabrera-Busto, M. Liébana, and M. G. Pennino. Modelling drivers of trawl fisheries discards using Bayesian spatio-temporal models. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002230.

Somov:2024:CJP

[SFYM24] Aleksey Somov, Edward V. Farley, Ellen M. Yasumiishi, and Megan V. McPhee. Comparison of juvenile Pacific salmon abundance, distribution, and body condition between Western and Eastern Bering Sea using spatiotemporal models. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001504.

Smukall:2021:ELT

[SGD^{+}21] Matthew J. Smukall, Tristan L. Guttridge, Félicie Dhellemmes, Andrew C. Seitz, and Samuel H. Gruber. Effects of leader type and gear strength on catches of coastal sharks in a longline survey around Bimini, The Bahamas. Fisheries Research, 240(??):Article 105989, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100117X.

Sulikowski:2020:OPR

[SGH^{+}20] James A. Sulikowski, Walt Golet, Eric R. Hoffmayer, William B. Driggers, Lisa J. Natanson, Amy Carlson, and Brett B. Sweezey. Observing post-release mortality for dusky sharks, Carcharhinus obscurus, captured in the U.S. pelagic longline fishery. Fisheries Research, 221(??):Article 105341, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783619301961.

Salcioglu:2020:PPD

[ŞGK^{+}20] Asli Şalcıoğlu, Chrysoula Gubili, Grigorios Krey, Adem Yavuz Sönmez, and Raşit Bilgin. Phylogeography and population dynamics of the Eastern Mediterranean whiting (Merlangius merlangus) from the Black Sea, the Turkish Straits System, and the North Aegean Sea. Fisheries Research, 229(??):Article 105614, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301314.

Steele:2024:ECM

[SGKAR24] Reid W. Steele, Jin Gao, Mariano Koen-Alonso, and Paul M. Regular. Exploring capelin (Mallotus villosus) population dynamics using Empirical Dynamic Modelling (EDM). Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001747.

Scheld:2020:QBE

[SGW^{+}20] Andrew M. Scheld, William M. Goldsmith, Shelby White, Hamish J. Small, and Susanna Musick. Quantifying the behavioral and economic effects of regulatory change in a recreational cobia fishery. Fisheries Research, 224(??): Article 105469, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783619303248.

Skov:2022:CRA

[SGW^{+}22] Christian Skov, Casper Gundelund, Marc Simon Weltersbach, Keno Ferter, Sissel K. Bertelsen, and Niels Jepsen. Catch and release angling for sea trout explored by citizen science: Angler behavior, hooking location and bleeding patterns. Fisheries Research, 255(??):Article 106451, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002284.

Sogn-Grundvaag:2020:FMA

[SGZD20] Geir Sogn-Grundvåg, Dengjun Zhang, and Bent Dreyer. Fishing methods for Atlantic cod and haddock: Quality and price versus costs. Fisheries Research, 230 (??):Article 105672, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620301892.

Sogn-Grundvaag:2021:CFA

[SGZD21] Geir Sogn-Grundvåg, Dengjun Zhang, and Bent Dreyer. Competition in a fish auction: the case of Atlantic cod in Northern Norway. Fisheries Research, 235(??):Article 105826, March 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578362030343X.

Sogn-Grundvaag:2022:FTF

[SGZH^{+}22] Geir Sogn-Grundvåg, Dengjun Zhang, Edgar Henriksen, Sjurdur Joensen, Bjørn-Inge Bendiksen, and Øystein Hermansen. Fishing tactics and fish quality: the case of the coastal fishery for Atlantic cod in Norway. Fisheries Research, 246(??):Article 106167, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002952.

Sagarese:2022:ERR

[SH22a] Skyler R. Sagarese and William J. Harford. Evaluating the risks of red tide mortality misspecification when modeling stock dynamics. Fisheries Research, 250(??):Article 106271, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000480.

Stohs:2022:BPF

[SH22b] Stephen M. Stohs and Karter M. Harmon. Bayesian prediction of fishery biological impacts from limited data: a deep-set buoy gear case study. Fisheries Research, 249(??):Article 106228, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622000054.

Sistiaga:2023:TLP

[SHB^{+}23] Manu Sistiaga, Bent Herrmann, Jesse Brinkhof, Roger B. Larsen, Juan Santos, Daniel Stepputtis, Ilmar Brinkhof, Nadine Jacques, Kristine Cerbule, Andrea Petetta, Elsa Cuende, and Liz Kvalvik. Is there a limit to the potential effects of shortening lastridge ropes on the size selectivity of diamond mesh codends? Fisheries Research, 262(??):Article 106671, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000644.

Stewart:2021:FSE

[SHC21] Ian J. Stewart, Allan C. Hicks, and Piera Carpi. Fully subscribed: Evaluating yield trade-offs among fishery sectors utilizing the Pacific halibut resource. Fisheries Research, 234(??):Article 105800, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303179.

Sandison:2021:EIP

[SHH^{+}21] Frances Sandison, Jon Hillier, Astley Hastings, Paul Macdonald, Beth Mouat, and C. Tara Marshall. The environmental impacts of pelagic fish caught by Scottish vessels. Fisheries Research, 236(??):Article 105850, April 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303672.

Simpson:2020:SER

[SHS20] Samantha J. Simpson, Nicolas E. Humphries, and David W. Sims. The spatial ecology of Rajidae from mark-recapture tagging and its implications for assessing fishery interactions and efficacy of Marine Protected Areas. Fisheries Research, 228(??):Article 105569, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300862.

Sarvala:2020:OSP

[SHV20] Jouko Sarvala, Harri Helminen, and Anne-Mari Ventelä. Overfishing of a small planktivorous freshwater fish, vendace (Coregonus albula), in the boreal lake Pyhäjärvi (SW Finland), and the recovery of the population. Fisheries Research, 230(??):Article 105664, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301818.

Simon:2023:DGE

[Sim23] Janek Simon. Do glass eels restocked in winter have a lower survival rate than glass eels restocked in spring? Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001777.

Shibata:2024:CLE

[SIM^{+}24a] Yasutoki Shibata, Yuka Iwahara, Masahiro Manano, Ayumi Kanaya, Ryota Sone, Satoko Tamura, Naoya Kakuta, Tomoya Nishino, Akira Ishihara, and Shungo Kugai. Corrigendum to “Length estimation of fish detected as nonoccluded using a smartphone application and deep learning method” [Fish. Res. 273 (2024) 106970]. Fisheries Research, 275(??):??, July 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000985. See [SIM^{+}24b].

Shibata:2024:LEF

[SIM^{+}24b] Yasutoki Shibata, Yuka Iwahara, Masahiro Manano, Ayumi Kanaya, Ryota Sone, Satoko Tamura, Naoya Kakuta, Tomoya Nishino, Akira Ishihara, and Shungo Kugai. Length estimation of fish detected as non-occluded using a smartphone application and deep learning method. Fisheries Research, 273(??):??, May 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783624000341. See corrigendum [SIM^{+}24a].

Sakamoto:2024:VHS

[SIMT24] Tatsuya Sakamoto, Toyoho Ishimura, Tomohiko Matsuura, and Motomitsu Takahashi. Vertical habitat selection of sardine juveniles inferred from oxygen stable isotope of otolith. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400170X.

Smith:2024:EDP

[SJ24] James A. Smith and Daniel D. Johnson. Evaluating drivers and predictability of catch composition in a highly mixed trawl fishery using stacked and joint species distribution models. Fisheries Research, 279(??):??, November 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002157.

Saemundsson:2020:DJC

[SJB^{+}20] Kristinn Saemundsson, Jónas P. Jonasson, Gavin A. Begg, Hjalti Karlsson, Gudrun Marteinsdottir, and Ingibjörg G. Jónsdóttir. Dispersal of juvenile cod (Gadus morhua L.) in Icelandic waters. Fisheries Research, 232(??):Article 105721, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302381.

Silburn:2020:ESW

[SJBT20] Jack Silburn, Daniel D. Johnson, David J. Booth, and Matthew D. Taylor. The effect of subsampling when monitoring bycatch in a penaeid trawl fishery. Fisheries Research, 224(??):Article 105459, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303145.

Schilling:2023:LTD

[SJH^{+}23] Hayden T. Schilling, Daniel. D. Johnson, Roshan Hanamseth, Iain M. Suthers, and Matthew D. Taylor. Long-term drivers of catch variability in south-eastern Australia’s largest portunid fishery. Fisheries Research, 260(??):Article 106582, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622003599.

Squires:2023:EVM

[SJTGAS23] Dale Squires, Ramón Jiménez-Toribio, Patrice Guillotreau, and Jimmy Anastacio-Solis. The ex-vessel market for tropical tuna in Manta, Ecuador. A new key player on the global tuna market. Fisheries Research, 262(??):Article 106646, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000395.

Staton:2022:AUW

[SJW^{+}22] Benjamin A. Staton, Casey Justice, Seth White, Edwin R. Sedell, Lauren A. Burns, and Matthew J. Kaylor. Accounting for uncertainty when estimating drivers of imperfect detection: an integrated approach illustrated with snorkel surveys for riverine fishes. Fisheries Research, 249(??):Article 106209, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003374.

Samy-Kamal:2021:FIP

[SK21a] Mohamed Samy-Kamal. Fishery Improvement Projects (FIPs): a global analysis of status and performance. Fisheries Research, 240(??):Article 105987, August 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001156.

Samy-Kamal:2021:PES

[SK21b] Mohamed Samy-Kamal. Prices in the Egyptian seafood market: insights for fisheries management and food security. Fisheries Research, 233(??):Article 105764, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302812.

Shimose:2022:IBC

[SK22] Tamaki Shimose and Minoru Kanaiwa. Influence of the body color and size on the market value of wild captured coralgroupers (Serranidae, Plectropomus): Implications for fisheries management. Fisheries Research, 248(??):Article 106223, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621003519.

Slaton:2023:EAM

[SKBA23] Catherine Slaton, Dieter Koemle, Max Birdsong, and Robert Arlinghaus. Explaining attitudes to management actions and beliefs about other user groups and conservation with angler characteristics: a case study in a coastal pike (Esox lucius) fishery in the southern Baltic Sea, Germany. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000620.

Skjaeraasen:2020:LAV

[SKD^{+}20] Jon E. Skjæraasen, Knut Korsbrekke, Gjert Endre Dingsør, Øystein Langangen, Anders Frugård Opdal, and Christian Jørgensen. Large annual variation in the amount of skipped spawning for female Northeast Arctic haddock Melanogrammus aeglefinus. Fisheries Research, 230(??):Article 105670, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620301879.

Sundin:2025:AMD

[SKJJ25] Josefin Sundin, Konrad Karlsson, Birgitta Jacobson, and Philip Jacobson. Assessment of mortality during trap and transport in adult European eel. Fisheries Research, 281(??): ??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000013.

Silas:2023:SCI

[SKS^{+}23] Mathew O. Silas, Mary A. Kishe, Masumbuko R. Semba, Bigeyo N. Kuboja, Benjamin Ngatunga, Said S. Mgeleka, Hans W. Linderholm, Martin Dahl, and Martin Gullström. Seascape configuration influences big blue octopus (Octopus cyanea) catches: Implications for a sustainable fishery. Fisheries Research, 264(??):??, August 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001091.

Sarr:2023:SCP

[SKST23] Ousmane Sarr, Richard Kindong, Fambaye Ngom Sow, and Siquan Tian. Standardized catch per unit effort and size compositions of Atlantic bonito, Sarda sarda (Bloch, 1793), harvested by artisanal fisheries in the Senegalese Exclusive Economic Zone (SEEZ). Fisheries Research, 261(??):Article 106626, May 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362300019X.

Scherrer:2021:EGP

[SKW^{+}21] Stephen R. Scherrer, Donald R. Kobayashi, Kevin C. Weng, Henry Y. Okamoto, Francis G. Oishi, and Erik C. Franklin. Estimation of growth parameters integrating tag-recapture, length-frequency, and direct aging data using likelihood and Bayesian methods for the tropical deepwater snapper Pristipomoides filamentosus in Hawaii. Fisheries Research, 233(??):Article 105753, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302708.

Saqib:2024:FED

[SKY^{+}24] Muhammad Saqib, Muhammad Rizwan Khokher, Xin Yuan, Bo Yan, Douglas Bearham, Carlie Devine, Candice Untiedt, Toni Cannard, Kylie Maguire, Geoffrey N. Tuck, L. Rich Little, and Dadong Wang. Fishing event detection and species classification using computer vision and artificial intelligence for electronic monitoring. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002054.

Smith:2024:IPG

[SL24] Jordan W. Smith and Chase C. Lamborn. The influence of population growth and weather on the value of recreational angling trips within Utah (USA). Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000407.

Signaroli:2024:NCV

[SLA24] Marco Signaroli, Arancha Lana, and Josep Alós. Novel computer vision tools applied to marine recreational fisheries spatial planning. Fisheries Research, 271(??):??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362300317X.

Sass:2023:DAC

[SLF23] Greg G. Sass, Samuel T. LaMarche, and Zachary S. Feiner. Do angler catch and harvest rates differ between open water and ice anglers in Wisconsin? Fisheries Research, 263(??): ??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000711.

Su:2020:ADJ

[SLW^{+}20] Nan-Jay Su, Yi-Sin Lu, Chia-Hui Wang, Cheng-Hsin Liao, Wei-Chuang Chiang, and Chen-Te Tseng. Age determination for juvenile fourfinger threadfin (Eleutheronema rhadinum) by using otolith microstructure and length data obtained from commercial fisheries off northwestern Taiwan. Fisheries Research, 227(??):Article 105560, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300771.

Sha:2021:NIP

[SLW^{+}21] Hang Sha, Xiang-Zhong Luo, Dan Wang, Xiao hui Li, Gui-Wei Zou, and Hong-Wei Liang. New insights to protection and utilization of silver carp (Hypophthalmichthys molitrix) in Yangtze River based on microsatellite analysis. Fisheries Research, 241(??):Article 105997, September 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001259.

Stock:2021:WHA

[SM21] Brian C. Stock and Timothy J. Miller. The Woods Hole Assessment Model (WHAM): a general state-space assessment framework that incorporates time- and age-varying processes via random effects and links to environmental covariates. Fisheries Research, 240(??):Article 105967, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000953.

Spence:2024:AEM

[SMA^{+}24] Michael A. Spence, James A. Martindale, Khatija Alliji, Hayley J. Bannister, Robert B. Thorpe, Nicola D. Walker, Peter J. Mitchell, Matthew R. Kerr, and Paul J. Dolder. Assessing the effect of multispecies interactions on precautionary reference points using an ensemble modelling approach: a North Sea case study. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002248.

Stoltenberg:2024:PBS

[SMC^{+}24] Ina Stoltenberg, Felix Mittermayer, Catriona Clemmesen, Jan Dierking, and Jamileh Javidpour. Predation on Baltic Sea yolk-sac herring larvae (Clupea harengus) by the invasive ctenophore Mnemiopsis leidyi. Fisheries Research, 273(??): ??, May 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000377.

Szynaka:2024:RIC

[SME^{+}24] Monika Jadwiga Szynaka, Pedro Monteiro, Karim Erzini, Jorge M. S. Gonçalves, and Aida Campos. Reducing invertebrate by-catch in a coastal fishery using a raised monofilament trammel net. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001048.

Schaub:2024:LLC

[SMK^{+}24] Michael Schaub, Mark N. Maunder, Marc Kéry, James T. Thorson, Eiren K. Jacobson, and André E. Punt. Lessons to be learned by comparing integrated fisheries stock assessment models (SAMs) with integrated population models (IPMs). Fisheries Research, 272(??):??, April 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003181.

Spoors:2021:ABU

[SMKJ21] Felicity Spoors, Tania Mendo, Nicola Khan, and Mark James. Assessing bait use by static gear fishers of the Scottish inshore fisheries: a preliminary study. Fisheries Research, 240(??):Article 105974, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001028.

Su:2024:DLB

[SML^{+}24] Shuyang Su, Qinghong Mao, Yangdong Li, Hongli Li, Juntai Leng, and Chengqian Lu. Deep learning-based fishing ground prediction for albacore and yellowfin tuna in the Western and Central Pacific Ocean. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400167X.

Sutton:2024:SAP

[SMLT24] Jolene T. Sutton, Jenni L. McDermid, Lysandre Landry, and François Turcotte. Spatiotemporal analysis provides solutions to mitigate bycatch of southern Gulf of St. Lawrence Atlantic cod in an expanding redfish fishery. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001024.

Strasburger:2023:UJS

[SNHM23] Wesley W. Strasburger, Dave Nicolls, Christine M. Hinds, and Kevin W. McNeel. The utility of juvenile sablefish otoliths in reconstructing life history and growth in the Gulf of Alaska. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002345.

Shanmughan:2024:OSR

[SNJ^{+}24] Ashna Shanmughan, Suman Nama, Ashok Kumar Jaiswar, Karankumar Ramteke, Binaya Bhusan Nayak, Annam Pavan-Kumar, and Shashi Bhushan. Otolith shape reveals the population structure of ’Malabar Labeo’, Labeo dussumieri (Valenciennes, 1842), in the four west-flowing rivers of India. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001486.

Santos:2025:BIF

[SO25] Régis Santos and Eduardo Ortega. Bibliometric insights into fisheries value chain research. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000062.

Shiode:2021:DBC

[SOS^{+}21] Daisuke Shiode, Jun Okamoto, Maika Shiozawa, Keiichi Uchida, Yoshinori Miyamoto, Fuxiang Hu, and Tadashi Tokai. Differences in the behavioral characteristics between green and loggerhead turtles in a setnet bycatch simulation. Fisheries Research, 242(??):Article 106036, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001648.

Sokolova:2022:TDS

[SOSK22] Maria Sokolova, Finbarr Gerard O’Neill, Esther Savina, and Ludvig Ahm Krag. Test and development of a sediment suppressing system for catch monitoring in demersal trawls. Fisheries Research, 251(??):Article 106323, July 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200100X.

Spanier:2024:MPA

[Spa24] Ehud Spanier. Marine protected areas – a review of their potential effects on lobster population size & structure and fisheries management. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000894.

Satumanatpan:2022:IFE

[SPC22] Suvaluck Satumanatpan, Richard Pollnac, and Ratana Chuenpagdee. Incorporating fishers’ evaluation of adaptive capacity in policy making in Thailand. Fisheries Research, 254(??):Article 106407, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001849.

Skovrind:2023:UNA

[SPC^{+}23] Mikkel Skovrind, George Pacheco, Emil Aputsiaq Flindt Christensen, Shyam Gopalakrishnan, Katharina Fietz, Tore Hejl Holm-Hansen, Filipe Garrett Vieira, Marcus Anders Krag, Henrik Carl, M. Thomas P. Gilbert, Morten Tange Olsen, and Peter Rask Møller. Uncovering neutral and adaptive genomic differentiation among European perch with brackish water and freshwater origin in the western Baltic Sea region. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002394.

Sepe:2025:IFI

[SPC^{+}25] Emily Sepe, Federica Poli, Federico Calì, Simone D’Acunto, Carlotta Mazzoldi, and Matteo Barbato. Implementing a fishery improvement programme to manage the common cuttlefish (Sepia officinalis) in artisanal sea and lagoon fisheries: the case study of the Chioggia’s fleet. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362400273X.

Spanou:2024:WSV

[SPD^{+}24] Despina S. Spanou, Panagiota Petroudi, Elli Dimou, Konstantinos Kokkinos, and Dimitris Klaoudatos. Walleye (Sander vitreus, Mitchill 1818) age and sex classification using innovative supervised and unsupervised machine learning and soft computing methodologies. Fisheries Research, 275 (??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400095X.

Silwal:2023:COP

[SPE^{+}23] Pratikshya Silwal, Neelam C. Poudyal, Augustin Engman, Xuqi Chen, and Kevin Cavasos. Catch orientation and public lake stocking preferences: a typology of freshwater anglers. Fisheries Research, 257(??):Article 106500, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002776.

Secci:2021:UTA

[SPG^{+}21] Marco Secci, Francesco Palmas, Ambra Angelica Giglioli, Viviana Pasquini, Jacopo Culurgioni, Andrea Sabatini, and Piero Addis. Underwater tagging of the Atlantic bluefin tuna in the trap fishery of Sardinia (W Mediterranean). Fisheries Research, 233(??):Article 105747, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302642.

Sion:2024:CNW

[SPM^{+}24] Letizia Sion, Alessio Pollice, Porzia Maiorano, Crescenza Calculli, Francesca Capezzuto, Angela Carluccio, Pasquale Ricci, and Gianfranco D’Onghia. Chondrichthyes in the north-western Ionian Sea (central Mediterranean): Species diversity, abundance and spatio-temporal changes. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000419.

Saygin:2022:SIO

[SPW^{+}22] Semra Saygın, Nazmi Polat, Malte Willmes, Levi S. Lewis, James A. Hobbs, Ataman Altuğ Atıcı, and Mahmut Elp. Strontium isotopes in otoliths reveal a diversity of natal origins for Tarek (Alburnus tarichi) in Lake Van, Turkey. Fisheries Research, 255(??):Article 106441, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002181.

Smallwood:2020:BRS

[SR20] Claire B. Smallwood and Karina L. Ryan. Benefits of a restricted spatial and temporal survey design for determining average weight of recreational catches. Fisheries Research, 232(??):Article 105735, December 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302526.

Sanchez:2021:AGM

[SR21] Phillip J. Sanchez and Jay R. Rooker. Age, growth, and mortality of threatened Warsaw grouper, Hyporthodus nigritus, in the Gulf of Mexico. Fisheries Research, 243(??):Article 106097, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002253.

Sulanke:2025:AEF

[SRB^{+}25] E. Sulanke, V. Rubel, J. Berkenhagen, M. Bernreuther, T. Stoeck, and S. Simons. Amending the European fishing fleet segmentation based on machine learning and multivariate statistics. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002546.

Smallwood:2024:SLR

[SRF^{+}24] Claire B. Smallwood, Karina L. Ryan, Ewan A. Flanagan, Jade Q. Maggs, Faith A. Ochwada-Doyle, and Sean R. Tracey. Spiny lobster recreational fisheries in Australia and New Zealand: an overview of regulations, monitoring, assessment and management. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002133.

Schilling:2022:EVR

[SRP^{+}22] Hayden T. Schilling, Indiana J. Riley, Aaron C. Puckeridge, Alexandra Milne-Muller, and Corey T. Callaghan. Economic value of regional spearfishing competitions. Fisheries Research, 250(??):Article 106289, June 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000662.

Sobrino:2020:API

[SRT^{+}20] Ignacio Sobrino, Lucia Rueda, Maria Pilar Tugores, Candelaria Burgos, Miguel Cojan, and Graham J. Pierce. Abundance prediction and influence of environmental parameters in the abundance of octopus (Octopus vulgaris Cuvier, 1797) in the Gulf of Cadiz. Fisheries Research, 221(??):Article 105382, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783619302371.

Satterthwaite:2023:MAR

[SS23] William. H. Satterthwaite and Andrew Olaf Shelton. Methods for assessing and responding to bias and uncertainty in U.S. West Coast salmon abundance forecasts. Fisheries Research, 257(??):Article 106502, January 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200279X.

Sampognaro:2024:BTP

[SS24] Lia Sampognaro and Angel M. Segura. Biogeographic trends in populations of the shrimp Artemesia longinaris (Decapoda:Penaeidae) on the Atlantic coast of South America. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000183.

Souza:2020:OID

[SSD^{+}20a] A. T. Souza, K. Soukalová, V. Děd, M. Šmejkal, P. Blabolil, M. Říha, T. Jůza, M. Vašek, M. Čech, J. Peterka, L. Vejřík, I. Vejříková, M. Tušer, M. Muška, M. Holubová, D. S. Boukal, and J. Kubečka. Ontogenetic and interpopulation differences in otolith shape of the European perch (Perca fluviatilis). Fisheries Research, 230(??):Article 105673, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301909.

Souza:2020:OSV

[SSD^{+}20b] A. T. Souza, K. Soukalová, V. Děd, M. Šmejkal, K. Moraes, M. Říha, M. Muška, J. Frouzová, and J. Kubečka. Otolith shape variations between artificially stocked and autochthonous pikeperch (Sander lucioperca). Fisheries Research, 231(??):Article 105708, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302253.

Schneider:2024:SHF

[SSFL24] Alexandra K. Schneider, Jeffrey D. Shields, Mary C. Fabrizio, and Romuald N. Lipcius. Spawning history, fecundity, and potential sperm limitation of female blue crabs in Chesapeake Bay. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001589.

Susanto:2022:BRR

[SSG^{+}22] Adi Susanto, Petri Suuronen, Saeid Gorgin, Ririn Irnawati, Mochammad Riyanto, Wahyudin, Hery Sutrawan Nurdin, Asep Hamzah, Fahresa Nugraheni Supadminingsih, and Hendrawan Syafrie. Behavioral response and retinal adaptation of blue swimming crab (Portunus pelagicus) exposed to LED lights — led light as a potential artificial attractant in trap fishing. Fisheries Research, 250(??):Article 106274, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622000510.

Shaff:2023:DHC

[SSI^{+}23] Jacquelyn F. Shaff, Irving Alexis Medina Santiago, Xochitl Elías Ilosvay, Javier Tovar-Ávila, Elena Ojea, Anne H. Beaudreau, Jennifer E. Caselle, and Eréndira Aceves Bueno. Documenting historical changes in shark fisheries near Islas Marías, Mexico, using fishers’ local ecological knowledge. Fisheries Research, 265(??):??, September 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001418.

Sziraki:2021:TEA

[SSJ^{+}21] Bence Sziráki, Ádám Staszny, Vera Juhász, András Weiperth, Gábor Nagy, Ferenc Fodor, Mihály Havranek, Tamás Koltai, Zsolt Szári, Béla Urbányi, and Árpád Ferincz. Testing the efficiency of artificial spawning nests for pikeperch (Sander lucioperca L.) under natural conditions (Lake Balaton, Hungary). Fisheries Research, 243(??):Article 106070, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001983.

Schrandt:2021:ASD

[SSKS21] Meagan N. Schrandt, Colin P. Shea, Benjamin N. Kurth, and Theodore S. Switzer. Amending survey design to improve statistical inferences: Monitoring recruitment of juvenile reef fish in the eastern Gulf of Mexico. Fisheries Research, 241(??):Article 106015, September 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001430.

Silas:2023:UFL

[SSM^{+}23] Mathew O. Silas, Masumbuko L. Semba, Said S. Mgeleka, Lisa Van Well, Hans W. Linderholm, and Martin Gullström. Using fishers’ local ecological knowledge for management of small-scale fisheries in data-poor regions: Comparing seasonal interview and field observation records in East Africa. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001145.

Schroeder:2022:SSB

[SSP^{+}22] Rafael Schroeder, Paulo R. Schwingel, Edgar Pinto, Agostinho Almeida, and Alberto T. Correia. Stock structure of the Brazilian sardine Sardinella brasiliensis from Southwest Atlantic Ocean inferred from otolith elemental signatures. Fisheries Research, 248(??):Article 106192, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003209.

Schrobback:2023:STF

[SSP^{+}23] Peggy Schrobback, Karsten Schrobback, Sean Pascoe, Stephanie McWhinnie, and Eriko Hoshino. Spatial and temporal fishery management assessment using DEA: Case study of spanner crabs in Queensland, Australia. Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001820.

Sellinger:2024:ROA

[SSP24] Emily L. Sellinger, Cody Szuwalski, and André E. Punt. The robustness of our assumptions about recruitment: a reexamination of marine recruitment dynamics with additional data and novel methods. Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002552.

Santos:2024:LMB

[SSRC24] Catarina C. Santos, Miguel N. Santos, Daniela Rosa, and Rui Coelho. Leader material and bait effects on target and bycatch species caught in an Atlantic Ocean pelagic longline fishery. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001577.

Schroeder:2023:CNA

[SSS^{+}23] Rafael Schroeder, Paulo R. Schwingel, Richard Schwarz, Felippe A. Daros, Taynara P. Franco, Natasha T. Hoff, Ana Méndez Vicente, Jorge P. Castro, André M. Vaz dos Santos, and Alberto T. Correia. Contribution of the nursery areas to the major fishing grounds of the Brazilian sardine (Sardinella brasiliensis) in Southeastern Brazilian Bight inferred from otolith fingerprints. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002187.

Saltalamacchia:2025:TSO

[SSSF25] Francesco Saltalamacchia, Martine Røysted Solås, Anne Gro Vea Salvanes, and Arild Folkvord. Time series of otolith growth patterns reflects oxygen depletion tolerance in an abundant mesopelagic fish species. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003229.

Sajikumar:2020:DAG

[SSV^{+}20] Kurichithara K. Sajikumar, Geetha Sasikumar, V. Venkatesan, R. Vidya, P. S. Alloycious, K. M. Jestin Joy, P. S. Karamathullah, G. D. Nataraja, and Kolliyil S. Mohamed. Distribution, age and growth of the diamondback squid, Thysanoteuthis rhombus (Cephalopoda: Thysanoteuthidae) from the tropical Arabian Sea. Fisheries Research, 224(??):Article 105478, April 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619303339.

Sato:2020:NMT

[SSY20] Taku Sato, Takuma Sugaya, and Hayate Yoshikawa. Novel method of tagging the kuruma prawn Penaeus japonicus with a trans-molting retentive external eye (TRAMORE) tag. Fisheries Research, 225(??):Article 105482, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303376.

Sablan:2025:EDN

[STH25] Leilani Sablan, Brett Taylor, and Peter Houk. Environmental drivers of non-commercial reef fisheries in Guam.

Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002443.

Syversen:2023:ARF

[SV23] Tore Syversen and Jørgen Vollstad. Application of radio frequency identification tags for marking of fish gillnets and crab pots: Trials in the Norwegian Sea and the Barents Sea, Norway. Fisheries Research, 259(??):Article 106557, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003344.

Simon:2020:LTR

[SW20] Janek Simon and Håkan Wickström. Long-term retention of alizarin red S marks and coded wire tags in European eels. Fisheries Research, 224(??):Article 105453, April 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361930308X.

Seung:2024:SDE

[SW24] Chang K. Seung and Edward Waters. Spatial distribution of economic multipliers for Southwest Alaska fisheries. Fisheries Research, 276(??):??, August 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000997.

Setyastuti:2024:NID

[SWH^{+}24a] Ana Setyastuti, Ismiliana Wirawati, Hadiyanto Hadiyanto, Nurjamin Nurjamin, Sandi Permadi, Tri Aryono Hadi, Bayu Prayudha, Muhammad Hafizt, Indra Bayu Vimono, Marindah Yulia Iswari, Ludi Parwadani Aji, Ardi Ardiansyah, I. Wayan Eka Dharmawan, Suratno Suratno, Muhammad Masrur Islami, Lisa Fajar Indriana, and Nurul Dhewani Mirah Sjafrie. New insight into the diversity, biometric distribution, and relationships of commercial sea cucumber species from Indonesia. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624001887.

Stewart:2024:EAM

[SWH24b] Nathan D. Stewart, Kyle L. Wilson, and Alex Hanke. Estimating age-at-maturity for individual Atlantic bluefin tuna (Thunnus thynnus) from back-calculated growth trajectories. Fisheries Research, 278(??):??, October 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001668.

Stoltenberg:2021:TNV

[SWIRF21] Ina Stoltenberg, Karl Michael Werner, Stefanie M. H. Ismar-Rebitz, and Heino O. Fock. Trophic niche variability influences body condition of female Atlantic cod (Gadus morhua) on organosomatic and biochemical levels. Fisheries Research, 239(??):Article 105921, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000497.

Schwartzkopf:2020:THD

[SWLH20] Brittany D. Schwartzkopf, Alison D. Whitman, Amy J. Lindsley, and Scott A. Heppell. Temporal and habitat differences in the juvenile demersal fish community at a marine-dominated northeast Pacific estuary. Fisheries Research, 227(??):Article 105557, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300746.

Song:2024:PCS

[SXK^{+}24] Dade Song, Ying Xiong, Zhongjie Kang, Long Liang, Xiaorui Wu, Dongjia Li, and Dongyan Liu. Population connectivity of small yellow croaker (Larmichthys polyactis) in the southern Yellow Sea: Implications from multiple otolithbased approaches. Fisheries Research, 276(??):??, August 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624001280.

Stock:2021:ITD

[SXM^{+}21] Brian C. Stock, Haikun Xu, Timothy J. Miller, James T. Thorson, and Janet A. Nye. Implementing two-dimensional autocorrelation in either survival or natural mortality improves a state-space assessment model for Southern New England–Mid Atlantic yellowtail flounder. Fisheries Research, 237(??):Article 105873, May 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000011.

Satoh:2021:SSS

[SXMV^{+}21] Keisuke Satoh, Haikun Xu, Carolina V. Minte-Vera, Mark N. Maunder, and Toshihide Kitakado. Size-specific spatiotemporal dynamics of bigeye tuna (Thunnus obesus) caught by the longline fishery in the eastern Pacific Ocean. Fisheries Research, 243(??):Article 106065, November 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621001934.

Shen:2025:BMB

[SYZ^{+}25] Duqing Shen, Jie Yin, Yunlei Zhang, Chongliang Zhang, Binduo Xu, Yupeng Ji, Yiping Ren, and Ying Xue. Bayesian modeling-based analysis on the shared habitat and species association between four Gobiidae in a marine bay ecosystem. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002765.

Sun:2024:YDS

[SZS^{+}24] Yueying Sun, Shengmao Zhang, Yongchuang Shi, Fenghua Tang, Junlin Chen, Ying Xiong, Yang Dai, and Lin Li. YOLOv7-DCN-SORT: an algorithm for detecting and counting targets on Acetes fishing vessel operation. Fisheries Research, 274(??):??, June 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362400047X.

Szuwalski:2022:ETV

[Szu22] Cody Szuwalski. Estimating time-variation in confounded processes in population dynamics modeling: a case study for snow crab in the eastern Bering Sea. Fisheries Research, 251(??):Article 106298, July 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000753.

Su:2025:VFC

[SZXC25] Li Su, Kui Zhang, Youwei Xu, and Zuozhi Chen. Variations in the fish community of the Beibu Gulf (South China Sea) following fishery resources protection measures. Fisheries Research, 283(??):??, March 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362500030X.

Tabatabaei:2020:MSA

[TAA^{+}20] Seyedeh Narjes Tabatabaei, Asghar Abdoli, Faraham Ahmadzadeh, Craig R. Primmer, Akarapong Swatdipong, and Iraj Hashemzadeh Segherloo. Mixed stock assessment of lakerun Caspian Sea trout Salmo caspius in the Lar National Park, Iran. Fisheries Research, 230(??):Article 105644, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301612.

Trijoulet:2023:MVC

[TAK^{+}23] Vanessa Trijoulet, Christoffer Moesgaard Albertsen, Kasper Kristensen, Christopher M. Legault, Timothy J. Miller, and Anders Nielsen. Model validation for compositional data in stock assessment models: Calculating residuals with correct properties. Fisheries Research, 257(??):Article 106487, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783622002648.

Thayer:2021:IAV

[TBF^{+}21] J. A. Thayer, Z. Burr, J. C. Field, R. D. Carle, and P. Warzybok. Inter-annual variability in forage fish population size structure: Comparison of selectivity of traditional vs. non-traditional sampling devices. Fisheries Research, 234(??):Article 105801, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303180.

Temming:2022:UHD

[TBH^{+}22] Axel Temming, Anne Bönisch, Wilhelm Hagen, Charlotte Brenneken, and Andreas Dänhardt. Unexpected high discard mortalities of juvenile brown shrimp (Crangon crangon) in the North Sea shrimp fishery. Fisheries Research, 252(??):Article 106354, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200131X.

Tray:2022:IPD

[TBÓ^{+}22] Elizabeth Tray, Deirdre Brophy, Niall Ó Maoiléidigh, Ross Finlay, Elvira de Eyto, Russell Poole, and Quentin G. Crowley. Investigating post-depositional alteration of trace elements in fish scales using tagged and recaptured wild salmon. Fisheries Research, 248(??):Article 106207, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621003350.

Taylor:2024:UML

[TC24] Matthew D. Taylor and Rowan C. Chick. Utility of a modified lightweight recreational scallop dredge for surveys of estuarine ark clams (cockles). Fisheries Research, 269(??):??, January 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002679.

Thorson:2021:SSI

[TCJ^{+}21] James T. Thorson, Curry J. Cunningham, Elaina Jorgensen, Andrea Havron, Peter-John F. Hulson, Cole C. Monnahan, and Paul von Szalay. The surprising sensitivity of index scale to delta-model assumptions: Recommendations for model-based index standardization. Fisheries Research, 233(??):Article 105745, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302629.

Trahan:2021:IAL

[TCL^{+}21] Alexandria Trahan, Auston D. Chhor, Luc LaRochelle, Andy J. Danylchuk, and Steven J. Cooke. Influence of artificial lure hook type on hooking characteristics, handling, and injury of angled freshwater gamefish. Fisheries Research, 243(??):Article 106056, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001843.

Tucker:2024:EAE

[TCL^{+}24] Caroline M. Tucker, Simone Collier, Geoffrey Legault, George E. Morgan, and Derrick K. de Kerckhove. Estimating angler effort and catch from a winter recreational fishery using a novel Bayesian methodology to integrate multiple sources of creel survey data. Fisheries Research, 272(??): ??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623003259.

Tulone:2020:CCA

[TCVG20] Antonio Tulone, Maria Crescimanno, Demetris Vrontis, and Antonino Galati. Are coastal communities able to pay for the protection of fish resources impacted by climate change? Fisheries Research, 221(??):Article 105374, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619302292.

Trudeau:2021:EFE

[TDI^{+}21] Ashley Trudeau, Colin J. Dassow, Carolyn M. Iwicki, Stuart E. Jones, Greg G. Sass, Christopher T. Solomon, Brett T. van Poorten, and Olaf P. Jensen. Estimating fishing effort across the landscape: a spatially extensive approach using models to integrate multiple data sources. Fisheries Research, 233(??):Article 105768, January 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362030285X.

Taylor:2021:BVC

[TDJ^{+}21] Ian G. Taylor, Kathryn L. Doering, Kelli F. Johnson, Chantel R. Wetzel, and Ian J. Stewart. Beyond visualizing catch-at-age models: Lessons learned from the r4ss package about software to support stock assessments. Fisheries Research, 239(??):Article 105924, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000527.

Tuffley:2024:DDI

[TdL24] Emma-Jade Tuffley and Simon de Lestang. Densitydependence inside a marine protected area increases natural mortality and stunts the growth of a spiny lobster. Fisheries Research, 280(??):??, December 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002169.

TenBrink:2022:DYS

[Ten22] Todd T. TenBrink. Delineating yellowfin sole (Limanda aspera) reproduction in the northern Bering Sea provides information across the eastern Bering Sea continental shelf. Fisheries Research, 252(??):Article 106335, August 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622001126.

Tokranov:2025:DLH

[TEO25] Alexei M. Tokranov, Pavel O. Emelin, and Alexei M. Orlov. Distribution, life history traits, and ecological significance of bigeye sculpin Triglops nybelini (Cottidae) in Siberian Arctic marine ecosystems. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783625000025.

Tolotti:2020:ADT

[TFC^{+}20] Mariana Travassos Tolotti, Fabien Forget, Manuela Capello, John David Filmalter, Melanie Hutchinson, David Itano, Kim Holland, and Laurent Dagorn. Association dynamics of tuna and purse seine bycatch species with drifting fish aggregating devices (FADs) in the tropical eastern Atlantic Ocean. Fisheries Research, 226(??):Article 105521, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300382.

Thomas:2024:TTQ

[TGG^{+}24] Rebecca E. Thomas, Stéphane Gauthier, Chris Grandin, Allan Hicks, and Sandy Parker-Stetter. To trawl or not to trawl: Questioning core assumptions of trawl placement choice in fisheries acoustics surveys. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002904.

Thorson:2019:PLS

[Tho19] James T. Thorson. Perspective: Let’s simplify stock assessment by replacing tuning algorithms with statistics. Fisheries Research, 217(??):133–139, September 2019. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016578361830033X. See corrigendum [Tho21].

Thorson:2021:CPL

[Tho21] James T. Thorson. Corrigendum to “Perspective: Let’s simplify stock assessment by replacing tuning algorithms with statistics” [Fish. Res. 219 (2019) 133–139]. Fisheries Research, 235(??):Article 105852, March 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303696. See [Tho19].

Tuohy:2023:MSB

[TJS23] Adrian M. Tuohy, Aaron T. Jorgenson, and John R. Skalski. Maximizing salmonid bycatch survival with passively operated commercial fish traps. Fisheries Research, 257(??):Article 106495, January 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002727.

Tate:2021:SBM

[TKB^{+}21] Rick D. Tate, Brendan P. Kelaher, Craig P. Brand, Christopher R. Gallen, Stephen D. A. Smith, and Paul A. Butcher. Shark behaviour and marine faunal assemblage beneath SMART drumlines. Fisheries Research, 243(??):Article 106102, November 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002307.

Trobbiani:2025:TLL

[TLAM25] Gastón Trobbiani, Getino Mamet Leandro, Irigoyen Alejo, and Parma Ana María. “Toki”, a light low-cost video system for seabed research: Performance and precision of Tehuelche scallop (Aequipecten tehuelchus) survey estimates in San José Gulf, Argentina. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624003126.

Tian:2022:ODE

[TLC^{+}22] Honglin Tian, Jinhu Liu, Liang Cao, Tao Zuo, and Shuozeng Dou. Otolith development and elemental incorporation in response to seawater acidification in the flounder Paralichthys olivaceus at early life stages. Fisheries Research, 252(??):Article 106359, August 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001369.

Tian:2021:TSE

[TLCD21] Honglin Tian, Jinhu Liu, Liang Cao, and Shuozeng Dou. Temperature and salinity effects on strontium and barium incorporation into otoliths of flounder Paralichthys olivaceus at early life stages. Fisheries Research, 239(??):Article 105942, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000709.

Thys:2023:BDI

[TLV23] Kelly Johanna Marie Thys, Laura Lemey, and Noémi Van Bogaert. Blondes do it better? A comparative study on the morphometry and life-history traits of commercially important skates blonde ray Raja brachyura, thornback ray Raja clavata, and spotted ray Raja montagui, with management implications. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000723.

Torres:2022:RAC

[TMDA22] Paulo Torres, David Milla i Figueras, Hugo Diogo, and Pedro Afonso. Risk assessment of coastal fisheries in the Azores (north-eastern Atlantic). Fisheries Research, 246(??):Article 106156, February 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002848.

Thorson:2023:DWI

[TMH23] James T. Thorson, Cole C. Monnahan, and Peter-John F. Hulson. Data weighting: an iterative process linking surveys, data synthesis, and population models to evaluate misspecification. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001558.

Takeshige:2021:EIB

[TMN^{+}21] Aigo Takeshige, Mai Miyamoto, Yoji Narimatsu, Shiroh Yonezaki, and Masashi Kiyota. Evaluation of impacts of bottom fishing on demersal habitats: a case study off the Pacific coast of north-eastern Japan. Fisheries Research, 238(??):Article 105916, June 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000448.

Thorson:2020:DST

[TMP20] James T. Thorson, Mark N. Maunder, and E. Punt. The development of spatio-temporal models of fishery catch-per-unit-effort data to derive indices of relative abundance. Fisheries Research, 230(??):Article 105611, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301284.

Tinlin-Mackenzie:2023:TEE

[TMSS^{+}23] Ashleigh Tinlin-Mackenzie, Heather Sugden, Catherine L. Scott, Robert Kennedy, and Clare Fitzsimmons. Trawling for evidence: an ecosystem-based multi-method trawling impact assessment. Fisheries Research, 268(??):??, December 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623002515.

Temperoni:2023:SVA

[TNDM23] Brenda Temperoni, Rocío Isla Naveira, Carla Derisio, and Agueda E. Massa. Spatial variations in age-0+ Merluccius hubbsi (Patagonian stock; 45^{o}–47^{o}S) feeding conditions through fatty acids composition. Fisheries Research, 260(??):Article 106596, April 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622003733.

Tran:2020:ETE

[TNTN20] Phu D. Tran, Luong T. Nguyen, Phuong V. To, and Khanh Q. Nguyen. Effects of the trap entrance designs on the catch efficiency of swimming crab Charybdis feriata fishery. Fisheries Research, 232(??):Article 105730, December 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302472.

Third:2024:PIS

[TP24] G. M. Third and D. M. Parsons. Population identification of snapper (Chrysophrys auratus) using body Geometric Morphometrics to inform sustainable fisheries management. Fisheries Research, 280(??):??, December 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002236.

Trumpickas:2020:EVS

[TPD20] Justin Trumpickas, Michael Pinder, and Erin S. Dunlop. Effects of vessel size and trawling on estimates of pelagic fish backscatter in Lake Huron. Fisheries Research, 224(??):Article 105430, April 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302851.

Tracey:2023:PRS

[TPW23] Sean R. Tracey, Julian Pepperell, and Barrett Wolfe. Post release survival of swordfish (Xiphias gladius) caught by a recreational fishery in temperate waters. Fisheries Research, 265(??):??, September 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001352.

Thompson:2023:RPH

[TRN^{+}23] Patrick L. Thompson, Christopher N. Rooper, Jessica Nephin, Ashley E. Park, James R. Christian, Sarah C. Davies, Karen Hunter, Devin A. Lyons, M. Angelica Peña, Beatrice Proudfoot, Emily M. Rubidge, and Amber M. Holdsworth. Response of Pacific halibut (Hippoglossus stenolepis) to future climate scenarios in the Northeast Pacific Ocean. Fisheries Research, 258(??):Article 106540, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003174.

Trotter:2024:STE

[TRS^{+}24] Adam W. Trotter, Larisa Rathjens, Sophie Schmiegel, Sina Mews, Paul D. Cowley, and Enrico Gennari. Short-term effects of standard procedures associated with surgical transmitter implantation on a benthic shark species requiring anaesthesia. Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002734.

Tenningen:2023:HOC

[TRWH23] Maria Tenningen, Shale Rosen, Taraneh Westergerling, and Nils Olav Handegard. How to obtain clear images from in-trawl cameras near the seabed? A case study from the Barents Sea demersal fishing grounds. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002497.

Talwar:2020:RFI

[TSC^{+}20] Brendan S. Talwar, Jeffrey A. Stein, Stephen M. H. Connett, Stephanie A. Liss, and Edward J. Brooks. Results of a fishery-independent longline survey targeting coastal sharks in the eastern Bahamas between 1979 and 2013. Fisheries Research, 230(??):Article 105683, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620302009.

Thompson:2022:NHB

[TSC^{+}22] Kevin A. Thompson, Theodore S. Switzer, Mary C. Christman, Sean F. Keenan, Christopher L. Gardner, Katherine E. Overly, and Matt D. Campbell. A novel habitatbased approach for combining indices of abundance from multiple fishery-independent video surveys. Fisheries Research, 247(??):Article 106178, March 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621003064.

TenBrink:2025:EUO

[TSG25] Todd T. TenBrink, Jane Y. Sullivan, and Christopher M. Gburski. Exploring the use of otolith shape analysis to identify the stock spatial structure of dusky rockfish (Sebastes variabilis). Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002534.

Tomiyasu:2021:STR

[TSI^{+}21] Makoto Tomiyasu, Hokuto Shirakawa, Yuki Iino, Daichi Oshiyama, Masahiro Ogawa, Takashi Kitagawa, Hiromichi Mitamura, Nobuaki Arai, Yoshinori Miyamoto, Keiichi Uchida, Kenji Minami, and Kazushi Miyashita. Sonic tagging reveals age and size-specific spatial variation during Pacific herring spawning migrations in northern Japan. Fisheries Research, 242(??):Article 106020, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100148X.

Theberge:2024:AEF

[TSPK24] Kaitlyn Theberge, Tonje K. Sørdalen, Tracy L. Pugh, and Holly K. Kindsvater. Assessing the effects of female protections on size structure and spawning potential in two clawed lobster fisheries subject to varying exploitation levels. Fisheries Research, 280(??):??, December 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002194.

Tesfaye:2023:SGP

[TSS^{+}23] Million Tesfaye, Allan T. Souza, Kateřina Soukalová, Marek Šmejkal, Josef Hejzlar, Marie Prchalová, Milan Říha, Milan Muška, Mojmír Vašek, Jaroslava Frouzová, Petr Blabolil, David S. Boukal, and Jan Kubečka. Somatic growth of pikeperch (Stizostedion lucioperca) in relation to variation in temperature and eutrophication in a Central Europe lake. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623002175.

Truong:2025:PGM

[TTC^{+}25] Oanh Thi Truong, Sang Quang Tran, Kent E. Carpenter, Quyen Dang Ha Vu, Thuy-Yen Duong, Mie Mie Kyaw, Chaiwut Grudpan, Van Ngo Thai Bich, and Binh Thuy Dang. Population genetics of Macrognathus siamensis (Synbranchiformes: Mastacembelidae): Implications for nonmigratory fishery resources in the Mekong River basin. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624002741.

Tanaka:2025:ALF

[TTK^{+}25] Kenzo Tanaka, Makoto Tomiyasu, Ryo Kusaka, Shin Sugiyama, Evgeny A. Podolskiy, and Yasuzumi Fujimori. Artisanal longline fishing for Greenland halibut (Reinhardtius hippoglossoides) operated under sea ice using a metal plate kite in northwest Greenland. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002674.

Thierry:2020:HPB

[TTL^{+}20] Nyatchouba Nsangue Bruno Thierry, Hao Tang, Xu Liuxiong, Xingxing You, Fuxiang Hu, Njomoue Pandong Achile, and Richard Kindong. Hydrodynamic performance of bottom trawls with different materials, mesh sizes, and twine thicknesses. Fisheries Research, 221(??):Article 105403, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302589.

Tanaka:2024:SVO

[TTYT24] Shota Tanaka, Shizuna Togoshi, Naotaka Yasue, and Akinori Takasuka. Seasonal variability in the otolith and somatic size relationship of Japanese anchovy larvae: Counter effects of somatic growth and temperature. Fisheries Research, 275 (??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000912.

Tomiyama:2021:SBD

[TYYK21] Takeshi Tomiyama, Manabu Yamada, Akibumi Yamanobe, and Yutaka Kurita. Seasonal bathymetric distributions of three coastal flatfishes: Estimation from logbook data for trawl and gillnet fisheries. Fisheries Research, 233 (??):Article 105733, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783620302502.

Tsagarakis:2024:FAV

[TZL^{+}24] Konstantinos Tsagarakis, Walter Zupa, Alessandro Ligas, Claudia Musumeci, George Tserpes, and Maria Teresa Spedicato. Factors affecting the variability of discards in Mediterranean bottom trawl fisheries. Fisheries Research, 274(??): ??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000432.

Uhlmann:2021:FST

[UAB^{+}21] Sven Sebastian Uhlmann, Bart Ampe, Noémi Van Bogaert, Christian Vanden Berghe, and Bart Vanelslander. Flatfish survivors have tales to tell: Cold seawater and reduced deployment duration contribute to the survival of European plaice (Pleuronectes platessa) discarded by Belgian beam trawlers. Fisheries Research, 240(??):Article 105966, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000941.

Uhlmann:2023:ECC

[UAG^{+}23] Sven Sebastian Uhlmann, Bart Ampe, Joanna Marie Goley, Esther Savina, and Matt K. Broadhurst. Effects of catch composition on the fate of European plaice (Pleuronectes platessa) discarded from Belgian beam trawlers. Fisheries Research, 261(??):Article 106616, May 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000097.

Uehara:2023:VOD

[UFYT23] Shinji Uehara, Yuichiro Fujinami, Tetsuo Yamada, and Takeshi Tomiyama. Validation of otolith daily ring deposition and hatch date estimations in larval and juvenile Japanese flounder Paralichthys olivaceus. Fisheries Research, 266(??): ??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001844.

Utne-Palm:2020:FRA

[UPBH^{+}20] Anne Christine Utne-Palm, André S. Bogevik, Odd-Børre Humborstad, Tone Aspevik, Michael Pennington, and Svein Løkkeborg. Feeding response of Atlantic cod (Gadus morhua) to attractants made from by-products from the fishing industry. Fisheries Research, 227(??):Article 105535, July 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300527.

Uhlmann:2020:DIA

[UVA20] Sebastian S. Uhlmann, Steven Verstockt, and Bart Ampe. Digital image analysis of flatfish bleeding injury. Fisheries Research, 224(??):Article 105470, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578361930325X.

Vasbinder:2020:ELH

[VA20] Kelly Vasbinder and Cameron Ainsworth. Early life history growth in fish reflects consumption-mortality tradeoffs. Fisheries Research, 227(??):Article 105538, July 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300552.

Velez-Arellano:2020:LTA

[VAVQGD^{+}20] Nurenskaya Vélez-Arellano, Fausto Valenzuela-Quiñonez, Federico Andrés García-Domínguez, Daniel Bernardo Lluch-Cota, José Luis Gutiérrez-González, and Raúl Octavio Martínez-Rincón. Long-term analysis on the spawning activity of green (Haliotis fulgens) and pink (Haliotis corrugata) abalone along the central west coast of Baja California. Fisheries Research, 228(??):Article 105588, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301053.

Vincent:2020:PEP

[VBB20] Matthew T. Vincent, Travis O. Brenden, and James R. Bence. Parameter estimation performance of a recaptureconditioned integrated tagging catch-at-age analysis model. Fisheries Research, 224(??):Article 105451, April 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303066.

Vu:2022:LHS

[VBL^{+}22] An V. Vu, Lee J. Baumgartner, Karin E. Limburg, Gregory S. Doran, Martin Mallen-Cooper, Bronwyn M. Gillanders, Jason D. Thiem, Julia A. Howitt, Cameron M. Kewish, Juliane Reinhardt, and Ian G. Cowx. Life history strategies of Mekong pangasiid catfishes revealed by otolith microchemistry. Fisheries Research, 249(??):Article 106239, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000169.

Vu:2024:DMS

[VBL^{+}24] An V. Vu, Lee J. Baumgartner, Karin E. Limburg, Bronwyn M. Gillanders, Martin Mallen-Cooper, Julia A. Howitt, Jason D. Thiem, Gregory S. Doran, Cameron M. Kewish, and Ian G. Cowx. Diverse migration strategies of ariid catfishes along a salinity gradient in the Mekong River. Fisheries Research, 279(??):??, November 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001978.

vanBeest:2023:VIE

[vBMP^{+}23] Gabrielle Shira van Beest, Francisco Esteban Montero, Francesc Padrós, Juan Antonio Raga, and Ana Born-Torrijos. Visible implant elastomers in gilthead seabream (Sparus aurata L.) for experimental research: Preferred injection sites to optimize tag retention and minimize histological effects. Fisheries Research, 262(??):Article 106651, June 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000449.

Vautier:2023:QEB

[VCG^{+}23] Marine Vautier, Cécile Chardon, Chloé Goulon, Jean Guillard, and Isabelle Domaizon. A quantitative eDNA-based approach to monitor fish spawning in lakes: Application to European perch and whitefish. Fisheries Research, 264(??): ??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001017.

Vasquez-Carrillo:2021:IES

[VCPO21] Catalina Vásquez-Carrillo and Manuela Peláez-Ossa. Insights into the ecology of sea turtles and the fisheries of eastern guajira from the traditional knowledge of fishermen. Fisheries Research, 238(??):Article 105915, June 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000436.

Villa-Diharce:2021:MIG

[VDCMRF^{+}21] Enrique R. Villa-Diharce, Miguel A. Cisneros-Mata, Demetrio Rodríguez-Félix, Evlin A. Ramírez-Félix, and Guillermo Rodríguez-Domínguez. Molting and individual growth models of Callinectes bellicosus. Fisheries Research, 239(??):Article 105897, July 2021. CODEN FISRDJ.

ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783621000254.

vandenHeuvel:2020:CRR

[vdHBBR20] Lotte van den Heuvel, Małgorzata Blicharska, Samuel Blyth, and Patrik Rönnbäck. Catch reporting in recreational fishing: Swedish anglers’ attitudes and preferences, and the effect of social factors. Fisheries Research, 223(??):Article 105444, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783619302991.

vanderHammen:2020:PRD

[vdHC20] Tessa van der Hammen and Chun Chen. Participation rate and demographic profile in recreational angling in The Netherlands between 2009 and 2017. Fisheries Research, 229(??):Article 105592, September 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301090.

vandenHeuvel:2023:WYS

[vdHR23] Lotte van den Heuvel and Patrik Rönnbäck. What you see isn’t always what you get: On how anglers’ fish stock perceptions are influenced by motivations, satisfaction and engagement. Fisheries Research, 258(??):Article 106519, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362200296X.

vanDeurs:2023:IBF

[vDJB^{+}23] Mikael van Deurs, Nis S. Jacobsen, Jane W. Behrens, Ole Henriksen, and Anna Rindorf. The interactions between fishing mortality, age, condition and recruitment in exploited fish populations in the North Sea. Fisheries Research, 267(??):??, November 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002151.

Vagenas:2021:CSS

[VK21a] Georgios Vagenas and Paraskevi K. Karachle. Corrigendum to “sBCB (standardized Burn–Crack–Burn): a standardized modification of burning method for otolith age estimation” [Fish. Res. 233 (January) (2021) 105767]. Fisheries Research, 234(??):Article 105818, February 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620303350. See [VK21b].

Vagenas:2021:SSB

[VK21b] Georgios Vagenas and Paraskevi K. Karachle. sBCB (standardized Burn–Crack–Burn): a standardized modification of burning method for otolith age estimation. Fisheries Research, 233(??):Article 105767, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0165783620302848. See corrigendum [VK21a].

Vaz:2025:RSG

[VKS^{+}25] Ana C. Vaz, Mandy Karnauskas, Kate Siegfried, Matthew W. Smith, LaTreese S. Denson, Christopher Gardner, and John F. Walter. Red snapper in the Gulf of Mexico: Can offshore populations export larval subsidies to nearshore fishing grounds? Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000086.

Veiga-Malta:2020:TSS

[VMFF^{+}20] T. Veiga-Malta, J. P. Feekings, R. P. Frandsen, B. Herrmann, and L. A. Krag. Testing a size sorting grid in the brown shrimp (Crangon crangon Linnaeus, 1758) beam trawl fishery. Fisheries Research, 231(??):Article 105716, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620302332.

Ventero:2021:AEB

[VMI21] Ana Ventero, Irene Marcos, and Magdalena Iglesias. Acoustic evidences of the beginning of anchovy (Engraulis encrasicolus) schooling in the Northern Alboran Sea (Mediterranean Sea). Fisheries Research, 239(??):Article 105950, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000783.

vanOverzee:2023:CCE

[vORP23] Harriët M. J. van Overzee, Adriaan D. Rijnsdorp, and Jan Jaap Poos. Changes in catch efficiency and selectivity in the beam trawl fishery for sole when mechanical stimulation is replaced by electrical stimulation. Fisheries Research, 260(??):Article 106603, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003800.

vanPoorten:2020:RTS

[vP20] Brett T. van Poorten. Recovery tactics for sockeye blocked from anadromy evaluated through decision analysis and value of information. Fisheries Research, 230(??):Article 105666, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620301831.

Venerus:2022:APS

[VP22] Leonardo A. Venerus and Ana M. Parma. An accesspoint survey approach to estimate recreational boat-fishing effort for stays of variable length. Fisheries Research, 254(??):Article 106429, October 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002065.

Vincent:2023:AIE

[VP23] Matthew T. Vincent and Graham M. Pilling. Assumptions influencing the estimation of natural mortality in a tagintegrated statistical model for western and central Pacific Ocean skipjack. Fisheries Research, 261(??):Article 106612, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362300005X.

vanPutten:2025:IPM

[vPDD^{+}25] Ingrid van Putten, Catherine M. Dichmont, Natalie A. Dowling, Roy A. Deng, Sean Pascoe, and André E. Punt. Interconnected partnerships: Mapping collaborations in Australian fisheries stock assessment. Fisheries Research, 282(??):??, February 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783625000189.

Vadziutsina:2020:RFT

[VR20] Maria Vadziutsina and Rodrigo Riera. Review of fish trap fisheries from tropical and subtropical reefs: Main features, threats and management solutions. Fisheries Research, 223(??):Article 105432, March 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302875.

Vincent:2022:EMF

[VRS^{+}22] B. Vincent, M. Robert, J. Simon, J. P. Vacherot, and R. Faillettaz. Exploring the mechanics of fish escape attempts through mesh. Fisheries Research, 248(??):Article 106195, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003234.

Opstal:2023:RBU

[VS23] Mattias Van Opstal and Maarten Soetaert. Reducing the bycatch of undersized whiting (Merlangius merlangus) in flyshooting fisheries using a square mesh escape panel. Fisheries Research, 260(??):Article 106591, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362200368X.

Vogelmann:2022:UZD

[VTS^{+}22] Christian Vogelmann, Maxim Teichert, Michael Schubert, Andreas Martens, Sabine Schultes, and Herwig Stibor. The usage of a zooplankton digitization software to study plankton dynamics in freshwater fisheries. Fisheries Research, 251(??):Article 106326, July 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001035.

Vainikka:2024:BSY

[VTSI^{+}24] Anssi Vainikka, Aatu Turunen, Andrés Salgado-Ismodes, Eliisa Lotsari, Mikko Olin, Jukka Ruuhijärvi, Hannu Huuskonen, Céline Arzel, Petri Nummi, and Kimmo K. Kahilainen.

Biomass and sustainable yields of Eurasian perch (Perca fluviatilis) in small boreal lakes with respect to lake properties and water quality. Fisheries Research, 271(??):??, March 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623003156.

vanZyl:2025:SMM

[vZvdHCA25] Danie van Zyl, Grant van der Heever, Andrew Cockcroft, and Lutz Auerswald. Seasonal migration of a male Jasus lalandii sub-population. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002650.

Waterhouse:2022:UGM

[WAA^{+}22] Lynn Waterhouse, Lisa Ailloud, Riley Austin, Walter J. Golet, Ashley Pacicco, Allen H. Andrews, Khady Diouf, Yacine Ndiour, Kyne Krusic-Golub, Guelson da Silva, and John M. Hoenig. Updated growth models for bigeye tuna (Thunnus obesus) in the Atlantic Ocean. Fisheries Research, 253(??):Article 106317, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000947.

Wetz:2020:ATV

[WASS20] Jennifer J. Wetz, Matthew J. Ajemian, Brooke Shipley, and Gregory W. Stunz. An assessment of two visual survey methods for documenting fish community structure on artificial platform reefs in the Gulf of Mexico. Fisheries Research, 225(??):Article 105492, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300096.

Wilson:2023:OSB

[WBA23] Lorna I. Wilson, Richard E. Brenner, and Beverly A. Agler. Overlapping scales: Bias and variability in Alaska Chinook salmon age estimates. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623001364.

Was-Barcz:2023:GSP

[WBBG^{+}23] Anna Wa̧s-Barcz, Rafał Bernaś, Martyna Greszkiewicz, Adam M. Lejk, and Dariusz P. Fey. Genetic structure of pike (Esox lucius Linnaeus, 1758) populations along the Polish coast of the southern Baltic Sea: Comparison to Danish brackish population. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001029.

Wiens:2021:GAI

[WBD^{+}21] Lauren N. Wiens, Robert Bajno, Jillian T. Detwiler, Muhammad Yamin Janjua, and Ross F. Tallman. Genetic assessment of inconnu (Stenodus leucichthys) in Great Slave Lake, Northwest Territories, Canada. Fisheries Research, 234(??):Article 105784, February 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303015.

Wood:2024:FSS

[WCC24] Madeline V. Wood, Felipe M. Carvalho, and Leandro Castello. Fishing shrinks the size structure of exploited coral reef fishes in Brazil. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624000936.

Walters:2022:RAM

[WCGB22] Erin A. Walters, Claire E. Crowley, Ryan L. Gandy, and Donald C. Behringer. A reflex action mortality predictor (RAMP) for commercially fished blue crab Callinectes sapidus in Florida. Fisheries Research, 247(??):Article 106188, March 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003167.

Wakefield:2020:LSS

[WCLN20] Corey B. Wakefield, Peter G. Coulson, Luke Loudon, and Stephen J. Newman. Latitudinal and sex-specific differences in growth and an exceptional longevity for the Maori snapper Lutjanus rivulatus from north-western Australia. Fisheries Research, 230(??):Article 105634, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362030151X.

Weerasekera:2024:TPS

[WCN^{+}24] S. J. W. W. M. M. P. Weerasekera, Noel G. Cadigan, Kunasekaran Nirmalkanna, Paul M. Regular, and Rick M. Rideout. Trends in population starvation mortality based on a spatiotemporal model of condition: Part 1: a case study of Atlantic cod on the Southern Grand Bank. Fisheries Research, 278(??):??, October 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001772.

Winker:2020:JSI

[WCT^{+}20] Henning Winker, Felipe Carvalho, James T. Thorson, Laurance T. Kell, Denham Parker, Maia Kapur, Rishi Sharma, Anthony J. Booth, and Sven E. Kerwath. JABBA-Select: Incorporating life history and fisheries’ selectivity into surplus production models. Fisheries Research, 222(??):Article 105355, February 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302103.

Westerbom:2025:CPF

[WEH^{+}25] Mats Westerbom, Camilla Ekblad, Juhani Hopkins, Toni Laaksonen, Mikko Olin, Antti Ovaskainen, and Veijo Jormalainen. Cormorant predation in fyke net fishing: the direct effects of a protected bird on coastal commercial fishing. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624002984.

Watson:2019:AEM

[WGC^{+}19] Jabin R. Watson, Harriet R. Goodrich, Rebecca L. Cramp, Matthew A. Gordos, and Craig E. Franklin. Assessment of the effects of microPIT tags on the swimming performance of small-bodied and juvenile fish. Fisheries Research, 218(??):22–28, October 2019. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016578361930116X. See corrigendum [WGC^{+}21].

Watson:2021:CAE

[WGC^{+}21] Jabin R. Watson, Harriet R. Goodrich, Rebecca L. Cramp, Matthew A. Gordos, and Craig E. Franklin. Corrigendum to “Assessment of the effects of microPIT tags on the swimming performance of small-bodied and juvenile fish” [Fish. Res. 218 October (2019) 22–28]. Fisheries Research, 241(??):Article 106005, September 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621001338. See [WGC^{+}19].

Wang:2021:EEF

[WGFM21] Yanjun Wang, Ali Gharouni, Kevin D. Friedland, and Donald C. Melrose. Effect of environmental factors and density-dependence on somatic growth of Eastern Georges Bank haddock (Melanogrammus aeglefinus). Fisheries Research, 240(??):Article 105954, August 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621000825.

Winger:2024:SDT

[WGNM24] Paul D. Winger, Alex Gardner, Truong X. Nguyen, and Liam McGregor. Spreading a demersal trawl without otter boards: a proof of concept using flexible foils. Fisheries Research, 275 (??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000778.

Williams:2023:FBR

[WGW^{+}23] Kresimir Williams, Pamela Goddard, Rachel Wilborn, David Bryan, and Chris Rooper. Fish behavior in response to an approaching underwater camera. Fisheries Research, 268(??):??, December 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623002163.

Wetzel:2023:APH

[WH23] Chantel R. Wetzel and Owen S. Hamel. Applying a probability harvest control rule to account for increased uncertainty in setting precautionary harvest limits from past stock assessments. Fisheries Research, 262(??):Article 106659, June 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000528.

Wang:2022:IOS

[WHCF22] Yan Wang, Peiwu Han, Xinjun Chen, and Zhou Fang. Interannual and ontogenetic stage differences on trophic ecology of neon flying squid Ommastrephes bartramii inferred from stable isotope analyses in beaks. Fisheries Research, 249(??):Article 106252, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000297.

Watanabe:2023:COC

[WHM23] Akira Watanabe, Yuta Hara, and Hiroyuki Matsuda. Combining output control and fishing pressure limitations improves the management of the Japanese sardine Sardinops melanostictus. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001807.

Werner:2024:OWF

[WHR^{+}24] Karl M. Werner, Holger Haslob, Anna F. Reichel, Antje Gimpel, and Vanessa Stelzenmüller. Offshore wind farm foundations as artificial reefs: the devil is in the detail. Fisheries Research, 272(??):??, April 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000018.

Wilhelm:2025:OSA

[WJN^{+}25] M. R. Wilhelm, C. E. Jagger, N. M. Nghipangelwa, B. A. Pringle, P. W. Shaw, W. M. Potts, R. Henriques, and N. J. McKeown. Otolith shape analysis as a tool for species identification and management of cryptic congeners in the northern Benguela ocean warming hotspot. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003266.

Wulfing:2024:ANT

[WKBMW24] Sophie Wulfing, Ahilya Kadba, Mez Baker-Médard, and Easton R. White. Assessing the need for temporary fishing closures to support sustainability for a small-scale octopus fishery. Fisheries Research, 276(??):??, August 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001097.

Wanzenbock:2020:HTS

[WKSF20] Josef Wanzenböck, Jan Kubecka, Zuzana Sajdlova, and Jaroslava Frouzova. Hydroacoustic target strength vs. fish length revisited: Data of caged, free-swimming European whitefish (Coregonus lavaretus L.) suggest a bi-phasic linear relationship under a limited range of tilt angles. Fisheries Research, 229(??):Article 105620, September 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301375.

Wolfenkoehler:2023:VSS

[WLG^{+}23] Wyatt Wolfenkoehler, James M. Long, Ryan Gary, Richard A. Snow, Jason D. Schooley, Lindsey A. Bruckerhoff, and Robert C. Lonsinger. Viability of side-scan sonar to enumerate paddlefish, a large pelagic freshwater fish, in rivers and reservoirs. Fisheries Research, 261(??):Article 106639, May 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623000322.

Wang:2021:AFR

[WLZ^{+}21] Xiaoyan Wang, Guoqing Lu, Linlin Zhao, Xiaoqin Du, and Tianxiang Gao. Assessment of fishery resources using environmental DNA: the large yellow croaker (Larimichthys crocea) in the East China Sea. Fisheries Research, 235(??):Article 105813, March 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303301.

Wirgin:2022:GPS

[WMSW22] Isaac Wirgin, Lorraine Maceda, Joseph Stabile, and John Waldman. Genetic population structure of summer flounder Paralichthys dentatus using microsatellite DNA analysis. Fisheries Research, 250(??):Article 106270, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000479.

Wirgin:2020:ACP

[WMT^{+}20] Isaac Wirgin, Lorraine Maceda, Matt Tozer, Joseph Stabile, and John Waldman. Atlantic coastwide population structure of striped bass Morone saxatilis using microsatellite DNA analysis. Fisheries Research, 226(??):Article 105506, June 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300230.

Walmsley:2025:ETA

[WOG^{+}25] S. F. Walmsley, J. F. Oaten, T. Götz, D. Rodmell, N. de Rozarieux, and C. J. Sweeting. Effectiveness of targeted acoustic startle technology on seal depredation in an inshore gillnet fishery. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002728.

Whitaker:2020:DFS

[WPB^{+}20] J. M. Whitaker, L. E. Price, J. C. Boase, L. Bernatchez, and A. B. Welsh. Detecting fine-scale population structure in the age of genomics: a case study of lake sturgeon in the Great Lakes. Fisheries Research, 230(??):Article 105646, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301636.

White:2022:DAF

[WPB22] Allison L. White, William F. Patterson, and Kevin M. Boswell. Distribution of acoustic fish backscatter associated with natural and artificial reefs in the Northeastern Gulf of Mexico. Fisheries Research, 248(??):Article 106199, April 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621003271.

Wilcox:2021:SVS

[WPGO21] Steven H. Wilcox, Tracy L. Pugh, Robert P. Glenn, and Kenneth Oliveira. Spatial variation in size and age at maturation and growth of the channeled whelk (Busycotypus canaliculatus) in Southern Massachusetts. Fisheries Research, 239(??):Article 105926, July 2021. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783621000540.

Wiley:2020:UMO

[WPLF20] John Wiley, Cassandra Pardee, Gwen Lentes, and Emma Forbes. Unaccounted mortality and overview of the Hawaiian Kona crab Ranina ranina (Linnaeus) fishery. Fisheries Research, 226(??):Article 105517, June 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300345.

Walton:2025:FIF

[WQGS25] Lydia N. Walton, Micah Quindazzi, Stéphane Gauthier, and Catherine Stevens. Fish ID face-off: a comparison of genetic barcoding and otolith shape analysis for streamlining species identification of mesopelagic fishes. Fisheries Research, 281(??):??, January 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624003187.

Watson:2020:GMG

[WS20] L. Cynthia Watson and Donald J. Stewart. Growth and mortality of the giant arapaima in Guyana: Implications for recovery of an over-exploited population. Fisheries Research, 231(??):Article 105692, November 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302095.

Williams:2024:ECW

[WS24] Erik H. Williams and Kyle W. Shertzer. Estimating catch within integrated stock assessments: Distinguishing truth from error in sparse-spike data. Fisheries Research, 274(??): ??, June 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000584.

Williams:2022:ELC

[WSB22] Erik H. Williams, Kyle W. Shertzer, and Nate Bacheler. Estimating length composition of fish observed with stereovideo cameras: a simulation study with application to red snapper (Lutjanus campechanus). Fisheries Research, 254(??):Article 106424, October 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002016.

Wolff:2024:CBF

[WSB24] François-Charles Wolff, Frédéric Salladarré, and Laurent Baranger. A classification of buyers in first-sale fish markets: Evidence from France. Fisheries Research, 275(??): ??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624000869.

Williams:2022:IRL

[WSF22] Benjamin Williams, S. Lynne Stokes, and John Foster. Investigating record linkage for combining voluntary catch reports with a probability sample. Fisheries Research, 251(??):Article 106301, July 2022. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622000789.

Wiley:2021:AST

[WSL21] John Wiley, Marlowe Sabater, and Brian Langseth. Aerial survey as a tool for understanding bigeye scad (Selar crumenophthalmus) dynamics around the island of O’ahu, Hawai’i. Fisheries Research, 236(??):Article 105866, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303830.

Wetzel:2024:OAS

[WSL^{+}24] Chantel R. Wetzel, Christine C. Stawitz, Bai Li, Kelli F. Johnson, and Giselle M. Schmitz. Ordering the alphabet soup: Strategies to improve consistency and develop a framework of tools for fisheries science vocabulary. Fisheries Research, 278(??):??, October 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001681.

Wolf:2023:UFO

[WST^{+}23] Skylar L. Wolf, Dusty A. Swedberg, Evan P. Tanner, Samuel D. Fuhlendorf, and Shannon K. Brewer. Using fiber-optic distributed temperature sensing in fisheries applications: an example from the Ozark Highlands. Fisheries Research, 258(??):Article 106542, February 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003198.

Woodward:2023:SDS

[WSUN^{+}23] Caroline Woodward, Michelle Schärer-Umpierre, Richard S. Nemeth, Richard Appeldoorn, and Laurent M. Chérubin. Spatial distribution of spawning groupers on a Caribbean reef from an autonomous surface platform. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016578362300187X.

Wang:2025:ECM

[WSWL25] Shun Wang, Liming Song, Wenxin Wang, and Yuwei Li. Effect of connection methods on the mechanical characteristics of tuna longline fishing gear. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578362400314X.

Wang:2021:ERF

[WW21a] Yixuan Wang and Nuo Wang. Exploring the role of the fisheries sector in China’s national economy: an input–output analysis. Fisheries Research, 243(??):Article 106055, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001831.

Weber:2021:BEM

[WW21b] Robert E. Weber and Michael J. Weber. Behavior, escapement, and mortality of adult Muskellunge in Midwestern reservoirs. Fisheries Research, 239(??):Article 105945, July 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621000734.

Wakefield:2020:VLH

[WWF^{+}20] Corey B. Wakefield, Ashley J. Williams, Emily A. Fisher, Norman G. Hall, Sybrand A. Hesp, Tuikolongahau Halafihi, Jeremie Kaltavara, Elodie Vourey, Brett M. Taylor, Joseph M. O’Malley, Simon J. Nicol, Brent S. Wise, and Stephen J. Newman. Variations in life history characteristics of the deep-water giant ruby snapper (Etelis sp.) between the Indian and Pacific Oceans and application of a data-poor assessment. Fisheries Research, 230(??):Article 105651, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620301685.

Williams:2020:IGS

[WWO20] Samuel M. Williams, Jamie Wyatt, and Jennifer R. Ovenden. Investigating the genetic stock structure of blue marlin (Makaira nigricans) in the Pacific Ocean. Fisheries Research, 228(??):Article 105565, August 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620300825.

Wei:2025:DTN

[WWT^{+}25] Xu Wei, Yan Wang, James R. Tweedley, Neil R. Loneragan, Tao Tian, Zhilin Wang, Yanchao Zhang, Wencong An, Longfei Xu, and Zhongxin Wu. Diet and trophic niches of sympatric Seriola species revealed by stomach content and multi-tissue stable isotope analyses. Fisheries Research, 282(??):??, February 2025. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783625000098.

Wo:2024:SPS

[WXJ^{+}24] Jia Wo, Binduo Xu, Yupeng Ji, Chongliang Zhang, Ying Xue, and Yiping Ren. Species portfolio schemes buffering the risk of overexploitation in mixed fisheries management. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000444.

Wong:2020:FRR

[WY20] Hoong Sang Wong and Chen Chen Yong. Fisheries regulation: a review of the literature on input controls, the ecosystem, and enforcement in the Straits of Malacca of Malaysia. Fisheries Research, 230(??):Article 105682, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301995.

Wang:2025:BCP

[WYM^{+}25] Zi Wang, Naizheng Yan, Tohru Mukai, Kohei Hasegawa, and Jun Yamamoto. Broadband characteristics of Pacific herring (Clupea pallasii) and Pacific chub mackerel (Scomber japonicus) target strength identified using the tether method. Fisheries Research, 283(??):??, March 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783625000438.

Wang:2022:RWS

[WZL^{+}22] Shuxian Wang, Shengmao Zhang, Yang Liu, Jiaze Zhang, Yongwen Sun, Yuhao Yang, Huijuan Hu, Ying Xiong, Wei Fan, Fei Wang, and Fenghua Tang. Recognition on the working status of Acetes chinensis quota fishing vessels based on a 3D convolutional neural network. Fisheries Research, 248(??):Article 106226, April 2022. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000030.

Wang:2021:FPL

[WZS^{+}21] Kun Wang, Chongliang Zhang, Ming Sun, Binduo Xu, Yupeng Ji, Ying Xue, and Yiping Ren. Fishing pressure and lifespan affect the estimation of growth parameters using ELEFAN. Fisheries Research, 238(??):Article 105903, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578362100031X.

Wang:2020:SOS

[WZX^{+}20] Kun Wang, Chongliang Zhang, Binduo Xu, Ying Xue, and Yiping Ren. Selecting optimal bin size to account for growth variability in Electronic LEngth Frequency ANalysis (ELE-FAN). Fisheries Research, 225(??):Article 105474, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303297.

Xing:2021:EIF

[XCB^{+}21] Lei Xing, Yong Chen, Robert Boenish, Kisei R. Tanaka, Nicolas Barrier, and Yiping Ren. Evaluating the impacts of fishing and migratory species in a temperate bay of China using the ecosystem model OSMOSE-JZB. Fisheries Research, 243(??):Article 106051, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362100179X.

Xuan:2023:YSO

[XDX^{+}23] Kui Xuan, Limiao Deng, Ying Xiao, Peidong Wang, and Juan Li. SO-YOLOv5: Small object recognition algorithm for sea cucumber in complex seabed environment. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001030.

Xu:2020:IRF

[XMCC20] Luoliang Xu, Mackenzie Mazur, Xinjun Chen, and Yong Chen. Improving the robustness of fisheries stock assessment models to outliers in input data. Fisheries Research, 230(??):Article 105641, October 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620301582.

Xu:2024:EIR

[XMLCMV24] Haikun Xu, Mark N. Maunder, Cleridy E. Lennert-Cody, and Carolina V. Minte-Vera. Evaluating the impacts of reduced longline fishing effort on the standardization of longline catch-per-unit-effort for bigeye tuna in the eastern Pacific Ocean. Fisheries Research, 278(??):??, October 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624001759.

Xu:2023:OMR

[XqRJ^{+}23] Qing Xu, Qing qiang Ren, Tao Jiang, Chen rui Jiang, Lü ping Fang, Ming zhe Zhang, Jian Yang, and Min Liu. Otolith microchemistry reveals various habitat uses and life histories of Chinese gizzard shad Clupanodon thrissa in the Min River and the estuary, Fujian Province, China. Fisheries Research, 264(??):??, August 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001169.

Xi:2023:SSH

[XSS^{+}23] Xiaoyu Xi, Peng Sun, Runlong Sun, Yongjun Tian, and Mikko Heino. Size-selective harvesting alters biological traits of marine medaka (Oryzias melastigma). Fisheries Research, 266(??):??, October 2023. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001686.

Xu:2021:AFD

[XWD^{+}21] Lei Xu, Xuehui Wang, Kay Van Damme, Delian Huang, Yafang Li, Lianggen Wang, Jiajia Ning, and Feiyan Du. Assessment of fish diversity in the South China Sea using DNA taxonomy. Fisheries Research, 233(??):Article 105771, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620302885.

Yassir:2023:AFS

[YAO^{+}23] Anas Yassir, Said Jai Andaloussi, Ouail Ouchetto, Kamal Mamza, and Mansour Serghini. Acoustic fish species identification using deep learning and machine learning algorithms: a systematic review. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783623001832.

Yilanci:2022:EKC

[YCC22] Veli Yilanci, Ibrahim Cutcu, and Bilal Cayir. Is the environmental Kuznets curve related to the fishing footprint? Evidence from China. Fisheries Research, 254 (??):Article 106392, October 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0165783622001692.

Yellapu:2025:GED

[YFJ^{+}25] Bhargavi Yellapu, Ahmad Farhadi, Andrew G. Jeffs, Gregory Smith, and Shane D. Lavery. Geospatial and environmental drivers of genetic divergence in the Indo-West Pacific spiny lobster Panulirus ornatus. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783624002601.

Yang:2020:GWP

[YGMJ20] Tian-Yan Yang, Tian-Xiang Gao, Wei Meng, and Yan-Lin Jiang. Genome-wide population structure and genetic diversity of Japanese whiting (Sillago japonica) inferred from genotyping-by-sequencing (GBS): Implications for fisheries management. Fisheries Research, 225(??):Article 105501, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620300187.

Yu:2021:CEM

[YH21a] Jinkai Yu and Qingchao Han. Corrigendum to “Exploring the management policy of distant water fisheries in China: Evolution, challenges and prospects” [Fish. Res. 236 (2021) 105849]. Fisheries Research, 242(??):Article 106058, October 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621001867. See [YH21b].

Yu:2021:EMP

[YH21b] Jinkai Yu and Qingchao Han. Exploring the management policy of distant water fisheries in China: Evolution, challenges and prospects. Fisheries Research, 236(??):Article 105849, April 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783620303660. See corrigen-

dum [YH21a].

Yu:2024:EST

[YHC^{+}24] Mengjie Yu, Bent Herrmann, Kristine Cerbule, Changdong Liu, Yilin Dou, Liyou Zhang, Linjie Li, and Yanli Tang. The effect of soak time on pot escape opening selectivity in swimming crab (Portunus trituberculatus) fishery. Fisheries Research, 275(??):??, July 2024. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624000845.

Yano:2022:YLT

[YHST22] Toshikazu Yano, Tsutomu Hattori, Yasutoki Shibata, and Sho Tanaka. Over 120 years of landing trends in Japan, for the commercially exploited shark species, Squalus suckleyi. Fisheries Research, 249(??):Article 106257, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000340.

Yamashita:2020:CSR

[YIM^{+}20] Yasunori Yamashita, Yuichi Iwasaki, Toshimitsu Matsubara, Kyuma Suzuki, Yuhei Kanzawa, Takehiro Okuda, Kazuya Nishina, and Carlos Augusto Strüssmann. Comparison of survival rates between domesticated and semi-native char using Bayesian multi-variate state-space model. Fisheries Research, 221(??):Article 105380, January 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302358.

Yoshiyama:2023:IBF

[YiTM23] Taku Yoshiyama, Jun ichi Tsuboi, and Takashi F. Matsuishi. Interactions between fish and angler behaviour affect species-specific catchability of sympatric native charr and two introduced trout. Fisheries Research, 259(??):Article 106547, March 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783622003241.

Yoon:2023:EST

[YLP^{+}23] Euna Yoon, Hyungbeen Lee, Cheol Park, Yong-Deuk Lee, Kangseok Hwang, and Doo Nam Kim. Ex situ target strength of yellow croaker (Larimichthys polyactis) in a seawater tank. Fisheries Research, 260(??):Article 106610, April 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000036.

Yu:2023:NCB

[YLS^{+}23] Mengjie Yu, Changdong Liu, Liyuan Sun, Liyou Zhang, and Yanli Tang. A new concept for bycatch reduction in small-scale accordion-shaped trap fisheries of the Yellow Sea, China. Fisheries Research, 263(??):??, July 2023. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623000784.

Yuyang:2024:TNP

[YLX^{+}24] Zhang Yuyang, Wang Linlong, Sun Xin, Dong Jianyu, Zhan Qipeng, and Zhang Xiumei. Trophic niche partitioning of golden cuttlefish (Sepia esculenta) during ontogeny revealed by stable isotope analysis and DNA metabarcoding. Fisheries Research, 271(??):??, March 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623003144.

Yoshimura:2021:AFC

[YMS21] Mayumi Yoshimura, Toshiya Matsuura, and Ken Sugimura. Attitudes to forest conditions and fishing activities in the mountain area in Japan. Fisheries Research, 244(??):Article 106125, December 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002538.

Yan:2020:ADB

[YMYH20] Naizheng Yan, Tohru Mukai, Jun Yamamoto, and Kohei Hasegawa. Acoustic discrimination between juvenile walleye pollock and pointhead flounder. Fisheries Research, 224(??):Article 105434, April 2020. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783619302899.

Young:2020:EVP

[YRTP20] Zaida Young, Hugo Robotham, Camilo Torres, and Elizabeth Palta. Ex-vessel price monitoring design in smallscale fisheries: an application in Chilean fisheries. Fisheries Research, 230(??):Article 105591, October 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620301089.

Yochum:2021:ERB

[YSB^{+}21] Noëlle Yochum, Michael Stone, Karsten Breddermann, Barry A. Berejikian, John R. Gauvin, and David J. Irvine. Evaluating the role of bycatch reduction device design and fish behavior on Pacific salmon (Oncorhynchus spp.) escapement rates from a pelagic trawl. Fisheries Research, 236(??):Article 105830, April 2021. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620303477.

Yamaguchi:2022:IMR

[YTH22] Tadanori Yamaguchi, Katsumi Takayama, and Naoki Hirose. Influence of migratory route on early maturation of swordtip squid, Uroteuthis edulis, caught off western Kyushu Island, Japan. Fisheries Research, 249(??):Article 106233, May 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622000108.

Yamaguchi:2020:SAP

[YTHM20] Tadanori Yamaguchi, Katsumi Takayama, Naoki Hirose, and Michiya Matsuyama. The Sea of Amakusa playing the role of a distributor of swordtip squid (Uroteuthis edulis) migrating from the East China Sea to the east and west sides of Japan. Fisheries Research, 225(??):Article 105475, May 2020. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783619303303.

Yamashita:2022:FCA

[YTSS22] Yasunori Yamashita, Yuya Takagi, Kyuma Suzuki, and Carlos Augusto Strüssmann. Factors contributing to anglers’ satisfaction and their requests concerning recreational salmonid fisheries management: Insights from a questionnaire survey in strictly regulated, catch-and-release stream fishing areas. Fisheries Research, 256(??):Article 106464, December 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622002417.

Yu:2021:TPM

[YWC^{+}21] Wei Yu, Jian Wen, Xinjun Chen, Yi Gong, and Bilin Liu. Trans-Pacific multidecadal changes of habitat patterns of two squid species. Fisheries Research, 233(??):Article 105762, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783620302794.

Zak:2021:DPC

[Žák21] Jakub Žák. Diel pattern in common carp landings from angling competitions corresponds to their assumed foraging activity. Fisheries Research, 243(??):Article 106086, November 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783621002149.

Zumpano:2023:RTF

[ZCFG23] Francisco Zumpano, Sofía Copello, Marco Favero, and Germán O. García. Research trends and future perspectives of recreational fisheries in South America. Fisheries Research, 258(??):Article 106546, February 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016578362200323X.

Zang:2023:TAI

[ZCM^{+}23] Xiaoqin Zang, Thomas J. Carlson, Jayson J. Martinez, Jun Lu, and Zhiqun Daniel Deng. Towards assessing the impact of anthropogenic sound on fishes: Gaps, perspectives, and a case study of a large floating bridge. Fisheries Research, 265(??):??, September 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783623001406.

Zander:2022:CPC

[ZDF^{+}22] Katrin Zander, Fabienne Daurès, Yvonne Feucht, Loretta Malvarosa, Claudio Pirrone, and Bertrand le Gallic. Consumer perspectives on coastal fisheries and product labelling in France and Italy. Fisheries Research, 246 (??):Article 106168, February 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0165783621002964.

Zakes:2022:III

[ZDZR^{+}22] Zdzisław Zakȩś, Krystyna Demska-Zakȩś, Maciej Rożyński, Piotr Gomułka, and Rafał Rożyński. Influence of intraperitoneal implantation of 12 mm PIT on the welfare of juvenile brown trout (Salmo trutta). Fisheries Research, 255(??):Article 106458, November 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622002351.

Zhang:2021:EKP

[ZF21] Zane Zhang and Ken Fong. Estimation of key population parameters and MSY-based reference points for sidestripe shrimp (Pandalopsis dispar) in the Fraser River Delta, British Columbia. Fisheries Research, 238(??): Article 105893, June 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783621000217.

Zhang:2025:OCT

[ZHC^{+}25] Ying Zhang, Xiaoshuang Huang, Xinjun Chen, Bilin Liu, Xianghong Kong, and Qiankai Deng. Optimal configuration and three-dimensional light field of fishing lights on a squid fishing boat obtained using a space matrix algorithm. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400287X.

Zaman:2023:FUI

[ZHM23] Atik Uz Zaman, Patrik J. G. Henriksson, and Abdullah-Al Mamun. Fuel use intensity of hilsa fisheries in the lower Meghna River estuary of Bangladesh. Fisheries Research, 263 (??):??, July 2023. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623000772.

Zhou:2025:EDR

[ZJJ^{+}25] Chunhua Zhou, Shaoqing Jian, Zhizhong Jiang, Jinping Chen, Shan Ouyang, and Xiaoping Wu. Environmental DNA reveals spatial and temporal variation in fish communities before the 10-year fishing ban in the Poyang Lake Basin. Fisheries Research, 281(??):??, January 2025. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578362400256X.

Zhang:2024:GWS

[ZJY^{+}24] Ruhua Zhang, Yanqing Ji, Kun Ye, Lingxiao Liu, Haiguang Zhang, and Yunguo Liu. Genome-wide SNP-based diversity analysis and phylogeographic inference in the ark shell (Anadara broughtonii). Fisheries Research, 270(??):??, February 2024. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783623002850.

Zhang:2020:PGE

[ZLXL20] Bai-Dong Zhang, Yu-Long Li, Dong-Xiu Xue, and Jin-Xian Liu. Population genomic evidence for high genetic connectivity among populations of small yellow croaker (Larimichthys polyactis) in inshore waters of China. Fisheries Research, 225(??):Article 105505, May 2020. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783620300229.

Zitek:2023:ESS

[ZOS^{+}23] Andreas Zitek, Johannes Oehm, Michael Schober, Anastassiya Tchaikovsky, Johanna Irrgeher, Anika Retzmann, Bettina Thalinger, Michael Traugott, and Thomas Prohaska. Evaluating ^{87}Sr/^{86}Sr and Sr/Ca ratios in otoliths of different European freshwater fish species as fishery management tool in an Alpine foreland with limited geological variability. Fisheries Research, 260(??):Article 106586, April 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783622003630.

Zhang:2022:SGS

[ZOZW22] Xiaoyu Zhang, Mei Ouyang, Futie Zhang, and Jianwei Wang. Study on the genetic structure of wild and hatchery populations of Procypris rabaudi Tchang, an endemic fish in the upper Yangtze River. Fisheries Research, 245(??):Article 106134, January 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783621002629.

Zhao:2024:OCR

[ZSDZ24] Zhen Zhao, Illia Slypko, Kostiantyn Demianenko, and Guoping Zhu. Otolith chemistry reveals ontogenetic movement of the Antarctic toothfish (Dissostichus mawsoni) in the Amundsen Sea polynya, Antarctica. Fisheries Research, 276(??):??, August 2024. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783624001103.

Zebro:2025:ECE

[ZSWK25] Logan R. Zebro, Greg G. Sass, Melissa R. Wuellner, and Keith D. Koupal. Evaluating the contributions and economic costs for walleye fry and fingerling stocking in a large midwestern reservoir. Fisheries Research, 281(??):??, January 2025. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783624002716.

Zhang:2023:ILH

[ZYZ^{+}23] Yunlei Zhang, Huaming Yu, Chongliang Zhang, Binduo Xu, Yupeng Ji, Yiping Ren, and Ying Xue. Impact of life history stages on fish species interactions and spatio-temporal distribution. Fisheries Research, 266(??):??, October 2023. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783623001856.

Zhu:2022:TLB

[ZZ22] Jiuyang Zhu and Guoping Zhu. Trophic linkage between mackerel icefish (Champsocephalus gunnari) and Antarctic krill (Euphausia superba) at South Georgia. Fisheries Research, 253(??):Article 106366, September 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783622001436.

Zhang:2021:HSK

[ZZC^{+}21] Jun Zhang, Kui Zhang, Zuozhi Chen, Junde Dong, and Yongsong Qiu. Hydroacoustic studies on Katsuwonus pelamis and juvenile Thunnus albacares associated with light fishaggregating devices in the South China Sea. Fisheries Research, 233(??):Article 105765, January 2021. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0165783620302824.

Zhang:2024:SCP

[ZZH^{+}24] Jian Zhang, Honglin Zhang, Pingguo He, Jiaojiao Fei, Wenhua Chu, and Jiangao Shi. Selectivity of crab pots with escape vents and the optimal vent height for reducing the catch of sublegal swimming crab in the East China Sea pot fishery. Fisheries Research, 274(??):??, June 2024. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783624000663.